A Complete Bibliography of Publications in Isis, 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.37

Title word cross-reference

$119 [1604]. $120 [1570]. $124.95 [1558]. $130 [1591]. $139 [1590]. $171 [1552]. 1839 [2651]. $216 [1564]. $220 [1546]. $24.95 [1585, 1535]. $245 [1549]. $25 [1587, 1603]. $25.95 [1576]. $26 [1677]. $27.50 [1595]. $28.95 [1544]. $29.95 [1588, 1547, 1534, 1599]. $30 [1543]. $30.95 [1542]. $34.95 [1607, 1589]. $35 [1598, 1605]. $37.95 [1566]. $38 [1580, 1610]. $39.95 [1579, 1548, 1593]. $40 [1537, 1583, 1596, 1565]. $45 [1601]. $47.96 [1554]. $49.50 [1676]. $49.95 [1600]. $50 [1568, 1586, 1572, 1606, 1573]. $55 [1602]. $56.54 [1561]. $59.95 [1609]. $60 [1545]. $74 [1557]. $74.95 [1582]. $79.95 [1555]. $85 [1590, 1314]. $99 [1559, 1536, 1584]. ῑ [267].

-April [2651].

. [1539]. ...

[2879].

/Max [1188].

1 [3118]. 1.III [866]. 100 [1188, 2952]. 101 [1995]. 13.-16 [2699]. 1300 [2698]. 1424 [39]. 1450 [2930]. 1450-1700 [2930]. 150 [288]. 1500 [2629]. 1542 [2505]. 1550 [2602]. 1590-1731 [2848]. 1598 [1939]. 1600 [1343, 2933, 2337]. 1600-1874 [2933]. 1610 [2183]. 1614-1672 [2637]. 1630 [1560, 731]. 1640 [1555]. 1660 [555]. 1668 [2794]. 1676 [375]. 1700 [44, 2013, 2930, 2574, 2016, 1290, 793]. 1709 [195]. 1713 [1705]. 1715 [2852, 559]. 1720 [938]. 1722-1790 [2643]. 1725 [2109]. 1730 [2640, 2338]. 1730-1800 [2640]. 1737-1823 [2858]. 1750 [2730, 193, 728]. 1750-1950 [2730]. 1760 [1131, 1353, 2647, 2785]. 1760-1815 [2647]. 1766 [2642]. 1766-1820 [2642]. 1767 [1709]. 1770 [2940]. 1770-1930 [2940]. 1780 [2710]. 1780-1840 [2710]. 1786 [2576]. 1789 [1797]. 1791 [875]. 1792 [1132]. 1794 [799]. 18 [990, 272]. 1800 [23, 2640, 2112, 2024, 2323, 43, 47, 53, 1204, 2732, 774, 479, 658]. 1800-1900 [2732]. 1806 [448]. 1810 [1210, 1271]. 1815 [2647, 1711]. 1820 [1704, 2642, 2018, 283, 983]. 1830 [1572, 678, 888]. 1836 [3192, 1419]. 1840 [2859, 2710, 196, 936]. 1840-1910 [2941]. 1841 [2116]. 1845 [2277]. 1845-2000 [2277]. 1848 [1065, 2278, 68]. 1848-1914 [2278]. 1850 [849, 1571, 2280, 2349, 2804, 1007]. 1850-1900 [2280]. 1850-1950 [2281]. 1850-1980 [2804]. 1851 [2434]. 1854 [2651]. 1854-1960 [2282]. 1857 [58]. 1859 [1719, 1370, 2028]. 1860 [2872, 2283, 884, 544]. 1860-1920 [2283]. 1860-1940.} [2872]. 1860s [2523]. 1860s-1970s [2523]. 1865 [216]. 1867 [2026]. 1870 [1566, 2525, 2284]. 1870-1890 [2525]. 1870-1930 [2284]. 1870s [2679]. 1874 [2933, 590]. 1876 [2777, 1890, 2285]. 1876-1980 [2285]. 1877 [1357]. 1879 [2579]. 1880 [2527, 356]. 1880-1940 [2875]. 1880-1945 [2527]. 1880s [1733, 3299]. 1881 [74]. 1882 [2718]. 1882-1952 [2718]. 1887 [817]. 1888 [2273]. 1889 [1375]. 1889-1979 [2287]. 1890 [2525, 69]. 1893-1913 [2721]. 1895 [811]. 1898 [2879, 2810, 136]. 1898-1935 [2879]. 1898/99 [2810]. 18th [3003]. 1900 [3008, 2811, 2095, 2280, 1354, 1301, 2732, 2290, 1361, 633]. 1900-1933 [2719]. 1904 [204]. 1910 [1218, 2869, 2872, 1001]. 1910-1950 [2872]. 1913 [2621]. 1914 [1424, 2447, 1362, 2278, 285, 444]. 1918 [2363, 1818, 1068, 1002]. 1919 [2203]. 1919-1948 [2884]. 1919-1960s [2658]. 1920 [1306, 2283, 1217, 71, 582]. 1920er [1597, 3111]. 1920er-Jahren [1597, 3111]. 1920s [1450, 2554, 3111, 1822, 290]. 1921 [1507, 2581]. 1922 [2885]. 1922-1923 [2885]. 1923 [1507, 2885]. 1924 [2591]. 1925 [2132, 2287]. 1926 [1928, 450, 949]. 1929 [1656, 2369, 923]. 1930 [2806, 2801, 948, 2284, 1740, 2940, 2521, 384, 960, 445]. 1930s [1960, 2679]. 1931 [812]. 1933 [1590, 2209, 940]. 1934 [2120]. 1935 [2879]. 1937 [854]. 1938 [2127]. 1939 [2032, 2373, 452, 1956, 666]. 1940 [2442, 2443, 1355, 440, 2364, 397, 1503, 1732, 253]. 1940.} [2872]. 1940s [3299, 1733]. 1942 [2306, 2133]. 1944 [2867]. 1945 [2527, 1664, 2221, 2777, 2535, 748, 680]. 1946 [947]. 1949 [1304]. 1950 [2439, 2730, 1658, 1358, 2872, 878, 815, 510, 303, 747]. 1950s [228]. 1950s-1970s [228]. 1952 [2718]. 1955 [2959]. 1955-1973 [2959]. 1956 [1418, 2137]. 1957 [395]. 1958 [1510]. 1959 [2298]. 1959-2014 [2298]. 1960 [1584, 2735, 2213, 2468, 2201, 2138, 2206, 507]. 1960s [1516, 2554, 619, 2658, 318]. 1964 [296]. 1965 [809]. 1970 [1158, 584, 672, 516]. 1970-1976 [2668]. 1970s [969, 466, 228, 2523, 1960, 304]. 1972 [896]. 1973 [2959]. 1975 [1968, 1241, 1745, 1157]. 1977 [1232]. 1978 [1972, 1233]. 1980 [2139, 1328, 2804, 2285, 311]. 1980s [399]. 1985 [1904]. 1989 [259]. 1990 [2226]. 1992 [255]. 19th [2030, 3150, 3130, 3166].

2 [1198, 866]. 20 [401]. 2000 [2042, 1082, 2277, 64, 215, 213, 315]. 2003 [2146]. 2005 [77]. 2007 [167]. 2008 [166, 626, 245, 953, 1431]. 2009 [11, 13, 12, 110, 246, 168]. 2010 [474, 473, 705, 349, 1959]. 2011 [707, 708, 847, 706, 846, 625, 1892, 1432, 857]. 2012 [1028, 1378]. 2013 [1393, 1255, 3218, 2506]. 2014 [2486, 2218, 2298]. 2015 [1928, 2173, 2008, 2806]. 2018 [2771, 2865]. 21 [593]. 250th [1648].

3033 [210]. 3033-3873 [210]. 33 [1798]. 3873 [210]. 3rd [3276].

4 [2920].

50 [396].

700 [2629]. 700-1500 [2629].

800-1700 [2914].

99 [2810].

= [3247].

Âge [1470, 36, 369, 1553, 988, 1289, 3250]. Être [1632].

A. [2045, 2946, 13, 1862, 1428, 1936, 2796]. à/Zu [593]. Aaron [2584, 1818, 1385]. Aaserud [1975, 1741]. Abandoned [2049, 3173]. Abandoning [840]. Abbate [1148]. Abeles [2281]. Abena [1687]. Abendstern [543]. abenteurliche [2453]. Abi [1238]. Abi-Rached [1238]. Abich [2680]. Abilities [1483]. Abolition [916]. Abortive [2680]. Above [1185]. Abraham [2372]. Abrahamic [263]. Abrupt [2156]. Abteilung [1145]. Abuses [1792]. Abusing [2311, 1011]. Abyss [737]. Academia [136]. Academiae [1802]. Academic [2032, 522, 1269, 222, 2574, 1561, 2231, 144, 2195]. Academically [2234]. Academician [2211]. Academics [2472, 2159]. Academies [2383].

Academisch [2777]. Academy [1563, 2535, 965]. Accademia [1367, 270, 1802, 2719, 3020]. accademiche [2353]. Accelerator [88]. Accident [111]. accidente [2567]. Accommodation [1302]. Accompanying [1698, 3252]. Accordance [1320]. according [2632]. Accountability [2550]. Accumulation [937]. Accuracies [2848]. Accuracy [1986]. Accurata [2901]. Accurate [2901]. Achievements [2298, 520]. Achinstein [1954, 1379, 833]. Acid [3167, 1581]. Ackerberg [64]. Ackerberg-Hastings [64]. Ackermann [1634]. Acknowledging [2615]. Acolytes [1007]. Acoustical [583]. Acoustics [1941, 209]. acque [784]. Across [1192, 1425, 1441, 3042, 1189, 3258, 3074, 2786, 3160, 97, 2957, 1747, 3188, 1434]. Act [1565, 969, 3170, 2234]. Acteurs [2295]. Action [78, 2727, 2138, 2450, 1170]. actions [3194, 1108]. Activism [840]. Activist [2368]. Actor [2761]. Actors [3183]. Acts [2166]. Actus [1562]. Aczel [2300]. Adam [1698, 22, 1403, 1793, 2742, 1207, 1366, 1872, 491]. Adamson [2701]. Adaptation [1910, 1833, 830]. Adapting [2480]. Adaptive [219]. Adelene [1300]. Adelheid [1486]. Adkins [1407]. Admiralty [1667]. Adolf [2121, 2680]. Adoption [830]. Adrenaline [1187]. Adrian [1453, 1140, 1697, 1736, 2467]. Adriana [1495]. Adrien [1210]. Adrienne [2144, 2923]. Advances [1038]. Advancing [1483, 1119]. Adventures [912, 2225]. Advertising [803]. Advisory [92]. Æmula [349]. Aerodynamics [960]. Aerofoil [960]. Aeronautics [1431, 3054]. Aerospace [1013]. Aesthetic [551]. Aesthetics [2007, 2038, 96, 2582, 909]. Afar [2905]. Affair [90, 952]. Affairs [1483, 2307, 730]. Affäre [740]. Affe [2280]. Affect [144, 611]. Afresh [2083]. Africa [2127, 2218, 2617, 3279, 2945, 2425, 807, 1687, 500, 747]. African [1573, 951, 3078, 7]. Africans [955]. Afrika [2944]. After [2992, 2907, 826, 1971, 238, 443, 2894, 1533, 2387, 3063, 2892, 1591, 1833, 3198, 1651, 2761, 3022, 1659, 2063, 2459, 2141]. Afterlives [772]. Aftermath [2387, 469]. Again [765, 2083]. against [1943, 2867, 2333]. Agamemnon [2416]. Agassi [1800]. Age [737, 1263, 1047, 391, 2814, 2529, 982, 1970, 1577, 3037, 1988, 1248, 469, 1081, 321, 803, 2223, 49, 1351, 1943, 756, 994, 798, 1548, 3255, 2577, 1556, 1209, 250, 433, 871, 584, 559, 1153, 2932, 913, 275, 501, 766, 1751, 3160, 2206, 2927, 2954, 1769, 216, 1728, 3286, 741, 90, 1570, 251, 2708, 693, 2192, 2409, 2016, 3267, 3085, 3163, 3287, 3224, 1425, 2461, 3145, 1534, 3271, 3015, 3218, 1718, 1909, 1868, 2223, 2531, 1376, 447, 1199, 57, 2587, 2044, 1703, 502, 685, 605, 279, 143, 630, 453, 111]. Agency [2769, 1992]. Agenda [843]. Agendas [2617]. Agent [1236]. Agentenmodelle [23]. Agents [1402, 3122]. Ages [262, 547, 1267, 1124, 2090, 1183, 3156, 2990, 3264, 3210, 1625, 2061, 1464, 2501, 1692, 1122, 726, 189, 985]. Aggression [1442, 3080]. Aging [1893]. Agnese [2505]. Agonistes [1798]. Agostino [2697, 2423]. Agricultural [290, 2736]. Agriculture [82, 1103, 2518]. Agustí [2369]. Ahead [1326]. Ahuja [2214]. Aid [227]. AIDS [2065, 957, 2961, 1982, 3002, 1433]. Ailing [878]. Aim [101]. Ainley [1328].

Air [2076, 599]. Air-Pump [2076]. Aires [878]. Airpower [2812]. Airs [795, 1681, 3245]. Airship [1228]. Airy [698]. Aitor [2213]. Akademie [1562, 2810, 309, 953]. Akademien [2591]. al [2257, 2564, 787, 495, 1340, 2430, 267, 3053, 3158, 1793, 1626, 1042]. al-buhran [1042]. al-Dīn [495]. al-Ghāfiqī [1793]. Al-Ghazāl [267]. al-Haytham [2564]. al-muntaha [2257]. al-Nafis [1340, 3053]. al-Qaeda [3158, 1626]. Alain [2330]. Alan [2952, 308, 1677, 1500, 1377, 2891, 2806, 1543, 2919, 2302, 2516, 3103, 1468, 2806]. Alarming [102]. Alaska [907]. Alba [1461]. Albert [1507, 2885, 2878, 578, 1507, 2450, 1623]. Alberti [2246]. Alberto [1084, 183, 2645, 2713]. Alberts [1906]. Albertson [1794]. Albrecht [370]. Albritton [1568]. Alchemical [2340, 1552, 1865, 3117]. alchemiczne [1865]. Alchemie [3058, 1212, 1631, 2257]. Alchemy [482, 1942, 415, 553, 189, 2011, 782, 414, 416, 413, 3016, 1884, 2106, 375, 3058, 2472, 1400, 1272, 781]. alchimie [1289, 2184]. Aldon [2586]. Aldrich [64, 2316]. Aldrovandi [2845]. Aleh [1333]. Alejandra [2872]. Aleksandr [1919]. Aleksandra [2569]. Alertz [39, 1795]. Alessandro [51, 2176]. Alesse [1338]. Alex [1495, 2044]. Alexa [2778]. Alexander [1880, 2510, 2578, 1067, 2432, 398, 743, 1077, 2039, 2386, 2894, 2870, 1402, 1132, 1214, 2773, 2494, 2481, 2931, 1716, 1297, 1219, 2626, 743]. Alexandra [228, 1514, 1218]. Alexandre [2023, 2289]. Alexandrescu [1130]. Alexandria [2329]. Alexandrian [551]. Alexis [1900]. Alf [1232]. alfabeti [595]. Alferov [601]. Alfred [1824, 1370, 2581, 2, 1370, 1086, 2041, 1871, 3123]. Algebra [2628, 968, 549]. Algebraic [1747, 3258]. algèbre [1126]. Alger [1888]. Algorithms [2829]. Alice [1432, 1867, 2437, 1358]. Alicia [2659]. Alina [1047]. Alisa [2952, 2175]. Alison [1600, 1059, 1328, 2404, 1660]. Alistair [497]. Alive [2713]. Alix [1866]. Alkhateeb [1549]. alla [1641, 1109, 3203, 727, 1197]. Allan [2803, 2408, 1606]. Alle [3039, 378, 1357]. allegorischen [2257]. Allemagne [371]. Allen [1135, 11]. Allen-Emerson [1135]. Allergy [2054]. Alley [77]. Allies [2131, 1195]. Allure [2890, 2009]. Almost [1605, 3051]. Alois [2035]. Alphabet [1183]. Alpine [2549]. Alteration [2156]. Altered [583]. Alternative [1753, 3253]. Althoff [1228]. Altitude [592]. Altmann [1136]. Aluminum [3169, 1599]. Alva [2514]. Alyssa [2746]. am [1561, 3105, 1310, 288, 1365]. Amanuenses [28]. Amato [2198]. Amber [2514]. Ambition [1850, 1057]. Ambitions [843]. Ambivalence [2890]. Amedeo [2353, 2353]. America [215, 1271, 196, 2598, 286, 920, 83, 1306, 572, 2746, 2138, 402, 1513, 2005, 506, 1217, 1935, 1750, 898, 751, 90, 2745, 69, 425, 445, 2872, 567, 2736, 375, 1026, 1586, 2992, 3237, 102, 1102, 1021, 1104, 1077, 308, 3255, 1191, 766, 410, 3095, 1103, 1105, 3018, 306, 1765, 1849, 2196, 3300, 1514, 1242, 1837, 2802, 1659, 1516, 2462, 2351, 1239, 1085, 2808, 1834, 2518, 1147, 2729, 896, 1889, 2648, 1728, 1752, 2288, 2537, 2542, 1825, 2375, 1313, 1011, 744, 92, 398, 858, 293, 515, 66, 390, 933, 455, 291]. American [71, 2191, 1177, 821, 850, 2530, 1572, 1119, 2063, 2142, 2463, 679, 1214, 1505, 684, 1231, 676, 134, 1663, 2141, 2296, 1229, 1496, 2536, 854, 2518, 1512, 1366, 289, 1488, 2583, 1350, 892, 2954, 1079, 2121, 951, 516, 295, 2374, 1878, 2718, 1510, 3219, 2469, 1593, 1157, 2097, 1088, 2189, 820, 1740, 2048, 64, 2816, 2806, 146, 1989, 2526, 1141, 3004, 840, 1887, 3081, 2910, 762, 103, 160, 619, 3293, 2911, 3121, 105, 1525, 1613, 2951, 737, 816, 939]. Americanarum [3081, 1350]. Americanist [1460]. Americanization [1758, 149, 80, 3269]. Americans [453]. Americas [44, 1219]. Amerika [947]. Amerindian [9]. Amherst [1576, 1677]. Amicorum [114]. Amir [2300, 2432]. Amit [1897, 1759]. among [974, 1128]. Amos [1712, 93, 3291, 2285]. Amparo [1983]. Ampère [2274]. Amphetamine [226]. Amphibious [1843]. Amster [1418]. Amsterdam [1375, 3131]. Amsterdamska [110]. Amy [1545, 456, 64, 1505]. Ana [2195]. Anagnostou [1475]. Analogy [2841, 2577]. analyse [564, 1459, 3036, 1561]. Analysis [1515, 1726, 2848, 2784, 2446, 2194, 2678, 2949, 256, 2828, 2546, 3036, 3257]. Analytical [2210]. Anarchism [1094]. Anatole [2639]. Anatomical [1788, 1270]. Anatomie [1477]. Anatomies [2730, 2646]. Anatomist [656]. Anatomists [1949, 1883]. Anatomy [783, 1788, 2632, 871, 1944, 1046, 2927, 2034, 2788, 1883, 2929, 797, 717]. Anaxagoras [1465]. Ancestors [294]. Ancestry [411]. Anciens [2014, 2072]. Ancient [482, 2059, 1848, 2841, 2091, 808, 1790, 2628, 2318, 2178, 2753, 924, 2923, 1456, 2169, 489, 913, 926, 2781, 1785, 483, 32, 429, 1458, 1937, 1986, 698, 485, 925, 1936, 1404, 2626, 2421, 2995, 2562, 1196, 2546, 2016]. Andean [2112, 2385]. Anders [1806]. Andersen [349]. Anderson [1419, 1647, 2093, 1506, 1948]. André [1817]. Andrea [1160, 2657, 1467, 1686, 1075]. Andreas [3249, 2632, 1622, 723, 2212, 2444, 2927, 2843, 2654, 2183]. Andreas-Holger [2444]. Andrei [400]. Andretta [1477]. Andrew [1954, 946, 2508, 2049, 2599, 1930, 1054, 610, 1933, 1946, 1751, 2371, 2702, 1809, 1748, 1749]. Andrews [2294]. Androids [1486]. Anduaga [2213]. Andy [2780]. Anew [1879]. Anfängen [2522]. Ángel [1479]. Angela [2252, 1518, 2332, 1490]. Angeli [1578]. Angelica [2566]. Angelis [45]. Angelo [2298, 1907]. Angewohnheiten [3008, 1361]. Anglais [785]. Anglicanus [2987, 1342]. Anglo [516, 1830]. Anglo-German [1830]. Angola [2268]. Angus [1581, 3167, 2452]. Animal [409, 2193, 172, 979, 2536, 4, 3099, 2794, 2284, 1638, 2250, 2113, 1066, 152]. Animalibus [2258]. Animalium [866, 866]. Animals [2198, 1949, 1959, 2934, 4, 772, 1549, 3223, 3107, 1403, 1931, 365]. animaux [350]. Animi [932]. Anindita [1585]. Anita [1569, 1949]. Anja [45, 2882]. Anja-Silvia [45]. Anke [1744, 1552, 1955]. Ann [1572, 1005, 1302, 1681, 1545, 1625, 2406]. Anna [656, 1476, 2249, 2047, 2345, 1489]. Anne [1160, 2295, 2095, 2518, 2064, 1683, 1968, 1692, 2572, 1458, 1937, 1421, 2348, 2856]. Anne-Lise [1683]. Annette [1005, 1592, 1069, 2279]. Anniversary [1648, 3207]. Annotated [186, 1370, 1622, 1432, 1827, 647, 3249, 3000, 2969, 3123]. Annual [704, 1028, 1532, 1859, 2165, 2482, 2771]. Anonymity [2661]. Antaki [2499].

Antarctic [2174, 741, 2467]. Antarctica [2142]. Antarktisreisen [2290]. Antebellum [1307]. Anthologie [3008, 1361]. Anthology [2174, 3008, 2840]. Anthony [24, 45, 886, 1964, 2317, 2740, 2055, 2375]. Anthrax [1369, 3112, 545]. Anthropocene [2227]. Anthropologie [2522, 2280, 380]. Anthropologies [89]. Anthropologists [2362]. Anthropology [3035, 2689, 1575, 1600, 2074, 334, 881, 3166, 1380, 1093, 1358, 791, 3028, 3040, 3079, 1355, 1866, 2664, 181, 996]. Anti [504]. Anti-Darwinism [504]. Antibodies [2301, 2055]. Antibody [2055]. Antievolution [840, 1366, 3219]. Antike [486]. Antikythera [2255, 2099, 2626]. Antiquarian [1168]. Antiquarianism [49, 2096]. Antique [2629, 571, 875]. Antiquity [262, 1942, 1868, 1464, 177, 1682, 2476, 2927, 1816, 1536, 1183, 977, 2168, 1679, 3232, 3193, 3196, 2494, 721, 1292]. Antitoxins [2055]. Antivaccine [2031]. Antoine [2428]. Antoine-Mahut [2428]. Anton [2938]. Antonella [1357, 2342]. Antoni [1842]. Antonio [1983, 1578, 2937, 1939, 3172]. Anxieties [1883]. Anxiety [90]. Anxious [1614]. Anya [2351]. Anything [2759]. Ape [59, 582]. Apel [2189]. Apes [294]. Aphorisms [2101]. Aping [834]. Apollo [1903, 2063, 2894]. Apollonius [476]. Apotheosis [185]. app [1575]. Appendice [1417, 3118]. Apple [1031]. Applications [304]. Applied [768, 2953, 766, 769, 767]. Appliquées [158]. Appreciation [2310]. Approach [1113, 1843, 2920, 2317]. Approaches [718, 2978, 2781]. Appropriating [2402]. Appropriation [817, 2809]. Appropriazione [784]. apps [1569, 1554, 1547, 369, 1589]. après [40]. April [2651, 288]. Aquarium [2837, 2941]. Arab [776, 2889]. Arabia [3207, 1648, 1709, 3230]. Arabian [3207, 1648]. Arabic [2499, 1198, 2426, 2696, 1042, 2101, 2843, 2423, 476, 2478, 2477, 2236]. Arabic-English [2101]. Arachnologist [1291]. Aragno [1556]. Arago [2030]. Aragon [1200]. Arapostathis [1583]. Aratus [1466]. Araujo [2438]. Arbeit [1087]. Arbeiten [1145]. Arbeitsforschung [1087]. Arbeitsphysiologie [1087]. Arbor [1545]. Arboretum [661]. Arch [929]. Arch-Conjuror [929]. Archaeological [1460]. Archaeology [1384, 1460]. Archibald [2649]. Archimede [1197]. Archimedeo [1197]. Archimedes [27, 70]. Architectural [2443]. Architecture [1047, 2447, 361, 193, 651]. Architette [1184]. Archiv [1188]. Archive [2514, 1668, 1772, 1768, 2091]. Archives [1026, 709, 1773, 1448, 1769, 1770, 1771, 3166, 2407, 1575]. Archiving [974]. archivio [1109]. Arctic [1730, 2371, 1228, 907, 1602, 3153, 3277]. Ardinghelli [966]. Ardrey [411]. Area [1109, 2402]. Areson [57]. Argentina [582, 1983, 1495]. Argentine [2074]. Argument [2094, 1046, 252, 2046]. Arguments [3109, 1378]. Ari [1839]. Arianna [2787]. Ariel [22]. Arielle [2631]. Ariew [1880, 1128]. aristocratici [1109]. Aristocrats [1109]. Aristoteles [866, 3211, 1405, 866]. Aristotelian [780, 2420, 2498, 1121]. Aristotelianism [1467]. aristotélicienne [3211, 1405]. Aristotelis [2103, 2258]. Aristotle [550, 641, 2254, 365, 1696, 27, 2319, 2151, 3272, 1870, 115]. Arithmetic [443, 1590, 1198, 3066]. arithmetica [1940]. arithmeticae [443]. Ark [2837, 554, 1668]. Arm [2469]. Armada [139]. Armand [1538, 2254]. armario [787]. Armed [1663]. Arming [3062, 1661]. Arms [1241, 1265]. Army [888, 1451]. Arnaud [2018]. Arnauld [2795]. Arnold [450, 2357, 3049, 1359]. Arose [2761]. Arpan [1736]. arpenteur [190]. Arrhenius [1247]. Arrow [1754, 3241]. Arsdall [1458]. Arsenic [965, 1116]. Art [1942, 2776, 446, 2091, 172, 49, 1640, 42, 138, 936, 2726, 1194, 2013, 2267, 2204, 319, 1504, 1411, 3086, 917, 2927, 2270, 776, 2633, 769, 299, 873, 2642, 686, 1377, 1345, 323, 3068, 3097, 3103, 1942, 2698, 1796, 2010, 995, 379]. arte [2502, 2110]. Arteries [2496]. Arthur [1295, 2133, 2835, 1827, 185, 2933]. arti [784]. Article [2574, 3141]. Articulating [2664]. Artifact [808]. Artifice [323]. Artificial [520, 768, 2890, 611]. artificielles [1123]. Artisan [2979, 1343]. Artisan/practitioners [2979, 1343]. Artisanal [842, 2641]. Artisans [1486]. Artistic [3304, 1278]. Arts [1790, 1209, 437, 2641, 2021, 1056, 1125]. Artstein [1677]. Arun [2090, 1473]. Ärzte [2292]. Asare [1687]. Ascendancy [481]. Ash [2700]. Ashgate [1569, 1581, 1558, 1567]. Ashkenaz [2424]. Asia [748, 1539, 830, 720, 1673, 1473, 2423]. Asian [1539, 1549, 3155]. Aside [1943]. Asklepios [117]. Aspects [1811]. Aspray [1330]. Asprem [1956]. Assemblage [1247]. Assembling [2930]. Assembly [1191]. Assessing [1615]. assistance [369]. Associate [1590]. Association [1119, 2808, 1766]. Associations [132]. Astbury [1985]. Asteroid [2515]. Asteroids [891]. ästhetischer [438]. Astounding [1077]. Astral [2783]. Astrid [349, 1684]. Astroculture [2894, 1077]. Astrolabes [184]. Astrologers [2506]. astrologiae [2183]. astrologicis [38]. Astrologie [2697, 2924, 543, 2183, 1470, 40, 120]. Astrology [796, 1347, 1200, 2105, 3086, 2898, 1120, 1639, 1345, 3032, 3239]. Astronauts [1436, 2889, 3213]. astronome [1184]. Astronomer [1563, 2111, 3077, 1133]. Astronomers [29]. astronomes [171, 2417]. Astronomia [992, 1877]. Astronomical [1726, 2737, 1234, 2099, 2695, 1466, 1375, 1122, 647, 2421, 2977, 3257, 3131]. Astronomie [1412, 2417]. astronomischen [1562, 2977]. Astronomy [980, 377, 1020, 1119, 1715, 1699, 1429, 2866, 2846, 1002, 489, 1546, 2153, 1937, 2756, 2782, 3046, 3276, 3283, 3294, 1182, 1890, 1155, 2169, 1298, 2565, 1465, 659, 302, 451]. Astrophysical [2959]. Astrophysics [302, 1604, 3039, 2030]. Astrophysik [2517]. Asymmetry [2579]. Athanase [2720]. Athanasius [1292]. Athene [53]. Atlantic [2710, 2112, 1573, 1205, 3078, 2507, 2635]. Atlas [2505, 723, 1089]. Atlases [2677]. Atmosphere [592, 2537, 297]. Atmospheric [2128, 1372, 3128, 218]. Atom [3200, 690, 1975, 2132, 3011, 1741, 207]. Atombilder [401]. atome [153, 153]. Atomenergie [2226]. Atomic [250, 2132, 3011, 1518, 1376, 3096, 453, 1823, 752, 3049, 2605, 1438, 3015, 892]. Atomism [119]. Atomphysiker [3049, 1359]. Atoms [401, 165, 690]. Atomstaat [2226]. Atomzeitalter [1310]. atoombom [2451]. Attachment [1313]. Attitudes [293]. attraction [1551, 3124]. aube [649]. Aubin [2131].

Aubry [2374]. Audiences [2159]. audio [3111]. Audiovisual [2352]. Audra [1242]. Aufbruch [2027]. Aufklärung [370, 281]. Aufrecht [296]. Aufzeichnung [635]. August [11, 473, 2270, 2868]. Auguste [1352, 2940, 3092]. Augustine [3205, 1624]. Augusto [1654]. Augustus [470]. aujourd’hui [26]. aura [2008]. Aurality [711]. Ausbildung [811]. Ausécache [2500]. Ausgabe [2651]. ausgehenden [2925]. Ausgewählte [2651]. Ausgleich [745]. ausrechnen [586]. Außereuropa [46]. Außeruniversitäre [2811]. Ausstellung [1955]. Ausstrahlung [1475]. Australia [2945, 138]. autel [2419]. Authenticity [2525, 103]. Authier [1817]. Author [2086]. Authority [324, 1943, 2942, 1200, 1319, 410, 2927, 956, 1491, 2905, 1421, 698, 2708, 2782, 3045]. Auto [1867]. Autobiographical [2060]. Autobiographies [2045]. Autobiography [433]. Autobiologies [1900]. Automaton [712, 1277]. Automatons [2296]. Automotive [1374, 3012]. Autonomous [2318]. Autonomy [2550, 1302]. Autopsy [1270]. Autour [992, 1470, 1352, 3092]. aux [1481, 587]. av [478]. avant [786]. Aventures [1459]. Averroe [2180]. Avery [1229]. aveugles [1551, 3124]. Avi [1413, 1729]. Avicenna [2843]. Aviva [2789]. Avogadro [2353]. Avramescu [497]. Awareness [2523]. Axel [2212]. Axworthy [2332]. Aya [2595]. Ayako [1365]. Ayurveda [2367, 1457]. ayyam [1042]. Azélina [1792]. Azeredo [2268]. Azzolini [1120]. Azzouni [593].

B [2586, 2512, 1726, 1341, 2852, 1110, 1070, 724, 1535, 2945, 1791, 19, 1032, 2203, 237, 2628, 2808, 2250, 2868, 1187, 1064, 2125, 1592, 2589, 2315, 2743, 1503, 1888, 1437, 228]. B.C [1870]. B.C.E [1467]. Baack [1709]. Babbage [2573]. Babel [129, 1118, 1971, 1503]. Babies [2452]. Babylonian [552, 2756]. Bach [1575, 53]. Bachelard [608]. Back [1451, 762, 1604, 2999, 957]. Background [2333, 116, 993]. Backhouse [1905]. Backlands [2895]. Backwards [694]. Bacon [1800, 1995, 2305, 1383, 1522, 1987, 872]. Bacteria [2286]. Bacteriology [2994, 1356, 387]. Bad [3007, 3008, 1397]. Badash [473]. Bæk [1648]. Baggott [1211]. Baghdad [776]. Bagliani [2697, 2423]. Baichun [736]. Bailey [1550]. bajo [1939]. Baker [2346]. Bakker [2180, 2916]. Bakterien [940]. Bakteriologie [940, 2449, 503]. Bala [2090, 1491, 1473, 20]. Balam [773]. Balance [1637, 2098, 3154]. Baldin [2513]. Baldwin [1969]. Ball [1895, 2009]. Ballistics [1476]. Balloon [2429, 2576]. Balmer [1983, 1149]. Balsoy [1354]. Baltimore [1535]. bambini [2324]. Bandinelli [1490]. Baneke [440]. Banerjee [1585, 1736]. Bang [1118]. Bankes [1962]. Banking [1837]. Banks [1704, 2574, 3189]. Baptist [2340]. Baptiste [561, 2227]. Baracca [2298, 1907]. Barak [1296]. Barbara [2799, 593, 2007, 2929, 1298, 2563]. Barbier [1032]. Barbieri [1462]. Barbour [1484, 41, 1202]. Barcelona [2369]. Bard [1545]. Barefoot [1151]. Barell [1839]. Bargheer [2860]. Bark [2112]. Barnabas [2425]. Barnard [2805]. Barnes [1112]. Barnum [1493, 567]. Baron [2526]. Baroque [2265, 1284]. Baroulkos [2331]. Barrels [2716]. Barres [1205]. Barrett [1517]. Barrier [569]. Barro [1668]. barroco [2110]. Barry [2605]. Bart [1796, 2021]. Barthel [2169]. Bartholin [1944]. Barton [2483, 2942]. Bartos [2098]. Bartusiak [2049]. Based [1937, 2284, 2761, 1990]. Bashford [1660]. Basic [2953]. Basil [1576]. Basis [2399]. Bateson [2740]. Batsaki [2639]. Battista [1943, 2505, 2850, 2505]. Battle [139]. Battled [1899]. Battlefields [173]. Baudhayana [2419]. Baudin [1136]. Baudin-Expedition [1136]. Bauer [2095]. Bauhaus [513]. Baumeister [868]. Baumgärtner [2505]. Baur [1721, 3227]. Baxter [2793]. Bay [3135, 356, 1566, 1297]. Baye [637]. Be [2456, 481]. Beach [1554]. Beagle [3192, 1419, 55, 2997]. Beam [2299]. Beanbag [596]. Beastly [2198]. Beasts [979, 1363, 2026, 3088]. Beat [1942]. Beattie [2356]. Beatty [1649]. Beaulieu [1160]. beautés [564]. Beautiful [2221]. Beauty [1737, 275, 3143, 3273]. Became [360, 6, 2594, 2049, 2122, 1079, 358, 3173]. Beck [1894, 288, 1385]. Becker [2742, 1298, 967]. Beckmann [2703]. Become [2757]. Becoming [636, 155]. Bede [724]. Bedrohte [719]. Bedwetting [102]. Beeckman [1293, 2760]. Beelden [2887]. Beeley [1049]. Beer [2437]. Beery [2808]. Bees [2208]. Beethoven [2352]. Before [1971, 2976, 630, 1465, 2852, 2503, 2115, 2418]. Begegnungen [46]. Beginning [19]. Beginnings [894, 406, 2691]. Begriffe [543]. Begriffsuntersuchung [486]. Behavior [1442, 228, 3080]. Behavioral [102]. Behaviorism [1767]. Behemoth [2080]. Behind [3152, 2863, 1607, 897]. Being [1843, 2620, 1248, 1787, 955, 98]. Beings [1139]. Beispiel [1310]. Beispielen [2124]. Beitraäge [1881]. Beitrage [1892, 53]. Bekennen [2226]. Belehrung [1498]. Belgique [2008]. Belgisch [2442]. Belief [2256]. Beliefs [1113]. believer [3280, 1724]. Belknap [2525]. Bell [1600, 1663, 2167, 2034, 2028]. Belle [1731, 3282]. Belles [1551]. Bellorini [2260]. Belon [779]. Below [94]. Beltrán [1752]. Ben [1839, 2887, 1362, 2837, 1879]. Ben-Gurion [1839]. Bench [162]. Bender [2536]. Benedetti [993]. benedica [2355]. Benjamin [24, 1192, 1733, 2186, 2543, 2858, 1791, 1319, 2597, 1885, 2096, 2582]. Bennett [1408, 1727]. Benoît [2561]. Bensaude [1225, 2585]. Bensaude-Vincent [1225, 2585]. Bentley [732]. Benutzer [1574, 3150]. Beobachten [1136]. Beobachtung [1807]. Beranek [514]. Berco [2265]. Beretta [1554, 2414, 2100, 2176, 1041]. Bergen [2871, 1252]. Berger [2633]. Berggren [2425]. Bergson [2291]. Bergwik [2326]. Beriberi [1297]. Berkel [953, 2777, 2021]. Berkowitz [2034, 2648]. Berlandier [1223]. Berlin [1590, 1507, 1562, 1413, 1597, 1087, 1604, 1508, 2363, 2773, 3013, 130]. Berliner [811, 23]. Berlivet [1463]. Bernadette [1225, 2585]. Bernard [1653, 1419, 1534, 2648, 1137, 1122, 1577, 1727, 2809, 2029, 909]. Bernardino [1201, 1873]. Bernard’s [558]. Bernd [593, 1472]. Bernhard [1361]. Bernhardt [1958, 2455]. Berni [1203]. Bernoulli [42]. Berns [1946]. Berquist [1708]. Bertillon [1582, 2998]. Bertoloni [2776]. Bertrand [813]. Bertsch [637]. Bertucci [2641]. besonders [486]. Bestandsaufnahme [2292, 589]. Bestiary [554]. Bestimmungen [806]. Bet [1757, 3216]. Bethe [1071]. Bethlehem [2169]. Betrayal [174]. Better [1069, 1390, 995]. Bettering [1483]. Bettina [1402, 2949]. Between [1176, 654, 1102, 468, 432, 2201, 340, 2396, 1805, 1251, 622, 2521, 3131, 117, 2879, 1047, 2329, 966, 2990, 1527, 2004, 2524, 1362, 1982, 808, 1782, 768, 2071, 2577, 2630, 1319, 1302, 2159, 458, 3083, 1167, 1168, 2954, 1201, 934, 2937, 2854, 830, 2492, 2018, 1375, 644, 524]. Betweens [283]. Bever [2487]. Bewell [2516]. Beyond [2837, 22, 825, 59, 2968, 912, 971, 3282, 1105, 1270, 3295, 2730, 1731, 2473, 3108, 1590, 1834, 1688, 1910, 2435, 1118, 2619, 327, 228, 818]. Beyond.} [2927]. Bezza [1471]. Biagioli [1033]. Bianchetti [2178]. Bianchini [931]. Biard [34]. bibl [1598, 1588, 1590, 1569, 1562, 1568, 1594, 1586, 1578, 1543, 1575, 1609, 1566, 1580, 1537, 1583, 1596, 1565, 1605, 1576, 1606, 1558, 1677, 1597, 1545, 1552, 1539, 1555, 1599, 1546, 1573, 1567, 1561, 1559, 1607, 1601, 1551, 1571, 1314, 1536, 1592, 1542, 1549, 1584, 369, 1593, 1540, 1570, 1564, 1608, 1574, 1595, 1603, 1582, 1544]. Bible [1930, 1946]. Biblical [2708, 277]. Bibliografía [351]. Bibliography [1, 333, 616, 1164, 1446, 1763, 2070, 2384, 1432, 1024, 3000, 2969, 2932]. Biblioteca [1795, 1560]. Bibliothekar [124]. bibls [1541, 1600, 1547, 1563, 1581, 1589]. Bicycle [1589]. Bidwell [2548]. Biehl [2135]. Biener [1876]. Bier [2667]. Biess [1833]. Big [1515, 2680, 1438, 1118, 1851, 1434, 1990, 3074, 3096]. Bigfoot [76]. Bigg [401]. Bihari [2367]. Bild [2024]. Bilder [2978, 1398]. bildnerische [1814]. Bildphysiologie [3210, 1691]. Bildschirmtechnik [2124]. Bildwerke [1136]. Bildwissenschaftliche [1398, 2978]. Bill [2028]. Billings [502]. Billion [31]. Bind [2740]. Binnema [1566]. Bio [340]. Biocultural [1038]. Biodefence [1229]. Biodiversity [2951]. Bioeconomy [1835]. Bioethics [1312, 2544, 2543]. biografico [1184]. Biografie [2273, 3049, 1359]. biografischen [61]. Biographical [1355, 29, 3035]. Biography [2662, 1644, 3075, 1750, 160, 1867, 3237, 3049, 1980, 1974, 1216, 673, 249, 819, 480]. bioinformatics [3113, 1439]. Bioinsecurities [2214]. Biologen [2453]. biologia [1490]. Biological [71, 2920, 212, 1391, 1241, 1229, 1190, 2294, 959, 2765, 1149, 2692]. Biologics [3122, 1402]. biologischen [2522, 67]. Biologist [2043]. biologistes [669]. Biology [2920, 2368, 1896, 2798, 2457, 3251, 2670, 2656, 1253, 2886, 2151, 1366, 449, 1016, 2302, 516, 2147, 1686, 791, 233, 1008, 3208, 3219, 3302, 1598, 1082, 1110, 2252, 1710, 254, 154, 828, 428, 1008]. Biomedical [142, 958]. Biomedicine [340, 336, 20, 1263]. biomedizinischen [441]. Biopolitical [152]. Biopolitics [2943, 185]. Biosocial [2206]. Biotech [1832, 894]. Biotechnologies [1159]. Biotechnology [2900, 2298, 530, 1914]. Birds [2728, 2860, 1146, 465, 1146, 2934, 1003]. Birdsong [2876]. Birgit [2002]. Birkenmajer [1842]. Birkhead [2341]. Birth [2586, 1661, 2947, 1909, 1625, 1010, 2672, 2574, 2935, 430, 1088, 2455, 2044, 2129, 1075, 3062, 3156]. Birthing [145]. bis [947, 2522, 811, 1597, 2504, 543, 3111, 2287]. Bishop [1708, 3231]. Bitter [3279, 1687]. Bix [1505]. Bizri [1198]. Bjerknes [2128, 2871]. Bjørn [1883]. Black [3173, 1094, 606, 1342, 2375, 2988, 2049, 1647]. Blacklock [2873]. Blackness [2710, 871]. Blackouts [455]. Blanchon [1419]. Blanckaert [1493]. Blancke [1784]. Blanco [1654]. Blaserna [2276]. Blay [1538]. błedów [1865]. Bleeding [715]. Bleichmar [43, 1050]. Blitz [811]. Bloch [1121]. Blockade [1666]. Blond [2703]. Blondel [651]. Blood [1837, 1067, 2095, 3196, 66, 1264, 2375, 1679]. Bloomberg [2373]. Bloomsbury [2726]. Blowing [2073]. Blue [1013]. Bluetooth [259]. Blum [1339]. Blume [2594, 1636]. Blur [2613]. Blyth [497]. Bo [2445]. Boantza [1558]. Board [2234]. Boas [1730, 13, 2115, 3277]. Bod [1932, 1682]. Boddice [2717]. Bode [2515]. Bodies [1004, 979, 1638, 454, 2152, 253, 2416, 1874, 2189, 682, 1469, 3099, 2730]. Bodily [1340, 3053]. Body [1837, 2814, 1139, 372, 2841, 717, 794, 1482, 2250, 2833, 1681, 96, 1729, 3089, 2927, 1627, 1622, 2265, 804, 2246, 185, 485, 3245, 3249, 3076, 3210, 3288, 2533, 1658, 185]. Boer [2699]. Boerhaave [1056]. Boethius [1940]. Boewe [1213]. Bogdanov [1067]. Bohr [1975, 2132, 1741, 3200, 3011]. Bois [2586, 1580, 3055]. Bois-Reymond [1580, 3055]. Bojana [2624]. Bokulich [2952, 2175]. Bolshevik [1828]. Bolton [1420]. Bolzano [2029]. Bomb [1265]. bomba [2084]. Bomber [217]. Bombing [217, 2955]. Bonah [2882]. Bondio [1782]. Bone [2196, 2202]. Boner [1347, 2509, 1560]. Bones [2528, 1893]. Bonn [1892]. Bonneuil [2227]. Bònoli [1471]. Bont [2730, 2284]. Book [2045, 78, 1308, 2281, 2620, 1598, 878, 1905, 2586, 2693, 2116, 2512, 757, 2879, 1585, 1228, 2956, 186, 597, 1114, 1154, 1515, 2573, 719, 1824, 79, 1424, 2835, 1940, 215, 1182, 941, 1052, 1486, 2466, 305, 71, 737, 1311, 1588, 2665, 297, 482, 795, 1031, 1263, 1964, 431, 1880, 1694, 1649, 2128, 1514, 2054, 521, 716, 1295, 1762, 1761, 2303, 2799, 2859, 570, 1502, 2408, 1877, 443, 1541, 1590, 2490, 1932, 2881, 266, 1296, 2249, 2527, 2839, 209, 1661, 2567, 2795, 2361, 1895, 805, 348, 2030, 2297, 38, 1744, 1653, 2058, 1507, 2947, 645, 2787, 2339, 1242]. Book [783, 2894, 1425, 1890, 1837, 1307, 89, 386, 665, 2275, 2660, 2789, 1185, 2802, 285, 435, 316, 992, 282, 1271, 2742, 196, 668, 1036, 2440, 2728, 2870, 2960, 2744, 1718, 1684, 749, 1903, 497, 1367, 2735, 262, 868, 1047, 885, 1726, 2486, 961, 2491, 494, 2510, 1804, 294, 574, 1419, 2199, 2578, 391, 735, 1569, 2172, 2643, 2848, 213, 1869, 202, 646, 792, 1272, 1650, 2289, 441, 2191, 1069, 2104, 2424, 550, 2329, 2562, 947, 296, 1227, 2226, 2454, 1344, 326, 1708, 950, 1499, 1218, 1554, 1942, 899, 1436, 2631, 2210, 257, 1341, 2499, 861, 1798, 2852]. Book [2110, 2710, 2439, 1755, 1897, 2700, 1562, 2598, 2895, 1697, 273, 690, 1400, 356, 594, 809, 1309, 1568, 2174, 2814, 298, 753, 2119, 780, 796, 153, 635, 2533, 1176, 422, 1115, 1659, 307, 1594, 1975, 362, 687, 2712, 1813, 2320, 214, 1984, 980, 1006, 436, 2790, 1511, 1900, 130, 423, 511, 593, 1981, 2815, 1866, 2193, 1730, 2001, 1225, 1634, 2529, 653, 1198, 498, 1641, 2776, 442, 227, 1441, 2262, 1910, 137, 286, 1360, 2032, 2299, 112, 2640, 2447, 733, 1155, 1193, 2356, 2689, 231, 545, 1135, 1001, 1177, 2783, 2858, 2838, 1333, 2622, 2571, 2926, 254, 229, 730].

Book [1481, 24, 1492, 920, 588, 1159, 1402, 1011, 681, 821, 52, 1139, 1586, 1110, 669, 2182, 2523, 816, 1460, 2064, 2360, 1067, 754, 982, 2047, 2595, 118, 377, 671, 2363, 954, 2810, 225, 450, 883, 1970, 1983, 1578, 1109, 850, 2112, 2514, 210, 385, 1577, 127, 129, 990, 1413, 2599, 2669, 1812, 2007, 1873, 726, 2103, 1038, 564, 1543, 1654, 1863, 2279, 2860, 1930, 207, 2585, 361, 1954, 1346, 788, 2570, 270, 656, 2530, 1222, 923, 1962, 1575, 1783, 2004, 1416, 2266, 942, 2059, 1347, 678, 1304, 2524, 608, 1184, 1698, 58, 2502, 121]. Book [862, 1334, 2813, 446, 522, 1719, 1643, 1969, 1160, 2352, 2925, 2415, 2730, 1048, 2019, 1192, 2727, 938, 1275, 1215, 1414, 394, 1864, 1282, 1070, 2198, 2580, 2650, 204, 502, 744, 880, 592, 1483, 1572, 1788, 995, 1909, 1080, 2050, 2784, 2560, 389, 205, 197, 1516, 2252, 427, 1881, 142, 2324, 403, 88, 1980, 815, 2432, 2179, 372, 2885, 2330, 2943, 576, 194, 774, 1005, 549, 1418, 2373, 565, 654, 1625, 2419, 1600, 2946, 261, 1579, 212, 577, 2143, 1609, 2841, 1566, 2412, 2127, 1949, 1004, 1003, 2368, 74, 1362, 1818, 2051, 2837, 724, 2420, 2188, 1127, 92, 823]. Book [1119, 2328, 2723, 1724, 1689, 2213, 2649, 1901, 2224, 1982, 1896, 1336, 2798, 2317, 986, 1129, 2091, 398, 743, 2557, 2592, 561, 171, 1652, 1081, 1426, 2063, 2142, 1408, 1868, 360, 2220, 1430, 321, 1819, 2588, 797, 2120, 51, 374, 803, 1130, 1470, 1014, 2056, 324, 1268, 77, 2503, 2286, 1646, 388, 1722, 1710, 1054, 553, 1535, 2519, 513, 172, 685, 1474, 2017, 2528, 717, 2020, 2945, 2463, 2322, 83, 300, 152, 1132, 1939, 1128, 1580, 1960, 2446, 3141, 123, 874, 1647, 1236, 1835, 2807, 1240, 1012, 1306, 2061, 1713, 1152, 2579, 404, 679, 2150, 1757, 203, 1207, 572]. Book [2522, 1656, 1501, 61, 677, 220, 520, 2223, 604, 518, 2461, 2346, 1356, 45, 2457, 808, 1223, 1464, 1487, 2177, 2458, 2878, 178, 1791, 641, 280, 563, 1112, 49, 1274, 1780, 568, 1146, 979, 1214, 1485, 893, 177, 22, 1715, 2966, 1624, 2591, 86, 1640, 585, 114, 1068, 2254, 139, 1370, 2532, 709, 1351, 1802, 1269, 2117, 2662, 1032, 1073, 1931, 199, 727, 1339, 1072, 2024, 1427, 2144, 500, 1060, 399, 2442, 191, 2295, 2053, 512, 781, 2731, 312, 750, 1505, 206, 42, 1631, 2338, 56, 573, 659, 1396, 1286, 1459, 2194, 2362, 855, 2749, 664]. Book [639, 2323, 2203, 317, 551, 81, 2699, 2513, 806, 684, 1094, 1977, 1957, 276, 1206, 853, 1699, 2426, 1943, 2276, 2719, 271, 554, 365, 1357, 2531, 154, 1337, 138, 2040, 1017, 378, 940, 657, 872, 756, 1287, 1462, 2269, 73, 638, 508, 1790, 1498, 1123, 1461, 610, 379, 1238, 395, 2211, 222, 488, 1782, 31, 1720, 2170, 2340, 36, 2344, 2746, 2942, 1735, 1429, 352, 632, 1086, 1683, 2952, 2291, 944, 2509, 367, 2180, 2866, 67, 510, 351, 2259, 493, 183, 2052, 1839, 2884, 2801, 432, 692, 1886, 731, 1495, 575, 2953, 824, 132, 259, 994]. Book [1738, 794, 798, 1077, 1189, 2526, 1537, 2108, 2698, 2875, 1231, 1967, 595, 503, 1966, 2707, 349, 869, 813, 1547, 2134, 125, 284, 62, 2114, 72, 1548, 2042, 509, 1010, 1712, 2218, 176, 755, 1338, 1714, 2468, 1280, 2697, 2539, 308, 936, 320, 1737, 1733, 2283, 1583, 140, 1638, 1882, 2014, 2632, 1261, 601, 1136, 2459, 2811, 1968, 2594, 2832, 2846, 1438, 173, 1734, 2936, 116, 721, 122, 2094, 2272, 2873, 91, 1299, 1727, 2726, 39, 629, 1107, 1695, 1480, 1262, 825, 832, 2540, 603, 955, 1241, 2955, 2000, 126, 1055, 2501, 310, 2146, 1644, 2411, 1037, 1074]. Book [1401, 1678, 1739, 833, 314, 2271, 2577, 2663, 676, 1141, 1051, 1908, 2541, 1758, 673, 547, 1131, 2201, 1959, 605, 1563, 856, 835, 2031, 1717, 2214, 120, 865, 1200, 2672, 2892, 2867, 858, 2630, 1904, 184, 303, 249, 1158, 135, 1556, 1633, 2095, 2847, 134, 434, 634, 1209, 1478, 1682, 2003, 2638, 2709, 2853, 133, 68, 523, 1145, 1329, 1822, 2740, 2138, 433, 1226, 1596, 1111, 1150, 2039, 2437, 2800, 1267, 402, 1663, 1591, 2871, 1148, 2414, 327, 2670, 1696, 1009, 2659, 2451, 2887, 2855, 2460, 2656, 871, 1194, 2141, 584, 948, 315, 1953, 2652, 2300]. Book [400, 1002, 2296, 779, 2022, 2805, 571, 1435, 1685, 2651, 2453, 2941, 1210, 1565, 1898, 2505, 2280, 1191, 2227, 541, 2443, 2863, 1642, 1945, 174, 736, 670, 2628, 2318, 2200, 2623, 2126, 2621, 1629, 1229, 1039, 2013, 811, 1049, 932, 1701, 2637, 1628, 1368, 1801, 1283, 2915, 1496, 2629, 1196, 2178, 1092, 1605, 2462, 2737, 998, 2365, 2536, 559, 2171, 946, 1477, 2696, 59, 1153, 817, 945, 1497, 1576, 598, 1581, 2351, 2132, 1373, 2038, 376, 1033, 496, 1285, 1335, 1602, 1482, 2435, 1203, 2263, 1239, 2575, 1781, 636, 739, 1013, 1237, 1085, 84, 2645, 2934, 1933, 293, 2267, 1290]. Book [457, 1273, 886, 1190, 1862, 1664, 2105, 43, 2692, 1509, 1233, 2517, 2186, 2511, 2844, 1606, 1944, 2779, 189, 1817, 1124, 712, 1626, 2932, 876, 30, 200, 2011, 2808, 2874, 1305, 1723, 1834, 1042, 2057, 1736, 44, 179, 637, 1952, 263, 2204, 2882, 1558, 2639, 854, 1035, 2175, 1053, 2494, 548, 2518, 2812, 1244, 642, 2456, 866, 2918, 1912, 1648, 1709, 1821, 587, 2261, 2929, 479, 1658, 2372, 1784, 1786, 131, 354, 714, 924, 1829, 268, 1056, 2508, 526, 2093, 2923, 319, 357, 1147, 1513, 2133, 290, 1302, 1828, 1058, 1397, 1456, 188, 2720, 1374, 2864, 27, 1071]. Book [1946, 252, 1504, 2590, 371, 1636, 2169, 607, 827, 454, 1518, 1442, 555, 1354, 1677, 2250, 606, 881, 1457, 1597, 1941, 560, 1050, 2027, 2325, 515, 452, 2369, 505, 1934, 35, 274, 784, 1197, 2427, 2049, 2671, 2497, 2464, 2893, 2504, 1632, 1062, 2951, 1833, 889, 2452, 383, 2066, 2060, 2868, 2886, 2958, 2739, 2343, 2833, 814, 2967, 2264, 2381, 182, 2417, 1681, 2101, 193, 562, 2584, 1545, 1517, 2006, 489, 926, 1552, 2115, 1467, 1662, 928, 115, 1349, 1512, 2781, 1332, 688, 486, 2618, 265, 275, 393, 2005, 1950, 2102, 223, 2450, 1366, 2358, 2948, 1348, 2190]. Book [264, 1756, 437, 725, 864, 1836, 2383, 2248, 2931, 2786, 1065, 2862, 501, 2597, 748, 1363, 1539, 2715, 384, 916, 76, 289, 506, 1420, 2167, 1751, 2208, 438, 1208, 2100, 2139, 449, 459, 1016, 1887, 1407, 2429, 1063, 1217, 1376, 2131, 2145, 487, 989, 1488, 3081, 2930, 1506, 2646, 2185, 96, 2729, 458, 1440, 461, 710, 26, 491, 1399, 1411, 1876, 2355, 2334, 1803, 1284, 224, 1138, 600, 1555, 1885, 1807, 2433, 368, 2487, 2410, 2257, 66, 782, 1716, 900, 2357, 277, 1753, 2122, 2206, 2444, 1893, 279, 1973, 48, 499, 2043, 2721, 2949, 2495, 2780, 918, 1061]. Book [2817, 609, 931, 198, 557, 884, 658, 2928, 313, 602, 1935, 2225, 1660, 141, 1792, 2500, 1355, 1405, 309, 1599, 1746, 1706, 1690, 1721, 1974, 2661, 1731, 745, 917, 1234, 2096, 2877, 2449, 2025, 826, 1403, 1729, 2197, 2778, 47, 1057, 1704, 2118, 2137, 2445, 2564, 1978, 21, 507, 143, 1610, 2559, 1785, 504, 1422, 2274, 2869, 683, 2293, 211, 890, 2861, 370, 1090, 849, 292, 902, 1265, 2809, 228, 896, 2582, 2655, 882, 477, 1789, 2350, 674, 1000, 1879, 1750, 1693, 1546, 2636, 2574, 1892, 2294, 2917, 630, 566, 1814, 1303, 2217, 2111, 599]. Book [2026, 2690, 1645, 1810, 287, 1034, 2515, 190, 302, 447, 859, 1455, 1573, 2583, 1046, 2015, 586, 1087, 1044, 2425, 147, 1567, 1350, 97, 405, 2099, 440, 1692, 2216, 943, 812, 660, 2187, 1082, 1832, 1221, 1484, 2496, 2927, 87, 390, 892, 1500, 2919, 2954, 772, 1142, 180, 984, 1889, 2349, 2489, 2876, 256, 1838, 2136, 53, 401, 1300, 2498, 1076, 1330, 773, 935, 1201, 1406, 1279, 2033, 2036, 1799, 2624, 1794, 2332, 2641, 581, 2221, 1806, 221, 301, 1264, 2282, 2938, 1951, 2791, 1820, 1297, 283, 558, 991, 1816, 1550, 2034, 82, 934, 993, 2270, 2924, 875]. Book [1490, 1364, 2937, 870, 3093, 1126, 2023, 2717, 857, 1187, 232, 2301, 2430, 2277, 325, 662, 2854, 933, 1894, 304, 2302, 807, 1079, 2572, 1276, 483, 1043, 1561, 2078, 1289, 2184, 978, 2596, 818, 1113, 962, 2507, 2635, 2538, 583, 2525, 2113, 234, 742, 2176, 155, 1118, 734, 2957, 1831, 2648, 1911, 640, 1725, 2843, 2121, 1091, 929, 2836, 128, 689, 2041, 802, 1294, 1423, 2319, 2371, 919, 1519, 2702, 800, 32, 887, 1204, 1745, 216, 542, 2516, 2695, 922, 956, 877, 40, 1559, 1805, 2298, 460, 851, 2202, 590, 267, 2423, 2914, 1557, 2335]. Book [852, 2062, 1607, 2327, 1884, 2667, 175, 192, 785, 2666, 672, 898, 2065, 666, 1298, 2345, 776, 1971, 93, 2173, 2090, 1747, 1902, 951, 475, 789, 1205, 1707, 2633, 2506, 997, 544, 1144, 1651, 888, 2002, 930, 829, 1741, 901, 897, 787, 2569, 358, 546, 1064, 1811, 694, 1140, 1728, 2037, 2644, 54, 579, 957, 181, 516, 1601, 1985, 2725, 278, 711, 1066, 2657, 2939, 350, 2701, 1551, 363, 1571, 272, 2222, 552, 429, 2331, 519, 1314, 1116, 2705, 1883, 2333, 582, 1752, 751, 2130, 1230, 448, 960, 1083, 1378, 1431, 591, 2125, 2738, 1536, 2788]. Book [867, 1410, 1702, 2107, 1907, 2794, 144, 2106, 879, 295, 1592, 667, 2544, 715, 2321, 2123, 2734, 1433, 2831, 2448, 136, 2029, 453, 1219, 2589, 382, 1059, 2733, 1288, 680, 543, 1120, 987, 1199, 1639, 2046, 1491, 1787, 937, 1458, 1937, 1473, 253, 2602, 2055, 1627, 732, 1542, 34, 2367, 439, 1216, 1899, 1301, 1549, 1793, 1622, 1183, 830, 1830, 2668, 860, 149, 2568, 1914, 2009, 525, 1809, 2857, 740, 1584, 1151, 1380, 723, 2374, 2260, 649, 2654, 1587, 2278, 722, 1891, 1310, 1293, 1878, 741, 2865, 2284, 540, 2935, 1220, 359, 2028, 90]. Book [75, 1800, 111, 1823, 2492, 373, 2718, 2347, 269, 728, 718, 353, 150, 1439, 738, 208, 578, 1359, 1510, 2251, 1045, 63, 2370, 1178, 1471, 2416, 2219, 95, 977, 2288, 2147, 2921, 790, 1808, 2713, 2850, 652, 1328, 2745, 151, 1266, 939, 1040, 2537, 775, 2359, 633, 1143, 1137, 2581, 752, 2889, 476, 2166, 2922, 2215, 1188, 2722, 2804, 455, 2961, 2265, 2469, 369, 306, 219, 299, 23, 430, 451, 366, 891, 985, 1089, 1421, 1432, 1278, 1748, 1493, 1593, 1604, 2653, 55, 1232, 1958, 1015, 2741, 260, 655, 949, 2256, 80, 921, 2520, 255, 2364, 2488, 1212, 69]. Book [1213, 1371, 2353, 799, 1680, 1874, 758, 1157, 2097, 2467, 2627, 291, 1088, 778, 2258, 1428, 1827, 611, 1312, 288, 1379, 1508, 953, 481, 113, 145, 746, 1093, 397, 1754, 2124, 2273, 2292, 1358, 33, 364, 720, 201, 873, 675, 1415, 589, 1352, 1270, 60, 490, 786, 958, 1871, 2418, 426, 804, 729, 983, 1211, 1635, 2109, 2092, 1759, 1979, 2797, 1224, 1475, 1976, 2534, 1589, 2135, 2287, 85, 894, 1243, 2441, 2253, 2246, 819, 2212, 1796, 29, 119, 648, 1637, 2706, 1078, 2619, 2950, 1865, 1156, 2803, 2207, 2642, 2694, 2196, 456, 2189, 2348]. Book [1331, 1906, 2851, 2436, 1466, 185, 1412, 2326, 281, 1353, 2647, 1417, 1815, 663, 1375, 2732, 217, 514, 820, 1540, 2455, 1743, 2336, 1570, 2098, 148, 828, 425, 2716, 713, 1084, 1861, 251, 686, 976, 480, 1867, 445, 2842, 2566, 527, 644, 187, 485, 925, 1195, 1938, 1472, 1687, 484, 2563, 2542, 2743, 226, 691, 1947, 1740, 1956, 1494, 495, 2565, 777, 94, 517, 999, 2048, 2872, 218, 70, 1007, 1377, 2247, 2625, 959, 1503, 65, 2845, 1826, 793, 1125, 1281, 1468, 1700, 1795, 2634, 64, 1122, 1465, 1564, 2792, 57, 396, 569, 1936]. Book [492, 1479, 1538, 2008, 2703, 2493, 2012, 2856, 1340, 2664, 1398, 1691, 2587, 927, 2342, 1277, 50, 1560, 2168, 2708, 2785, 1235, 2465, 1948, 567, 1825, 428, 2891, 834, 647, 524, 2035, 2195, 2777, 195, 1292, 1369, 124, 650, 2782, 556, 651, 952, 2209, 2205, 2593, 2888, 1404, 1342, 1608, 2849, 2366, 2658, 37, 1476, 1961, 2434, 981, 2375, 392, 682, 693, 1186, 1679, 1574, 2192, 41, 2422, 2089, 1688, 1345, 1553, 2181, 996, 1595, 822, 1955, 2290, 2796, 2018, 631, 322, 1149, 2140, 258, 230, 2816, 2711, 810, 1108, 1365, 1489, 2021, 1872]. Book [1469, 2880, 1603, 2576, 895, 1434, 2044, 2704, 1582, 1544, 1121, 380, 831, 1041, 1711, 2129, 2691, 2736, 1133, 2940, 318, 444, 1181, 1202, 1870, 1797, 2543, 2724, 1361, 1291, 1749, 2806, 1117, 596, 28, 1372, 1972, 1657, 1686, 375, 20, 387, 747, 1343, 1630, 2409, 2407, 1075, 1913, 1772, 2341, 2626, 323, 146, 406, 1705, 355, 643, 2775, 2285, 3129, 25, 2521, 2561, 1963, 98, 328, 863, 1965, 381, 801, 1888, 2010, 2438, 791, 988, 2916, 661, 580, 1742, 1463, 1134, 2268, 233, 1437, 2714, 311, 2413, 46, 1409, 2428, 2535, 2793, 2558, 478, 2421]. Book [2431, 1655, 2337, 2016, 2354, 1703, 1875, 117, 1313, 424, 1732, 1008, 2183, 3184, 2895, 2946, 3141, 2536, 2967, 3081, 2927, 2954, 3093, 2544, 2469, 2872, 2940, 2543, 3129, 177, 430]. Bookchin [2135]. Booklet [2792]. Books [2728, 1718, 1344, 2612, 2401, 1999, 1576, 1677, 2427, 1667, 278, 1314, 3038, 3218, 3287, 1790, 2949]. Boran [2709]. bord [1419, 2997]. Borderlands [994]. Borders [2555, 2730, 1731, 3282]. Borel [1459]. Borghero [1203]. Boris [2211, 1078]. Born [364]. Borrelli [2787]. Borrowed [97]. Boston [1546, 1564, 2162]. Botanic [1060, 2805]. Botanical [967, 1319, 1050, 2015, 1350, 3081]. Botanik [272]. botanischen [2343]. Botanist [1878, 1232]. Botany [464, 2920, 967, 2639, 278, 1143, 2647, 2562, 1046, 2260]. Botelho [2746]. Both [755]. Bots [2785]. Boudet [1470, 2697]. Boudia [1664]. Bouillaud [2579]. Bouk [2122]. Bouley [2788]. Boulogne [571]. Boulter [2726]. Boulton [1211]. Boumans [2327]. Boumediene [2261]. Bound [1618, 2246, 981]. Boundaries [88, 239, 1619, 2554, 1550, 1770, 54, 2492, 3026, 3074, 1867, 1473, 1434]. Boundary [716]. Bounded [1617]. Boundless [2724]. Bourdeau [2940]. Bourgeois [1838]. Bourguet [2578]. Bourke [1864]. Bow [241]. Bowditch [2191]. Bowen [1468]. Bowers [1348, 2563]. Bowler [1299]. Bowman [2075]. Box [606, 228, 2619]. Boyd [1731]. Boyle [1414, 932, 1811, 1212, 2706, 3082, 1811, 1800, 273, 654]. Boys [597]. BPA [1240]. Bracco [2878]. Brad [1017]. Bradford [2788]. Bradley [1536, 1656]. Bragg [810, 2045]. Braided [2367]. Brain [2322, 1169, 610, 1238, 2652, 2038, 2590, 2250, 2978, 1569, 1152, 1153]. Brains [2620, 2414, 908]. Brainwashing [2296]. Branching [496]. Brand [91]. Brandie [1801]. Brandt [593, 2521]. Brant [2576]. Braun [2257, 1725]. Brave [4]. Brazil [2275, 2368, 2588]. Brazilian [2895]. bread [3187]. Break [569, 1086]. Breakfast [746]. Breaking [2225, 1867]. Breakthrough [853]. Breakthroughs [176, 527]. Breasted [2879, 939]. Breathing [3233, 1725]. Breathless [1037]. Bredekamp [789, 2002]. Breeding [995, 1959, 2518, 699, 1035]. Breeds [2732]. Breidbach [1329]. Brenda [1005, 74]. Brendan [2486, 1639, 2566]. Brenna [349]. Brenni [2796]. Brent [2015]. Brentjes [2629, 2914]. Bret [1650]. Breuil [942]. Breve [3203, 1641]. Brian [1983, 2514, 2510, 2439, 24, 1187, 2664, 2168, 2888, 1149]. Bricker [1182]. Bridge [2004, 934]. Bridger [2220]. Bridging [1670, 2106]. Brief [1755, 2320, 1456, 489, 48, 3285]. Briefe [947, 125]. Briefwechsel [743, 2651]. Bright [854, 2090]. Brigitte [1069, 2326]. Brill [1552, 1539, 1546, 1549, 1564, 2178]. Brilliant [447]. Bringing [762]. Brita [349]. Britain [1424, 1649, 3025, 665, 1176, 2640, 2595, 2860, 2338, 2801, 2042, 2539, 1968, 2138, 1581, 2444, 2721, 1265, 2648, 544, 949, 1759, 2196, 2647, 2576, 2975, 101, 1450, 838, 2976, 3244, 1583, 1241, 3072, 3281, 3167, 530, 1309, 2032, 1135, 2858, 2266, 2040, 1734, 1482, 1821, 1058, 2452, 2430, 1140, 2251, 131, 881, 674, 87, 877, 959, 392, 318]. Britannica [355]. britannici [1109]. British [2956, 795, 1964, 285, 2032, 2356, 1109, 2994, 815, 1818, 1356, 2344, 2862, 2809, 888, 1536, 2544, 2804, 2732, 661, 2134, 4, 335, 1667, 1951, 1766, 666]. Brittleness [2814]. Broadcast [599]. Broca [2579, 380]. Brock [1314, 1137, 2089]. Broderip [838]. Brodman [2751]. Brogan [2150]. Brogi [988]. Brokered [283]. Bromberg [1775]. Bronk [21]. Brook [1875]. Broomhall [2323]. Brosco [1437]. Brotons [1877]. Browman [1460, 1355]. Brown [1600, 2091, 849, 807, 2234, 1199, 2653, 481, 567]. Browne [277, 1484, 1059, 41]. Bruce [2036, 2106, 2816, 44, 2062, 578]. Brückle [789]. Bruderer [1230]. Brummelen [1194, 33]. Bruno [2700, 1339, 2713, 2989, 1980, 1380]. Brunton [1308]. Brussels [1060]. Brute [908]. Bruun [1341]. Bruyninckx [2876]. Bruzzaniti [2207]. Bryant [1844]. Bubbles [1619]. Buchalter [1432]. Buchanan [3010, 1567, 517]. Buchwald [1507, 808, 2049, 1455, 1676]. Buckland [1300]. Buckley [2588]. Budapest [2060]. Buddhism [2840, 182]. Buddhist [1787]. Buenos [878]. Buffon [1485, 369, 3]. Bühler [1561]. buhran [1042]. Build [2891]. Building [1523, 2158, 2650, 2368, 2479, 2884, 2741, 2540, 1839]. Bulldog [59]. Bulmahn [2339]. Bund [2668]. Bundesrepublik [2226, 1310]. Bunk [1172]. Burdick [44]. Bureau [4]. Bureaucracy [1351, 3085]. Bürgertum [1498]. Buridan [2103]. Burke [2639, 1888]. Burkhardt [1720, 1059]. Burndy [24]. Burned [2713]. Burnett [1546]. Burney [2448]. Burnham [1659, 2967, 2005, 2405, 2912, 2908, 2909, 2911, 2904]. Burton [2167]. Burtt [1214]. Burwell [2749]. Bush [2225]. Business [2191, 1051, 884, 828, 2728]. búsqueda [2722]. Butrica [1748]. Butterfly [2489]. Buzjani [2499]. Byrne [1517]. Byways [179]. byzantinischer [2422]. Byzantium [2330, 1690, 3275].

C [1964, 1531, 2894, 2210, 1341, 2032, 2912, 1416, 378, 1077, 1734, 2146, 1074, 2039, 1701, 946, 2696, 1237, 1933, 2186, 1305, 1784, 1056, 1513, 2005, 1950, 2908, 1411, 1061, 1721, 1889, 1806, 2302, 2635, 1423, 2909, 1571, 2537, 2469, 2653, 1213, 1353, 2911, 1468, 2856, 1711, 2840, 1705, 3227, 478, 2571, 3043, 2698, 2042, 3263, 1555, 3161, 2173, 2016, 443, 452, 849]. C. [2210, 2846, 1399, 1693, 2816]. C.-E [1806]. ca [2787, 2932]. Cabibbo [1463]. Cabinet [1485, 369, 2642]. Cabinets [3029, 1408]. Cable [135]. caché [478]. Cadwallader [2344]. Cage [2447]. Cagle [2930]. Cahalan [2639]. Cahier [909]. Cain [1829]. Calculating [2573]. Calculation [1129, 2993]. calculs [478]. Calculus [2761]. Calculus-Based [2761]. Calendar [2846, 1693, 2821, 3181, 3185]. Calendaria [2821]. Calendars [1546, 773, 1937, 3294, 489]. Calendrical [2846]. calendrier [478]. California [82, 1013]. Calinger [2114]. Call [2223]. Callaway [1827]. Calling [1738, 3242]. Calmthout [2451]. Calvin [724]. Calvinists [3246]. Calvino [595]. Cambodia [805]. Cambridge [574, 422, 1193, 1416, 1579, 1580, 154, 1537, 1548, 1583, 721, 1596, 1602, 1599, 287, 2982, 1571, 1536, 1542, 3218, 63, 1589, 1570, 396, 1623, 424, 3262]. Cambridgeshire [1423, 3005]. Cambrosio [1084]. Came [853]. Camera [2564, 299, 3034, 1329]. Cameralism [128]. Camilleri [56]. Camilo [1146]. Camouflage [1072]. Campbell [2267, 1330]. Campbell-Kelly [1330]. Campe [1129]. Campenot [2250]. Camper [2021]. Campi [45]. Campos [1902]. Camprubí [1596]. Can [1331, 3059]. Canada [2945, 479, 3050, 2968, 1430, 1261]. Canadian [1229, 1328, 2984]. Canal [198]. Canales [1983, 2291]. Canalis [2927]. Canals [2153]. Canani [1203]. Cancer [314, 1746, 2217, 640, 3297, 1084]. Candolle [1319]. Canham [1328]. Cannibalism [497]. Cannibals [881, 165]. Canonical [2317]. Cantabrigiensis [1423, 3005]. Cantemir [1130]. Cantor [1497, 2882, 1833]. Capacity [2817]. Capecchi [2427, 2416]. Capellini [1109]. Capital [1159, 152, 1571, 3070]. Capitale [875]. Capitalism [531]. Capitanucci [1641]. Cappelli [1075]. Capsule [2913]. Capsules [2163]. Captain [569]. Capturing [2388]. Car [3012, 1374]. Cardano [2898, 2334]. Cards [923, 1888, 3150]. Care [467, 265, 2005, 3203]. Career [133, 1605, 1725, 495, 517, 238, 3233, 3051]. Careers [700, 1309, 144]. Caribbean [2951, 1350, 2507, 3081, 1103]. Carin [2034, 2648]. Caring [2036]. Carl [1441, 2273]. Carla [2095]. Carlat [1064]. Carleton [2450]. Carlo [1203]. Carlos [1654, 1752]. Carlsberg [2445]. Carlson [1534]. carnet [2578]. Carnino [2569, 2703]. Carola [1233, 1972]. Carole [1469]. Carolina [1593]. Caroline [2879, 576, 1360, 3067]. Carolus [367]. Carolyn [2318]. Caroti [1344]. C̈arrara [2926]. carrefour [875]. Carrier [2381, 2842]. Carroll [2437]. Cars [1374, 3012]. Carson [250]. Carsten [1783, 3207, 1648, 1709, 1188, 3230, 1746]. carteggio [3039, 1357]. Carter [60]. Cartesian [2006]. Cartesianism [1880]. Cartesians [1880]. cartésiens [1203]. cartográfica [1939]. Cartographer [1563, 3077]. Cartographic [920]. Cartographies [850]. Cartography [2386, 2571, 52, 1464, 994, 2459, 1147, 2278, 1986, 2724, 262]. Cartwright [1964]. Casanova [1646, 3003, 3003]. Case [386, 1004, 2211, 2866, 916, 882, 97, 2760, 105]. Cases [757]. Caso [75]. Caspar [2694]. Casper [2042, 2652, 1220]. Cassiano [1197]. Cassirer [2203]. Castellani [1075]. Castle [1715, 3276]. Casualties [938]. Cat [1899]. Catalin [497]. Catalog [2219, 186]. Cataloging [2044]. Catalogs [923]. Catalogue [1424, 1416, 1564, 3072, 2982, 3129]. Catastrophe [2814, 916, 2135]. Catastrophic [1661, 3062]. Cataudella [2178]. Catch [2526]. Catches [2640]. Catching [3170, 1565]. Categorization [1027]. Category [2018]. Catelijne [1685]. Caterina [2934]. Catherine [443, 1835, 2131, 1380]. Catholic [1268, 2788, 709]. Catholicism [1501]. Catholics [3246, 1766]. Cathryn [250]. Cathy [2619]. Cats [1995]. Cattell [72, 160]. Cattle [699]. Causality [3071, 2763, 1645]. Cause [1397, 3007]. Cautionary [352]. Cavendish [436, 1414, 2022, 3082, 1801, 2706]. Cavert [2262, 2217]. Cayleff [2288]. ce [669]. Ceccarelli [1512]. Cecilia [1341]. Cédric [1485, 369]. Celebrating [2808]. Celebrity [1970, 1057]. céleste [992, 478]. celesti [493]. Celestial [2848, 1464, 641, 3157, 2898, 552, 1937, 1178, 2409, 2421, 1559, 1560]. Celeyrette [1344]. Cell [2752, 2656, 1435, 1440, 3199, 3116]. Cells [1009, 358]. Celtic [1501]. Censorship [2677]. Census [2330]. Centennial [1119, 2808]. Center [77, 1563, 1545]. Central [2447, 2584, 2046]. Centres [937]. Centric [2670]. Centuries [1726, 2094, 637, 2489, 2195, 3257, 1102, 2490, 2519]. Centuries-Long [2094]. Century [1523, 1649, 2030, 2802, 2191, 1247, 2499, 1309, 1568, 1659, 214, 1910, 2032, 821, 377, 1983, 1109, 1577, 129, 1575, 1788, 389, 1005, 1004, 2188, 2142, 1408, 2588, 797, 2056, 1580, 1715, 86, 2662, 1073, 500, 1917, 853, 554, 1495, 1737, 3150, 1438, 629, 1644, 2071, 1739, 1591, 2659, 2628, 1581, 376, 1013, 2267, 200, 2518, 131, 1147, 290, 1058, 188, 607, 881, 515, 2427, 1062, 2066, 1467, 1349, 2102, 725, 2786, 2862, 2646, 66, 2928, 1706, 1721, 2294, 1046, 390, 1500, 1300, 2430, 2572, 583, 2113, 734, 2648, 877, 3166, 2231, 1918, 1144, 278, 2130, 1491, 151]. Century [2537, 1976, 2073, 1865, 2196, 2436, 2336, 2743, 959, 567, 392, 2021, 1133, 1870, 117, 2975, 3029, 2987, 2992, 3238, 3242, 3055, 3057, 3001, 3013, 3273, 3091, 3167, 3183, 3121, 3301, 3227, 3130, 3122, 965, 966, 1942, 3204, 1618, 102, 1021, 3003, 2547, 1565, 1606, 3075, 3010, 2478, 3198, 2088, 3276, 1667, 3096, 1105, 1593, 3220, 3170, 1024, 339, 3172, 1515, 2835, 1402, 1413, 2019, 1896, 2557, 1426, 2503, 2346, 2194, 1977, 2953, 1738, 1077, 2134, 2459, 2855, 2915, 1496, 2171, 2186, 2882, 2639, 1829, 1833, 1951, 1911, 1707, 2333, 1830, 999, 1342, 2413, 861, 142]. Century [575, 523, 221, 852, 930, 977, 820, 976, 556, 381, 661]. Century-Long [86]. Century/La [1109]. Cerebral [2620]. Cerezo [2659]. Certainty [1138, 70, 62, 425]. Ceruzzi [77, 1076]. Cesi [2015, 2791, 1700, 3248]. Cetaceans [976]. Cevolini [2645]. Chabás [1122]. Chairman [1314]. Chakrabarti [1356]. Challenge [434, 477, 97, 448, 2520].

Challenger [147]. Challenges [2825, 291, 1867, 1108, 3194]. Challenging [2161]. Chamber [2642]. Chambers [1701, 884, 1704, 592]. Chamisso [2799]. Champion [369]. Chance [1965]. Chaney [2740]. Chang [525]. Change [1516, 1370, 2238, 2769, 2351, 1374, 2302, 216, 233, 3012, 3123, 525]. Changed [2361, 2191, 821, 2291, 368, 558, 746, 1826, 2796]. Changes [546]. Changing [2598, 197, 1060, 1148, 1679, 3196]. Channell [2492]. Chantal [1481]. Chao [2902, 2233]. Chaos [173, 97]. Chapel [1593]. Chard [2746]. Chariot [2605]. Charis [1440]. Charles [1419, 2172, 1900, 2858, 1720, 2411, 1546, 1144, 1059, 2028, 168, 1965, 2622, 3181, 3182, 3185, 3191, 2363, 1988, 2723, 1715, 1174, 2034, 1774, 1767, 1213, 55]. Charlotte [401, 1984]. Charly [1181]. Charm [2394]. Charra [1553]. Chart [1875]. Chartes [1553]. Charting [1812]. Chartres [1940, 1794]. Charts [2784, 1612]. Chasing [2753, 1504, 2754, 980]. Chatto [1577, 1559, 1584]. Che [2355]. Checkered [598]. Chela [22]. Chela-Flores [22]. Chemical [2870, 1783, 874, 1241, 1558, 1056, 1149, 3031]. Chemicals [1740, 3293, 1240]. Chemie [2449, 1188]. Chemist [1581, 279, 3167, 3019, 2881, 1655]. Chemistry [1869, 1247, 2120, 1886, 621, 2472, 2552, 1449, 945, 2435, 1056, 2060, 47, 2554, 857, 2277, 2553, 2473, 2106, 1948, 2089, 2881, 945, 882]. Chemists [2555]. Chemla [2321, 2166, 1404]. Chen [174, 2264, 1372]. Chen-Morris [2264]. Chen-Pang [1372]. Cheng [1077]. chercheurs [2008]. Chernobyl [2140]. Cherp [1333]. Chertok [1078]. Cheryl [1224]. Chesapeake [356]. Chess [507]. Chet [2602]. Cheung [23]. Chiang [2001]. Chicago [1598, 1586, 1543, 1565, 1605, 1573, 1601, 1595]. Chicago/London [1598, 1586, 1543, 1565, 1605, 1573, 1601, 1595]. Chickens [995]. Chilam [773]. Childbirth [1697, 775, 3226, 2527]. Childhood [3237, 2293, 1750]. Children [1083, 445, 2366]. Chile [3217, 1723]. chimiche [1184, 1490]. Chimico [508]. Chimie [2184, 587]. chimistes [587]. China [1848, 2050, 2091, 1733, 2283, 1107, 1642, 670, 223, 230, 1875, 3043, 1850, 2071, 3064, 2154, 1314, 2237, 969, 2905, 3289, 3299, 2783, 2224, 2855, 2628, 1151, 1743, 2782, 1961, 280, 276, 734, 519, 151, 775]. China.} [2920]. Chinatown [1070]. Chine [2342]. Chinese [2001, 2059, 1304, 549, 688, 2695, 1266, 647, 2421, 2612, 2684, 1390, 1112, 2479, 969, 1765, 105, 2233]. Chinese/American [105]. ching [174]. Chinnici [1357]. chinois [478]. chinoise [478]. Chiropractic [2530]. Chlorine [232]. Choice [107, 2238, 2244, 428, 3048, 1588]. choisir [3250, 1553]. Cholera [1496]. Chomsky [1767]. Choosing [516]. Choreography [1440, 3116]. Chorlay [2321]. Chosen [1738, 3242]. Chotkowski [1231]. Chris [2455]. Christelijke [3246, 1732]. Christensen [1181]. Christi [2187]. Christiaan [1564, 3129]. Christian [947, 1459, 1946, 1693, 723, 1181, 2340, 618, 2878, 3246, 1966, 2882, 1303, 2023, 2209]. Christiane [1419]. Christianity [294, 1921, 2400, 265]. Christina [593, 1834, 2521, 1898]. Christine [48, 1553, 1812, 2644, 2251]. Christof [739]. Christoph [1511, 990, 1049, 1412, 1623, 1393, 1698, 1206, 3252, 2916]. Christophe [1594, 2227]. Christopher [1036, 82, 1081, 1943, 1678, 1912, 1488, 2257, 1747, 2792, 1825, 2782, 2421]. Chronic [3301, 1739]. Chronology [1399, 2821, 2981]. Chrysopoeia [2472]. Chuanbo [736]. Chunjuan [1314]. Church [1901, 709, 2788]. Churches [1669]. Churchill [2868, 2483, 1265]. Chymist [2080]. Chymistry [415, 1261]. Ciampoli [2185]. Ciancio [1109]. Ciavolella [2927]. ciel [371, 190]. Cieli [2850]. ciencia [2110, 2279, 859]. Ciencias [136]. científica [566]. Cigarette [515, 916]. Cimento [270]. Cinchona [2112]. Cinema [1967, 2204, 822]. Cinematicus [2654]. Cinquecento [2324]. Cinzia [1471]. Cipriani [1461]. circa [101, 1674, 4, 702, 2677]. Circle [1309, 2844, 2655]. Circulation [792, 966, 2514, 2630, 817, 971, 2222]. Circuli [1197]. circulo [1873]. Cirella [2355]. Citations [169, 417, 1255]. Cities [2542, 1469]. Citizenship [1510]. Citti [2176, 1041]. City [878, 1917, 2928, 1536, 1742, 3193, 3240, 2262, 2704]. Civic [376]. Civil [2156, 1177, 361, 132, 2157, 2593]. Civilisation [2920]. Civility [2085]. Civilization [737, 1930, 1676, 3202, 174, 2717]. Civilizations [853, 1642, 3064]. Civita [2947]. Ciyuan [2695]. Claims [1086, 2151]. Claire [1137, 41, 1424, 2336]. Clancey. [1150]. Clandestine [839, 1143]. clandestinité [649]. Clare [2576]. Clarifications [1997]. Clark [57]. Clarke [2062, 1705, 3263]. Class [303, 2073]. Classic [2082, 2403, 2480, 2175]. Classical [1457, 1627, 2416, 2921, 3076]. Classics [1585, 1790]. Classification [1824, 1989, 1429, 519, 3046]. Classifying [459]. classique [1289]. Classroom [1790]. Claude [909, 478, 1499, 1493]. Claudia [2916, 2324]. Claus [756, 1815, 1679, 2916]. Clavelin [2170]. Clear [2045, 2236]. Clemens [2847]. Clergue [2316]. Clerk [3023, 1954, 1379]. Clifford [2575, 2515]. Climate [795, 518, 2769, 598, 2351, 1613, 1267]. Clinic [2036]. Clinical [225, 2547, 2604, 1835]. cliniques [753]. Clinton [1712, 3291]. Clio [605]. Clipping [1744, 3296]. Clock [2094]. Clocks [1990, 1997]. Clodius [1448]. Closed [897]. Closet [467]. cloth [1590, 1569, 1562, 1568, 1586, 1577, 1575, 1600, 1579, 1609, 1580, 1537, 1547, 1548, 1583, 1563, 1596, 1591, 1565, 1605, 1534, 1576, 1581, 1602, 1606, 1558, 1677, 1545, 1552, 1539, 1555, 1599, 1610, 1546, 1573, 1567, 1561, 1559, 1557, 1607, 1601, 1314, 1536, 1592, 1549, 1584, 369, 1593, 1604, 1676, 1589, 1570, 1564, 1560, 1608, 1574, 1582]. Cloud [752]. Club [2457, 2942, 746]. Clubs [1510]. Clusius [367]. Coal [2283, 69]. Coast [670]. Coat [206, 1985]. Coat-Tails [206]. Coats [822]. Cobb [1985]. Code [637, 2460, 1985]. Codex [1201, 1795]. Codices [1182]. Coeckelbergh [2466]. Coen [1215]. Coffin [569]. Cognition [2117, 985, 693, 2916]. Cognitive [2823, 448, 160]. Cohen [1866, 1662, 185]. Cohen-Cole [1662]. Coil [2246]. Colavita [1471]. Colbert [561]. Cold [2956, 1242, 1177, 92, 2142, 2220, 893, 2540, 2955, 1908, 2541, 1663, 2141, 1605, 2462, 1239, 1662, 2957, 898, 751, 952, 895, 1313, 2240, 2550, 3137, 108, 107, 3051, 907, 3004, 109, 106, 410, 2222, 2307, 105, 2616, 73, 1012, 149]. Colden [2344]. Cole [1662]. coleccionismo [2110]. Coles [2095]. Colin [1538, 1888]. Collaboration [2297, 469, 810, 1434, 3074, 829]. collaborations [1463]. Collaborative [1069]. Collapse [1131]. Collard [1470, 2181]. Collected [1507, 1517, 742]. Collecting [2835, 2532, 2715, 2267]. Collection [89, 2385]. Collections [508, 1882, 1319, 2163, 1788, 1788]. Collector [951]. Collectors [572]. College [908, 2277, 2736, 824]. Colleges [1501]. Collet [46]. collezioni [1461]. Collider [2144]. Collins [1515, 2837, 2468, 2671, 1064]. Colloquia [1556]. Colombia [1146, 465]. Colomo [1075]. Colonial [2527, 805, 1708, 2514, 1222, 1418, 2412, 1818, 1689, 991, 2516, 922, 1491, 3024, 3100, 3231]. Colonialism [500, 948, 2371]. Colonies [922, 1235]. colonisation [2261]. colonne [205]. Colony [2805]. Color [1414, 2929, 640, 490, 456, 3082]. Colorado [1668]. colore [1341, 3093]. Colors [1201]. Colour [1341, 3093]. Colp [166]. Colton [626]. Combe [1368, 2973]. Combeau [2581]. Combeau-Mari [2581]. Come [2685]. Comet [1698, 1806, 3252]. cometas [1873, 1479]. cometis [1873]. Comets [27, 891]. Comfort [1306, 1079]. Coming [1505, 2312, 481, 1081, 519]. Coming-to-Be [481]. Comisión [75]. Command [217]. comme [689]. Commelin [255]. Comment [2082, 1850, 2084, 2077, 2079, 1773, 2083, 2078, 2081, 2086, 2085, 2080]. commentaire [2419, 3141, 2177]. commentaires [648]. commentaries [3005, 2584, 1423]. commentario [493]. Commentary [1940, 1621, 284, 1042, 2497, 2972, 2446, 1517, 1183]. Commented [1795]. commentée [785]. Comments [2158, 2313]. Commerce [1586, 1111, 2647, 3095]. Commercial [2213, 3271, 1703]. Commercialization [1872]. Commission [2605]. Committee [92, 1452]. Commodification [588, 522]. Commoditization [803]. Commodity [854]. Common [1326, 2152, 758]. Commoner [2605]. Communal [1469]. Communicating [3224, 1556, 320, 817]. Communication [88, 2809, 65]. Communications [135]. Communist [601]. Communities [1127, 2131, 1456, 376]. Community [1653, 809, 112, 1155, 2020, 2540, 1237, 2265, 3138, 2789]. Companies [1080]. Companion [574, 154, 2178, 3262, 2495, 1639, 1623, 424, 3239]. Companions [2244]. Company [1566, 2345, 2933, 3135, 1918]. Comparable [2841]. Comparative [1529, 2317, 2584, 264, 1539, 1925, 2090, 3155]. compared [3043, 1107]. Comparison [789, 969, 361]. Compass [1047, 193]. Compatible [2902]. Compendious [470]. Compete [1331, 3059]. Competing [2617, 1516, 1242]. Compétitions [1463]. Complaint [782]. Complementary [372]. completa [3147, 1654]. Complete [2883, 354, 269]. complètes [1485, 1564, 3129]. Complex [1989, 1025, 2194, 1734, 1451, 2976, 3244]. Complexity [460]. Complicated [1328, 2984, 386]. Composing [2702]. compositorum [2500]. Compound [2859]. Comprehensive [1672]. Compulsory [2031]. computability [3303, 2216]. Computation [2631, 2587, 220]. Computational [2829, 914, 2825]. Computer [597, 518, 86, 1230, 2455, 1330]. Computer-and [86]. Computerizing [2751]. Computers [597, 2307, 1990, 1997]. Computing [186, 1082, 1378, 958, 1906, 3109, 1076, 1154, 2539, 1148]. Computus [1871]. Comte.} [2940]. Con [493, 1325, 1197]. Conan [2948]. Conant [1844]. Concept [1524, 2464, 934, 159, 2416, 300]. Conception [1919]. Conceptions [1389, 2213]. Concepts [2848, 1670, 1679, 3196, 2749]. Conceptual [2004, 2833, 934, 993, 2302]. Concerning [365]. Concerted [2609]. Conches [867]. Conchis [35]. concile [40]. Concise [2958, 1076]. Concordance [1564, 3129]. Conditional [2474]. Conditioning [102]. Conditions [2903]. Conevery [1420]. conference [3194, 1975, 439]. Confession [2789]. Confidence [2151]. Confidentiality [2444]. Conflict [1523, 1697, 3226]. Conflicting [728]. Conflicts [1492]. conflits [1463]. Confluence [603]. Conformist [1482]. Conforti [1554, 2414]. Confronting [251, 2000, 795]. Confucianism [1539, 3155]. Congregations [709]. Conics [476]. Conis [2598]. Conjecturing [42]. Conjuring [2269]. Conjuror [929]. Conklin [1358]. Conlin [1821]. Conn [1585, 1568, 1610]. Conn./London [1568, 1610]. connaissance [786]. Connect [2032]. Connection [372]. Connections [2445, 1168]. Connective [2090]. Connors [1201]. Conodont [1066]. Conquest [1602, 3153, 556]. Conscience [1312]. Consciousness [2740]. consensus [3046, 1429]. Consequences [2158, 2156, 2592, 1267, 2157, 715, 898]. Conservation [1989, 305, 196, 286, 1146, 1038, 2837]. Conserving [2860]. Consider [1768]. consoles [3111]. Consolidating [1953]. Conspiracies [1766]. Conspiracy [957]. Constance [57]. Constant [22]. Constantine [2255]. Constantinople [1346, 3017, 925]. Constellations [2990]. Constitution [641, 4]. Constraints [2764]. Constructing [2604, 2153, 2234]. Construction [1304, 1014, 1487, 1495, 1253, 1283, 969, 2048, 478]. Consultant [2909]. consultation [1797]. Consulting [69, 1797]. Consumer [2198]. Contact [1600, 2328, 3028]. Contagion [2000, 2943, 1111]. Contagionism [2338, 2640]. contemporaines [1243]. contemporanea [954]. Contemporaries [1723, 3217, 3267, 1577]. Contemporary [899, 883, 1391, 2396, 175]. Contentious [919, 85]. Contents [331, 614, 905, 1163, 1445, 1762, 2068, 2377, 2674, 2963]. Contestable [1583, 3025]. Contested [584, 20, 314]. Contesting [2444, 1491]. Context [367, 1567, 57, 2021, 270]. Contexts [1362, 2503, 1539, 3151, 3155, 3083, 3097, 1640, 20, 1091]. Contextual [2760]. Contextualized [250]. Conti [1641, 1630]. Continent [2174, 2361]. Continental [1262, 933, 2041, 1666]. continu [478]. Continuation [2628]. Continue [2218]. Continuités [2023]. Continuity [2922, 3138, 1653]. Continuous [1897]. continus [3141]. Contours [751]. contrôlés [753]. Contraception [826]. contraddizione [2850]. Contradictions [862]. contre [1243]. contre-science [1243]. Contributions [1963]. Contributors [99, 156, 235, 329, 407, 462, 528, 612, 695, 759, 836, 903, 963, 1018, 1095, 1161, 1245, 1315, 1381, 1443, 1760, 1520, 1611, 1665, 1840]. contro [879]. Control [2669, 1625, 1241, 2318, 2590, 2217, 969, 2537, 1593, 949, 1525, 3156, 3121, 598]. controverses [2627]. Controversial [2890, 517, 2594]. Controversies [1887, 989]. Controversy [617, 808, 1429, 72, 1822, 319, 87, 993, 232, 2189, 3046, 3146, 1569, 2795, 637, 1262]. Conventional [1753, 3253]. Conversation [1779]. Conversations [691]. Convict [1976]. Convincing [22]. Conway [2145]. Cook [2712, 1545, 1409]. Cooley [1937]. Cooper [1835, 1042]. Cooperation [1523]. Coopmans [1685]. Copenhagen [1944]. Copenhaver [2168]. Copernican [1178, 1929, 2757]. Copernicana [1563]. copernicani [727]. copernicienne [371]. Copernico [493]. Copernicus [780, 730, 2503, 1943, 1563, 1815, 2409, 1842, 2506, 2757, 780]. Copies [789]. Coppola [2430]. Copy [789]. Corbeil [2500]. Cordelia [1692]. Cordilleras [1219]. Corine [2295]. Corinna [1874]. Cormack [2512]. Cornelius [492]. Corner [77]. cornue [1289]. Coroners [425]. corpo [2324]. Corporate [531, 969]. Corporeal [2519]. Corporis [1622, 3249]. Corpses [1690, 3275]. Corpus [1197, 1552, 3117]. Corpuscular [1798]. Corpuscular-Mechanism [1798]. correspondance [2991, 1286, 1650]. Correspondence [2795, 1507, 2643, 2210, 2858, 210, 58, 2946, 1647, 1720, 1049, 1860, 1952, 1704, 1000, 1059, 1705, 2988, 3181, 3182, 3185, 3191, 3189, 3263, 2114]. Correspondences [1878]. correspondent [3077, 1563]. Correspondents [2946]. Corrigan [2510]. Corrigible [1976]. Corry [2320]. Cos [2496]. Cosmas [1608, 3047]. Cosmetic [1737, 3273]. Cosmic [451]. cosmo [493]. Cosmogonies [364]. Cosmography [282]. cosmológica [1479]. Cosmological [1347, 1396, 3032, 3065]. cosmologie [2170]. Cosmology [2052, 492, 1987, 3177, 1937, 1487, 2509, 961, 641]. Cosmonauts [1436, 1909, 3213, 1909]. Cosmopolitan [2270]. Cosmopolitanism [2364]. Cosmopolitics [526]. Cosmos [2789, 398, 2105, 2494, 2893, 1234, 930, 2626, 2702]. Cosmoscope [3150]. Cost [215, 1396, 3065]. Costa [2199, 1370, 1834]. Coste [1797]. costituzioni [1802]. Cotta [743]. Coulston [1774]. Counseling [1514]. Count [1378, 3109, 2300]. Countercultural [2743]. Counterculture [91, 2374]. Counterfeit [482]. Counterfeits [1554, 3134]. Countries [2996, 2932, 1409, 792]. Countryside [289]. Couples [1069]. cour [1632]. cours [2697]. Court [42]. Courtier [1949]. Courtly [1640, 3097]. Courtney [2329, 2188, 2518]. Courtroom [965]. Couturat [2441]. Cover [2600, 2675, 2747, 2818, 2896, 2964]. Cowan [514, 514]. Cowdry [2656]. Crack [1985]. Cracked [637, 535]. Cradle [1948]. Craft [1978, 117]. Crafting [1384, 734, 450]. Craftsman [2305]. Craftsmen [2630]. Craig [1696]. Crapanzano [1803]. Crashes [1374, 3012]. Crater [1960]. Cravens [2556]. Craver [1441]. Crawford [2112]. Creager [1518]. Crease [176]. Create [1899]. Creating [2984, 995, 2051, 2390, 854, 2507, 829, 1328, 361, 1008]. Creation [1492, 1887, 507, 773, 2745, 939, 375, 98, 2744]. Creationism [1784, 59]. Creationists [755]. Creative [298, 1033, 782, 527]. Creaturely [1403, 3223]. Credit [938, 3241, 1754]. Credulity [2861]. Creese [1963]. Crémière [1485, 369, 1485]. Crescents [2752]. Crick [403]. Crim [2888]. Crime [2618]. Crimes [2658]. Crisis [410, 579]. Cristian [2265]. Cristina [727, 48, 1202, 1414, 2260]. critica [2182, 351]. Critical [1198, 730, 1318, 1054, 2458, 2203, 2800, 1042, 1552, 2564, 993, 887, 2695, 1219, 1793, 1183, 2215, 2258, 1415, 490, 3087, 3117]. Criticism [2708]. Criticizing [2472]. Critics [2733]. Critique [1173, 2177, 779, 648, 1123]. critiques [608]. Cronache [2355, 1802]. Cross [1027, 361, 487, 555]. Cross-Cultural [487, 555]. Cross-Cutting [1027]. Cross-Disciplinary [361]. Crossed [640]. Crossing [1473]. Crossroads [439]. Crowded [532]. Crowe [2948]. Crown [1200, 2105]. Cruises [158]. cruz [1939]. Cryle [2800]. Cryptography [2714]. Crystal [2045]. Crystallography [1817]. CSI [2448]. CT [321]. Cuba [887, 2298, 1907]. Cuban [604, 384, 2359]. Cueto [1935]. Cullen [2782, 2421]. Cult [831, 117]. Cultivating [633, 2647, 82, 922]. culto [2176]. Cultura [932, 2110, 1578, 3195, 1745, 3172]. Cultural [1, 333, 616, 1164, 1446, 1763, 2070, 2384, 1006, 2007, 1930, 1575, 2524, 1909, 220, 756, 1735, 367, 2095, 1591, 1033, 1062, 1545, 1836, 487, 141, 2661, 1731, 2861, 87, 3282, 1314, 2765, 2714, 3049, 3209, 2356, 1460, 555]. culturale [2100]. Culture [878, 1031, 1326, 805, 1569, 257, 809, 1568, 423, 681, 883, 1970, 1577, 2352, 1282, 2528, 107, 2344, 2134, 1591, 1581, 376, 2844, 1558, 1821, 1504, 1050, 1545, 864, 1506, 2646, 2928, 47, 2293, 849, 1279, 2635, 1884, 544, 2705, 1702, 1278, 949, 644, 959, 2168, 37, 375, 1100, 661, 2933, 1703, 3028, 3243, 3198, 3268, 3271, 3099, 3290, 3172, 1600, 2374, 2198, 1638, 1727, 2618, 2854, 1143, 1874, 2246, 377, 446, 638, 327, 278, 738, 1266]. Cultures [1486, 2486, 2503, 1474, 972, 756, 1037, 2604, 1779, 87, 221, 2648, 1559, 2251, 1906, 2647, 3302, 2569, 3188, 699, 1686, 1189, 1789]. Culturing [358]. Cuneiform [2418]. Cunningham [2508, 2515]. Cura [2324, 1641, 3203]. Curating [2560]. Cure [604]. Curie [950, 1970, 673]. Curiosities [877, 2933, 2934]. Curiosity [653, 1967, 1709, 2715, 234, 3230]. Curious [222, 2108]. Curran [24]. Current [1, 333, 616, 1164, 1446, 1763, 2070, 2384]. Currents [360]. Curriculum [1790]. Curry [2518]. Curtain [2957, 1747, 3258]. Curtis [2204, 1842, 1965]. Curve [1607, 3152]. Curved [1615]. Curves [1820]. Curzio [1461]. Cusa [1794]. Custred [2689]. Cutting [1027, 321]. Cuvier [1816]. Cyberculture [91]. Cybernetic [610]. Cybernetics [2223, 2233]. Cyberspace [1025]. Cyborg [130]. Cyborgs [2466]. Cycles [937]. Cyclone [2679]. Cyclonic [2679]. Cynthia [2535]. Cypess [2108]. Cyrus [2469, 1237]. Cytology. [2656].

D [2586, 1031, 1569, 1416, 2415, 2340, 2656, 2300, 1701, 1605, 946, 1558, 1946, 2618, 47, 2954, 1550, 2121, 1971, 2125, 525, 2278, 1823, 686, 517, 2796, 2140, 2413]. D. [1155]. D.C [1600]. Dīn [495]. d’Abano [1470]. d’acclimatation [2282]. Dacome [2646]. Dada [130]. Daddy [145]. Dag [2426]. Dagmar [1132, 1457]. d’Ailly [2417]. Daily [464, 3187]. Dairy [699, 1686, 3302]. d’Albrecht [1352, 3092]. d’Alessandro [1197]. d’Alexandrie [2177]. d’algèbre [2502]. dalle [1461]. Damerow [736, 914]. Damian [1608, 3047]. Dampier [1337, 3126]. Dan [2122, 1181]. Dana [2033]. Dance [1138]. Dancing [2208]. Dänemark [39]. Dangerous [2432, 2009]. Daniel [2746, 2536, 1833, 2107, 1622, 1742, 2783, 1141, 2877, 1043, 1465, 1703]. Daniela [2949, 1483, 43, 1050, 2495]. Daniele [1197]. Danielle [1632, 2179]. Danilo [2427, 2416]. Danish [3230, 3207, 1648, 1709, 1065]. dans [2578, 1594, 564, 942, 3141, 2697, 2014, 122, 1203, 371, 3211, 1405, 988]. Dao [2855]. D’Arcy [1826]. Darden [1441]. Darin [2105]. Dario [2509, 1560]. Dark [313, 2090]. Darkness [1141, 447]. darkroom [3034]. Darnton [1995]. Darren [2019]. Darrigol [2006]. Darryl [1314]. Darwin [1661, 1419, 2199, 1900, 1193, 446, 394, 1724, 2649, 568, 138, 351, 2280, 59, 1723, 1821, 714, 889, 1221, 1831, 2319, 694, 1728, 55, 2275, 464, 2232, 3182, 3191, 2007, 2580, 967, 1299, 3255, 2271, 2066, 2980, 2899, 3280, 3217, 1958, 1843, 3126, 1965, 2532, 1337, 1720, 2272, 1299, 2726, 1059, 1965, 574, 139, 512, 206, 755, 449, 590, 382, 299]. Darwin-Rezeption [2280]. Darwinian [1724, 2733, 1595, 3280, 3101]. Darwinism [59, 582, 1390, 2358, 206, 504]. Darwinisms [446]. Darwinistas [1495]. Darwinists [755]. Darwinitis [834]. Darwins [2577]. Dassow [1716]. Daston [1605, 2175, 2407]. Data [2827, 961, 2826, 518, 2751, 2221, 1439, 3113, 2670]. Data-Centric [2670]. Data-Driven [1439, 3113]. Database [2219]. Daten [635]. Dathan [1232]. Dating [2981, 1399]. Daughters [1280, 1884]. Dauser [272]. Davenport [1377, 3103]. David [1590, 1460, 1724, 2649, 1336, 123, 1487, 1839, 2623, 1701, 2882, 2464, 1833, 1785, 2294, 2015, 1064, 579, 2374, 245, 686, 64, 1202, 1531, 1590, 2275, 2960, 1903, 1804, 2424, 1115, 1155, 2810, 1306, 1427, 1206, 2531, 2953, 2526, 1712, 184, 104, 3066, 1191, 1368, 636, 1217, 2131, 2357, 1355, 2574, 1573, 440, 1794, 1831, 1519, 1144, 2705, 2492, 2653, 2959, 2803, 168, 2366, 1121, 1133, 2793]. Davies [2259, 2434, 1742]. Davis [1214, 2659]. Davy [2118]. Dawes [1750]. Dawn [348, 1718, 301, 3218, 1823, 752, 3287]. Dawson [1653, 1137, 1979, 2196]. Dax [2579]. Daykeeping [773]. Days [1817, 2122, 892]. DDT [821]. Dead [572]. Deadliest [187]. Deadly [515]. Dealing [1831]. Dean [1755]. Dear [2879]. Death [1073, 1706, 3238, 545, 425]. Debate [808, 177, 2117, 2291, 1085, 2444, 1312, 22, 294]. Debates [1413, 1045, 2627, 3013, 175]. Debating [1086, 2272]. Debora [2895]. Deborah [1215, 948, 2706]. Debreu [1754, 3241]. Debt [2305]. Debus [11]. Debye [133, 740]. Decade [1152]. Decades [1089]. December [1507, 2506]. Deception [1586, 3134, 3095]. Deceptions [1554]. Deceptive [1865]. Decimals [1888]. Decision [636]. Decoding [2099, 86]. Decolonisation [2950]. décompte [2417]. Deconstruction [606]. découverte [2350]. decretoriis [1042]. DeCroes [1984]. Dedo [1942]. Deduction [2587]. Dee [929]. Deep [1274, 2862, 9, 1748, 1172, 2196, 3234]. Deep-Space [1748, 3234]. Defection [952]. Defects [2672]. Defending [2621, 185, 305, 730]. Defends [409]. défendues [274]. Defense [315, 2234, 1378, 1468, 3109]. Deficits [160]. Define [1080, 1240]. Defined [515]. Defining [1384, 1007]. Defrance [2295]. Degrees [2622]. Deirdre [2949, 2352]. Dekker [1464]. Dekonstruktion [2422]. Del [787, 2567, 498, 1109, 270, 205, 2324, 51, 199, 2513, 1461, 493, 595, 1197, 931, 993, 75, 2850]. DeLacy [2640, 2338]. Delay [2232]. Delbourgo [2715]. Delessert [1319]. Deleted [1299]. Deleuze [1603, 3098]. Delft [73, 2887]. Delia [2652]. Delicate [2098]. Delight [2489]. delineatio [2901]. Delire [2419]. Della [2787, 1038, 2355, 1641, 3203, 3204, 205, 2324, 51, 1130, 1802, 2929, 1197, 1490]. dell’archeologia [205]. dell’Arte [2929, 205]. dell’astrofisica [3039, 1357]. dell’astronomo [2355]. dell’Istituto [2276]. dell’Ottocento [3172, 1578, 1109]. dell’Universita [1461]. Delphine [2428]. Demands [400]. Demarcation [1527, 2906]. Demented [2655]. Dementia [816]. Demeter [2705]. Demidov [2211]. d’Émile [2585]. Democracy [668, 231, 1177, 1012, 2543]. Democratic [902]. Demokratie [1186]. Demon [1901]. demone [879]. démonologique [645]. Demoscenes [1906]. DeNardis [1610]. Dengjian [2325]. Denied [2586]. Denis [2125, 1713, 1060, 3197]. Denise [1123]. denken [440]. Dennis [2872]. Department [315, 2277]. deren [811]. Derksen [2693]. Derry [1959]. Descarte [2428]. Descartes [2712, 438, 1880, 1798, 1643, 2974, 1128, 197, 987]. Descartes-Agonistes [1798]. Descelier [2602]. Descent [2717]. Description [1026, 501, 1485, 369]. Desert [2465]. Design [172, 2452, 1589, 1117, 323, 3012, 1374, 513]. Designed [65]. Desire [1073, 368]. Désirée [2953]. Desires [588]. destin [1499]. Destiny [2590]. Destroyed [1057]. Destruction [506, 298]. d’Étaples [2844]. Detection [2671]. Detectives [2802]. Detector [941]. Determann [2147, 2889]. Determinants [175]. Determinate [856]. Determinism [2765]. Detlev [739]. Deutsche [2292, 589, 2190]. Deutschen [2657, 2504]. Deutschland [2226, 2522, 940, 2280, 1310]. Devastating [2146]. Developed [1032]. Development [861, 227, 2588, 2322, 2461, 676, 315, 2868, 159, 1911, 2125, 2416, 1415, 322, 1372, 747, 3034, 2976, 3087, 3111, 3128, 2302]. Developmental [1266]. Developments [304]. Deviance [2658]. Device [2099]. Devices [1967]. devient [2878]. Devin [2577]. Devlin [2175]. DeVorkin [1336]. DeVorkin. [2959]. Devotions [2188]. Devozione [790]. Dewar [1655, 3019]. Dewey [1888]. Dewhirst [1487]. DeWitt [1421, 1712, 3291]. Dezsö [2190]. DGGTB [1892]. d’Hélène [2703]. d’histoire [1882]. Dhombres [274, 114]. d’Humanisme [45]. diable [645]. Diagnosing [2658]. Diagnosis [321, 2621, 2751, 2672]. Diagram [2901, 423]. Diagrams [2420]. Diagre [1060]. Diagre-Vanderpelen [1060]. Dialogue [1047, 2569, 2092, 1480]. Diana [1507, 2049, 2884]. Diane [1336, 808, 2945, 1437]. Diaries [2885]. Diarmid [2192]. Diary [1419, 3010, 2997]. diavolo [2355]. Dibattista [1061]. Dibner [2506]. Dice [1899, 1965]. Dick [1429, 1211]. dictadura [3195, 1745]. dictae [2182]. Dictatorships [1983]. Dictionary [249, 1961]. Did [927, 2757, 580, 694]. Didier [2184]. diebus [1042]. Dieter [723, 1302, 1636, 1508]. Dietetics [2098]. Dietrich [1598, 39, 1506, 2917, 1795]. Dieu [3030, 1538]. Difference [2710, 969, 2012]. Differential [2720]. Differenz [510]. Differing [1435, 3199]. diffusion [36]. Digby [787]. Digital [1027, 1855, 1548, 91, 832, 605, 913, 1751, 1854, 2828, 2614, 3286, 2587, 3163]. Digitization [1769]. Digitizing [958]. Dilemmas [1022]. dimensio [1197]. Dimension [2064, 2873]. Dimensions [1341, 647, 3093, 284]. Dimitrie [1130]. Dinckal [739]. Diner [2542]. Dinge [2925]. d’ingénieurs [2703]. Diniz [2895]. Dinkova [1341]. Dinkova-Bruun [1341]. Dinosaur [567]. Dinosaurs [762, 2664]. Diogène [2627]. Diogenes [2627]. Diogo [2195]. Dioramen [2657]. Diplomacy [1448]. Diplomat [743]. Dippel [2863]. Dipper [739]. Dirac [207, 2052, 3177]. Dirk [73, 1556, 2887, 2708]. Dirt [1536, 3193]. Disappearance [447]. Disarmament [1241]. Disaster [1215, 858, 2129, 147]. Discarded [2539]. Disciple [512]. Disciplinary [361, 2604]. Disciplinas [1877]. Discipline [2397, 3221, 1722, 1519, 509]. Disciplines [97, 1192, 2370, 487]. Discontents [2370, 461]. Discourse [1698, 3252, 652, 1956]. Discourses [2659, 2096]. Discover [86]. Discovered [275, 2196, 567]. Discoverers [576]. Discoveries [1441, 1962, 3042, 1806, 137]. Discovering [1966, 1267, 113]. Discovers [22]. Discovery [1515, 681, 2373, 3046, 2108, 523, 1150, 2460, 2208, 2515, 525, 1765, 793, 2409, 3125, 1429, 1336, 2559, 823, 1274]. Discreet [2394]. discret [478]. Discrete [1897]. Disease [1649, 2776, 2752, 2595, 2599, 142, 797, 1104, 1739, 2214, 1053, 2456, 2559, 1536, 715, 2265, 387, 2975, 3193, 3007, 3094, 3112, 3297, 3301, 2373, 2000, 1074, 1111, 1397, 1746, 1297, 1369, 1437, 310, 978]. Diseases [1141, 1053]. Disenchantment [1956]. Diser [2442]. Disharmony [1043]. Disknowledge [2011]. Diskurs [2183]. Dismantling [239]. Disorder [1694, 2975, 3274, 1649]. Disordered [128]. Dispatches [2174]. Display [204, 1829, 1403, 2648, 3228, 3223, 1634, 1270]. Disputationen [2354]. Disputationes [2182]. Dispute [2744, 2854]. Disputed [1595, 3101]. Disquisitiones [443]. Disquisitions [2792]. Disrepair [535]. Disrupted [2113]. Disruptors [454, 682]. Dissection [783, 1883, 1270]. Dissemination [692, 2857]. Dissent [2892]. Dissidents [1913]. Distinction [768]. Distinguishing [1931]. Divarci [1476]. Divergence [2687]. Divergences [1107, 3043]. Diverging [1629, 3159]. diverse [2427]. diversité [779]. Diversity [1527]. Divided [1139]. Divides [2604]. Divination [552, 2697]. Divine [1628, 3225, 657]. Divining [650]. division [2350]. Divisions [1527]. Dixie [2593]. Dixon [2344]. Dizionario [1184]. Dizziness [1614]. Dmitri [2016]. Dmitry [1333]. DNA [1082, 85, 466, 394]. Do [692, 242, 1865]. Dock [2621]. Doctoring [2367]. Doctors [2047, 1151, 25, 2524, 2621]. Documentary [1497, 683, 244]. Documenti [51, 199]. Documenting [2463, 2491]. Documents [709, 2746]. Dodd [1403]. d’OEnoanda [2627]. Dog [1515, 2474, 2891]. Dogs [807]. Doig [2025]. Doing [855]. Dolev [2884]. doliorum [2716]. Dollond [1569, 3146]. Domain [2521]. Domenico [1471, 2776]. Domesticity [2326]. Dominik [46]. Dominique [1575, 1541, 2564, 1243]. Don [1501, 1068]. Donald [1069, 2628, 1229, 182, 2326, 2743]. Donato [1706, 1463]. Done [1971].

Dongen [2663, 2222]. Donkey [2050]. Donna [1824, 1904]. donnée [274]. Donner [811]. Donors [1835]. Don’t [1115]. Doody [1624]. Doogab [1914]. Dooley [1639, 2566]. Doors [897]. Dóra [2957]. Dorn [2325]. Dorothee [2337]. Dorothy [1962]. Dorr [60]. Dortmund [1087]. dottoresse [1184]. Double [2740, 1985]. Double-Helix [1985]. Doubt [521]. Doubting [1004]. Doubts [1922]. Doug [1909, 2464]. Douglas [1880, 2735, 1428, 2061, 1701, 2804]. douloureux [1499]. Dove [1687]. Dowd [2061, 1937]. Down [637]. Downey [1031]. Doyle [2948]. Draft [223]. Dragon [453]. Draht [1303]. Drake [2061]. Drawings [2596]. Dream [672]. Dreamers [2917]. Dreams [823, 712, 2923, 109, 1599, 2219, 3169]. Dresser [2728]. Drews [2799]. Dreyfus [78]. Drift [1262, 2041]. Drifting [1524]. Drinker [1792]. Dritten [806, 739, 740]. Driven [1439, 3113]. Dronamraju [2589]. Dronin [1333]. Drought [2588]. Drouin [1540]. Drucker [1824, 1904]. Drug [2112, 300, 898]. Drugs [1836, 1080]. Drunk [1792]. Dry [1633]. Dryden [825]. Dual [1755, 3285]. Duara [2090]. Duarte [2368]. Dublin [195]. Duchenne [571]. Ducheyne [1051]. Duda [1742]. Dudley [639]. Dueck [2495]. Duel [348]. Duffin [1608, 25]. Dugatkin [308, 2891]. Duke [1120, 1544]. Dumit [1080]. d’un [1477, 478]. Duncan [1663, 2544, 2271, 2107]. d’une [36, 944, 2350]. Dünkel [2002]. Dunkelkammer [3034, 1329]. Dunn [1068, 2349]. Düppe [1754]. Dupré [1942, 1409]. Duran [2104]. durch [573]. Durée [2774, 1858]. Durham [1544]. during [2240, 436, 1109, 2540, 1663, 106, 2583, 937, 619, 2593, 105, 1666]. Duris [2014, 26]. Durst [2872]. Dust [2489]. Dutch [1880, 2571, 2946, 1287, 559, 800, 2950, 2708, 2021, 1703, 3246, 2479, 3271, 2677, 311]. Duties [728]. Duzer [2602]. Dyar [2440]. Dyck [1588, 2171]. Dying [662]. Dylan [2520]. Dynamics [2156, 468]. Dynastic [2695].

E. [237, 2267, 1353, 2010]. E.-J [1061]. Each [2395]. Eadweard [2193]. Ear [1218]. Earl [1715, 3276]. Early [2512, 1585, 1052, 1694, 2339, 668, 534, 2848, 792, 1247, 1554, 1697, 796, 214, 980, 2790, 2262, 2783, 1812, 415, 2266, 1414, 1282, 2996, 1020, 2, 2188, 1689, 2091, 1868, 1474, 2017, 2020, 2150, 1021, 45, 280, 2117, 639, 1699, 271, 756, 1790, 2509, 432, 1967, 1712, 1714, 1280, 308, 1737, 547, 1556, 2709, 1696, 932, 1860, 2351, 496, 2645, 1290, 839, 876, 2204, 1053, 131, 1528, 560, 2893, 2381, 1552, 265, 1348, 2931, 1208, 1885, 277, 2780, 2096, 1168, 2636, 1645, 1044, 2305, 2507, 2635, 1770, 2758, 1559, 1557, 2335, 2231, 1884, 1918, 2345, 2633]. Early [278, 1883, 2788, 1702, 2106, 1787, 937, 1627, 2935, 451, 1593, 2256, 160, 33, 804, 2797, 1796, 2189, 1815, 2842, 1947, 2012, 1948, 650, 2782, 2849, 1679, 1024, 2704, 355, 863, 801, 2916, 2714, 1409, 1008, 3134, 3033, 3243, 3196, 3071, 3076, 3266, 3082, 3272, 3273, 3274, 3278, 3283, 3100, 3291, 3121, 3226, 3224, 2571, 3036, 838, 2977, 3207, 2547, 1817, 1648, 1671, 1973, 2152, 1672, 1204, 850, 670, 2830]. Earth [1067, 377, 1757, 2509, 91, 2227, 2863, 2915, 2518, 33, 729, 650, 981, 1603, 3133, 3098, 1248, 1669, 2062, 3216, 2196, 1660, 435, 1016]. Earth-Bound [981]. Earthquake [1215]. earthquakes [3120, 1420]. Earths [1447]. Easily [535]. East [2281, 2885, 2218, 2169, 2584, 489, 2248, 1539, 2345, 1937, 2950, 2933, 2752, 3155, 1918, 2147, 2889, 489, 364]. East-Central [2584]. Easterby [2647]. Easterby-Smith [2647]. Eastern [982, 555, 718, 1927, 1926, 184, 1924, 2743]. Eating [1593, 3021, 3121, 1353]. Eaton [1712, 3291]. eaux [1123]. Eberhard [723]. Ebert [1960]. Ebola [2895]. ebree [890]. Eccentric [2440]. Ecclesiastical [1872]. Echenberg [2864]. Echterhölter [1489]. Eckart [2591]. Eckert [1359]. Eckes [1594]. Eclecticism [2020]. Eclipse [980, 2526, 1138, 2564]. Eclipses [698]. Eco [2356]. Eco-Cultural [2356]. Ecohouse [513]. Ecole [1553, 1553]. Ecological [513, 2740, 1525]. Ecology [1295, 1309, 2605, 3148, 2062, 2679, 686, 691, 1525, 2135, 251]. Economic [215, 1619, 1253]. Economics [21, 2327, 2416]. Economies [1032, 336]. économique [689]. Economist [21, 287]. Economists [2075]. Economization [2665]. Economy [1146, 1918, 1571, 3070, 2060]. Ecosystem [862]. écrire [3250, 1553]. écrits [2181, 1797]. écriture [36]. Ecstasy [386]. Ecumenical [2685]. Ed [185, 2879, 2835, 1052, 2894, 1367, 2486, 1419, 1227, 1499, 1942, 1436, 1659, 1866, 2001, 1198, 1641, 1193, 1135, 2858, 1159, 2182, 2279, 2004, 1184, 1572, 1909, 1881, 2885, 2946, 2841, 2412, 1868, 1722, 2463, 1713, 1715, 1286, 1459, 2323, 2513, 2719, 1782, 1683, 2180, 1077, 1189, 1714, 2632, 2832, 1556, 2847, 2013, 2637, 1628, 2696, 1497, 636, 2645, 1933, 1053, 2494, 2584, 2781, 1332, 2190, 2248, 2862, 1208, 2334, 2495, 1405, 1731, 2096, 1704, 1090, 2809, 1789, 1810, 1567, 1799, 1816, 2302, 1204, 1805, 2423, 2062, 2173, 2939, 2222, 1536, 2107, 1639, 1491, 1473, 2568, 2251]. ed [1040, 2922, 2215, 1493, 2097, 1078, 2642, 1417, 1861, 2625, 1700, 1564, 1479, 1948, 1404, 1342, 1186, 1361, 1749, 2806, 1630, 2409, 2407, 2840, 2341, 2129]. ed.m [1460]. Eddington [1827]. Eddy [47]. Eden [2839, 1834]. Edens [1958]. Edens-Meier [1958]. Edgar [2803]. Edge [2539, 1544, 3110]. Edges [2945, 2817]. Edinburgh [47, 1948]. Edison [1504, 1064]. Edited [1569, 369, 1604, 1622, 1623]. Edith [639]. Editing [1860]. Editio [2100, 2182]. Edition [1198, 1698, 1370, 781, 554, 284, 1042, 2257, 2564, 1793, 1183, 2258, 490, 1871, 777, 1795, 3252, 3123, 1123, 785, 648, 779, 2101, 1219, 1827, 1476, 813, 354, 887]. Editions [1341, 1552, 1693, 1622, 3206, 3249, 3117, 1541, 1540, 1123]. Editor [1590, 1845, 26, 1776, 1778, 29, 28, 20, 1258, 1920, 2149, 2155, 2471, 2681, 2755, 420, 770, 343, 345, 247, 771, 1256, 975, 344, 1395, 248, 15, 16, 2380, 2382, 421, 1030, 2229, 418, 628, 627, 2820, 419, 342, 2230, 347, 2304, 848, 14, 1106, 17]. Editore [1556]. Editorial [2750]. Editors [1590, 31, 43, 1314, 34, 41, 2967]. edizione [1197, 3118]. Edmond [1344]. Edmund [1833]. Édouard [992, 2417]. eds [2045, 2281, 1598, 1905, 2512, 1154, 1311, 1031, 2799, 2859, 1541, 2490, 1932, 2249, 1653, 1507, 2787, 2339, 992, 2735, 2491, 2172, 2643, 1650, 1069, 947, 2454, 1344, 1554, 1942, 1341, 1975, 2790, 1511, 1225, 1634, 2356, 1481, 1402, 1139, 1110, 1983, 1577, 2007, 1038, 1654, 1954, 1783, 1416, 1698, 1160, 2730, 2019, 1192, 1788, 1600, 1362, 2837, 2317, 1408, 1470, 2503, 1710, 1054, 2519, 1474, 2945, 1132, 2446, 1647, 2061, 2346, 1112, 1624, 2591, 1068, 1802, 2117, 2295, 2699, 1790, 2211, 1720, 2746, 1086, 2952, 2509, 2953, 1338, 2697, 1733, 1727, 1695, 1644, 2411, 1401, 2663]. eds [1908, 2201, 1563, 2630, 2095, 1478, 2709, 1663, 1591, 2414, 2659, 2656, 2652, 1435, 1685, 2505, 2623, 1049, 1701, 1801, 2629, 2178, 2171, 1153, 1033, 1781, 2934, 2267, 1862, 1664, 2692, 1233, 2186, 2808, 1834, 2882, 2639, 2175, 1648, 1784, 2508, 2255, 1302, 2720, 2169, 2369, 1833, 2066, 2417, 1681, 1545, 1517, 928, 2450, 1836, 2100, 2131, 1506, 1876, 2355, 2487, 1716, 1973, 2500, 1403, 2778, 1785, 1422, 1546, 1892, 2294, 2917, 1455, 1087, 2425, 1692, 2216, 2927, 2349, 1201, 1126, 2023, 2176, 2648, 1911, 1423, 1745, 1557, 2090, 1144, 2002, 2569, 1140, 2701, 2331, 1752, 1702]. eds [1907, 2106, 1592, 2321, 1219, 1059, 1458, 1937, 1542, 1793, 1830, 2374, 1891, 2865, 1220, 1471, 2921, 1328, 1137, 2166, 1188, 1432, 2653, 1958, 2627, 2124, 1211, 2797, 2441, 2212, 1796, 1906, 2326, 1815, 1867, 1195, 2563, 2845, 2703, 1560, 2708, 2195, 2434, 1679, 2192, 1553, 1955, 2796, 2021, 1434, 1041, 2940, 1202, 1972, 1657, 1075, 2521, 2916, 1463, 1409, 2428, 2535, 2354]. eds. [1641, 1064]. eds.Penser [1210]. Eduard [1981]. Eduardo [1752]. Educated [1506]. Education [780, 74, 1505, 967, 2164, 2882, 2381, 1099, 2234, 1510, 622, 2542, 3243, 2234, 1702, 2743]. Educational [158]. Edward [2356, 2186, 2650, 1214, 1032, 2487]. Edwin [2548, 1236]. Eeuwen [2777]. Effizienz [745]. Effort [2609, 2476]. Efram [2862]. Efremenko [1333]. Efron [1738]. Efthymios [2255]. Eggert [2011]. Eggleson [2066]. Eghigian [1976]. Egil [1956]. Egle [2541]. Egypt [1340, 3053, 1296, 2877]. Egyptian [2841, 808, 1292]. Egyptologist [1962, 939]. Ehren [593]. Ehrenfest [1022]. Ehrlich [1757, 3216]. Eickhoff [133]. Eigen [2621]. Eighteenth [1649, 1942, 1568, 1109, 1413, 2019, 1788, 2188, 1408, 2346, 500, 2855, 1565, 2267, 2639, 290, 1058, 1349, 2646, 1706, 2430, 2113, 1707, 1865, 556, 2021, 2975, 3029, 3238, 3001, 3013, 966, 3010, 3220, 3170]. Eighteenth-Century [1649, 1568, 1788, 2188, 1408, 500, 2267, 1058, 1349, 2646, 1706, 2430, 2113, 2021, 2975, 3029, 3238, 3001, 3010]. Eijk [1195]. Eileen [1699]. Einheit [1881]. Einleitung [2257, 39]. Einstein [1507, 2885, 2291, 183, 714, 2049, 1063, 1364, 2734, 1623, 3173, 3175, 1822, 3262, 2049, 2980, 3014, 2123, 2734, 1899, 1428, 1508, 3114, 1623, 2947, 2878, 578, 307, 998, 313, 211, 540]. Eivind [2865]. Ekphrasis [2329]. Eksteins [1138]. El-Bizri [1198]. Elaine [1305]. Elbows [2073]. Eldredge [1910]. Eleanor [32]. Electric [1274, 2250, 531]. Electrical [3025, 3037, 1583, 1534]. Electricity [2250, 2046, 2718, 2436]. Electroacoustics [2529]. Electrodynamics [2274]. Electromagnetic [1576, 3144]. Electron [1502]. Electronic [1795]. Elegant [1788]. Elektrisiermaschinen [1134]. elektronenmicroscoop [2887]. Element [232]. elementarmathematische [3036, 1561]. elementary [3036]. Elemente [1814]. Elements [2744, 2843, 2247, 1509]. Elena [1075, 2567, 2598, 1578]. Elephants [1572, 2983]. Eli [1587]. eliocentrico [493]. Elisa [1477]. Élisabeth [2874]. Elite [1235]. Elites [487]. Elixir [2501]. Elizabeth [2800, 2217, 1694, 2266, 2373, 1758, 2623, 1234]. Elizabethanne [2709]. Ellen [2434, 1418]. Ellenbogen [1582]. Ellinor [2944]. Elliott [2015]. Ellis [1830, 2481]. Ellison [2714]. Elly [1464]. Elman [1733, 2096]. Elmer [1482]. Eloge [1393, 1392, 1394, 1531, 1675, 1454, 1453, 707, 708, 11, 847, 538, 1998, 706, 846, 13, 1775, 12, 167, 110, 166, 626, 474, 2087, 1174, 473, 1774, 537, 1175, 1327, 2772, 625, 2556, 245, 246, 168, 1928, 705]. Elshakry [1137]. Elusive [768]. Embalming [2385]. Embodied [2117, 804]. Embryo [2543]. Embryologie [678]. Embryos [3159, 2037, 1629, 141]. Emerged [637]. Emergence [71, 2352, 2873, 174, 1399, 1473, 1914, 2766, 1637, 485, 3154, 2981]. Emergency [826]. Emerson [1135, 123]. Emidio [45]. Emigration [2807]. Emil [1580, 3055]. Émile [1225]. Emilia [2355]. Emily [2356, 1708, 2861]. EMLO [2830]. Emma [1403, 1466, 1711]. Emmanuelle [2581]. Emotion [2322]. Emotions [2323, 1833]. Empire [391, 2356, 2524, 1722, 2945, 1146, 893, 2344, 2214, 2385, 1663, 2805, 2351, 2639, 881, 1050, 335, 3211, 748, 1363, 2429, 2930, 1405, 2877, 1205, 1251, 2679, 741, 2804, 1371, 2959, 1358, 2732, 747, 3040, 3004, 3160, 3088, 3221, 3127, 2032, 3084, 1167, 2026, 1425, 2381, 2780, 2653, 1635, 2842, 1938, 2192, 2724, 52, 994]. Empires [43, 2349, 2467, 1388, 2283]. Empirical [1545, 3290]. Empiricism [405, 804, 1870, 381, 1876, 49]. Empiricist [2188]. empirisme [2170]. Enchantment [8]. Encounter [1418, 2192, 3024, 3064, 1642]. Encountering [3145, 1570]. Encounters [2524, 1362, 2463, 1648, 3207, 3083, 2983]. Encyclopedia [1193, 2097, 29, 643, 501, 600, 355]. Encyclopedic [1199]. Encyclopédie [2025]. End [2466, 2895, 121, 1945, 189, 2011, 752, 1371, 1865, 3127]. Endeavor [2893]. Endocrinology [672]. Ends [2863]. Enduring [2084, 2961]. Enemy [402, 827, 187]. Enenkel [1702, 2797]. Energy [2262, 385, 2605, 313, 1825, 2463]. Engagements [1831]. Engelmann [2961]. Engelmeier [2280]. Engelstein [2727]. Engineer [1839, 886, 562]. Engineer-King [1839]. Engineering [2693, 1505, 824, 2145, 198, 390, 1889, 1914, 456, 148, 828, 3103, 2928, 1377, 676]. Engineers [3136, 1909, 2812, 1083, 1516, 1596]. Engines [37]. England [1812, 2150, 817, 193, 277, 849, 929, 278, 1270, 2849, 2658, 375, 2016, 2717, 3226, 1052, 1697, 2188, 200, 2011, 2572, 2336, 1872, 865, 560, 581]. English [1198, 1048, 938, 2017, 2446, 49, 2698, 1717, 2696, 2101, 1552, 2448, 1183, 2935, 1328, 1277, 1469, 2714, 2984, 1483, 968, 3270, 3117, 1971]. English-Arabic [2696]. English-Canadian [1328]. Englishing [2608]. Englishman [1962]. englobement [2342]. Enigma [1730, 2784, 637, 1066, 960, 3277]. Enigmatic [1206]. Enlightened [617, 2189, 3135, 1566]. Enlightening [3162, 1547]. Enlightenment [1486, 671, 1413, 123, 1207, 2344, 68, 3006, 3013, 501, 1407, 2641, 1806, 2705, 1765, 1270, 729, 983, 1353, 663, 2856, 2431, 3063, 3132, 1568, 3145, 2114, 3021, 1570, 2480, 2020, 19, 1886, 1209, 1050, 2115, 2633, 1651, 1211, 2168, 203, 126, 871, 732]. Enlightenments [2429]. Enrico [2531, 3176, 2207]. Enrique [2510, 1883, 2636]. Enros [1430]. ensayo [427]. Enslaved [1065]. Ensmenger [1330]. Enspaña [1745]. Entangled [893]. Entanglement [1425, 3160]. Entanglements [2860, 1722, 1156, 3221]. Enterprise [674, 397, 1503, 1832]. Enterprises [2349]. Enterprising [1211]. Entertainment [2269]. Enthusiasts [2812]. Entomology [2156, 2157]. Entre [374, 1493, 1470, 984, 1479]. Entrepreneurial [530, 2120]. Entrepreneurs [702]. Entry [2283]. Entschlüsselung [723]. Entstehung [67, 1212]. Entwicklung [1597, 208, 3111, 3034, 1329]. Entwicklungsarbeit [28]. Entwurfs [635]. Envelope [360]. Environment [716, 2262, 1430, 3050, 2805, 2724, 3188, 1430, 3050, 2125]. Environmental [1989, 795, 2356, 1333, 821, 2523, 676, 1581, 1933, 1034, 922, 2125, 366, 2467, 3143, 3073, 3167]. environmentalism [3062, 3006, 1568, 1661, 1088, 301]. Environments [548]. environnement [1863]. Envisioning [3231, 1708]. Envy [2141]. Enzyme [1158]. Ephemera [1910]. Ephraim [2690]. Epic [2526, 2655, 1838]. Epicureanism [2627]. Épicurisme [2627]. Epicuro [2176]. Epidemic [1074, 3237, 1750, 2957, 2961, 384]. Epidemic.} [2895]. Epidemics [1609, 978, 3009]. Epigenesis [1415, 3087]. Epigenetic [2596]. Epigenetics [2136, 2457, 2886]. Epinomide [2985, 1338]. Epinomis [2985]. Epiphany [2605]. Episcience [911]. Epistemic [2823, 324, 468, 972, 2071, 2151, 1476, 2663]. epistemische [503]. Epistemological [2084, 1614]. Epistemology [2001, 1556, 2659, 2624, 2625, 2085, 3224, 1017, 921]. EPISTLES [1198]. Epistolary [1127]. Epistolography [1556, 3224]. Epitaxy [2299]. época [1877]. Epochal [1086]. Époque [3282, 1731]. Epple [756]. épreuve [564]. Epstein [2440, 1734]. Equation [2061]. Equations [176, 334]. Equilibria [1910]. Equilibrium [993, 1754, 1637, 3241, 3154]. Equity [1099]. Era [1436, 2769, 1057, 2477, 2234, 2741, 3213, 2583, 2106, 2048, 2593]. Erasing [2492]. Eratosthene [488]. Eratosthenes [1447]. Erfindung [2347]. erfolgreichen [2287]. erfunden [1230]. Erhard [1589]. Eric [1710, 1054, 1876, 1628, 1509, 2779, 1889, 1796]. Erich [2210]. Erickson [2058, 1605]. Erik [2457, 2145]. Erika [1588, 2171, 2927]. Erin [1609]. Erkenntnis [2024]. Erlaüterung [1871]. Erlenmeyer [2473]. Erling [1232]. Ernährung [1087]. Ernan [625]. Ernst [2795, 498, 2887, 2126, 1721, 2173, 1878, 2203, 2314, 2651, 3227]. Errata [170, 346, 1260]. Erratum [18, 539]. Error [1618, 2846, 1986]. Errors [1865]. erste [1303]. Erudition [651, 224]. erudizione [51]. Eruption [1826]. Erwin [186, 664, 186, 1941, 1175]. erzeugen [635]. Erziehung [1301]. Escape [2391]. Escaping [2051]. Esoteric [1956]. espada [1939]. Espagne [1632]. España [2567, 3195, 351]. español [2110]. española [1877]. espèces [669]. Espionage [106]. espiritismo [2279]. ESS [1434, 3074]. Essais [2179, 753]. Essay [1533, 19, 250, 1534, 2433, 887, 1676, 1623, 577]. Essays [1031, 229, 206, 639, 2509, 2411, 1834, 1539, 1785, 1645, 546, 694, 579, 1458, 524, 3252, 3071, 3155, 3184, 2268, 1698, 1054, 2637, 1793, 2215, 123].

Establishing [2806]. Esteve [1126]. esthétiques [564]. Estimating [2061]. estructura [2722]. Estudio [75, 2279]. Età [2845, 2570]. Etablierung [2209]. Eternal [1248, 2928, 1536, 3193, 1910]. Ethan [2642]. Ethic [286]. Ethical [1440, 897, 3116]. Ethics [1159, 2220, 1782, 2218, 1916, 1322, 1457, 422]. Ethnic [519]. Ethnobotanical [2920]. Ethnobotany [2839]. Ethnography [2259, 2115]. Ethnology [2115]. Ethology [2]. Etiquette [1457]. etologia [954]. Ette [2799, 1219]. Ettore [447]. étude [779]. Etudes [1553, 608, 872]. Étudier [297]. etwa [2287]. Euclid [2247]. Euclidis [642]. Eugenic [2048]. Eugenics [305, 1588, 1390, 383, 2886, 2584, 60, 3048, 2945, 2872, 1268, 918]. Euler [2114, 127]. Eun [2524]. EUR [1541, 1562, 1594, 1578, 1575, 1563, 1556, 1597, 1539, 1551, 1592, 1549, 1540, 1538, 1560, 1574, 1553]. Eurasian [1919]. Eureka [2803]. Europäische [2290]. europäischen [272]. Europe [2297, 2447, 1207, 2323, 271, 432, 994, 1107, 1591, 1290, 268, 1885, 2425, 2349, 558, 875, 2423, 2633, 1473, 830, 1906, 2703, 2342, 2195, 3026, 3029, 3043, 2071, 3196, 3198, 1671, 1167, 1672, 3283, 1673, 2512, 1408, 1474, 45, 2662, 1699, 2846, 2709, 2038, 2645, 1784, 2931, 2429, 1550, 2335, 2788, 2256, 1679, 118, 547]. europea [498, 1281]. European [213, 270, 1077, 2013, 3070, 712, 2584, 437, 579, 1571, 2333, 2107, 149, 1122, 2231, 2677]. Europese [2785]. Eva [1225, 1110, 2585, 1970]. Evan [2736]. evangelischen [2226]. Evangelista [1641]. Evans [1059]. Évariste [944]. Eve [2133, 491, 1728, 1560, 2588, 3255]. Evelleen [2580]. Evelyn [2790, 686, 3235, 1701]. Évelyne [2581]. Evening [1021, 3283, 1699]. Event [2153]. Everett [1517, 1183]. Evert [2201]. Everyday [459, 1294]. Everything [2531]. Evidence [3023, 961, 2232, 22, 2152, 1612, 2285, 3076, 1954, 1379, 1627, 639]. Evolution [294, 464, 362, 446, 2837, 1724, 2649, 1656, 2532, 138, 813, 2645, 1677, 840, 3268, 2862, 1887, 2869, 2717, 2302, 2062, 1728, 2037, 2147, 57, 1749, 3229, 3061, 3254, 3255, 3280, 31, 1727, 2518, 1821, 669, 380, 2868, 2358, 1831, 2891, 139, 860, 150, 299]. Evolutionary [1193, 750, 1017, 1495, 1519, 590, 516, 1266, 428, 1595, 3101, 1016, 919]. Evolutionist [2809]. évolutionniste [1894]. Evolutionstheorie [67]. Evolving [2275, 2000, 135]. Ewald [1590]. Exact [2932, 2655]. Examination [2695]. Examiners [425]. Example [2156, 1048]. Exceptional [2519]. Exceptionalism [911]. Excercitium [2182]. Exchange [2362, 2071, 1731, 3282, 969, 2192, 3209, 1735]. Exchanges [555]. Exercise [606]. Exhibiting [762]. Exhibition [761, 1497, 1363, 2306, 65, 2192, 3088, 2983]. Exhibitions [1001, 717, 881]. Exile [2210, 531]. Existence [2918, 1380, 3079, 22]. Exoticism [1885]. Expanding [113]. Expansion [748, 991, 2025]. Expedition [3230, 2810, 2532, 3207, 1136, 1648, 1709, 1350, 729, 3081]. Expeditions [1600, 1648, 1050, 3028, 3207, 1036]. Experience [1127, 172, 2801, 524, 1613, 1870, 323]. Experiencing [2013]. Experiential [2071, 2507]. Experiment [2408, 2360, 797, 172, 2218, 2171, 1411, 2430, 2305, 1587, 1666, 3068, 3069].

Experimental [1218, 214, 1408, 237, 290, 2118, 2564, 1614, 2474, 160, 2521, 3029, 1494]. Experimentalphysik [733]. Experimentalphysiker [2287]. Experimentation [409, 753]. Experimenter [336, 242]. Experimenters [1614, 2558]. Experimenting [1160]. Experiments [1218, 225, 1274, 855, 2083, 624, 1606, 1828, 1063, 2274, 2293, 1099, 1091, 3168, 3057, 1684, 2543, 2056]. Expert [1168]. Experten [2504]. Expertise [231, 432, 2201, 1663, 2152, 956, 829, 366, 597, 676, 827]. Expertos [859]. Experts [858, 2169, 402]. Explained [2759]. Explaining [2770, 2758, 2756, 2764]. Explanation [2769, 2760, 2761, 2765, 2763, 550]. Explanations [1311]. Explicit [609]. Exploitation [1032, 1512]. explorateur [2581]. Exploration [1052, 899, 1223, 1249, 1150, 1512, 600, 2041, 2028, 1089, 3265, 2145, 2172, 2863, 1756]. Explorations [1109, 2952]. Exploratory [2274]. Explored [2521]. Explorers [2361]. Exploring [2064, 757, 935]. Exposed [318]. Expositions [204, 2074]. Exposure [2817]. Exposures [2193, 2592]. Extended [1462, 2093]. Extensions [2828]. External [2625]. Extinction [2026, 251]. Extraordinary [2022, 447, 810, 1113]. Extraterrestrial [2328, 2061, 177]. Extremes [756]. Eye [1020, 74, 1349, 1506, 3001, 986]. Eysenck [517].

F [2281, 1228, 443, 2361, 1307, 2643, 1441, 2813, 2503, 2061, 1351, 2623, 1781, 2115, 1555, 1721, 228, 1820, 2507, 1205, 1083, 1823, 2492, 675, 1825, 1369, 754]. F. [1793, 3227]. Fabian [1807]. Fabien [2569]. Fabri [934]. Fabric [1622, 3249]. Fabrica [1622, 2927, 3249]. Fabricating [103]. fabrication [944]. fabrique [571]. Fabrizio [1471]. Face [2871, 2678, 1264, 453, 1157, 1203, 1551, 3124]. Face-to-Face [2678]. Faces [392]. Facetten [1188]. Facil [2249]. Facing [3048, 2884, 1588]. Faconner [153]. Facsimile [1370, 1793, 3123, 490]. Facsimiles [2773]. Fact [352, 2251]. Factish [831]. Factory [1869]. Facts [668, 227, 1304, 692]. Faidra [1998]. Failed [690]. Failure [1658, 997, 1224, 663, 3089, 1117]. Fair [2306]. Fairchild [683]. Fairs [2133, 1510]. Fairy [2040, 1266]. Faith [2104, 724, 1793, 693, 3187]. Fakes [3134, 1554]. Fakultäten [2504]. Falchetta [1406]. Falcon [1467]. Fall [816, 194, 2144, 402, 2453, 515, 925, 238]. Fallacy [1323, 1324]. Fallout [950]. Fallwell [1584]. False [2245]. falso [890]. Familienkorrespondenz [2651]. Families [1694, 2431, 3274]. Fancy [436, 1340]. Fang [1151]. Fantasy [1322, 2990, 1500]. Fantoni [1461]. Far [2885, 227]. Far-Fetched [227]. Fara [2725]. Faraday [2274, 3144, 1576, 210, 1000]. Farewell [2750, 1317]. fari [498]. Farmacopea [2570]. Farmelo [207]. Farming [2224]. Farnham [1569, 1581, 1558, 1567]. Fascism [2807, 2306, 2534]. Fascist [2534]. Fashionable [1649, 2975]. Fashioned [1287]. Faster [1364, 3014]. Fatal [2146, 1488]. Fate [1788, 2456]. Fates [585]. Father [2531, 810]. Fathers [982]. Fault [2687]. Fauna [1423, 3005, 3005, 1423]. Favino [2185]. Fay [2246]. fazhan [223]. Fco [1983]. Fe [514, 1939]. Fear [1070, 1192, 900].

Fearful [3026, 2579, 1550]. Feathery [1003]. Features [2871]. Februar [2651]. February [3218]. Federal [619]. Federica [2185]. Federico [2015, 2791, 1417, 1947]. Fee [2373]. Feeble [2048]. Feeble-Mindedness [2048]. Feeding [3292, 1711]. Fehige [2248]. Feichtinger [2940, 2535]. Feingold [2709, 1676]. Feist [1562]. Feldhay [2503]. Feldman [2463]. Felfe [1814]. Felix [1375, 3207, 3131, 1648]. Female [1601, 428, 3102]. Feminism [1737, 2139, 1156, 3273]. Feminisms [94]. Fengler [1233, 1972]. Fenny [1478]. fenómenos [2279]. Ferber [1139, 2544]. Ferdinand [2013]. Ferguson [1207, 1431]. Feringa [2887]. Fermat [1459]. Fermi [88, 2531, 2207, 3176]. Fermilab [84]. Fernández [1883]. Fernando [2927, 2620]. Ferngren [1535]. Feros [1790]. Ferrari [1338]. Ferriello [2331]. Fersman [1919]. Fertility [1625, 3156]. Féry [1459]. Fetched [227]. Feuerhahn [1575]. Fever [384, 2189, 381, 3256, 1678]. Feverish [2189]. Few [428]. Feynman [2731]. Fiasco [2768]. Fichant [2441]. Fichte [675]. Fiction [1585, 1077, 1786, 109, 162, 1300, 2251, 2986, 1828]. Fictions [930]. Fiege [1034]. Field [2835, 1036, 2829, 2549, 1576, 548, 319, 1325, 2876, 2327, 2284, 2764, 3144, 2583]. Fields [1305, 290]. Fieldwork [1036, 334, 1460]. Fifteenth [2835, 861, 2557, 2503, 629, 725]. Fifteenth-Century [629, 725]. Fifth [188, 1043, 117]. Fifth- [188]. Fifth-Century [117]. Fifty [2175, 891]. Figala [167]. Fight [2111]. Fighter [2618]. Fights [957]. figs [1569, 1554, 1677]. Figure [1206]. Figuren [2024]. Figures [2485, 2660, 1264]. Filippo [1075]. Fillafer [2940]. Filling [1011]. Film [2491]. Films [2882, 1786]. filo [2570]. filosofare [3248, 1700]. Filosofia [2100, 1877, 727, 2185, 870]. Fin [2873, 2038, 171]. Fin-de [2038]. Final [2912, 2086, 2691]. Financial [938, 1619, 2362]. Finches [714, 2980]. Finding [3241, 268, 1771, 306, 1754]. Findlen [2934, 1204]. Fine [494, 2678, 2332]. Fine-Grained [2678]. Fingerprinting [85]. Finkelstein [1580]. Finland [1002]. Finlay [1327]. Finn [2660, 1308, 1975, 1741]. Finnegan [2192]. Finocchiaro [1480]. Fiorenza [1474]. Fire [22, 517, 258]. Firebrace [2666]. Firenze [1461]. Fireworks [437]. First [2835, 1880, 1600, 31, 1591, 1642, 2915, 2066, 1467, 2225, 2564, 1265, 2954, 2725, 1571, 1431, 1089, 1291, 1515, 3028, 861, 3054, 3064, 523, 3070, 221, 3198, 1024]. Fischer [2339]. Fish [1274, 997]. Fisher [2046]. Fisheries [1253, 997]. física [3195, 1745]. fisiche [1184]. fisico [2276]. Físicos [1745, 3195]. fisiologia [1061]. Fiss [2133]. Fit [1857, 2804]. Five [1999, 1693, 1089]. fixe [2184]. Fixierung [573]. Fixing [598]. Flach [2007]. Flachowsky [2454]. Flament [1541]. Flamsteed [2409]. Flandreau [2362]. Flask [2473]. Flatland [284, 284]. Flaubert [2660]. Flavia [2850]. Fleck [2209, 2391]. Flecken [2290]. Fleeing [585]. Fleming [2128, 1681]. Flesh [66]. Flight [2812]. Flights [2225, 2576]. Flipse [1732]. Flood [1341, 1954, 1140, 1115]. Flora [2015]. Floracrats [663]. Florence [1346, 2807, 1560, 3017, 646, 776]. Florentine [1201]. Flores [22]. Florian [2454, 1498, 2292]. Flow [2823]. Floyd [2952, 1410]. Floyd-Wilson [1410]. Fluid [360]. Fluids [1110].

Flüssigkristalle [2124]. Focaccia [1184, 2276]. Focus [137]. Folding [2836]. Folk [2530]. Folkerts [2425]. Folklore [2051]. Follies [1550, 3026]. fondazione [493]. Fong [174]. Fong-ching [174]. font [1683, 3215]. Fontaine [1905]. Fontana [1695, 928]. Food [2054, 3187, 1593, 1353, 2375, 1711, 3021, 3292, 3121]. Footsteps [2864, 1508]. Forbes [1576]. Force [2898]. Forces [1071]. Forecasting [2403]. Foreign [2306, 253]. Forest [366]. Forests [956]. Forge [1109]. Forged [1155]. Forgery [1138]. Forget [1005, 969]. Forgetting [2645]. Forging [896, 3018]. Forgive [1117]. Forgotten [286, 2047, 1194, 840, 1985]. Form [2686, 678, 1208, 164]. Formal [2587]. Formalization [1672]. Formalizing [459]. Forman [883]. Formation [303]. Formative [2701, 2734]. Formen [2343, 438]. Forms [2346]. Formulas [2071]. Forrester [2334]. Fors [1886]. Forschung [1145, 811]. Forschungs [28]. Forschungs- [28]. Forschungsfeldes [1233, 1972]. Forschungspraxis [2454]. Forschungstechnologien [1332]. Forster [2799, 2348]. Forster-Humboldt-Chamisso [2799]. fortificazioni [1276]. Fortschrittsfeinde [1310]. Forward [1860, 2729]. Fossil [703, 1519, 1066]. fossile [942]. Fossils [2154, 58]. Foundation [1555, 619]. Foundations [2128, 1590, 362, 194, 1098, 1283, 2263, 1239, 710, 2655, 405, 1613, 1913, 2438, 2421, 3066]. Founded [1214]. Founders [1838]. Founding [1460]. Four [139, 2401, 853, 31, 542, 746, 1816]. Fourteenth [2102]. Fourteenth-Century [2102]. Fourth [2873, 188, 1870]. Fourth-Century [188]. Fox [1052, 1062, 1455, 2364, 2891]. Foxon [2299]. fra [2324, 1406, 2972, 191, 725]. Fragile [234]. Fragmented [2905]. Fragments [1275, 2389]. Frames [688]. Framework [1615, 1529, 2823]. Frameworks [2004, 934]. Framing [2048]. francais [669, 191, 875]. francaise [2177, 1882, 380]. France [442, 1812, 2875, 122, 68, 817, 505, 1632, 2131, 812, 1894, 949, 1358, 2647, 1711, 1797, 3040, 3273, 2761, 3292, 3115, 2269, 1737, 1062, 2033, 2641, 2794, 2856, 2431, 494, 660]. Frances [1305]. Francesca [1338, 2418]. francescana [3041, 1630]. Franceschi [1867]. Francesco [931, 2176, 1041, 1109, 1803, 2937]. Francia [1652, 3115]. Francine [2281]. Francis [872, 1800, 2341, 1569, 3146, 2305, 1383, 1987, 403]. Franciscan [3041]. Franciscans [2501]. Francisco [2620, 1070]. Franck [296, 1470, 1061, 2181]. Franco [1578, 1338]. Francois [1061, 2030, 651]. Francois-Franck [1061]. Francoist [1596, 3136]. Franconi [2298]. Frängsmyr [2484]. Frank [210, 22, 1795, 1124, 1833, 267, 705]. Frankenstein [652]. Frankfurt [1561, 1365, 3105]. Franklin [2408, 1606, 1199, 2460]. Franklin-Brown [1199]. Frankreich [743]. Fransen [2797]. Franz [1730, 272, 1592, 2035, 2940, 3277]. Franzmann [2212]. Fraser [2172]. Frate [1641]. Fraud [352, 2037]. Fred [619, 2959]. Frédéric [2697]. Frederick [2483, 1720, 1448, 2868, 1059]. Frederickson [1957]. Fredrik [1568]. Free [404, 3009, 1609, 1176]. Freedom [1273, 2447, 3248]. Freezing [73]. Freiheit [2668]. Freiherr [238]. Freire [1913]. Freitas [2438]. Fremde [740, 2209, 2337]. French [2527, 805, 2030, 2852, 1572, 1407, 991, 818, 799, 3132, 3030, 2997, 3139, 3197, 2154, 2983, 2239]. Fresh [262]. Fressoz [2227]. Freud [1659, 1930, 379, 2949, 2992, 2874, 2949]. Freudenthal [1789]. Freudian [2660, 832]. Freuds [2449]. Freund [1141]. Frieden [940]. Friedman [2836]. Friedrich [498, 2107, 288, 2274]. Friend [42]. Friends [746]. Frietsch [1631]. Friis [1648]. Fringe [2659, 1010]. Frisch [2208]. Frisius [190]. Fritz [2275, 2533, 857]. froide [2295]. frondeur [785]. Front [2600, 2675, 2747, 2818, 2896, 2964, 352, 295]. Frontier [1568, 88, 84, 1512, 892, 891, 3006, 1223]. Frontiers [1032, 2364, 2142]. Frontispiece [100, 157, 236, 332, 408, 463, 529, 615, 696, 760, 837, 906, 964, 1019, 1096, 1097, 1165, 1246, 1316, 1382, 2936]. Frozen [2467]. FRS [2341]. Frühen [46, 3036, 1561, 1310, 2925]. Früher [1472]. Frühneuzeitliche [2354]. frühneuzeitlichen [1562, 2977, 1807]. Fruits [2704]. Frumer [2737]. Fuchs [1081]. fucina [1109]. Fuechtner [2735]. Fuel [1009]. Fueling [2283]. Fuller [2281]. Fullilove [2518]. Fulvio [2845]. Fundamental [1897, 101]. Fungal [2595]. Funkenstein [93]. füreinander [2053]. Fürstenwillkür [124]. Furukawa [2881]. Fusang [1765]. Füssel [2699]. Füßl [2657]. Future [2829, 832, 605, 2485, 3216, 2360, 1757, 1207, 610, 1861, 1235, 234, 148]. Futures [1021, 162, 2407]. Futurism [1009]. Fuzzification [304]. Fuzzy [304]. Fye [2036].

G [1502, 2795, 2440, 2210, 1543, 1127, 2945, 1647, 1306, 1435, 1862, 2919, 3103, 2130, 1431, 1510, 2653, 1827, 397, 2292, 675, 60, 686, 495, 1377, 1564, 1948, 686]. G. [2795, 1407]. Gábor [1973]. Gabriel [2119, 1580]. Gabrielle [2704]. Gacek [1793]. Gad [1789]. Gadebusch [1782]. Gaffarel [1805]. Gaia [1334, 3104, 79]. Gaian [2062]. Gain [690]. Galan [2369]. Galantuomo [2276]. Galápagos [2532]. Galbraith [1701]. Galen [2632, 2696, 1183, 1635, 1195, 1042, 3084, 722]. Galilaeana [1560]. Galilée [2170, 786]. Galilei [199, 1417, 3118, 727]. galileiana [3118, 1417]. Galileo [1047, 730, 1048, 199, 727, 2513, 2108, 1480, 2915, 2511, 1803, 75, 1808, 2713, 1417, 644, 1281, 1813, 3248, 3085, 3118, 1125, 1351, 1943, 2791, 1700, 2792, 565, 798, 562, 475, 789, 75, 260]. Galina [2734]. Gallaccini [1047]. Galluzzi [2791, 1700]. Galois [944]. Galton [1582, 2998]. Galvani [790]. Gamble [1757, 3216]. Gambled [2049, 3173]. Game [2058, 389, 1129, 2536, 507, 623, 2993]. Games [1325]. Ganot [817]. Gantt [2474]. ganzen [990]. gao [223]. García [2636, 1082]. García-Sancho [1082]. Garden [196, 1060, 2013, 633, 1865]. Gardens [290, 2339, 2805]. Gareth [1478]. Gargano [2355]. Garner [951]. Garrison [1622]. Garrod [2931]. Gartenlaube [1498]. Gary [1953, 1535, 1953]. Gas [2592]. Gascoigne [1570, 1133]. Gaslight [983]. Gasper [1341]. Gasperini [1357]. Gaston [608]. Gastric [815]. Gatekeepers [1450]. Gates [426]. Gatterer [990]. Gatto [2331]. gauche [1894]. Gaudillière [1911]. Gaukroger [2116, 1131, 2428]. Gausemeier [593]. Gauss [443]. Gautrey [846]. Gave [933]. Gavin [2562]. Gavroglu [2195]. Gavrus [2652]. Gay [1309, 2277].

Gaycken [1967]. Gaze [1349, 3001]. Gdańsk [1799]. Geburt [2517]. Geburtsjahr [2187]. Geburtstag [288]. Gee [1569, 1466]. Geert [1591]. gegen [2226]. Gegenstücke [2143]. Geheimcode [2517]. geheimen [1145]. Gehirn [2978, 1398]. Gehrke [2178]. geistes [3151]. geistes- [3151]. Geisthövel [2778]. Geld [2126]. Gelehrte [53, 2035]. Gelehrten [2591, 2354]. Gelehrtenkultur [2699]. gemalte [2990, 1636]. Gemma [492, 190]. Gen [2881]. Gen-itsu [2881]. GenBank [336]. Gender [1714, 1148, 1410, 2804, 2073, 1867, 3056, 3266, 1189, 864]. Gendered [1247, 214, 1836]. Gene [394, 2065, 2738, 1832]. Genealogical [2727, 921]. Genealogie [1631, 3058]. Genealogies [7]. Genealogische [1489]. Genealogy [2319, 2065, 2370, 3058, 1780, 2800]. Genentech [894]. General [2947, 1615, 2540, 2656, 531]. Générale [158, 1485, 369, 2703]. Generality [2321]. Generation [307, 2019, 1014, 275]. Generic [2057]. Generizität [1332]. Generizität-Interstitialität-Transfer [1332]. Genes [1514, 685, 1079]. Genese [2200, 2350]. Genesis [22, 81, 2115, 1887, 304, 1838]. Genetic [71, 1514, 2599, 310, 2460, 256, 1985, 1914, 1437, 3094, 1311, 256]. Genetics [2752, 2387, 3060, 591, 2738, 3233, 1542, 2599, 1725, 596]. Genetik [510]. Geneviève [2008]. génie [1499]. Genius [3176, 2114, 2161, 2372, 447, 527, 1508, 2207]. genome [3102, 1601]. Genomes [312]. Genomic [1391]. Genomics [2669, 1752, 3300]. Genre [624, 1867]. Gent [1375]. Gentleman [2966, 409]. Genus [2488]. géo [171]. géo-hèliocentrique [171]. Geodetic [2784, 1612]. Geoff [1328]. Geoffrey [1822, 2267, 1497, 2186, 2525]. Geographic [201]. Geographical [2125]. Géographie [178, 1412]. Geographies [2622, 893, 2780, 1144, 667]. Geography [2386, 577, 869, 2178, 2900, 2397, 1885, 1573, 2396, 54, 201, 3150, 3078, 3017, 2172, 1346, 488]. geologia [205]. Geological [359]. Geologist [1115]. geologo [1109]. Geology [1712, 1743, 2400, 3289, 3291, 1300, 180, 353]. geometriam [642]. géométrie [191]. Geometry [1198, 1868, 2628, 2263, 2720, 2716, 856]. Geomorphology [353]. Geophysical [2213]. geophysics [3177, 2213, 2052]. Geopolitics [1074, 1660, 1388, 3133]. Georg [498, 2449, 2287, 1412]. George [11, 1724, 512, 509, 1368, 2169, 2060, 2029, 369, 514, 1722, 1739, 968, 1057, 369, 2973, 2902]. George-Louis [369, 369]. Georges [2920, 1485, 3280]. Georges-Louis [1485]. Georgia [1609, 3009, 3178]. Georgian [2858]. Georgiana [2340]. Georgina [2051]. Geppert [2894, 1077]. Gerald [2066, 1893, 2048]. Gerali [1109]. Gerard [1906]. Gere [2619]. gereformeerden [3246, 1732]. Gerhard [1201]. Gerhardt [2939]. Gerlach [2287]. Gerlandi [1871]. Germ [2338]. German [1425, 2798, 2117, 1302, 2864, 2115, 128, 1584, 2278, 201, 873, 2704, 2990, 3034, 3036, 3049, 3052, 2977, 3058, 621, 3150, 3151, 3003, 3008, 3210, 2978, 2971, 3160, 2473, 1830, 3111, 2593, 3130, 391]. Germana [2314]. Germania [723]. Germans [2278, 2888]. Germany [2524, 1960, 585, 2203, 1435, 607, 452, 3199, 1721, 3166, 2251, 1007, 617, 1575, 1580, 1524, 3055, 133, 3157, 2152, 1559, 3288, 3227, 2860, 2630, 2204, 2444, 1729, 2654, 1312, 1976, 129, 212, 1280, 650, 996]. Gero [2711]. Gerontologists [2206]. Geroulanos [2814]. Gerovitch [1909]. Gersonides [2102]. Gertraud [2810]. gesammte [2190]. Geschichte [940, 1597, 1332, 2778, 1301, 3111, 46, 2522, 1188, 2124, 990]. Geschiedenis [309, 953]. Geschlechterordnungen [510]. Gesellschaft [658, 1188, 288, 2035]. gesellschaftlicher [745]. Gesta [1802]. Gestalt [2948]. Gestation [2798]. Gesundheitspolitik [378]. Get [755]. Getting [1326]. Gezeigt [1212]. Ghāfiqī [1793]. Ghana [1099]. Ghazāl [267]. Ghost [1515, 388, 523, 2136]. Ghosts [251]. Giacomo [1646, 3003]. Giacomotto [1553]. Giacomotto-Charra [1553]. Giambattista [2787]. Giancarlo [1474]. Gianfrancesco [1813]. Giant [308, 506]. Giard [2]. Gibbs [2346]. Gibson [2225, 2113]. Gideon [800, 1208]. Gielen [2927]. Gierl [990]. Gifford [1827]. Gift [2200]. Gifted [2234]. Gilain [2023, 1459]. Gilded [502, 766, 1728, 3255]. Giles [1341]. Gillen [1826]. Gilles [2500]. Gillette [2584]. Gillian [2437]. Gillin [2650]. Gillispie [1774]. Gilmore [2217]. Gimbel [2123]. Gindhart [2354]. Ginés [706]. Gingerich [2003]. Gingras [2092]. Ginoux [2875]. Ginther [1341]. Giocondo [191]. Giora [2787]. Giordano [2713, 2989, 2700, 1339]. Giovanni [1109, 1943, 2185, 2850, 2929, 2216]. Girls [202, 2739, 1505]. Girolamo [2334]. Gisela [2523]. Gissis [1110]. Giuseppe [1471, 2207, 2845]. Giuseppina [2331]. Glare [395]. Glasner [2102]. Glass [482, 1449]. Glassmaking [482]. Glazer [2045]. Gleitsmann [1310]. Glen [1194, 1042, 33]. Glenn [2317]. Glimpses [606]. Global [1523, 1389, 521, 3133, 2735, 2439, 2895, 653, 1159, 821, 225, 1674, 2317, 1835, 604, 518, 893, 585, 1966, 955, 676, 1908, 1781, 548, 2518, 335, 1610, 2603, 912, 283, 971, 1607, 175, 1971, 9, 10, 1088, 2212, 7, 2687, 2493, 1369, 2192, 2940, 381, 3044, 3152, 3112, 1660, 2445, 2809, 2383, 1863]. Globalism [1885]. Globalization [2412, 2629, 982]. Globalizing [1104]. Gloria [1038]. Glory [2526, 2863]. Glosae [35]. Glow [2205]. Glyn [1337]. Glynn [2689]. Go [283]. Go-Betweens [283]. God [796, 654, 1268, 22, 547, 1157, 2708, 3030, 2003, 1801, 2066, 57]. God-or [57]. Gödel [2591, 710, 2215, 2247]. Godin [2561]. Godless [1501]. Gods [364, 2923, 831]. Goeing [45]. Goering [1341]. Goes [260]. Goethe [433, 873]. Gogh [1138]. Gold [1604]. Golden [559, 2932, 3271, 2445, 2708, 1703, 916]. Goldstein [443, 1624, 2131, 1122]. Golinski [2118]. Gómez [1983, 1836, 2507]. Gonzaga [731]. Gonzales [202]. González [2510, 1654]. Gonzalo [1038]. Good [2408, 809, 2619, 3116, 428, 1440]. Goodale [1931]. Gooday [2801, 1583]. Gooding [168]. Goodrich [1432]. Goods [2823]. Goodstein [2947]. Gorbachev [2815]. Gordin [1605, 1971]. Gordon [2193, 1600]. Gorgeous [979]. Gorham [2186]. Gorilla [57]. Gorlæus [1206]. Gosse [1492]. Gosselin [2502]. Gosztonyi [1328]. Gotta [242]. Götter [811]. Gottfried [124]. Gotthilf [1878]. Göttingen [370, 1574]. Gottorf [1448]. Goudsmit [2451]. Gould [150, 687]. Gouverner [252]. Governance [2859, 1159, 3044, 1610, 1957]. Governing [1114]. Government [2214, 2137]. Governor [800]. Governs [2395]. Govoni [1867]. Gowan [1653, 1137, 2196]. Goy [1710]. Goyens [2927]. Graboyes [2218]. Grace [1743, 143]. Gradual [527]. Graduate [1545]. Graeme [1583, 2801]. Graff [1977]. Grafton [24, 2317, 45]. Graham [2271, 2136, 1458, 1331, 1465, 207]. Grained [2678]. Granada [1698, 2509, 1479]. Grandidier [2581]. granduchessa [727]. Graney [1943, 2792]. Grant [2806, 305]. Granting [647]. Graphics [2411]. Grattan [179, 1675]. Grattan-Guinness [179, 1675]. Grave.J [274]. Gravid [2413]. Gravitation [2540]. Gravitational [2671]. Gravity [961, 2052, 2915, 2225, 3177, 1515, 2671, 523]. Gravity-Breaking [2225]. Gray [2194, 1216]. Graz [789]. Great [885, 2814, 761, 176, 1734, 1107, 1074, 2022, 1497, 2325, 2312, 1066, 1899, 2687, 2587, 3043, 2976, 3244]. Greater [2619]. Greatest [1985]. Greatrakes [1482]. Greco [2841, 808, 2494, 2169, 1322, 3211, 1405]. gréco-romain [3211, 1405]. Greco-Roman [2841, 2494, 2169]. Greece [117, 188, 926, 485]. Greek [2329, 982, 2330, 551, 2178, 1042, 483, 2843, 429, 977, 476, 2922, 364, 1195, 2546]. Green [2224, 301]. Greene [2057, 626, 2041]. Greenwald [2808]. Greenwood [2047]. greffe [571]. Greg [1931, 1976]. Gregg [2491]. Grégoire [274]. Gregor [1086, 70]. Gregorio [2513]. Gregory [1031, 2740, 60, 2203]. Grell [1481, 2508]. Greti [1341]. Grevsmühl [1863]. Griffel [267]. Griffith [2277]. Griffiths [2577, 1542]. Grit [2613]. Grob [1893]. Groningen [2777]. Groot [1870]. Grootenboer [1349]. Groovy [2374]. Gross [1543, 1833, 2919]. Grosseteste [1341, 3206, 3093]. Großherzogtum [2343]. Ground [289, 978]. Group [919, 1235]. groupes [1594, 3141]. Grove [2966]. Growing [134]. Growth [229, 1519, 1088, 355]. Gruber [1311]. Grundlagenforschung [441]. Gruter [39]. Grzesik [2182]. Guardo [1802]. Guarnieri [2807]. Guatemala [773]. Guenter [1070]. Guenther [2370]. guerre [171, 2295]. Guerrini [1949, 2015]. Guicciardini [2853]. Guide [2826, 182, 2983]. Guidebook [1480, 2060]. Guidobaldo [993]. Guilbaud [2023]. Guillaume [2502, 2569, 2703, 785, 867]. Guillelmi [35]. Guillou [2502]. Guillou-Kouteynikoff [2502]. Guinea [254]. Guinness [179, 1675]. Guinter [2632]. Gulf [2147]. Gülhan [1354]. Gundlach [1656]. Guns [2131]. Günter [2571, 1145]. Günther [1310]. Güremen [2627]. Gurion [1839]. Gurka [2190]. Gustav [2446]. Gutachten [1132, 2778]. Gutenberg [426]. Gutfreund [2734]. Güttler [1574]. Gyatso [1787]. Gyung [2429].

H [1036, 2643, 1227, 1416, 2946, 2798, 1214, 2460, 2863, 946, 1781, 1190, 2372, 2133, 2169, 1518, 1545, 2954, 1423, 1557, 1651, 1622, 2959, 1827, 397, 1375, 2732, 2012, 2089]. Håkon [349]. Habana [136]. Haber [857]. Habinek [2935]. habits [3008]. Hächler [272]. Hachtmann [2454]. Hacking [1906]. Hadass [2849]. Hadland [1589]. Haeckel [2651, 2037]. Haffner [1185, 1636]. Hagar [1897]. Hagner [1301, 1186]. Hahn [1286, 789, 1188].

Haigh [1154]. Hairy [202]. Halcyon [1587, 3069]. Haldane [2589, 596]. Hale [2717]. Hall [1735, 13, 1985, 3209]. Hallam [1439]. Halle [498, 2020]. Haller [1753, 2949, 370, 1352, 3092]. Hallett [1590]. Halleux [2008]. Hallucinogen [1152]. Halo [891]. Halpern [2731, 2221, 1899]. Hamanaka [2024]. Hamblin [1661]. Hamilton [1749, 2556]. Hamish [2945]. Hamlin [1678, 1728]. Hammer [1043, 269]. Hammerstaedt [2627]. Han [2115]. Hancock [2267]. Hand [1426, 811, 3091]. Handbook [2790, 261, 1933, 918, 1455, 2321, 2921, 1015, 3151]. Handbuch [3151, 1422]. Hände [28]. Handful [521]. Hands [2286, 445, 2796]. Hands-On [445]. Haney [1555]. Hanford [892]. hangtian [223]. Hanna [1072, 2280]. Hannah [1341, 1309, 2277]. Hanneke [1349]. Hanoch [2734]. Hans [593, 1086, 1435, 2178, 1784, 1071, 2715, 1087, 579, 1589, 517, 708, 2738, 2950, 2785, 1181]. Hans-Erhard [1589]. Hans-Joachim [2178]. Hans-Jörg [593, 1435, 579, 2738]. Hans-Ulrich [1087]. Hansen [1651, 2278]. Hanson [525]. Hanspeter [2354]. Happened [2753]. Happiness [1306]. Happy [83]. Harald [2422]. Harbsmeier [1036, 1648]. Harden [1433]. Hardesty [2025]. Harding [1090]. Hardy [2562, 1968]. Häresie [1631, 3058]. Hargittai [2060, 1867]. Harland [2464]. Harley [1900]. Harman [1598, 2917]. Harmon [1543]. Harmonic [483]. Harmony [2789, 180, 2702]. Harness [464]. Harold [2325, 1545, 2712, 1409]. Harper [2537]. Harrassowitz [1575]. Harriet [1363]. Harriot [1052, 1048]. Harris [619]. Harrison [2440, 1111, 1934]. Harry [1515, 2671]. Harshad [2047]. Hart [1642]. Hartlib [1448]. Hartmut [1657]. Haru [2024]. Harvard [1579, 1548, 1602, 1355, 2642, 3035, 3090]. Harvesting [82]. Harvey [1182, 1977]. Harwit [1396]. Haskell [1053]. Hasok [525]. Hasse [2426]. Hast [1622]. Haster [1795]. Hastings [64]. Hatch [2375]. Hated [2049, 3173]. Hauslehrer [1301]. haut [1863]. Haven [1568, 1610]. Hawking [2049, 1093, 3173, 1093]. Hayden [2409]. Haydn [2352]. Haynes [2735, 2618, 2804]. Haytham [2564]. Hayton [2105]. Hazel [1362]. Healers [1280, 1458, 1938]. Healing [2249, 2530, 1112, 2443, 1482, 2507, 2288, 1687, 2563, 25, 117, 3279, 718]. Health [878, 805, 2617, 821, 2373, 2592, 1306, 1104, 638, 1968, 2443, 1664, 479, 265, 1348, 864, 1065, 1935, 1729, 852, 175, 2741, 1156, 2493, 1469, 3201, 3033, 3288, 2373, 2882, 2005, 1080, 2373, 2098, 588, 204]. Healthcare [2301]. Hearing [2801, 159]. Hearnshaw [1726]. Heart [531, 2036, 1079, 2593]. Heartland [2530]. Hearts [2890, 318]. Heat [2146]. Heather [2117, 452, 1649, 1830]. Heaven [2509, 268]. Heavenly [1194, 2782, 641]. Heavens [377, 1487, 86, 33, 3067, 1360]. Hebrew [2884]. Hedesan [2340]. Heereswaffenamtes [1145]. Heesen [1744, 1955]. Hegels [806]. Hegemonic [2235]. Hegemony [2298]. Heidarzadeh [27]. Heidegger [2918]. Heidelberg [1269]. Heidemarie [2535]. Heike [73]. Heiko [1402, 1158, 2200, 1134]. Heilbron [1741, 2089]. Heinrich [2810, 1862, 1878, 943]. Heisenberg [56, 250, 176]. Heit [1422]. Helen [2518, 1679, 1442, 1627, 1593]. Hélène [1093]. Helge [1975, 1973, 1907, 50, 2744, 2412, 2052, 2132, 2435]. hèliocentrique [171]. Heliocentrism [1842, 780]. Helix [1985]. Hell [532]. Heller [1422, 1043]. Heller-Roazen [1043]. Helmholtz [511, 856, 1941, 1820, 70, 1528, 2481, 671]. Helmholtz-Kurven [511]. Helmont [2340]. Helmut [1422]. Helped [2895]. Helsvig [2865]. Helvoort [2887]. Hemispheres.} [2834]. Hemmungs [1970]. Hemophilia [715]. Henderson [2694]. Hendriksen [1788]. Hendryk [2699]. Henk [2519, 2708]. Henke [82]. Henneveld [1339]. Hennig [401]. Henning [593, 1980, 1820]. Henri [1650, 2883, 1216]. Henrietta [2934]. Henrik [2490, 1784]. Henrika [1392]. Henrique [2630]. Henry [2879, 2022, 1756, 742, 1549, 939, 2728, 1117]. Hentschel [2446, 1302, 1332, 2124, 28]. Herbal [1793, 1342, 2926, 2987]. Herbert [2535, 1230]. herbes [1123]. Herbs [1625, 3156, 1458]. Hercberg [1677]. Herds [2732]. Hereditary [2872]. Heredity [2868, 2886, 1224, 2521, 2644, 1006]. Hereford [2698]. Heresy [2031, 2898, 3058, 1468]. Heritage [2163, 174]. Herman [1318, 2663, 1056]. Hermann [1594, 3141, 1974, 1500, 70]. Hermaphrodite [1004]. Hermaphrodites [967]. hermetisch [2257]. hermetisch-allegorischen [2257]. Herodotus [1626, 3158]. Heroes [348]. Heroic [1436, 2161, 741, 3213]. Héron [2177, 2331]. Herran [1745]. herrliche [586]. Herrmann [1646]. Herrschaftssystem [2454]. Herschel [1360, 1487, 3067, 2866, 576]. Hertha [1005]. Hertz [943]. Hesketh [2661]. Hess [2778]. Hessen [2795]. Heumann [2143]. Heuristic [1620]. heute [1597, 3111]. Hevelius [1563, 1799, 3077]. Heyck [2461, 104]. Heywood [886]. Heyworth [1790]. Hibner [1625]. Hicks [2539, 2702]. Hidden [207, 1547, 3162]. Hide [1072]. Hiebert [1941, 1175]. Hier [26]. Hieronymus [38]. Higgitt [2349]. High [77, 1938]. Higher [2873, 2542]. Highland [773]. Highlands [1568, 3006]. Highways [179]. Hilaire [2569, 2703, 2703]. Hilbert [1590, 1590, 3066, 3108]. Hilgartner [2669]. Hilger [2117]. Hill [2186, 1057, 1593, 481]. Hill-Brown [481]. Hillard [1123]. Hillel [1118]. Hilster [2848]. Himmel [2990, 1636]. Himmels [2517]. Hindenburg} [2129]. Hindsight [2762, 2912, 2403]. Hinrichs [1112]. hipnotismo [2279]. Hippocrates [1195]. Hippocratic [2098, 188]. Hirai [1805, 1045]. Hiro [1805, 1045]. Hirsch [1735]. Hispanic [1050]. Hispánica [1939]. Hiss [1888]. Histoire [1485, 2938, 1551, 753, 564, 1683, 2261, 3215, 2569, 369, 350]. Histoires [34, 777]. Historia [2110, 75, 1350, 3081, 859]. Historiality [468]. Historian [1848, 974, 1676, 2310, 164]. Historians [467, 2075, 1173, 1857, 2613]. Historic [2217]. histórica [427]. Historical [536, 899, 1173, 1619, 1535, 2390, 979, 1267, 1628, 2692, 2485, 548, 1681, 864, 1836, 600, 2515, 1279, 993, 1113, 1091, 842, 363, 2544, 830, 2416, 1680, 2625, 1503, 57, 50, 2765, 1149, 98, 2763, 3034, 3142, 3245, 3284, 3225, 2001, 1961]. Historically [2770, 768]. Historicism [2396]. Historicized [2761]. Historicizing [2862]. historiens [1683, 3215]. Histories [2859, 1529, 2599, 1545, 2931, 2357, 2721, 163, 341, 1204, 175, 9, 2961, 1759, 7, 2687, 1369, 258, 2016, 3281, 3112, 2693, 1154, 2249, 1460, 2695, 1004, 632]. Historiographical [1615, 2823, 6]. Historiographie [990]. Historiographies [912]. Historiographische [2422]. Historiography [1905, 700, 874, 1916, 340, 414, 2554, 2490, 2721]. historique [252]. Historisch [2651]. Historisch-Kritische [2651]. historischen [3034, 1329]. historischer [1472]. History [1308, 186, 1114, 2827, 1989, 215, 941, 911, 467, 1025, 1031, 1263, 1694, 2054, 1, 333, 616, 704, 1164, 1446, 1763, 2070, 2384, 1502, 665, 2802, 196, 668, 1036, 2870, 2960, 497, 2735, 1869, 1227, 1554, 2439, 1755, 1309, 1568, 161, 1176, 422, 2320, 1527, 980, 1006, 1511, 1900, 2776, 2689, 545, 1333, 254, 1492, 1402, 681, 2612, 2162, 1171, 1930, 2266, 2059, 1848, 974, 2813, 1969, 1282, 338, 1572, 815, 2373, 2826, 1530, 261, 1579, 74, 1119, 1430, 1102, 1850, 513, 300, 1580, 874, 1236, 701, 1524, 520, 1104, 2458, 280, 1146, 2117, 1931, 1505, 2194]. History [2362, 855, 1169, 1206, 972, 1852, 2774, 2269, 73, 508, 379, 2829, 2509, 367, 1966, 1547, 471, 2042, 308, 2769, 2832, 2936, 721, 2094, 825, 1844, 2411, 1678, 1563, 2031, 2672, 1682, 2609, 1663, 1591, 1696, 2652, 3, 1565, 2227, 2443, 2682, 2915, 1846, 945, 1773, 598, 376, 293, 1305, 1736, 1847, 44, 179, 1035, 354, 1829, 1147, 290, 1397, 1456, 27, 1518, 334, 1528, 2325, 2369, 968, 2893, 1451, 914, 2825, 814, 2584, 926, 1552, 2908, 223, 2389, 840, 2397, 2597, 1539, 1420, 1751, 1016, 2159, 1488, 1854, 472, 2410, 1893, 2398, 2949, 918, 1660, 141, 1355]. History [1746, 826, 683, 2861, 292, 2603, 599, 1034, 1455, 1567, 1082, 416, 915, 180, 256, 1330, 1279, 1856, 2221, 1816, 2277, 2854, 807, 1672, 2836, 2396, 1103, 919, 800, 971, 1918, 1747, 2002, 546, 2644, 181, 3285, 711, 762, 1116, 1851, 1907, 2734, 3286, 680, 1787, 2399, 860, 2284, 90, 111, 2492, 1439, 1387, 2416, 977, 2288, 2166, 455, 763, 451, 1432, 1748, 2256, 255, 69, 799, 2097, 2467, 1843, 1172, 2245, 33, 720, 2418, 2534, 85, 336, 1849, 2246, 1637, 2196, 456, 185, 820, 7, 2911, 925, 2613, 70, 2634, 64, 1564, 2493, 1404, 2366, 2089, 1595]. History [2140, 2816, 2940, 1772, 2561, 1437, 2714, 311, 2558, 1623, 1990, 3133, 3232, 3035, 3234, 3186, 2995, 3050, 3000, 3256, 3002, 3258, 3004, 3073, 3154, 3007, 3155, 2969, 3272, 3162, 3274, 3090, 3094, 3178, 2983, 3101, 3111, 3113, 3170, 3222, 3223, 3297, 3120, 3301, 3122, 169, 417, 1028, 1532, 1859, 2165, 2482, 2771, 24, 2528, 2875, 1995, 103, 1589, 1503, 2299, 2356, 1139, 2047, 2007, 1864, 1112, 1231, 1739, 1497, 1781, 1933, 1190, 1912, 1513, 2958, 2833, 2005, 2433, 1935, 1403, 1076, 1406, 32, 2516, 1433, 2831, 1220, 2359, 716, 570, 209, 1275, 1002, 541, 264]. History [437, 477, 542, 175, 353, 219, 758, 218, 643, 791]. Hitchhiker [2826]. Hitler [1595, 3101, 1895, 2451]. HIV [2065]. HIV/AIDS [2065]. Hjalmar [1886]. Hjermitslev [1784]. Hoarding [1170]. Hoax [327]. Hoaxes [1554, 3134]. Hobbes [2513]. Hochadel [2369]. Hochfelder [1217, 1064]. Hochschulen [739]. Hochschullehrer [61]. Hochschulreform [2668]. Hockett [1767]. Hockey [29]. Hoddeson [2144]. Hodson [2797]. Hoffman [1305, 1187]. Hoffmann [498, 1302, 1508]. Hogan [2599]. Hogarth [2710]. Hogg [2134]. Hoi-Eun [2524]. Holding [2286]. Hole [2049, 3173]. Holger [2444].

Holistic [2689]. Holland [2571, 1205]. Hollings [1747]. Holly [2530]. Hollywood [822]. Holmes [1507, 1834, 2948]. Holocaust [916, 677]. Holt [2739]. Holtorf [1303]. Holy [709, 3143, 2677]. Home [2643, 1247, 295]. Homei [2595]. homeland [3172]. Hommage [593]. Homme [1541, 3030, 1538, 988]. Homo [2590, 2654]. Homocentric [641]. homöopathischer [2292]. Homuncular [2533]. Hon [2787]. Honeybee [2208]. Hong [670, 2099]. Hong-Sen [2099]. Honnecourt [490]. honneur [1459, 2703]. Honor [593, 639, 1458]. Honoratus [2104]. Honoré [369, 934]. Hoock [2703]. Hoogendoorn [2932]. Hoogleraar [2126]. Hooke [1038]. Hopkins [2373, 1535, 44, 69, 64]. Hopper [143]. Hopwood [2037]. Horizon [1937]. Horizon-Based [1937]. Horizons [710]. Hormone [1158, 454, 682]. Hormonen [2126]. Hormones [1224, 672]. Hormonforschung [510]. Horning [1504]. Hornsby [1205]. horologiis [928]. horreur [1551, 3124]. Horse [2050]. Horsley [2474]. Horst [39, 1695, 2630, 928, 789, 2002, 1188]. Horstmanshoff [1679]. Horta [2368]. Horticultural [908]. Hoskin [1360, 1487]. Hösle [2851]. Hospital [2443, 2366]. Hoßfeld. [2522]. Hostile [2137]. Hostility [180]. Hot [1678, 3256]. hottentote [1493]. Hough [1307]. House [315, 1944]. Household [2572]. Houses [2650]. Housni [1549]. Howard [2790, 2001, 1456, 2046]. Howe [1607]. Howkins [2467]. Hoyningen [1244]. Hoyningen-Huene [1244]. Hsiang [2050]. Hsiang-Lin [2050]. HSS [1029, 1255, 1859]. Hubbard [2294]. Hubble [172, 2464, 1234]. Huber [2287]. Hubert [78]. Hubertus [2339]. Hubris [106]. Hudecek [2059]. Hudson [1566, 3135]. Huene [1244]. Huffine [43]. Huggins [659, 1952, 1298, 1952]. Hugh [1517, 2930, 2097]. Hughes [2425]. Hui [1218]. Huib [1375]. huitong [736]. Hülsenberg [1132]. humaine [2179, 350]. Human [2693, 2116, 2814, 1176, 130, 1139, 225, 469, 2056, 172, 685, 1012, 568, 2259, 547, 1166, 1628, 3070, 703, 1442, 2886, 1662, 96, 141, 2217, 1046, 2927, 1079, 962, 2202, 852, 1601, 1571, 1105, 2307, 591, 1787, 1622, 151, 1170, 1371, 2246, 2012, 693, 2543, 323, 98, 791, 3137, 3249, 2970, 3080, 3102, 3225, 3127, 1073, 2296]. Humani [1622, 3249]. Humanism [687, 2926, 2011, 1045]. Humanist [2030]. Humanistic [2630]. Humanists [2902]. Humanities [1027, 1529, 1526, 1527, 1160, 1855, 1682, 221, 2828, 3232, 3151, 1932, 1932, 2663]. Humanity [2373, 123, 2066, 600, 2219, 834]. Humans [1949, 1931, 2171, 2862]. Humboldt [2799, 2578, 398, 743, 1132, 2773, 1716, 1219, 2386, 2864]. Humboldtian [1891]. Humboldts [2027]. Hume [123, 2705]. Humfrey [781]. humido [788]. Humor [2392]. Humphry [2118]. hun [440]. Hunayn [1042, 2696]. Hundred [1615, 1975, 2049, 857]. Hungary [2957]. Hunger [2106]. Hunt [2915, 313, 578]. Hunted [637]. Hunter [403, 2936, 2267, 1411, 1811, 2461, 104]. Hunting [650]. Huntington [310]. Hurlbut [2543]. Husserl [579]. Hutchinson [686, 686, 1202]. Hutchison [1362]. Hutton [2858, 1951]. Huxley [1901, 698]. Huxtable [1341]. Huygens [1564, 3129]. Hybrid [1035, 2345, 862]. Hybridity [697]. Hyde [1950]. Hyder [579]. Hygiene [2373, 1536, 3193, 505]. Hynek [2098].

Hypothesis [1334, 2784, 1618, 2883, 429, 3104]. Hypothetical [109]. Hysteria [2018, 386]. Hyung [2206].

Iacopo [1197]. Iain [1207]. Ian [497, 2934, 1363, 2661, 2448]. Iannini [1488]. Iannucci [2176]. Ib [1648]. Iberia [2104]. Iberian [2412]. ibérico [1939]. Ibn [2257, 2564, 1340, 1098, 2696, 1042, 3053]. icône [944]. Ice [1267, 741]. ich [586]. Icon [146]. Iconografia [1417, 2176, 3118]. iconographicus [3247, 1695]. Icons [141]. Icy [1336, 3125]. Ida [397]. Idea [2338, 755, 1663, 1039, 1397, 2049, 1407, 1800, 1965, 3132, 3173, 3004, 3007, 1896, 1114, 1081, 2561]. Ideal [1534, 404]. Ideas [520, 2216, 2553, 1331, 3059, 3303, 750]. Ideen [1212]. Idées [1538]. identités [2585]. Identities [1583, 3025]. Identity [1025, 1653, 2104, 2557, 1904, 2074, 849, 151, 3138, 1909, 1083]. Ideology [2640, 1751, 2705, 3286, 72]. Idiom [2610]. Idiosyncrasy [396]. Ido [1938]. Idol [1813]. ihre [3150, 1574]. ihrer [3034, 1329]. II [2592, 2697, 825, 580, 2103, 1823, 252, 295]. III [866]. Iina [2893]. IJdelheid [2126]. Ikonografie [401]. Ilana [2672]. Ileana [1357]. Iliffe [2638]. Ilkley [382]. Illicit [1124]. Illness [1307, 815, 638]. illus [1598, 1585, 1588, 1541, 1590, 1569, 1554, 1562, 1568, 1594, 1578, 1577, 1543, 1572, 1600, 1579, 1609, 1566, 1580, 1537, 1547, 1548, 1583, 1563, 1556, 1596, 1591, 1565, 1605, 1534, 1576, 1581, 1602, 1606, 1558, 1677, 1597, 1545, 1599, 1610, 1546, 1573, 1561, 1557, 1607, 1601, 1571, 1536, 1592, 1542, 1549, 1587, 369, 1604, 1676, 1589, 1540, 1570, 1564, 1560, 1608, 1574, 1553, 1595, 1582, 1544]. Illusion [1534, 2657]. Illusory [716]. Illustrate [1543, 3149]. Illustrated [186, 1112, 2832, 1589]. Illustrating [1464]. Illustration [2533, 1168]. Illustrations [1020, 2413, 2776]. illustrées [274]. illustriertes [3247, 1695]. Illustrious [2864]. Illy [1507, 1063]. ilustrada [351]. im [2799, 296, 2454, 1562, 990, 2925, 2143, 2977, 120, 1373, 739, 1597, 2027, 2504, 2343, 2190, 658, 2449, 370, 2657, 272, 740, 1310, 3111, 2292, 2944, 1955, 1972, 3130, 1134]. Image [1268, 2936, 1290, 2978, 1168, 1046, 2610, 2002, 3190, 2916, 393]. image-thinking [3190]. Imagery [2002]. Images [1436, 2523, 1909, 1169, 1898, 293, 917, 1234, 1978, 2037, 57, 1582, 2998, 3213, 2978, 563, 984]. Imaginary [1053]. Imagination [1818, 395, 1053, 782, 21, 799, 201, 37, 2576, 2916, 675, 2803, 712, 393, 917, 985]. Imagined [2309, 1021, 3064, 2429, 2310, 2313, 2308, 1642]. Imaging [1108, 3194]. imaginibus [38]. Imagining [1077, 1299, 1808, 1077, 638]. Imago [1562]. Imitation [482]. immagini [2845]. Immanuel [1010]. Immediate [536]. Immigration [1717, 3270]. immobile [1061]. Immortality [1828]. Immortalized [2792]. Immunity [185]. Immunization [2594, 2598]. Immunology [1326, 185]. Impact [1960, 1854, 2216, 2423, 1637, 3154, 3303, 585]. Imperfect [2672, 324]. Imperial [2783, 52, 2112, 2524, 880, 1722, 276, 1363, 1729, 3088, 2277, 3281, 465, 775, 2782, 20, 1963, 3221, 1818, 2237, 3288, 1759]. Imperialism [381]. Impersonality [1898]. Impetus [934]. Implementing [1617]. Implications [135]. Imponderable [2039]. Imported [1834]. Impossibilities [93]. Impossible [198]. Impressions [536, 2342]. Imprinting [1313]. Improbable [2415, 808]. Improvement [3231, 1708, 257]. Impure [2553]. In-between [2937]. In-House [315]. In/visible [966]. Ina [2143, 1710]. inattuale [954]. Inca [2385]. Including [2060, 1564, 3129]. Inclusive [1391]. Incognito [2289]. Incommensurability [405]. Incorporated [1093]. Incorrigible [1976]. Indecent [2193]. Independence [384]. Indeterminable [1525]. Indétermination [478]. Index [330, 613, 904, 1162, 1444, 1761, 2067, 2376, 2673, 2962, 1888, 1598, 1585, 1588, 1590, 1569, 1554, 1568, 1594, 1586, 1578, 1577, 1543, 1575, 1572, 1600, 1579, 1609, 1566, 1535, 1580, 1537, 1547, 1548, 1583, 1556, 1596, 1591, 1565, 1605, 1534, 1576, 1581, 1602, 1606, 1558, 1677, 1545, 1552, 1539, 1555, 1599, 1610, 1546, 1573, 1567, 1561, 1559, 1607, 1601, 1551, 1571, 1314, 1536, 1592, 1542, 1549, 1584, 1587, 1593, 1604, 1676, 1589, 1540, 1570, 1538, 1560, 1608, 1574, 1553, 1595, 1603, 1582, 1544, 709]. indexes [1590, 1597, 1557, 369]. India [1890, 2994, 1356, 357, 1294, 1918, 2345, 2933, 973, 974, 8, 3096, 3281, 3302, 1222, 1438, 30, 2357, 1491, 1759, 1686, 749]. Indian [2249, 1425, 1488, 1704, 2047, 3189, 972, 3160, 2679]. Indien [2337]. Indies [1065, 2950]. Indigenous [1219, 1491]. Indigo [1222]. Individual [585, 2122]. Individuality [2692]. Indonesia [2950, 2741, 663]. Industrial [862, 1734, 1451, 557, 82, 1592, 2976, 3244, 3298]. Industrialist [1211]. industrielle [745, 122]. Industrious [2688]. Industry [2213, 123, 2552, 1039, 390, 672, 983, 514, 1740, 3293, 2119, 2634, 2140, 737]. inédit [571]. inédites [1225]. Inequality [2539]. Infanticide [319]. infanzia [2324]. Infections [1968]. Infertility [2410]. Infinite [861, 781]. Infinitesimal [2432]. Infinity [62, 742]. Influences [1, 333, 616, 1164, 1446, 1763, 2070, 2384, 3086, 1345]. Information [2466, 71, 1818, 561, 1081, 2223, 143, 630, 1330, 3282, 2677, 2044, 2233, 1731, 1888, 1970]. Infrastructure [2928]. infusoires [2350]. ingénierie [689]. Ingenuity [2931]. Ingo [1132]. Ingrid [2505]. Inhabited [2178]. Initial [304]. Initiative [895]. Injektion [2449]. Ink [1626, 3158]. Inkpot [280]. Innes [1059, 2028]. Innocence [560, 2954]. Innocent [2293]. Innovation [1598, 2573, 298, 77, 315, 2518, 2464, 1451, 234, 216, 2374, 2561, 442, 3208, 487]. Innovations [2843]. Innovationsmanagement [1310]. Inquiry [1380, 927, 3079, 1783, 2705, 326]. Inquisition [1351, 876, 1555, 3161, 2898, 3085, 2713]. Insanity [1930]. Insatiable [234]. Inscription [909]. Inscriptions [1004, 2255, 2972]. Insect [1290, 969, 3139]. insecte [3139]. Insel [723]. Insider [2816]. Insight [1543, 3149]. Insights [1318]. Insolubilia [778]. Inspiration [464]. Inspired [2656]. Instinct [1957, 411]. Institut [1087, 1188]. Institute [824, 1563, 1556, 1838, 857, 60, 514, 1100, 939]. Institute/Turin [1556]. Instituting [956]. Institution [1600, 74]. Institutional [827]. Institutions [231, 290, 144, 2295]. Instituts [46]. Instrument [1707, 804, 3220]. Instrumental [2108, 3131, 1237]. instrumentorum [3247, 1695].

Instruments [536, 494, 534, 2172, 2848, 1634, 855, 1462, 535, 533, 1476, 2796, 3228]. Insurance [285]. Insurrection [2954]. Inszenierung [2343]. Integrating [2692]. Integration [2814, 736]. Integrative [1171]. Intellectual [497, 1341, 2119, 1970, 1980, 664, 2344, 116, 1033, 1974, 1731, 3282, 93, 1787, 2399, 426, 495, 2856, 2940, 2775, 3206, 653]. Intellectuals [1425, 80, 3160]. intellectuele [2785]. intelletto [2343]. Intelligence [561, 520, 1950, 1411, 283, 1666, 3068, 2328, 611]. Intelligent [1020]. Intelligibility [22]. Intentional [2322]. Interaction [1672, 1688, 3295]. Interactional [829]. Interdisciplinarity [1977]. Interdisciplinary [1563, 2169]. interdisziplinären [1233, 1972]. interdisziplinäres [2699]. Interessenverband [2668]. Interest [2109, 466]. Interests [2617, 2213, 1448, 827, 2364]. Interfacing [2773]. Interiority [1883]. Intermediary [1167]. International [1964, 749, 1155, 204, 1267, 2074, 4, 607, 1731, 2153, 3282, 2307, 1089]. internationalen [441, 1972]. internationalisation [2295]. Internationalism [948]. Internet [1610, 3044, 78, 1308, 2597, 77]. interpretacão [566]. Interpretation [874, 56, 639, 1517, 1383, 2292]. Interpreting [1054, 2215]. Intersection [998, 857, 691]. Intersections [1591, 1702, 3243, 3198]. Interstitialität [1332]. Intersubjective [2851]. Intervention [2084]. Interventions [2214, 2493, 3304]. Interwar [1077, 2452, 1224]. Intimate [1349, 3001]. intorno [1641, 3203]. Intrigue [1888]. Introduction [2989, 1526, 764, 1845, 1920, 2149, 2155, 2471, 2606, 2681, 2755, 1616, 1670, 841, 2475, 2177, 2824, 2235, 104, 779, 2967, 489, 910, 2159, 1101, 1321, 412, 970, 337, 2551, 1250, 1386, 5, 533, 2565, 620, 1023, 1768, 3142, 3060, 1535, 1542, 2089, 1339, 2325, 2258, 835, 541]. Intuition [172]. Inuit [1730, 2371, 3277]. Inutility [2319]. Invasions [384]. Invented [2254]. inventeurs [442]. Inventing [2976, 3244, 1734, 2128, 1056, 1885, 1068]. Invention [77, 1931, 2259, 2161, 1534, 1335, 143, 1300, 2760, 1979, 686, 793, 2018, 1863, 3119, 1039]. inventioni [2427]. Inventions [2108, 2900, 1063]. Inventor [3037, 1583, 1534, 3025]. Inventors [2305]. Investigate [1442, 3080]. Investigates [1115]. Investigating [812]. Investigation [2336, 425]. Invisibility [2264]. Invisible [126, 1626, 1525, 3158, 2009]. Iranian [303]. Iraq [32]. IRBs [897]. Ireland [376, 3276, 1715]. Irene [789]. Irina [2343]. Iris [1507, 586]. Irish [1501]. Irmas [2449]. Iron [2957, 1747, 3258]. irraffigurabile [1130]. Irrational [159]. Isaac [1954, 1287, 1261, 1051, 1633, 2638, 1663, 1212, 1379, 3023, 3140, 2760, 1676, 2853, 961, 1293]. Isabel [2643]. Isabelle [2392, 1785]. Isador [531]. Isenberg [1933]. Ishaq [2696, 1042]. Isis [1317, 1445, 1444, 1762, 1761, 1024, 2067, 905, 330, 613, 904, 1163, 1162, 2068, 2377, 2376, 2674, 2673, 2963, 2962]. Islam [719, 1418, 2877, 3107, 3024, 1549, 495, 2701, 477]. Islamic [563, 2565]. Islamicate [2914]. Island [887, 1668]. Islands [2955, 1103]. Isn’t [1249]. Isolation [2943, 2146, 1385, 255]. Israel [2667, 1064]. Israelite [1937]. Israelowich [1938]. Issue [2684]. Issues [521]. István [2060]. Italian [2947, 2927, 1278, 1345, 2985, 3039, 3203, 3204, 3041, 2807, 3248, 2383, 3086, 3115, 3172]. italiana [1578, 3039, 1357, 3172]. italiane [1184, 890]. italiano [211]. Italie [2578, 1632, 1894]. Italien [39]. Italo [595]. Italy [2926, 1983, 1946, 1555, 1248, 3161, 3199, 3274, 1694, 2631, 2108, 1435, 2646, 1884, 2566, 1282, 553, 644]. Itinerari [3204, 211, 1641]. Itineraries [9]. Itineris [2937]. Ito [1719]. Itself [1601, 2968, 3102]. itsu [2881]. Iurii [886]. iuvenalia [928]. IV [2502, 1963]. Ivan [1834]. IVe [178]. IVe-XVIe [178]. Ivey [1587]. Ivor [1675, 179]. Iwan [2966, 2832]. ix [1568, 1586, 1556, 1591, 1571, 1676, 1608, 1603]. Ixtlilxochitl [2514].

J [1824, 1031, 1393, 1502, 1242, 1036, 2172, 2643, 1869, 1436, 2439, 2712, 2622, 210, 2599, 1347, 2650, 1418, 1600, 2946, 1819, 677, 1112, 2532, 1977, 1429, 2509, 2180, 1738, 2272, 1299, 1904, 133, 846, 1150, 948, 1049, 1781, 2844, 200, 2808, 1305, 2175, 1912, 1709, 2464, 2958, 1545, 2450, 2948, 2931, 1061, 2661, 2350, 2294, 2515, 2425, 1500, 2717, 2202, 1298, 1205, 1144, 1741, 1066, 2125, 1702, 1219, 2589, 2733, 2046, 2028, 578, 1089, 1748, 397, 2196, 1375, 2732, 2542, 1560, 2205, 1595, 2796, 2543, 1409, 478, 675, 517]. J. [2852, 2415, 1724, 2649, 19]. J.-C [478]. Jaap [1932]. Jablonka [1110]. Jaboulet [1792]. Jaboulet-Vercherre [1792]. Jacalyn [1608, 25]. jacht [2451]. Jackson [946, 1716, 2065, 63]. Jacob [457, 1571, 1661, 2662, 42]. Jacobs [1984]. Jacobson [2031]. Jacquart [2179, 1632]. Jacqueline [2808]. Jacques [441, 2166, 397, 2499, 1805, 1195]. Jacquin [1350, 3081]. Jacyna [1220]. Jaghmini [2565]. Jahnavi [1438]. Jahre [1188]. Jahren [1597, 2668, 3111]. Jahrestagung [1892]. Jahrhundert [593, 2699, 3130, 1373, 2944, 1134]. Jahrhunderts [990, 2190, 3150, 3003, 401, 272, 1574, 1646, 2200, 2504, 1814]. Jail [260]. Jakob [140]. Jamaica [1387]. James [2879, 296, 1341, 2112, 210, 1954, 2837, 1370, 2532, 1720, 2325, 2217, 1573, 54, 3019, 1059, 578, 939, 2520, 1379, 1655, 3023, 2128, 2030, 1718, 2199, 2356, 1954, 2142, 2463, 1844, 1397, 3078, 1681, 2450, 2715, 2043, 1879, 160, 1369, 2434, 1688, 79, 279]. Jamie [1662, 2729, 1705]. Jamil [2503, 1793]. Jan [2340, 2126, 2865, 2008, 2940, 1192, 2699, 1136, 2118, 2587]. Jan-Hendryk [2699]. Jana [2504]. Jane [2837, 2656, 1629]. Janeiro [2867]. Janelle [1226]. Janet [1685, 1059, 1148, 2808, 1978, 1787, 2135, 2658]. Janiak [1054, 1809]. Jansen [2212]. Janssen [1623]. January [1507]. Japan [2881, 1297, 2524, 2737, 1947]. Japanese [2439, 748, 2096, 2653, 2479]. Jar [2762]. Jardine [1698]. Jared [1363]. Jarosław [1563]. Jarrell [1453]. Jas [1664]. Jason [2137, 2278, 2724]. Jaszi [1033]. Jaume [1502]. Jay [687, 150]. Jazz [57]. Jean [1344, 561, 1647, 1223, 114, 2227, 274, 478, 1499, 2419, 1470, 2875, 2697, 1911, 1540, 1870]. Jean-Baptiste [561, 2227]. Jean-Claude [1499]. Jean-Marc [2875, 1540]. Jean-Michel [2419]. Jean-Patrice [1470, 2697]. Jean-Paul [1911]. Jeanne [1185, 2443]. Jed [808, 1455, 1676]. Jeff [2918]. Jefferson [308, 251]. Jeffery [1330]. Jeffrey [1517, 1937, 2192, 1437, 1887]. Jehanne [2581]. Jehanne-Emmanuelle [2581]. Jelved [946]. Jena [2190, 53]. Jenell [1153]. Jenna [2346]. Jennifer [2749, 2294, 2718, 2359, 1415]. Jens [1190]. Jensen [1460, 2808, 1065, 2410]. Jensen-Vallin [2808]. Jenyns [1423, 3005]. Jérôme [38]. Jeremiad [2911]. Jeremy [1311, 951, 2194, 2057, 2583, 1216]. Jeroen [2663, 2222]. Jerome [2775, 1548]. Jerusalem [2807]. Jess [2667]. Jesse [12]. Jessica [2094]. Jesuit [1271, 1945, 2154]. Jesuits [276, 731]. Jesus [1399, 2661, 1836, 2981]. Jet [2145]. Jetze [2708]. Jewell [2871]. Jewish [2104, 2807, 1735, 3209, 998, 1946, 1789, 1693, 1546, 921, 3294]. Jews [1738, 3242, 839]. Ji [2720, 230]. Jian [2099]. Jian-Liang [2099]. jiaoyi [736]. Jigu [2628]. Jim [1408]. Jimena [2291]. Jin [2325]. Jing [1733]. Jiri [2059]. jishu [223]. Joachim [443, 2339, 2178, 1892]. Joan [2020, 287, 800, 1979, 706, 2241, 1775]. Joanna [1864, 88]. João [447]. Jochen [401, 2703]. Jockeys [1832]. Joël [1797, 1663, 2621, 1637, 34]. Joella [1564]. Joelle [1238]. Joeri [2876]. Johan [1806, 1891, 1806]. Johann [990, 2632, 1862, 873, 2348, 2703, 2449]. Johannes [2445, 2111, 2940, 1695, 1563, 3077, 928, 1721, 2187, 1799, 2716, 2535, 373]. John [2512, 2510, 754, 389, 2945, 512, 2731, 639, 1790, 2866, 1055, 1908, 1049, 1701, 189, 2967, 1690, 47, 1057, 1785, 1567, 929, 1458, 2028, 738, 2921, 2089, 375, 2775, 1394, 1454, 2297, 1726, 1341, 1798, 1659, 2299, 2103, 2405, 2912, 3235, 1416, 2841, 838, 2798, 2063, 1624, 2117, 2344, 1077, 1241, 2822, 2863, 2637, 2915, 2696, 1056, 3010, 2833, 2005, 2908, 2334, 1753, 626, 474, 2982, 2909, 1116, 245, 1570, 2911, 2247, 1603, 1121, 2775, 754, 502, 677]. Johns [2373, 1535, 44, 69, 64]. Johnson [1153, 1681, 1142, 1965, 2406]. Johnston [1083]. Jolley [1645]. Jonas [53, 1984]. Jonathan [2192, 1073, 2134, 1401, 1821, 2559, 2125, 1680, 37]. Jones [2573, 1424, 2735, 1647, 2494, 2026, 1825, 2626]. Jonsson [1568]. Jordan [1142, 2680]. Jordanus [993]. Jörg [593, 1435, 579, 2738, 2147, 2889]. Jørgen [1648]. Jorgensen [2205]. Jorink [1796]. Joris [2201]. José [2110, 2931, 1122, 2268]. Josef [2124, 2287]. Josefowicz [808]. Joseph [1341, 298, 1543, 3189, 1647, 31, 2833, 3081, 1704, 1350, 2954, 1201, 2988, 2119, 1080, 1547, 2682, 1800]. Josephson [1333, 1602, 1891]. Josh [1582]. Joshua [1600, 1607]. Jost [947]. Jouanna [1195]. Journal [2680, 2997, 2822, 2574, 1419, 1969]. Journalism [1699, 3283]. Journals [679]. Journey [3207, 1648, 531, 145, 2694]. jours [350]. joy [3187]. Joyce [2420]. Joydeep [1890]. József [1507, 1063]. Jr [182, 1913, 2440, 375]. Juan [2433, 2722]. Juanita [1790]. Judgment [2079, 222]. Jüdische [61]. Judith [2947]. Judy [1605, 2409]. Jugendwerke [928]. Julia [593]. Julian [1757, 2799, 1318, 22, 3216]. Juliet [2952]. Juliette [2013, 2215]. Julius [2435, 2781]. julliet [274]. Jump [2891]. Jump-Started [2891]. June [12]. Jung [2210]. Jungle [2839, 602]. Juno [2515]. Jurassic [567]. Jurek [1903]. Jürgen [1511, 736, 2629, 1907, 2517, 2734, 1310, 2627]. Jürgensen [618]. Jurisprudence [271]. Justice [2375, 3187]. Justifier [1541]. Justin [1557, 1137, 2012]. Jutta [2799, 2558]. Jütte [1861].

K. [1017, 2932]. Kaasch [1892]. Kaat [2519, 2201, 2730]. Kagan [1650]. Kahn [2533, 2184]. Kaiser [636, 1087, 2374, 1188, 104]. Kaiser-Wilhelm- [1188]. Kaiser-Wilhelm-/Max-Planck-Institut [1188]. Kaiser-Wilhelm-Institut [1087]. Kaiserfeld [1434]. Kaiserlichen [2810]. Kaji [1973]. Kaldewey [2953]. Kaliapparat [1449]. Kaliardos [1640]. Kamerlingh [73]. Kammerer [2453]. Kant [856, 1188, 1866, 1710, 3251, 3087, 1415, 1283, 802, 791]. Kara [1837]. Kardel [1644]. Karen [2373, 2801, 2133, 1403, 1638, 2031, 946, 1829, 2225, 807, 2106]. Kargon [2133]. Karin [2107, 2872, 167]. Karine [2321, 2166, 1404]. Karl [2208, 2197, 1142, 2797, 2656, 2485, 2655, 1702, 405]. Karlsohn [1891]. Karola [1542]. Karsten [2925]. Karten [3150, 1574]. Kasseler [2505]. Kassler [1705]. Katharina [1640]. Katharine [1419]. Katherine [1734, 2011, 2714]. Kathleen [2066, 2879, 1957, 2025]. katholieken [3246, 1732]. Kathryn [1740]. Katja [2370]. Katy [1364]. Katz [2425]. Katzir [1511]. Kaushik [1159]. Kaye [1637]. Keating [1084]. Keene [2040]. Keeve [2751]. Keighren [2028]. Keith [1864, 308, 2127, 2346, 2955]. Kekulé [393]. Keller [2146, 2109]. Kelley [2491]. Kelley. [1143]. Kellie [2498]. Kelly [707, 1330, 2020]. Kempe [272]. Kempf [2799]. Ken [2300, 1559]. Kendall [724]. Kendrick [1157]. Kenelm [787]. Kennaway [1397]. Kennedy [754, 2215]. Kenneth [2186, 1536, 1823, 1211]. Kenya [2047]. Kepler [2789, 1373, 2264, 2111, 2187, 373, 2716, 3130, 992, 3032, 1347, 2901, 3130]. Kepler-Rezeption [1373, 3130]. Kerne [1972]. Kernenergiepläne [1310]. Kernforschung [1972, 1233, 1972]. Kerr [2943]. Kerrin [1329]. Kerssenbrock [1942]. Kerssenbrock-Krosigk [1942]. Kersten [1985]. Kessler [1234]. Kevin [43]. Kew [2869]. Key [2301, 873]. Keyser [2921, 2563]. Kidd [1336]. Kidwell [64]. Kiechle [2802]. Kieran [2458]. Killen [2654]. Killer [411]. Kim [2524, 1624, 1539, 2429, 30, 2865]. Kimball [2411]. Kimberly [2095, 1728]. Kinds [1988, 829]. Kinematics [183]. King [2602, 1627, 187, 1679, 1839, 1116]. Kingdom [1200, 935]. Kingsbury [1035]. Kinsey [1824]. Kinship [163]. Kirchberger [1830]. Kirchen [2226]. Kircher [935, 1292]. Kirchhoff [2446]. Kirkham [1692]. Kirschner [1344]. Kisacky [2443]. Kiss [2671]. Kita [2881]. Kitab [1042, 2257]. Kitchen [149]. Kjærgaard [1784]. Klaas [953, 2777, 2777, 2021]. Klaassen [1124]. Klassieke [2777]. Klaus [2446, 2453, 1332, 1040, 2124, 28]. Kléber [1209]. Kleeberg [1361]. Klein [1225, 2585, 1783, 1605, 12, 2027, 2347]. Kleineberg [723]. Kleist [2993, 1129]. Klemm [1691]. Klep [397]. Kline [2223]. Klinger [1329]. Knapp [2520]. Knegtmans [2126]. Knell [1066]. Knew [2531]. Knobloch [723]. Knoeff [1788]. Knossos [266]. Knot [2935]. Know [630, 927, 402]. Knowing [1545, 667, 1093, 3290, 936, 548, 691]. Knowledge [2512, 2158, 1025, 2297, 2339, 792, 966, 690, 161, 2514, 2669, 2266, 1160, 2650, 774, 1674, 2412, 123, 1012, 45, 2977, 1640, 1269, 1977, 2340, 1968, 2071, 2630, 2629, 2645, 839, 188, 2325, 2825, 2786, 884, 2928, 2152, 1978, 1168, 97, 962, 2507, 2081, 734, 1672, 1103, 2345, 1571, 1410, 9, 1473, 2374, 2418, 804, 1947, 2434, 2192, 1024, 747, 98, 1409, 2996, 3056, 3070, 3097, 3290, 1496, 1545, 2109, 1824, 1813, 2575, 2918, 2667, 2222, 2568, 2253, 803, 500, 317, 692, 496, 277, 609, 54, 830, 722]. Knowledgeable [303]. Knowledges [581, 158]. Kobiljski [2569]. Koblitz [1625]. Koby [1822]. Koch [1609, 387]. Kochan [2918]. Kochs [503]. Koenig [583]. Kohonen [2893]. Koischwitz [2668]. Kolb [2144]. Kölbl [1960]. Kölbl-Ebert [1960]. Kolchinskii [1981]. końca [1865]. Konflikt [2504]. Konflikte [2226]. Kong [670]. König [39, 2035]. Koninklijke [309, 953]. Konkurrenz [1972]. Konrad [39, 1230, 954]. Kontext [1972, 740]. Kontexte [3151]. Konzepte [1475]. Kooperation [1972]. Kooten [2169]. Kopp [393, 1500, 2608]. Kopperman [1567]. Korea [2557]. Kormos [1507, 2049]. Körper [3210, 1691]. Korrespondenznetzen [272]. Kosmann [1713]. Kosmann-Schwarzbach [1713]. Kosmoskop [3150, 1574]. Kostas [2195]. Kostelnick [2411]. Kouteynikoff [2502]. Kox [2946]. Koyré [2289, 1803]. Kozodoy [2104]. Kraft [2799]. Kragh [2490, 2744, 1975, 2052, 2132, 2435, 1973, 50]. Krämer [1807]. Kranz [39, 1695, 928]. Kraybill [2743]. Kreft [1795]. Krementsov [1067, 1828]. Kremer [1634, 1563]. Krieg [2591, 940, 1145]. Krige [2297, 1908]. Kriminalfalls [1301]. Krimsky [1311]. Kris [1425]. Krischel [2053]. Krishna [2589]. Krista [1717]. Krister [2520]. Kristian [1036, 56]. Kristie [1626]. Kristin [43, 1142]. Kristine [2537]. Kristy [1348]. Kritik [2292]. Kritische [2651]. Krosigk [1942]. Kruse [2373]. Kubaczek [2535]. Kuhn [1092, 1190, 2721, 405, 2722, 1017, 2485, 2175, 2624, 962]. Kuklick [1392]. Kulturbote [3049, 1359]. Kulturen [1498, 2799]. Kulturgeschichte [678]. Kulturwissenschaft [23]. Kumar [1222, 1189]. Kunst [1814]. Kunstkammer [1640, 3097]. Kunstkammern [46]. Kupfer [2698]. Kurihara [1559]. Kursell [593]. Kurt [710]. Kurven [511]. Kusukawa [1046]. Kutzinski [1219]. Kwaschik [2295]. Kyle [2926]. Kyoto [2881].

L [78, 2879, 2573, 1424, 1031, 2859, 1069, 1634, 1460, 210, 2943, 2592, 123, 2457, 1112, 1427, 1563, 2031, 2095, 1605, 2808, 1784, 1751, 2729, 1355, 2294, 951, 1741, 1937, 2718, 2745, 2359, 1358, 2326, 2872, 2465, 2089, 2736, 2940, 886]. L. [2087, 1220]. Lab [316, 907, 2725, 2816, 822]. L’abbé [942]. Labor [464, 1835, 1955]. Laborassistenten [28]. Laboratoria [2442]. Laboratories [602, 1942]. Laboratorio [508]. Laboratory [1318, 1356, 4, 2145, 228, 1614, 244, 1099, 2327, 909, 160, 255, 747, 1963, 2994, 88, 1869, 2816, 432, 387]. Labors [560]. Labyrinth [2731, 292]. Lachapelle [2269]. Lacour [1882]. Láctea [1873]. Láctea/De [1873]. lacteo [1873]. Ladies [1963]. Ladoff [531]. Lady [2045, 656]. LaFollette [1231]. Lagoon [2254]. Lalli [2540]. Lam [1304]. Lamarck [1490]. Lamarckism [1110]. Lamb [2559, 2121]. Lambe [2359]. Lambert [1573, 1862]. Lambright [1756]. Lammer [2843]. Lamont [1113]. Lamoreaux [2696]. Land [907, 276, 2806, 2968]. Land-Grant [2806]. Landgrave [2795]. Landry [1221]. Lands [2333, 2371]. Landscape [546]. Landscapes [2596, 2667, 2195]. Landschaft [2517]. Laney [2593]. Lang [1561]. langage [2441]. Langlitz [1152]. Language [129, 2548, 2684, 2557, 2579, 618, 2238, 2749, 2953, 621, 2186, 3013, 2096, 47, 2236, 470, 2237, 2239, 2233, 1413, 2833, 2208]. Languages [2240, 2235]. Lanham [1314]. l’Antiquité [350]. Lantz [1744]. Laplace [1286, 573]. Laqueur [2126]. Lara [2301]. Largent [1085]. Larin [1333]. Larink [1398]. Larry [2171]. Larson [2007]. Larvor [2486]. Lasser [2642]. Last [2531, 2079]. Late [1649, 2104, 1309, 2926, 588, 726, 276, 721, 1200, 189, 1349, 2780, 2928, 1792, 2877, 1279, 1550, 2633, 2237, 775, 119, 2566, 644, 1122, 1469, 3026, 2975, 3001, 1383]. Late-Victorian [588]. lateinischen [2183]. Later [2501, 1124]. Latin [920, 118, 2180, 1834, 1935, 1693, 1752, 1183, 1102, 1104, 2238, 2584, 2476, 1922, 1103, 1105, 1849, 3300]. Latino [1197]. Latinus [866]. Latium [935]. Latour [1980, 1380]. Laubichler [2656]. laudibus [2500]. Laue [947]. Laukötter [2882]. Laura [1716, 2662, 2139, 1610, 1750, 2690, 2857]. Laureate [601]. Laurence [2562, 988, 40]. Lauri [349]. Lauterbur [1979]. Lavoisier [1650, 1490, 2991, 2184]. Law [880, 1370, 271, 2515, 1667, 2469, 2970, 3123]. Lawrence [2045, 2816, 810, 2413, 1400, 2328, 1709, 473, 1808, 1961, 2413]. Laws [1351, 271, 1558, 3031, 3085]. Lawson [1223]. Lay [2764]. Lazier [1192]. Leah [1512]. Learned [2032, 1127, 2300, 1124, 2250, 414, 2614]. Learning [1483, 2503, 179, 487, 97, 2914, 2364, 2775, 855]. Least [2460]. Leben [140, 2453]. Lebenskraft [2117]. Lebenskraft-Debate [2117]. Leblanc [2579]. Leck [2197]. Leclerc [1485, 369]. Lecon [2502]. lecteurs [1894]. lecturam [2103]. Lecture [1735, 2506, 3209]. Lectures [1590, 1866, 1827, 3066, 3218, 2971, 2734]. Lee [308, 2538, 2869, 2891]. Leemans [866]. Leemans.II [866]. Leemans.II-III [866]. Leeuwen [2420]. Lefèvre [2844]. Left [1659, 2907, 2992]. Legacies [1819, 1891, 922]. Legacy [305, 1171, 2952, 2052, 454, 1941, 405, 2624, 1794, 93, 2130, 682, 3177, 935, 552]. Legal [1033]. Legami [1490]. Legend [187, 76]. Leggett [1068]. Legitimization [2487]. Legitimizing [3074, 2212, 1434]. Lehleiter [2644, 2251]. Lehmann [1405]. Lehner [1511, 1623]. Lehrbücher [2183]. Lei [2050]. Leibniz [2795, 1881, 657, 3212, 1087, 2854, 124, 1657, 2263, 431, 794]. Leiden [1483, 1788, 1552, 1539, 1546, 1549, 1564]. Leiden/Boston [1546, 1564]. Leistung [1087]. Leitão [2630]. Leitenberg [1190]. Lektüre [1807]. Lelgemann [723]. Lemelson [77]. Lemmerich [947]. Lemov [1605, 2219]. Length [1897]. Lenin [601]. Lennart [2425]. L’enquête [122]. Leo [514, 1560, 2320]. Léon [999]. Leonard [1401, 1423, 1423]. Leonardi [1374]. Leonardo [172, 1795]. Leonelli [2670]. Leonhard [2114, 2114]. Leonid [1044]. Leprosy [1690, 3275]. Lequeux [2030]. Lerch [2183]. L’eredità [954]. Leroi [2254]. Leroux [386]. Lesley [2512]. Leslie [2447, 2634]. Lessing [1589]. Lessons [1993, 1374, 1816, 164, 3012]. Let [2476, 1856]. Letter [1258, 420, 343, 345, 42, 247, 1256, 975, 344, 1395, 248, 421, 418, 2820, 419, 342, 347, 848, 1797, 1106]. lettera [727, 48]. Letterbooks [3235, 1701]. Lettere [2353]. Letteren [2785]. Letters [2879, 210, 770, 771, 349, 1548, 1556, 15, 946, 16, 2380, 2382, 1030, 2229, 628, 627, 2718, 2230, 2830, 2304, 1666, 14, 17, 3039, 3163, 3224, 1878, 2677, 512, 2610, 55]. Lettevall [1591]. Lettow [1714]. Lettres [1551]. letzte [2290]. L’Europe [2569]. Leven [2777]. Levi [2947, 89]. Levi-Civita [2947]. Leviathan [2084, 2076]. Levine [1495]. Levitin [2016]. Lëvshin [2211]. Lewis [1568, 2437, 2496, 1318]. Lexell [1806]. Lexington [1314]. Leyden [2762]. L’histoire [779]. Li [1881, 1657, 2603]. Liang [2099]. Lianne [2935]. Liber [648, 3247, 1695, 2500, 114]. Liberal [1012, 1957]. Liberalism [619]. Libertà [3248, 1700]. Libertas [727]. libertinage [649]. Liborio [1061]. Library [186, 24, 1552, 2506, 1604, 1564, 2773, 1047]. Libri [2103, 1622, 2845, 2570, 2502, 3249]. Libro [493, 2258]. libros [2103]. Lichtenberg [1412]. Lichtenbergischen [2024]. Liddel [2107]. Lidgard [2692]. Lie [941, 1115, 1594, 3141]. lieber [947]. Lieberman [2718]. Liebig [1449]. liegt [811]. Lies [147]. Life [1295, 285, 2440, 2510, 1441, 2669, 207, 1954, 656, 415, 1080, 403, 1005, 1819, 2286, 152, 2061, 2457, 177, 512, 664, 2531, 1714, 2607, 601, 2501, 1478, 400, 1229, 2549, 1581, 2435, 1305, 2372, 2060, 1950, 2450, 2715, 76, 1016, 1506, 1399, 1660, 141, 228, 2917, 447, 2583, 1832, 1142, 1806, 2270, 2899, 232, 3098, 800, 1557, 2474, 2062, 672, 358, 1811, 579, 762, 2589, 860, 2718, 1045, 939, 1213, 958, 3113, 2135, 514, 187, 1494, 959, 3133, 3174, 3042, 3266, 3159, 3278, 2981, 3167, 2599, 2346, 1829, 1518, 1985, 1439, 1603, 2665, 2960, 2191, 1984, 1629]. Life [1484, 1902, 316, 362, 657, 115, 459, 828]. Life-World [579]. Lifestyle [1247]. Lifschitz [1413]. Light [2446, 2753, 1599, 1364, 578, 2754, 1936, 3014, 3169, 2754, 1141]. Lighthouse [2877]. Lightman [1653, 1577, 1727, 2809, 2648, 1137]. Lightning [258]. Lights [455]. lignes [1862]. Like [1978]. Likely [2460]. Liliane [2569, 2703]. Lille [2821]. Lillian [2144]. Lim [2628]. límites [2279]. Limiting [2894]. Limitless [1235]. Limits [2615, 1886, 384, 2679]. L’immagine [1130]. L’impossible [2092]. Lin [2050, 2099]. Linares [1654]. Lincei [1367, 1802, 2719, 3020]. Lincoln [889]. Linda [1112, 2563]. Lindberg [1531]. Linden [316]. Lindley [1441]. Lindsay [64]. Lindy [1476]. Line [456, 1191, 640]. linéaire [478]. linéaires [3141]. Linear [549, 368]. Lines [22, 352, 2718, 2687]. Linguerri [1184]. Linguistic [2684, 405]. Linh [2527]. Links [951]. Lino [1641, 1641, 1596, 1630]. l’insecte [1540]. Lioba [1212]. Lipking [1808]. Lire [988, 3250, 1553]. Lisa [1801, 1775, 1422]. Lisbon [1215]. Lise [1683]. Lissa [2859]. Listening [2876]. Lister [1291]. Listing [844]. Lists [845, 842]. Literary [1818, 1898, 2251, 652, 320]. Literature [3200, 2329, 2309, 161, 2790, 385, 1362, 1129, 798, 2577, 2892, 2039, 2011, 2358, 2636, 87, 2498, 163, 930, 1741, 2935, 201, 2436, 164, 2993, 3083, 2409, 2117, 1488, 2582, 1137, 1494, 1277]. Littérature [645]. Little [1107, 3043]. Littlefield [1153]. Liu [1961, 2695]. Liv [1403]. Lived [1729, 481, 3288]. Lively [1614, 1159]. Livermore [2816]. Liverpool [2120]. Lives [285, 2217, 244, 1328, 1867, 226, 2493, 620, 2984, 144]. Living [215, 1402, 2094, 712, 747, 3122]. Livingstone [1831, 1144]. livres [3092, 1352]. Lizhen [2720]. Local [2557, 1960, 1648, 2383, 912, 1369, 3207, 3112]. Localization [2579, 2370]. Locating [2162]. l’Occident [1538]. Lock [781, 2301]. Locke [932, 2775]. Lockhart [2434]. Locomotion [2193]. Loeb [441]. Log [1667]. Logan [1224]. Logic [2281, 1590, 203, 220, 2395, 3066]. Logical [405]. Logodaedalus [2931]. Logos [1597]. Logsdon [754, 2063]. Lohr [1871]. Lohrmann [39, 1795]. Lomazzo [2929]. Lomonosov [1335, 886, 3119]. London [1598, 2486, 1568, 2262, 1586, 1577, 1543, 58, 1579, 1580, 1548, 1583, 1556, 1596, 1591, 1565, 1605, 1602, 1807, 1599, 1610, 1573, 2277, 1559, 1607, 1707, 1601, 1584, 1587, 1589, 2634, 2434, 1595, 1603, 1544, 1719, 3068, 3259, 3260, 3261, 1448, 3220, 2943, 1411, 588, 359]. Lonely [3059, 1331]. Long [24, 2019, 2346, 86, 2134, 2094, 1558, 2639, 2928, 1668, 1830, 2469, 1343, 3031]. Long-Term [1668]. Longing [1858, 1037]. Longino [1442]. Longitude [2161, 1986, 306]. Longo [2413]. Longue [1858, 2774]. Look [2682, 2905]. Looking [2729, 720, 428]. Looks [924]. L’opera [1109]. Lopes [2630]. Lopez [182, 2110, 1752]. Lordship [1872]. Loren [2136, 1331]. Lorena [727, 48]. Lorena/Brief [48]. Lorentz.} [2946]. Lorenz [954, 288]. l’organisme [431]. Lori [1744]. Lorraine [1605, 2175, 2407]. Loses [2907]. Loss [2801, 70]. Lost [2361, 2560, 183, 2539, 1605, 1420, 1820, 2219, 1742, 1875, 3240, 3051, 3120, 2903, 2908, 2967]. Loten [800]. Lothringen [48]. lotta [879]. Lotze [1974]. Loughridge [2352]. Louis [1949, 1223, 1485, 1201, 1064, 369, 369, 2441]. Louvain [274, 492]. Love [3200, 2643, 950, 2302, 623, 1313, 1741, 2940]. loved [3173, 2049]. Lovedale [2127]. Lovelock [79]. Lover [782]. Lovers [1626, 3158]. Loving [3014, 1364]. Low [1569, 792, 2996, 2932, 2133, 2796, 1409]. Lowell [390]. Löwy [2672]. LP [1504]. Lubar [2560]. Lubenow [2032]. Luc [1463]. Luca [1109]. Lucas [2657, 1872]. Lucia [1041, 2646]. Luciana [1461]. Lucier [69]. Luckhurst [1137]. Lucrezio [2100]. Lucy [1202]. Ludger [1730]. Ludic [551]. Ludwik [1842]. Luglia [2282]. Luigi [2015, 790]. Luis [1902]. Luisa [1461]. Luise [1735]. Lukas [2961]. Lumières [369, 2239, 875, 649, 374]. Lunar [1903]. Lunbeck [1758]. Lund [525]. Lundy [1725]. lune [1481]. Lung [1746, 3297]. lunga [879]. L’uroscopie [988]. Lüthy [1206, 2916]. Luzin [2211]. Luzzini [2937]. lv [1606]. Lvov [1742, 3240]. Lydia [967]. Lykknes [1069]. Lyn [2628]. Lynceae [1802]. Lynceorum [1802]. Lynch [1685]. Lynda [1780]. Lynn [2692]. Lynne [1584]. Lynx [2791]. Lysenko [2387, 2136]. Lyudmila [2891]. Lyvia [2442].

M [2045, 1182, 2735, 2643, 1341, 1400, 2119, 2262, 754, 1788, 2051, 1689, 2063, 2346, 1223, 2117, 1943, 1238, 2344, 2180, 2955, 2271, 1153, 1237, 1042, 1736, 1829, 1302, 1374, 2464, 1833, 2145, 2334, 2197, 1785, 1064, 1219, 1059, 1458, 2028, 1143, 2804, 2469, 1432, 397, 185, 686, 2792, 2493, 1963, 705, 3285, 1755]. M-Theory [3285, 1755]. M. [1979]. M.D. [3010, 1567]. Maarseveen [397]. Maarten [2693]. Maat [1932]. Maatsch [53]. maatschappij [440]. MacDonald [2172, 2869]. MacGregor [2835]. MacGregor. [2933].

Machine [941, 2329, 1533, 1996, 469, 518, 798, 2547, 991, 1725, 111, 85, 3022, 1990, 3233, 2466, 1330, 37]. Machine-Based [1990]. Machinery [2587]. Machines [2573, 1991, 1533, 923, 1674, 1617, 657, 1904, 2645, 712, 2923, 109, 1294, 2250]. machinis [39]. MacIntyre [1916]. Mackensen [2925]. MacQueen [1573, 3078]. Macrakis [1626]. Mad [3274, 3015, 1694, 2621, 1376, 807]. Mad-Doctors [2621]. Made [1402, 2518, 2954, 3122, 2058, 112]. Madhouse [1930, 2359]. Madison [305]. Madlen [2504]. Madman [2618]. Madness [816, 2892, 2867, 2399, 2576, 2831, 1930]. Madrid [1420, 1795, 3120]. Madriñán [1350]. Maehle [2444]. Maestre [2104]. Maestro [2137]. Magdolna [2060, 1867]. Magellan [2643]. Magi [2169]. Magic [121, 2269, 1124, 200, 1834, 2060, 1805, 2256, 2849, 2010, 2499, 2487, 2168]. magica [870]. magie [1470, 2697, 984]. Magistri [2182]. Maglen [1717]. magna [2502]. Magne [349]. Magnetic [335, 381]. magnética [787]. Magnetical [1987]. magnétique [1551, 3124]. Magnetism [2046]. Magnum [2758]. Magueijo [447]. mahameleth [648]. Mahon [1576]. Mahoney [1154]. Mahut [2428]. Maienschein [2837, 2656, 1629]. Maimonides [2101]. Main [1051, 1561, 3105, 1365]. Maison [1541]. Majestic [1320]. Major [409]. Majorana [447]. Make [312, 2385, 2476, 1856, 2480, 2537, 145]. Makers [435, 2646, 1707, 3220, 683]. Makes [2408, 2094]. Making [2620, 1585, 795, 1932, 2240, 2075, 665, 2710, 594, 1309, 161, 2001, 2580, 1909, 1917, 1537, 310, 1717, 1596, 1009, 584, 2351, 1071, 1677, 968, 1545, 688, 487, 602, 1599, 2118, 1879, 1168, 244, 287, 1297, 2078, 962, 1294, 1668, 2236, 2231, 1205, 897, 2544, 2448, 2654, 359, 2804, 2326, 1815, 1369, 2593, 693, 444, 146, 530, 158, 3229, 2986, 3136, 3270, 3165, 3169, 3290, 3112, 1970, 1969, 2866, 2737, 1033, 2844, 1118, 2865, 2253, 1810, 316, 646, 854, 1293]. Maladies [2268]. Malady [2167]. Malaquias [2643]. Malaria [142, 1525]. malattie [2324]. Malcolm [1622, 1211]. Malcolmson [1414]. Male [1601, 3102]. maleficarum [269]. Males [428]. Malinda [1432]. Malleable [2646]. Malleus [269]. Malpighi [797]. Malthus [2717, 1819]. Malthusian [1088]. Mamluk [1340, 3053]. Mamluks [3107, 1549]. Man [2275, 2191, 207, 194, 1931, 2531, 2460, 2365, 1057, 2118, 1651, 1985, 582, 151, 1213, 1358, 567, 411, 1008, 3030, 3040, 3063, 2428, 925]. Management [1238, 2604, 2645, 1451, 2130, 997]. Manager [2035]. Managing [965, 1609, 630, 3009, 2801]. Mancall [2635]. Manchester [301]. Manchu [2603]. Manchurian [1074]. Mandate [2806]. Manfred [2656, 1679]. Mangle [155]. Manhattan [514]. Manhood [410]. Manifest [1847]. Manifesto [1848, 1850, 1846, 1847, 1851, 1849]. Manilius [116]. Manjapra [1425]. Mankind [1058, 430, 834]. Mannered [1457]. Manning [1208]. Mantua [731]. manual [3010]. Manuals [2714]. Manuel [2643]. Manuella [2867]. Manuscript [1288, 2847, 629]. Manuscripts [2330, 3140, 1790, 813, 865, 1564, 3129, 1633, 2263, 2015]. Many [284, 2237, 226, 647, 2616]. Manzolini [656]. Mao [1314]. Map [52, 2698, 2459, 2603, 1406, 2602, 1986, 2694, 1875, 2972]. Mappa [2698, 725]. Mapped [2465]. Mapping [2667, 1812, 1147, 2667, 2278, 2961, 920, 994]. Maps [2259, 1889, 978, 1791]. Maquet [1644]. Mara [1634]. Marc [2440, 2362, 2875, 2365, 2350, 1540]. Marcacci [2850]. Marcaida [2110, 2931]. Marcel [1231, 2327]. Marcello [797]. March [1507, 11, 12]. Marchitello [2790]. Marcia [2698, 2049]. Marco [1791, 2100, 1554, 1802, 2414, 2100, 2176, 2319, 1041]. Marcon [1947]. Marconi [2365]. Marcos [1935]. Marcovich [2064]. Marcus [2806]. Mareike [2941]. Marelene [1328]. Mares [739]. Marey [1061, 2998, 1582]. Marga [1313]. Margaret [436, 659, 1801, 896, 1298, 2706, 2640, 1336, 2338, 1959, 1571, 819, 146]. Margarete [1955]. Margins [2836]. Margócsy [1703]. Margot [2538]. Mari [2581]. Maria [1554, 2414, 1126, 1471, 2195, 1202, 1836, 2815, 1058, 1706, 1463]. Mariacarla [1782]. Mariafranca [1798]. Marian [2699, 1799]. Mariangela [966]. Marianne [2528, 1328]. Marie [1419, 950, 1970, 1812, 1005, 2254, 13, 2569, 2627, 2578, 2539, 673]. Marie-Ann [1005]. Marie-Thérèse [1419]. Marieke [1788]. Marília [2630]. Marilyn [1463, 36, 1632]. Marine [1523, 2445, 2724]. Mario [1033]. Marion [2354]. Mariss [2348]. Maritime [2571, 1068, 629, 2294, 1667, 1843]. Marius [2584, 2584]. Marjorie [707]. Marjory [2286]. Mark [2281, 1954, 1302, 1716, 1936, 1114, 2466, 2873, 2241, 1111, 1945, 1239, 1085, 1034, 1327, 1536, 1125, 2880]. Market [1837, 1609, 432, 3009]. Marketing [646, 1911, 1903]. Marketplace [588]. Markets [1159]. Marks [2301]. Märkte [2504]. Marli [1307]. Marques [1834]. Marr [2931]. Marren [2489]. Marriage [509, 1143]. Marrone [2256]. Mars [1150, 1756, 667, 1089, 3265, 2153, 2739, 2145, 1978]. Marsden [1362]. Marseus [2711]. Marshall [1804, 2955, 2460]. Mart [684]. Martha [1033]. Marti [2354]. Martian [1067]. Martijn [133, 2451]. Martin [990, 1737, 1696, 2954, 272, 2196, 1291, 1396, 12, 1330]. Martina [2490, 1960]. Martine [2697]. Martínez [183, 2713]. Martini [1236]. Martis [2901]. Martyrs [348]. Martzloff [478]. Marvelous [202]. Marvels [2694]. Marwa [1137]. Marx [1273, 723]. Mary [2940, 1565, 1410, 1199, 1345, 1963]. Maryland [356]. Maryvonne [1126]. Masanori [1973]. Masawayh [1098]. Maschinenbuch [39, 3247, 1695]. Masculinity [410]. Mason [2433]. Masonic [1766]. Mass [1554, 1579, 1580, 1537, 1548, 1583, 1596, 1602, 293, 1599, 1589, 506]. Mass./London [1579, 1580, 1548, 1583, 1596, 1602, 1599, 1589]. Massa [1126]. Massachusetts [2031]. Massimo [2355, 2927]. Mast [2808]. Master [305, 561, 2724]. Mastering [858, 3078, 1573]. Masterminding [1959]. Mastery [52]. Masthead [2069, 1764, 1841, 1915, 2378, 2148, 2228, 2470, 2545, 2601, 2676, 2748, 2819, 2897, 2965, 2303]. matemáticas [1877]. Matematiche [1184, 784]. materia [2513, 2567]. Material [792, 1139, 1545, 2646, 3290]. Materialforschung [2124]. Materialien [733, 1412]. Materialism [1387]. Matérialisme [1289]. Materialist [2582]. Materially [2684]. Materialprüfung [2124]. Materialprüfungsanstalt [2124]. Materials [2859, 214, 432, 556]. Maternity [2527]. Math [1912, 1320]. Mathemata [2546]. Mathematica [813]. Mathematical [534, 2848, 2858, 389, 2432, 107, 1547, 62, 2114, 2844, 2808, 193, 2131, 2761, 644, 1404, 3036, 2995, 3162, 2512, 2486, 44, 1794, 2792]. Mathematician [1478]. Mathematicians [2947, 585]. mathématicien [2332]. Mathematics [467, 1052, 1590, 2490, 3229, 494, 2611, 861, 980, 2548, 2059, 1848, 1579, 468, 551, 2707, 471, 1194, 3258, 2628, 2808, 1305, 44, 968, 2131, 710, 472, 2425, 2854, 1320, 2836, 470, 2321, 33, 2247, 1125, 64, 1798, 3240, 3197, 3090, 30, 1677, 32, 1747, 1140, 179, 2922, 1742, 348, 261, 388, 292, 363]. mathématique [478, 2502, 944]. Mathématiques [2441, 2419, 1713, 3197, 2332, 2023, 34, 1541, 786]. Mathematization [2186]. Mathematizing [2836, 1868]. Mathes [2949]. Mathew [2060]. Matis [2535]. Matlin [2656]. Matteo [1476, 2603, 2568, 1476]. Matter [2573, 1869, 3031, 415, 2686, 1886, 313, 1293, 1045, 1558, 1208, 2954, 801]. Mattern [1635]. Matters [833, 801, 2196, 675]. Matthei [947]. Matthew [2061, 1407, 47, 1985, 525, 2573, 2054, 2112, 1516, 2943, 1901, 2532, 2167, 1411, 1211]. Matthias [1329, 736, 2657, 2147, 2889, 1048, 1878, 2625]. Matthis [2053]. Matus [2501]. Maud [2104]. Mauelshagen [272]. Maura [2808]. Maurice [2170, 1480]. Mauritius [2679]. Maurizio [2331, 2886]. Mauro [725, 1406, 2972, 2355]. Maushart [1005]. Maverick [2909]. Max [947, 1087, 857, 879, 1188, 288, 46, 1100, 288]. Max-Planck-Gesellschaft [1188, 288]. Max-Planck-Institut [1087, 1188]. Max-Planck-Instituts [46]. Maximilian [679, 2699, 2105]. Maximum [997]. Maxwell [1954, 1901, 1576, 1273, 3023, 1954, 3144, 2046, 1379]. May [2949, 474, 473]. Maya [1182, 773, 773]. Mayer [1351, 1555]. Mayhew [1819]. Mayo [2036]. Mayor [2923]. Mayoral [2722]. Mayr [596]. Mazet [1344]. Mazie [1307]. Mazower [1114]. Mazur [1547]. Mazzarello [2414]. Mazzolini [1435]. McBurney [2934]. McCall [1474]. McCarthy [2117]. McCartney [1954]. McClellan [2325]. McClelland [2363]. McConnell [1569]. McCormmach [1055, 2022]. McCray [2374]. McCray.The [1235]. McCulloch [2372]. McDaniel [1207]. McGann [1548]. McGhie [2728]. McGrath [1345]. McGrayne [637]. McKeen [160]. McKellar. [2890]. McKenny [2066]. McKenzie [1754, 3241]. McKillop [2137]. McKinnell [2217]. McKinnon [1305]. McLaren [2452]. McLeish [1341]. McMillen [1966]. McMullin [625]. McNally [2458]. McRae [1823]. McVeigh [2439]. Md [1314]. Me [1005, 2196]. Mead [819, 146]. Meaning [2742, 767, 856, 96, 2734, 2733, 1686, 3149, 3302, 1543, 328]. Meanings [536, 101, 1629, 3159]. Means [1156]. Measure [729]. Measured [2631]. Measurement [2915, 2737, 2327]. Measuring [2264]. Meat [852]. mécanique [274]. Mechanical [2641, 216, 429, 1293, 2760, 2793]. mechanicis [39]. Mechanics [1486, 2852, 883, 56, 2427, 1517, 2072, 993, 1707, 2331, 1870, 1913, 3220, 2420, 2416, 1048]. Mechanikern [28].

Mechanism [1798, 1347, 1131, 70, 2626, 2010, 3032, 2255, 797]. Mechanisms [1441, 2099, 3042]. Mechanization [1285]. Mechthild [1636]. Medal [1305]. médecin [1632]. médecine [753, 1470, 1463]. médecins [1632, 3164]. Media [1818, 1072, 832, 293, 2678, 318, 3111]. medialen [1597, 3111]. medica [1477, 2567]. Medical [1424, 536, 2527, 422, 2640, 588, 2599, 2524, 2330, 3047, 123, 2042, 2218, 948, 293, 1658, 1457, 2751, 2781, 2444, 2096, 47, 849, 1046, 2890, 877, 715, 2253, 425, 2849, 2018, 387, 3072, 3010, 3089, 1477, 2119, 3194, 1782, 2101, 1045, 1861, 1608, 1797, 25, 1108]. médicale [2179, 36, 1797]. médicales [1632, 3164]. Medicalizing [2710]. medicaminum [2500]. Medici [2343]. medicina [498, 1641, 3204, 2324, 787]. Medicinal [2071]. médicinales [2261]. Medicine [3024, 2249, 2839, 2617, 646, 121, 204, 502, 338, 2050, 815, 774, 1127, 1098, 2091, 2557, 102, 553, 2519, 2150, 1356, 1104, 1112, 49, 3142, 1073, 1917, 639, 1782, 314, 948, 541, 2682, 1496, 1457, 2685, 1552, 688, 1836, 1065, 2930, 1753, 1706, 917, 1729, 849, 1567, 2927, 1264, 232, 2301, 2572, 341, 2231, 852, 888, 2701, 1491, 1787, 1549, 1151, 2260, 728, 718, 2921, 775, 2804, 763, 2653, 1874, 2097, 1635, 1976, 2950, 925, 1195, 1564, 57, 492, 556, 1620, 444, 339, 2840, 117, 2994, 3204, 3238, 3253, 3003, 3084, 3288, 3107, 2926, 1418, 1535, 1401]. Medicine [2508, 340, 1935, 3214, 2849, 2658, 2001, 2841, 1196, 2057, 1518, 1079, 795, 1930, 880, 372, 265]. Medicines [1964, 805]. Medien [2941, 1301]. Medieval [2104, 2424, 2926, 118, 2784, 2091, 1790, 1669, 2846, 865, 1200, 1552, 1792, 1690, 1789, 2425, 2498, 1550, 2702, 1458, 1549, 2256, 119, 2563, 1872, 1469, 2916, 3026, 3184, 3107, 3275, 1693, 2010]. médiévale [785, 1551, 3124]. médiévales [2179]. Medikale [1498]. Medina [1834]. Meditation [985]. Mediterranean [1346, 2629, 555, 1458, 1924, 3017]. Medizin [1646, 2053, 3003, 2504]. Medizinbuch [2924]. Medizinische [1498]. medizinischen [940]. Medizinisches [2778]. Medusa [2711]. Meerkats [807]. Meerman [1903]. Meeting [704, 1028, 1532, 1859, 2165, 2482, 2771, 400]. Meetings [2486]. Meets [2902]. Meeusen [2497]. Megan [2951]. Megascience [84]. Megaterio [427]. Megatherium [2433]. Meghan [2431]. Mehl [992, 2417]. Meier [1958]. Meiji [2524]. Mein [947]. Meis [1471]. Melamed [1810]. Melancholia [2167]. Melancholy [2167]. Mélanges [2703, 1225, 1459]. Melanie [1790, 2802, 2040]. Melbourne [885]. Meli [2776]. Melillo [2356]. Melinda [686, 1969, 1835]. Melissa [2218, 1153, 2465]. Mellyn [1694]. Meloni [2886]. Melvin [2000]. Melvyn [1950]. memoir [2999, 1604]. Memoirs [1572, 514]. Memories [1909]. Memory [379, 1548, 914, 3163, 1275]. Men [123, 1576, 506, 1421, 3045, 3144]. Menace [1524]. Menagerie [2711, 772]. Meneghello [2662]. Mennonite [2743]. Mensch [2280, 1415]. Menschen [281, 802]. Menschengeschlechts [281]. Menschheit [2944]. Menschheitswohl [124]. Menso [2425]. mensuris [928]. mensuris/Jugendwerke [928]. Mental [1694, 1991, 2604, 3274]. Merchant [2318]. Merchants [521]. Meredith [1884]. Meridian [2622]. Meritis [1375, 3131]. Merrill [2596]. Merrills [2780]. Mesmer [372].

Mesmerism [2861]. Mesoamerica [1937]. Mesopotamian [2841, 1937]. Message [192]. messenger [3049]. Messengers [792, 311]. Messungen [928]. Mestizo [3300, 1752]. metà [1197]. Métailié. [2920]. Metalepsis [1620]. Metallurgy [1476]. Metaphern [503]. Metaphilosophy [2624]. Metaphor [2841, 775, 2596]. Metaphorik [503]. Metaphors [1996]. Metaphysical [1283]. Metaphysician [1810]. Metaphysics [617, 801]. Metapsychics [812]. Meteor [2723]. Meteorology [2479, 670, 1667, 987, 81, 2128, 2871, 599]. Methanol [2060]. Method [757, 961, 1798, 1954, 1319, 1558, 2474, 1379, 2558, 3023, 3031]. Méthode [3215, 942, 1683]. Methodenhandbuch [2699]. Méthodes [587]. méthodique [2025]. Methodists [1620]. Methodological [1091]. Methodology [2705, 801, 1051]. Methods [1460, 1988, 2829, 10, 262]. Metodo [2513]. Metrica [2177]. Metropolis [1013]. Metzger [1636]. Mexican [602, 956]. Mexico [1223, 2864, 2872, 2514]. Meyer [2867, 2121, 2680]. Meyer-Abich [2680]. Meyers [2814, 1545]. Meyerson [1225, 2585]. Mezzogiorno [51]. Mi [2429]. Mialet [1093]. Miao [736]. Michael [1598, 186, 1590, 1036, 1341, 2595, 210, 1577, 1119, 1896, 2346, 629, 1685, 1605, 2038, 1648, 2450, 1506, 2949, 849, 1000, 2917, 2106, 60, 1154, 2361, 2660, 2226, 1436, 2533, 1360, 1193, 2838, 1586, 1334, 2322, 1487, 2144, 31, 2936, 2726, 1200, 1781, 2958, 2358, 2948, 2786, 1892, 1550, 2836, 2335, 1971, 1811, 1301, 1359, 1186, 1179, 1014]. Michale [272]. Michalis [2922]. Michel [2703, 1623, 2419, 2441, 1538, 2940]. Michele [2178, 2316, 2927]. Micheletti [2139]. Michell [1055]. Michelle [1544, 2665, 1135, 1156]. Michetti [1463]. Michiel [2497]. Michigan [1545]. Microchip [683]. Microelectronics [2469]. Microscope [1629, 3159]. Microscopic [2678]. Microscopy [126, 1237]. Mid [1368, 2333, 2973, 102]. Mid-Sixteenth [2333]. Mid-Twentieth-Century [102]. Mid-Victorian [1368, 2973]. Midcentury [1320]. Middle [262, 726, 1464, 2501, 547, 189, 1124, 489, 1552, 1692, 1183, 2147, 2889, 985, 1122, 2990, 2752, 3264, 3210, 3117]. Middleton [2141]. Middletown [1585]. Midi [198]. Midwife [2704]. Midwifery [1584, 3052]. Miert [1556, 2708]. Migl [2563]. Migration [2749]. Miguel [1698, 2509, 1082, 1479]. Mikhail [1335, 3119]. Mikroskop [503]. Milan [1578, 1632, 3164, 1120]. Milbrath [1937]. Mildenberger [1498, 2292]. Milena [1822]. Miles [2411]. Milgram [410]. Milieu [1341, 2, 3206, 1551, 3124]. Militärärzte [811]. Militarization [1663, 888]. Military [1734, 1451, 925, 2976, 3244, 314]. Military-Industrial [1734, 2976, 3244]. Milk [1837, 1686, 3302]. Millard [1909, 2464]. Millennium [257]. Miller [1804, 1363, 1690, 2217, 1793, 2857, 1432, 2217]. Milling [1872]. Milsom [2915]. Milton [1190]. Miltown [83]. Mimi [1599]. Mind [1176, 372, 2652, 932, 1605, 2590, 2833, 1662, 2582, 1645, 2121, 1172, 397, 3137, 3051, 3071, 2973, 1238, 1368, 2093, 1181, 197]. Mind/Body [372]. Mindedness [2048]. Minds [2834, 1899, 2189]. Mineral [2113]. Mineralogen [2190]. mineralogiche [1461]. Mineralogie [2190]. Mineralogy [1919, 47]. Mines [506]. Miniati [1634]. miniatures [3001, 1349]. Minimum [2033]. Mining [2386, 1886, 1889]. Minna [1514]. Minneapolis [1587, 1603]. Minneapolis/London [1587, 1603]. Minnesota [1587, 1603]. Minority [2234]. Minteer [2837]. Minuzzi [2570]. mirabilibus [870]. Miracles [25]. Miraculous [1482]. Mirage [458]. Miranda [2091, 2093]. Mireille [2500]. Miriam [1184, 2276, 1352, 2253]. Mirror [368]. Mirrors [379]. Mischpulte [1597, 3111]. Misfits [111]. Misrepresentation [834]. Miss [2879]. Missile [1378, 3109]. Missiles [2739]. Missing [951]. Mission [1228, 121, 2127]. Missionary [1271]. Missions [276, 2154, 2225]. Missionspharmazie [1475]. Misticall [781]. Mitchell [1494]. Mitchinson [1658]. Mithradates [187]. Mitman [2491]. Mitte [2200]. Mittelalter [2925, 3210, 1472, 1691, 120]. Mittelalters [2990, 1636]. Mittler [743]. Mitton [1604]. Mixed [805]. mixer [3111]. mixing [3111, 750]. Mixture [1752, 3300]. Mladenović [2624]. Mnatsakanian [1333]. Mobile [158]. Mobilités [2703]. Mobilities [2477]. Mobility [1599]. Mobilizing [317]. möchte [586]. Model [2132, 3011, 1637, 2736, 3154]. Modeling [103]. modélisation [689]. Models [1755, 1617, 107, 518, 2646, 1889, 3285, 2561]. Modern [1905, 2586, 2512, 1585, 1052, 1694, 2128, 1932, 1296, 348, 1744, 2339, 1837, 534, 2848, 792, 1554, 1897, 1697, 594, 2119, 796, 2533, 2790, 2001, 2262, 112, 2447, 229, 1011, 3036, 1139, 1812, 1930, 415, 2266, 522, 1450, 2415, 1275, 1282, 2996, 815, 2432, 1020, 1323, 1003, 2091, 1868, 1474, 2017, 2322, 2150, 2461, 45, 808, 3052, 280, 2977, 709, 1427, 750, 81, 853, 1699, 271, 756, 2269, 2509, 432, 1537, 2108, 2042, 509, 1010, 1280, 1733, 2283, 2769, 547, 858, 303, 1556, 2709, 1696, 2443, 174, 932, 1860, 2575, 1085, 2645, 1290, 839, 876, 2057, 1053, 924]. Modern [2169, 560, 1552, 1348, 2397, 1671, 2931, 1208, 2729, 1885, 2487, 2357, 277, 2398, 2152, 2096, 292, 1168, 2636, 630, 1645, 660, 2033, 301, 1297, 159, 2507, 2635, 341, 1770, 1672, 2758, 1204, 1559, 1767, 1557, 2335, 1884, 1651, 278, 711, 519, 1314, 1883, 2788, 1702, 2106, 1787, 937, 1473, 1627, 2654, 2865, 2935, 1800, 738, 1266, 1593, 2653, 2256, 291, 397, 804, 2797, 1796, 2694, 185, 2326, 1815, 1743, 686, 1947, 2830, 1936, 2012, 3121, 1825, 1369, 650, 1961, 1679, 2704, 831, 444, 2806, 25, 863, 801, 2916, 2714, 1409, 1990, 2986, 3134, 3033, 3243]. modern [3063, 3196, 3071, 3076, 3272, 3274, 3278, 3165, 3283, 3106, 3289, 3112, 3299, 3226, 3296, 3224, 2281, 1584, 1593, 212]. moderna [1877, 2567, 2570, 2845]. Moderne [806, 374, 440]. modernen [719, 573]. Modernes [2014, 2072]. Modernism [2038, 2880, 1898, 266]. Modernist [388]. Modernities [94]. Modernity [2116, 1131, 2133, 2369, 201, 444, 1022, 2986, 3193, 3169, 3190, 1585, 2659, 1599, 1536, 2595, 2727, 2050, 2520, 173, 1294, 524]. Modernization [584]. Modernizing [2233]. Moderns [496, 3079, 1380]. Modes [1380, 3079]. Modifiable [162]. Modris [1138]. Mody [1237, 2469]. moitié [1203]. Mole [2909]. Molecular [1110, 685, 1082, 1500, 1008, 2299]. Molecule [2055]. Molecules [2528]. molekulare [1087]. Molella [2133]. Moleschott [2662]. Mollan [1715]. Molluscan [1423, 3005]. Molnar [2949]. molteplicità [931].

Moment [2223, 1088]. Moments [636]. Monad [794]. Monads [617]. Monarchy [1771]. Monarquía [1939]. Mond [1562, 2977]. Monde [2261, 89, 2578, 153, 786, 2342]. Mondes [2569]. mondi [3248, 1700]. mondo [2850, 595]. Money [700, 701]. Mongolian [2154]. Monica [1120]. Monique [2593]. Monism [1227]. Monkey [280]. Monkeynut [1985]. Monks [865]. Monnier [2581]. Monochord [483]. Monoclonal [2301]. Monod [1209]. Monogamy [2252]. Monopoly [2119, 752]. Monsanto [2900]. Monsters [2196, 2129, 2259]. Monte [993]. Montgomery [1115, 2051]. Montreal [1766]. Monuments [1219]. Moon [1903, 2526, 2739, 552, 481, 50, 2465, 754, 2977]. Moore [2360, 2469]. Moose [308]. Moquegua [3100, 1689]. Moral [3045, 986, 324, 271, 2549, 2733, 1593, 336, 3121, 2860, 1421]. Morality [2717]. Moralizing [2910]. Moran [1887, 2106]. Morandi [656]. Morbid [877]. Mordechai [2709, 1676]. Morel [2627]. morfología [427]. Morgan [2783, 470]. Morgen [586]. Morgenstern [507]. Moritz [756]. Morning [826]. Morocco [1418, 3024]. Moron [2048]. Morris [2586, 1869, 1005, 2133, 2264, 1371, 2285, 3127]. Morris-Reich [2285]. Morrison [1569, 495, 2796]. Morrison-Low [1569, 2796]. Morrisson [2880]. Mort [2938]. Mortal [2246]. Morus [2966, 2832]. Moses [569, 2101]. Mosley [1698]. Mosquitoes [2440]. Most [2317, 1039, 810]. Mother [1661, 2363, 3062, 2111]. Mothers [2704]. Moths [2440]. Motion [1048, 2322, 2654, 1293, 2416, 481]. Motions [1468]. moto [2513]. Mott [2041]. Motu [866]. motuum [2901]. Moulinier [988]. Moulinier-Brogi [988]. Mount [2884]. Mountain [322]. Mountaineering [1651, 3063]. mouvement [2170, 34]. Move [9]. Movement [286, 2746, 1366, 3219]. Movements [317]. movimento [1061]. Moving [2730, 2552]. Moyen [1470, 36, 1553, 988, 3250]. Moynahan [2203]. MPA [2124]. Mr [308]. Mr. [1314, 1875]. MRI [1759, 3281, 1979]. Much [630]. Muhandis [868]. Mühlenberg [1878]. Mukharji [2367]. Mulakhkhas [2565]. Mülberger [2279]. Müller [1730, 2810, 1329, 2738, 2521, 2275]. Müller-Wille [1730]. multiples [2585]. Multispecies [1452]. Mummy [2385]. Mundane [406]. Mundi [2698, 725, 870]. mundo [1939]. Mungo [2267]. Munich [1640, 3097]. Munns [1155]. Münster [789]. muntaha [2257]. Münte [2212]. Munz [2208]. Murder [2448]. Murdoch [474, 639]. Murphy [2665, 1156]. Murray [2028, 2135]. Muse [2636, 1853, 1125]. Museo [1461]. Museum [2560, 2368, 2161, 772, 1918, 762, 103, 1358, 589, 336, 65, 1955, 3040, 2657, 2745, 1493]. Museums [536, 2160, 2162, 2164, 1829, 2163, 2159, 1403, 2648, 877, 567, 3223, 2202]. Music [2117, 2108, 1397, 2834, 159, 3165, 2434, 3007, 1537, 1058]. Musical [1218, 2352, 3036]. Musicology [1528]. Musik [2925]. Musikalische [3036, 1561]. Musique [40]. Muspratts [2120]. Must [2218]. Mustard [2592]. Mutant [1021]. Mutilating [2480]. Mutio [644]. Mutualism [2717]. Muybridge [2193]. My [606, 2879, 2300]. Mycoses [2595]. Myhre [2865]. Myles [2065]. Myron [2864]. Mysteries [1791, 2919]. Mysterious [814, 447]. Mystery [1942, 553]. Mystic [207]. Mysticism [62]. Myth [1896, 2691, 3119, 1335]. Mythologies [1909]. Mythologist [924]. Mythology [924]. Myths [2440, 2923, 773, 260, 2980, 714].

N [2019, 2531, 1518, 1893, 1831, 1144, 2803, 525]. N.C [1544]. N.C./London [1544]. N.J [1547, 1534, 1676]. N.J./Oxford [1547, 1534]. N.S [24]. N.Y [1576, 1677]. Nīsābūrī [495]. Nīzām [495]. Nabonnand [1541]. nach [2944, 2209]. Nachkriegszeit [2522]. Nachrufen [1489]. Nachtomy [1557]. Nacional [1795]. Nader [1198]. Nadis [1579]. Nafis [1340, 3053]. Nagel [1145]. Nahrung [2200]. Nahyan [1340]. Naissance [36]. Naked [685]. Name [133]. Named [430]. Names [781]. Naming [2557, 62]. Nancy [1594, 1314, 686, 1961, 1127, 1576, 1506]. Nanette [386]. Nanoscale [2064]. Nanotechnologies [1235]. Nanotechnology [1237]. Nanovision [148]. Naomi [1908, 1894]. Napier [2849, 2510]. Naples [966]. Napoleon [3022, 1666, 1533]. Napoleone [2934]. napoletane [2355]. Napolitani [1197]. narcissism [3269, 1758]. Narrative [1419, 2761, 3192, 2839, 1937]. Narratives [2134, 2397, 1851, 9]. NASA [1756, 1748, 218, 3234, 3054, 3265, 1431]. nascita [2276]. Nashville [1609]. Nasim [1426]. Nasser [2691]. Nathalia [2739]. Nathalie [1210, 1664]. Nathan [1330, 2813]. Nathaniel [2191, 1079]. Nation [749, 2598, 129, 2266, 1304, 2368, 1306, 1146, 1839, 2526, 2805, 1147, 2597, 933, 519, 1752, 2278, 2741, 1743, 3289, 3300, 1060]. Nation-Building [2741]. Nation-State [1304]. National [2762, 88, 2368, 1102, 134, 2462, 1335, 2139, 602, 1297, 465, 2364, 2212, 1369, 2888, 2816, 3112, 3119, 2234, 619]. Nationale [1553]. Nationalism [2889, 1743, 3289, 2142]. Nationalsozialismus [2053, 2292, 589, 3212, 1657]. Native [2514, 2732]. Natur [3151]. Natur- [3151]. Natura [3041, 2324, 2100, 931, 1630]. naturae [51]. Natural [2116, 2512, 2835, 1052, 196, 668, 436, 1900, 1919, 1492, 726, 2266, 1572, 1102, 203, 280, 1032, 271, 2180, 2707, 308, 1141, 1051, 2853, 433, 3, 1319, 1565, 1283, 376, 1285, 2934, 2186, 1829, 2508, 2970, 1946, 968, 2497, 2971, 1488, 2433, 1792, 1403, 1729, 221, 1816, 2430, 2113, 842, 2516, 1918, 516, 762, 429, 2125, 103, 1045, 619, 799, 1843, 336, 1007, 492, 643, 1877, 2110, 3151, 2983, 3288, 3170, 3223, 271, 354, 693]. Natural-Philosophical [1051]. Naturale [1461, 2185, 2845, 870]. Naturales [2497]. naturali [1041]. naturalibus [3248, 1700]. Naturalism [1653, 1577, 1956, 3138, 3267]. Naturalist [1223, 1142, 367]. naturaliste [1882, 1184]. Naturalisti [1109]. Naturalistic [1323, 1901, 1322, 1324]. naturalistica [51]. Naturalists [1109, 3126, 2835, 1337]. Naturally [244]. Nature [1661, 2424, 809, 1176, 286, 726, 862, 212, 724, 203, 1640, 1957, 271, 308, 936, 603, 676, 547, 1959, 634, 2638, 1565, 1945, 2318, 2013, 1628, 1801, 1283, 559, 1290, 2186, 263, 1244, 2066, 1363, 458, 2690, 1034, 1046, 2937, 933, 1320, 465, 956, 544, 2733, 1288, 3188, 1383, 799, 1827, 873, 2418, 445, 691, 1947, 1007, 2012, 1688, 1181, 1291, 2341, 2010, 2933, 1313, 3030, 564, 3137, 2179, 3143, 3248, 779, 2968, 2970, 3088, 2282, 3097, 1551, 3180, 3184, 3183, 3295, 3170, 1538, 3225, 3124, 988, 3179, 3171, 1969, 3041, 22, 3062, 2805, 2498, 2635, 2288]. Nature [2012, 1662, 1716, 2939, 550, 1012, 263, 818, 251, 98]. naturel [1210]. naturelle [273, 564, 1485, 1882, 779, 369]. Natures [2516]. Naturform [1814]. Naturforschung [2811, 1807]. Naturgeschichte [593, 2348]. Naturgesetz [486]. Naturlehre [733, 1412, 2971, 208]. Naturopathic [2288]. Naturphilosophie [806]. Naturwissenschaft [1212, 1472]. Naturwissenschaften [719]. Naturwissenschaftler [1489]. Naturwissenschaftlern [3036, 1561]. natuurwetenschap [3246, 1732]. Natuurwetenschappers [440]. Naval [361, 2976]. Navarro [1502, 1877]. Navicula [865]. Navigating [2786]. Navigation [2848, 2459, 3234, 2848, 1748]. Navigational [2848, 2349]. Navigators [1748, 3234]. Navy [2724]. Nazi [585, 133, 591, 2593]. Nazified [2680]. Nazionale [1367, 2719, 3020, 3118]. Neandertal [2944]. Near [2169, 489, 1937, 364]. Nebulae [1426, 3091]. Necessity [2006]. Nederlandse [3246, 309, 953, 1732]. Need [1173, 1854, 2033]. Needham [789, 2684, 2686, 1850, 2682, 2685, 2683]. Neel [2214]. Neelakantan [2741]. Neelam [1189]. Negotiations [1460, 1782]. Nehruvian [973]. Neil [2448, 1195, 1704]. Neill [948]. Neither [2050, 945]. nel [1641, 3204, 51]. nella [3041, 205, 727, 2845, 1630]. nell’antica [2100]. Nellen [2708]. nell’Italia [51]. néolamarckisme [669]. Neptune [1205]. Nervous [2975, 1649]. Nesbitt [1690]. Néstor [1745]. Neswald [2623]. Netherlands [1287, 73, 584, 937, 397, 2472, 699, 1796]. Network [91, 135, 2597, 2828, 2761]. Networked [2365]. Networks [1523, 700, 749, 2032, 2356, 2858, 1566, 1818, 1102, 1968, 2383, 1452, 2107, 3286, 1830, 2147, 50, 3135, 948, 1751, 981]. Netz [272]. Neuber [1815]. neue [2517]. Neuen [1303, 2449]. neuer [1212]. Neuerschließung [2422]. Neufeld [1436, 2958]. Neumann [2210, 389, 507, 2210]. Neural [1173]. Neurasthenic [1306]. Neurath [65]. Neuro [1173, 2370, 1238]. Neuro-Turn [1173]. Neurohistory [1169, 1170, 1172]. Neurological [1220]. Neurologists [2042]. Neuromania [834]. Neuropsychedelia [1152]. Neuropsychiatry [2322]. Neuroscience [1171, 1580, 1169, 3055, 1786, 2935, 671]. Neuroscientific [1153]. Neurotic [2474]. Neutrality [3198, 1591]. Neutron [1083]. Neuzeit [3036, 1561, 2925, 1472, 46]. neuzeitlichen [2778]. Newcomb [698]. Newman [639]. News [1699, 2153, 3283]. Newsom [2943]. Newspaper [1744, 3296]. Newton [961, 1954, 803, 1054, 19, 1287, 1261, 1051, 1633, 2709, 2853, 1203, 2854, 1676, 1379, 1705, 3023, 3202, 3140, 19, 3263, 1876, 1676, 2638, 2417, 1809, 1212, 1287, 1273, 95, 655]. Newtonian [2852, 2857]. Newtonianism [2705, 2857]. Newtonians [2049, 3173]. Nexus [1239, 693]. Nguyen [2527]. Niall [2797]. Niccolò [2853, 493]. Nicholas [1698, 2117, 1794, 1842, 1645, 1183, 434]. Nick [2267, 2535, 1813, 2037]. Nicolai [1333, 2784]. Nicolaidis [2255]. Nicolas [992, 1470, 2417, 1152, 1832, 1551, 2848]. Nicolaus [1644, 3075]. Nicole [1344, 3038, 642]. Nicoli [1352]. Nicolò [1476]. Nicoud [36, 1632, 1463]. Niebuhr [3207, 1648, 1709, 3230]. Niels [1741, 3200, 2132, 3011]. Nielsen [1036]. Nieto [1654, 2369]. Nieto-Galan [2369]. Nietzsche [3151, 1422, 2733, 3151, 504]. Niger [1573, 3078]. Niklas [1065]. Nikolaevich [2211]. Nikolai [2211, 1067, 1828]. Nikolas [1238]. Nikolaus [2426, 1350, 3081]. Nile [2780]. Niles [1910]. Nils [1820, 1574]. NIMH [532]. Nineteenth [1523, 2490, 2802, 2191, 1910, 2032, 377, 1577, 129, 1575, 1004, 1426, 2519, 1580, 1715, 2662, 1073, 2194, 1495, 575, 2659, 1496, 1581, 376, 1147, 881, 1062, 2862, 1721, 390, 1500, 1300, 583, 2648, 877, 1918, 1144, 1491, 1830, 977, 2073, 2196, 2436, 2195, 392, 381, 661, 2413, 3055, 3091, 3167, 3227, 965, 1618, 1102, 3276, 1667, 1105, 339]. Nineteenth- [1491, 1105]. Nineteenth-Century [1523, 2802, 2191, 2032, 377, 129, 1004, 1580, 1715, 2662, 1073, 2659, 1581, 376, 1147, 881, 1062, 2862, 1721, 390, 1500, 1300, 583, 2648, 877, 1144, 2073, 2196, 2436, 392, 3055, 3167, 965, 3276]. Ning [2446]. Nino [1556]. Ninth [3172]. Nirenberg [2460]. Nixon [2063]. Njåstad [349]. No [2391]. Noah [1115, 1738]. Noam [1767]. Nobel [2060]. Nobility [2712, 699]. Noble [2664]. Noblewomen [1280]. Noel [63, 1737, 1035, 3273]. Noëlle [2578]. Noémi [2817]. noetics [3143]. Noise [1118, 451]. nombres [252]. nome [890]. non [2355, 785, 3188]. non-western [3188]. Noninvasive [321]. Nonlinear [2875]. Nonmathematical [3]. Nonproliferation [2297]. Nonsense [1311]. Noorden [2777]. nor [2050, 945]. Norbert [443]. Norbrook [1202]. Nordmann [1086]. Normal [965]. Normality [2800]. Norms [1318, 479]. North [1228, 2571, 679, 1229, 2026, 2425, 306, 1593, 245, 445, 1705, 3263]. Northern [1223, 2472, 2107]. Northumberland [554]. Norwegian [349]. nos [350]. Not-So-Deep [1172]. Notation [1547, 264, 3162]. Note [1528]. Notebook [1370, 3123]. Notebooks [2017]. Notes [99, 156, 235, 329, 407, 462, 528, 612, 695, 759, 836, 903, 963, 1018, 1095, 1161, 1245, 1315, 1381, 1443, 1760, 1520, 1611, 1665, 1840, 1081, 813, 284, 993, 2614, 1705, 2177, 3263, 289]. Nothaft [2846, 1399, 1693]. Notieren [635]. Notion [2305]. Notions [1391]. Notizen [733, 1412]. Notorious [1057]. Nouveau [2261, 371]. nouvelle [371]. Nova [992, 1476, 566, 2716]. Novalis [873]. Novara [1471]. Novas [1479]. novel [3045, 1300, 1221, 1421, 457]. Novelties [1560]. November [210]. Novoa [1495]. Noyan [739]. NS [2454]. NS-Herrschaftssystem [2454]. Nuclear [2463, 399, 2531, 2134, 955, 1241, 2462, 827, 900, 1265, 447, 2954, 1083, 952, 2140, 2463, 1071, 1823]. Nucleus [749]. Nudes [2193]. Number [159]. Numbered [2122]. Numbers [2191, 81, 2411, 1784, 2782, 2320]. Numerical [264]. Nuncius [789, 192]. Nunes [2095]. nuova [51]. Nurture [818, 1313, 3295, 1688, 458]. Nurturing [2950]. Nutritional [2623]. Nutton [2632, 1196]. Nützliches [2347]. Nye [1191]. Nyhart [2692]. Nystrom [1889].

O [24, 1790, 2928, 147, 1343, 789, 566]. O-Rings [147]. Oaks [165]. Obedience [410]. Obedient [3176, 2207]. Obelisk [1962, 24]. Obesity [1750, 3237]. Object [1993, 804, 3296, 1744]. Objective [1094]. Objectivity [237]. Objects [792, 1868, 468, 618, 641, 575, 1204, 2563, 1783]. Obra [3147, 1654]. obras [351]. O’Brien [2048]. Obscura [1329, 3034, 2564]. Obscured [521]. Observation [2020, 2151, 525, 632]. Observations [1556, 3224]. Observatories [377, 666, 2889]. Observatory [2877, 2869, 2959, 1375, 3131]. Observers [1215, 1119]. Observing [3091, 1426]. Obsessed [244]. Occasion [1648, 3207]. Occident [178, 3030, 984]. Occult [3056, 1209, 1410]. occulte [1551, 3124]. Occultism [812]. Occupational [456]. Occupied [2667]. Ocean [2249, 360, 1251, 2956, 2679]. Oceania [253]. Oceans [1252]. O’Connor [1362, 1137]. octo [2103]. October [210, 1327]. Oddi [644]. O’Dell [1434]. Odile [2502]. Odyssey [3140, 1633]. Oedipus [1292]. Oetzel [1310]. oeuvre [1594, 3141]. Oeuvres [2991, 1485, 1650, 1564, 3129, 1485]. Ofer [2849]. Off [496, 1604, 2999]. Öffentlichkeit [120, 401]. Office [709, 1317, 2677, 2073]. Official [2134, 397]. officielle [478]. Ogden [2861]. oggi [1461]. O’Gorman [2356]. ogród [1865]. Ohad [1557]. Oikumene [723]. Oil [69]. Oinoanda [2627]. Okholm [1883]. Ökonomische [745]. Okzident [868]. Ola [1460]. Olaf [1329]. Olah [2060]. Old [86, 2641, 991, 2206]. Oldenziel [1906]. Oldfield [2125]. Ole [2508]. O’Leary [1501]. Olfactory [2802]. Olga [110]. Olival [1913]. Oliveira [2416]. Oliver [789, 1157, 1967, 2369]. Olivier [2723, 2006]. Olmi [2845]. Olostro [2355]. Olschki [1560]. Olshin [1791]. Olson [2130]. Omar [1426]. Omens [634]. Omodeo [2107]. Omond [2137]. Oncology [1084]. Ondernemer [2126]. One [1326, 1615, 2401, 853, 2725, 1627, 3076, 1975, 857]. One-Sex [1627, 3076]. Ones [2612]. Onion [2293]. Online [2830]. onlookers [3183]. Only [2032]. Onnes [73]. Ono [2300]. Ontology [2394]. Oosterhoff [2844, 2931]. Oostvaardersplassen [2768]. op [2442, 2451]. Open [1662, 3286, 3137, 1751]. Opened [2541]. Opening [823]. Opera [928]. Operational [217]. opere [1417, 3118]. Opiate [949]. Opinion [2632]. Opitz [1069, 2326]. Opponents [1822]. Opportunities [2825, 2760]. Optical [68, 1898]. Optics [2787, 1699, 2698, 2264, 3086, 2636, 977, 219, 1345, 3283, 1936]. Optik [2422]. Opus} [2758, 38]. Oracle [324]. Oral [188]. Orange [1236]. Orbit [2464]. Orbits [891]. Orchids [1958]. Order [2283, 173, 1628, 2518, 1320, 1288, 2677, 2940, 3031, 3225, 1558, 2113, 1178]. Ordered [972, 2706]. Ordering [2323, 1142]. Ordnung [2925, 2343, 23, 1892]. ordre [1210]. Orel [1928]. Oren [1598, 2917]. Oreskes [1908]. Oresme [1344, 642, 3038]. Oreste [51]. Organ [1462, 66, 2978]. Organic [2457, 1370, 1740, 3293, 3123]. Organicism [1415, 3087]. organique [587]. Organisation [1460, 1892, 1475, 442]. organischer [23]. Organisms [1402, 2, 2528, 2534, 3122]. Organismus [1892]. Organization [1824, 1025, 1026, 1674]. Organizational [1166, 1374, 3012]. organizations [3304]. Organizing [1024]. Ori [2855]. Orient [868, 984]. Oriental [939]. Orientalische [2337]. Orientalism [391]. Origin [2199, 1910, 2784, 1842, 1612, 694, 1676, 3202, 2784]. Original [1644, 3075]. originalité [3030, 1538]. originality [3030]. origine [171]. origines [587]. origini [3039, 1357, 1461, 1490]. Origins [435, 735, 2848, 2852, 1568, 2119, 307, 1575, 2091, 1014, 568, 183, 703, 916, 766, 2444, 279, 2197, 2274, 2121, 3166, 2644, 1891, 921, 983, 2587, 863, 3039, 3006, 2649]. Orinoko [743]. Orioli [1802]. Orit [2221]. Orly [2496]. Ornithology [2728, 1214]. Oronce [494, 2332]. Orphanage [2020]. Ørsted [1181, 1181, 946]. Ortega [2620]. Orthodox [1922, 1921, 1923]. Orthodoxies [1925]. Orthodoxy [796, 982, 1927, 1926]. Ortiz [1836]. Ortiz-Gómez [1836]. Orton [2299]. ortus [1342, 2987]. Oscillations [2875]. Oslo [2865]. Osram [586]. Osselaer [2519]. Osseo [1687]. Osseo-Asare [1687]. Osteoporosis [1893]. Österreich [1233, 1972]. östling [1891]. Ostorero [2697]. Oswald [1416]. Other [1931, 2906, 2934, 714, 789, 260, 2493, 2980, 688]. Other-Worldly [688]. Otherwise [430]. Otlet [2044]. Otmar [238]. Ottaviani [51]. Ottmar [2799, 1219]. Otto [1188, 2711, 65]. Otto-Hahn-Institut [1188]. Ottoman [1354, 2877, 3027, 1167]. Oude [2777]. Ought [1324]. Our [1080, 360, 1757, 2223, 2291, 2000, 368, 2122, 1254, 3216, 2888, 827]. Ours [2874]. Outer [2894, 2142, 1077]. Outsider [3208, 1598]. Outstanding [2298, 355]. Outward [2582]. Ouverte [1540]. Overcoming [907, 3190]. Overcontextualization [1852]. overheidsbeleid [2442]. Overpotential [1009]. Overseas [991]. Overtook [1265]. Overture [508]. Overview [1921]. Oviedo [1038]. Owen [449, 838, 2003, 1496, 59]. Own [827, 606, 2725]. Ownership [466]. Owning [2088]. Owns [1513, 2088, 466]. Oxford [1590, 294, 261, 1547, 2832, 1534, 1933, 918, 1455, 1557, 2321, 2921, 1015, 2097, 1608]. Oxymoron [911]. Oyster [356].

P [1502, 1155, 1005, 2837, 176, 2805, 2812, 2133, 1456, 1887, 1488, 2277, 1607, 2256, 1635, 2565, 2168, 1437, 2413, 133]. P. [1416]. Pa [1606]. Pablo [2507]. Pacific [1704, 2026, 1570, 3189, 3145]. Package [545]. Packard [2493]. Packer [1786]. Padrón [1939]. Padua [2926]. paesaggio [2176]. Paetrow [2273]. Paffenroth [1624]. Pagan [1334, 3104]. Paget [2450]. Pain [1864, 1864, 2619]. Painkillers [1864]. painted [2990]. Painting [368]. Palace [2650]. Paleoanthropology [1384]. Paleobiological [229]. Paleobiology [1519]. Paleontological [2154]. Paleontology [838, 567, 229, 835]. Palestine [2885, 2667]. Palestrina [1528]. Palgrave [2790]. Pallas [53]. Palló [1973]. Palmer [1935]. Pamela [24, 1793, 1545, 2928, 1343]. Pampu [2488]. Panacea [1009]. Panaceia [1280]. Panama [1668]. Pancino [2324]. Pang [1372]. Panisperna [2276]. Panopticon [2619]. Panoramic [2160]. Panti [1341]. Paola [1867, 2289, 2641]. Paolo [1641, 2414, 2929, 1197, 2796, 1075, 2791, 1700]. Papadopoulos [1862, 2720]. Papal [1351, 3085]. Papanelopoulou [1998]. Paper [1744, 2547, 1008, 1598, 1585, 1588, 1541, 1554, 1594, 1578, 1543, 1572, 1566, 1535, 1556, 1597, 1551, 1571, 1542, 1587, 1540, 1538, 1553, 1595, 1603, 1544, 3296, 923]. Paperclip [2888]. Papers [1507, 3140, 1644, 1633, 3075, 1064, 883, 1195, 195]. Paperwork [842]. Parable [227]. parabolae [1197]. Paracelse [2184]. Paracelsian [2106]. paracelsischen [1631, 3058]. Paracelsus [1212, 3058, 121]. Parade [881]. Paradigm [1953]. Paradigmatic [789]. Paradigms [2769]. Paradise [633]. Paradox [1437, 3094]. Paradoxes [531, 1388]. Parak [2523]. Parallel [2696, 2263, 2101, 2791, 3248]. parallèles [1862]. paralleli [3248, 1700]. paranormales [2279]. Parapsychology [452]. Parascandola [1116]. Parasites [1264]. Paravicini [2697, 2423]. Paredes [2659]. Paris [1541, 58, 2579, 808, 2146, 789, 1551, 369, 1353, 1540, 1538, 1553, 966, 3021, 1949, 1073, 583]. Park [2206, 1582]. parla [211]. Parliament [2650]. Parmenides [1465]. Parot [1470, 1551]. Parra [2567]. Parshall [2106]. Parsons [1715, 2955, 3276, 962]. Part [1432]. Participation [317, 1148, 2764]. Particle [1606, 3057]. particulière [1485, 369]. partir [478]. Parts [838, 2450]. Parva [2821]. Pascal [2573, 1129, 2008, 2993, 2014, 26, 2290]. Pascale [1575]. Paschoud [1210]. Pascual [2680]. Pasetti [1041]. Passage [398, 1936]. Passion [1304, 509, 308, 1399, 2981]. Past [2311, 665, 2309, 1460, 2560, 1207, 618, 605, 1513, 2006, 2397, 2312, 3284, 1170, 698, 1861, 2593, 1680, 1040]. Pasteur [293]. Pasts [2310, 2308, 2407, 2313]. Patent [1569, 3146, 2073]. Patenting [531]. Patently [3025, 1583]. Patents [1063, 702, 2065]. Path [780, 1237, 552, 2288, 527]. Pathogens [1229]. Pathological [2776, 2129, 1874]. Pathologies [2856]. Pathologist [2121]. Pathology [532, 1270]. Pathway [2734]. Pathways [179, 1785]. Patience [2045, 1723]. Patient [2949, 1220]. Patients [1938]. patria [1578, 3172]. Patrice [1650, 1470, 2697]. Patricia [274, 2725]. Patrick [2783, 2509, 2374, 1432, 1560, 1235, 1347, 2488]. patrioti [1367, 3020]. Patriotism [1960, 132]. Patrizia [2807]. Patrizio [1462]. Patronage [700, 1812, 1239, 702]. Patterns [2599, 874, 1682, 3232]. Patterson [1953]. Paul [883, 207, 77, 2945, 1757, 677, 1644, 2663, 2453, 789, 1064, 1702, 2029, 1059, 69, 397, 2044, 380, 2916, 1437, 2058, 2895, 1333, 2605, 1160, 1757, 1339, 2731, 2180, 1209, 1605, 2462, 1373, 1602, 1244, 1374, 1567, 1076, 1911, 3216, 1542, 1899, 1587, 2921, 1979, 1022]. Paula [2934, 43, 2195, 1737, 1204]. Paulian [1081]. Paulin [2270]. Pax [2307]. Paying [703]. Payne [1047]. Peace [2137]. Peacock [968]. Pearce [2267, 2087]. Pearls [3240, 1742]. Pearn [1059]. Peasants [602]. Pechenkin [2870]. Peck [2535]. Peck-Kubaczek [2535]. Peculiarities [2022]. Pedagogical [1844]. Pedagogy [1506, 244]. Peer [2550, 1450]. Peeters [2201]. Peggy [64]. Pegler [2015]. Peirce [1988]. Peking [151]. Pemberton [2448]. Penal [2658]. Pending [1851]. Peninsula [2467]. Penn [2736]. Pennsylvania [1555, 1582]. Pensée [1632, 3164, 1289]. Pensées [89]. People [1585, 2895, 844, 2746, 2333, 151, 230, 2986, 1078]. Peoples [881, 1219, 2493]. perception [3210]. Pérez [2569, 2703]. Perfect [3220, 1707]. Perfection [1079]. Performance [2201, 240, 243]. Performativity [244]. Performing [849]. Peril [828]. Period [756, 2701, 3036, 2339, 1053, 937, 801]. Periodic [2870, 1973]. Periodical [2525]. Periodicals [1450]. Periodization [1994]. Periphery [213]. permanente [478]. Perplexed [182]. Persistence [1896, 515]. Person [2054, 1716]. Personal [334, 800, 1604, 238, 2999]. Personality [2022, 1100]. Perspective [212, 1850, 1960, 2929, 1922, 368, 2546, 3245, 1833, 2584, 1033, 1681, 2544, 899, 653, 979, 607]. Perspectives [262, 2439, 883, 1954, 1689, 2171, 548, 1648, 2169, 914, 487, 2302, 2125, 1680, 3207, 3284, 3100, 3225, 1628, 2692, 1836, 2090, 864, 325]. Perspektiven [806]. Pertinence [664]. Peru [3231, 1708]. Perusing [2311]. Peruvian [2385]. Pesic [1537, 2834]. peste [3203, 1641]. Pesticide [821]. Pestre [1243]. Peter [1339, 833, 2271, 736, 2621, 1033, 1784, 2508, 1517, 2433, 2701, 1958, 2212, 1869, 2360, 1954, 1982, 2120, 1537, 2468, 1299, 846, 2800, 2141, 2126, 1581, 1482, 2169, 1934, 914, 2834, 2489, 1113, 2635, 1651, 2701, 1752, 2794, 2733, 2258, 1379, 1084]. Peters [2597]. Peterson [2457, 2095, 1125]. Petite [2811]. Petites [1225]. Petr [1373]. Petroski [1117]. Petrus [2021]. Petteri [2831]. Pettit [1586]. Petto [1812]. Petty [1288]. Pflanzengeographie [3150, 1574]. Pflanzentransfer [272]. Phaenomena [1464]. Phalkey [1438]. Phantasie [2990, 1636]. Pharmaceutical [2119, 1080, 1098, 672]. pharmaciens [587]. Pharmacological [188]. Pharmacology [2375]. Pharmacopoeia [1964]. Pharmacy [1917, 1183]. Pharmazeutische [378]. Phenomena [2523, 2519, 660]. Philadelphia [1572, 936, 1555]. Philanthropy [1427, 3106]. Philip [2949, 2653, 1195, 1895, 1430, 1049, 2721, 2009, 1089]. Philipp [2846, 1399, 1693]. Philippe [1905, 1541]. Phillip [2066]. Phillips [1912]. Philological [2855]. Philologies [1530]. Philology [129, 1575, 2096, 3166, 1702, 3243]. philosophandi [727]. Philosopher [690, 2291, 1563, 2102, 2305, 3077, 56, 95, 655]. Philosophers [2793]. Philosophical [326, 1572, 1051, 835, 2670, 2692, 2781, 325, 2302, 267, 694, 2399, 2627, 746, 195, 1666, 2438, 2431, 2952, 2574]. philosophicum [438]. Philosophie [3139, 1203, 867, 273, 1540, 2441]. philosophizing [3248]. Philosophy [1052, 1326, 1880, 1227, 690, 780, 436, 726, 2798, 1408, 203, 1206, 2426, 154, 2180, 1714, 721, 833, 1051, 327, 1696, 2186, 2779, 2508, 27, 1946, 2971, 1854, 405, 1951, 2430, 2633, 54, 1542, 1293, 2760, 1045, 201, 119, 2706, 2642, 999, 70, 492, 1108, 2940, 1623, 2016, 2340, 3029, 3194, 3143, 3060, 3071, 3266, 3272, 3087, 3278, 3139, 1557, 2701, 2098, 1782, 2707, 2853, 1208, 1792, 1645, 2498, 1415, 2012, 271, 833, 1285, 429]. Photograph [2525]. Photographs [2523]. Photography [2491, 1072, 379, 299, 1582, 1544, 2285, 2998, 3110]. Phrase [767]. Phylloxera [662]. Phylogenetic [1749, 3254]. Physica [1472]. Physical [1247, 2866, 856, 1302, 27, 1827, 485, 1008]. Physicam [1344, 3038]. Physician [1792]. Physicians [2812, 1946]. Physicien [1499]. Physicist [1005, 56, 2291, 1071, 3049, 127]. Physicists [1302]. Physico [1798]. Physico-mathematics [1798]. Physicorum [2182, 2103]. Physics [1590, 1895, 1450, 73, 2039, 998, 817, 945, 84, 1606, 2388, 2554, 2954, 857, 2843, 2898, 1378, 1907, 578, 2416, 2761, 1324, 820, 999, 233, 3144, 3057, 3109, 1462, 2141, 1576, 2006, 2089, 2742, 1897, 1511, 1455, 1899, 607, 952, 406]. physikalischen [28]. Physiker [296]. Physikunterrichts [208].

Physiognomy [392]. Physiological [2038, 1941]. Physiologie [1087]. Physiology [2798, 3210, 2245, 1679, 3196]. physionomie [571]. Physique [3, 992]. Physis [2680]. Physischen [1412]. Phytopathology [255]. Pia [1706, 1463]. Picart [558]. Picatrix [984]. Pickering [1577, 610, 1559, 1584, 2940]. Pickstone [1454]. Picture [1990, 3210]. Pictures [1779, 590, 2654]. Picturing [2893, 1234, 1046, 293]. Pieces [1980, 1225]. Pier [1197]. Pierce [2840]. Piero [1406]. Pierre [950, 1286, 779, 2417, 2602, 2627, 1882, 2821]. Pierre-Marie [2627]. Pierre-Yves [1882]. Piers [2717]. Piet [2708]. Pieter [866, 1891]. Pietikäinen [2831]. Pietism [2020]. Pietro [1470, 2276, 2107]. Pietruska [2729]. Pietsch [2032, 1816]. Piezoelectricity [406]. Pig [254]. Pigs [2534]. Pilaski [1640]. Pill [602]. Pills [83, 898]. Pilots [2812]. Pimentel [2433]. Pingree [184, 1785, 1785]. Pinto [2268]. Pioneer [2030]. Pioneering [1962]. Pioneers [2360]. Piot [1982]. Pious [755, 2788]. Pipes [1462]. Piracy [426]. Pirot [2008]. Pisa [58]. Pisano [2004, 2427]. Piśmiennictwo [1865]. Pitfalls [2760]. Pittsburgh [1606]. Pittura

[2929]. PKU [1437, 3094]. Place [415, 1681, 277, 1831, 465, 1668, 2399, 2284, 3245, 2649]. Place-Based [2284]. Placebo [1753, 3253]. Places [2616]. Placing [240, 581]. Plague [3033, 1074, 3203, 1070, 1348]. Plakins [2191]. Plamper [1192]. Planck [1087, 857, 879, 1188, 288, 46]. Planet [360, 1681, 3245, 3104, 1334, 2003, 2062, 506, 667]. Planetarium [2666]. Planetary [1960]. Planets [1468, 2465]. Planning [2884]. Plant [2386, 1035, 2518, 3150]. Plantation [1725, 3233]. Plantations [1222]. plantes [2261]. Planting [633]. Plants [134, 1350, 278, 2260, 1687, 3081, 3279, 577]. Plato [2587, 388]. Platone [1803]. Platonem [35]. Platonic [2702]. Playing [623, 517, 258]. Pleasure [2619]. Pleins [1724]. Plesser [1087]. Pliny [1278, 643]. Ploeger [88]. Plofker [30]. Ploy [1957]. Pluralism [2527, 1994, 1861]. Pluralité [1126]. Plutarch [2497]. Pluto [1336, 3125]. Pneuma [2496]. poder [3195, 1745]. Poe [2803]. Poetic [1937]. Poetics [161, 873, 2336]. Poetry [2188, 2141, 1552, 732, 3117, 2270, 63]. Poggi [1461]. Poincaré [1216]. Poincaré. [2883]. Points [3124, 1551]. Poison [2054, 2056, 2357, 949, 187]. Poisonous [882]. Poisons [1116, 2181]. Poisson [1713, 3197]. poissons [779]. Polanyi [1014]. Polar [2863]. Pole [335]. Polemic [2908]. polemiche [211]. Poles [2396]. Police [128]. Policies [1983, 2623]. Policy [1430, 404, 2953, 2541, 2138, 898, 619, 2306, 322, 3050, 460]. Polio [293, 2957]. Poliquin [1037]. Polish [1563, 2333]. Politic [1482]. política [1479]. Political [215, 1864, 68, 1253, 1912, 1918, 1702, 1853, 1874, 758, 691, 2664, 238, 3243, 2886, 2717, 887]. Politically [1391]. Politics [597, 1588, 1813, 821, 2059, 2813, 1070, 2252, 2368, 2588, 2528, 1236, 2150, 518, 2662, 399, 756, 1904, 2095, 134, 1591, 1253, 998, 1482, 1239, 2105, 2074, 1821, 607, 1354, 1062, 1662, 1756, 2139, 277, 2690, 87, 2033, 2717, 1831, 1767, 1607, 591, 1120, 150, 2359, 949, 921, 1740, 999, 2872, 2940, 2543, 117, 3027, 3137, 3048, 3152, 3265, 3198, 3293, 1982]. Politik [2053]. politique [122, 2008]. Politiques [1243]. Pollution [2357, 1536, 3193]. Polo [1791]. Pols [2950]. polvo [787]. Polycentered [972]. Polymathy [2107]. Polymer [1953, 1953]. Polyphonic [2834]. Polytechnic [74]. Polyvalente [2354]. Pompa [2343]. Pomponazzi [987]. Pont [1499]. Pontecorvo [952]. Pontificia [75]. Poole [2637]. Poonam [1491, 20]. Popescu [2488]. Popescu-Pampu [2488]. Popper [405, 2485]. Popplestone [1394]. Popular [1600, 1967, 320, 293, 924, 864, 2293, 2919, 2857, 151, 2904, 3028, 1001]. Populäres [2143]. Popularization [2662, 131, 622, 2911, 2904]. Popularizer [286]. Popularizing [2910, 411, 2589, 213]. Population [1660, 532, 1088, 3133]. Populist [2530]. Pormann [2701]. Porsild [1232]. Port [1717, 3270]. Porta [2787]. Portable [2626]. Portal [2773]. Porter [2271, 1380]. Portfolio [490]. Portolan [2784, 1612]. Portrait [2102]. Portugal [2630, 2460, 2930, 2163]. Portuguese [43]. Portulan [2505]. Portulan-Atlas [2505]. Positivism [2396, 2940]. Possessions [2198]. Possibilities [2252]. Post [1609, 2629, 2508, 1513, 2060, 686, 3009]. Post-Antique [2629]. Post-Nobel [2060]. Post-Reformation [2508]. Post-Soviet [1609, 3009]. Postclassical [1172]. Postcolonial [973, 974, 2817, 1090]. Postcolonialism [971]. Postcolonialities [94]. Postgenomics [2738]. Postmodern [3047, 2900, 1608]. Postmortems [2788]. Postpositivism [971]. Postulate [2263]. Posturing [238]. Postwar [2599, 87, 80, 238]. potere [1652, 3115]. Poulsen [2445]. pour [2282, 2938, 252]. Power [749, 2191, 1812, 592, 2942, 1712, 2541, 1009, 1065, 82, 465, 956, 898, 578, 1388, 1825, 2140, 1872, 3291, 3115, 2718]. Powered [1274, 884]. Powerful [1039]. Powers [1056, 2167, 3162, 1547]. Poynter [1219]. Pozzuoli [205]. pp [1598, 1585, 1588, 1541, 1590, 1569, 1554, 1562, 1568, 1594, 1586, 1578, 1577, 1543, 1575, 1572, 1600, 1579, 1609, 1566, 1535, 1580, 1537, 1547, 1548, 1583, 1563, 1556, 1596, 1591, 1565, 1605, 1534, 1576, 1581, 1602, 1606, 1558, 1677, 1597, 1545, 1552, 1539, 1555, 1599, 1610, 1546, 1573, 1567, 1561, 1559, 1557, 1607, 1601, 1551, 1571, 1314, 1536, 1592, 1542, 1549, 1584, 1587, 369, 1593, 1604, 1676, 1589, 1540, 1570, 1564, 1538, 1560, 1608, 1574, 1553, 1595, 1603, 1582, 1544]. Prachtexemplar [2505]. Practical [101, 1063, 2568]. prácticas [1877]. Practice [450, 1916, 1685, 1565, 1449, 193, 2949, 1567, 2937, 155, 2473, 2804, 160, 339, 3236, 3168, 3304, 3170, 3010, 1684, 756, 1084, 2849, 2916, 128]. Practices [214, 1460, 2317, 2475, 459, 2131, 2106, 1375, 3131, 2623]. Practitioners [2512, 2979, 1343]. praeclarum [38]. Praecox [816]. Praelectio [2502]. Praescriptiones [1802]. Pragmatism [2624]. Praise [1724, 3280]. Prakash [1222]. Praktiken [2941, 2348, 1489]. Prasad [1759]. Prasenjit [2090]. pratiche [2570]. Pratik [1356]. pratique [191]. pratiques [587, 1632, 3164]. Pratt [2436]. Pratt-Smith [2436]. Praxagoras [2496]. Praxis [2124, 1475, 2778]. Prayer [1864]. präzisierte [990]. Pre [2660, 2140, 122]. Pre-Chernobyl [2140]. Pre-Freudian [2660]. pré-industrielle [122].

Precancer [713]. Precious [2234]. Preclassical [1048]. Predicting [698]. Prediction [2318, 2729]. Preece [1423]. Prefiguring [652]. Pregnancies [2672]. Prehistoric [1384, 2196]. préhistorien [942]. Prehistory [2762, 1953, 2202, 762, 1522, 396, 2775]. Prejudice [1105, 1100]. Premedieval [1612]. première [571, 1203]. Premodern [1782, 2840]. Prenatal [2672]. Prentice [1328]. Prescribed [1011, 479]. Prescription [1011]. Present [2162, 1861, 3232, 2560, 1966, 1682, 2006, 717, 263]. Presents [2407]. preservation [3203]. preservatione [1641, 3203]. Preserving [2835]. President [92]. presidenti [1367, 2719, 3020]. Press [1598, 1585, 1588, 1590, 1568, 1586, 1543, 1600, 1579, 1609, 1566, 1535, 1580, 1537, 1547, 1548, 1583, 1596, 1565, 1605, 1534, 1602, 1606, 1545, 1555, 1599, 1610, 1573, 2525, 1557, 1607, 1601, 1571, 1536, 1542, 1587, 1593, 1676, 1589, 1570, 1608, 1595, 1603, 1582, 1544]. Presses [1594]. Preston [1416, 2336, 41]. Presumptions [933]. Pretzel [2901]. Preußen [743, 2027]. preußischen [61]. preuve [942]. preuves [753]. Prevalence [2061]. Prevention [2599, 2217]. Preventive [713]. prévoir [297]. Price [2401, 703, 1364]. Pride [465, 1105]. Priest [2638]. Priestess [1360, 3067]. Priestley [2174, 1376]. Prima [1802, 2324, 2845]. Primate [2051, 951]. Primates [2051]. Primatology [319]. Prime [2622]. P̈rimi [2937]. primo [498, 51]. Primroses [1021]. Prince [1572, 1098, 1635, 1632, 3084, 3164]. princeps [2100]. Princeton [1656, 1547, 1534, 2734, 1676, 3061]. Principe [1400]. Principia [2857, 813]. Principles [1682, 2828, 3232, 2632]. Prinke [1865]. Print [1031, 494, 2844, 278, 2857, 2028, 1666, 1031]. Printed [1790, 44]. Printing [2927, 3017, 1346]. Priority [2854]. Prisoners [3158, 1626]. Private [2815, 1909, 466]. Privatizing [684]. Prize [169, 417, 1255, 2060]. Prizes [1859, 1029]. Probability [1129, 2993]. Probe [1237]. Probing [3128, 1372]. Problem [2906, 2161, 2812, 1513, 2817, 2416, 1754, 1956, 2765, 747, 3241, 1113]. problematica [954]. Probleme [806]. Problems [2318, 2497, 2828, 1452]. Proceedings [1975, 1648, 3207, 3194]. Process [3061, 1656, 681]. Processes [720, 2870]. processo [199]. Procreative [873]. prodigieuses [777]. Producing [1673, 2140]. Product [515]. Production [2859, 1012, 817, 1033, 4, 699, 3111, 1812, 1269]. Products [861, 1686, 3302]. Produktion [378]. Produktionsapparate [2354]. Produktionsverfahren [1597, 3111]. proef [2442]. Profane [1157]. profanos [859]. Profayt [2104]. Profession [915, 2581]. Professional [320, 238]. Professionalism [588]. Professionalization [1503]. Professor [553]. professore [879]. Professors [222]. Profit [2518]. Prögler [1483]. Prognostication [1178]. Program [1909, 1190, 1157, 1903, 2063]. Programm [441]. Programmed [2539]. Programmers [597]. Programmes [564, 1241]. Programming [2296]. Programs [1590, 3108]. Progress [761, 293, 397, 2940, 715]. Progressive [983]. Progressu [866]. Project [1261, 514, 2888, 2605]. Projects [602]. projeto [566]. Projit [2367]. Proletarian [1067]. Proliferation [827]. Prometheus [1576, 1677]. Promise [828]. Promotion [2468]. Prone [111]. pronostici [1471]. Proof [469, 551, 789, 1404, 2995]. Propelled [2739]. Properties [1449, 2775]. Property [2119, 1970, 1033, 426, 2976]. Prophecy [189]. Prophet [1819, 298]. Prophets [266, 796, 1780]. Prophylactic [713]. Proposal [619]. Proposed [41]. Propriété [442]. Propriety [1536, 3193]. Propulsion [2145]. Prospective [915]. Prospero [375]. protéger [2282]. Proteins [1082]. Protest [1310]. Protestant [491, 2872]. Protestantism [1926]. Protestants [1766]. Protevi [1603]. Proust [2660]. Providence [761, 1656, 3061]. Provincial [849]. Provine [2315]. Provocative [287]. Provoked [808]. Prozac [83]. Prozesse [1814]. Prudence [1689]. Prussia [448]. Pseudo [2257]. Pseudo-Ibn [2257]. Pseudoscience [1010]. psicologia [566]. Psychiatrist [677]. Psychiatry [2447, 2892, 2867, 2474, 2173, 2399, 2359, 2121]. Psychic [660]. Psychical [452, 812, 2520]. Psychoanalysis [1659, 2370, 2992, 2721]. Psychoanalytic [2949]. Psychological [1113, 1452]. Psychology [2210, 2439, 1586, 2807, 237, 2604, 1614, 111, 160, 623, 2438, 3095, 899, 1176, 863]. Psychopathy [1524]. Psychophysical [1218]. Psychotechnik [745]. Ptolemaic [2565]. Ptolemaio [723]. ptolemäische [2422]. Ptolémée [178]. Ptolemy [1986, 869]. Puaca [2139]. Public [2550, 2617, 617, 1909, 2373, 717, 1104, 638, 1968, 250, 1822, 335, 1348, 1935, 1729, 2741, 80, 2109, 959, 466, 2521, 3033, 3288]. Publication [1350, 3081]. Publications [1554, 24]. públicos [859]. Publics [2212]. publiques [505]. Publishing [884, 2028, 311]. Pugh [2736]. Pulcinella [2355]. Pulmonary [1340, 3053]. Pulpit [554]. Pulse [2038, 2496]. Pump [2076]. Pumped [2081]. Punctuated [1910]. Pungetti [1038]. Pupils [2]. Pure [101, 768, 766, 2553, 1950]. Pures [158]. Purpose [2838]. Pursued [1235]. Pursuing [578]. Pursuit [2789, 394, 425]. Pursuits [2006]. Putting [1451]. Puzzle [1643, 2649, 2974]. Pyrotechnic [437]. Pythagoras [176, 1043, 1044, 268]. Pythagoreans [1044].

Q [2434]. Qaeda [3158, 1626]. Qian [1850]. Qing [2603]. Qiqi [736]. Quadratura [1197]. Quaestiones [2497, 2103]. Qualifying [2158, 2157]. Quammen [2960]. Quand [2878]. Quantification [689, 252]. Quantifying [459]. Quantitative [2156, 2546, 61]. Quantum [3200, 2742, 1511, 450, 883, 1081, 2731, 56, 945, 2132, 3011, 1517, 2006, 1741, 1899, 1428, 1913, 3114, 2749, 439]. Quarantine [1717, 3270]. Quarks [2089]. Quaternary [353]. Quattrami [1641]. Quattrocento [1197]. que [669]. Queen [1368, 2973]. Quelle [1683, 3215, 2014]. Quellen [2699, 288]. Quellen- [2699]. querelle [171, 2014, 2072]. Quesiti [2427]. Quest [2742, 1897, 520, 73, 2340, 126, 1141, 496, 839, 1828, 335, 2145, 2655, 2219, 1428, 328, 3114, 669]. Qu’est-ce [669]. Question [356, 1656, 1178, 669, 3061]. Questioned [2708]. Questiones [642, 3038, 1344, 2258]. Questioning [3155, 1539]. Questions [1851, 10, 1595, 3101, 757, 365]. Quickenden [1211]. qu’il [2008]. Quinlan [1345]. Quinlan-McGrath [1345]. Quintero [1146]. Quintessence [2089]. quotidien [478].

R [1598, 186, 1182, 1649, 2490, 2947, 2339, 2660, 1341, 1115, 2926, 1215, 1119, 2322, 2223, 639, 1241, 1801, 1602, 1781, 2066, 1545, 1366, 1974, 2661, 2917, 1330, 1297, 82, 2055, 525, 2416, 1432, 1122, 1961, 1963, 2010, 951]. Rabault [1575]. Rabault-Feuerhahn [1575]. Raber [1638]. Rabies [807]. Rabouin [2321]. Raboy [2365]. Raby [2951]. Race [305, 697, 391, 2752, 754, 1390, 1318, 2127, 1896, 572, 1391, 2526, 72, 1714, 881, 889, 2862, 1725, 1985, 1752, 253, 1823, 2804, 1358, 1224, 2872, 2012, 996, 3233, 3040, 3266, 3300, 380, 1896, 2285, 2065, 568, 224, 1265, 291, 456]. Races [750]. Rached [1238]. Rachel [1037]. Racial [1389, 2710, 1388, 2375, 2285]. Racism [750, 2202, 291]. Radder [1086]. Radelet [274]. Radelet-de [274]. Rader [1829, 1403, 928]. Radiance [3069, 1587]. Radiation [2205]. Radical [2960, 2117, 2746]. radicali [788]. Radio [1155, 451, 1372, 3128]. Radioactive [322, 950]. Radioactivity [814]. Radioisotopes [1518]. radiology [3186, 3178, 3222, 1736]. Radium [214, 3174, 1376, 3015, 1902]. Raedschelders [2008]. RAF [217, 2730, 2284]. Rafał [1865]. Raffaele [2004, 2427]. Raffaella [1367, 1643, 2719]. Rafinesque [1213]. Ragep [2503, 1793, 2565]. Rahmen [2190]. Rahn [1815]. Railroad [2583]. Railways [886]. Raimondo [1463]. Rain [1581, 2537, 3167]. Rainbow [2489]. Rainer [2591]. Rainger [2242]. raison [1798]. Rajan [1159]. Ralegh [1279]. Ralph [1005, 1362, 2095, 1137, 2871, 2197, 166]. Ramakers [1796, 2021]. Ramanujan [2300]. Ramón [2110, 2931]. Ramsden [1833]. Rana [2710]. Randall [2487, 2493]. Randomness [1899]. Rangeland [2813]. Raniero [1802]. Rankin [2459]. Ransom [2879]. Raphael [2511, 644]. Raphaële [2931]. Rappaport [246]. Rareza [2567]. Rasche [53]. Rasmussen [1832]. Ratcliff [2350]. Rational [107, 2138]. Rationalist [2006]. rationalité [1551, 3124]. Rationality [71, 1619, 1617, 2390, 405, 3051, 1605]. Rats [532]. Rawcliffe [1469]. Ray [1416, 1884, 1817, 2982]. Raymond [1954, 1190, 2019, 2847, 1140]. Raynaud [2564]. Rayner [1328]. Rayner-Canham [1328]. Rays [2446]. Rayward [1731]. Raz [2264]. Re [1068]. Re-inventing [1068]. Reactions [2463]. Read [363, 1849]. Reader [1090, 920, 357]. Readers [1718, 1364, 3014, 3218, 3287]. Reading [780, 1848, 1283, 2683, 2898, 2566, 1181, 2709, 2511, 1199, 2857, 559]. Ready [7]. Real [2742, 2051, 2194, 1677, 136, 3229]. Realism [2213]. Reality [1649, 2117, 393, 1375, 2975, 3131, 2731, 851]. Réanimation [2938]. Reappraisal [2438]. Reason [436, 604, 400, 2907, 1605, 2221, 2307, 524, 3051, 711]. Reasoned [2998, 1550, 3026, 1582]. Reasoning [730, 2867]. Reasons [2408, 1312]. Reassessed [1986]. Reassessing [2186]. Reassessment [2232, 2787]. Réaumur [1565, 3170]. Rebecca [1605, 2174, 2108, 2892, 1376, 2293, 1378, 2219, 2732]. Rebecka [1591]. Rebekah [2349]. Rebel [2372]. Rebuilding [1913]. rebus [39]. Recalcitrant [1746, 3297]. Recasting [1226]. Reception [2180, 1267, 643, 2985, 3130, 3211, 1405, 2428]. Receptor [300]. Recherches [2179]. Recipes [188, 851]. Reckoning [2421, 2573, 980]. Recoding [1148]. Recollections [825]. Recombinant [1914, 466]. Reconceptualizing [557]. Reconciling [681]. Reconfiguring [547]. Reconnaissance [1072]. Reconsidering [415, 439, 619]. Reconstructing [2196]. Reconstructions [1040]. Record [2949, 1519, 2491]. Recorde [1478]. Recording [2835, 1504, 2876]. Records [2695, 647]. Recounted [2912]. Recreating [3028, 1040, 1600]. Recreation [2549]. Red [2752, 395, 667, 2224, 1094, 752]. Redefined [2812]. Redefining [2462]. rédemptrices [89]. Redivivus} [1676]. Redman [2202]. Reed [2120, 1581]. réel [786]. Reengaging [533]. Rees [2055]. Reeves [1699]. Refashioning [238]. Reference [334, 2078]. Refik [2627]. Reflection [1060, 368]. Reflections [1460, 974, 2750, 1996, 6, 2656, 1847, 2175, 334, 2685, 2060, 2908, 163, 1925, 1867, 164, 693, 1108, 3194]. Reflex [2474]. Reflexes [1937, 2245]. réflexion [2179]. Reflexive [2398]. Reform [1135, 967, 2846, 1056, 898, 2872, 2821, 647, 2034]. Reformation [2508, 1559, 3157, 491]. réforme [478, 371]. Reformulating [2806]. Refracted [2636]. Reframed [2325]. Reframing [138]. Regards [2170]. Régent [122]. Regime [2641, 2236, 3136, 1596, 991]. Regimen [2098]. Regimens [932]. Regimental [1567, 3010]. régimes [36]. Regina [272, 2368]. Regionalizing [581]. Regis [2129]. règle [1798]. Regulated [285]. Regulating [622]. Regulation [1625, 1664, 3201, 3156]. Rehabilitation [744]. Reibungsfläche [1212]. Reich [1895, 2892, 739, 1302, 2043, 740, 2285, 591]. Reicheneder [2449]. Reid [2707, 1484, 41, 1202]. Reidy [1577]. Reinarz [1401, 1680]. Reindeer [1232]. Reinhardt [1783, 1188]. Reinhold [2348]. reinterpretation [3132, 1407]. Reinvent [2397]. Reinventing [685]. Reinvention [1615]. Reisen [2337]. Rejoinder [1180]. Rekontextualisierung [2422]. Related [1672]. Relations [1697, 1269, 1614, 465, 619, 3226, 1146]. Relationship [2598, 1384]. Relativism [2765]. Relativity [1615, 307, 2540, 1822, 2734, 2947, 2720, 2049, 2734, 183]. Relazioni [2353]. Reliable [692]. Relics [2414]. Relies [1391]. Religion [391, 1227, 2530, 3043, 2188, 1535, 2519, 808, 22, 2338, 2501, 1696, 2508, 2358, 2248, 1539, 1706, 2661, 1924, 180, 2733, 1587, 728, 2872, 57, 1340, 2849, 693, 631, 3238, 3053, 3142, 3148, 3069, 1203, 3272, 3172, 1463, 1934, 1107, 998, 889, 260, 424]. religione [1578, 727, 3172, 931]. Religions [1927, 263, 2092]. religiosa [1479]. Religious [1492, 62, 2638, 1831, 1559, 1702, 3243]. Relocating [2603]. Reluctant [702]. Remains [1073, 2414]. Remaking [317, 603]. Remarkable [1677, 746, 3229]. Remedy [672]. Remember [1981]. Remembrance [2424]. Rémi [2282]. Remmert [2490, 2339]. Ren [2902, 2233]. Renaissance [783, 494, 646, 2631, 1346, 374, 1470, 553, 45, 1790, 367, 1638, 2011, 268, 2093, 1946, 368, 2927, 776, 2633, 1120, 2260, 1045, 366, 2566, 644, 1125, 492, 1277, 37, 1929, 2413, 3017, 1248, 3239, 3250, 2259, 2630, 3210, 3086, 1167, 3099, 1383, 1553, 2330, 1127, 2426, 1124, 2927, 1816, 1126, 2107, 1639, 1691, 1345, 869, 1279, 987, 1278]. renaissant [2332]. Renate [1592]. Renato [1435]. Renaud [2321]. Rencontres [1553].

Rendering [152, 244]. Renée [2511]. Renn [1511, 736, 2629, 1907, 2734]. renovación [1479]. Rens [1932, 1682]. Reordering [2669]. Reply [1259, 1257, 1777, 2379, 1179, 2243]. Report [2075]. representación [1939]. Representation [3236, 1640, 2625, 2974, 3097, 3141, 1685, 1643, 231]. Representing [1804]. Repression [2487]. Reproducing [775]. Reproduction [1588, 2019, 1354, 3266, 1156, 3027, 3048, 3163, 1548, 1714, 2452]. Republic [1548, 2780, 1034, 230, 3163, 2677, 2189, 2021, 668, 1287]. Republican [1743, 3289]. Republiek [2785]. Republik [745]. République [1882, 442]. Reputations [58]. Rereading [2082, 2392, 1519]. rerum [2100]. Res [23]. Rescue [2774]. Research [2156, 2870, 899, 214, 2363, 522, 101, 142, 823, 1960, 1835, 1152, 756, 2953, 2218, 2395, 314, 1904, 2547, 1663, 315, 1435, 843, 607, 452, 845, 812, 1561, 1452, 1668, 2284, 2520, 217, 1525, 2543, 580, 3199, 3304, 3116, 3298, 1592, 2220, 1440, 2515, 1963, 897]. Researching [2147]. réseaux [2295]. Reshaped [1843, 729]. Resistance [604]. Resonance [711]. Resource [1032]. Resources [2614]. Response [2157, 2086, 2754, 2313]. Responses [1973, 1491]. Responsibility [80]. Responsible [326, 2456]. Ressourcen [2053]. Ressourcenmobilisierung [2454]. Restivo [1094]. Restless [2094]. Restoration [435, 2936, 1482, 1411, 3068]. Restored [413]. Restraints [919]. Restrepo [1752]. Resurgent [807]. resurrection [3053, 1340]. Retardation [2604]. Retha [1958]. Rethinking [2839, 2374]. Retrospective [915]. Reuleaux [2035]. Reunited [471]. Reuter [2456]. Rev [1423]. Rev. [3005]. Revealed [1790]. Revealing [2626]. Revelations [20]. Reverend [1055]. Review [2045, 78, 1308, 2281, 2620, 1598, 878, 1905, 2586, 2693, 2116, 2512, 757, 2879, 1585, 1228, 2956, 186, 597, 1114, 1154, 1515, 2573, 719, 1824, 79, 1424, 2835, 1940, 215, 1182, 941, 1052, 1486, 2466, 305, 71, 737, 1311, 1588, 2665, 297, 482, 795, 1031, 1263, 1964, 431, 1880, 1694, 1649, 2128, 1514, 2054, 521, 716, 1295, 1762, 1761, 2303, 2799, 2859, 570, 1502, 2408, 1877, 443, 1541, 1590, 2490, 1932, 2881, 266, 1296, 2249, 2527, 2839, 209, 1661, 2567, 2795, 2361, 1895, 805, 348, 2030, 2297, 38, 1744, 1653, 2058, 1507, 2947, 645, 2787, 2339, 2550]. Review [1242, 783, 2894, 1425, 1890, 1837, 1307, 89, 386, 665, 2275, 2660, 2789, 1185, 2802, 285, 435, 316, 992, 282, 1271, 2742, 196, 668, 1036, 2440, 2728, 2870, 2960, 2744, 1718, 1684, 749, 1903, 497, 1367, 2735, 262, 868, 1047, 885, 1726, 2486, 961, 2491, 494, 2510, 1804, 294, 574, 1419, 2199, 2578, 391, 735, 1569, 2172, 2643, 2848, 213, 1869, 202, 646, 792, 1272, 1650, 2289, 441, 2191, 1069, 2104, 2424, 550, 2329, 2562, 947, 296, 1227, 2226, 2454, 1344, 326, 1708, 950, 1499, 1218, 1554, 1942, 899, 1436, 2631, 2210, 257, 2499, 861, 1798, 2852]. Review [2110, 2710, 2439, 1755, 1897, 2309, 2700, 1562, 2598, 2895, 1697, 273, 690, 1400, 356, 594, 809, 1309, 1568, 2174, 2814, 298, 753, 2119, 780, 796, 153, 635, 2533, 1176, 422, 1115, 1659, 307, 1594, 1975, 362, 687, 2712, 1813, 2320, 214, 1984, 980, 1006, 436, 2790, 1511, 1900, 130, 423, 511, 593, 1533, 1981, 2815, 1866, 2193, 1730, 2001, 1225, 1634, 2529, 653, 1198, 498, 1641, 2776, 442, 227, 1441, 2262, 1910, 137, 286, 1360, 2032, 2299, 112, 2640, 2447, 733, 1155, 1193, 2356, 2689, 231, 545, 1135, 1001, 1177, 2783, 2858, 2838, 1333, 2622, 2571, 2926, 254]. Review [229, 730, 1481, 24, 1492, 920, 588, 1159, 1402, 1011, 681, 821, 52, 1139, 1586, 1110, 669, 2182, 2523, 816, 1460, 2064, 2360, 1067, 754, 982, 2047, 2595, 118, 377, 671, 2363, 954, 2810, 225, 450, 883, 1970, 1983, 1578, 1109, 850, 2112, 2514, 210, 385, 1577, 127, 129, 990, 1413, 2599, 2669, 1812, 2007, 1873, 726, 2103, 1038, 564, 1543, 1654, 1863, 2279, 2860, 1930, 207, 2585, 361, 1346, 788, 2570, 270, 656, 2530, 1222, 923, 1962, 1575, 1783, 2004, 1416, 2266, 942, 2059, 1347, 678, 1304, 2524, 608, 1184, 1698, 58, 2502]. Review [121, 862, 1334, 2813, 446, 522, 1719, 1643, 1450, 1969, 1160, 2352, 2925, 2415, 2730, 1048, 2019, 1192, 2727, 938, 1275, 1215, 1414, 394, 1864, 1282, 1070, 2198, 2580, 2650, 204, 502, 744, 880, 592, 1483, 1572, 1788, 995, 1909, 1080, 2050, 2784, 2560, 389, 205, 197, 1516, 2252, 427, 1881, 142, 2324, 403, 88, 1980, 815, 2432, 2179, 372, 2885, 2330, 2943, 576, 194, 774, 1005, 549, 1418, 2373, 565, 654, 1625, 2419, 1600, 2946, 261, 1579, 212, 577, 2143, 1609, 2841, 1566, 2412, 2127, 1949, 1004, 1003, 2368, 74, 1362, 1818, 2051, 2837, 724, 2420, 2188, 1127]. Review [92, 823, 1119, 2328, 2723, 1724, 1689, 2213, 2649, 1901, 2224, 1982, 1896, 1336, 2798, 2317, 986, 1129, 2091, 398, 743, 2557, 2592, 561, 171, 1652, 1081, 1426, 2063, 2142, 1408, 1868, 360, 2220, 1430, 321, 1819, 2588, 797, 2120, 51, 374, 803, 1130, 1470, 1014, 2056, 324, 1268, 77, 2503, 2286, 1646, 388, 1722, 1710, 1054, 553, 1535, 2519, 513, 172, 685, 1474, 2017, 2528, 717, 2020, 2945, 2463, 2322, 83, 300, 152, 1132, 1939, 1128, 1580, 1960, 2446, 3141, 123, 874, 1647, 1236, 1835, 2807, 1240, 1012, 1306, 2061, 1713, 1152, 2579, 404, 679, 2150, 1757, 203]. Review [1207, 572, 2522, 1656, 1501, 61, 677, 220, 520, 2223, 604, 518, 2461, 2346, 1356, 45, 2457, 808, 1223, 1464, 1487, 2177, 2458, 2878, 178, 1791, 641, 280, 563, 1112, 49, 1274, 1780, 568, 1146, 979, 1214, 1485, 893, 177, 22, 1715, 2966, 1624, 2591, 19, 86, 1640, 585, 114, 1068, 2254, 139, 1370, 2532, 709, 1351, 1802, 1269, 2117, 2662, 1032, 1073, 1931, 199, 727, 1339, 1072, 2024, 1427, 2144, 500, 1060, 399, 2442, 191, 2295, 2053, 512, 781, 2731, 312, 750, 1505, 206, 42, 1631, 2338, 56, 573, 659, 1396, 1286, 1459, 2194, 2362]. Review [855, 2749, 664, 639, 2323, 2203, 317, 551, 81, 2699, 2513, 806, 684, 1094, 1977, 1957, 276, 1206, 853, 1699, 2426, 1943, 2276, 2719, 271, 554, 365, 1357, 2531, 154, 1337, 138, 2040, 1017, 378, 940, 657, 872, 756, 1287, 1462, 2269, 73, 638, 508, 1790, 1498, 1123, 1461, 610, 379, 1238, 395, 2211, 222, 488, 1782, 31, 1720, 2170, 2773, 2340, 36, 2344, 2746, 2942, 1735, 1429, 352, 632, 1086, 1683, 2952, 2291, 944, 2509, 367, 2180, 2866, 67, 510, 351, 2259, 493, 183, 2052, 1839, 2884, 2801, 432, 692, 1886, 731, 1495, 575, 2953]. Review [824, 132, 259, 994, 1738, 794, 798, 1077, 1189, 2526, 1537, 2108, 2698, 2875, 1231, 1967, 595, 503, 1966, 2707, 349, 869, 813, 1547, 2134, 125, 284, 62, 2114, 72, 1548, 2042, 509, 1010, 1712, 2218, 176, 755, 1338, 1714, 2468, 1280, 2697, 2539, 308, 936, 320, 1737, 1733, 2283, 1583, 140, 1638, 1882, 2014, 2632, 1261, 601, 1136, 2459, 2811, 1968, 2594, 2832, 2846, 1438, 173, 1734, 2936, 116, 721, 122, 2094, 2272, 2873, 91, 1299, 1727, 2726, 39, 629, 1107, 1695, 1480, 1262, 825, 832, 2540, 603, 955, 1241, 2955, 2000, 126, 1055, 2077, 2501, 310]. Review [2146, 1644, 2411, 1037, 1074, 1401, 1678, 1739, 833, 314, 2271, 2577, 2663, 676, 1141, 1051, 1908, 2541, 1758, 673, 547, 1131, 2201, 1959, 605, 1563, 856, 835, 2031, 1717, 2214, 120, 865, 1200, 2672, 2892, 2867, 858, 2630, 1904, 184, 303, 249, 1158, 135, 1556, 1633, 2095, 2847, 134, 434, 634, 1209, 1478, 1682, 2003, 2638, 2709, 2853, 133, 68, 250, 523, 1145, 1329, 1822, 2740, 2138, 433, 1226, 1596, 1111, 1150, 2039, 2437, 2800, 1267, 402, 1663, 1591, 2871, 1148, 2414, 327, 2670, 1696, 1009, 2659, 2451, 2887, 2855, 2460, 2656, 871, 1194, 2141, 584]. Review [948, 315, 1953, 2652, 2300, 400, 1002, 2296, 779, 2022, 2805, 571, 1435, 1685, 2651, 2453, 2941, 1210, 1565, 1898, 2505, 2280, 1191, 2227, 541, 2443, 2863, 1642, 1945, 174, 736, 670, 2628, 2318, 2200, 2623, 2126, 2621, 1629, 1229, 1039, 2013, 811, 1049, 932, 1701, 2637, 1628, 1368, 1801, 1283, 2915, 1496, 2629, 1196, 2178, 1092, 1605, 2462, 2737, 998, 1534, 2365, 2536, 559, 2171, 946, 1477, 2696, 59, 1153, 817, 945, 1497, 1576, 598, 1581, 2351, 2132, 1373, 2038, 376, 1033, 496, 1285, 1335, 1602, 1482, 2435, 1203, 2263, 1239, 2575, 1781, 636, 739, 1013, 1237, 1085, 84]. Review [2645, 2934, 1933, 293, 2267, 1290, 457, 1273, 886, 1190, 1862, 1664, 2105, 43, 2692, 1509, 1233, 2517, 2186, 2511, 2844, 1606, 1944, 2779, 189, 1817, 1124, 712, 1626, 2932, 876, 30, 200, 2011, 2808, 2874, 1305, 1723, 1834, 1042, 2057, 1736, 44, 179, 637, 1952, 263, 2204, 2882, 1558, 2639, 854, 1035, 2175, 1053, 2494, 548, 2518, 2812, 1244, 642, 2456, 866, 2918, 1912, 1648, 1709, 1821, 587, 2261, 2929, 479, 1658, 2372, 1784, 1786, 131, 354, 714, 924, 1829, 268, 1056, 2508, 526, 2093, 2923, 319, 357, 1147, 1513, 2133, 290, 1302, 1828, 1058, 1397, 1456]. Review [188, 2720, 1374, 2864, 27, 1071, 1946, 252, 1504, 2590, 371, 1636, 2169, 607, 827, 454, 1518, 1442, 555, 1354, 1677, 2250, 606, 881, 1457, 1597, 1941, 560, 1050, 2027, 2325, 515, 452, 2369, 505, 1934, 35, 274, 784, 1197, 2427, 2049, 2671, 2497, 2464, 2893, 2504, 1632, 1062, 2951, 1833, 889, 2452, 383, 2066, 2060, 2868, 2886, 2958, 2739, 2343, 2833, 814, 2967, 2264, 2381, 182, 2417, 1681, 2101, 193, 562, 2584, 1545, 1517, 2006, 489, 926, 1552, 2115, 1467, 1662, 928, 115, 1349, 1512, 2781, 1332, 688, 486, 2618, 265, 275, 393, 2005, 1950, 2102, 223]. Review [2450, 1366, 2358, 2948, 1348, 2190, 264, 1756, 437, 725, 864, 1836, 2383, 2248, 2931, 2786, 1065, 2862, 501, 2597, 748, 1363, 1539, 2715, 384, 916, 76, 289, 506, 1420, 2167, 1751, 2208, 438, 1208, 2100, 2139, 449, 459, 1016, 1887, 1407, 2429, 1063, 1217, 1376, 2131, 2145, 487, 989, 1488, 3081, 2930, 1506, 2646, 2185, 96, 2729, 458, 1440, 461, 710, 26, 491, 1399, 1411, 1876, 2355, 2334, 1803, 1284, 224, 1138, 600, 1555, 1885, 1807, 2433, 368, 2487, 2410, 2257, 66, 782, 1716, 900, 2357, 277, 1753, 2122, 2206, 2444, 1893, 279, 1973, 48, 499, 2043]. Review [2721, 2949, 2495, 2780, 918, 1061, 2817, 609, 931, 198, 557, 884, 658, 2928, 313, 602, 1935, 2225, 1660, 141, 1792, 2500, 1355, 1405, 309, 1599, 1746, 1706, 1690, 1721, 1974, 2661, 1731, 745, 917, 1234, 2096, 2877, 2449, 2025, 826, 1403, 1729, 2197, 2778, 47, 1057, 1704, 2118, 2137, 2445, 2564, 1978, 21, 507, 143, 1610, 2559, 1785, 504, 1422, 2274, 2869, 683, 2293, 211, 890, 2861, 370, 1090, 849, 292, 902, 1265, 2809, 228, 896, 2582, 2655, 882, 477, 1789, 2350, 674, 1000, 1879, 1750, 1693, 1546, 2636, 2574, 1892, 2294, 2917, 630, 566]. Review [1814, 1303, 2217, 2111, 599, 2026, 2690, 1645, 1810, 287, 1034, 2515, 190, 302, 447, 859, 1455, 1573, 2583, 1046, 2015, 586, 1087, 1044, 2425, 147, 1567, 1350, 97, 405, 2099, 440, 1692, 2216, 943, 812, 660, 2187, 1082, 1832, 1221, 1484, 2496, 2927, 87, 390, 892, 1500, 2919, 2954, 772, 1142, 180, 984, 1889, 2349, 2489, 2876, 256, 1838, 2136, 53, 401, 1300, 2498, 1076, 1330, 773, 935, 1201, 1406, 1279, 2033, 2036, 1799, 2624, 1794, 2332, 2641, 581, 2221, 1806, 221, 301, 1264, 2282, 2938, 1951, 2791, 1820, 1297, 283, 558, 991, 1816, 1550, 2034]. Review [82, 934, 993, 2270, 2924, 875, 1490, 1364, 2937, 870, 3093, 1126, 2023, 857, 1187, 232, 2301, 2430, 2277, 325, 662, 2854, 933, 1894, 304, 2302, 807, 1079, 2572, 1276, 483, 1043, 1561, 1289, 2184, 978, 2596, 818, 1113, 962, 2507, 2635, 2538, 583, 2525, 2113, 234, 742, 2176, 155, 1118, 734, 2957, 1831, 2648, 1911, 640, 1725, 2843, 2121, 1091, 929, 2836, 128, 689, 2041, 802, 1294, 1423, 2319, 2371, 919, 1519, 2702, 800, 32, 887, 1204, 1745, 216, 542, 2516, 2695, 922, 956, 877, 40, 1559, 1805, 2298, 460, 851, 2202, 590, 267]. Review [2423, 2914, 1557, 2335, 852, 2062, 1607, 2327, 1884, 2667, 175, 192, 785, 2666, 672, 898, 2065, 666, 1298, 2345, 776, 1971, 93, 2173, 2090, 1747, 1902, 951, 475, 789, 1205, 1707, 2633, 2506, 997, 544, 1144, 1651, 888, 2002, 930, 829, 1741, 901, 897, 787, 2569, 358, 546, 1064, 1811, 694, 1140, 1728, 2037, 2644, 54, 579, 957, 181, 516, 1601, 1985, 2725, 278, 711, 1066, 2657, 2939, 350, 2701, 1551, 363, 1571, 272, 2222, 552, 429, 2331, 519, 1314, 1116, 2705, 1883, 2333, 582, 1752, 751, 2130, 1230, 448, 960, 1083, 1378, 1431, 591]. Review [2125, 2738, 1536, 2788, 867, 1410, 1702, 2107, 1907, 2794, 144, 2106, 879, 295, 1592, 667, 2544, 715, 2321, 2123, 1433, 2831, 2448, 136, 2029, 453, 1219, 2589, 382, 1059, 2733, 1288, 680, 543, 1120, 987, 1199, 1639, 2046, 1491, 1787, 937, 1458, 1937, 1473, 253, 2602, 2055, 1627, 732, 1542, 34, 2367, 439, 1216, 1899, 1301, 1549, 1793, 1622, 1183, 830, 1830, 2668, 860, 149, 2568, 1914, 2009, 525, 1809, 2857, 740, 1584, 1151, 1380, 723, 2374, 2260, 649, 2654, 1587, 2278, 722, 1891, 1310, 1293, 1878, 741, 2865, 2284, 540, 2935, 1220]. Review [359, 2028, 90, 75, 1800, 111, 1823, 2492, 373, 2718, 2347, 269, 728, 718, 353, 150, 1439, 738, 208, 578, 1359, 1510, 2251, 1045, 63, 2370, 1178, 1471, 2416, 2219, 95, 977, 2288, 2147, 2921, 790, 1808, 2713, 2850, 652, 1328, 2745, 151, 1266, 939, 1040, 2537, 775, 2359, 633, 1143, 1137, 2581, 752, 2889, 476, 2166, 2922, 2215, 1188, 2722, 2804, 455, 2961, 2265, 2469, 369, 306, 219, 299, 23, 430, 451, 366, 891, 985, 1089, 1421, 1432, 1278, 1748, 1493, 1593, 1604, 2653, 55, 1232, 1958, 1015, 2741, 260, 655, 1676, 949, 2256, 80, 921, 2520, 255]. Review [2364, 2488, 1212, 69, 1213, 1371, 2353, 799, 1680, 1874, 758, 1157, 2097, 2467, 2627, 291, 1088, 778, 2258, 1428, 1827, 611, 1312, 288, 1379, 1508, 953, 481, 113, 145, 746, 1093, 397, 1754, 2124, 2273, 2292, 1358, 33, 364, 720, 201, 873, 675, 1415, 589, 1352, 1270, 60, 490, 786, 958, 1871, 2418, 426, 804, 729, 983, 1211, 1635, 2109, 2092, 1759, 1979, 2797, 1224, 1475, 1976, 2534, 1589, 2135, 2287, 85, 894, 1243, 2441, 2253, 2246, 819, 2212, 1796, 29, 119, 648, 1637, 2706, 1078, 2619, 2950, 1865, 1156, 2803, 2207, 2642, 2694]. Review [2196, 456, 2189, 2348, 1331, 1906, 2851, 2436, 1466, 185, 1412, 2326, 281, 1353, 2647, 1417, 1815, 663, 1375, 2732, 217, 514, 820, 1540, 2455, 1743, 2336, 1570, 2098, 148, 828, 425, 2716, 713, 1084, 1861, 251, 686, 976, 480, 1867, 445, 2842, 2566, 527, 644, 187, 485, 925, 1195, 1938, 1472, 1687, 484, 2563, 2542, 2743, 226, 691, 1947, 1740, 1956, 1494, 495, 2565, 777, 94, 517, 999, 2048, 2872, 218, 70, 1007, 1377, 2247, 2625, 959, 1503, 65, 2845, 1826, 793, 1125, 1281, 1468, 1700, 1795, 2634, 64, 1122, 1465, 1564, 2792]. Review [57, 396, 569, 1936, 492, 1479, 1538, 2008, 2493, 2012, 2856, 1340, 2664, 1398, 1691, 2587, 927, 2342, 1277, 50, 1560, 2168, 2708, 2785, 1235, 2465, 1948, 567, 1825, 428, 2891, 834, 647, 524, 2035, 2195, 195, 1292, 1369, 124, 650, 2782, 556, 651, 952, 2209, 2205, 2593, 2888, 1404, 1342, 1608, 2849, 2366, 2658, 37, 1476, 1961, 2434, 981, 2375, 392, 682, 693, 1186, 1679, 1574, 2192, 41, 2422, 2089, 1688, 1345, 1553, 2181, 996, 1595, 822, 1955, 2290, 2796, 2018, 631, 322, 1149, 2140, 258, 230, 2816, 2711, 810, 1108, 1365, 1489]. Review [2021, 1872, 1469, 2880, 1603, 2576, 895, 1434, 2044, 2704, 1582, 1544, 1121, 380, 831, 1041, 1711, 2129, 2691, 2736, 1133, 2940, 318, 444, 1181, 1202, 1870, 1797, 2543, 2724, 1361, 1291, 1749, 1117, 596, 28, 1372, 1972, 1657, 1686, 375, 20, 387, 747, 1343, 1630, 2409, 2407, 1075, 1913, 2341, 2626, 323, 146, 406, 1705, 355, 643, 2775, 2285, 3129, 25, 2521, 2561, 1963, 98, 328, 863, 1965, 381, 801, 1888, 2010, 2438, 791, 988, 2916, 661, 580, 1742, 1463, 1134, 2268, 233, 1437, 2714, 311, 2413, 46, 1409, 2428, 2535, 2793, 2558, 478]. Review [2421, 2431, 1655, 1623, 2337, 2016, 2354, 1703, 1875, 117, 1313, 424, 1732, 1008, 2183]. Reviewing [2077, 2614]. Reviews [2076, 1341, 1954, 2717, 2544, 2734, 2703, 2777, 2806]. Revision [2025]. Revisited [3200, 2080, 3236, 1685, 1741]. Revisiting [1844, 2903]. Revival [1152, 516]. Reviving [2059]. Revolución [1479]. Revolution [653, 229, 938, 338, 2224, 874, 1999, 1166, 1558, 2216, 1314, 219, 799, 3115, 1560, 233, 1882, 3031, 3041, 3303, 3224, 2014, 371, 1556, 1804, 2669, 2318, 1808, 651, 307, 436, 989, 557]. Revolutionaries [2917]. Revolutionary [886, 2429, 799, 1828, 592]. Revolutionized [2731, 1576, 3144]. Revolutionizing [1829]. Revolutions [2175, 1488, 2922, 1092]. revoluzioni [493]. Revue [158]. Rey [1683]. Reymond [1580, 3055]. Reynolds [2413]. Rezeption [2280, 1373, 3130]. Rezvykh [1373]. Rhône [603]. Rheinberger [1435, 2738, 593, 579]. Rheinfels [2795]. Rheinprovinz [208]. Rheinsberg [1132]. Rhetoric [88, 1512, 1831, 1375, 3131, 200]. Rhetoric/Communication [88]. Rhetorical [1780, 1931]. Rhinoceros [2433]. Rhoda [246]. Rhodes [629]. Rhys [2966, 2832]. Ricardo [1752]. Ricci [2947, 2603]. Riccioli [1943, 2850]. Rice [2510, 1689, 1140]. ricerca [51]. ricezione [2985, 1338]. Rich [951]. Richard [1903, 1634, 1575, 2063, 1068, 1339, 2731, 2931, 1423, 2097, 2849, 1453, 838, 2723, 2017, 2579, 2146, 1563, 2844, 2381, 21, 2349, 1707, 2130, 2842, 2793, 449, 732]. Richards [2895, 2580, 2272, 2175, 1595]. Richardson [1601]. Richter [1215]. Rickles [1755]. Riddle [1458]. Ridgway [1003]. Riding [514]. Ridler [2137]. Riedler [2035]. Riehm [1305]. Riemann [2720]. Ries [1036]. Rifle [1995]. Right [2912, 2456]. Righting [2745]. Rights [2119, 1728, 2593, 2976, 3255]. Rigor [456]. Rik [2008]. Riki [1392]. Rima [1031]. Rina [1788]. Rinaldo [2927]. Rindzeviciute [2541]. Rings [147]. Rinoceronte [427]. Rio [2567, 2867]. Riordan [2144]. Rise [348, 2550, 816, 212, 2723, 2144, 659, 91, 1131, 2740, 402, 948, 1581, 2325, 515, 900, 2122, 1832, 2036, 933, 1298, 1593, 37, 1343, 2979, 3167, 3121, 2443, 2739]. Risi [1868, 2263]. Risk [858, 2122]. Riskin [2094]. Risse [1070]. Ristretto” [1802]. Ritter [873, 873]. Ritual [3226, 1697]. Ritvo [1363]. Ritz [1499]. Rival [960]. Rivka [2503]. rivoluzione [3041, 1652, 2100, 1490, 3115, 1630, 790]. Road [2075, 1985, 322]. Roazen [1043]. Rob [2717, 1375, 2638]. Robert [1341, 273, 1788, 1600, 1003, 2446, 2966, 503, 2955, 1581, 2294, 1455, 1811, 1800, 1212, 387, 2354, 1052, 2839, 1341, 847, 1896, 1819, 1647, 3206, 176, 2272, 1478, 2038, 2175, 1513, 2133, 2250, 1062, 2167, 1814, 2217, 3093, 3167, 2506, 1431, 2364, 1861, 1494, 495, 1503, 2008, 1948, 411, 1595, 1797]. Roberto [2540]. Roberts [2859, 1346, 2127, 1474, 1478, 1377, 64, 2431]. Robertson [2498, 1088]. Robilliard [2704]. Robin [1140, 2410]. Robinson [2956, 2361, 2049, 287]. Robot [832]. Robots [2923, 2452, 1978, 2010]. Robson [32]. Roby [2329]. Rocca [2781]. Rochberg [2418]. Rocke [1500]. Rockefeller [1427, 3106]. Rocket [395, 2739]. Rocketeers [2593]. Rockets [1078]. Rocks [1115]. Rod [258]. Rodent [532]. Roderick [517, 2643]. Rodger [2128, 1681]. Rodolphe [2697]. Roel [2784]. Rogacheva [2815]. Roger [123, 1137, 1705, 1905, 1880, 1492, 1128, 1286, 1642, 3263, 2954, 2270, 2736]. Rogers [569]. Roi [1485, 2425, 1485]. rol [440]. Roland [2529]. Role [1460, 1790, 315, 670, 1677, 292, 738, 3229]. Roles [2408]. Rolf [2029, 1310]. Rolf-Jürgen [1310]. Rolfe [2934]. Roling [1472]. Rolle [503, 438, 28]. Rom [1562, 2977]. Roma [2100, 1477, 2276]. romain [3211, 1405]. Roman [2329, 2841, 1207, 709, 1351, 2178, 2494, 2169, 2381, 1555, 1635, 2842, 1938, 1322, 3084, 3085, 3161, 2780, 2898, 2677, 1742]. Romance [284, 1364, 3014]. Romanes [1724, 512, 3280]. Romano [2331, 2342]. Romans [927]. Romantic [2466, 1533, 275, 21, 63, 1143, 3022, 1494]. Romanticism [2466, 433, 2516, 63, 873, 2644, 2352, 1226, 201]. Romanus [1405, 3211]. Rome [875, 1536, 728, 2306, 187, 3193, 3238, 2977, 1536, 2329, 2928, 1706, 374]. Rommevaux [1126, 34]. Ronald [1784, 2242, 2223, 2114]. Room [2390, 145, 693, 145]. Rooms [2202, 3246, 1732]. rooms-katholieken [3246, 1732]. Roots [372, 549, 2457, 1844, 2951, 1687, 3279]. Roqué [1745]. Rosa [1643, 1126]. Rosalind [1371]. Röschner [2657]. Rose [1072, 1238]. Rosella [2298]. Rosemblatt [2872]. Rosen [1039]. Rosenfeld [999]. Rosenkranz [1507, 2885]. Roses [1572, 2983]. Roslynn [2618]. Rossby [2128]. Rosse [1715, 3276]. Rossiter [896]. Roth [22]. Rothman [2789].

Rothmann [1698, 3252]. rotis [928]. Roudet [992, 2417]. Roudinesco [2874]. Rouge [909]. Round [1669, 2732]. Rousseau [433, 1057]. Route [2233]. Routes [1598, 1435, 1488, 3208, 3199, 3204, 179, 1825]. Routledge [1480, 1591, 2495]. Roux [2441]. Rover [1978]. Rovers [1150]. Rowlinson [1655]. Roy [1754, 369]. Royal [1414, 74, 2150, 3207, 349, 2468, 2936, 1648, 1709, 1707, 3230, 3082, 3220]. Rubber [134]. Rubbing [2073]. Ruben [1333]. Rubinson [2462]. Rublack [2111]. Rüdiger [2454, 1129]. Rudolf [2924]. Rudolph [583, 1657]. Rudwick [2196]. Rule [1857, 637, 193]. Rumor [2712]. Runaway [2740]. Runge [586]. Rupert [13]. Rupescissa [189]. ruptures [2023]. Ruse [1193, 2838, 1334, 31, 2272, 2358]. Rush [567]. Ruska [2887]. Rusnock [2029]. Russel [1370]. Russell [813, 1751, 1879, 1327, 1223, 1055, 2022]. Russia [132, 886, 1331, 1963, 3059, 3073, 1333, 2136, 880]. Russian [1585, 209, 994, 1335, 1602, 637, 2125, 2240, 2986, 621, 1923, 3153, 3119]. Russians [2026]. Ruth [1906, 2942, 2102]. Rutherford [228]. Ruthless [1655, 3019]. Ruys [1790]. Ryan [2735, 1200, 1219, 2375, 2659, 200, 2026].

S [1492, 1577, 88, 2211, 2746, 2114, 1822, 2656, 1092, 1781, 2372, 1071, 182, 1753, 1690, 2506, 1601, 2589, 1937, 2722, 1213, 1827, 397, 2196, 2563, 1560, 2880, 1686, 1963, 2793, 2179, 2237, 274]. S. [2121, 1655]. sabbat [645]. saberes [1877]. Sabin [1757]. Sabina [2670]. Sabine [2007, 34, 1646, 1126, 1475, 2018]. Sabrina [2570]. Sacha [1546]. Sachen [586]. Sachiko [1046]. Sachs [2456]. Sachse [1233, 1972]. sacra [1472]. Sacred [1157, 2563, 1038]. sacro [1130]. sacro-santa [1130]. Safe [1240]. Safety [1240]. Safia [593]. Saga [50]. Sagamore [1554]. Sagredo [1813]. Sahagún [1201]. Sahlins [2794]. Saiber [2631]. Saint [274, 2775]. Saint-Vincent [274]. Saints [1418, 3047, 1608, 25, 3024]. Sakharov [400]. Sakurai [1365]. Sal [1094]. Saladyga [1119]. Salguero [2840]. Salisbury [1121]. Salk [1984, 1838]. Sally [1139, 1160, 1211, 2565]. Salmonella [1968]. salubrité [505]. Salvo [1471]. Sam [2451]. Samantha [1059]. Samir [2261]. Sämtliche [125]. Samuel [1205, 2772, 1352, 1705, 2956, 3263, 1448, 3092, 2202]. Samuel-Auguste [1352, 3092]. San [1070, 1197]. Sánchez [1939]. Sancho [1082]. Sanctity [2788]. Sandars [3218]. Sandra [1184, 1090]. Sanitary [1717, 3270, 1135]. Sanskrit [1575, 3166]. Santa [514, 1130]. santé [36]. Santer [2000]. Santesmases [1836]. Santiago [1350]. Santo [2636]. Santos [2630, 1752]. sapere [931]. saperi [931]. Sappol [2533]. Sara [2796, 1463]. Sarah [2808, 2639, 2544, 2193, 2926, 1139, 2198, 2220, 1240, 1633, 2039, 1951, 1601, 2647]. Saraiva [2534]. Sarasohn [1801]. Sarton [509, 2902, 1024]. Sasquatch [681]. Satellite [981, 2464, 50]. Satire [2472]. Saturn [935]. Sauer [1507]. Saul [2117]. Saulo [2438]. Sausman [1137]. Savagery [737, 2074]. Savannah [569]. Savant [1062, 297, 2414]. Savants [2431, 3092, 2282, 1352]. Save [2767]. saviors [1289]. savoir [2261, 2332]. savoirs [3250, 2569, 1243, 1553]. Saw [1808]. Sawday [37]. Sayer [2801]. Sayre [2813]. Scale [2813]. scales [3036]. Scaling [1254]. Scandalous [2846]. Scandalously [541]. Scandinavia [2508]. Scarborough [2921]. Scarcity [1032]. Scarpellini [1461]. Scarred [506]. Scavans [2574]. Scents [1680, 3284]. Sceptical [2080]. Scerri [1509, 2779]. Schappacher [443]. Scharnhorst [1160]. Schattengefechte [1489]. Schauz [2953]. Schechner [2796]. Scheidegger [2811]. Schell [1723]. Schelling [1373, 3130, 3130, 2798]. Schemes [1027]. Schemmel [1048, 736, 2625]. Schickore [2558]. Schiemann [1086, 70]. Schilder [2571]. Schillings [2290]. Schizophrenia [2458]. Schlechte [3008, 1361]. Schlegel [2270]. Schliesser [1054, 1876]. Schmaltz [2454]. Schmalzer [2224, 2746]. Schmaus [2940]. Schmid [2140]. Schmidgen [593, 1980, 1820]. Schmidt [2056, 1504, 1885, 2445, 2337]. Schmiedel [2343]. Schmitt [1485, 369]. Schneider [2799, 2490, 2002]. Schneider-Kempf [2799]. Schofield [847]. Scholar [2586]. Scholarly [1600, 2317, 1917, 1167, 630, 1830, 45]. Scholars [1036, 2032, 2630, 2855, 2856, 2711]. Scholarship [112, 73, 1548, 1923, 277, 1503, 3163, 391]. Scholastic [1199]. Scholastics [1128]. Scholten [255]. Schönhofer [1878]. School [2881, 2571, 2373, 2871, 47, 1536, 1742, 3240]. schools [3219, 1366]. Schott [1820]. Schreiben [635]. Schrieck [2711]. Schriften [125]. Schriftsinn [1472]. Schrödinger [1899, 664]. Schuh [2699]. Schulten [1147]. Schultz [2812]. Schumpeter [298]. Schüring [2226]. Schuster [2512, 1798, 1306]. Schütte [2504]. Schwartz [2531, 1226, 1118]. Schwarz [1684, 1132, 1401]. Schwarzbach [1713]. Schweber [1071, 2772]. Schweighöfer. [2944]. Schweitzer [2450]. Schweiz [2811, 1230]. Schwerin [1402]. Schwermer [443]. Science [2116, 1523, 1585, 2956, 911, 467, 765, 1025, 71, 697, 1031, 1026, 2128, 1, 169, 333, 417, 616, 704, 1164, 1028, 1446, 1763, 1532, 2070, 1859, 2384, 2165, 2482, 2771, 2770, 973, 2075, 1242, 665, 2275, 1185, 2617, 282, 1271, 1718, 2735, 1047, 1569, 2643, 213, 2160, 2191, 1247, 550, 2329, 1227, 2920, 1554, 2895, 690, 1529, 2814, 161, 687, 1533, 2529, 1155, 1927, 1001, 767, 681, 1586, 2612, 377, 2162, 2112, 761, 1577, 1222, 2684, 974, 1390, 1334, 522, 101, 1215, 1282, 2650, 502, 995, 1173, 2252, 774, 1418, 2680, 2826, 1530, 2127]. Science [838, 1362, 2392, 2188, 92, 239, 1901, 1982, 1430, 803, 2056, 2883, 172, 2475, 2528, 1580, 1236, 1012, 6, 572, 108, 1501, 618, 3053, 568, 1146, 22, 1068, 709, 2117, 2662, 399, 3246, 2238, 855, 639, 2203, 317, 1094, 1169, 1391, 853, 1943, 972, 1852, 2774, 2269, 73, 508, 2829, 2344, 1839, 768, 2953, 1384, 824, 132, 1738, 2235, 798, 1077, 1189, 471, 72, 176, 2468, 936, 320, 3149, 2811, 1438, 2936, 2726, 1923, 1844, 2501, 833, 2577, 1131, 1926, 1563, 2385, 2164, 1249, 303, 133, 250, 2609, 433, 1591, 2871, 1916, 327, 871, 2141, 1953, 1898]. Science [1945, 2682, 2907, 174, 2013, 2462, 998, 1773, 1581, 376, 1239, 2267, 8, 200, 843, 1448, 1834, 2204, 1558, 1035, 2175, 1821, 1786, 131, 354, 924, 1829, 357, 2133, 1828, 1058, 27, 2900, 1518, 1995, 2325, 2903, 452, 2369, 1934, 2427, 2497, 1062, 889, 383, 335, 914, 2886, 2967, 182, 913, 2476, 1552, 109, 2678, 1512, 2908, 2389, 840, 2163, 1539, 289, 766, 1208, 1922, 2159, 2930, 1440, 1854, 1411, 472, 1779, 2980, 277, 2206, 1721, 2877, 162, 47, 2137, 2445, 507, 2869, 2293, 2477, 1090, 2603, 2582, 477, 1924, 2294, 912, 2583, 405, 1921, 660, 1832]. Science [87, 1500, 2919, 2954, 915, 2876, 1300, 2033, 2641, 581, 908, 163, 991, 82, 2910, 2717, 232, 1079, 2525, 2113, 341, 155, 1770, 128, 1771, 2041, 2371, 1103, 1925, 2298, 460, 2236, 971, 1607, 2327, 1971, 1251, 930, 546, 1728, 54, 957, 2725, 2222, 1314, 1752, 1851, 1410, 1702, 144, 2679, 2589, 9, 3105, 1937, 1473, 253, 2237, 769, 2374, 2654, 2278, 2402, 1800, 2492, 728, 738, 622, 2147, 2921, 652, 1328, 151, 1040, 1137, 2889, 763, 1421, 1593, 1604, 260, 2256, 80, 619, 921, 2306, 2097, 291, 746, 698, 873, 60, 1976, 2212, 1796, 2073, 2642, 2189]. Science [2851, 3115, 663, 217, 514, 1540, 7, 1388, 2687, 976, 1867, 445, 2911, 691, 3022, 1494, 495, 3299, 2613, 2872, 218, 70, 164, 1007, 959, 1564, 57, 2008, 927, 50, 2764, 650, 105, 1961, 2434, 693, 2904, 996, 822, 1149, 2616, 1613, 2021, 2880, 1434, 2691, 2940, 444, 339, 1772, 2233, 2521, 1963, 3172, 1100, 661, 530, 311, 1623, 1990, 2016, 1703, 424, 2984, 3024, 2986, 3134, 2990, 3041, 3045, 3205, 374, 3242, 3243, 3050, 3056, 3058, 2170, 3255, 3151, 3152, 3068, 3069, 1203, 3074, 3155, 3197, 2978, 3211, 3198, 26, 3271, 3083, 3272, 1405, 3165]. science [3095, 3096, 689, 867, 3104, 3218, 3287, 3106, 1243, 3116, 3121, 3300, 3128, 3227, 1031, 1554, 1983, 1543, 22, 2040, 2746, 1086, 1537, 1231, 1733, 1908, 1953, 43, 2918, 2325, 1833, 2381, 2248, 1539, 1789, 2648, 2327, 1587, 1510, 2653, 2741, 2364, 1465, 1340, 1365, 2313, 2308, 2806, 2407, 2408, 2174, 2790, 2689, 1067, 2813, 2415, 2051, 2723, 1901, 2017, 2461, 808, 2966, 1624, 2254, 1427, 2362, 1206, 2942, 1967, 2832, 1696, 2779, 1244, 2372, 2951, 1887, 2197, 2118, 2655, 2690, 1805, 1144, 2029, 2857, 2166, 2469, 2418, 2797, 2436, 2326, 1815, 2455, 1121, 1372, 1713, 809]. Science [231, 1177, 982, 118, 385, 389, 194, 654, 360, 1014, 324, 874, 404, 604, 563, 684, 756, 352, 601, 634, 1283, 131, 714, 319, 814, 275, 1284, 902, 483, 542, 776, 901, 579, 295, 830, 649, 741, 63, 720, 804, 693, 631, 810, 895, 328, 233, 608]. Science-Language [2236]. Science-Mart [684]. Sciences [1541, 2339, 1526, 1176, 1527, 2783, 1575, 2004, 1957, 2426, 657, 1238, 349, 122, 2663, 2541, 1563, 2138, 1166, 2652, 2511, 2932, 843, 1723, 1662, 437, 1407, 1785, 221, 1816, 818, 962, 2423, 2914, 1557, 579, 1105, 680, 34, 3217, 10, 2092, 1353, 1711, 1343, 863, 791, 158, 2535, 3132, 3137, 3042, 2295, 1683, 2014, 3151, 3266, 2979, 3277, 3278, 3098, 3215, 2569, 3021, 3292, 2023, 2195, 1603, 1905, 1069, 1730, 1441, 1160, 1714, 2862, 1506, 2917, 2321, 2367, 548, 94]. Sciences.} [2271]. scientia [1476, 2183]. Scientific [757, 3228, 1515, 2156, 911, 700, 1653, 2550, 961, 2491, 1804, 441, 1309, 2533, 436, 653, 2523, 118, 1577, 1954, 58, 1969, 2650, 338, 1988, 2680, 2946, 1566, 823, 2224, 2503, 2020, 1274, 750, 2749, 276, 1337, 2040, 2269, 632, 575, 237, 2395, 173, 249, 1556, 1999, 523, 1150, 2822, 1916, 1685, 2318, 1860, 2549, 1092, 1273, 2692, 1448, 2175, 2918, 1648, 2864, 555, 2060, 393, 2163, 2139, 989, 1399, 2225, 243, 1168, 2919, 1769, 1099, 2302, 2081, 1911, 2202, 1884, 2667, 2002, 2130, 2125, 1216, 535, 525, 969, 2251, 1808, 160, 291, 1379, 397, 1754, 2803, 2310]. Scientific [2336, 2687, 1956, 1560, 651, 2796, 747, 2341, 2626, 2285, 1100, 3023, 3135, 3236, 3138, 3241, 3207, 3267, 2981, 3183, 3126, 3224, 1634, 1971, 325, 580]. Scientifically [2558]. scientifiche [2355]. scientifique [2014, 1551, 3124]. scientifiques [564]. Scientism [2130]. Scientist [326, 690, 409, 2022, 3075, 2618, 244, 738, 1815, 2310, 2842, 2201, 1644, 2102]. Scientists [1598, 521, 2361, 435, 1271, 749, 356, 2815, 1543, 974, 2726, 1229, 2462, 2779, 1442, 2393, 2154, 896, 295, 1867, 1235, 2891, 105, 822, 580, 3036, 3149, 3208, 3080, 3018, 1036, 2220, 1780, 2746, 2414, 69]. Scienza [1652, 3172, 3041, 51, 1130, 2513, 595, 2100, 790, 3115, 1630, 1578, 2306, 727]. scienze [1041]. Scienziate [890, 1184]. Scienziati [3020, 1367]. Scienziato [2276, 954]. SciTech [1847]. scoperta [1109]. Scopes [1366, 1887, 3219]. Scot [1214]. Scotland [3186, 376]. Scott [1903, 741, 675, 2296, 2692, 2204]. Scottish [1568, 123, 203, 1207, 2705, 3006]. Screen [2751]. Scriba [1393, 1049]. Scribal [2511]. Scribner [675]. Scriptural [2708]. Scripture [263]. Scritti [727]. Scull [1930]. Scurvy [2559]. Sea [2191, 1337, 310, 135, 2026, 997, 366, 2724, 3126, 1875]. Sean [1154, 1474, 2137, 1346, 2050, 1083]. Search [79, 435, 1441, 767, 225, 2328, 1306, 107, 86, 1396, 1682, 2300, 1946, 405, 1601, 1066, 921, 2694, 1687, 1133, 3232, 3042, 3065, 3279, 3102]. Searches [1510]. Searching [2536, 2139]. Sears [2605]. Seasons [647]. Seattle [1607]. Seattle/London [1607]. Sebastian [1863]. Secchi [3039, 1357, 3039]. secoli [1184]. Second [570, 316, 22, 813, 2682, 2954, 567]. Secondary [622]. Secord [1718, 1720, 1059]. Secrecy [1474, 1149]. Secret [2104, 1109, 1336, 561, 1209, 1902, 1985, 620, 3174, 3125, 2056, 282]. Secrets [1272, 1400, 2019, 553, 839, 714, 3016, 1292, 2980, 1289, 1945, 403, 774, 86]. Secular [2563]. secundum [2103]. Security [2462, 2139, 2888, 134]. Seduction [1221]. See [2360, 2912]. Seed [2767]. Seeds [2518]. Seeing [1252, 667, 2754, 1978, 1889]. Seek [1072]. Seeking [3263, 1705, 203]. Seelenorgan [2978, 1398]. Seeley [2661]. Seelig [2711]. Seen [172]. Segal [1456]. Segregation [383, 60]. segreta [1109]. segreti [2570]. Seicento [3204, 1641, 51]. Seidel [2354]. Seigneurial [1872]. Seim [1427]. sein [441, 2287]. seinen [947]. seiner [2453, 1188]. Seizing [1156]. Sela [2855]. Selbstaufzeichnung [635]. Selbstmobilisierung [739]. Selden [1875]. Selectarum [3081, 1350]. Selected [1047, 883, 1785, 1195, 1952]. Selection [1221, 516, 2580, 919]. Selections [2795, 2427]. Selectivity [456]. Seleucia [1467]. Self [1326, 1674, 1580, 2822, 2118, 1593, 3055, 3121, 1486, 1900, 606]. Self-Control [1593, 3121]. Self-Deconstruction [606]. Self-Organization [1674]. Selisker [2296]. Sellers [1133]. Selling [1903, 803]. selon [89, 2332]. Selves [1004]. Semi [1058, 3141]. semi-simples [3141]. Semiconductor [683]. semigroups [3258, 1747]. Sen [1890, 2099]. Senator [619]. Seneca [1041]. Senegal [2817]. Sensation [243, 63]. Sensations [583, 2481]. Sense [1311, 758, 2593]. Sensibility [1131]. Sensory [1643, 2974]. Sentimental [2431]. Sepkoski [1519]. Septem [1622, 3249]. September [474, 1327]. Sequence [1439, 3113]. Sequencing [1082]. Sera [2862]. Sera-Shriar [2862]. Serapide [205]. Serdeczny [2938]. Serena [2178]. Serendipity [1522]. Sergei [2211]. Seriality [575]. Serialization [838]. Series [861, 1568, 1555, 1549, 60]. Seriously [1173, 924, 2393]. Serpentis [2182]. Service [2806, 1713, 3197, 2806]. Serving [1895]. ses [1123, 1894]. Sesiano [2499]. Set [292, 304]. Sets [42]. Settecento [1109, 2324, 498]. Setting [1943, 2623]. Seuil [1540]. Sevan [1510]. Seven [2174, 22, 1509, 2779]. Seventeenth [797, 1917, 853, 1644, 2071, 2171, 2186, 200, 1951, 2572, 734, 2231, 930, 2336, 1133, 3204, 3075, 2478]. Seventeenth-Century [797, 1917, 853, 1644, 2071, 200, 2572, 734, 2231, 2336, 3075]. Seventh [2628]. Seventh-Century [2628]. Sevigny [2465]. Seville [3033, 1348]. Sewage [862]. Sex [1824, 1004, 1904, 3156, 2452, 3102, 1627, 3076, 1307, 1625, 1452, 1601]. Sexual [2735, 2580, 2197, 1221, 2658]. Sexualis [2197]. Sexualität [1301]. Sexuality [1957, 3080, 1442]. Seyler [1962]. Shackleton [741]. Shadow [1308, 129, 2352, 92, 677, 2526, 3253, 1252, 68, 1753]. Shadows [980, 2264]. Shakespeare [798, 782]. Shakespearean [1410, 3056]. Shaky [1239]. Shank [2852, 19, 1179]. Shape [1669, 2204, 459, 2564, 915, 2667]. Shaped [2432, 2213, 720, 2806]. Shapes [361]. Shaping [2116, 443, 2297, 398, 1396, 1131, 2575, 1083, 3065, 1295]. Shapiro [1366]. Shared [1448, 889]. Sharing [2297]. Sharma [1729]. Sharon [637, 1786]. Shaul [1511]. Shaw [502, 2125]. Shawn [1544]. Shayler [2464]. Sheehan [43]. Shehada [1549]. Sheldon [1311]. Shell [1072]. Shellen [2283]. Sheller [1599]. Shelley [2890, 1059, 2456]. Shelves [1025]. Shen [1743]. Sheppard [2879]. Sherlock [2948]. Shetterly [2538]. shi [223]. Shift [2948]. Shifting [2617, 3057, 1606]. Shing [1579]. Shing-Tung [1579]. Shinn [2064]. Ship [1068]. Shirilan [2167]. Shock [2227, 1833]. Shocking [1274, 410]. Short [2299, 545, 300, 1547, 1678, 541, 1893, 481, 820, 2089, 1437, 3256, 3162, 3094, 3203]. Short-Lived [481]. Shot [2732]. Should [467]. Show [2196]. Showmanship [1001]. Shriar [2862]. Shtetl [1735, 3209]. shu [230]. Shuttle [147, 1432, 3000, 2969]. Shyovitz [2424]. Sialaros [2922]. Siberian [2891, 531]. Sibling [2727]. sichern [635]. Sick [2366]. SickKids [2366]. Sickle [2752]. Sickness [1307]. Sidereal [192]. Sidereus [789, 192]. Sideways [2437]. Sidrat [2257]. Siebert [2422]. Siebold [1476]. Siècle [2038, 2239, 564, 178, 36, 2697, 779, 1632, 785, 2569, 1551, 2703, 2342, 3124, 2170, 1477, 1203, 2569, 2873, 297, 942, 171, 505, 867]. siècles [2014, 34, 1463, 1481]. Sieg [1590]. Siegfried [1801]. Sienell [2535]. Sietske [2797]. Sight [1543, 1936, 1544, 3149, 3110]. Sigmund [2655]. Sign [2519]. Significance [2769, 2485, 2661]. Signifier [2238]. Sigrid [2224, 2746]. Sik [1539]. Silent [792]. Silke [1972, 1634, 1233, 1972]. Silvan [1071, 2772]. Silverstein [185]. Silvi [1553]. Silvia [45]. Silwood [1309]. Sima [1850]. Siméon [1713, 3197]. Siméon-Denis [1713, 3197]. Simili [1367, 2719]. Simmons [2033]. Simões [2195]. Simon [2859, 1757, 1286, 2569, 1823, 1604, 1621, 1066, 3216]. Simone [45, 2383]. Simoni [2845]. Simonsen [1648]. simpático [787]. simples [3141]. Simplicius [1468]. Simulation [1374, 3012, 461]. Sin [2150]. since [356, 3201, 2557, 1152, 2468, 1664, 479, 896, 2221, 3018, 680]. Singapore [2805]. Single [1155]. singularidad [2567]. singularités [2939]. Sino [2096]. Sino-Japanese [2096]. Siobhan [1377]. Sir [2045, 2866, 1057, 1704, 742, 1549, 3189, 1484, 3019, 41, 1655]. Siraisi [1127]. Sirens [1621]. Siskin [2575]. sistema [2850]. Sisters [202]. Site [20]. Sites [1038, 2563]. Sivasundaram [1137]. Sixteenth [2427, 2786, 2928, 1046, 2231, 278, 2333]. Sixteenth- [2231]. Sixteenth-Century [2786, 2928, 1046, 278]. Skaarup [1883]. Skalen [3036, 1561]. Skeleton [467]. Skelly [686]. Skepticism [1178]. Sketches [610]. Sketching [1426, 341, 3091]. Skies [2465, 565]. Skills [981]. Skin [1414, 3082]. Skinner [228, 2619]. Skizzen [61]. Skizzieren [635]. Skordoulis [2255]. Skull [572, 1368, 2973]. Sky [1155, 598, 1013, 1372, 2990, 3128, 825]. Slack [686]. Slaton.S [456]. Slava [1909, 1909]. Slavery [1307, 871, 1488, 1387, 3078, 1573]. Slaves [1065]. Slayton [1378]. Sloan [2066]. Sloane [2715]. Slotkin [2949]. Slotten [2097]. Slowik [2186]. Smaele [2519]. Small [545, 2367]. smell [3284, 2802, 1680]. Smith [2054, 2188, 2592, 1478, 2623, 1545, 1557, 1702, 2436, 2647, 686, 1936, 2012, 1544, 2724, 1581, 3167]. Smithson [1341]. Smithsonian [1600, 1513, 2959]. Smoke [521, 2262, 2073]. Smuts [967]. Smyrek [1597]. Snait [1110]. Snake [2558]. Snedegar [2127]. Soberanía [1939]. Sobre [1873]. Sociable [1723, 3217, 2266]. Social [1905, 1515, 941, 1185, 1697, 809, 1460, 1575, 1304, 1390, 1160, 1014, 1012, 108, 2461, 1427, 206, 1017, 1663, 1239, 2918, 2886, 109, 840, 507, 221, 818, 32, 175, 680, 80, 619, 397, 2048, 2872, 1024, 1100, 3151, 3106, 3226, 949]. Socialism [259]. Socialist [2224, 969]. sociedad [1877]. Societät [2190]. Société [89, 505, 2282]. sociétés [2569, 1243]. Societies [2032, 376, 2914, 1365, 3105]. Society [704, 1028, 2771, 285, 2004, 1414, 1572, 815, 1580, 2338, 349, 2468, 2936, 2201, 303, 1834, 1302, 1354, 383, 289, 857, 155, 2173, 1707, 359, 2265, 3220, 397, 60, 1375, 959, 195, 2021, 3027, 3055, 3131, 169, 417, 1532, 1859, 2165, 2482, 3082, 132, 902]. Socio [1836]. Socio-Cultural [1836]. Sociocultural [334]. Sociological [2546]. sociologie [252]. Sociology [1094, 2918, 2081, 3221, 1149, 1722, 2586]. Socks [882]. Socrates [1465]. Soekarno [2741]. Soekarno-Era [2741]. Sofia [1408]. Sofie [2269]. Sogno” [1803]. Sohn [947]. soigner [505]. Soil [289]. Sojourners [276]. Sokhieng [2730]. Solandt [2137]. Solar [1138]. Soldiers [1909]. Solid [2716, 2954]. Solinas [2319]. Solomon [2253, 1209]. Solovey [1239]. Solution [102]. Solvay [439]. Solving [2649]. Some [1996, 1922, 1851, 20, 2268, 1997]. Somers [2535]. Sommaruga [2216]. Sommer [2528]. Sommerfeld [450, 3049, 1359, 3049]. Somsen [1591]. Son [810, 2419, 2332]. Sona [2286]. Sonar [2854]. Sonja [2629, 2914, 2140]. Sonne [1562, 2977]. Sonnets [782]. Sophie [2441]. Soraya [1664]. sorcellerie [645]. Sören [2454]. Sørensen [2490]. Sorger [2569]. Sorrenson [1707]. Sotto [890]. souffrance [1797]. Sought [2360]. Soul [1895, 2910, 863, 3032, 2978, 1347, 1045]. Soule [1708]. Sound [1504, 1015, 514, 2434, 2529]. Sounding [976]. Sounds [1218]. Source [177]. Sourcebook [2425]. Sources [2584, 2843, 720, 1123, 2840, 861, 315]. South [1389, 1271, 2127, 2945, 500, 335, 1875, 1525, 1307]. Southern [1013, 807]. Sovereignty [2074]. Soviet [2815, 1909, 1609, 395, 402, 1190, 290, 2893, 2597, 2140, 230, 1909, 2387, 3009]. Sovieticus [2590]. Soviets [1242]. sowie [2124]. Sowina [2333]. Soyoung [2557]. sozialwissenschaftliche [3151]. Space [71, 2894, 1804, 899, 1615, 592, 1909, 823, 2063, 2142, 1868, 172, 1077, 2437, 2464, 2893, 223, 1756, 458, 2225, 1234, 147, 3100, 2123, 891, 1432, 1748, 1604, 1157, 1235, 3234, 3175, 3265, 1909, 1689, 3000, 2969, 2889, 1185, 2625, 600]. Space-time [1615, 3100, 1689]. Spacefarers [1436, 3213]. Spaceflight [395, 2958, 3213, 1436]. Spaces [1917, 1251, 65, 2192, 508]. Spain [1983, 2885, 2659, 302, 2258, 1883, 2636, 1883, 2265]. Spallanzani [1798]. Spam [1308]. Spanagel [1712]. Spanish [282, 1271, 2112, 2157, 43, 1221, 2156, 1771]. Spannungsfeld [2799, 658]. Spark [1274]. Sparks [1423]. Spary [1353, 1711]. Spätantike [2422]. späten [2343]. Spatial [2873, 2625]. Spatio [1269]. Spatio-Temporal [1269]. Speak [623, 665]. Speaks [2690, 2498]. Special [334, 2734]. Speciality [2042]. Specialization [2036]. Specialty [948]. Species [1989, 1910, 1038, 2649, 1370, 3123, 2199, 2214]. Specimen [842]. Specimen. [2937]. Spectacle [717]. Spectacular [1224]. Spectatorship [2204]. spectroscopy [3257, 1726]. Speed [226]. Spencer [2945, 1894]. Spencerism [2809]. Sperling [1312]. Sperm [1837]. spezzino [1109]. Sphere [250]. Spheres [641]. Spherical [1194]. Spice [1875]. Spies [1626, 3158]. Spiesser [1126]. Spiller [2142]. Spinning [441]. Spinoza [1810, 1880]. Spirit [196, 1273, 2388, 675]. Spiritism [812]. Spirits [660, 1550, 3026]. Spirituality [693]. Spirometer [1725, 3233]. Splendid [1385, 255, 396]. Spoken [825]. Sponer [1005]. Sponsored [663]. Sprachen [1881]. Sprachforschung [1881]. Sprat [2936]. Spread [676, 1111]. Springer [1590, 1604]. Spur [511]. Spuren [635]. Spurious [50]. Sputnik [92]. Squares [2499]. Squeri [2328]. Squier [2596]. St [2294]. Staaten [947]. Stack [1368]. Staffan [2738, 2326, 2521]. Stage [2269, 1555, 3056, 3161, 1410]. Staging [2430]. Stahl [2795, 498, 2798]. Stalin [2815, 752, 2892]. Stamhuis [397]. Stamos [2803]. Stamp [2319]. Standards [1964, 2623, 1606, 4, 1751, 3286, 3057, 459, 851]. Stanford [1914]. Stanley [1901, 44, 410]. Star [1119, 2169, 2666]. Stargazers [1200]. Stark [200, 1369]. Starlight [1726, 659, 1298, 3257]. Stars [2866, 1120]. Started [2891]. State [1242, 1304, 880, 1483, 561, 2773, 1438, 2201, 2867, 2462, 1062, 702, 2954, 128, 2537, 663, 895, 1582, 2736, 3096, 2892, 2956, 1964, 2888]. State-Sponsored [663]. States [2595, 204, 1146, 1734, 479, 2444, 1265, 933, 216, 922, 535, 2147, 1759, 2872, 1503, 1686, 2550, 2976, 3244, 2968, 3281, 969, 2905, 3302, 2592, 2031, 2293, 1034, 826, 80, 958, 322]. States-Colombia [1146]. Stathis [1583]. Station [2294]. Stations [2284]. statistica [879]. Statistical [2411, 2122, 2278, 397]. Statistics [215, 397]. statistique [252]. statut [2332]. statuti [1802]. Staubermann [1040]. Steam [1039, 279, 884, 569]. Steam-Powered [884]. Steamship [569]. Steele [1785]. Steen [1740]. Steenbakkers [2708]. Stefaan [1784]. Stefan [1344, 2535, 2860, 272, 1312]. Stefani [2727]. Stefanie [2924]. Stefano [1344]. Stefanos [2814]. Steffen [1051]. Steiner [1592]. Steingutfertigung [1132]. Steinicke [1487]. Steinle [2274]. Steinmetz [1722]. Stella [2436]. stellae [2901]. Stem [1435, 1440, 309, 3199, 953, 3116]. Stén [1806]. Stenger [2939]. Stengers [2392]. Stenhouse [2945]. Steno [1644, 434, 3075]. Step [2395]. Stepchildren [452]. Stephan [2273]. Stéphane [1485, 369, 2703]. Stephanie [2117, 2167]. Stephanson [2019]. Stephen [1031, 687, 1716, 1355, 1220, 1093, 2428, 2116, 2669, 2150, 1429, 2866, 2042, 1131, 2652, 1205, 150]. Stephens [2800]. Stephenson [2294, 2286]. Stepping [22]. Stepping-Stones [22]. stereometria [2716]. Sterility [672]. Sterilization [1588, 3048]. Stern [1514, 1546]. Sternbilder [2990, 1636]. Sterne [2517]. Steve [1685, 1579]. Steven [2512, 1935, 2560, 1335, 2123, 2256, 2542]. Stevens [1439]. Stevenson [1371, 3127]. Stewart [91, 2171]. Stillwell [2247]. Stirpium [1350, 3081]. Stoa [486]. Stobart [2572]. Stock [2362]. Stockholm [1917]. Stoczkowski [89]. Stoff [1402, 1158, 2200]. Stoic [2388]. Stolberg [2335]. Stolz [2877]. Stomach [815]. Stone [690, 1428]. Stones [22]. Stooke [1089]. Stoppani [1578, 3172]. Storia [1578, 51, 1461, 2845, 205, 931, 1281]. Stories [459, 2225, 2948]. Storms [1667]. Story [1514, 2744, 1864, 592, 2373, 62, 2871, 1039, 1626, 2666, 1765, 2205, 2691, 3158]. Stotz [1542]. Stow [1536]. Strabo [2495]. Strahm [2216]. Stranded [1067]. Strange [3140, 276, 1633, 1605, 2026, 3051, 2205]. Strangest [207]. Strategic [134]. Strategies [1460, 1954, 2487, 1379, 3023]. Strauss [1073, 89]. Strbánová [2286]. Street [1855]. Streit [2944]. Stress [1833]. Strick [2043]. Strikes [713]. String [1755, 3285, 1075]. Strong [303]. Strongly [1274]. Structural [2764]. Structure [2727, 1092, 2175, 3011, 478]. Structures [2132, 2568, 216]. Struggle [1895, 2050, 1268, 1240, 1496, 1573, 898, 1688, 3078, 3295]. Stuart [2849, 2594, 2305]. Stube [46]. Stuber [272]. Student [2234]. Studentenbewegung [2668]. Students [783, 1483]. Studi [1338, 2985]. Studia [1563]. Studie [2053]. Studies [1523, 2985, 1036, 415, 1575, 88, 239, 77, 1580, 639, 1189, 1563, 1591, 44, 27, 1552, 3082, 1599, 1090, 2515, 2927, 1811, 1536, 830, 63, 69, 70, 64, 50, 3206, 3117, 1341, 1414, 2004, 1015, 691]. Studio [1504, 3111]. Study [2823, 286, 1003, 360, 781, 2497, 1779, 1325, 2564, 1785, 1693, 445, 647, 406, 1100, 2670]. Studying [1529, 3080, 1442]. Stuewer [2954]. Stuff [641]. Stuhl [2371]. Sturm [296, 2810, 1605]. Stuttgart [1592, 2124]. Styers [2487]. Style [1084]. Styles [2762, 106, 2002]. sua [2985, 1338]. suanjing [2628]. Suarez [1781]. Subalternity [2298]. Subject [485, 2620, 2533, 1093]. Subjecthood [2399]. Subjectivity [2398]. Subjects [1835, 1614, 225]. Sublime [2919, 1234, 712]. Submarine [1228]. Submarines [637]. Substance [794, 2954]. Subtilitate [2334]. Subtle [1110, 2935]. Subversive [1221]. Subverting [3272, 1696]. Succeeded [690]. Success [2426, 2464]. südarabische [2810]. Sudden [3238, 1706, 527]. Sue [1505, 1572]. Suffering [2856]. Suffrage [2725]. Sugar [2679, 2375]. Suh [2557]. Suite [321]. suivi [2502, 2170]. sujet [171]. Sujit [1137]. Sul [2570]. ’Sulbadipika [2419]. ’Sulbasutra [2419]. Sulian [230]. sulle [1276]. Sullivan [2117, 1961]. sull’opera [2985, 1338]. Sum [1579, 3090]. Sumer [925]. Sumio [2720]. Summary [1846]. Summers [1074]. Summits [1651, 3063]. Sun [2977]. Sunder [1159]. Sundials [865]. Sunrise [2352]. Sunshine [1141]. Sunspots [499, 2610]. suo [879]. Super [2144, 1344, 3038, 2103, 642, 35, 2258]. Superconducting [2144]. Superheavy [2744]. Supernatural [2424, 812]. Superstition [2903, 2967, 2908, 1550, 3026]. Supper [2700]. supplement [3185]. Support [607]. Suppression [2426]. Supraleitung [2124]. Surekha [2259]. Surgery [1737, 3273, 713]. Surman [2940]. Surprising [1725, 3233]. Surrey [1569, 1581, 1558, 1567]. Survey [1122, 1963, 2076]. Surveying [1843]. Surveyors [1205]. Surveys [1304, 163]. Surviving [1790]. Susan [1937, 2592, 2323, 1147, 1504, 2596, 2288, 2745, 1635]. Susanna [2633]. Susannah [2113]. Susanne [1714]. Sussman [1896]. Sustainable [997]. Suzanne [1459, 1737, 3273, 1838]. Svante [1247]. Svea [2668]. Sven [1942, 1591, 1409, 1942]. Swabian [3114, 1428]. Swan [2916]. Swanley [908]. Swanson [1837]. Sweat [1679, 3196]. Sweden [796, 895]. Swedenborg [372]. Swift [3082, 1414]. Swindlers [69]. Sygkepleriazein [1373, 3130]. Sylla [639]. Sylvie [1481]. Symbiosis [591]. Symbols [1547, 3162]. Sympathy [2717]. Symposium [3207, 1648]. Symptom [2519, 485]. Symptome [484]. synchronicity [3190]. Synchrotron [233]. Synergies [2687]. Synthesis [1347, 1391, 1767, 3032]. Synthetic [1740, 3293]. Synthetisch [440]. Syrische [2924, 2924]. System [2461, 1717, 2062, 2658, 3270, 2575, 1973, 561]. Systematicity [1244]. systematics [3254, 1749]. système [171, 1477]. Systems [2541, 315, 304, 2421]. Sytsma [2793].

T [2894, 2199, 1869, 45, 1370, 1077, 2042, 2241, 2652, 1801, 1716, 2869, 2041, 1985, 1385, 2106, 1220, 2921, 1865]. T. [1112]. Tabac [2938]. Tabery [1688]. table [1545]. Tables [1122, 1590, 1569, 1554, 1547, 1563, 1606, 1677, 1610, 1546, 1601, 1676, 1589]. Tacchini [3039, 1357, 3039]. Tacit [609]. Tact [2693]. Tadeusz [2182]. Taibi [2723]. Tail [453]. Tails [206]. Tak [1981]. Takashi [1719]. Take [597, 1173]. Taken [2075]. Taking [2684, 2999, 2083, 2393, 241, 1604]. Talas [1408]. Tale [950, 1336, 399, 1169, 2552, 624, 1509, 2779, 2900, 2891, 3125]. Talent [1510]. Talented [2234]. Tales [2040, 352, 882, 1266]. Talk [2678, 165]. Tamara [2191]. Tamás [2705]. Tambora [1826]. Tame [2891]. Taming [796, 881]. Tamson [2032]. Tangled [2960]. Tania [2208]. Tanja [1691]. Tansley [1295]. Tant [2008]. Tara [2372]. Targets [2552]. Tartaglia [1476, 2427]. Tarter [2816]. Taschwer [2453]. Taussig [1481]. Taxidermy [1037]. Tay [2456]. Tchikine [2639]. te [1744, 1955]. Teachers [783]. Teaching [2642, 64, 855, 2914, 445]. Teams [1978]. Tears [1679, 3196]. Tech [1505]. Technical [597, 2002, 3111, 2329]. Technik [2027, 2035]. Techniken [2941]. Technikwissenschaften [2347]. Technique [2652, 959]. techniques [1683, 122, 3215, 2569]. Technische [739, 1597, 3111]. Technocracy [2130]. Technocratic [2139]. Technocrats [2588]. Technologica [2307]. Technological [2297, 259, 2518, 1451, 216, 546, 675]. technologie [2703]. Technologies [594, 1818, 1583, 2822, 845, 2367, 3025, 358]. Technologietransfer [868]. Technologists [2539]. Technology [765, 2466, 716, 1296, 1242, 2617, 1569, 2172, 213, 1942, 257, 1533, 2004, 1516, 823, 239, 77, 1580, 1068, 824, 1733, 1583, 1107, 603, 1908, 1904, 2652, 2682, 1581, 2511, 1834, 1558, 2133, 1374, 27, 1504, 2325, 223, 198, 47, 1090, 228, 82, 2899, 232, 341, 734, 1672, 1294, 1314, 149, 111, 578, 1040, 2653, 69, 828, 3022, 70, 524, 1961, 230, 895, 444, 1372, 323, 466, 3043, 3012, 3299, 3128, 2693, 2923, 2890, 2492, 2097, 2880, 2129, 584, 748, 546]. Technology-Related [1672]. Technology.}Part [2920]. Technophilic [106]. Technopolitics [893]. Technoscience [2492, 1759, 153, 3281, 1156]. Technosciences [702]. Technoscientific [2534]. Ted [1566]. Tega [1652]. Teichmann [2517]. Teil [733]. Telecommunications [748]. Telefunken [586]. Telegrafenkabels [1303]. Telegraph [1217]. Telegraphic [2153]. Teleology [550, 2319, 2781]. Telepathic [2590]. Telescope [1047, 1569, 2464, 368, 313, 1234, 2791, 3146, 885, 735, 172]. Telescopes [981]. telescopio [1281]. Telesio [1873]. Television [1231]. Telkes [1225, 2585]. Telkes-Klein [1225, 2585]. Telliamed [649, 649]. Telling [1514]. Tells [2358]. tempête [297]. Temperate [2351]. tempio [205]. Temple [2884]. tempo [205]. Temporal [1269]. Temporality [1296, 843]. temprano [2567]. temprano-moderna [2567]. Temps [1538, 297, 2417, 875, 478]. Ten [734]. Tendency [969]. Tending [2704]. Tenn [1609]. Tennis [42]. Tense [2162]. tentations [564]. Tenth [570, 2499]. Teofilo [1047]. Teologia [727]. terapeutiche [2570]. Teresa [1836]. Term [1668, 2410]. Terms [2312, 519]. Terrall [1565]. Terre [1863, 371, 190]. Territoriality [198]. Territories [1934]. territorio [1109]. Territory [2459]. Terror [275]. Terry [2064]. Terzian [1510]. terzo [2850]. Tesfaye [2249]. Tesla [3037, 1534]. Tessicini [2509, 1560]. Test [2452]. Test-Tube [2452]. Testa [2383]. Testi [51, 2176]. Testing [2456]. testo [493]. Texas [1223]. Text [2420, 798, 1046, 1183, 2695, 2883, 2696]. Textbooks [45, 621, 817, 624, 1366, 622, 620, 623, 3219]. texte [2177, 1123]. textes [191]. textliche [2422]. Texts [1020, 2317, 1790, 1693, 873, 2166, 2563]. Textual [2827, 2166]. Textualism [2855]. Thalassaemia [480]. Thamer [1087]. Thate [2450]. Theater [1533, 2430, 241, 3183]. Theaters [783]. Theatre [2666]. Theatrum [51, 438]. Thébaud [2569]. Their [1694, 2848, 202, 2, 276, 1642, 2383, 459, 2206, 993, 1204, 922, 2542, 1468, 1963, 3150, 3064, 3274]. Theistic [1901]. Them [312]. Theo [1087]. Theodor [947]. Theodore [1816]. Theologies [1794]. Theology [271, 1922, 1702, 3227, 3243, 1721, 1880, 267, 119]. Theoretical [2457]. théoricien [1499]. théorie [1862, 34, 3141]. Theories [2007, 2753, 27, 960, 2587]. Theorists [2058]. Theory [2361, 1755, 450, 2432, 1710, 2446, 641, 139, 2875, 3251, 1822, 2623, 637, 2949, 507, 292, 304, 2041, 590, 1747, 3285, 2473, 2734, 439, 1899, 10, 299, 481, 2916, 3258, 3101, 3304, 2937, 2029, 1595, 1755, 1075, 97, 694]. Therapeutic [588, 2055, 2374]. Therapeutics [556]. thérapeutique [753]. therapies [3253, 1753]. Therapy [2678, 2652]. Theresa [1064, 1143]. Thérèse [1419]. Théret [1862]. ThermoPoetics [385]. Theses [2792, 274]. Thesis [883]. Thierry [1794, 1940]. Thijs [1932]. Thing [1521]. Things [724, 2094, 742, 734, 1204, 165, 241, 2348]. Think [222]. Thinking [78, 2573, 765, 2346, 2873, 2655, 2633, 2237, 1266, 2625, 3190, 256, 93]. Third [1715, 1302, 1528, 591]. Thirlmere [301]. Thirteenth [554]. Thirty [1433]. Thode [1065]. Thomas [1154, 1048, 2373, 308, 2936, 2138, 1605, 1736, 2060, 277, 2216, 1484, 53, 1064, 1891, 1815, 1795, 41, 1963, 1052, 1119, 1351, 2707, 2630, 1092, 1944, 2485, 1555, 2854, 2722, 1604, 1088, 29, 2189, 1434]. Thomaz [2643]. Thompson [677, 1440]. Thoms [2623, 1911]. Thomsen [2435, 618]. Thomson [2045, 308, 1502]. Thoreau [433]. Thornhill [1055]. Thornton [2191]. Thorny [277]. Thorsen [1403]. Thought [1495, 400, 1801, 2450, 2786, 292, 1811, 2125, 2310, 1965, 3065, 3154, 1193, 1396, 1637, 1091]. Thousand [86, 917, 734, 542]. Threat [2895]. Three [2489, 773, 221, 2421, 1008]. Three-Man [1008]. Thule [723]. Thuy [2527]. Tiago [2534]. Tian [736]. Tibet [1787]. Tick [2094]. Tiger [1501]. Tiggelen [1069, 2326]. Till [1754]. Tilman [1507]. Tim [1423, 2341]. Time [71, 1296, 2913, 2392, 2649, 701, 2731, 2238, 3, 1496, 2737, 189, 2874, 489, 2862, 2163, 1854, 1546, 3294, 1615, 3100, 1689, 2494, 1546, 2291, 1820, 121]. Timelessness [468]. Times [724, 2531, 1478, 2318, 1305, 76, 2655, 3175, 2123, 152]. Timmermann [1552, 1746]. Timothy [1458, 1474, 2805, 2812, 1690, 2374, 2205, 1875]. Tina [2628]. Tine [2519]. Tissot [1352, 3092]. Tissue [1835, 959]. Titchener [237]. Tobacco [521, 854, 2909]. Tobias [2799, 2811, 23]. Tobies [1592]. today [3111]. Todd [2814, 1227]. Tofigh [27]. Together [341, 42]. Tognoni [1417]. Tokugawa [2737]. Tokyo [1363, 3088]. Tom [1341, 2299, 1434]. Tomás [2636]. Tomash [186]. Tomaso [1203]. Tomory [2634]. Ton [2887]. Tone [2481]. Tong [1304]. Tongue [2236]. Tongues [2237]. Tonmischpults [1597, 3111]. Tonstudio [1597, 3111]. Tony [1589]. Too [630]. Took [1909]. Tools [1991, 1396, 2614, 3065, 64]. Topiwala [2047]. Topography [2928]. Topp [2447]. Tore [2484]. Toro [1146]. Toronto [1588, 1566]. Torpedo [3244, 2976, 1734]. Torrella [38]. tot [2887]. Totelin [2562]. Touber [2708]. Touch [2150, 1157]. Tousignant [2817]. Touwaide [2330]. Towns [2333, 1469]. Townsend [1503]. Toxic [3245, 2592, 1681, 2357, 454, 682]. Toxicants [3201, 1664]. Toxicology [965, 2817]. tra [51, 931, 1490, 870]. Trabucco [51]. trace [942]. Tracing [1527, 2410, 1820]. Track [2903]. Tracking [594]. Tractatus [788]. Trade [1424, 703, 556, 1703, 1875, 3072, 3271, 955]. Trading [839, 1671, 465, 829]. Tradition [780, 2762, 1098, 1891, 3294, 1546, 3115, 2178, 1142, 1466, 932]. Traditional [2920]. Traditions [783, 2106, 2367, 2995, 1511, 1404, 476]. Tradizione [3115, 870, 1652]. traduction [3141, 2177, 785]. Traduire [26]. traduzione [493]. Trafficking [214]. Tragedy [2955]. Trail [700]. Training [1020, 622]. Traité [2502, 1123]. Traktats [2257]. Tramelli [2929]. Tranquilizers [90]. trans [2696, 2949]. Transactions [2574]. Transatlantic [1833, 2222]. transatlantischen [2143, 1303]. Transcendence [2325]. Transcript [1423, 3005]. Transcription [1370, 3123]. Transdisciplinarity [1153]. Transdisciplinary [2372]. Transfer [2823, 230, 1332]. Transformation [2512, 388, 2459, 840, 2301, 2371, 3111, 1942]. Transformations [1511, 1580, 280, 1124, 2554, 1815, 3141, 1110, 2436]. Transformative [2167, 2410]. Transformed [1086, 662, 746]. Transforming [2529]. transformisme [649]. Transfusion [66]. Transfusions [1067]. Transit [674, 1340, 3053]. Transition [2312, 2648]. Transitional [554]. Translated [118, 1575, 1677, 497]. Translating [2611, 2612, 2996, 2479, 621, 2609, 2610, 2613, 2480, 2797, 1409]. Translation [1198, 1698, 2475, 2446, 554, 2607, 1042, 2481, 2477, 2603, 2516, 1622, 1183, 269, 647, 3249, 3252]. Translations [118, 2696, 2972, 1693]. Transmission [736, 188, 2153, 830]. Transmutation [482, 1552, 3117]. Transnational [2662, 688, 1759, 105, 3281]. Transparency [68]. Transplantation [66]. Transplants [1264, 318]. Transuranic [2744]. trasformazione [784]. T̈rattato [2929, 1641, 3203, 1276]. Traum [2449]. Travaux [45]. Travel [225, 2885, 692, 946]. Travelers [1319]. Traveling [1674]. Traveller [2864]. Travellers [500, 1337, 3126]. Travels [2028]. través [859]. Travis [31, 2374, 1221]. Treasure [650]. Treasuring [3001, 1349]. Treatise [986, 2446, 781, 174, 2046, 3203, 2428, 1288]. Treatises [2499, 118, 2101]. Treatment [862]. Tree [2960, 362]. Trees [2452, 661, 860]. Trente [40]. Tria [2821]. Trial [1618, 1583, 1366, 1627, 1084, 3025, 3076, 3219]. Trials [225, 2231]. Triangulation [2827]. Tribble [2790]. Tribe [2361, 1003]. Trickster [2912]. Trigonometry [1194, 33]. Trilogy [3200, 1741]. Trimble [2322]. Triumph [464, 2415, 908, 1371, 3127]. Triumphant [637]. Triumphator [1181]. Troels [1644]. Troeltsch [3227, 1721]. Trollinger [2745]. Tropes [2161]. Tropical [948]. Tropics [2951, 2930, 2994, 1356]. Trotsky [259]. Troubled [2031, 447]. Troubling [1896]. Trout [2415]. True [3140, 1396, 62, 1633, 3065, 457]. Truitt [2010]. Truman [752]. Trust [926]. Trut [2891]. Truth [941, 521, 2863, 714, 1705, 3263, 2980, 1946, 434, 710, 147, 405, 85]. Truths [2415, 460]. Trying [1366, 3219]. Trysh [2374]. Tsar [290]. Tsarist [132]. Tsu [1733]. Tsun [2059]. Tube [2452]. Tuberculosis [878, 1609, 1966, 3009]. Tübingen [373]. Tucker [2026]. Tucson [2465]. Tudor [1478]. Tung [1579]. Tunnel [592, 2144]. Turbulent [90, 1305]. Turda [2584]. Turek [1799]. Turin [1556]. Turing [2952, 2216, 3303, 2455]. Turkish [2478]. Turn [2684, 1173, 2855, 1153, 405, 567]. Turner [705]. Turning [961]. Tuscans [1694, 3274]. Tuscany [2260]. tushuo [736]. Twelfth [1342, 2987]. Twentieth [2490, 1309, 214, 1402, 1983, 389, 142, 1005, 1896, 2588, 102, 2519, 1977, 2953, 1738, 1077, 2134, 1737, 2459, 1438, 1739, 2547, 1591, 2171, 1606, 2882, 2518, 131, 1829, 607, 1833, 66, 1911, 852, 2130, 1491, 151, 2537, 1593, 1976, 976, 999, 959, 567, 2195, 3242, 3057, 3273, 3183, 3121, 3301, 3122, 1247, 1102, 1021, 3198, 2088, 3096, 1105]. Twentieth-Century [1309, 1983, 2588, 1737, 1438, 1591, 2518, 1491, 2537, 1976, 3273, 3183, 1021, 3198, 3096, 1105, 1247]. Twenty [1515, 2835, 861, 523, 2915, 2066, 221, 1816, 1024]. Twenty-First [2835, 2915, 2066, 1515, 861, 523, 221, 1024]. Twenty-four [1816]. Two [1726, 2499, 730, 1447, 86, 1855, 1169, 2604, 2552, 1576, 624, 2511, 637, 2900, 917, 87, 1201, 2396, 1925, 1899, 3144, 3257]. Two-Thousand-Year-Old [86]. Two-Way [1855]. Tyndall [1577, 738, 2688, 2822, 3267]. Tysons [77].

U [1962, 2591, 456]. U.K [1241]. U.S [2297, 1829, 2724]. U.S. [1451, 465, 969]. Übersetzung [2257, 1871]. Uexkülls [140]. Ugaritic [1937]. Uhl [2535]. Uhren [928]. uirtutibus [2500]. Ulf [2056]. Ulinka [2111]. Ulisse [2845]. Ulrich [39, 1087, 53, 1795]. Ulrichs [2197]. Ulrike [2623, 1911, 1830, 1562, 2949]. ultimam [2103]. Ultra [755]. Ultra-Darwinists [755]. Umanisti [2719]. Umbruch [658]. Umfang [990]. umstrittensten [2453]. Umwelt [140]. Unbranding [2057]. Unburied [572]. Uncertain [3004, 1663]. Uncertainty [965, 2729, 1236]. Uncommon [3083, 1213, 1362]. Unconscious [2660, 832]. Undercontextualization [1852]. Underground [1889]. Understand [1016, 1688, 3295]. Understanding [1483, 1020, 2461, 2291, 2000, 547, 1267, 279, 2217, 325, 1117, 96]. Underwater [209]. Undisciplining [1977]. Undoing [2252]. Undying [3230, 1709]. Unearthing [3289, 1743]. Uneasy [2597, 163]. Unemployed [2387]. Unending [843]. Unfinished [2742, 501]. Unfreedom [2296]. Unfreezing [2371]. ungarische [2190]. ungüento [787]. Unification [540]. Unified [1899]. Unifying [2509]. Unintended [715]. Union [230]. Unità [931]. United [2550, 2595, 204, 2592, 1146, 1734, 2031, 479, 2444, 826, 2293, 1265, 1034, 933, 216, 80, 958, 1759, 2872, 1503, 322, 1686, 2976, 3244, 2968, 3281, 969, 2905, 3302]. Uniting [1016, 1369, 3112]. Unity [496]. Universal [1102, 2340, 2239]. Universale [2306]. Universe [576, 1396, 2706, 3065, 3014, 2919, 1364, 368, 313, 113]. Universitaires [1594]. Universitäre [2699]. Universität [658, 53, 658]. Universitäten [61]. Universitäts [53]. Universitäts- [53]. Universitätsstädten [2504]. Universiteit [2777, 2777]. Universities [2032, 2363, 1328, 2195, 2806, 2984, 2542]. University [1598, 1585, 1588, 1590, 1568, 1177, 1586, 1543, 1483, 1579, 1609, 1566, 74, 1535, 1269, 2884, 731, 1547, 1548, 1565, 1605, 1534, 1602, 2844, 1606, 1545, 1555, 1610, 1573, 1557, 1607, 666, 1601, 1571, 1536, 1542, 1914, 1587, 2865, 3218, 1593, 1676, 1570, 2743, 1948, 1608, 1595, 1603, 1582, 1544, 522].

Unmaking [1033, 1767]. Unmasked [1896]. Unofficial [2134]. Unpublished [2263, 1288]. Unpuzzling [1613]. Unravelling [1298, 659]. Unravels [2919]. Unreasoned [1582, 2998]. Unruly [1524, 660]. Unscientific [1896]. Unseen [172, 3110, 1544, 2009]. Unsichtbare [28]. Unsimple [460]. Untersuchung [61]. until [2333, 3111, 1865, 568]. Untilled [196]. Untimely [1819]. Upon [2903]. Uranium [955]. Urban [2802, 2369, 532, 2542, 1469]. urina [785]. Urologie [2053]. uroscopie [785]. Uroscopy [2335]. Ursprung [2944]. Ursula [1783, 2027, 2347]. Urszula [2333]. USA [2523, 2209]. Use [4]. users [3150, 149]. Uses [2171, 1792]. Using [2827, 2311, 1011]. Usitalo [1335]. Usselman [1950]. USSR [1909]. Ute [1631]. Uterus [2413]. Utopia [1708, 532, 3231]. Utopian [1533, 1587, 3022, 3069]. Utopianism [259, 91]. Utopias [1021, 1456]. Uwe [2522].

V [1454, 2211, 2863, 2301, 2048, 2031, 2234, 2080]. Vaccination [2031]. Vaccine [2598, 1085]. Vaccines [2594]. vacío [2084]. Vague [768]. Vakoch [2061]. Valencius [1420]. Valentine [1482]. Valiant [1428, 3114]. Valle [2355]. Valleriani [2568, 1476]. Vallin [2808]. Vallisneri [2937]. Value [404, 1525]. Value-Free [404]. Values [1252, 2886]. Vance [2745]. Vandendriessche [2201]. Vanderbilt [1609]. Vanderpelen [1060]. Vandersmissen [2008]. Vargha [2957]. Variable [1119]. Varietäten [281]. Variety [568]. Varying [3177, 2052]. Vast [518]. vaticani [199]. Vaticano [75]. Vaughan [1031]. Vecchia [1075]. Vecchio [1803]. védique [2419]. Vegetable [2113]. Veils [1173]. Veit [1593]. Velasco [2510]. Velikovsky [1010]. Velminski [2590]. Velvet [233]. Venereal [2265]. veneti [1109]. Veneto [1109]. Venezia [2570]. Venice [783, 876, 725, 366]. Vennen [2941]. Venom [2558]. Ventura [1752]. Ventures [2120]. Venu [50]. Venus [1562, 2977, 674, 1493, 2977]. Vera [2002, 1219, 2109]. Verbraucherpolitik [2200]. Vercherre [1792]. Vereinigten [947]. Vererbung [510]. Verfahren [635]. Vergleich [2143]. Vérin [2703]. Verlag [1562, 1575, 1597, 1592, 1574]. Verlegung [1303]. verlorenen [511]. Vermarktung [378]. Vermeulen [2115]. Vernacular [118, 7]. Verne [1371, 3127]. Vernet [706]. Vernunft [1881]. Veröffentlichungen [1188, 46]. Verona [1109]. Veronese [1109]. Veronika [2735]. verre [988]. Versa [164]. Verschiedenheit [484]. Verschuer [238]. Verse [3117, 1342, 2987, 1552]. Versus [1688, 1994, 2908, 2473, 3295, 2613]. Vertebrate [1423, 3005]. Vertesi [1685, 1978]. Vertical [2386]. vertiges [153]. Vertriebene [2209]. Vertus [1123]. Very [1800, 2089]. Vesalius [3249, 2927, 1622, 2632]. Vesuvius [1282]. Veterinary [1549, 3107]. Vetter [2583]. Via [2276, 1873, 784]. Vibrations [1397, 3007]. Vice [164]. Vicedo [1313]. Vicente [2722]. Vico [2851]. Victor [1558, 2425, 1877]. Victoria [1182, 1368, 1829, 1433, 2973]. Victorian [665, 2728, 1718, 1135, 588, 385, 2198, 2650, 2943, 2966, 2362, 1957, 2040, 2942, 1727, 2437, 1368, 376, 2869, 2582, 674, 581, 2717, 1140, 738, 1421, 1653, 409, 838, 3138, 3045, 2039, 3268, 2661, 243, 3218, 3287, 1137, 2973].

Victorians [1719, 1821, 881, 3259, 3260, 3261]. Vidal [2620]. vide [1551, 3124]. vie [2014]. Vielfalt [1881]. Vienna [2655, 2535, 1224, 214]. Vier [2777]. Vietnam [2527]. View [1185, 2160, 2416, 2413]. Views [2356, 1658, 590, 3179, 3171, 2968, 3089, 3183, 3188, 1716, 1219]. VII [1344, 2427, 1543, 1539, 1601, 2991, 3038]. viii [1585, 1562, 1537, 1563, 1605, 1602, 1610, 1573, 1542, 2427]. Vila [2856]. Vilhelm [2871]. Villard [490]. vinariorum [2716]. Vincent [1225, 1863, 2585, 274]. Vincenzo [1868, 2263]. Vinci [172, 1795]. Vindicated [2532]. Vine [662]. Violaine [1553]. Violence [799]. Virbel [2166]. Virginia [1507, 60, 383]. Virol [2703]. Virtual [316, 1456, 1160]. Virtue [1916]. Virtues [2663, 1916]. Virtuosi [2017]. Virtuoso [49, 2341, 800]. visa [785]. visage [571]. visible [966, 3143, 2411, 1050]. Vision [1521, 1850, 140, 1844, 1349, 368, 1254, 2221, 738, 2455, 3001, 1273, 368]. Visionaries [2917]. Visionary [2891]. Visioneers [1235]. Visions [130, 1516, 2144, 889, 2206, 3218, 3287, 1703, 3271, 1718, 2656]. Visual [446, 638, 1050, 2893, 1506, 1411, 1779, 2633, 860, 2961, 1703, 3068, 3271, 1474, 2110]. Visualisierungen [1562, 2977]. Visualization [2278]. visualizations [2977]. Visualizing [1290, 1452, 2776]. Vítêzslav [1928]. vita [1490, 2197]. Vital [2033]. Vitalism [415, 1494]. Vitamin [378]. Vitamine [1158]. Vittorio [2851]. vivant [431]. Vivek [2741]. vivens [23]. Vivian [2632, 1196]. Vivisecting [409]. Vlad [1130]. Vladislav [1333]. Voeks [2839]. Vogel [1240]. Vogt [1005, 1592]. Vöhringer [1955]. Voices [1909]. voir [274]. Vol [2651]. volatile [2184]. Volker [2339, 2778, 2490, 1597]. Volksschulen [208]. Voltaire [2852]. Volume [331, 330, 614, 613, 905, 904, 1163, 1162, 1445, 1444, 1762, 1761, 2068, 2067, 2377, 2376, 2674, 2673, 2963, 2962, 1932, 1507, 2920, 2991, 210, 1184, 2946, 1485, 756, 39, 896, 369, 953, 397, 1417, 2777, 789, 3118]. Voluntary [132]. Vom [1087, 2944, 2699, 2978, 802, 281, 1398, 2187]. von/In [593]. Vorlesungen [2971, 1412, 733]. Vorlesungprogramm [53]. Vorstellung [486]. Vos [43]. Voskuhl [1486]. Voyage [697, 1419, 2786, 3192, 2997]. Voyages [139, 1150]. Voynich [2847]. vs [22, 2298]. vspominaetsja [1981]. vue [1863]. Vuillemin [1210]. Vulcan [1109]. Vulcano [1109]. vulgarisation [3250, 1553].

W [2586, 1694, 2795, 1837, 2172, 2643, 1341, 2622, 2127, 2317, 1426, 2463, 1647, 2144, 1966, 2271, 133, 1545, 1716, 1690, 896, 1816, 1205, 1144, 2374, 2028, 2642, 1465, 1948, 2724, 677]. W. [1534, 1756, 1731, 2036, 2474, 1235]. Waddell [1945]. Wade [349, 1752]. Wagner [2019, 2628, 2425, 1212, 1874]. wahren [2187]. Wahrheit [719]. Wahrig [1402]. Wahrnehmung [3210, 1691]. Wahsiya [2257]. Wailoo [1864]. Waiting [145, 2328]. Wake [2523, 2726, 1253]. Wald [1896]. Waldby [1835]. Wäldern [743]. Waldman [1201]. Walked [310]. Walker [1241, 1302, 47]. Walking [3275, 1492, 1690]. Wallace [1370, 2199, 3123]. Wallis [1049, 1793, 724]. Walloch [2031]. Walls [1716]. Wallstein [1574]. Walpole [1568, 1522]. Walsh [1780]. Walter [1898, 1652, 1561, 1279]. Walther [1499, 2287]. Walton [2512, 2106]. Waltraud [2173]. Wandel [1597, 3111]. Wandlungen [1233, 1972]. Wang [2628]. War [2956, 2550, 1242, 1177, 744, 92, 2592, 2142, 2220, 1012, 2976, 907, 19, 2955, 2541, 1663, 2141, 1229, 1605, 2462, 1239, 2157, 2133, 1662, 2131, 2137, 1610, 2957, 898, 2222, 448, 295, 149, 1823, 1740, 952, 996, 1603, 895, 580, 1313, 3044, 3098, 3293, 2156, 2240, 3137, 108, 107, 3051, 1591, 3004, 109, 3198, 106, 410, 2307, 105, 2616, 2814, 2712, 2540, 1908, 1663, 2725, 1177, 744, 142, 893, 232, 888, 751, 580]. Warburg [1556]. Ward [1790]. Wardhaugh [2858]. Warfare [2056, 173, 895, 2976, 1149]. warming [3152, 1607, 521, 518]. Warr [1692]. Warren [2372, 2940]. Wars [2392, 19, 1010, 426]. Wars.} [2954]. Warsaw [1563]. Was [2912, 1287, 1971, 3101, 2196, 1595]. Washington [1600, 1607]. Wasn’t [50]. Waste [744, 322]. Watch [1604, 2999]. Watcher [565]. Watching [1282]. Water [676, 2634, 2333]. Watermen [356]. Waterpower [390]. Waters [2186]. Watkins [1569, 1710, 1628, 3146]. Watt [279, 1879, 578]. Wave [2146]. Waves [2590, 514, 1372, 3128, 2671]. Way [1855, 2607, 173, 1860, 2465]. Ways [3290, 1545]. Wazeck [1822]. Weak [1100]. Weapon [1100]. Weapons [2220, 1241, 315, 1229, 2462, 1190, 827, 1823]. Wear [259]. Weather [2360, 2871, 598, 81, 599]. Weathered [2402]. Web [441, 1291]. Webel [2124]. Weber [2799, 1134]. Websites [2614]. Webster [707]. Wednesday [2700]. Wee [2841]. Weeden [74]. weerga [2887]. Weg [2287]. Wegener [2041]. Wei [1314]. Weighing [1055]. Weights [2427]. Weill [1470, 1551]. Weill-Parot [1470, 1551]. Weimar [130, 883]. Weimarer [745]. Weindling [677]. Weiner [1307, 1174]. Weinreich [1100]. Weinstein [2734]. Weintraub [1754]. Weir [1227]. Weismann [2868]. Weiss [24, 2188]. weiße [2290]. Weisz [1739]. Weitekamp [1336]. Welcoming [1856]. Welfare [895]. Well [972, 692, 2706, 1457]. Well-Mannered [1457]. Well-Ordered [972, 2706]. Wellcome [742, 1549]. Welt [2591, 1303, 46]. Weltbilder [1212]. Welten [2337]. Weltreisende [2799]. Wen [2059]. Wenchao [1881, 1657]. Wende [2190]. Wendt [2412, 1907]. Wendy [1658, 1232]. Went [455]. Wer [1230]. Werden [39]. Werdens [678]. Were [364]. wereld [2785]. Werk [140]. Werrett [2859]. Wesleyan [1585]. West [2275, 1642, 1065, 1488, 2583, 1458, 3030, 112, 3064, 3275, 2281, 2180, 2248, 1690, 892]. Western [2848, 257, 172, 174, 354, 2325, 2618, 2167, 2131, 368, 1151, 718, 2168, 3188, 1098, 6]. Westman [2506]. Westminster [74]. Weston [2658, 1797]. Weststeijn [1932]. Wetenschap [309, 3246, 953, 1732, 2442]. Wetenschappen [953, 309]. Wexler [2128]. Weyl [1594, 3141]. Whale [103, 976]. Wheeler [2731]. Whewell [1988]. Which [2612]. Whipple [2959, 27]. Whitaker [1954]. White [2361, 1059]. Whitewashing [238]. Whitmer [2020]. Who [2360, 1909, 1214, 2531, 310, 2365, 2739, 2088, 746, 567, 466, 1513, 2690, 303]. Whole [91]. Whooley [1496]. Whose [2907]. Wibke [1398]. Wicked [3068, 1411]. Widmalm [1591]. Wied [679, 679]. wieder [586]. wieku [1865]. Wiesbaden [1575]. Wife [714, 2980]. Wiktor [89]. Wilde [1139].

Wilder [2491]. Wilderness [594]. Wilding [1813]. Wildish [2294]. Wildlife [594]. Wildness [2536]. Wiley [1686]. Wilfried [1590, 1590]. Wilhelm [566, 1087, 2270, 1188, 873, 124, 2043, 2657]. Wilkins [2637]. Will [1176, 1173]. Wille [1730, 2738, 2521]. Willet [1575]. Willey [2252]. William [1590, 1962, 576, 1487, 1214, 659, 639, 1952, 2175, 1950, 884, 1330, 2277, 1298, 1985, 2745, 1133, 1228, 2262, 1988, 838, 1715, 2966, 2459, 1074, 2138, 1150, 1039, 2637, 2267, 1952, 1974, 3276, 2666, 2315, 2520, 2089, 54, 1288, 810]. Williams [2879, 186, 1119, 1337, 1355, 1371, 2087]. Willie [255]. Willinsky [2775]. Willughby [2341]. Wils [2730, 2519, 2201]. Wilson [1697, 1214, 1348, 1140, 1410, 2544, 2548, 1842]. Wimmer [2273]. Win [2526]. Wind [592, 1462, 3103, 1377, 1377]. Window [368]. Wine [1792, 2716, 662]. Winnewisser [1005]. Wins [2907]. Winston [1342]. Winter [399, 2949, 2404]. Winterbottom [2249, 2345]. Winthrop [375]. Wintroub [2786]. Wire [306]. Wired [605, 506, 594]. Wireless [1372, 3128]. Wirklichkeit [2657]. Wirkstoffe [1158]. Wirtén [1970]. Wisdom [2016]. Wismer [1942]. Wisnioski [1516]. Wissen [2143, 2347, 2337, 272]. Wissens [1562, 2977, 2343, 2354]. Wissenschaft [2990, 990, 1631, 3058, 739, 1636, 2027, 658, 2449, 53, 401, 2668, 2035, 1145, 1186, 1955, 438]. Wissenschaften [2810, 503, 3151, 802, 281, 1422]. Wissenschaftler [740, 2209]. wissenschaftliche [1475]. Wissenschaftlichkeit [2183]. Wissenschaftsatlas [1269]. Wissenschaftsgeschichte [2024, 1158, 53]. Wissenschaftsideals [573]. Wissenschaftspolitik [2454]. Wissensgeschichte [1814]. Wissensproduktion [2941]. Wissenstransfer [2190]. Witch [2111]. Witchcraft [876]. Witches [269]. Withers [2172, 2622, 1144, 2028]. Within [2528, 2617, 989]. without [2555, 1299, 3, 1374, 3012, 449, 405, 2364]. Wittje [2529]. Wittmann [593]. Wittner [2653]. Wizard [1377, 3103]. Wladimir [2590]. Włodarczyk [1563]. Wolf [1201]. Wolfe [1242]. Wolffram [452]. Wolfgang [1487, 1636, 2273, 2591, 1815, 2035]. Wolfhechel [1460]. Wollaston [1950]. Wolschke [2339]. Wolschke-Bulmahn [2339]. Woman [966, 1005, 310, 430]. Womanpower [2139]. Womb [2704]. Wombs [2413]. Women [1460, 1625, 1735, 2539, 1148, 479, 1658, 2739, 908, 1884, 1728, 295, 3018, 652, 1328, 775, 713, 3298, 1963, 2984, 3255, 3209, 3156, 3089, 2808, 2225, 1592, 1867, 1505, 357, 896]. Won [2903, 2967, 2908]. Wonder [2112, 1077, 275, 243, 2626]. Wonderful [1449]. Wonderland [2040]. Wonders [2424, 1559, 2694, 3157]. Wondrous [2415]. Won’t [1005]. Wood [2848, 1826]. Wooden [2848]. Woodford [2733]. Woodmansee [1033]. Woods [2732]. Woodson [60]. Woodward [1974]. Wook [2206]. Woolgar [1685]. Worboys [2595]. Word [796, 825, 2931, 2708]. Wordes [781]. Words [2631, 112, 2912, 2479, 416, 773]. Work [656, 2059, 1048, 469, 321, 2286, 1226, 1581, 2949, 2217, 2474, 2589, 2416, 65, 2704, 2985, 3167, 2173, 397, 1954]. workhouse [3214, 1401]. Working [2386, 1150, 962]. Works [2510, 44, 1517, 416, 2587, 1202]. Workshop [583]. World [1114, 2835, 1052, 2386, 2240, 2058, 435, 2491, 2710, 422, 2815, 2032, 1927, 1919, 1346, 744, 1483, 389, 1248, 2051, 1336, 2592, 3047, 1068, 853, 222, 2283, 1299, 1055, 547, 1563, 856, 2437, 1591, 1039, 2178, 2267, 843, 4, 2169, 1677, 2325, 1545, 2781, 3198, 2715, 1922, 2131, 2487, 2603, 896, 1879, 1500, 2954, 1406, 1279, 283, 2702, 1559, 470, 579, 2725, 295, 2679, 1199, 1787, 2602, 2260, 722, 1823, 2889, 2364, 2306, 1827, 2348, 2732, 7, 2566, 925, 2821, 41, 996, 1613, 2044, 2806, 580, 3229, 2972, 3077, 3017, 3018, 3184, 3125, 3127, 2976, 2133, 1608, 2249]. World [2432, 2412, 2526, 2541, 2365, 2786, 1885, 1043, 2345, 2921, 1371, 2851, 1826, 2626, 25, 282, 821, 742, 746, 729, 631]. Worldly [688]. Worldmaking [927]. Worlds [316, 494, 2479, 2259, 2638, 243, 773, 1201, 2791, 951, 2940, 2341, 3248, 202, 112]. Worldview [1227]. Worm [1226]. Worse [1069]. Worst [827]. Worth [1818, 1619, 2088, 185]. Wotton [1492]. Woudhuysen [1781]. Would [637, 259]. Wound [925]. Wounds [3264, 1692]. Wouters [1160]. Wray [1017]. Wrestling [634]. Wright [2366, 2192, 2044]. Write [694]. Writing [2631, 1011, 2769, 842, 1199, 2028, 2558, 1867, 1263, 924]. Writings [1507, 1047, 1456, 1865, 686, 1462]. Written [2329, 188]. Wrong [755]. Wu [2059, 2283]. Wuidar [40]. Wujastyk [1457]. Wunder [1472]. Wundt [566, 2438]. Wußing [708]. Wyatt [1160].

X [1880, 2942, 1817]. X-ray [1817]. Xavier [1745]. Xenarchus [1467]. Xhayet [2008]. xi [1598, 1588, 1545, 1559, 1582]. xiang [230]. Xiao [2283]. Xiaoping [1151]. Xiaotong [2628]. XII [1463, 1590, 1600, 1535, 1599, 1604, 1564]. XIIe [2179, 867]. xiii [1596, 1534, 1557, 1584, 1587, 1593, 1589]. XIIIe [36, 2697, 1632, 785, 1551, 3124]. XIIIe-milieu [1551, 3124]. XIV [1949, 1609, 1558, 1552, 1567]. XIVe [34]. XIVe-XVIe [34]. XIXe [297, 505]. xv [1554, 1577, 1583, 1581, 1592]. XVe [2179, 36, 1551, 2703, 2569, 3124]. xvi [1566, 1560]. XVIe [171, 178, 2014, 779, 1477, 2569, 34, 2342]. XVII [866, 1544]. XVIIe [1481, 2697, 2342]. XVIII [1184, 1865, 1579, 1607]. XVIIIe [1481, 564, 2014, 1203, 1632, 2569, 2703]. XX [1184, 1463, 1536]. XXe [2170]. xxi [1546]. xxii [1549]. xxiii [1547]. xxix [1597]. xxv [1590]. xxvi [1569, 1572]. xxxii [1314]. xxxv [1570].

Yaakov [2787]. Yale [1568, 2266, 1610]. Yamada [2720]. Yan [2446, 2099]. yanjiu [736]. Yasmin [1053]. Yasu [2881]. Yau [1579]. Yeang [1372]. Year [86, 542, 2794]. Years [1507, 1615, 1975, 1579, 824, 2175, 2049, 2060, 489, 917, 857, 2734, 891, 1948, 3090, 31]. Yellow [384, 2189]. Yen [1743]. Yeo [2017]. Yi [1914, 230]. Yiddish [1100]. Yield [997]. Yiftach [2248]. Yingshi [2603]. Yitzhak [1810]. Yoder [1564]. York [1591, 789, 1317, 3291, 2807, 2344, 1712]. York/London [1591]. Yost [1330]. Yota [2639]. Young [2712, 1810]. Youth [1510]. yu [736]. Yucca [322]. Yudell [1896]. Yuen [2233, 2902]. Yuhanna [1098]. Yulia [2737]. Yung [1539]. Yves [1882, 1405, 2092]. Yvette [1713].

Z [2841, 808, 2456, 1455, 1676]. Zaballa [2955]. Zaca [697]. Zach [2355]. Zachary [2501]. Zachhuber [1721]. Zadig [942]. Zagan [373]. Zakariya [2691]. Zambelli [2289]. Zammito [2798]. Zanoni [1578]. Zeal [1566, 3135, 1213]. Zealand [2945, 1376, 3015]. Ze’ev [1507, 2885]. Zeichentheorie [1881]. Zeichnen [635, 1136]. Zeichnern [28]. Zeiss [2273]. Zeit [296, 3212, 589, 281, 2422, 1657, 2453, 511]. Zeitschrift [1498]. Zelda [1867]. Zentaur [1807]. Zentralverein [2292]. Zerbino [1202]. Zero [2622]. Zhang [736]. Zhmud [1044]. Zhongguo [223, 230]. Zhores [601]. Zhu [2446]. zhuan [230]. Ziche [1373]. Zijn [2451]. Zik [2787]. Zika [2895]. Zilberstein [2351]. Zilinskas [1190]. Zita [2095]. Zittel [756, 1815, 1679, 2916]. Zodiac [808]. Zoe [1593]. Zon [1727]. zonder [2887]. zone [3270, 1717]. Zones [1671, 829]. Zoo [1719, 2837, 1037, 3259, 3260, 3261, 3088, 2536, 1363]. Zoology [3243, 1702]. Zoos [1403, 3223]. Zugänge [2978, 1398]. Zuidervaart [1375]. Zum [288, 2978, 1087, 2287, 1398]. Zur [2200, 28, 733, 1132, 2971, 2257, 1892, 1303, 53, 1412, 1332]. zusammengestellt [288]. Zuse [1230]. Zvi [1677, 1876]. zweier [1212]. Zwijnenberg [1788]. Zwillingsproblem [543]. zwischen [868, 2990, 743, 1636, 2668, 1472, 124]. Zwodniczy [1865].

References

Anonymous:2010:CBH

[1] Anonymous. Current bibliography of the history of science and its cultural influences, 2010. Isis, 101(S1):xxx + 327, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

      www.jstor.org/stable/10.1086/660768.

DeBont:2010:OTM

[2] Raf De Bont. Organisms in their milieu: Alfred Giard, his pupils, and early ethology, 1870–1930. Isis, 101(1):1–29, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

      /www.jstor.org/stable/10.1086/652687.

Hoquet:2010:HTB

[3] Thierry Hoquet. History without time: Buffon’s natural history as a nonmathematical physique. Isis, 101(1):30–61, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

      www.jstor.org/stable/10.1086/652688.

Kirk:2010:BNA

[4] Robert G. W. Kirk. A brave new animal for a brave new world: The British Laboratory Animals Bureau and the Constitution of International Standards of Laboratory Animal Production and Use, circa 1947–1968.

Isis, 101(1):62–94, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
652689.

Sivasundaram:2010:I

[5] Sujit Sivasundaram. Introduction. Isis, 101(1):95–97, March 2010. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/652690.

Elshakry:2010:WSB

[6] Marwa Elshakry. When science became Western: Historiographical reflections. Isis, 101(1):98–109, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652691.

Tilley:2010:GHV

[7] Helen Tilley. Global histories, vernacular science, and African genealogies; or, is the history of science ready for the world? Isis, 101(1):110– 119, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652692.

Kapila:2010:ESI

[8] Shruti Kapila. The enchantment of science in India. Isis, 101(1):120– 132, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/652700.

Safier:2010:GKM

[9] Neil Safier. Global knowledge on the move: Itineraries, Amerindian narratives, and deep histories of science. Isis, 101(1):133–145, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/652693.

Sivasundaram:2010:SGM

[10] Sujit Sivasundaram. Sciences and the global: On methods, questions, and theory. Isis, 101(1):146–158, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/652694.

Chang:2010:EAG

[11] Ku-Ming “Kevin” Chang and Karen Hunger Parshall. Eloge: Allen George Debus, 16 August 1926–6 March 2009. Isis, 101(1):159–162, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/652695.

Kox:2010:EMJ

[12] A. J. Kox. Eloge: Martin Jesse Klein, 25 June 1924–28 March 2009. Isis, 101(1):163–165, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653843.

James:2010:ERH

[13] Frank A. J. L. James. Eloge: A. Rupert Hall, 1920–2009, and Marie Boas Hall, 1919–2009. Isis, 101(1):165–171, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/653844.

Wulff:2010:LE

[14] Karl Wulff. Letters to the Editor. Isis, 101(1):172, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652696.

Huff:2010:LE

[15] Toby E. Huff. Letters to the Editor. Isis, 101(1):172–173, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653845.

Kubbinga:2010:LE

[16] Henk Kubbinga. Letters to the Editor. Isis, 101(1):173, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653846.

vanBerkel:2010:LE

[17] Klaas van Berkel. Letters to the Editor. Isis, 101(1):173–174, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/653847.

Anonymous:2010:Ea

[18] Anonymous. Erratum. Isis, 101(1):174, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/653848.

Feingold:2010:ERW

[19] Mordechai Feingold. Essay review: The war on Newton: J. B. Shank: The Newton Wars and the Beginning of the Enlightenment. Isis, 101 (1):175–186, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
652697.

Worboys:2010:BRP

[20] Michael Worboys. Book review: Poonam Bala (Editor): Biomedicine as a Contested Site: Some Revelations in Imperial Contexts. Isis, 101(1): 187, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/653849.

Mirowski:2010:BRR

[21] Philip Mirowski. Book review: Richard Bronk: The Romantic Economist: Imagination in Economics. Isis, 101(1):187–189, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/653850.

Fayter:2010:BRJ

[22] Paul Fayter. Book review: Julian Chela-Flores: A Second Genesis: Stepping-Stones toward the Intelligibility of Nature. Adam Frank: The Constant Fire: Beyond the Science vs Religion Debate. Ariel A. Roth: Science Discovers God: Seven Convincing Lines of Evidence for His Existence. Isis, 101(1):189–190, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653851.

Smith:2010:BRT

[23] Justin E. H. Smith. Book review: Tobias Cheung: Res vivens: Agentenmodelle organischer Ordnung 1600–1800, (Berliner Kulturwissenschaft, 8.). Isis, 101(1):190–191, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/653852.

Buchwald:2010:BRB

[24] Jed Buchwald. Book review: Brian A. Curran; Anthony Grafton; Pamela O. Long; Benjamin Weiss: Obelisk: A History, (Burndy Library Publications, N.S., 2.). Isis, 101(1):191–192, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653853.

York:2010:BRJ

[25] William H. York. Book review: Jacalyn Duffin: Medical Miracles: Doctors, Saints, and Healing in the Modern World. Isis, 101(1):192–193, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/653854.

Mandelbrote:2010:BRP

[26] Scott Mandelbrote. Book review: Pascal Duris (Editor): Traduire la science: Hier et aujourd’hui. Isis, 101(1):194, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653855.

Kragh:2010:BRT

[27] Helge Kragh. Book review: Tofigh Heidarzadeh A History of Physical Theories of Comets, from Aristotle to Whipple (Archimedes: New Studies in the History and Philosophy of Science and Technology, 19.). Isis, 101(1):195–196, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653856.

Wittje:2010:BRK

[28] Roland Wittje. Book review: Klaus Hentschel (Editor): Unsichtbare Hände: Zur Rolle von Laborassistenten, Mechanikern Zeichnern und Amanuenses in der physikalischen Forschungs- und Entwicklungsarbeit. Isis, 101(1):196–197, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/653857.

Swerdlow:2010:BRT

[29] N. M. Swerdlow. Book review: Thomas Hockey (Editor): The Biographical Encyclopedia of Astronomers. Isis, 101(1):197–198, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653858.

Keller:2010:BRK

[30] Agathe Keller. Book review: Kim Plofker: Mathematics in India. Isis, 101(1):199–200, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653859.

Ghiselin:2010:BRM

[31] Michael T. Ghiselin. Book review: Michael Ruse; Joseph Travis (Editors): Evolution: The First Four Billion Years. Isis, 101(1):200–201, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/653860.

Proust:2010:BRE

[32] Christine Proust. Book review: Eleanor Robson: Mathematics in Ancient Iraq: A Social History. Isis, 101(1):201–202, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653861.

Steele:2010:BRG

[33] John Steele. Book review: Glen Van Brummelen: The Mathematics of the Heavens and the Earth: The Early History of Trigonometry. Isis, 101 (1):203, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/653862.

Sarnowsky:2010:BRJ

[34] Jürgen Sarnowsky. Book review: Joël Biard; Sabine Rommevaux (Editors): Mathématiques et théorie du mouvement (XIVe-XVIe siècles) (Histoires des Sciences). Isis, 101(1):204, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/653863.

Lagerlund:2010:BRG

[35] Henrik Lagerlund. Book review: Guillelmi de Conchis: Glosae super Platonem. Isis, 101(1):205, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/653864.

Gil-Sotres:2010:BRM

[36] Pedro Gil-Sotres. Book review: Marilyn Nicoud: Les régimes de santé au Moyen Âge: Naissance et diffusion d’une écriture médicale (XIIIe– XVe siècle). Isis, 101(1):205–207, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653865.

Walton:2010:BRJ

[37] Steven A. Walton. Book review: Jonathan Sawday: Engines of the Imagination: Renaissance Culture and the Rise of the Machine. Isis, 101(1):207–208, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653866.

Azzolini:2010:BRJ

[38] Monica Azzolini. Book review: Jérôme Torrella (Hieronymus Torrella): Opus praeclarum de imaginibus astrologicis. Isis, 101(1):208–209, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/653867.

Hall:2010:BRK

[39] Bert Hall. Book review: Konrad Gruter von Werden; Dietrich Lohrmann; Horst Kranz; Ulrich Alertz: De machinis et rebus mechanicis: Ein Maschinenbuch aus Italien für den König von Dänemark, 1393–1424, Volume 1: Einleitung. Isis, 101(1):209–211, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/653868.

Rabin:2010:BRL

[40] Sheila J. Rabin. Book review: Laurence Wuidar: Musique et astrologie après le concile de Trente. Isis, 101(1):211–212, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653869.

Webster:2010:BRR

[41] Charles Webster. Book review: Reid Barbour; Claire Preston (Editors): Sir Thomas Browne: The World Proposed. Isis, 101(1):212–213, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/653870.

Franklin:2010:BRJ

[42] James Franklin. Book review: Jacob Bernoulli: The Art of Conjecturing, together with Letter to a Friend on Sets in Court Tennis. Isis, 101(1):213– 214, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/653871.

Kagan:2010:BRD

[43] Richard L. Kagan. Book review: Daniela Bleichmar; Paula De Vos; Kristin Huffine; Kevin Sheehan (Editors): Science in the Spanish and Portuguese Empires, 1500–1800. Isis, 101(1):214–215, March 2010. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653872.

Kidwell:2010:BRB

[44] Peggy Aldrich Kidwell. Book review: Bruce Stanley Burdick: Mathematical Works Printed in the Americas, 1554–1700. (Johns Hopkins Studies in the History of Mathematics.). Isis, 101(1):215–216, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/653873.

Eskildsen:2010:BRS

[45] Kasper Risbjerg Eskildsen. Book review: Scholarly Knowledge: Textbooks in Early Modern Europe, by Emidio Campi; Simone De Angelis; Anja-Silvia Goeing; Anthony T. Grafton. (Travaux d’Humanisme et Renaissance, 447.). Isis, 101(1):216–217, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/653874.

teHeesen:2010:BRW

[46] Anke te Heesen. Book review: Die Welt in der Stube: Begegnungen mit Außereuropa in Kunstkammern der Frühen Neuzeit, by Dominik Collet. (Veröffentlichungen des Max-Planck-Instituts für Geschichte, 232.). Isis, 101(1):217–218, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653875.

Miller:2010:BRL

[47] David Philip Miller. Book review: The Language of Mineralogy: John Walker, Chemistry, and the Edinburgh Medical School, 1750–1800, by Matthew D. Eddy. (Science, Technology, and Culture, 1700–1945.). Isis, 101(1):218–219, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653876.

Mayer:2010:BRL

[48] Thomas F. Mayer. Book review: Lettera a Cristina di Lorena/Brief an Christine von Lothringen. Isis, 101(1):219–220, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653877.

Fara:2010:BRE

[49] Patricia Fara. Book review: The English Virtuoso: Art, Medicine, and Antiquarianism in the Age of Empiricism. Isis, 101(1):220–221, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/653878.

Vermij:2010:BRM

[50] Rienk Vermij. Book review: The Moon That Wasn’t: The Saga of Venus’ Spurious Satellite, by Helge Kragh. (Science Networks: Historical Studies, 37). Isis, 101(1):221–222, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653879.

Dooley:2010:BRT

[51] Brendan Dooley. Book review: Theatrum naturae: La ricerca naturalistica tra erudizione e nuova scienza nell’Italia del primo Seicento, by Alessandro Ottaviani; Oreste Trabucco. (Testi e Documenti per la Storia della Scienza nel Mezzogiorno, 4.). Isis, 101(1):222–223, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653880.

Buisseret:2010:BRI

[52] David Buisseret. Book review: The Imperial Map: Cartography and the Mastery of Empire. Isis, 101(1):223–224, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/653881.

Oldham:2010:BRG

[53] Kalil Oldham. Book review: “Gelehrte” Wissenschaft: Das Vorlesungprogramm der Universität Jena um 1800, by Thomas Bach; Jonas Maatsch; Ulrich Rasche. (Pallas Athene: Beiträge zur Universitäts- und Wissenschaftsgeschichte, 26.). Isis, 101(1):224–225, March 2010. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653882.

Richardson:2010:BRW

[54] Alan Richardson. Book review: William James at the Boundaries: Philosophy, Science, and the Geography of Knowledge. Isis, 101(1):225–227, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/653883.

Smocovitis:2010:BRC

[55] Vassiliki Betty Smocovitis. Book review: Charles Darwin: The Beagle Letters. Isis, 101(1):227–228, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653884.

Frappier:2010:BRK

[56] Mélanie Frappier. Book review: Kristian Camilleri: Heisenberg and the Interpretation of Quantum Mechanics: The Physicist as Philosopher. Isis, 101(1):228–229, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/653885.

VanRiper:2010:BRG

[57] A. Bowdoin Van Riper. Book review: God-or Gorilla: Images of Evolution in the Jazz Age, by Constance Areson Clark. (Medicine, Science, and Religion in Historical Context.). Isis, 101(1):229–230, March 2010.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653886.

Ciancio:2010:BRF

[58] Luca Ciancio. Book review: Fossils and Reputations: A Scientific Correspondence: Pisa, Paris, London, 1853–1857. Isis, 101(1):230–231, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/653887.

Jacyna:2010:BRO

[59] Stephen Jacyna. Book review: Owen’s Ape and Darwin’s Bulldog: Beyond Darwinism and Creationism. Isis, 101(1):231–232, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653888.

Stern:2010:BRS

[60] Alexandra Minna Stern. Book review: Segregation’s Science: Eugenics and Society in Virginia, by Gregory Michael Dorr. (Carter G. Woodson Institute Series.). Isis, 101(1):232–233, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/653889.

Engstrom:2010:BRJ

[61] Eric J. Engstrom. Book review: Jüdische Hochschullehrer an preußischen Universitäten (1870–1924): Eine quantitative Untersuchung mit biografischen Skizzen. Isis, 101(1):233–234, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/653890.

Gray:2010:BRN

[62] Jeremy Gray. Book review: Naming Infinity: The True Story of Religious Mysticism and Mathematical Certainty. Isis, 101(1):234–235, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/653891.

Shaffer:2010:BRS

[63] Elinor Shaffer. Book review: Science and Sensation in Romantic Poetry, by Noel Jackson. (Cambridge Studies in Romanticism, 73.). Isis, 101 (1):235–236, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653892.

VanBrummelen:2010:BRT

[64] Glen Van Brummelen. Book review: Tools of American Mathematics Teaching, 1800–2000, by Peggy Aldrich Kidwell; Amy Ackerberg-Hastings; David Lindsay Roberts. (Johns Hopkins Studies in the History of Mathematics.). Isis, 101(1):236–237, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/653893.

Uebel:2010:BRO

[65] Thomas Uebel. Book review: Otto Neurath: Museum and Exhibition Work: Spaces (Designed) for Communication. Isis, 101(1):237–238, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/653894.

Martin:2010:BRF

[66] Aryn Martin. Book review: Flesh and Blood: Organ Transplantation and Blood Transfusion in Twentieth–Century America. Isis, 101(1):238– 239, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/653895.

Gliboff:2010:BRE

[67] Sander Gliboff. Book review: Die Entstehung der biologischen Evolutionstheorie. Isis, 101(1):239–240, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653896.

Hentschel:2010:BRS

[68] Klaus Hentschel. Book review: The Shadow of Enlightenment: Optical and Political Transparency in France, 1789–1848. Isis, 101(1):240–241, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/653897.

Spanagel:2010:BRS

[69] David Spanagel. Book review: Scientists and Swindlers: Consulting on Coal and Oil in America, 1820–1890, by Paul Lucier. (Johns Hopkins Studies in the History of Technology.). Isis, 101(1):241–243, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653898.

Turner:2010:BRH

[70] R. Steven Turner. Book review: Hermann von Helmholtz’s Mechanism: The Loss of Certainty, by Gregor Schiemann. (Archimedes: New Studies in the History and Philosophy of Science and Technology.). Isis, 101 (1):243–244, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653899.

Allen:2010:BRE

[71] Garland E. Allen. Book review: The Emergence of Genetic Rationality: Space, Time, and Information in American Biological Science, 1870–1920. Isis, 101(1):244–246, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653900.

Green:2010:BRC

[72] Christopher Green. Book review: The Cattell Controversy: Race, Science, and Ideology. Isis, 101(1):246–247, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/653901.

Gavroglu:2010:BRF

[73] Kostas Gavroglu. Book review: Freezing Physics: Heike Kamerlingh Onnes and the Quest for Cold, by Dirk van Delft. (History of Science and Scholarship in The Netherlands, 10.). Isis, 101(1):247–249, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/653902.

Dawson:2010:BRE

[74] Gowan Dawson. Book review: The Education of the Eye: History of the Royal Polytechnic Institution, 1838–1881, by Brenda Weeden. (History of the University of Westminster, 1.). Isis, 101(1):249, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653903.

Segre:2010:BRG

[75] Michael Segre. Book review: Galileo y el Vaticano: Historia de la Comisión Pontificia de Estudio del Caso Galileo (1981–1992). Isis, 101 (1):249–250, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653904.

Lucier:2010:BRB

[76] Paul Lucier. Book review: Bigfoot: The Life and Times of a Legend. Isis, 101(1):250–251, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653905.

Downey:2010:BRI

[77] Greg Downey. Book review: Internet Alley: High Technology in Tysons Corner, 1945–2005, by Paul E. Ceruzzi. (Lemelson Center Studies in Invention and Innovation.). Isis, 101(1):251–252, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653906.

Abbate:2010:BRI

[78] Janet Abbate. Book review: On the Internet, by Hubert L. Dreyfus. (Thinking in Action.). Isis, 101(1):252–253, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/653907.

Alagona:2010:BRJ

[79] Peter S. Alagona. Book review: James Lovelock: In Search of Gaia. Isis, 101(1):253–254, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653908.

Solovey:2010:BRA

[80] Mark Solovey. Book review: The Americanization of Social Science: Intellectuals and Public Responsibility in the Postwar United States. Isis, 101(1):254–255, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653909.

Friedman:2010:BRW

[81] Robert Marc Friedman. Book review: Weather by the Numbers: The Genesis of Modern Meteorology. Isis, 101(1):255–257, March 2010. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653910.

Palladino:2010:BRC

[82] Paolo Palladino. Book review: Cultivating Science, Harvesting Power: Science and Industrial Agriculture in California, by Christopher R. Henke. (Inside Technology.). Isis, 101(1):257–258, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653911.

Dyck:2010:BRH

[83] Erika Dyck. Book review: Happy Pills in America: From Miltown to Prozac. Isis, 101(1):258–259, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653912.

Jones-Imhotep:2010:BRF

[84] Edward Jones-Imhotep. Book review: Fermilab: Physics, the Frontier, and Megascience. Isis, 101(1):259–260, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/653913.

Strasser:2010:BRT

[85] Bruno J. Strasser. Book review: Truth Machine: The Contentious History of DNA Fingerprinting. Isis, 101(1):260–261, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653914.

Feke:2010:BRD

[86] Jacqueline Feke. Book review: Decoding the Heavens: A Two-Thousand-Year-Old Computer-and the Century-Long Search to Discover Its Secrets. Isis, 101(1):261–262, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/653915.

Nye:2010:BRT

[87] Mary Jo Nye. Book review: The Two Cultures Controversy: Science, Literature, and Cultural Politics in Postwar Britain. Isis, 101(1):262– 263, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/653916.

Crease:2010:BRJ

[88] Robert P. Crease. Book review: Joanna S. Ploeger, The Boundaries of the New Frontier: Rhetoric and Communication at Fermi National Accelerator Laboratory (Studies in Rhetoric/Communication). Isis, 101 (1):263–264, March 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.
1086/653917;   http://www.jstor.org/stable/pdfplus/10.1086/
653917.pdf.

Barberis:2010:BRA

[89] Daniela S. Barberis. Book review: Anthropologies rédemptrices: Le monde selon Lévi–Strauss, by Wiktor Stoczkowski. (Collection Société et Pensées). Isis, 101(1):264–266, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653918.

Segal:2010:BRA

[90] Judy Z. Segal. Book review: The Age of Anxiety: A History of America’s Turbulent Affair with Tranquilizers. Isis, 101(1):266–267, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653919.

Haigh:2010:BRC

[91] Thomas Haigh. Book review: From Counterculture to Cyberculture: Stewart Brand, the Whole Earth Network, and the Rise of Digital Utopianism. Isis, 101(1):267–268, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653920.

DeVorkin:2010:BRS

[92] David H. DeVorkin. Book review: In Sputnik’s Shadow: The President’s Science Advisory Committee and Cold War America. Isis, 101(1):268– 270, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/653921.

Raz-Krakotzkin:2010:BRT

[93] Amnon Raz-Krakotzkin. Book review: Thinking Impossibilities: The Intellectual Legacy of Amos Funkenstein. Isis, 101(1):270–271, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/653922.

Tuana:2010:BRS

[94] Nancy Tuana. Book review: Sciences from Below: Feminisms, Postcolonialities, and Modernities. Isis, 101(1):271–272, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653923.

Shapiro:2010:BRN

[95] Alan E. Shapiro. Book review: Newton as Philosopher. Isis, 101(1):272– 273, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/653924.

Magnani:2010:BRM

[96] Lorenzo Magnani. Book review: The Meaning of the Body: Aesthetics of Human Understanding. Isis, 101(1):273–274, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653925.

Nickles:2010:BRB

[97] Thomas Nickles. Book review: Borrowed Knowledge and the Challenge of Learning across Disciplines: The Case of Chaos Theory. Isis, 101(1):274– 276, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/653926.

Zammito:2010:BRB

[98] John H. Zammito. Book review: Being Human: Historical Knowledge and the Creation of Human Nature. Isis, 101(1):276–277, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653927.

Anonymous:2010:NCa

[99] Anonymous. Notes on contributors. Isis, 101(1):278–283, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/653928.

Anonymous:2010:Fa

[100] Anonymous. Frontispiece. Isis, 101(1):vii, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/653929.

Clarke:2010:PSP

[101] Sabine Clarke. Pure science with a practical aim: The meanings of fundamental research in Britain, circa 1916–1950. Isis, 101(2):285–311, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/653094.

Doroshow:2010:ASB

[102] Deborah Blythe Doroshow. An alarming solution: Bedwetting, medicine, and behavioral conditioning in Mid-Twentieth-Century America. Isis, 101(2):312–337, June 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
653095.

Rossi:2010:FAM

[103] Michael Rossi. Fabricating authenticity: Modeling a whale at the American Museum of Natural History, 1906–1974. Isis, 101(2):338–361, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/653096.

Heyck:2010:HHD

[104] Hunter Heyck and David Kaiser. Hunter Heyck and David Kaiser: Introduction. Isis, 101(2):362–366, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653097.

Wang:2010:TSD

[105] Zuoyue Wang. Transnational science during the Cold War: The case of Chinese/American scientists. Isis, 101(2):367–377, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653098.

Macrakis:2010:THE

[106] Kristie Macrakis. Technophilic hubris and espionage styles during the Cold War. Isis, 101(2):378–385, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/653104.

Erickson:2010:MMR

[107] Paul Erickson. Mathematical models, rational choice, and the search for Cold War culture. Isis, 101(2):386–392, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/653105.

Engerman:2010:SSC

[108] David C. Engerman. Social science in the Cold War. Isis, 101(2):393– 400, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/653106.

Lemov:2010:HMS

[109] Rebecca Lemov. “Hypothetical machines”: The science fiction dreams of Cold War social science. Isis, 101(2):401–411, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653107.

Lowy:2010:EOA

[110] Ilana Löwy. Eloge: Olga Amsterdamska, 1953–2009. Isis, 101(2):412– 415, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/653108.

Seidel:2010:BRA

[111] Robert W. Seidel. Book review: Accident Prone: A History of Technology, Psychology, and Misfits of the Machine Age. Isis, 101(2):416–417, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/655702.

Brockliss:2010:BRW

[112] Laurence Brockliss. Book review: Worlds Made by Words: Scholarship and Community in the Modern West. Isis, 101(2):417–418, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655703.

Stanley:2010:BRD

[113] Matthew Stanley. Book review: Discovering the Expanding Universe. Isis, 101(2):418, June 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
655704.

Ferraro:2010:BRL

[114] Giovanni Ferraro. Book review: Liber Amicorum Jean Dhombres. Isis, 101(2):418–420, June 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
655705.

Lennox:2010:BRA

[115] James G. Lennox. Book review: Aristotle on Life. Isis, 101(2):420– 421, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/655706.

Hahm:2010:BRM

[116] David E. Hahm. Book review: Manilius and His Intellectual Background. Isis, 101(2):421–422, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/655707.

vanderEijk:2010:BRA

[117] Philip van der Eijk. Book review: Asklepios, Medicine, and the Politics of Healing in Fifth-Century Greece: Between Craft and Cult. Isis, 101(2): 422–423, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/655708.

Cadden:2010:BRS

[118] Joan Cadden. Book review: Science Translated: Latin and Vernacular Translations of Scientific Treatises in Medieval Europe. Isis, 101(2):423– 424, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/655709.

Sylla:2010:BRA

[119] Edith Sylla. Book review: Atomism in Late Medieval Philosophy and Theology. Isis, 101(2):424–425, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655710.

Hayton:2010:BRA

[120] Darin Hayton. Book review: Astrologie und Öffentlichkeit im Mittelalter. Isis, 101(2):425–426, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/655711.

Cislo:2010:BRP

[121] Amy Eisen Cislo. Book review: Paracelsus: Medicine, Magic and Mission at the End of Time. Isis, 101(2):426–427, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/655712.

Hahn:2010:BRL

[122] Roger Hahn. Book review: L’enquête du Régent 1716–1718: Sciences, techniques et politique dans la France pré-industrielle. Isis, 101(2):427– 428, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/655713.

Eddy:2010:BRE

[123] M. E. Eddy. Book review: Essays on David Hume, Medical Men and the Scottish Enlightenment: “Industry, Knowledge and Humanity”, by Roger L. Emerson. Isis, 101(2):428–429, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655714.

Wakefield:2010:BRZ

[124] Andre Wakefield. Book review: Zwischen Fürstenwillkür und Menschheitswohl: Gottfried Wilhelm Leibniz als Bibliothekar. Isis, 101(2): 429–430, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/655715.

Grave:2010:BRS

[125] Patricia Radelet de Grave. Book review: Sämtliche Schriften und Briefe. Isis, 101(2):430–431, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/655716.

Hankins:2010:BRQ

[126] Thomas L. Hankins. Book review: The Quest for the Invisible: Microscopy in the Enlightenment. Isis, 101(2):431–432, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/655717.

Caparrini:2010:BRE

[127] Sandro Caparrini. Book review: Euler as Physicist. Isis, 101(2):432– 433, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/655718.

Porter:2010:BRD

[128] Theodore M. Porter. Book review: The Disordered Police State: German Cameralism as Science and Practice. Isis, 101(2):433–434, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655719.

Carhart:2010:BRB

[129] Michael C. Carhart. Book review: In Babel’s Shadow: Language, Philology and the Nation in Nineteenth-Century Germany. Isis, 101(2):434– 436, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/655720.

Brain:2010:BRD

[130] Robert Brain. Book review: The Dada Cyborg: Visions of the New Human in Weimar Berlin. Isis, 101(2):436–437, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/655721.

Knight:2010:BRS

[131] David Knight. Book review: Science For All: The Popularization of Science in Early Twentieth–Century Britain. Isis, 101(2):437–438, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/655722.

Gordin:2010:BRV

[132] Michael D. Gordin. Book review: Voluntary Associations in Tsarist Russia: Science, Patriotism, and Civil Society. Isis, 101(2):438–439, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/655723.

Hentschel:2010:BRN

[133] Klaus Hentschel. Book review: In the Name of Science? P, by Martijn Eickhoff. J. W. Debye and his Career in Nazi Germany. Isis, 101(2):439– 440, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/655724.

Henke:2010:BRG

[134] Christopher R. Henke. Book review: Growing American Rubber: Strategic Plants and the Politics of National Security. Isis, 101(2):440, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/655725.

Hempstead:2010:BRC

[135] Colin A. Hempstead. Book review: Communications Under the Sea: The Evolving Cable Network and Its Implications. Isis, 101(2):440–441, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/655726.

Ruiz-Castell:2010:BRR

[136] Pedro Ruiz-Castell. Book review: La Real Academia de Ciencias de la Habana, 1861–1898. Isis, 101(2):441–442, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/655727.

Brock:2010:BRF

[137] William H. Brock. Book review: A Focus of Discoveries. Isis, 101(2): 442–443, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/655728.

Gates:2010:BRR

[138] Barbara T. Gates. Book review: Reframing Darwin: Evolution and Art in Australia. Isis, 101(2):443–444, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655729.

Fichman:2010:BRD

[139] Martin Fichman. Book review: Darwin’s Armada: Four Voyages and the Battle for the Theory of Evolution. Isis, 101(2):444–445, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655730.

Guenther:2010:BRU

[140] Katja Guenther. Book review: Umwelt als Vision: Leben und Werk Jakob von Uexkülls (1864–1944). Isis, 101(2):445–446, June 2010. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655731.

McTavish:2010:BRI

[141] Lianne McTavish. Book review: Icons of Life: A Cultural History of Human Embryos. Isis, 101(2):446–447, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655732.

Cravens:2010:BRW

[142] Hamilton Cravens. Book review: War and Disease: Biomedical Research on Malaria in the Twentieth Century. Isis, 101(2):447–448, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655733.

Misa:2010:BRG

[143] Thomas J. Misa. Book review: Grace Hopper and the Invention of the Information Age. Isis, 101(2):448–449, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655734.

Rosser:2010:BRL

[144] Sue V. Rosser. Book review: Lives in Science: How Institutions Affect Academic Careers. Isis, 101(2):449–450, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/655735.

Stanley:2010:BRM

[145] Heather Stanley. Book review: Make Room for Daddy: The Journey from Waiting Room to Birthing Room. Isis, 101(2):450–451, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655736.

Yans:2010:BRM

[146] Virginia Yans. Book review: Margaret Mead: The Making of an American Icon. Isis, 101(2):451–452, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655737.

Neufeld:2010:BRT

[147] Michael J. Neufeld. Book review: Truth, Lies, and O-Rings: Inside the Space Shuttle Challenger Disaster. Isis, 101(2):452–453, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655738.

Thurs:2010:BRN

[148] Daniel Patrick Thurs. Book review: Nanovision: Engineering the Future. Isis, 101(2):453–454, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/655739.

Schatzberg:2010:BRC

[149] Eric Schatzberg. Book review: Cold War Kitchen: Americanization, Technology, and European Users. Isis, 101(2):454–455, June 2010. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655740.

Sepkoski:2010:BRS

[150] David Sepkoski. Book review: Stephen Jay Gould and the Politics of Evolution. Isis, 101(2):455–456, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655741.

Shen:2010:BRP

[151] Grace Shen. Book review: The People’s Peking Man: Popular Science and Human Identity in Twentieth–Century China. Isis, 101(2):456–457, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/655742.

Dyer:2010:BRA

[152] Ruthann Dyer. Book review: Animal Capital: Rendering Life in Biopolitical Times. Isis, 101(2):457–458, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655743.

Bontems:2010:BRV

[153] Vincent K. Bontems. Book review: Les vertiges de la technoscience: Façonner le monde atome par atome. Isis, 101(2):458–459, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655744.

Garvey:2010:BRC

[154] Brian Garvey. Book review: The Cambridge Companion to the Philosophy of Biology. Isis, 101(2):459–460, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655745.

Pinch:2010:BRM

[155] Trevor Pinch. Book review: The Mangle in Practice: Science, Society, and Becoming. Isis, 101(2):460–462, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655746.

Anonymous:2010:NCb

[156] Anonymous. Notes on contributors. Isis, 101(2):463–466, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655747.

Anonymous:2010:Fb

[157] Anonymous. Frontispiece. Isis, 101(2):vi, June 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/655748.

dellaDora:2010:MMK

[158] Veronica della Dora. Making mobile knowledges: The educational cruises of the Revue Générale des Sciences Pures et Appliquées, 1897–1914. Isis, 101(3):467–500, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/655789.

Pesic:2010:HIM

[159] Peter Pesic. Hearing the irrational: Music and the development of the modern concept of number. Isis, 101(3):501–530, September 2010. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655790.

Sokal:2010:SBC

[160] Michael M. Sokal. Scientific biography, cognitive deficits, and laboratory practice: James McKeen Cattell and early American experimental psychology, 1880–1904. Isis, 101(3):531–554, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/655791.

Bono:2010:MKH

[161] James J. Bono. Making knowledge: History, literature, and the poetics of science. Isis, 101(3):555–559, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655792.

Milburn:2010:MFS

[162] Colin Milburn. Modifiable futures: Science fiction at the bench. Isis, 101 (3):560–569, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
655793.

Otis:2010:SSH

[163] Laura Otis. Science surveys and histories of literature: Reflections on an uneasy kinship. Isis, 101(3):570–577, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/655794.

Turner:2010:LLH

[164] Henry S. Turner. Lessons from literature for the historian of science (and vice versa): Reflections on “form”. Isis, 101(3):578–589, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/655795.

Walls:2010:AOC

[165] Laura Dassow Walls. Of atoms, oaks, and cannibals; or, more things that talk. Isis, 101(3):590–598, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/655796.

Moore:2010:ERC

[166] James Moore. Eloge: Ralph Colp, 1924–2008. Isis, 101(3):599–602, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/655798.

Kubbinga:2010:EKF

[167] Henk Kubbinga. Eloge: Karin Figala, 1938–2007. Isis, 101(3):603–606, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/657171.

Tweney:2010:EDC

[168] Ryan D. Tweney. Eloge: David Charles Gooding, 1947–2009. Isis, 101 (3):607–609, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657172.

Anonymous:2010:HSS

[169] Anonymous. 2009 History of Science Society prize citations. Isis, 101 (3):610–614, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657173.

Anonymous:2010:Eb

[170] Anonymous. Errata. Isis, 101(3):614, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/657174.

DiLiscia:2010:BRG

[171] Daniel A. Di Liscia. Book review: La guerre des astronomes: La querelle au sujet de l’origine du système géo-hèliocentrique à la fin du XVIe siècle. Isis, 101(3):615–617, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/655799.

Dupre:2010:BRH

[172] Sven Dupré. Book review: The Human Animal in Western Art and Science Leonardo Leonardo da Vinci: Experience, Experiment, and Design Seen — Unseen: Art, Science, and Intuition from Leonardo to the Hubble Space Telescope. Isis, 101(3):618–622, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657175.

Hacker:2010:BRS

[173] Bart Hacker. Book review: The Scientific Way of Warfare: Order and Chaos on the Battlefields of Modernity. Isis, 101(3):623–624, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/655800.

Hu:2010:BRF

[174] Minghui Hu. Book review: Fong-ching Chen, Heritage and Betrayal: A Treatise on the Emergence of Modern Science in Western Civilization. Isis, 101(3):624–625, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/657176.

Raphael:2010:BRH

[175] Dennis Raphael. Book review: History of the Social Determinants of Health: Global Histories, Contemporary Debates. Isis, 101(3):625–626, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/657177.

Gregory:2010:BRR

[176] Andrew Gregory. Book review: Robert P. Crease: The Great Equations: Breakthroughs in Science from Pythagoras to Heisenberg. Isis, 101(3): 626–627, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657178.

Fayter:2010:BRE

[177] Paul Fayter. Book review: The Extraterrestrial Life Debate, Antiquity to 1915: A Source Book. Isis, 101(3):627–628, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657179.

Falchetta:2010:BRG

[178] Piero Falchetta and Sara Scharf. Book review: La Géographie de Ptolémée en Occident (IVe-XVIe siècle). Isis, 101(3):628–629, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/657180.

Kidwell:2010:BRI

[179] Peggy Aldrich Kidwell. Book review: Ivor Grattan-Guinness, Routes of Learning: Highways, Pathways, and Byways in the History of Mathematics. Isis, 101(3):629–630, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657181.

OConnor:2010:BRG

[180] Ralph O’Connor. Book review: Geology and Religion: A History of Harmony and Hostility. Isis, 101(3):630–631, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657182.

Richmond:2010:BRN

[181] Jesse Richmond. Book review: A New History of Anthropology. Isis, 101(3):632, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657183.

Lazenby:2010:BRD

[182] Jill Lazenby. Book review: Donald S. Lopez, Jr.: Buddhism and Science: A Guide for the Perplexed. Isis, 101(3):632–633, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/657184.

Goenner:2010:BRA

[183] Hubert Goenner. Book review: Alberto A. Martínez: Kinematics: The Lost Origins of Einstein’s Relativity. Isis, 101(3):633–635, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/657185.

Hehmeyer:2010:BRD

[184] Ingrid Hehmeyer. Book review: David Pingree: Eastern Astrolabes. Isis, 101(3):635–636, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/657186.

Tauber:2010:BRA

[185] Alfred I. Tauber. Book review: Arthur M. Silverstein: A History of Immunology. Ed Cohen: A Body Worth Defending: Immunity, Biopolitics, and the Apotheosis of the Modern Body. Isis, 101(3):636–637, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/657187.

Agar:2010:BRE

[186] Jon Agar. Book review: Erwin Tomash; Michael R. Williams: The Erwin Tomash Library on the History of Computing: An Annotated and Illustrated Catalog. Isis, 101(3):637–639, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/657188.

Totelin:2010:BRP

[187] Laurence Totelin. Book review: The Poison King: The Life and Legend of Mithradates, Rome’s Deadliest Enemy. Isis, 101(3):639–640, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/657189.

Kosak:2010:BRH

[188] Jennifer Clarke Kosak. Book review: Hippocratic Recipes: Oral and Written Transmission of Pharmacological Knowledge in Fifth- and Fourth-Century Greece. Isis, 101(3):640–641, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657190.

Kassell:2010:BRP

[189] Lauren Kassell. Book review: Prophecy, Alchemy, and the End of Time: John of Rupescissa in the Late Middle Ages. Isis, 101(3):641–642, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/657191.

Navarro-Brotons:2010:BRG

[190] Víctor Navarro-Brotons. Book review: Gemma Frisius, arpenteur de la terre et du ciel. Isis, 101(3):642–643, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/657192.

Folkerts:2010:BRF

[191] Menso Folkerts. Book review: Fra Giocondo et les textes français de géométrie pratique. Isis, 101(3):643–644, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/657193.

Raphael:2010:BRS

[192] Renée Raphael. Book review: Sidereus nuncius; or, A Sidereal Message. Isis, 101(3):644–645, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/657194.

Lefevre:2010:BRC

[193] Wolfgang Lefèvre. Book review: Compass and Rule: Architecture as Mathematical Practice in England, 1500–1750. Isis, 101(3):645–646, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/657195.

Crowther:2010:BRF

[194] Kathleen Crowther. Book review: The Fall of Man and the Foundations of Science. Isis, 101(3):646–647, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657196.

Waddell:2010:BRP

[195] Mark A. Waddell. Book review: Papers of the Dublin Philosophical Society, 1683–1709. Isis, 101(3):647–648, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657197.

Barrow:2010:BRU

[196] Mark V. Barrow, Jr. Book review: The Untilled Garden: Natural History and the Spirit of Conservation in America, 1740–1840. Isis, 101(3): 648–649, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657198.

Cottingham:2010:BRD

[197] John Cottingham. Book review: Descartes’s Changing Mind. Isis, 101 (3):649–650, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657199.

McClellan:2010:BRI

[198] James E. McClellan III. Book review: Impossible Engineering: Technology and Territoriality on the Canal du Midi. Isis, 101(3):651–652, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/657200.

Finocchiaro:2010:BRD

[199] Maurice A. Finocchiaro. Book review: I documenti vaticani del processo di Galileo Galilei (1611–1741). Isis, 101(3):652–653, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/657201.

Keller:2010:BRR

[200] Eve Keller. Book review: Ryan J. Stark. Rhetoric, Science, and Magic in Seventeenth-Century England. Isis, 101(3):653–654, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/657202.

Steigerwald:2010:BRG

[201] Joan Steigerwald. Book review: The Geographic Imagination of Modernity: Geography, Literature, and Philosophy in German Romanticism. Isis, 101(3):654–655, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/657203.

Benzaquen:2010:BRM

[202] Adriana S. Benzaquén. Book review: The Marvelous Hairy Girls: The Gonzales Sisters and Their Worlds. Isis, 101(3):655–656, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/657204.

Emerson:2010:BRS

[203] Roger Emerson. Book review: Seeking Nature’s Logic: Natural Philosophy in the Scottish Enlightenment. Isis, 101(3):656–657, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/657205.

Connor:2010:BRH

[204] J. T. H. Connor. Book review: Health and Medicine on Display: International Expositions in the United States, 1876–1904. Isis, 101(3):657–658, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/657206.

Corsi:2010:BRC

[205] Pietro Corsi. Book review: Le colonne del tempo: Il “tempio di Serapide” a Pozzuoli nella storia della geologia, dell’archeologia e dell’arte (1750– 1900). Isis, 101(3):658–660, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657207.

Francis:2010:BRD

[206] Mark Francis. Book review: Darwin’s Coat-Tails: Essays on Social Darwinism. Isis, 101(3):660, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657208.

Cassidy:2010:BRG

[207] David C. Cassidy. Book review: Graham Farmelo, The Strangest Man: The Hidden Life of Paul Dirac, Mystic of the Atom. Isis, 101(3):661, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/657209.

Seth:2010:BRN

[208] Suman Seth. Book review: Naturlehre: Die Entwicklung des Physikunterrichts an den Volksschulen der Rheinprovinz (1815–1968). Isis, 101 (3):662, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657210.

Aronova:2010:BRH

[209] Elena Aronova and Naomi Oreskes. Book review: History of Russian Underwater Acoustics. Isis, 101(3):662–663, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657211.

Cantor:2010:BRC

[210] Geoffrey Cantor. Book review: The Correspondence of Michael Faraday, by Frank A. J. L. James. Volume 5: November 1855–October 1860, Letters 3033-3873. Isis, 101(3):664, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657212.

Monaldi:2010:BRE

[211] Daniela Monaldi. Book review: Einstein parla italiano: Itinerari e polemiche. Isis, 101(3):665, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657213.

Daum:2010:BRM

[212] Andreas W. Daum. Book review: Modern Nature: The Rise of the Biological Perspective in Germany. Isis, 101(3):665–667, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/657214.

Bensaude-Vincent:2010:BRP

[213] Bernadette Bensaude-Vincent. Book review: Popularizing Science and Technology in the European Periphery, 1800–2000. Isis, 101(3):667–668, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/657215.

Boudia:2010:BRT

[214] Soraya Boudia. Book review: Trafficking Materials and Gendered Experimental Practices: Radium Research in Early Twentieth Century Vienna. Isis, 101(3):668–669, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/657216.

Alborn:2010:BRC

[215] Timothy Alborn. Book review: The Cost of Living in America: A Political History of Economic Statistics, 1880–2000. Isis, 101(3):669–670, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/657217.

Pursell:2010:BRS

[216] Carroll Pursell. Book review: Structures of Change in the Mechanical Age: Technological Innovation in the United States, 1790–1865. Isis, 101 (3):670–671, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657218.

Thomas:2010:BRS

[217] William Thomas. Book review: The Science of Bombing: Operational Research in RAF Bomber Command. Isis, 101(3):671–672, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/657219.

Turner:2010:BRA

[218] Roger Turner. Book review: Atmospheric Science at NASA: A History. Isis, 101(3):672–673, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/657220.

Smith:2010:BRA

[219] Robert W. Smith. Book review: The Adaptive Optics Revolution: A History. Isis, 101(3):673–674, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657221.

Ensmenger:2010:BRC

[220] Nathan Ensmenger. Book review: The Cultural Logic of Computation. Isis, 101(3):674–675, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/657222.

Ortolano:2010:BRT

[221] Guy Ortolano. Book review: The Three Cultures: Natural Sciences, Social Sciences, and the Humanities in the Twenty-first Century. Isis, 101 (3):675–676, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657223.

Gerson:2010:BRH

[222] Elihu M. Gerson. Book review: How Professors Think: Inside the Curious World of Academic Judgment. Isis, 101(3):676–677, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/657224.

Lewis:2010:BRD

[223] John W. Lewis and Xue Litai. Book review: A Draft History of Space Technology in China [Zhongguo hangtian jishu fazhan shi gao]. Isis, 101 (3):677–678, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657225.

Marchand:2010:BRR

[224] Suzanne Marchand. Book review: Race and Erudition. Isis, 101(3): 678–679, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657226.

Cambrosio:2010:BRW

[225] Alberto Cambrosio. Book review: When Experiments Travel: Clinical Trials and the Global Search for Human Subjects. Isis, 101(3):679–680, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/657227.

Tracy:2010:BRS

[226] Sarah W. Tracy. Book review: On Speed: The Many Lives of Amphetamine. Isis, 101(3):681–682, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657228.

Breyman:2010:BRF

[227] Steve Breyman. Book review: Far-Fetched Facts: A Parable of Development Aid. Isis, 101(3):682–683, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657229.

Morawski:2010:BRB

[228] Jill G. Morawski. Book review: Beyond the Box: B, by Alexandra Rutherford. F. Skinner’s Technology of Behavior from Laboratory to Life, 1950s-1970s. Isis, 101(3):683–684, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657230.

Brysse:2010:BRP

[229] Keynyn Brysse. Book review: The Paleobiological Revolution: Essays on the Growth of Modern Paleontology. Isis, 101(3):684–685, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/657231.

Westad:2010:BRS

[230] Odd Arne Westad. Book review: Sulian ji shu xiang Zhongguo de zhuan yi: 1949–1966 [Technology Transfer from the Soviet Union to the People’s Republic of China, 1949–1966]. Isis, 101(3):685–686, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/657232.

Brown:2010:BRS

[231] Matthew J. Brown. Book review: Science in Democracy: Expertise, Institutions, and Representation. Isis, 101(3):686–687, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/657233.

Park:2010:BRE

[232] Buhm Soon Park. Book review: An Element of Controversy: The Life of Chlorine in Science, Medicine, Technology, and War. Isis, 101(3): 687–688, September 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657234.

deChadarevian:2010:BRV

[233] Soraya de Chadarevian. Book review: Velvet Revolution at the Synchrotron: Biology, Physics, and Change in Science. Isis, 101(3):688–689, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/657235.

Pickstone:2010:BRI

[234] John Pickstone. Book review: Insatiable Curiosity: Innovation in a Fragile Future. Isis, 101(3):689–690, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657236.

Anonymous:2010:NCc

[235] Anonymous. Notes on contributors. Isis, 101(3):691–695, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/657237.

Anonymous:2010:Fc

[236] Anonymous. Frontispiece. Isis, 101(3):vi, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657170.

Green:2010:SOB

[237] Christopher D. Green. Scientific objectivity and E. B. Titchener’s experimental psychology. Isis, 101(4):697–721, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657473.

Weiss:2010:AFP

[238] Sheila Faith Weiss. After the fall: Political whitewashing, professional posturing, and personal refashioning in the postwar career of otmar freiherr von verschuer. Isis, 101(4):722–758, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/657474.

Dear:2010:DBS

[239] Peter Dear and Sheila Jasanoff. Dismantling boundaries in science and technology studies. Isis, 101(4):759–774, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/657475.

Morus:2010:PP

[240] Iwan Rhys Morus. Placing performance. Isis, 101(4):775–778, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/657476.

Wintroub:2010:TBT

[241] Michael Wintroub. Taking a bow in the theater of things. Isis, 101 (4):779–793, December 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
657477.

Heering:2010:EGD

[242] Peter Heering. An Experimenter’s gotta do what an Experimenter’s gotta do—but how? Isis, 101(4):794–805, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657478.

Morus:2010:WWS

[243] Iwan Rhys Morus. Worlds of wonder: Sensation and the Victorian scientific performance. Isis, 101(4):806–816, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/657479.

Myers:2010:PPR

[244] Natasha Myers. Pedagogy and performativity: Rendering laboratory lives in the documentary naturally obsessed: The making of a scientist. Isis, 101(4):817–828, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/657480.

Snedegar:2010:EJD

[245] Keith Snedegar. Eloge: John David North, 1934–2008. Isis, 101(4):829– 832, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659744.

Taylor:2010:ERR

[246] Kenneth L. Taylor and Alice Stroup. Eloge: Rhoda Rappaport, 1935– 2009. Isis, 101(4):833–837, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657514.

Gasman:2010:LE

[247] Daniel Gasman. Letter to the Editor. Isis, 101(4):838, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/657515.

Hopwood:2010:LE

[248] Nick Hopwood. Letter to the Editor. Isis, 101(4):838, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659658.

Heilbron:2010:BRN

[249] J. L. Heilbron. Book review: New Dictionary of Scientific Biography. Isis, 101(4):839–842, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/657516.

Hentschel:2010:ERH

[250] Klaus Hentschel. Essay review: Heisenberg contextualized: Cathryn Carson: Heisenberg in the Atomic Age: Science and the Public Sphere. Isis, 101(4):843–845, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/657517.

Tjossem:2010:BRN

[251] Sara Tjossem. Book review: Nature’s Ghosts: Confronting Extinction from the Age of Jefferson to the Age of Ecology. Isis, 101(4):846–847, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659659.

Kreager:2010:BRP

[252] Philip Kreager. Book review: Pour une sociologie historique de la quantification: L’argument statistique I Gouverner par les nombres: L’argument statistique II. Isis, 101(4):847–848, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659660.

Samson:2010:BRF

[253] Jane Samson. Book review: Foreign Bodies: Oceania and the Science of Race, 1750–1940. Isis, 101(4):848–849, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659661.

Brysse:2010:BRG

[254] Keynyn Brysse. Book review: A Guinea Pig’s History of Biology. Isis, 101(4):849–850, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/659662.

Somsen:2010:BRS

[255] Geert Somsen. Book review: In Splendid Isolation: A History of the Willie Commelin Scholten Phytopathology Laboratory, 1894–1992. Isis, 101(4):850–851, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/659663.

Olby:2010:BRG

[256] Robert Olby. Book review: Genetic Analysis: A History of Genetic Thinking. Isis, 101(4):851–852, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659664.

Bix:2010:BRC

[257] Amy Sue Bix. Book review: A Culture of Improvement: Technology and the Western Millennium. Isis, 101(4):852–854, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659665.

Werrett:2010:BRP

[258] Simon Werrett. Book review: Playing with Fire: Histories of the Lightning Rod. Isis, 101(4):854–855, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659666.

Gordin:2010:BRW

[259] Michael D. Gordin. Book review: Would Trotsky Wear a Bluetooth? Technological Utopianism under Socialism, 1917–1989. Isis, 101(4):855– 856, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659667.

Snobelen:2010:BRG

[260] Stephen D. Snobelen. Book review: Galileo Goes to Jail and Other Myths about Science and Religion. Isis, 101(4):856–857, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659668.

Dauben:2010:BRO

[261] Joseph W. Dauben. Book review: The Oxford Handbook of the History of Mathematics. Isis, 101(4):857–859, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659669.

Baumgartner:2010:BRC

[262] Ingrid Baumgärtner. Book review: Cartography in Antiquity and the Middle Ages: Fresh Perspectives, New Methods. Isis, 101(4):859–861, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659670.

Killeen:2010:BRN

[263] Kevin Killeen. Book review: Nature and Scripture in the Abrahamic Religions: Up to 1700 Nature and Scripture in the Abrahamic Religions: 1700–Present. Isis, 101(4):861–864, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659671.

Lloyd:2010:BRN

[264] G. E. R. Lloyd. Book review: Numerical Notation: A Comparative History. Isis, 101(4):864–865, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659672.

Leven:2010:BRM

[265] Karl-Heinz Leven. Book review: Medicine and Health Care in Early Christianity. Isis, 101(4):865–866, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/659673.

Armstrong:2010:BRK

[266] Richard H. Armstrong. Book review: Knossos and the Prophets of Modernism. Isis, 101(4):866–867, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659674.

Ragep:2010:BRF

[267] F. Jamil Ragep. Book review: Frank Griffel. Al-Ghazālῑ’s Philosophical Theology. Isis, 101(4):867–868, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659675.

Knobloch:2010:BRP

[268] Eberhard Knobloch. Book review: Pythagoras and Renaissance Europe: Finding Heaven. Isis, 101(4):869, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659676.

Seitz:2010:BRH

[269] Jonathan Seitz. Book review: The Hammer of Witches: A Complete Translation of the Malleus maleficarum. Isis, 101(4):870–871, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/659677.

Cavazza:2010:BRA

[270] Marta Cavazza. Book review: The Accademia del Cimento and Its European Context. Isis, 101(4):871–872, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/659678.

Garber:2010:BRN

[271] Daniel Garber. Book review: Natural Law and Laws of Nature in Early Modern Europe: Jurisprudence, Theology, Moral and Natural Philosophy. Isis, 101(4):872–873, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659679.

Robin:2010:BRW

[272] Nicolas Robin. Book review: Wissen im Netz: Botanik und Pflanzentransfer in europäischen Korrespondenznetzen des 18, by Regina Dauser; Stefan Hächler; Michale Kempe; Franz Mauelshagen; Martin Stuber. Jahrhunderts. Isis, 101(4):873–875, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/659680.

Boantza:2010:BRP

[273] Victor Boantza. Book review: La philosophie naturelle de Robert Boyle. Isis, 101(4):875, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/659681.

Laird:2010:BRM

[274] Roy Laird. Book review: Une mécanique donnée à voir: Les théses illustrées défendues à Louvain en julliet 1624 par Grégoire de Saint-Vincent S, by Jean Dhombres; Patricia Radelet-de Grave.J. Isis, 101 (4):875–877, December 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
659682.

Levere:2010:BRA

[275] Trevor Levere. Book review: The Age of Wonder: How the Romantic Generation Discovered the Beauty and Terror of Science. Isis, 101(4): 877–878, December 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
659683.

Furth:2010:BRS

[276] Charlotte Furth. Book review: Sojourners in a Strange Land: Jesuits and Their Scientific Missions in Late Imperial China. Isis, 101(4):878–879, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659684.

Matthews:2010:BRB

[277] Steven Matthews. Book review: Biblical Scholarship, Science, and Politics in Early Modern England: Thomas Browne and the Thorny Place of Knowledge. Isis, 101(4):880–881, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659685.

Rickman:2010:BRB

[278] Melissa Rickman. Book review: Of Books and Botany in Early Modern England: Sixteenth-Century Plants and Print Culture. Isis, 101(4):881– 882, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659686.

Mauskopf:2010:BRJ

[279] Seymour Mauskopf. Book review: James Watt, Chemist: Understanding the Origins of the Steam Age. Isis, 101(4):882–883, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659687.

Fan:2010:BRM

[280] Fa ti Fan. Book review: The Monkey and the Inkpot: Natural History and Its Transformations in Early Modern China. Isis, 101(4):883–884, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659688.

Teo:2010:BRV

[281] Thomas Teo. Book review: Varietäten des Menschengeschlechts: Die Wissenschaften vom Menschen in der Zeit der Aufklärung. Isis, 101 (4):884–885, December 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
659689.

Barrera-Osorio:2010:BRS

[282] Antonio Barrera-Osorio. Book review: Secret Science: Spanish Cosmography and the New World. Isis, 101(4):885–886, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659690.

Outram:2010:BRB

[283] Dorinda Outram. Book review: The Brokered World: Go-Betweens and Global Intelligence, 1770–1820. Isis, 101(4):886–888, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659691.

Gray:2010:BRF

[284] Jeremy Gray. Book review: Flatland: An Edition with Notes and Commentary Flatland: A Romance of Many Dimensions. Isis, 101(4):888– 889, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659692.

Barnett:2010:BRR

[285] Richard Barnett. Book review: Regulated Lives: Life Insurance and British Society, 1800–1914. Isis, 101(4):889–890, December 2010. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659693.

Brock:2010:BRN

[286] Emily K. Brock. Book review: The Nature Study Movement: The Forgotten Popularizer of America’s Conservation Ethic. Isis, 101(4):890–891, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659694.

Naldi:2010:BRP

[287] Nerio Naldi. Book review: The Provocative Joan Robinson: The Making of a Cambridge Economist. Isis, 101(4):891, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659695.

Staley:2010:BRM

[288] Richard Staley. Book review: Max Planck und die Max-Planck-Gesellschaft: Zum 150, by Lorenz Friedrich Beck. Geburtstag am 23. April 2008 aus den Quellen zusammengestellt. Isis, 101(4):891–892, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/659696.

Lucier:2010:BRN

[289] Paul Lucier. Book review: Notes from the Ground: Science, Soil, and Society in the American Countryside. Isis, 101(4):892–893, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/659697.

Kojevnikov:2010:BRT

[290] Alexei Kojevnikov. Book review: From the Tsar’s Gardens to Soviet Fields: A History of Agricultural Experimental Institutions, Eighteenth Century to the 1920s. Isis, 101(4):893–894, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659698.

Spiro:2010:BRR

[291] Jonathan Spiro. Book review: Race and Science: Scientific Challenges to Racism in Modern America. Isis, 101(4):894–895, December 2010.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659699.

Moore:2010:BRL

[292] Gregory H. Moore. Book review: Labyrinth of Thought: A History of Set Theory and Its Role in Modern Mathematics. Isis, 101(4):895–896, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659700.

Jordanova:2010:BRP

[293] Ludmilla Jordanova. Book review: Picturing Medical Progress from Pasteur to Polio: A History of Mass Media Images and Popular Attitudes in America. Isis, 101(4):896–897, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659701.

Bellon:2010:BRA

[294] Richard Bellon. Book review: Of Apes and Ancestors: Evolution, Christianity, and the Oxford Debate. Isis, 101(4):897–898, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659702.

Rossiter:2010:BRS

[295] Margaret W. Rossiter. Book review: Science on the Home Front: American Women Scientists in World War II. Isis, 101(4):898–900, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/659703.

Beyler:2010:BRA

[296] Richard Beyler. Book review: Aufrecht im Sturm der Zeit: Der Physiker James Franck 1882–1964. Isis, 101(4):900–901, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659704.

Anderson:2010:BRS

[297] Katharine Anderson. Book review: Le savant et la tempête: Étudier l’atmosphère et prévoir le temps au XIXe siècle. Isis, 101(4):901–902, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659705.

Boehm:2010:BRP

[298] Stephan Boehm. Book review: Prophet of Innovation: Joseph Schumpeter and Creative Destruction. Isis, 101(4):902–903, December 2010.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659706.

Smith:2010:BRD

[299] Jonathan Smith. Book review: Darwin’s Camera: Art and Photography in the Theory of Evolution. Isis, 101(4):903–904, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659707.

Dyck:2010:BRS

[300] Erika Dyck. Book review: A Short History of the Drug Receptor Concept. Isis, 101(4):904–905, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/659708.

Osborn:2010:BRD

[301] Matthew Osborn. Book review: The Dawn of Green: Manchester, Thirlmere, and Modern Environmentalism. Isis, 101(4):905–906, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/659709.

Navarro:2010:BRA

[302] Jaume Navarro. Book review: Astronomy and Astrophysics in Spain (1850–1914). Isis, 101(4):906–907, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/659710.

Heidarzadeh:2010:BRW

[303] Tofigh Heidarzadeh. Book review: Who Is Knowledgeable Is Strong: Science, Class, and the Formation of Modern Iranian Society, 1900– 1950. Isis, 101(4):907–908, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659711.

Peckhaus:2010:BRF

[304] Volker Peckhaus. Book review: The Fuzzification of Systems: The Genesis of Fuzzy Set Theory and Its Initial Applications: Developments up to the 1970s. Isis, 101(4):908–909, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659712.

Allen:2010:BRD

[305] Garland E. Allen. Book review: Defending the Master Race: Conservation, Eugenics, and the Legacy of Madison Grant. Isis, 101(4):909–911, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659713.

Slotten:2010:BRL

[306] Hugh Richard Slotten. Book review: Longitude by Wire: Finding North America. Isis, 101(4):911, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659714.

Borrelli:2010:BRE

[307] Arianna Borrelli. Book review: Einstein’s Generation: The Origins of the Relativity Revolution. Isis, 101(4):912–913, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659715.

Gronim:2010:BRP

[308] Sara S. Gronim. Book review: A Passion for Nature: Thomas Jefferson and Natural History Mr, by Keith Thomson; Lee Alan Dugatkin. Jefferson and the Giant Moose: Natural History in Early America. Isis, 101 (4):913–914, December 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
659716.

Mehos:2010:BRS

[309] Donna C. Mehos. Book review: De Stem van de Wetenschap: Geschiedenis van de Koninklijke Nederlandse Akademie van Wetenschappen. Isis, 101(4):914–915, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/659717.

Hannaway:2010:BRW

[310] Caroline Hannaway. Book review: The Woman Who Walked into the Sea: Huntington’s and the Making of a Genetic Disease. Isis, 101(4):915–916, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/659718.

deRidder:2010:BRD

[311] Matthijs de Ridder. Book review: Dutch Messengers: A History of Science Publishing, 1930–1980. Isis, 101(4):916–917, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659719.

Fortun:2010:BRG

[312] Mike Fortun. Book review: Genomes and What to Make of Them. Isis, 101(4):917–918, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/659720.

McConnell:2010:BRE

[313] Craig McConnell. Book review: Einstein’s Telescope: The Hunt for Dark Matter and Dark Energy in the Universe. Isis, 101(4):918–919, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/659721.

Harkness:2010:BRC

[314] Jon M. Harkness. Book review: Contested Medicine: Cancer Research and the Military. Isis, 101(4):919–921, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659722.

Holbrook:2010:BRS

[315] Daniel Holbrook. Book review: Sources of Weapons Systems Innovation in the Department of Defense: The Role of In-House Research and Development, 1945–2000. Isis, 101(4):921–922, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659723.

Barnick:2010:BRM

[316] Heather Barnick. Book review: Making Virtual Worlds: Linden Lab and Second Life. Isis, 101(4):922, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659724.

Frickel:2010:BRM

[317] Scott Frickel. Book review: Mobilizing Science: Movements, Participation, and the Remaking of Knowledge. Isis, 101(4):923, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659725.

Wilson:2010:BRH

[318] Duncan Wilson. Book review: Hearts Exposed: Transplants and the Media in 1960s Britain. Isis, 101(4):923–924, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659726.

Kohler:2010:BRI

[319] Robert E. Kohler. Book review: The Infanticide Controversy: Primatology and the Art of Field Science. Isis, 101(4):924–926, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659727.

Gross:2010:BRC

[320] Alan Gross. Book review: Communicating Science: Professional, Popular, Literary. Isis, 101(4):926–927, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/659728.

Dolan:2010:BRC

[321] Brian Dolan. Book review: CT Suite: The Work of Diagnosis in the Age of Noninvasive Cutting. Isis, 101(4):927–928, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659729.

Wellerstein:2010:BRR

[322] Alex Wellerstein. Book review: The Road to Yucca Mountain: The Development of Radioactive Waste Policy in the United States. Isis, 101 (4):928–929, December 2010. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
659730.

Yaneva:2010:BRA

[323] Albena Yaneva. Book review: Artifice and Design: Art and Technology in Human Experience. Isis, 101(4):929–930, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659731.

Douglas:2010:BRI

[324] Heather Douglas. Book review: Imperfect Oracle: The Epistemic and Moral Authority of Science. Isis, 101(4):931–932, December 2010. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/659732.

Patton:2010:BRS

[325] Lydia Patton. Book review: Scientific Understanding: Philosophical Perspectives. Isis, 101(4):932–933, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659733.

Biddle:2010:BRR

[326] Justin Biddle. Book review: The Responsible Scientist: A Philosophical Inquiry. Isis, 101(4):933, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659734.

Hilgartner:2010:BRB

[327] Stephen Hilgartner. Book review: Beyond the Hoax: Science, Philosophy, and Culture. Isis, 101(4):934–935, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659735.

Zammito:2010:BRS

[328] John H. Zammito. Book review: Science and the Quest for Meaning. Isis, 101(4):935–936, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/659736.

Anonymous:2010:NCd

[329] Anonymous. Notes on contributors. Isis, 101(4):937–942, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/664964.

Anonymous:2010:IIV

[330] Anonymous. Index to Isis, volume 101, 2010. Isis, 101(4):943–955, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/664962.

Anonymous:2010:CV

[331] Anonymous. Contents, volume 101. Isis, 101(4):959–961, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664963.

Anonymous:2010:Fd

[332] Anonymous. Frontispiece. Isis, 101(4):vi, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/659866.

Anonymous:2011:CBH

[333] Anonymous. Current bibliography of the history of science and its cultural influences, 2011. Isis, 102(S1):i–327, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664935.

Kuklick:2011:PER

[334] Henrika Kuklick. Personal equations: Reflections on the history of fieldwork, with special reference to sociocultural anthropology. Isis, 102(1): 1–33, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/658655.

Larson:2011:PSG

[335] Edward J. Larson. Public science for a global empire: The British quest for the south magnetic pole. Isis, 102(1):34–59, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/658656.

Strasser:2011:EMG

[336] Bruno J. Strasser. The experimenter’s museum: GenBank, natural history, and the moral economies of biomedicine. Isis, 102(1):60–96, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/658657.

Pickstone:2011:I

[337] John V. Pickstone and Michael Worboys. Introduction. Isis, 102(1):97– 101, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/658658.

Cook:2011:HMS

[338] Harold J. Cook. The history of medicine and the scientific revolution. Isis, 102(1):102–108, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/658659.

Worboys:2011:PSM

[339] Michael Worboys. Practice and the science of medicine in the Nineteenth Century. Isis, 102(1):109–115, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/658660.

Lowy:2011:HBB

[340] Ilana Löwy. Historiography of biomedicine: “Bio,” “Medicine,” and in between. Isis, 102(1):116–122, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/658661.

Pickstone:2011:STM

[341] John V. Pickstone. Sketching together the modern histories of science, technology, and medicine. Isis, 102(1):123–133, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657506.

Shank:2011:LE

[342] J. B. Shank. Letter to the Editor. Isis, 102(1):134–135, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/658662.

Feingold:2011:LE

[343] Mordechai Feingold. Letter to the Editor. Isis, 102(1):135–136, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/658663.

Green:2011:LE

[344] Judy Green and Jeanne LaDuke. Letter to the Editor. Isis, 102(1):136– 137, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/658664.

Fetzer:2011:LE

[345] James H. Fetzer. Letter to the Editor. Isis, 102(1):137–138, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/658859.

Anonymous:2011:Ea

[346] Anonymous. Errata. Isis, 102(1):138, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/660266.

Swerdlow:2011:LE

[347] N. M. Swerdlow. Letter to the Editor. Isis, 102(1):138, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/660265.

Aubin:2011:BRD

[348] David Aubin. Book review: Duel at Dawn: Heroes, Martyrs, and the Rise of Modern Mathematics. Isis, 102(1):139–140, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660206.

Grandin:2011:BRH

[349] Karl Grandin. Book review: Håkon With Andersen, Brita Brenna, Magne Njåstad, and Astrid Wade. Æmula Lauri: The Royal Norwegian Society of Sciences and Letters, 1760–2010. Isis, 102(1):140–141, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/660207.

Ritvo:2011:BRH

[350] Harriet Ritvo. Book review: Histoire humaine des animaux, de l’Antiquité à nos jours. Isis, 102(1):141–142, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/660208.

Glick:2011:BRB

[351] Thomas F. Glick. Book review: Bibliografía crítica ilustrada de las obras de Darwin en España (1857–2008). Isis, 102(1):142, March 2011. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/660209.

Gingras:2011:BRF

[352] Yves Gingras. Book review: On Fact and Fraud: Cautionary Tales from the Front Lines of Science. Isis, 102(1):142–143, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660210.

Sengor:2011:BRH

[353] A. M. Celâl Sengör. Book review: History of Geomorphology and Quaternary Geology. Isis, 102(1):143–145, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660211.

Knight:2011:BRN

[354] David Knight. Book review: Natural Science in Western History: Complete Edition. Isis, 102(1):145–146, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660212.

Yeo:2011:BRE

[355] Richard Yeo. Book review: The Early Britannica: The Growth of an Outstanding Encyclopedia. Isis, 102(1):146–147, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660213.

Bocking:2011:BRO

[356] Stephen Bocking. Book review: The Oyster Question: Scientists, Watermen, and the Maryland Chesapeake Bay since 1880. Isis, 102(1):147–148, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/660214.

Kohlstedt:2011:BRW

[357] Sally Gregory Kohlstedt. Book review: Women and Science in India: A Reader. Isis, 102(1):148–149, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660215.

Reynolds:2011:BRCa

[358] Andrew Reynolds. Book review: Culturing Life: How Cells Became Technologies. Isis, 102(1):149–150, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660216.

Secord:2011:BRM

[359] James A. Secord. Book review: The Making of the Geological Society of London. Isis, 102(1):150–151, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660217.

Doel:2011:BRF

[360] Ronald E. Doel. Book review: The Fluid Envelope of Our Planet: How the Study of Ocean Currents Became a Science. Isis, 102(1):151–153, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/660218.

Castro:2011:BRC

[361] Filipe Castro. Book review: Creating Shapes in Civil and Naval Architecture: A Cross-Disciplinary Comparison. Isis, 102(1):153–154, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/660219.

Borrello:2011:BRN

[362] Mark E. Borrello. Book review: The New Foundations of Evolution: On the Tree of Life. Isis, 102(1):154–155, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660220.

Roberts:2011:BRH

[363] David Lindsay Roberts. Book review: How to Read Historical Mathematics. Isis, 102(1):155–156, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/660221.

Steele:2011:BRW

[364] John Steele. Book review: When the Gods Were Born: Greek Cosmogonies and the Near East. Isis, 102(1):156–157, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660222.

Garbutt:2011:BRQ

[365] Gwyndaf Garbutt. Book review: Questions Concerning Aristotle’s On Animals. Isis, 102(1):157–158, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660223.

Smith:2011:BRF

[366] Pamela H. Smith. Book review: A Forest on the Sea: Environmental Expertise in Renaissance Venice. Isis, 102(1):158–159, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/660224.

Giuliani:2011:BRC

[367] Erika Giuliani. Book review: Carolus Clusius in a New Context: Towards a Cultural History of a Renaissance Naturalist. Isis, 102(1):159– 160, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/660225.

Marr:2011:BRV

[368] Alexander Marr. Book review: Vision, Reflection, and Desire in Western Painting The Mirror, the Window, and the Telescope: How Renaissance Linear Perspective Changed Our Vision of the Universe. Isis, 102(1):160– 162, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/660226.

Sloan:2011:BRG

[369] Phillip R. Sloan. Book review: George-Louis Leclerc de Buffon. Histoire naturelle, générale et particulière, avec la description du Cabinet du roy. Volume 2. Edited by Stéphane Schmitt with the assistance of Cédric Crémière. (L’Âge des Lumières, 45.) 805 pp., illus., apps., bibl., indexes.

Paris: Honoré Champion, 2008. 130 (cloth); George-Louis Leclerc de Buffon. Histoire naturelle, générale et particulière, avec la description du Cabinet du roy. Volume 3. Edited by Stéphane Schmitt with the assistance of Cédric Crémière. (L’Âge des Lumières, 46.) 767 pp., illus., apps., bibl., indexes. Paris: Honoré Champion, 2009. 120 (cloth). Isis, 102(1):162–164, March 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
660227.

Monti:2011:BRA

[370] Maria Teresa Monti. Book review: Albrecht von Haller im Göttingen der Aufklärung. Isis, 102(1):164–165, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660228.

Kremer:2011:BRN

[371] Richard L. Kremer. Book review: Nouveau ciel, nouvelle terre: La révolution copernicienne dans l’Allemagne de la Réforme (1530–1630). Isis, 102(1):165–166, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/660229.

Croce:2011:BRS

[372] Paul J. Croce. Book review: Swedenborg, Mesmer, and the Mind/Body Connection: The Roots of Complementary Medicine. Isis, 102(1):166– 167, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/660230.

Seidengart:2011:BRJ

[373] Jean Seidengart. Book review: Johannes Kepler: From Tübingen to Zagan. Isis, 102(1):167–168, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/660231.

Dooley:2011:BRR

[374] Brendan Dooley. Book review: Rome et la science moderne: Entre Renaissance et Lumières. Isis, 102(1):168–170, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660232.

Woodward:2011:BRP

[375] Walter Woodward. Book review: Prospero’s America: John Winthrop, Jr, by Walter Woodward., Alchemy, and the Creation of New England Culture, 1606–1676. Isis, 102(1):170–171, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660233.

Jarrell:2011:BRC

[376] Richard A. Jarrell. Book review: Communities of Science in Nineteenth-Century Ireland Natural History Societies and Civic Culture in Victorian Scotland. Isis, 102(1):171–173, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660234.

Cahan:2011:BRH

[377] David Cahan. Book review: The Heavens on Earth: Observatories and Astronomy in Nineteenth-Century Science and Culture. Isis, 102(1):173– 174, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/660235.

Gaudilliere:2011:BRV

[378] Jean-Paul Gaudillière. Book review: Vitamin C für alle! Pharmazeutische Produktion, Vermarktung und Gesundheitspolitik (1933–1953). Isis, 102(1):175, March 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
660236.

Gere:2011:BRM

[379] Cathy Gere. Book review: Mirrors of Memory: Freud, Photography, and the History of Art. Isis, 102(1):175–176, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660237.

Williams:2011:BRR

[380] Elizabeth A. Williams. Book review: De la race à l’évolution: Paul Broca et l’anthropologie française (1850–1900). Isis, 102(1):176–178, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/660238.

Zeller:2011:BRM

[381] Suzanne Zeller. Book review: Magnetic Fever: Global Imperialism and Empiricism in the Nineteenth Century. Isis, 102(1):178–179, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/660239.

Ruse:2011:BRD

[382] Michael Ruse. Book review: Darwin in Ilkley. Isis, 102(1):179, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/660240.

Larson:2011:BRS

[383] Edward J. Larson. Book review: Segregation’s Science: Eugenics and Society in Virginia. Isis, 102(1):180, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660241.

Lowy:2011:BRE

[384] Ilana Löwy. Book review: Epidemic Invasions: Yellow Fever and the Limits of Cuban Independence, 1878–1930. Isis, 102(1):180–181, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/660242.

Cantor:2011:BRT

[385] Geoffrey Cantor. Book review: ThermoPoetics: Energy in Victorian Literature and Science. Isis, 102(1):181–182, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/660243.

Barberis:2011:BRH

[386] Daniela S. Barberis. Book review: Hysteria Complicated by Ecstasy: The Case of Nanette Leroux. Isis, 102(1):182–183, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660244.

Worboys:2011:BRL

[387] Michael Worboys. Book review: Laboratory Disease: Robert Koch’s Medical Bacteriology. Isis, 102(1):183–184, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/660245.

Drucker:2011:BRP

[388] Thomas Drucker. Book review: Plato’s Ghost: The Modernist Transformation of Mathematics. Isis, 102(1):185–186, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660246.

Corry:2011:BRW

[389] Leo Corry. Book review: The World as a Mathematical Game: John von Neumann and Twentieth Century Science. Isis, 102(1):186–187, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/660247.

Nye:2011:BRW

[390] David E. Nye. Book review: Waterpower in Lowell: Engineering and Industry in Nineteenth-Century America. Isis, 102(1):187–188, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/660248.

Benes:2011:BRG

[391] Tuska Benes. Book review: German Orientalism in the Age of Empire: Religion, Race, and Scholarship. Isis, 102(1):188–189, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/660249.

Wassmann:2011:BRA

[392] Claudia Wassmann. Book review: About Faces: Physiognomy in Nineteenth-Century Britain. Isis, 102(1):190–191, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660250.

Levere:2011:BRI

[393] Trevor H. Levere. Book review: Image and Reality: Kekulé, Kopp, and the Scientific Imagination. Isis, 102(1):191–192, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660251.

Comfort:2011:BRP

[394] Nathaniel Comfort. Book review: In Pursuit of the Gene: From Darwin to DNA. Isis, 102(1):192–193, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660252.

Gerovitch:2011:BRR

[395] Slava Gerovitch. Book review: The Red Rockets’ Glare: Spaceflight and the Soviet Imagination, 1857–1957. Isis, 102(1):193–194, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/660253.

VanRiper:2011:BRSa

[396] A. Bowdoin Van Riper. Book review: “A Splendid Idiosyncrasy”: Prehistory at Cambridge, 1915–50. Isis, 102(1):195, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660254.

Stapleford:2011:BRS

[397] Thomas A. Stapleford. Book review: The Statistical Mind in Modern Society: The Netherlands, 1850–1940, by Jacques G. S. J. van Maarseveen; Paul M. M. Klep; Ida H. Stamhuis. Volume 1: Official Statistics, Social Progress, and Modern Enterprise. Volume 2: Statistics and Scientific Work. Isis, 102(1):195–197, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660255.

Dettelbach:2011:BRP

[398] Michael Dettelbach. Book review: The Passage to Cosmos: Alexander von Humboldt and the Shaping of America. Isis, 102(1):197–198, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/660256.

Fleming:2011:BRN

[399] James Rodger Fleming. Book review: A Nuclear Winter’s Tale: Science and Politics in the 1980s. Isis, 102(1):198–199, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660257.

Holloway:2011:BRM

[400] David Holloway. Book review: Meeting the Demands of Reason: The Life and Thought of Andrei Sakharov. Isis, 102(1):199–200, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/660258.

Oldham:2011:BRA

[401] Kalil Oldham. Book review: Atombilder: Ikonografie des Atoms in Wissenschaft und Öffentlichkeit des 20, by Charlotte Bigg; Jochen Hennig. Jahrhunderts. Isis, 102(1):200–201, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660259.

Heyck:2011:BRK

[402] Hunter Heyck. Book review: Know Your Enemy: The Rise and Fall of America’s Soviet Experts. Isis, 102(1):201–202, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660260.

Creager:2011:BRF

[403] Angela N. H. Creager. Book review: Francis Crick: Hunter of Life’s Secrets. Isis, 102(1):202–204, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660261.

Elliott:2011:BRS

[404] Kevin Elliott. Book review: Science, Policy, and the Value-Free Ideal. Isis, 102(1):204–205, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/660262.

Nickles:2011:BRK

[405] Thomas Nickles. Book review: Karl Popper’s Philosophy of Science: Rationality without Foundations Thomas Kuhn’s “Linguistic Turn” and the Legacy of Logical Empiricism: Incommensurability, Rationality, and the Search for Truth. Isis, 102(1):205–207, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/660263.

Yeang:2011:BRB

[406] Chen-Pang Yeang. Book review: The Beginnings of Piezoelectricity: A Study in Mundane Physics. Isis, 102(1):207–208, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660264.

Anonymous:2011:NCa

[407] Anonymous. Notes on contributors. Isis, 102(1):209–213, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/660307.

Anonymous:2011:Fa

[408] Anonymous. Frontispiece. Isis, 102(1):vi, March 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/660356.

Boddice:2011:VMV

[409] Rob Boddice. Vivisecting major: A Victorian Gentleman scientist defends animal experimentation, 1876–1885. Isis, 102(2):215–237, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/660128.

Nicholson:2011:SMS

[410] Ian Nicholson. “Shocking” masculinity: Stanley Milgram, “obedience to authority,” and the “crisis of manhood” in Cold War America. Isis, 102(2):238–268, June 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
660129.

Weidman:2011:PAM

[411] Nadine Weidman. Popularizing the ancestry of man: Robert ardrey and the killer instinct. Isis, 102(2):269–299, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/660130.

Moran:2011:I

[412] Bruce T. Moran. Introduction. Isis, 102(2):300–304, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/660138.

Principe:2011:AR

[413] Lawrence M. Principe. Alchemy restored. Isis, 102(2):305–312, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/660139.

Newman:2011:WWL

[414] William R. Newman. What have we learned from the recent historiography of alchemy? Isis, 102(2):313–321, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/660140.

Chang:2011:ASL

[415] Ku ming (Kevin) Chang. Alchemy as studies of life and matter: Reconsidering the place of vitalism in early modern chymistry. Isis, 102(2): 322–329, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/660141.

Nummedal:2011:WWH

[416] Tara E. Nummedal. Words and works in the history of alchemy. Isis, 102(2):330–337, June 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
660142.

Anonymous:2011:HSS

[417] Anonymous. 2010 History of Science Society prize citations. Isis, 102(2): 338–342, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/660143.

Mueller-Wille:2011:LE

[418] Staffan Mueller-Wille. Letter to the Editor. Isis, 102(2):343, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/660145.

Schwartz:2011:LE

[419] James Schwartz. Letter to the Editor. Isis, 102(2):343, June 2011. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/660144.

Day:2011:LE

[420] Dwayne A. Day. Letter to the Editor. Isis, 102(2):343–344, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/660146.

Macrakis:2011:LE

[421] Kristie Macrakis. Letter to the Editor. Isis, 102(2):344–345, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/661665.

Borgerson:2011:BRC

[422] Kirstin Borgerson. Book review: The Cambridge World History of Medical Ethics. Isis, 102(2):346–347, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/661666.

Brain:2011:BRC

[423] Robert M. Brain. Book review: The Culture of Diagram. Isis, 102(2): 347–348, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/661667.

vanderMeer:2011:BRC

[424] Jitse M. van der Meer. Book review: The Cambridge Companion to Science and Religion. Isis, 102(2):348–349, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/661668.

Tighe:2011:BRD

[425] Janet A. Tighe. Book review: Death Investigation in America: Coroners, Medical Examiners, and the Pursuit of Medical Certainty. Isis, 102(2): 349–350, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/661669.

Stewart:2011:BRP

[426] Larry Stewart. Book review: Piracy: The Intellectual Property Wars from Gutenberg to Gates. Isis, 102(2):350–351, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/661670.

Cowie:2011:BRR

[427] Helen Cowie. Book review: El Rinoceronte y el Megaterio: Un ensayo de morfología histórica. Isis, 102(2):351–352, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/661671.

Vicedo:2011:BRL

[428] Marga Vicedo. Book review: Looking for a Few Good Males: Female Choice in Evolutionary Biology. Isis, 102(2):352–353, June 2011. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/661672.

Roby:2011:BRM

[429] Courtney Roby. Book review: The Mechanical Hypothesis in Ancient Greek Natural Philosophy. Isis, 102(2):353–354, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/661673.

Smith:2011:BRB

[430] Lisa Wynne Smith. Book review: The Birth of Mankind: Otherwise Named, The Woman’s Book. Isis, 102(2):354–355, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/661674.

Andrault:2011:BRL

[431] Raphaële Andrault. Book review: Leibniz: Le vivant et l’organisme. Isis, 102(2):355–356, June 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
661675.

Golinski:2011:BRM

[432] Jan Golinski. Book review: Materials and Expertise in Early Modern Europe: Between Market and Laboratory. Isis, 102(2):356–357, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/661676.

Heringman:2011:BRA

[433] Noah Heringman. Book review: Autobiography and Natural Science in the Age of Romanticism: Rousseau, Goethe, Thoreau. Isis, 102(2):357– 358, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/661677.

Henry:2011:BRN

[434] John Henry. Book review: Nicholas Steno’s Challenge for Truth. Isis, 102(2):358–359, June 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
661678.

Barnett:2011:BRW

[435] Lydia Barnett. Book review: The World Makers: Scientists of the Restoration and the Search for the Origins of the Earth. Isis, 102(2): 359–360, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/661679.

Boyle:2011:BRN

[436] Deborah Boyle. Book review: The Natural Philosophy of Margaret Cavendish: Reason and Fancy during the Scientific Revolution. Isis, 102(2):360–361, June 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
661680.

Long:2011:BRF

[437] Pamela O. Long. Book review: Fireworks: Pyrotechnic Arts and Sciences in European History. Isis, 102(2):361–362, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/661681.

Luthy:2011:BRT

[438] Christoph Lüthy. Book review: Theatrum philosophicum: Descartes und die Rolle ästhetischer Formen in der Wissenschaft. Isis, 102(2):362– 364, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/661682.

Sauer:2011:BRQ

[439] Tilman Sauer. Book review: Quantum Theory at the Crossroads: Reconsidering the 1927 Solvay Conference. Isis, 102(2):364–365, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/661683.

Nieuwland:2011:BRD

[440] Ilja Nieuwland. Book review: David Baneke, Synthetisch denken: Natuurwetenschappers over hun rol in een moderne maatschappij, 1900–1940. Isis, 102(2):365–366, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/661684.

Berger:2011:BRS

[441] Silvia Berger. Book review: Spinning the Scientific Web: Jacques Loeb (1859–1924) und sein Programm einer internationalen biomedizinischen Grundlagenforschung. Isis, 102(2):366–367, June 2011. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/661685.

Bret:2011:BRR

[442] Patrice Bret. Book review: La république des inventeurs: Propriété et organisation de l’innovation en France (1791–1922). Isis, 102(2):367– 368, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/661686.

Archibald:2011:BRS

[443] Thomas Archibald. Book review: The Shaping of Arithmetic after C, by Catherine Goldstein; Norbert Schappacher; Joachim Schwermer. F. Gauss’s Disquisitiones arithmeticae. Isis, 102(2):368–369, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/661687.

Wilson:2011:BRM

[444] David B. Wilson. Book review: The Making of Modern Science: Science, Technology, Medicine, and Modernity, 1789–1914. Isis, 102(2):369–371, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/661688.

Tolley:2011:BRT

[445] Kim Tolley. Book review: Teaching Children Science: Hands-On Nature Study in North America, 1890–1930. Isis, 102(2):371–372, June 2011.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/661689.

Clark:2011:BRA

[446] Constance Areson Clark. Book review: The Art of Evolution: Darwin, Darwinisms, and Visual Culture. Isis, 102(2):372–373, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/661690.

Navarro:2011:BRJ

[447] Jaume Navarro. Book review: João Magueijo: A Brilliant Darkness: The Extraordinary Life and Mysterious Disappearance of Ettore Majorana, the Troubled Genius of the Nuclear Age. Isis, 102(2):373–374, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/661691.

Roland:2011:BRC

[448] Alex Roland. Book review: The Cognitive Challenge of War: Prussia, 1806. Isis, 102(2):374, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/661692.

Lynch:2011:BRR

[449] John M. Lynch. Book review: Richard Owen: Biology without Darwin. Isis, 102(2):374–375, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/661693.

Camilleri:2011:BRC

[450] Kristian Camilleri. Book review: Crafting the Quantum: Arnold Sommerfeld and the Practice of Theory, 1890–1926. Isis, 102(2):375–376, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/661694.

Smith:2011:BRC

[451] Robert W. Smith. Book review: Cosmic Noise: A History of Early Radio Astronomy. Isis, 102(2):377–378, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/661695.

Lachapelle:2011:BRS

[452] Sofie Lachapelle. Book review: The Stepchildren of Science: Psychical Research and Parapsychology in Germany, c, by Heather Wolffram.

1870–1939. Isis, 102(2):378, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/661696.

Rumiel:2011:BRD

[453] Lisa Rumiel. Book review: The Dragon’s Tail: Americans Face the Atomic Age. Isis, 102(2):378–379, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/661697.

Krimsky:2011:BRT

[454] Sheldon Krimsky. Book review: Toxic Bodies: Hormone Disruptors and the Legacy of DES. Isis, 102(2):379–380, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/661698.

Sinclair:2011:BRW

[455] Bruce Sinclair. Book review: When the Lights Went Out: A History of Blackouts in America. Isis, 102(2):380–381, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/661699.

Tang:2011:BRR

[456] Joyce Tang. Book review: Race, Rigor, and Selectivity in U, by Amy E. Slaton.S. Engineering: The History of an Occupational Color Line. Isis, 102(2):381–382, June 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
661700.

Josefowicz:2011:BRT

[457] Diane Greco Josefowicz. Book review: True Jacob: A Novel. Isis, 102(2): 382–383, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/661701.

Maienschein:2011:BRM

[458] Jane Maienschein. Book review: The Mirage of a Space between Nature and Nurture. Isis, 102(2):383–384, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/661702.

Lynch:2011:BRS

[459] Michael Lynch. Book review: Standards and Their Stories: How Quantifying, Classifying, and Formalizing Practices Shape Everyday Life. Isis, 102(2):384–385, June 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
661703.

Radder:2011:BRU

[460] Hans Radder. Book review: Unsimple Truths: Science, Complexity, and Policy. Isis, 102(2):385–386, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/661704.

Malaby:2011:BRS

[461] Thomas Malaby. Book review: Simulation and Its Discontents. Isis, 102(2):387–388, June 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
661705.

Anonymous:2011:NCb

[462] Anonymous. Notes on contributors. Isis, 102(2):389–391, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/661707.

Anonymous:2011:Fb

[463] Anonymous. Frontispiece. Isis, 102(2):vi, June 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/661706.

Bellon:2011:IHD

[464] Richard Bellon. Inspiration in the harness of daily labor: Darwin, botany, and the triumph of evolution, 1859–1868. Isis, 102(3):393–420, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/661591.

Quintero:2011:TBI

[465] Camilo Quintero. Trading in birds: Imperial power, national pride, and the place of nature in U.S.–Colombia relations. Isis, 102(3):421–445, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/661592.

Yi:2011:WOW

[466] Doogab Yi. Who owns what? Private ownership and the public interest in recombinant DNA technology in the 1970s. Isis, 102(3):446–474, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/661619.

Alexander:2011:SCS

[467] Amir Alexander. The skeleton in the closet: Should historians of science care about the history of mathematics? Isis, 102(3):475–480, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/661620.

Epple:2011:BTH

[468] Moritz Epple. Between timelessness and historiality: On the dynamics of the epistemic objects of mathematics. Isis, 102(3):481–493, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/661622.

Dick:2011:AWP

[469] Stephanie Dick. AfterMath: The work of proof in the age of human– machine collaboration. Isis, 102(3):494–505, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/661623.

Richards:2011:CLM

[470] Joan L. Richards. “This compendious language”: Mathematics in the world of Augustus De Morgan. Isis, 102(3):506–510, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/661624.

Gray:2011:HMH

[471] Jeremy Gray. History of mathematics and history of science reunited? Isis, 102(3):511–517, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/661625.

Mann:2011:HMH

[472] Tony Mann. History of mathematics and history of science. Isis, 102 (3):518–526, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
661626.

Osborne:2011:ELB

[473] Michael Osborne. Eloge: Lawrence Badash, 8 May 1934–23 August 2010. Isis, 102(3):527–529, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/661627.

Newman:2011:EJM

[474] William Newman. Eloge: John Murdoch, 10 May 1927–16 September 2010. Isis, 102(3):529–532, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/663021.

Reeves:2011:BRG

[475] Eileen Reeves. Book review: Galileo. Isis, 102(3):533–536, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/661628.

Sidoli:2011:BRA

[476] Nathan Sidoli. Book review: Apollonius’s Conics: The Greek and Arabic Traditions. Isis, 102(3):537–542, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/661629.

Morrison:2011:BRI

[477] Robert G. Morrison. Book review: Islam, Science, and the Challenge of History. Isis, 102(3):543–544, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/663022.

vanDalen:2011:BRC

[478] Benno van Dalen. Book review: Le calendrier chinois: Structure et calculs (104 av J.-C.–1644), by Jean–Claude Martzloff: Indétermination céleste et réforme permanente: La construction chinoise officielle du temps quotidien discret à partir d’un temps mathématique caché, linéaire et continu. Isis, 102(3):544–545, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/663023.

Kline:2011:BRP

[479] Wendy Kline. Book review: Prescribed Norms: Women and Health in Canada and the United States since 1800. Isis, 102(3):545–546, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/663024.

Tobbell:2011:BRT

[480] Dominique A. Tobbell. Book review: Thalassaemia: The Biography. Isis, 102(3):546–547, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/663025.

Standish:2011:BRH

[481] Myles Standish. Book review: The Hill-Brown Theory of the Moon’s Motion: Its Coming-to-Be and Short-Lived Ascendancy (1877–1984). Isis, 102(3):547–548, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/663026.

Anderson:2011:BRA

[482] Robert G. W. Anderson. Book review: The Alchemy of Glass: Counterfeit, Imitation, and Transmutation in Ancient Glassmaking. Isis, 102 (3):548–549, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
663027.

Pesic:2011:BRM

[483] Peter Pesic. Book review: The Monochord in Ancient Greek Harmonic Science. Isis, 102(3):549–550, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/663028.

Touwaide:2011:BRV

[484] Alain Touwaide. Book review: Über die Verschiedenheit der Symptome. Isis, 102(3):550–551, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/663029.

Totelin:2011:BRS

[485] Laurence Totelin. Book review: The Symptom and the Subject: The Emergence of the Physical Body in Ancient Greece. Isis, 102(3):551–552, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663030.

Leunissen:2011:BRN

[486] Mariska Leunissen. Book review: Naturgesetz in der Vorstellung der Antike, besonders der Stoa: Eine Begriffsuntersuchung. Isis, 102(3): 552–553, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
663031.

Machamer:2011:BRD

[487] Peter Machamer. Book review: Disciplines in the Making: Cross-Cultural Perspectives on Elites, Learning, and Innovation. Isis, 102(3): 553–554, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
663032.

Geus:2011:BRE

[488] Klaus Geus. Book review: Eratosthenes’ Geography. Isis, 102(3):554, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663033.

Lehoux:2011:BRC

[489] Daryn Lehoux. Book review: Calendars and Years: Astronomy and Time in the Ancient Near East A Brief Introduction to Astronomy in the Middle East. Isis, 102(3):554–555, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/663034.

Stevens:2011:BRP

[490] Wesley M. Stevens. Book review: The Portfolio of Villard de Honnecourt: A New Critical Edition and Color Facsimile. Isis, 102(3): 555–557, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
663035.

Mandelbrote:2011:BRA

[491] Scott Mandelbrote. Book review: Adam and Eve in the Protestant Reformation. Isis, 102(3):557–558, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/663036.

VandenBroecke:2011:BRC

[492] Steven Vanden Broecke. Book review: Cornelius Gemma: Cosmology, Medicine, and Natural Philosophy in Renaissance Louvain. Isis, 102 (3):558–559, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
663037.

Goddu:2011:BRN

[493] André Goddu. Book review: Niccolò Copernico e la fondazione del cosmo eliocentrico: Con testo, traduzione e commentario del Libro I de Le revoluzioni celesti. Isis, 102(3):559–560, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/663038.

Belle:2011:BRW

[494] Riccardo Bellé. Book review: The Worlds of Oronce Fine: Mathematics, Instruments, and Print in Renaissance France. Isis, 102(3):560–561, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663039.

Trigg:2011:BRR

[495] Scott Trigg. Book review: Robert G. Morrison. Islam and Science: The Intellectual Career of Nīz.ām al-Dīn Nīsābūrī. Isis, 102(3):561–562, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663040.

Jenkins:2011:BRB

[496] Jane Jenkins. Book review: Branching Off: The Early Moderns in Quest for the Unity of Knowledge. Isis, 102(3):562–563, September 2011. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/663041.

Bates:2011:BRC

[497] David W. Bates. Book review: Cătălin Avramescu; translated by Alistair Ian Blyth. An Intellectual History of Cannibalism. Isis, 102(3):563–565, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663042.

Bresadola:2011:BRF

[498] Marco Bresadola. Book review: I fari di Halle: Georg Ernst Stahl, Friedrich Hoffmann e la medicina europea del primo Settecento. Isis, 102 (3):565–566, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
663043.

Mayer:2011:BRS

[499] Thomas F. Mayer. Book review: On Sunspots. Isis, 102(3):566, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/663044.

Fleischer:2011:BRK

[500] Alette Fleischer. Book review: Knowledge and Colonialism: Eighteenth-Century Travellers in South Africa. Isis, 102(3):567–568, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/663045.

Loveland:2011:BRU

[501] Jeff Loveland. Book review: The Unfinished Enlightenment: Description in the Age of the Encyclopedia. Isis, 102(3):568–569, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/663046.

Connor:2011:BRJ

[502] J. T. H. Connor. Book review: John Shaw Billings: Science and Medicine in the Gilded Age. Isis, 102(3):569–570, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/663047.

Gradmann:2011:BRM

[503] Christoph Gradmann. Book review: Metaphern unter dem Mikroskop: Die epistemische Rolle von Metaphorik in den Wissenschaften und in Robert Kochs Bakteriologie. Isis, 102(3):570–571, September 2011. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/663048.

Mitchell:2011:BRN

[504] Benjamin Mitchell. Book review: Nietzsche’s Anti-Darwinism. Isis, 102 (3):571–572, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
663049.

Lacy:2011:BRS

[505] Cherilyn Lacy. Book review: Une société à soigner: Hygiène et salubrité publiques en France au XIXe siècle. Isis, 102(3):572–573, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/663050.

Lucier:2011:BRM

[506] Paul Lucier. Book review: Mass Destruction: The Men and Giant Mines That Wired America and Scarred the Planet. Isis, 102(3):573–574, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663051.

Mirowski:2011:BRN

[507] Philip E. Mirowski. Book review: Von Neumann, Morgenstern, and the Creation of Game Theory: From Chess to Social Science, 1900–1960. Isis, 102(3):574–575, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/663052.

Gay:2011:BRS

[508] Hannah Gay. Book review: Spaces and Collections in the History of Science: The Laboratorio Chimico Overture. Isis, 102(3):575–576, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/663053.

Greene:2011:BRP

[509] Mott Greene. Book review: The Passion of George Sarton: A Modern Marriage and Its Discipline. Isis, 102(3):576–577, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/663054.

Gliboff:2011:BRD

[510] Sander Gliboff. Book review: Differenz und Vererbung: Geschlechterordnungen in der Genetik und Hormonforschung 1890–1950. Isis, 102(3): 577–578, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
663055.

Brain:2011:BRH

[511] Robert M. Brain. Book review: Die Helmholtz-Kurven: Auf der Spur der verlorenen Zeit. Isis, 102(3):578–579, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/663056.

Forsdyke:2011:BRD

[512] Donald R. Forsdyke. Book review: Darwin’s Disciple: George John Romanes, a Life in Letters. Isis, 102(3):579–580, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/663057.

Dunaway:2011:BRB

[513] Finis Dunaway. Book review: From Bauhaus to Ecohouse: A History of Ecological Design. Isis, 102(3):580–581, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/663058.

Thomas:2011:BRR

[514] William Thomas. Book review: Riding the Waves: A Life in Sound, Science, and Industry Manhattan Project to the Santa Fe Institute: The Memoirs of George A, by Leo Beranek; George A. Cowan. Cowan. Isis, 102(3):581–582, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/663059.

LaBerge:2011:BRC

[515] Ann La Berge. Book review: The Cigarette Century: The Rise, Fall, and Deadly Persistence of the Product That Defined America. Isis, 102 (3):582–584, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
663060.

Richmond:2011:BRC

[516] Marsha L. Richmond. Book review: Choosing Selection: The Revival of Natural Selection in Anglo–American Evolutionary Biology, 1930–1970. Isis, 102(3):584–585, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/663061.

Tucker:2011:BRP

[517] William Tucker. Book review: Playing with Fire: The Controversial Career of Hans J, by Roderick D. Buchanan. Eysenck. Isis, 102(3): 585–586, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
663062.

Erickson:2011:BRV

[518] Paul Erickson. Book review: A Vast Machine: Computer Models, Climate Data, and the Politics of Global Warming. Isis, 102(3):586–587, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663063.

Rocha:2011:BRC

[519] Leon Antonio Rocha. Book review: Coming to Terms with the Nation: Ethnic Classification in Modern China. Isis, 102(3):587–588, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/663064.

Ensmenger:2011:BRQ

[520] Nathan Ensmenger. Book review: The Quest for Artificial Intelligence: A History of Ideas and Achievements. Isis, 102(3):588–589, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/663065.

Anker:2011:BRM

[521] Peder Anker. Book review: Merchants of Doubt: How a Handful of Scientists Obscured the Truth on Issues from Tobacco Smoke to Global Warming. Isis, 102(3):589–590, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/663066.

Clark:2011:BRC

[522] William Clark. Book review: The Commodification of Academic Research: Science and the Modern University. Isis, 102(3):590–591, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663067.

Hentschel:2011:BRG

[523] Klaus Hentschel. Book review: Gravity’s Ghost: Scientific Discovery in the Twenty-first Century. Isis, 102(3):591–592, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/663068.

Voskuhl:2011:BRB

[524] Adelheid Voskuhl. Book review: Between Reason and Experience: Essays in Technology and Modernity. Isis, 102(3):592–593, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/663069.

Schickore:2011:BRM

[525] Jutta Schickore. Book review: N, by Matthew D. Lund; Hasok Chang. R. Hanson: Observation, Discovery, and Scientific Change. Isis, 102 (3):593–594, September 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
663070.

Kochan:2011:BRC

[526] Jeff Kochan. Book review: Cosmopolitics I. Isis, 102(3):594–595, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663071.

Topper:2011:BRS

[527] David R. Topper. Book review: Sudden Genius? The Gradual Path to Creative Breakthroughs. Isis, 102(3):595–596, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/663072.

Anonymous:2011:NCc

[528] Anonymous. Notes on contributors. Isis, 102(3):597–600, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/661631.

Anonymous:2011:Fc

[529] Anonymous. Frontispiece. Isis, 102(3):vi, September 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/663020.

deChadarevian:2011:MES

[530] Soraya de Chadarevian. The making of an entrepreneurial science: Biotechnology in Britain, 1975–1995. Isis, 102(4):601–633, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/663596.

Miller:2011:PPG

[531] David Philip Miller. The paradoxes of patenting at General Electric: Isador Ladoff’s journey from Siberian exile to the heart of corporate capitalism. Isis, 102(4):634–658, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/663597.

Ramsden:2011:RUU

[532] Edmund Ramsden. From rodent utopia to urban hell: Population, pathology, and the crowded rats of NIMH. Isis, 102(4):659–688, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/663598.

Taub:2011:IRI

[533] Liba Taub. Introduction: Reengaging with instruments. Isis, 102(4):689– 696, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663605.

Bennett:2011:EMM

[534] Jim Bennett. Early modern mathematical instruments. Isis, 102(4):697– 705, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663607.

Schaffer:2011:ECS

[535] Simon Schaffer. Easily cracked: Scientific instruments in states of disrepair. Isis, 102(4):706–717, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/663608.

Arnold:2011:MIM

[536] Ken Arnold and Thomas Söderqvist. Medical instruments in museums: Immediate impressions and historical meanings. Isis, 102(4):718–729, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663613.

Principe:2011:E

[537] Lawrence M. Principe. Eloge. Isis, 102(4):730–733, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664916.

Gattei:2011:E

[538] Stefano Gattei. Eloge. Isis, 102(4):734–737, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664837.

Anonymous:2011:Eb

[539] Anonymous. Erratum. Isis, 102(4):738, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/663616.

Schweber:2011:BRE

[540] S. S. Schweber. Book review: Einstein’s Unification. Isis, 102(4):739– 742, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/663617.

Howell:2011:BRH

[541] Joel Howell. Book review: History of Medicine: A Scandalously Short Introduction. Isis, 102(4):743, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/664917.

Pyenson:2011:BRS

[542] Lewis Pyenson. Book review: Science: A Four Thousand Year History. Isis, 102(4):744–745, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664838.

Rutkin:2011:BRB

[543] H. Darrel Rutkin. Book review: Begriffe der Astrologie: Von Abendstern bis Zwillingsproblem. Isis, 102(4):746, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664839.

Reidy:2011:BRC

[544] Michael S. Reidy. Book review: The Culture of Nature in Britain, 1680– 1860. Isis, 102(4):746–748, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/664840.

Brown:2011:BRD

[545] Karen Brown. Book review: Death in a Small Package: A Short History of Anthrax. Isis, 102(4):748, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/664841.

Reynolds:2011:BRCb

[546] Terry S. Reynolds. Book review: Changes in the Technological Landscape: Essays in the History of Science and Technology. Isis, 102(4): 748–749, December 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
664842.

Harrison:2011:BRR

[547] Peter Harrison. Book review: Reconfiguring the World: Nature, God, and Human Understanding from the Middle Ages to Early Modern Europe. Isis, 102(4):749–750, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664843.

Kingsland:2011:BRK

[548] Sharon Kingsland. Book review: Knowing Global Environments: New Historical Perspectives on the Field Sciences. Isis, 102(4):750–751, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/664844.

Cullen:2011:BRC

[549] Christopher Cullen. Book review: The Chinese Roots of Linear Algebra. Isis, 102(4):751–752, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664845.

Berryman:2011:BRE

[550] Sylvia Berryman. Book review: Explanation and Teleology in Aristotle’s Science of Nature. Isis, 102(4):752–753, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664846.

Fried:2011:BRL

[551] Michael N. Fried. Book review: Ludic Proof: Greek Mathematics and the Alexandrian Aesthetic. Isis, 102(4):753–754, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664847.

Robson:2011:BRP

[552] Eleanor Robson. Book review: In the Path of the Moon: Babylonian Celestial Divination and Its Legacy. Isis, 102(4):754–755, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664848.

Duffin:2011:BRP

[553] Jacalyn Duffin. Book review: The Professor of Secrets: Mystery, Medicine, and Alchemy in Renaissance Italy. Isis, 102(4):756, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/664849.

Garbutt:2011:BRA

[554] Gwyndaf Garbutt. Book review: From the Ark to the Pulpit: An Edition and Translation of the “Transitional” Northumberland Bestiary (Thirteenth Century). Isis, 102(4):757–758, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664850.

Kucuk:2011:BRC

[555] B. Harun Küçük. Book review: Cross-Cultural Scientific Exchanges in the Eastern Mediterranean, 1560–1660. Isis, 102(4):758–759, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/664851.

Walker:2011:BRM

[556] Timothy Walker. Book review: Materials and Medicine: Trade, Conquest, and Therapeutics in the Eighteenth Century. Isis, 102(4):759–760, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664852.

McClellan:2011:BRR

[557] James E. McClellan III. Book review: Reconceptualizing the Industrial Revolution. Isis, 102(4):760–761, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/664853.

Outram:2011:BRB

[558] Dorinda Outram. Book review: The Book That Changed Europe: Picart and Bernard’s. Isis, 102(4):761–762, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/664854.

Jacob:2011:BRR

[559] Margaret C. Jacob. Book review: Reading the Book of Nature in the Dutch Golden Age, 1575–1715. Isis, 102(4):763, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664855.

Kusukawa:2011:BRL

[560] Sachiko Kusukawa. Book review: Labors of Innocence in Early Modern England. Isis, 102(4):763–765, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/664856.

Dew:2011:BRI

[561] Nicholas Dew. Book review: The Information Master: Jean-Baptiste Colbert’s Secret State Intelligence System. Isis, 102(4):765, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/664857.

Lefevre:2011:BRG

[562] Wolfgang Lefèvre. Book review: Galileo Engineer. Isis, 102(4):766, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/664858.

Fancy:2011:BRI

[563] Nahyan Fancy. Book review: Images of Islamic Science. Isis, 102(4):766– 768, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664859.

Carteret:2011:BRB

[564] Xavier Carteret and Sara Scharf. Book review: Les beautés de la nature à l’épreuve de l’analyse: Programmes scientifiques et tentations esthétiques dans l’histoire naturelle du XVIIIe siècle (1744–1805). Isis, 102(4):768– 769, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664860.

DalPrete:2011:BRG

[565] Ivano Dal Prete. Book review: Galileo: Watcher of the Skies. Isis, 102 (4):769–770, December 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
664861.

Mulberger:2011:BRP

[566] Annette Mülberger. Book review: O projeto de uma psicologia científica em Wilhelm Wundt: Uma nova interpretação. Isis, 102(4):770–771, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/664862.

Vetter:2011:BRS

[567] Jeremy Vetter. Book review: The Second Jurassic Dinosaur Rush: Museums and Paleontology in America at the Turn of the Twentieth Century Barnum Brown: The Man Who Discovered. Isis, 102(4):771–772, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664863.

Farber:2011:BRU

[568] Paul Lawrence Farber. Book review: Until Darwin: Science, Human Variety, and the Origins of Race. Isis, 102(4):772–773, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664864.

VanRiper:2011:BRSb

[569] A. Bowdoin Van Riper. Book review: Steam Coffin: Captain Moses Rogers and the Steamship Savannah Break the Barrier. Isis, 102(4):773– 774, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664865.

Arabatzis:2011:BRT

[570] Theodore Arabatzis. Book review: A Tenth of a Second: A History. Isis, 102(4):774–775, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664866.

Hoppe:2011:BRF

[571] Brigitte Hoppe. Book review: La fabrique du visage: De la physionomie antique à la première greffe, avec un inédit de Duchenne de Boulogne. Isis, 102(4):775–776, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664867.

Endersby:2011:BRS

[572] Jim Endersby. Book review: The Skull Collectors: Race, Science, and America’s Unburied Dead. Isis, 102(4):776–777, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664868.

Fraser:2011:BRF

[573] Craig Fraser. Book review: Die Fixierung des modernen Wissenschaftsideals durch Laplace. Isis, 102(4):777–778, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664869.

Bellon:2011:BRC

[574] Richard Bellon. Book review: The Cambridge Companion to Darwin. Isis, 102(4):778–779, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664870.

Gooday:2011:BRS

[575] Graeme Gooday. Book review: Seriality and Scientific Objects in the Nineteenth Century. Isis, 102(4):779–780, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/664871.

Crowe:2011:BRD

[576] Michael J. Crowe and Stephen Case. Book review: Discoverers of the Universe: William and Caroline Herschel. Isis, 102(4):780–781, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/664872.

Daum:2011:BRE

[577] Andreas W. Daum. Book review: Essay on the Geography of Plants. Isis, 102(4):781–782, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664873.

Seth:2011:BRB

[578] Suman Seth. Book review: Bruce J. Hunt: Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein. Isis, 102 (4):782–783, December 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
664874.

Richardson:2011:BRS

[579] Alan Richardson. Book review: Science and the Life-World: Essays on Husserl’s “Crisis of European Sciences”, by David Hyder; Hans-Jörg Rheinberger. Isis, 102(4):783–784, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/664875.

Zimmerman:2011:BRS

[580] David Zimmerman. Book review: Scientific Research in World War II: What Scientists Did in the War. Isis, 102(4):784–785, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664876.

Opitz:2011:BRR

[581] Donald L. Opitz. Book review: Regionalizing Science: Placing Knowledges in Victorian England. Isis, 102(4):785–786, December 2011. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664877.

Rodriguez:2011:BRM

[582] Julia Rodriguez. Book review: From Man to Ape: Darwinism in Argentina, 1870–1920. Isis, 102(4):786–787, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/664878.

Phillipson:2011:BRA

[583] Tacye Phillipson. Book review: Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris. Isis, 102(4):787–788, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664879.

Hogselius:2011:BRT

[584] Per Högselius. Book review: Technology and the Making of The Netherlands: The Age of Contested Modernization, 1890–1970. Isis, 102(4): 788–789, December 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
664880.

Fenster:2011:BRM

[585] Della D. Fenster. Book review: Mathematicians Fleeing from Nazi Germany: Individual Fates and Global Impact. Isis, 102(4):789–790, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/664881.

Neswald:2011:BRM

[586] Elizabeth Neswald. Book review: “Morgen möchte ich wieder 100 herrliche Sachen ausrechnen”: Iris Runge bei Osram und Telefunken. Isis, 102(4):790–791, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664882.

Klein:2011:BRO

[587] Ursula Klein. Book review: Aux origines de la chimie organique: Méthodes et pratiques des pharmaciens et des chimistes (1785–1835). Isis, 102(4):791–793, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664883.

Bud:2011:BRH

[588] Robert Bud. Book review: Health in the Marketplace: Professionalism, Therapeutic Desires, and Medical Commodification in Late-Victorian London. Isis, 102(4):793–794, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/664884.

Steinhauser:2011:BRD

[589] Thomas Steinhauser. Book review: Das Deutsche Museum in der Zeit des Nationalsozialismus: Eine Bestandsaufnahme. Isis, 102(4):794–795, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664885.

Radick:2011:BRD

[590] Gregory Radick. Book review: Darwin’s Pictures: Views of Evolutionary Theory, 1837–1874. Isis, 102(4):795–796, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664886.

Roll-Hansen:2011:BRN

[591] Nils Roll-Hansen. Book review: The Nazi Symbiosis: Human Genetics and Politics in the Third Reich. Isis, 102(4):796–797, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664887.

Conway:2011:BRR

[592] Erik M. Conway. Book review: Revolutionary Atmosphere: The Story of the Altitude Wind Tunnel and the Space Power Chambers. Isis, 102 (4):797–798, December 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
664888.

Brain:2011:BRN

[593] Robert Michael Brain. Book review: Eine Naturgeschichte für das 21, by Safia Azzouni; Christina Brandt; Bernd Gausemeier; Julia Kursell; Henning Schmidgen; Barbara Wittmann. Jahrhundert: Hommage à/Zu Ehren von/In Honor of Hans-Jörg Rheinberger. Isis, 102(4):798–799, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664889.

Bocking:2011:BRW

[594] Stephen Bocking. Book review: Wired Wilderness: Technologies of Tracking and the Making of Modern Wildlife. Isis, 102(4):799–800, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/664890.

Govoni:2011:BRI

[595] Paola Govoni. Book review: Italo Calvino e la scienza: Gli alfabeti del mondo. Isis, 102(4):800–801, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/664891.

Witteveen:2011:BRH

[596] Joeri Witteveen. Book review: Haldane, Mayr, and Beanbag Genetics. Isis, 102(4):801–802, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664892.

Agar:2011:BRC

[597] Jon Agar. Book review: The Computer Boys Take Over: Computers, Programmers, and the Politics of Technical Expertise. Isis, 102(4):802– 803, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664893.

Jankovic:2011:BRF

[598] Vladimir Jankovic. Book review: Fixing the Sky: The Checkered History of Weather and Climate Control. Isis, 102(4):803–804, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664894.

Myers:2011:BRW

[599] Greg Myers. Book review: Weather on the Air: A History of Broadcast Meteorology. Isis, 102(4):805, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/664895.

Marche:2011:BRS

[600] Jordan D. Marché II. Book review: Space Exploration and Humanity: A Historical Encyclopedia. Isis, 102(4):805–806, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664896.

Gurshtein:2011:BRL

[601] Alexander A. Gurshtein. Book review: Lenin’s Laureate: Zhores Alferov’s Life in Communist Science. Isis, 102(4):807–808, December 2011. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664897.

McCook:2011:BRJ

[602] Stuart McCook. Book review: Jungle Laboratories: Mexican Peasants, National Projects, and the Making of the Pill. Isis, 102(4):808–809, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/664898.

Hamblin:2011:BRC

[603] Jacob Darwin Hamblin. Book review: Confluence: The Nature of Technology and the Remaking of the Rhône. Isis, 102(4):809–810, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/664899.

Eraso:2011:BRC

[604] Yolanda Eraso. Book review: The Cuban Cure: Reason and Resistance in Global Science. Isis, 102(4):810–811, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664900.

Hatch:2011:BRC

[605] Robert Alan Hatch. Book review: Clio Wired: The Future of the Past in the Digital Age. Isis, 102(4):811–812, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/664901.

Kuklick:2011:BRG

[606] Henrika Kuklick. Book review: Glimpses into My Own Black Box: An Exercise in Self-Deconstruction. Isis, 102(4):812–813, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664902.

Krige:2011:BRP

[607] John Krige. Book review: Physics and Politics: Research and Research Support in Twentieth Century Germany in International Perspective. Isis, 102(4):813–814, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664903.

Chimisso:2011:BRG

[608] Cristina Chimisso. Book review: Gaston Bachelard et les études critiques de la science. Isis, 102(4):814–815, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664904.

Mazzotti:2011:BRT

[609] Massimo Mazzotti. Book review: Tacit and Explicit Knowledge. Isis, 102 (4):815–816, December 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
664905.

Gere:2011:BRA

[610] Cathy Gere. Book review: Andrew Pickering: The Cybernetic Brain: Sketches of Another Future. Isis, 102(4):816–818, December 2011. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664906.

Stahnisch:2011:BRA

[611] Frank W. Stahnisch. Book review: Affect and Artificial Intelligence. Isis, 102(4):818–819, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664907.

Anonymous:2011:NCd

[612] Anonymous. Notes on contributors. Isis, 102(4):820–824, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/663619.

Anonymous:2011:IIV

[613] Anonymous. Index to Isis, volume 102, 2011. Isis, 102(4):825–835, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/664835.

Anonymous:2011:CV

[614] Anonymous. Contents, volume 102. Isis, 102(4):839–841, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664933.

Anonymous:2011:Fd

[615] Anonymous. Frontispiece. Isis, 102(4):vi, December 2011. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664834.

Anonymous:2012:CBH

[616] Anonymous. Current bibliography of the history of science and its cultural influences, 2012. Isis, 103(S1):i–317, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670150.

Broman:2012:MEP

[617] Thomas Broman. Metaphysics for an enlightened public: The controversy over monads in Germany, 1746–1748. Isis, 103(1):1–23, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/664974.

Eskildsen:2012:LOC

[618] Kasper Risbjerg Eskildsen. The language of objects: Christian Jürgensen Thomsen’s science of the past. Isis, 103(1):24–53, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664975.

Solovey:2012:SFH

[619] Mark Solovey. Senator Fred Harris’s National Social Science Foundation Proposal: Reconsidering Federal science policy, natural science–social science relations, and American liberalism during the 1960s. Isis, 103(1): 54–82, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664976.

Vicedo:2012:ISL

[620] Marga Vicedo. Introduction: The secret lives of textbooks. Isis, 103(1): 83–87, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664979.

Gordin:2012:TTR

[621] Michael D. Gordin. Translating textbooks: Russian, German, and the language of chemistry. Isis, 103(1):88–98, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/664980.

Shapiro:2012:BTP

[622] Adam R. Shapiro. Between training and popularization: Regulating science textbooks in secondary education. Isis, 103(1):99–110, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/664981.

Vicedo:2012:PGP

[623] Marga Vicedo. Playing the game: Psychology textbooks speak out about love. Isis, 103(1):111–125, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/664982.

Kaiser:2012:TTT

[624] David Kaiser. A tale of two textbooks: Experiments in genre. Isis, 103(1):126–138, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
664983.

Sloan:2012:EEM

[625] Phillip R. Sloan. Eloge: Ernan McMullin, 1924–2011. Isis, 103(1):139– 143, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664984.

Moore:2012:EJC

[626] James Moore. Eloge: John Colton Greene, 1917–2008. Isis, 103(1):144– 148, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/664985.

Push:2012:LE

[627] Stephen Push. Letters to the Editor. Isis, 103(1):149, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664986.

Nicholson:2012:LE

[628] Ian Nicholson. Letters to the Editor. Isis, 103(1):149–150, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/666394.

Hall:2012:BRB

[629] Bert S. Hall. Book review: The Book of Michael of Rhodes: A Fifteenth-Century Maritime Manuscript. Isis, 103(1):151–154, March 2012. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664988.

Mukerji:2012:BRT

[630] Chandra Mukerji. Book review: Too Much to Know: Managing Scholarly Information before the Modern Age. Isis, 103(1):155–156, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/666395.

Weldon:2012:BRS

[631] Stephen P. Weldon. Book review: Science and Religion around the World. Isis, 103(1):156–157, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666396.

Gingras:2012:BRH

[632] Yves Gingras. Book review: Histories of Scientific Observation. Isis, 103(1):157–158, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666397.

Shteir:2012:BRP

[633] Ann Shteir. Book review: Planting Paradise: Cultivating the Garden, 1501–1900. Isis, 103(1):158–159, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666398.

Henry:2012:BRW

[634] John Henry. Book review: Wrestling with Nature: From Omens to Science. Isis, 103(1):159–160, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/666399.

Borck:2012:BRD

[635] Cornelius Borck. Book review: Daten sichern: Schreiben und Zeichnen als Verfahren der Aufzeichnung Spuren erzeugen: Zeichnen und Schreiben als Verfahren der Selbstaufzeichnung Notieren, Skizzieren: Schreiben und Zeichnen als Verfahren des Entwurfs. Isis, 103(1):160– 161, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/666400.

Johnson:2012:BRD

[636] Ann Johnson. Book review: David Kaiser, ed., Becoming MIT: Moments of Decision. Isis, 103(1):161–162, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666401.

Kidwell:2012:BRS

[637] Peggy Aldrich Kidwell and Mark E. Kidwell. Book review: Sharon Bertsch McGrayne, The Theory That Would Not Die: How Bayes’ Rule Cracked the Enigma Code, Hunted Down Russian Submarines, and Emerged Triumphant from Two Centuries of Controversy. Isis, 103 (1):162–163, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666402.

Gavrus:2012:BRI

[638] Delia Gavrus. Book review: Imagining Illness: Public Health and Visual Culture. Isis, 103(1):163–164, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666403.

Freudenthal:2012:BRE

[639] Gad Freudenthal. Book review: Evidence and Interpretation in Studies on Early Science and Medicine: Essays in Honor of John E, by Edith Dudley Sylla; William R. Newman. Murdoch. Isis, 103(1):164– 166, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/666404.

Pollock:2012:BRH

[640] Anne Pollock. Book review: How Cancer Crossed the Color Line. Isis, 103(1):166–167, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666405.

Falcon:2012:BRH

[641] Andrea Falcon. Book review: Heavenly Stuff: The Constitution of the Celestial Objects and the Theory of Homocentric Spheres in Aristotle’s Cosmology. Isis, 103(1):167, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666406.

Kirschner:2012:BRN

[642] Stefan Kirschner. Book review: Nicole Oresme: Questiones super geometriam Euclidis. Isis, 103(1):168, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666407.

Yeo:2012:BRP

[643] Richard Yeo. Book review: Pliny’s Encyclopedia: The Reception of the Natural History. Isis, 103(1):169, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666408.

Topper:2012:BRB

[644] David R. Topper. Book review: Between Raphael and Galileo: Mutio Oddi and the Mathematical Culture of Late Renaissance Italy. Isis, 103 (1):169–171, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666409.

Bailey:2012:BRD

[645] Michael D. Bailey. Book review: Le diable au sabbat: Littérature démonologique et sorcellerie (1440–1460). Isis, 103(1):171–172, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/666410.

Beretta:2012:BRM

[646] Marco Beretta. Book review: Making and Marketing Medicine in Renaissance Florence. Isis, 103(1):172–173, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666411.

Volkov:2012:BRG

[647] Alexei Volkov and Yi-Long Huang. Book review: Granting the Seasons: The Chinese Astronomical Reform of 1280, with a Study of Its Many Dimensions and an Annotated Translation of Its Records. Isis, 103(1): 173, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/666412.

Sylla:2012:BRL

[648] Edith Dudley Sylla. Book review: Le Liber mahameleth: Édition critique et commentaires. Isis, 103(1):173–174, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/666413.

Schmitt:2012:BRS

[649] Stéphane Schmitt. Book review: Science, libertinage et clandestinité à l’aube des Lumières: Le transformisme de Telliamed Telliamed. Isis, 103(1):174–175, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666414.

Wakefield:2012:BRD

[650] Andre Wakefield. Book review: Divining Science: Treasure Hunting and Earth Science in Early Modern Germany. Isis, 103(1):175–177, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/666415.

Walker:2012:BRF

[651] Matthew Walker. Book review: François Blondel: Architecture, Erudition, and the Scientific Revolution. Isis, 103(1):177–178, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/666416.

Sheffield:2012:BRN

[652] Suzanne Le-May Sheffield. Book review: The New Science and Women’s Literary Discourse: Prefiguring Frankenstein. Isis, 103(1):178–179, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/666417.

Brentjes:2012:BRI

[653] Sonja Brentjes. Book review: Intellectual Curiosity and the Scientific Revolution: A Global Perspective. Isis, 103(1):179–180, March 2012.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/666418.

Daniel:2012:BRB

[654] Dane T. Daniel. Book review: Boyle: Between God and Science. Isis, 103(1):180–181, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666419.

Snobelen:2012:BRN

[655] Stephen D. Snobelen. Book review: Newton as Philosopher. Isis, 103 (1):181–182, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666420.

Cavazza:2012:BRL

[656] Marta Cavazza. Book review: The Lady Anatomist: The Life and Work of Anna Morandi Manzolini. Isis, 103(1):182–183, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/666421.

Gaukroger:2012:BRD

[657] Stephen Gaukroger. Book review: Divine Machines: Leibniz and the Sciences of Life. Isis, 103(1):184–185, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666422.

McClelland:2012:BRU

[658] Charles E. McClelland. Book review: Universität im Umbruch: Universität und Wissenschaft im Spannungsfeld der Gesellschaft um 1800. Isis, 103(1):185, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666423.

Fraser:2012:BRU

[659] Craig Fraser. Book review: Unravelling Starlight: William and Margaret Huggins and the Rise of the New Astronomy. Isis, 103(1):186–187, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/666424.

Normandin:2012:BRU

[660] Sebastian Normandin. Book review: Unruly Spirits: The Science of Psychic Phenomena in Modern France. Isis, 103(1):187–188, March 2012.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/666425.

Zilberstein:2012:BRB

[661] Anya Zilberstein. Book review: The British Arboretum: Trees, Science, and Culture in the Nineteenth Century. Isis, 103(1):188–189, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/666426.

Paul:2012:BRD

[662] Harry Paul. Book review: Dying on the Vine: How Phylloxera Transformed Wine. Isis, 103(1):189–190, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666427.

Theunissen:2012:BRF

[663] Bert Theunissen. Book review: The Floracrats: State-Sponsored Science and the Failure of the Enlightenment in Indonesia. Isis, 103(1):190– 191, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/666428.

French:2012:BRW

[664] Steven French. Book review: What Is Life? The Intellectual Pertinence of Erwin Schrödinger. Isis, 103(1):191, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/666429.

Barberis:2012:BRS

[665] Daniela Barberis. Book review: The Science of History in Victorian Britain: Making the Past Speak. Isis, 103(1):191–192, March 2012. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/666430.

Ratcliff:2012:BRB

[666] Jessica Ratcliff. Book review: British University Observatories, 1772– 1939. Isis, 103(1):192–193, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666431.

Rothenberg:2012:BRG

[667] Marc Rothenberg. Book review: Geographies of Mars: Seeing and Knowing the Red Planet. Isis, 103(1):193–194, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/666432.

Barrow:2012:BRD

[668] Mark V. Barrow, Jr. Book review: A Democracy of Facts: Natural History in the Early Republic. Isis, 103(1):194–195, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/666433.

Burkhardt:2012:BRQ

[669] Richard W. Burkhardt, Jr. Book review: Qu’est-ce que le néolamarckisme? Les biologistes français et la question de l’évolution des espèces. Isis, 103(1):195–196, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666434.

Hu:2012:BRE

[670] Danian Hu. Book review: Early China Coast Meteorology: The Role of Hong Kong. Isis, 103(1):196–197, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666435.

Cahan:2012:BRH

[671] David Cahan. Book review: Helmholtz: From Enlightenment to Neuroscience. Isis, 103(1):197–199, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666436.

Rasmussen:2012:BRH

[672] Nicolas Rasmussen. Book review: Hormones of Life: Endocrinology, the Pharmaceutical Industry, and the Dream of a Remedy for Sterility, 1930–1970. Isis, 103(1):199–200, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666437.

Harris:2012:BRM

[673] Martha Harris. Book review: Marie Curie: A Biography. Isis, 103(1): 200, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/666438.

Morus:2012:BRT

[674] Iwan Rhys Morus. Book review: The Transit of Venus Enterprise in Victorian Britain. Isis, 103(1):200–202, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666439.

Steigerwald:2012:BRM

[675] Joan Steigerwald. Book review: Matters of Spirit: J, by F. Scott Scribner. G. Fichte and the Technological Imagination. Isis, 103(1):202–203, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/666440.

Harper:2012:BRE

[676] Kristine C. Harper. Book review: Engineering Nature: Water, Development, and the Global Spread of American Environmental Expertise. Isis, 103(1):203–204, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666441.

Engstrom:2012:BRJ

[677] Eric J. Engstrom. Book review: John W, by Paul J. Weindling. Thompson: Psychiatrist in the Shadow of the Holocaust. Isis, 103(1):204–205, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/666442.

Cheung:2012:BRF

[678] Tobias Cheung. Book review: Die Form des Werdens: Eine Kulturgeschichte der Embryologie, 1760–1830. Isis, 103(1):205–206, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/666443.

Elliott:2012:BRN

[679] Clark A. Elliott. Book review: The North American Journals of Maximilian of Wied The North American Journals of Maximilian of Wied. Isis, 103(1):206–207, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666444.

Rutherford:2012:BRH

[680] Alexandra Rutherford. Book review: The History of the Social Sciences since 1945. Isis, 103(1):207–208, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666445.

Buhs:2012:BRDa

[681] Joshua Blu Buhs. Book review: The Discovery of the Sasquatch: Reconciling Culture, History, and Science in the Discovery Process. Isis, 103(1):208–209, March 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
666446.

Watkins:2012:BRT

[682] Elizabeth Siegel Watkins. Book review: Toxic Bodies: Hormone Disruptors and the Legacy of DES. Isis, 103(1):209–210, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/666447.

Mody:2012:BRM

[683] Cyrus C. M. Mody. Book review: Makers of the Microchip: A Documentary History of Fairchild Semiconductor. Isis, 103(1):210–211, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/666448.

Fuller:2012:BRS

[684] Steve Fuller. Book review: Science-Mart: Privatizing American Science. Isis, 103(1):211–212, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/666449.

Dupre:2012:BRN

[685] John Dupré. Book review: Naked Genes: Reinventing the Human in the Molecular Age. Isis, 103(1):212–213, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666450.

Tjossem:2012:BRG

[686] Sara Tjossem. Book review: G, by Nancy G. Slack; David K. Skelly; David M. Post; Melinda D. Smith. Evelyn Hutchinson and the Invention of Modern Ecology The Art of Ecology: Writings of G. Evelyn Hutchinson. Isis, 103(1):213–214, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/666451.

Borrello:2012:BRS

[687] Mark E. Borrello. Book review: The Science and Humanism of Stephen Jay Gould. Isis, 103(1):214–215, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666452.

Leung:2012:BRO

[688] Angela Ki Che Leung. Book review: Other-Worldly: Making Chinese Medicine through Transnational Frames. Isis, 103(1):215–216, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/666453.

Porter:2012:BRS

[689] Theodore M. Porter. Book review: La science économique comme ingénierie: Quantification et modélisation. Isis, 103(1):216–217, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/666454.

Boantza:2012:BRS

[690] Victor D. Boantza. Book review: The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms. Isis, 103(1):217–218, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/666455.

Tracy:2012:BRK

[691] Sarah Tracy. Book review: Knowing Nature: Conversations at the Intersection of Political Ecology and Science Studies. Isis, 103(1):218–219, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/666456.

Golinski:2012:BRH

[692] Jan Golinski. Book review: How Well Do Facts Travel? The Dissemination of Reliable Knowledge. Isis, 103(1):219–220, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/666457.

Watts:2012:BRN

[693] Fraser Watts. Book review: Natural Reflections: Human Cognition at the Nexus of Science and Religion Science and Spirituality: Making Room for Faith in the Age of Science. Isis, 103(1):220–222, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/666458.

Richards:2012:BRD

[694] Richard A. Richards. Book review: Did Darwin Write the Origin Backwards? Philosophical Essays on Darwin’s Theory. Isis, 103(1):222–223, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/666459.

Anonymous:2012:NCa

[695] Anonymous. Notes on contributors. Isis, 103(1):224–228, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/664990.

Anonymous:2012:Fa

[696] Anonymous. Frontispiece. Isis, 103(1):vi, March 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/666392.

Anderson:2012:HRS

[697] Warwick Anderson. Hybridity, race, and science: The voyage of the Zaca, 1934–1935. Isis, 103(2):229–253, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666354.

Stanley:2012:PPA

[698] Matthew Stanley. Predicting the past: Ancient eclipses and Airy, Newcomb, and Huxley on the authority of science. Isis, 103(2):254–277, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/666355.

Theunissen:2012:BNP

[699] Bert Theunissen. Breeding for nobility or for production? Cultures of dairy cattle breeding in The Netherlands, 1945–1995. Isis, 103(2):278– 309, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/666356.

Andersen:2012:MTN

[700] Casper Andersen, Jakob Bek-Thomsen, and Peter C. Kjærgaard. The money trail: A new historiography for networks, patronage, and scientific careers. Isis, 103(2):310–315, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/666357.

Edgerton:2012:TMH

[701] David Edgerton. Time, money, and history. Isis, 103(2):316–327, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/666358.

MacLeod:2012:REP

[702] Christine MacLeod. Reluctant entrepreneurs: Patents and state patronage in new technosciences, circa 1870–1930. Isis, 103(2):328–339, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/666359.

Kjaergaard:2012:FTP

[703] Peter C. Kjærgaard. The fossil trade: Paying a price for human origins. Isis, 103(2):340–355, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/666365.

Anonymous:2012:HSS

[704] Anonymous. History of science society annual meeting, 2011. Isis, 103(2): 356–362, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667467.

vonArx:2012:EFM

[705] Jeffrey von Arx, S.J. and Bernard Lightman. Eloge: Frank M. Turner, 1944–2010. Isis, 103(2):362–365, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/667468.

Glick:2012:EJV

[706] Thomas F. Glick. Eloge: Joan Vernet Ginés, 1923–2011. Isis, 103(2): 365–367, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667469.

Bolt:2012:EMK

[707] Marvin Bolt. Eloge: Marjorie Kelly Webster, 1915–2011. Isis, 103(2): 367–371, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667470.

Brentjes:2012:EHW

[708] Sonja Brentjes. Eloge: Hans Wußing, 1927– 2011. Isis, 103(2):371– 375, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667471.

Findlen:2012:BRC

[709] Paula Findlen and Hannah Marcus. Book review: Catholic Church and Modern Science: Documents from the Archives of the Roman Congregations of the Holy Office and the Index. Isis, 103(2):376–382, June 2012.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/666367.

Mancosu:2012:BRK

[710] Paolo Mancosu. Book review: Kurt Gödel and the Foundations of Mathematics: Horizons of Truth. Isis, 103(2):383–384, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/667472.

Rieger:2012:BRR

[711] Matthias Rieger. Book review: Reason and Resonance: A History of Modern Aurality. Isis, 103(2):384, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/667473.

Keating:2012:BRS

[712] Jessica Keating. Book review: Sublime Dreams of Living Machines: The Automaton in the European Imagination. Isis, 103(2):385–386, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/667474.

Timmermann:2012:BRP

[713] Carsten Timmermann. Book review: Preventive Strikes: Women, Precancer, and Prophylactic Surgery. Isis, 103(2):386–387, June 2012. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667475.

Knight:2012:BRS

[714] David Knight. Book review: Science Secrets: The Truth about Darwin’s Finches, Einstein’s Wife, and Other Myths. Isis, 103(2):387, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667476.

Rothstein:2012:BRB

[715] William G. Rothstein. Book review: The Bleeding Disease: Hemophilia and the Unintended Consequences of Medical Progress. Isis, 103(2):387– 388, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667477.

Anker:2012:BRI

[716] Peder Anker. Book review: The Illusory Boundary: Environment and Technology in History. Isis, 103(2):388–389, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/667478.

Durbach:2012:BRA

[717] Nadja Durbach. Book review: Anatomy as Spectacle: Public Exhibitions of the Body from 1700 to the Present. Isis, 103(2):389–390, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667479.

Selin:2012:BRW

[718] Helaine Selin. Book review: What Is Medicine? Western and Eastern Approaches to Healing. Isis, 103(2):390–391, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/667480.

Akasoy:2012:BRB

[719] Anna Akasoy. Book review: Bedrohte Wahrheit: Der Islam und die modernen Naturwissenschaften. Isis, 103(2):391–392, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/667481.

Steele:2012:BRL

[720] John Steele. Book review: Looking at It from Asia: The Processes That Shaped the Sources of History of Science. Isis, 103(2):392–393, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/667482.

Hahm:2012:BRC

[721] David E. Hahm. Book review: The Cambridge History of Philosophy in Late Antiquity. Isis, 103(2):393–395, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/667483.

Schubert:2012:BRG

[722] Charlotte Schubert. Book review: Galen and the World of Knowledge. Isis, 103(2):395–396, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/667484.

Schmidl:2012:BRG

[723] Petra G. Schmidl. Book review: Germania und die Insel Thule: Die Entschlüsselung von Ptolemaios’ “Atlas der Oikumene”, by Andreas Kleineberg; Christian Marx; Eberhard Knobloch; Dieter Lelgemann.

Isis, 103(2):396, June 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
667485.

DeGregorio:2012:BRN

[724] Scott DeGregorio. Book review: “On the Nature of Things” and “On Times”, by Bede; Calvin B. Kendall; Faith Wallis. Isis, 103(2):396– 397, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667486.

Long:2012:BRF

[725] Pamela O. Long. Book review: Fra Mauro’s Mappa Mundi and Fifteenth-Century Venice. Isis, 103(2):397–398, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/667487.

Caroti:2012:BRN

[726] Stefano Caroti. Book review: The Nature of Natural Philosophy in the Late Middle Ages. Isis, 103(2):398–399, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/667488.

Finocchiaro:2012:BRS

[727] Maurice A. Finocchiaro. Book review: Scienza e religione: Scritti copernicani Libertas philosophandi: Teologia e filosofia nella lettera alla granduchessa Cristina di Lorena di Galileo Galilei. Isis, 103(2):400, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/667489.

Seitz:2012:BRC

[728] Jonathan Seitz. Book review: Conflicting Duties: Science, Medicine, and Religion in Rome, 1550–1750. Isis, 103(2):401–402, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/667490.

Stewart:2012:BRM

[729] Larry Stewart. Book review: Measure of the Earth: The Enlightenment Expedition That Reshaped the World. Isis, 103(2):402–403, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667491.

Bucciantini:2012:BRD

[730] Massimo Bucciantini. Book review: Defending Copernicus and Galileo: Critical Reasoning in the Two Affairs. Isis, 103(2):403–404, June 2012.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667492.

Gomez:2012:BRU

[731] Susana Gómez. Book review: The University of Mantua, the Gonzaga, and the Jesuits, 1584–1630. Isis, 103(2):404–405, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/667493.

Sarasohn:2012:BRR

[732] Lisa T. Sarasohn. Book review: Richard Bentley: Poetry and Enlightenment. Isis, 103(2):405–406, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/667494.

Broman:2012:BRV

[733] Thomas Broman. Book review: Vorlesungen zur Naturlehre: Notizen und Materialien zur Experimentalphysik, Teil I. Isis, 103(2):406–407, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/667495.

Pingyi:2012:BRC

[734] Chu Pingyi. Book review: The Crafting of the Ten Thousand Things: Knowledge and Technology in Seventeenth-Century China. Isis, 103(2): 407–408, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667496.

Bennett:2012:BRO

[735] Jim Bennett. Book review: The Origins of the Telescope. Isis, 103(2): 408–410, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667497.

Hu:2012:BRC

[736] Minghui Hu. Book review: Chuanbo yu huitong: Qiqi tushuo yanjiu yu jiaoyi, by Baichun Zhang; Miao Tian; Matthias Schemmel; Jürgen Renn; Peter Damerow. [Transmission and Integration: “Qiqi tushuo.”]. Isis, 103(2):410–411, June 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
667498.

Allen:2012:BRA

[737] Garland E. Allen. Book review: American Abyss: Savagery and Civilization in the Age of Industry. Isis, 103(2):411–412, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/667499.

Sera-Shriar:2012:BRV

[738] Efram Sera-Shriar. Book review: A Vision of Modern Science: John Tyndall and the Role of the Scientist in Victorian Culture. Isis, 103(2): 412–413, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667500.

Johnson:2012:BRS

[739] Jeffrey Allan Johnson. Book review: Selbstmobilisierung der Wissenschaft: Technische Hochschulen im “Dritten Reich”, by Noyan Dinçkal; Christof Dipper; Detlev Mares. Isis, 103(2):413–414, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667501.

Schirrmacher:2012:BRF

[740] Arne Schirrmacher. Book review: “Fremde” Wissenschaftler im Dritten Reich: Die Debye–Affäre im Kontext. Isis, 103(2):414–415, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667502.

Schwach:2012:BRE

[741] Vera Schwach. Book review: An Empire of Ice: Scott, Shackleton, and the Heroic Age of Antarctic Science. Isis, 103(2):415–416, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667503.

Pickstone:2012:BRI

[742] John Pickstone. Book review: An Infinity of Things: How Sir Henry Wellcome Collected the World. Isis, 103(2):416–417, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/667504.

Dettelbach:2012:BRA

[743] Michael Dettelbach. Book review: Alexander von Humboldt und Cotta: Briefwechsel Ein “Diplomat aus den Wäldern des Orinoko”: Alexander von Humboldt als Mittler zwischen Preußen und Frankreich. Isis, 103(2): 417, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667505.

Connor:2012:BRW

[744] J. T. H. Connor. Book review: War’s Waste: Rehabilitation in World War I America. Isis, 103(2):419–420, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/667506.

Meskill:2012:BRO

[745] David Meskill. Book review: Ökonomische Effizienz und gesellschaftlicher Ausgleich: Die industrielle Psychotechnik in der Weimarer Republik. Isis, 103(2):420, June 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
667507.

Stanley:2012:BRP

[746] Matthew Stanley. Book review: The Philosophical Breakfast Club: Four Remarkable Friends Who Transformed Science and Changed the World. Isis, 103(2):421, June 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
667508.

Worboys:2012:BRA

[747] Michael Worboys. Book review: Africa as a Living Laboratory: Empire, Development, and the Problem of Scientific Knowledge, 1870–1950. Isis, 103(2):421–423, June 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
667509.

Low:2012:BRT

[748] Morris Low. Book review: Technology of Empire: Telecommunications and Japanese Expansion in Asia, 1883–1945. Isis, 103(2):423–424, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/667510.

Bassett:2012:BRN

[749] Ross Bassett. Book review: Nucleus and Nation: Scientists, International Networks, and Power in India. Isis, 103(2):424–425, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667511.

Fortun:2012:BRM

[750] Mike Fortun. Book review: Mixing Races: From Scientific Racism to Modern Evolutionary Ideas. Isis, 103(2):425, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/667512.

Rohde:2012:BRC

[751] Joy Rohde. Book review: The Contours of America’s Cold War. Isis, 103(2):425–426, June 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
667513.

Siddiqi:2012:BRR

[752] Asif A. Siddiqi. Book review: Red Cloud at Dawn: Truman, Stalin, and the End of the Atomic Monopoly. Isis, 103(2):426–427, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667514.

Bonah:2012:BRM

[753] Christian Bonah. Book review: La médecine des preuves: Une histoire de l’expérimentation thérapeutique par essais cliniques contrôlés. Isis, 103(2):427–428, June 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
667515.

Butrica:2012:BRJ

[754] Andrew J. Butrica. Book review: John F, by John M. Logsdon. Kennedy and the Race to the Moon. Isis, 103(2):428–429, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/667516.

Gregory:2012:BRD

[755] Frederick Gregory. Book review: Darwin’s Pious Idea: Why the Ultra-Darwinists and Creationists Both Get It Wrong. Isis, 103(2):429–431, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/667517.

Gaukroger:2012:BRS

[756] Stephen Gaukroger. Book review: Science as Cultural Practice, by Moritz Epple; Claus Zittel. Volume 1: Cultures and Politics of Research from the Early Modern Period to the Age of Extremes. Isis, 103(2):431– 432, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667518.

Abounader:2012:BRE

[757] Brooke Abounader. Book review: Exploring the Scientific Method: Cases and Questions. Isis, 103(2):432–433, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/667519.

Spencer:2012:BRC

[758] Mark G. Spencer. Book review: Common Sense: A Political History. Isis, 103(2):433–434, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/667520.

Anonymous:2012:NCb

[759] Anonymous. Notes on contributors. Isis, 103(2):435–437, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/666369.

Anonymous:2012:Fb

[760] Anonymous. Frontispiece. Isis, 103(2):vi, June 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/667466.

Cantor:2012:SPP

[761] Geoffrey Cantor. Science, providence, and progress at the great exhibition. Isis, 103(3):439–459, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/667968.

Rieppel:2012:BDB

[762] Lukas Rieppel. Bringing dinosaurs back to life: Exhibiting prehistory at the American museum of natural history. Isis, 103(3):460–490, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/667969.

Siraisi:2012:MHS

[763] Nancy G. Siraisi. Medicine, 1450–1620, and the history of science. Isis, 103(3):491–514, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/667970.

Bud:2012:I

[764] Robert Bud. Introduction. Isis, 103(3):515–517, September 2012. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667971.

Alexander:2012:TAA

[765] Jennifer Karns Alexander. Thinking again about science in technology. Isis, 103(3):518–526, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/667975.

Lucier:2012:OPA

[766] Paul Lucier. The origins of pure and applied science in gilded age America. Isis, 103(3):527–536, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/667976.

Bud:2012:ASP

[767] Robert Bud. “Applied science”: A phrase in search of a meaning. Isis, 103 (3):537–545, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
667977.

Gooday:2012:VAH

[768] Graeme Gooday. “Vague and artificial”: The historically elusive distinction between pure and applied science. Isis, 103(3):546–554, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/667978.

Schatzberg:2012:AAS

[769] Eric Schatzberg. From art to applied science. Isis, 103(3):555–563, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/667979.

DePace:2012:LE

[770] Anna De Pace. Letters to the Editor. Isis, 103(3):564, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/667980.

Goddu:2012:LE

[771] André Goddu. Letters to the Editor. Isis, 103(3):564–565, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/668916.

Nyhart:2012:BRA

[772] Lynn K. Nyhart. Book review: The Afterlives of Animals: A Museum Menagerie. Isis, 103(3):566–567, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/668967.

Olshin:2012:BRM

[773] Benjamin B. Olshin. Book review: Maya Creation Myths: Words and Worlds of the Chilam Balam Maya Daykeeping: Three Calendars from Highland Guatemala. Isis, 103(3):567–568, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/668968.

Crowther:2012:BRS

[774] Kathleen Crowther. Book review: Secrets and Knowledge in Medicine and Science, 1500–1800. Isis, 103(3):568–569, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/668969.

Shinno:2012:BRR

[775] Reiko Shinno. Book review: Reproducing Women: Medicine, Metaphor, and Childbirth in Late Imperial China. Isis, 103(3):569–570, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/668970.

Raynaud:2012:BRF

[776] Dominique Raynaud. Book review: Florence and Baghdad: Renaissance Art and Arab Science. Isis, 103(3):570–572, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/668971.

Truitt:2012:BRH

[777] E. R. Truitt. Book review: Histoires prodigieuses (Edition de 1561). Isis, 103(3):572–573, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/668972.

Spruyt:2012:BRI

[778] Joke Spruyt. Book review: Insolubilia. Isis, 103(3):573, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/668973.

Holtz:2012:BRL

[779] Grégoire Holtz. Book review: L’histoire naturelle au XVIe siècle: Introduction, étude et édition critique de La nature et diversité des poissons (1555) de Pierre Belon. Isis, 103(3):573–574, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/668974.

Boner:2012:BRC

[780] Patrick J. Boner. Book review: Copernicus and the Aristotelian Tradition: Education, Reading, and Philosophy in Copernicus’s Path to Heliocentrism. Isis, 103(3):574–575, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/668975.

Forshaw:2012:BRM

[781] Peter Forshaw. Book review: Misticall Wordes and Names Infinite: An Edition and Study of Humfrey Lock’s Treatise on Alchemy. Isis, 103 (3):575–577, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
668976.

Martin:2012:BRS

[782] Catherine Gimelli Martin. Book review: Shakespeare, Alchemy, and the Creative Imagination: The Sonnets and A Lover’s Complaint. Isis, 103(3):577, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
668977.

Bamji:2012:BRT

[783] Alexandra Bamji. Book review: Theaters of Anatomy: Students, Teachers, and Traditions of Dissection in Renaissance Venice. Isis, 103(3): 577–578, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
668978.

Laird:2012:BRD

[784] W. R. Laird. Book review: La via delle acque (1500–1700): Appropriazione delle arti e trasformazione delle matematiche. Isis, 103(3): 578–580, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
668979.

Rapisarda:2012:BRG

[785] Stefano Rapisarda. Book review: Guillaume l’Anglais, le frondeur de l’uroscopie médiévale (XIIIe siècle): Édition commentée et traduction du De urina non visa. Isis, 103(3):580, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/668980.

Stevens:2012:BRM

[786] Wesley M. Stevens. Book review: Mathématiques et connaissance du monde réel avant Galilée. Isis, 103(3):580–582, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/668981.

Rey-Bueno:2012:BRM

[787] Mar Rey-Bueno and Maria M. Portuondo. Book review: La medicina magnética: Del ungüento armario al polvo simpático de Kenelm Digby. Isis, 103(3):582, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/668982.

Cavalli:2012:BRT

[788] Fabio Cavalli. Book review: Tractatus de humido radicali. Isis, 103 (3):583, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
668983.

Reeves:2012:BRG

[789] Eileen Reeves. Book review: Galileo’s O, by Horst Bredekamp; Irene Brückle; Oliver Hahn; Paul Needham. Volume 1: The Sidereus Nuncius of 1610: A Comparison of the Proof Copy (New York) and Other Paradigmatic Copies (Graz, Paris, Münster). Isis, 103(3):583–585, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/668984.

Shea:2012:BRL

[790] William R. Shea. Book review: Luigi Galvani: Devozione, scienza e rivoluzione. Isis, 103(3):585–586, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/668985.

Ziegler:2012:BRK

[791] Rafael Ziegler. Book review: Kant and the Human Sciences: Biology, Anthropology, and History. Isis, 103(3):586–587, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/668986.

Beretta:2012:BRS

[792] Marco Beretta. Book review: Silent Messengers: The Circulation of Material Objects of Knowledge in the Early Modern Low Countries. Isis, 103 (3):587–588, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
668987.

Valleriani:2012:BRI

[793] Matteo Valleriani. Book review: The Invention of Discovery, 1500–1700. Isis, 103(3):588–589, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/668988.

Gorham:2012:BRL

[794] Geoffrey Gorham. Book review: Leibniz: Body, Substance, Monad. Isis, 103(3):589–590, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/668989.

Anderson:2012:BRC

[795] Katharine Anderson. Book review: Confronting the Climate: British Airs and the Making of Environmental Medicine. Isis, 103(3):590–591, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/668990.

Boner:2012:BRT

[796] Patrick J. Boner. Book review: Taming the Prophets: Astrology, Orthodoxy, and the Word of God in Early Modern Sweden. Isis, 103(3): 591–592, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
668991.

Donato:2012:BRM

[797] Maria Pia Donato. Book review: Mechanism, Experiment, Disease: Marcello Malpighi and Seventeenth-Century Anatomy. Isis, 103(3):592–593, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/668992.

Gossin:2012:BRM

[798] Pamela Gossin. Book review: The Machine in the Text: Science and Literature in the Age of Shakespeare and Galileo. Isis, 103(3):593–594, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/668993.

Spary:2012:BRN

[799] E. C. Spary. Book review: A Natural History of Revolution: Violence and Nature in the French Revolutionary Imagination, 1789–1794. Isis, 103 (3):594–595, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
668994.

Protschky:2012:BRL

[800] Susie Protschky. Book review: The Life of Governor Joan Gideon Loten (1710–1789): A Personal History of a Dutch Virtuoso. Isis, 103(3): 595–596, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
668995.

Zepeda:2012:BRM

[801] Joseph Zepeda. Book review: Matter Matters: Metaphysics and Methodology in the Early Modern Period. Isis, 103(3):596–597, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/668996.

Pozzo:2012:BRK

[802] Riccardo Pozzo. Book review: Kant und die Wissenschaften vom Menschen. Isis, 103(3):597–598, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/668997.

Dooley:2012:BRS

[803] Brendan Dooley. Book review: Selling Science in the Age of Newton: Advertising and the Commoditization of Knowledge. Isis, 103(3):598–599, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/668998.

Stewart:2012:BRB

[804] Ian Stewart. Book review: The Body as Object and Instrument of Knowledge: Embodied Empiricism in Early Modern Science. Isis, 103 (3):599–600, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
668999.

Aso:2012:BRM

[805] Mitch Aso. Book review: Mixed Medicines: Health and Culture in French Colonial Cambodia. Isis, 103(3):600–601, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/669000.

Fritscher:2012:BRH

[806] Bernhard Fritscher. Book review: Hegels Naturphilosophie in der Dritten Moderne: Bestimmungen, Probleme und Perspektiven. Isis, 103(3): 601–602, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
669001.

Pemberton:2012:BRK

[807] Neil Pemberton. Book review: Karen Brown. Mad Dogs and Meerkats: A History of Resurgent Rabies in Southern Africa. Isis, 103(3):602–603, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/669002.

Evans:2012:BRJ

[808] James Evans. Book review: Jed Z. Buchwald; Diane Greco Josefowicz. The Zodiac of Paris: How an Improbable Controversy over an Ancient Egyptian Artifact Provoked a Modern Debate between Religion and Science. Isis, 103(3):603–604, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/669003.

Bocking:2012:BRS

[809] Stephen Bocking. Book review: Science and the Social Good: Nature, Culture, and Community, 1865–1965. Isis, 103(3):604–605, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/669004.

Wheaton:2012:BRW

[810] Bruce R. Wheaton. Book review: William and Lawrence Bragg, Father and Son: The Most Extraordinary Collaboration in Science. Isis, 103 (3):605–606, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
669005.

Huntelmann:2012:BRG

[811] Axel C. Hüntelmann. Book review: “Götter in deren Hand Donner und Blitz liegt”: Ausbildung und Forschung der Berliner Militärärzte von 1870 bis 1895. Isis, 103(3):606–607, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/669006.

Normandin:2012:BRI

[812] Sebastian Normandin. Book review: Investigating the Supernatural: From Spiritism and Occultism to Psychical Research and Metapsychics in France, 1853–1931. Isis, 103(3):607–608, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/669007.

Grattan-Guinness:2012:BRE

[813] I. Grattan-Guinness. Book review: The Evolution of Principia Mathematica: Bertrand Russell’s Manuscripts and Notes for the Second Edition. Isis, 103(3):608–609, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/669008.

Lavine:2012:BRR

[814] Matthew Lavine. Book review: Radioactivity: A History of a Mysterious Science. Isis, 103(3):609–610, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/669009.

Crenner:2012:BRM

[815] Christopher Crenner. Book review: A Modern History of the Stomach: Gastric Illness, Medicine, and British Society, 1800–1950. Isis, 103(3): 610–611, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
669010.

Burnham:2012:BRA

[816] John C. Burnham. Book review: American Madness: The Rise and Fall of Dementia Praecox. Isis, 103(3):611–612, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/669011.

James:2012:BRC

[817] Frank A. J. L. James. Book review: Communicating Physics: The Production, Circulation, and Appropriation of Ganot’s Textbooks in France and England, 1851–1887. Isis, 103(3):612–613, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/669012.

Pettit:2012:BRN

[818] Michael Pettit. Book review: Nature and Nurture in French Social Sciences, 1859–1914 and Beyond. Isis, 103(3):613–614, September 2012.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/669013.

Sullivan:2012:BRM

[819] Gerald Sullivan. Book review: Margaret Mead: A Biography. Isis, 103 (3):614, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
669014.

Thomas:2012:BRS

[820] William Thomas. Book review: A Short History of Physics in the American Century. Isis, 103(3):614–615, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/669015.

Buhs:2012:BRDb

[821] Joshua Blu Buhs. Book review: DDT and the American Century: Global Health, Environmental Politics, and the Pesticide That Changed the World. Isis, 103(3):615–616, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/669016.

Weingart:2012:BRL

[822] Peter Weingart. Book review: Lab Coats in Hollywood: Science, Scientists, and Cinema. Isis, 103(3):616–617, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/669017.

DeVorkin:2012:BRO

[823] David DeVorkin. Book review: Opening Space Research: Dreams, Technology, and Scientific Discovery. Isis, 103(3):617–618, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/669018.

Goodstein:2012:BRC

[824] Judith Goodstein. Book review: College of Science and Engineering: The Institute of Technology Years (1935–2010). Isis, 103(3):618–619, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/669019.

Hallion:2012:BRS

[825] Richard P. Hallion. Book review: The Spoken Word II: Recollections of Dryden History, Beyond the Sky. Isis, 103(3):619–620, September 2012.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/669020.

Milam:2012:BRM

[826] Erika Lorraine Milam. Book review: The Morning After: A History of Emergency Contraception in the United States. Isis, 103(3):620–621, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/669021.

Krige:2012:BRO

[827] John Krige. Book review: Our Own Worst Enemy? Institutional Interests and the Proliferation of Nuclear Weapons Expertise. Isis, 103 (3):621–622, September 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
669022.

Thurtle:2012:BRB

[828] Phillip Thurtle. Book review: Biology Is Technology: The Promise, Peril, and New Business of Engineering Life. Isis, 103(3):622, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/669023.

Rentetzi:2012:BRT

[829] Maria Rentetzi. Book review: Trading Zones and Interactional Expertise: Creating New Kinds of Collaboration. Isis, 103(3):623–624, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/669024.

Schaefer:2012:BRS

[830] Dagmar Schaefer. Book review: Science between Europe and Asia: Historical Studies on the Transmission, Adoption, and Adaptation of Knowledge. Isis, 103(3):624–625, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/669025.

Williams:2012:BRM

[831] Rosalind Williams. Book review: On the Modern Cult of the Factish Gods. Isis, 103(3):625–626, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/669026.

Halpern:2012:BRF

[832] Orit Halpern. Book review: The Freudian Robot: Digital Media and the Future of the Unconscious. Isis, 103(3):626–627, September 2012.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/669027.

Harker:2012:BRP

[833] David Harker. Book review: Philosophy of Science Matters: The Philosophy of Peter Achinstein. Isis, 103(3):627–628, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/669028.

Vidal:2012:BRA

[834] Fernando Vidal. Book review: Aping Mankind: Neuromania, Darwinitis, and the Misrepresentation of Humanity. Isis, 103(3):628–629, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/669029.

Haufe:2012:BRP

[835] Chris Haufe. Book review: Paleontology: A Philosophical Introduction. Isis, 103(3):629–630, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/669030.

Anonymous:2012:NCc

[836] Anonymous. Notes on contributors. Isis, 103(3):631–634, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/667982.

Anonymous:2012:Fc

[837] Anonymous. Frontispiece. Isis, 103(3):vi, September 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/669031.

Dawson:2012:PPR

[838] Gowan Dawson. Paleontology in parts: Richard Owen, William John Broderip, and the serialization of science in Early Victorian Britain. Isis, 103(4):637–667, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/668961.

Jutte:2012:TSJ

[839] Daniel Jütte. Trading in secrets: Jews and the early modern quest for clandestine knowledge. Isis, 103(4):668–686, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/668962.

Lienesch:2012:AEF

[840] Michael Lienesch. Abandoning evolution: The forgotten history of antievolution activism and the transformation of American social science. Isis, 103(4):687–709, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/668963.

Delbourgo:2012:I

[841] James Delbourgo and Staffan Müller-Wille. Introduction. Isis, 103(4): 710–715, December 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
669045.

Pugliano:2012:SLA

[842] Valentina Pugliano. Specimen lists: Artisanal writing or natural historical paperwork? Isis, 103(4):716–726, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/669049.

Keller:2012:NWS

[843] Vera Keller. The “new world of sciences”: The temporality of the research agenda and the unending ambitions of science. Isis, 103(4):727–734, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/669047.

Delbourgo:2012:LP

[844] James Delbourgo. Listing people. Isis, 103(4):735–742, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/669046.

Muller-Wille:2012:LRT

[845] Staffan Müller-Wille and Isabelle Charmantier. Lists as research technologies. Isis, 103(4):743–752, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/669048.

Herbert:2012:EPJ

[846] Sandra Herbert. Eloge: Peter J. Gautrey, 1925–2011. Isis, 103(4):753– 755, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/669051.

Cravens:2012:ERS

[847] Hamilton Cravens. Eloge: Robert E. Schofield, 1923–2011. Isis, 103 (4):756–759, December 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
669053.

Wilding:2012:LE

[848] Nick Wilding. Letter to the Editor. Isis, 103(4):760, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/669050.

Mooney:2012:BRP

[849] Graham Mooney. Book review: Performing Medicine: Medical Culture and Identity in Provincial England, c, by Michael Brown. 1760–1850. Isis, 103(4):761, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670070.

Canizares-Esguerra:2012:BRE

[850] Jorge Cañizares-Esguerra. Book review: Early American Cartographies. Isis, 103(4):762, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670071.

Radder:2012:BRS

[851] Hans Radder. Book review: Standards: Recipes for Reality. Isis, 103 (4):762–763, December 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670072.

Ramsingh:2012:BRM

[852] Brigit Ramsingh. Book review: Meat, Medicine, and Human Health in the Twentieth Century. Isis, 103(4):763–764, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670073.

Gal:2012:BRH

[853] Ofer Gal. Book review: How Modern Science Came into the World: Four Civilizations, One Seventeenth-Century Breakthrough. Isis, 103(4):764– 766, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670074.

Kimmelman:2012:BRM

[854] Barbara Kimmelman. Book review: Making Tobacco Bright: Creating an American Commodity, 1617–1937. Isis, 103(4):766–767, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670075.

Freiburger:2012:BRL

[855] Dana A. Freiburger. Book review: Learning by Doing: Experiments and Instruments in the History of Science Teaching. Isis, 103(4):767–769, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670076.

Hatfield:2012:BRD

[856] Gary Hatfield. Book review: The Determinate World: Kant and Helmholtz on the Physical Meaning of Geometry. Isis, 103(4):769–770, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670077.

Paradowski:2012:BRO

[857] Robert J. Paradowski. Book review: One Hundred Years at the Intersection of Chemistry and Physics: The Fritz Haber Institute of the Max Planck Society, 1911–2011. Isis, 103(4):770–771, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670078.

Hecht:2012:BRD

[858] Gabrielle Hecht. Book review: The Disaster Experts: Mastering Risk in Modern America. Isis, 103(4):771–772, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670079.

Navarro:2012:BRP

[859] Jaume Navarro. Book review: Los públicos de la ciencia: Expertos y profanos a través de la historia. Isis, 103(4):772–773, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670080.

Scharf:2012:BRT

[860] Sara Scharf. Book review: Trees of Life: A Visual History of Evolution. Isis, 103(4):773–774, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670081.

Blay:2012:BRS

[861] Michel Blay. Book review: Sources in the Development of Mathematics: Infinite Series and Products from the Fifteenth to the Twenty-first Century. Isis, 103(4):774–775, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/670082.

Cittadino:2012:BRH

[862] Eugene Cittadino. Book review: Hybrid Nature: Sewage Treatment and the Contradictions of the Industrial Ecosystem. Isis, 103(4):775–776, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670083.

Zammito:2012:BRS

[863] John H. Zammito. Book review: The Sciences of the Soul: The Early Modern Origins of Psychology. Isis, 103(4):776–777, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670084.

Long:2012:BRG

[864] Vicky Long. Book review: Gender, Health, and Popular Culture: Historical Perspectives. Isis, 103(4):777–778, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/670085.

Hayton:2012:BRM

[865] Darin Hayton. Book review: Monks, Manuscripts, and Sundials: The Navicula in Medieval England. Isis, 103(4):778–779, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670086.

Kitchell:2012:BRA

[866] Kenneth F. Kitchell, Jr. Book review: Aristoteles Latinus XVII 2, by Pieter de Leemans; Pieter de Leemans.II-III: De Progressu Animalium, De Motu Animalium Aristoteles Latinus XVII 1.III: De Motu Animalium. Isis, 103(4):779–780, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/670087.

Rommevaux:2012:BRG

[867] Sabine Rommevaux. Book review: Guillaume de Conches: Philosophie et science au XIIe siècle. Isis, 103(4):780, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670088.

Becchi:2012:BRB

[868] Antonio Becchi. Book review: Baumeister und Muhandis: Technologietransfer zwischen Orient und Okzident. Isis, 103(4):781, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/670089.

Grasshoff:2012:BRP

[869] Gerd Graßhoff. Book review: Ptolemy’s Geography in the Renaissance. Isis, 103(4):781–782, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670090.

Panti:2012:BRM

[870] Cecilia Panti. Book review: Il De mirabilibus mundi tra tradizione magica e filosofia naturale. Isis, 103(4):782–783, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670091.

Hogarth:2012:BRA

[871] Rana Asali Hogarth. Book review: The Anatomy of Blackness: Science and Slavery in an Age of Enlightenment. Isis, 103(4):783–784, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/670092.

Gaukroger:2012:BRE

[872] Stephen Gaukroger. Book review: Études sur Francis Bacon. Isis, 103 (4):784–785, December 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670093.

Steigerwald:2012:BRK

[873] Joan Steigerwald. Book review: Key Texts of Johann Wilhelm Ritter (1776–1810) on the Science and Art of Nature German Romanticism and Science: The Procreative Poetics of Goethe, Novalis, and Ritter. Isis, 103(4):785–786, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670094.

Eddy:2012:BRH

[874] Matthew Daniel Eddy. Book review: The Historiography of the Chemical Revolution: Patterns of Interpretation in the History of Science. Isis, 103 (4):786–787, December 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670095.

Pancaldi:2012:BRR

[875] Giuliano Pancaldi. Book review: La Rome des Français au temps des Lumières: Capitale de l’antique et carrefour de l’Europe, 1769–1791. Isis, 103(4):787–788, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670096.

Keitt:2012:BRW

[876] Andrew Keitt. Book review: Witchcraft and Inquisition in Early Modern Venice. Isis, 103(4):788–789, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/670097.

Qureshi:2012:BRM

[877] Sadiah Qureshi. Book review: Morbid Curiosities: Medical Museums in Nineteenth-Century Britain. Isis, 103(4):789–790, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670098.

Ablard:2012:BRA

[878] Jonathan Ablard. Book review: The Ailing City: Health, Tuberculosis, and Culture in Buenos Aires, 1870–1950. Isis, 103(4):790, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/670099.

Rossi:2012:BRP

[879] Arcangelo Rossi. Book review: Il professore e il suo demone: La lunga lotta di Max Planck contro la statistica (1896–1906). Isis, 103(4):790– 791, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670100.

Conroy:2012:BRM

[880] Mary Schaeffer Conroy. Book review: Medicine, Law, and the State in Imperial Russia. Isis, 103(4):791–792, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/670101.

Kuklick:2012:BRT

[881] Henrika Kuklick. Book review: Taming Cannibals: Race and the Victorians Peoples on Parade: Exhibitions, Empire, and Anthropology in Nineteenth-Century Britain. Isis, 103(4):793–794, December 2012. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670102.

Morris:2012:BRC

[882] Peter Morris. Book review: The Case of the Poisonous Socks: Tales from Chemistry. Isis, 103(4):794, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/670103.

Camilleri:2012:BRW

[883] Kristian Camilleri. Book review: Weimar Culture and Quantum Mechanics: Selected Papers by Paul Forman and Contemporary Perspectives on the Forman Thesis. Isis, 103(4):794–796, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670104.

McClellan:2012:BRS

[884] James E. McClellan III. Book review: Steam-Powered Knowledge: William Chambers and the Business of Publishing, 1820–1860. Isis, 103 (4):796–797, December 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670105.

Becker:2012:BRG

[885] Barbara J. Becker. Book review: The Great Melbourne Telescope. Isis, 103(4):797–798, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670106.

Josephson:2012:BRE

[886] Paul Josephson. Book review: Engineer of Revolutionary Russia: Iurii L, by Anthony Heywood. Lomonosov (1876–1952) and the Railways. Isis, 103(4):798–799, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670107.

Pruna-Goodgall:2012:BRP

[887] Pedro M. Pruna-Goodgall. Book review: Political Essay on the Island of Cuba: A Critical Edition. Isis, 103(4):799–800, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670108.

Reinarz:2012:BRW

[888] Jonathan Reinarz. Book review: War and the Militarization of British Army Medicine, 1793–1830. Isis, 103(4):800–801, December 2012. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670109.

Largent:2012:BRL

[889] Mark Largent. Book review: Lincoln and Darwin: Shared Visions of Race, Science, and Religion. Isis, 103(4):801–802, December 2012. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670110.

Monaldi:2012:BRS

[890] Daniela Monaldi. Book review: Sotto falso nome: Scienziate italiane ebree (1938–1945). Isis, 103(4):802–803, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/670111.

Smith:2012:BRF

[891] Robert W. Smith. Book review: Fifty Years on the Space Frontier: Halo Orbits, Comets, Asteroids, and More. Isis, 103(4):803–804, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/670112.

Nye:2012:BRA

[892] David E. Nye. Book review: Atomic Frontier Days: Hanford and the American West. Isis, 103(4):804–805, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/670113.

Farish:2012:BRE

[893] Matthew Farish. Book review: Entangled Geographies: Empire and Technopolitics in the Global Cold War. Isis, 103(4):805–806, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670114.

Strasser:2012:BRG

[894] Bruno J. Strasser. Book review: Genentech: The Beginnings of Biotech. Isis, 103(4):806–807, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670115.

Widmalm:2012:BRS

[895] Sven Widmalm. Book review: Science for Welfare and Warfare: Technology and State Initiative in Cold War Sweden. Isis, 103(4):807–808, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670116.

More:2012:BRW

[896] Ellen S. More. Book review: Women Scientists in America, by Margaret W. Rossiter. Volume 3: Forging a New World since 1972. Isis, 103 (4):808–810, December 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670117.

Reverby:2012:BRB

[897] Susan M. Reverby. Book review: Behind Closed Doors: IRBs and the Making of Ethical Research. Isis, 103(4):810–811, December 2012. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670118.

Rasmussen:2012:BRP

[898] Nicolas Rasmussen. Book review: Pills, Power, and Policy: The Struggle for Drug Reform in Cold War America and Its Consequences. Isis, 103(4):811, December 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670119.

Bimm:2012:BRP

[899] Jordan Bimm. Book review: Psychology of Space Exploration: Contemporary Research in Historical Perspective. Isis, 103(4):812, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/670120.

Masco:2012:BRR

[900] Joseph Masco. Book review: The Rise of Nuclear Fear. Isis, 103(4):813– 814, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670121.

Renwick:2012:BRS

[901] Chris Renwick. Book review: Science. Isis, 103(4):814–815, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/670122.

Moore:2012:BRS

[902] Alfred Moore. Book review: Science in a Democratic Society. Isis, 103 (4):815–816, December 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670123.

Anonymous:2012:NCd

[903] Anonymous. Notes on contributors. Isis, 103(4):817–819, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/669055.

Anonymous:2012:IIV

[904] Anonymous. Index to Isis, volume 103, 2012. Isis, 103(4):821–832, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/670012.

Anonymous:2012:CIV

[905] Anonymous. Contents of ISIS, volume 103. Isis, 103(4):835–836, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/670014.

Anonymous:2012:Fd

[906] Anonymous. Frontispiece. Isis, 103(4):vi, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670069.

Farish:2013:LLO

[907] Matthew Farish. The lab and the land: Overcoming the Arctic in Cold War Alaska. Isis, 104(1):1–29, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/669881.

Opitz:2013:TBB

[908] Donald L. Opitz. “A triumph of brains over brute”: Women and science at the Horticultural College, Swanley, 1890–1910. Isis, 104(1):30–62, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/669882.

Sattar:2013:ALI

[909] Atia Sattar. The aesthetics of laboratory inscription: Claude Bernard’s cahier rouge. Isis, 104(1):63–85, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/669883.

Lightman:2013:I

[910] Bernard Lightman. Introduction. Isis, 104(1):86–87, March 2013. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/669893.

Alder:2013:HSO

[911] Ken Alder. The history of science as oxymoron: From scientific exceptionalism to episcience. Isis, 104(1):88–101, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/669889.

Nappi:2013:GBA

[912] Carla Nappi. The global and beyond: Adventures in the local historiographies of science. Isis, 104(1):102–110, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/669894.

Lehoux:2013:ASD

[913] Daryn Lehoux. Ancient science in a digital age. Isis, 104(1):111–118, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/669892.

Laubichler:2013:CPH

[914] Manfred D. Laubichler, Jane Maienschein, and Jürgen Renn. Computational perspectives in the history of science: To the memory of Peter Damerow. Isis, 104(1):119–130, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/669891.

Nyhart:2013:SHS

[915] Lynn K. Nyhart. The shape of the history of science profession, 2038: A prospective retrospective. Isis, 104(1):131–139, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/669895.

Lowy:2013:BRG

[916] Ilana Löwy. Book review: Golden Holocaust: Origins of the Cigarette Catastrophe and the Case for Abolition. Isis, 104(1):140–144, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/669888.

Messbarger:2013:BRA

[917] Rebecca Messbarger. Book review: The Art of Medicine: Over Two Thousand Years of Images and Imagination. Isis, 104(1):145, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/670864.

Mazumdar:2013:BRO

[918] Pauline M. H. Mazumdar. Book review: The Oxford Handbook of the History of Eugenics. Isis, 104(1):146–147, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670865.

Princehouse:2013:BRE

[919] Patricia Princehouse. Book review: Evolutionary Restraints: The Contentious History of Group Selection. Isis, 104(1):147–148, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670866.

Buckley:2013:BRM

[920] Eve Buckley. Book review: Mapping Latin America: A Cartographic Reader. Isis, 104(1):148–149, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/670867.

Sommer:2013:BRG

[921] Marianne Sommer. Book review: The Genealogical Science: The Search for Jewish Origins and the Politics of Epistemology. Isis, 104(1):149– 150, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670868.

Quintero:2013:BRC

[922] Camilo Quintero. Book review: Cultivating the Colonies: Colonial States and Their Environmental Legacies. Isis, 104(1):150–151, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670869.

Chamayou:2013:BRP

[923] Gregoire Chamayou. Book review: Paper Machines: About Cards and Catalogs, 1548–1929. Isis, 104(1):151–152, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670870.

Knight:2013:BRA

[924] David Knight. Book review: The Ancient Mythology of Modern Science: A Mythologist Looks (Seriously) at Popular Science Writing. Isis, 104 (1):152–153, March 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670871.

Totelin:2013:BRM

[925] Laurence Totelin. Book review: Man and Wound in the Ancient World: A History of Military Medicine from Sumer to the Fall of Constantinople. Isis, 104(1):153–154, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670872.

Lehoux:2013:BRH

[926] Daryn Lehoux. Book review: A History of Trust in Ancient Greece. Isis, 104(1):154–155, March 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670873.

Verelst:2013:BRW

[927] Karin Verelst. Book review: What Did the Romans Know? An Inquiry into Science and Worldmaking. Isis, 104(1):155–156, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670874.

Leng:2013:BRJ

[928] Rainer Leng. Book review: Johannes Fontana; Horst Kranz, eds. Opera iuvenalia de rotis horologiis et mensuris/Jugendwerke über Räder, Uhren und Messungen. Isis, 104(1):156–157, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/670875.

Popper:2013:BRA

[929] Nicholas Popper. Book review: The Arch-Conjuror of England: John Dee. Isis, 104(1):157–158, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670876.

Reiss:2013:BRF

[930] Timothy Reiss. Book review: Fictions of the Cosmos: Science and Literature in the Seventeenth Century. Isis, 104(1):158–159, March 2013.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670877.

Mazzotti:2013:BRU

[931] Massimo Mazzotti. Book review: Unità del sapere molteplicità dei saperi Francesco Bianchini (1662–1729) tra natura, storia e religione. Isis, 104(1):159–160, March 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670878.

Hunter:2013:BRR

[932] Michael Hunter. Book review: Regimens of the Mind: Boyle, Locke, and the Early Modern Cultura Animi Tradition. Isis, 104(1):160–161, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/670879.

Pawley:2013:BRN

[933] Emily Pawley. Book review: The Nation’s Nature: How Continental Presumptions Gave Rise to the United States of America. Isis, 104(1):161– 162, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670880.

Palmerino:2013:BRH

[934] Carla Rita Palmerino. Book review: Honoré Fabri and the Concept of Impetus: A Bridge between Conceptual Frameworks. Isis, 104(1):162– 163, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670881.

Olshin:2013:BRE

[935] Benjamin B. Olshin. Book review: Exploring the Kingdom of Saturn: Kircher’s Latium and Its Legacy. Isis, 104(1):163–164, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670882.

Gronim:2013:BRK

[936] Sara Stidstone Gronim. Book review: Knowing Nature: Art and Science in Philadelphia, 1740–1840. Isis, 104(1):164–165, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670883.

Sacks:2013:BRC

[937] Benjamin J. Sacks. Book review: Centres and Cycles of Accumulation in and around the Netherlands during the Early Modern Period. Isis, 104(1):165–166, March 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670884.

Coffman:2013:BRC

[938] D’Maris Coffman. Book review: Casualties of Credit: The English Financial Revolution, 1620–1720. Isis, 104(1):166–167, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670885.

Sheppard:2013:BRA

[939] Kathleen L. Sheppard. Book review: American Egyptologist: The Life of James Henry Breasted and the Creation of His Oriental Institute. Isis, 104(1):167–168, March 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670886.

Gaudilliere:2013:BRB

[940] Jean-Paul Gaudillière. Book review: Bakterien in Krieg und Frieden: Eine Geschichte der medizinischen Bakteriologie in Deutschland, 1890– 1933. Isis, 104(1):168–170, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/670887.

Alder:2013:BRT

[941] Ken Alder. Book review: The Truth Machine: A Social History of the Lie Detector. Isis, 104(1):170–171, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/670888.

Chazan:2013:BRM

[942] Michael Chazan. Book review: La méthode de Zadig: La trace, le fossile, la preuve L’abbé Breuil: Un préhistorien dans le siècle. Isis, 104(1):171– 172, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670889.

Nordmann:2013:BRH

[943] Alfred Nordmann. Book review: Heinrich Hertz. Isis, 104(1):172–173, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/670890.

Gispert:2013:BRE

[944] Hélène Gispert. Book review: Évariste Galois: La fabrication d’une icône mathématique. Isis, 104(1):173–174, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/670891.

James:2013:BRN

[945] Jeremiah James. Book review: Neither Physics nor Chemistry: A History of Quantum Chemistry. Isis, 104(1):174–175, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670892.

Jacobsen:2013:BRK

[946] Anja Skaar Jacobsen. Book review: Karen Jelved; Andrew D. Jackson. The Travel Letters of H. C. Ørsted. Isis, 104(1):175–176, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670893.

Beyerchen:2013:BRJ

[947] Alan Beyerchen. Book review: Jost Lemmerich; Christian Matthei, eds. Mein lieber Sohn! Die Briefe von Max von Laue an seinen Sohn Theodor in den Vereinigten Staaten von Amerika 1937 bis 1946. Isis, 104(1):176– 177, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670894.

Hokkanen:2013:BRD

[948] Markku Hokkanen. Book review: Deborah J. Neill. Networks in Tropical Medicine: Internationalism, Colonialism, and the Rise of a Medical Specialty, 1890–1930. Isis, 104(1):177–178, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670895.

Snow:2013:BRS

[949] Stephanie Snow. Book review: Social Poison: The Culture and Politics of Opiate Control in Britain and France, 1821–1926. Isis, 104(1):178– 179, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670896.

Bigg:2013:BRR

[950] Charlotte Bigg. Book review: Radioactive: Marie and Pierre Curie: A Tale of Love and Fallout. Isis, 104(1):179–180, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670897.

Rees:2013:BRM

[951] Amanda Rees. Book review: Missing Links: The African and American Worlds of R, by Jeremy Rich. L. Garner, Primate Collector. Isis, 104(1): 180, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670898.

Walker:2013:BRP

[952] Mark Walker. Book review: The Pontecorvo Affair: A Cold War Defection and Nuclear Physics. Isis, 104(1):180–181, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670899.

Stamhuis:2013:BRS

[953] Ida Stamhuis. Book review: De stem van de wetenschap: Geschiedenis van de Koninklijke Nederlandse Akademie van Wetenschappen, by Klaas van Berkel. Volume 2: 1914–2008. Isis, 104(1):181–182, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670900.

Caianiello:2013:BRK

[954] Silvia Caianiello. Book review: Konrad Lorenz e l’etologia contemporanea: L’eredità problematica di uno scienziato inattuale. Isis, 104 (1):182–183, March 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670901.

Hamblin:2013:BRB

[955] Jacob Darwin Hamblin. Book review: Being Nuclear: Africans and the Global Uranium Trade. Isis, 104(1):183–184, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670902.

Quintero:2013:BRI

[956] Camilo Quintero. Book review: Instituting Nature: Authority, Expertise, and Power in Mexican Forests. Isis, 104(1):184–185, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670903.

Richert:2013:BRA

[957] Lucas Richert. Book review: The AIDS Conspiracy: Science Fights Back. Isis, 104(1):185–186, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/670904.

Stevens:2013:BRB

[958] Hallam Stevens. Book review: Biomedical Computing: Digitizing Life in the United States. Isis, 104(1):186–187, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/670905.

Tybjerg:2013:BRT

[959] Karin Tybjerg and Adam Bencard. Book review: Tissue Culture in Science and Society: The Public Life of a Biological Technique in Twentieth Century Britain. Isis, 104(1):187–188, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/670906.

Roland:2013:BRE

[960] Alex Roland. Book review: The Enigma of the Aerofoil: Rival Theories in Aerodynamics, 1909–1930. Isis, 104(1):188–189, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670907.

Belkind:2013:BRI

[961] Ori Belkind. Book review: Isaac Newton’s Scientific Method: Turning Data into Evidence about Gravity and Cosmology. Isis, 104(1):189–190, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/670908.

Pettit:2013:BRW

[962] Michael Pettit. Book review: Working Knowledge: Making the Human Sciences from Parsons to Kuhn. Isis, 104(1):190–191, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670909.

Anonymous:2013:NCa

[963] Anonymous. Notes on contributors. Isis, 104(1):192–195, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/669897.

Anonymous:2013:Fa

[964] Anonymous. Frontispiece. Isis, 104(1):vi, March 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/670921.

Bertomeu-Sanchez:2013:MUA

[965] José Ramón Bertomeu-Sánchez. Managing uncertainty in the Academy and the courtroom: Normal arsenic and Nineteenth-Century toxicology. Isis, 104(2):197–225, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670945.

Bertucci:2013:IVW

[966] Paola Bertucci. The in/visible woman: Mariangela Ardinghelli and the circulation of knowledge between Paris and Naples in the Eighteenth Century. Isis, 104(2):226–249, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/670946.

Gianquitto:2013:BSH

[967] Tina Gianquitto. Botanical smuts and hermaphrodites: Lydia Becker, Darwin’s botany, and education reform. Isis, 104(2):250–277, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670947.

Lambert:2013:NHM

[968] Kevin Lambert. A natural history of mathematics: George Peacock and the making of English algebra. Isis, 104(2):278–302, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670948.

Schmalzer:2013:ICS

[969] Sigrid Schmalzer. Insect control in socialist China and the corporate United States: The act of comparison, the tendency to forget, and the construction of difference in 1970s U.S.–Chinese scientific exchange. Isis, 104(2):303–329, June 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
670949.

Phalkey:2013:I

[970] Jahnavi Phalkey. Introduction. Isis, 104(2):330–336, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/670950.

Raj:2013:BPL

[971] Kapil Raj. Beyond postcolonialism … and postpositivism: Circulation and the global history of science. Isis, 104(2):337–347, June 2013.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670951.

Ganeri:2013:WOS

[972] Jonardon Ganeri. Well-ordered science and Indian epistemic cultures: Toward a polycentered history of science. Isis, 104(2):348–359, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/670953.

Arnold:2013:NSP

[973] David Arnold. Nehruvian science and postcolonial India. Isis, 104(2): 360–370, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/670954.

Chowdhury:2013:HAS

[974] Indira Chowdhury. A historian among scientists: Reflections on archiving the history of science in postcolonial India. Isis, 104(2):371–380, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/670955.

Grattan-Guinness:2013:LE

[975] I. Grattan-Guinness. Letter to the Editor. Isis, 104(2):381, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670956.

Tjossem:2013:BRS

[976] Sara Tjossem. Book review: The Sounding of the Whale: Science and Cetaceans in the Twentieth Century. Isis, 104(2):382–383, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/672167.

Shapiro:2013:BRH

[977] Alan E. Shapiro. Book review: A History of Optics: From Greek Antiquity to the Nineteenth Century. Isis, 104(2):383–384, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/672168.

Petersen:2013:BRD

[978] Katrina Petersen. Book review: Disease Maps: Epidemics on the Ground. Isis, 104(2):384–385, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/672169.

Farber:2013:BRG

[979] Paul Lawrence Farber. Book review: Gorgeous Beasts: Animal Bodies in Historical Perspective. Isis, 104(2):385, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/672170.

Bowen:2013:BRC

[980] Alan C. Bowen. Book review: Chasing Shadows: Mathematics, Astronomy, and the Early History of Eclipse Reckoning. Isis, 104(2):385–387, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/672171.

Warner:2013:BRE

[981] Deborah Warner. Book review: From Earth-Bound to Satellite: Telescopes, Skills, and Networks. Isis, 104(2):387, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/672172.

Byden:2013:BRS

[982] Börje Bydén. Book review: Science and Eastern Orthodoxy: From the Greek Fathers to the Age of Globalization. Isis, 104(2):387–388, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/672173.

Stewart:2013:BRP

[983] Larry Stewart. Book review: Progressive Enlightenment: The Origins of the Gaslight Industry, 1780–1820. Isis, 104(2):388–389, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/672174.

Obrist:2013:BRI

[984] Barbara Obrist. Book review: Images et magie: Picatrix entre Orient et Occident. Isis, 104(2):389–390, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/672175.

Smith:2013:BRI

[985] A. Mark Smith. Book review: Imagination, Meditation, and Cognition in the Middle Ages. Isis, 104(2):390–391, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/672176.

Denery:2013:BRM

[986] Dallas G. Denery II. Book review: The Moral Treatise on the Eye. Isis, 104(2):391–392, June 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
672177.

Rutkin:2013:BRR

[987] H. Darrel Rutkin. Book review: Renaissance Meteorology: Pomponazzi to Descartes. Isis, 104(2):392–393, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/672178.

Ziegler:2013:BRL

[988] Joseph Ziegler. Book review: L’uroscopie au Moyen Âge: “Lire dans un verre la nature de l’homme”, by Laurence Moulinier-Brogi. Isis, 104(2): 393–394, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/672179.

Machamer:2013:BRC

[989] Peter Machamer and Benjamin Goldberg. Book review: Controversies within the Scientific Revolution. Isis, 104(2):394–395, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/672180.

Carhart:2013:BRG

[990] Michael C. Carhart. Book review: Geschichte als präzisierte Wissenschaft: Johann Christoph Gatterer und die Historiographie des 18, by Martin Gierl. Jahrhunderts im ganzen Umfang. Isis, 104(2):395–396, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/672181.

Outram:2013:BRC

[991] Dorinda Outram. Book review: The Colonial Machine: French Science and Overseas Expansion in the Old Regime. Isis, 104(2):396–398, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/672182.

Barreca:2013:BRE

[992] Francesco Barreca. Book review: Édouard Mehl; Nicolas Roudet, eds. Kepler: La physique céleste: Autour de l’Astronomia Nova (1609). Isis, 104(2):398–399, June 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
672183.

Palmieri:2013:BRE

[993] Paolo Palmieri. Book review: The Equilibrium Controversy: Guidobaldo del Monte’s Critical Notes on the Mechanics of Jordanus and Benedetti and Their Historical and Conceptual Background. Isis, 104(2):399–400, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/672184.

Gordin:2013:BRM

[994] Michael D. Gordin. Book review: Mapping Europe’s Borderlands: Russian Cartography in the Age of Empire. Isis, 104(2):400–401, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/672185.

Cooke:2013:BRA

[995] Kathy J. Cooke. Book review: Art and Science in Breeding: Creating Better Chickens. Isis, 104(2):401–402, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/672186.

Weindling:2013:BRA

[996] Paul Weindling. Book review: Anthropology at War: World War I and the Science of Race in Germany. Isis, 104(2):402–403, June 2013. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/672187.

Reid:2013:BRA

[997] Chris Reid. Book review: All the Fish in the Sea: Maximum Sustainable Yield and the Failure of Fisheries Management. Isis, 104(2):403–404, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/672188.

Israel:2013:BRE

[998] Giorgio Israel. Book review: Einstein’s Jewish Science: Physics at the Intersection of Politics and Religion. Isis, 104(2):404–405, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/672189.

Turchetti:2013:BRL

[999] Simone Turchetti. Book review: Léon Rosenfeld: Physics, Philosophy, and Politics in the Twentieth Century. Isis, 104(2):405–406, June 2013.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/672190.

Morus:2013:BRC

[1000] Iwan Rhys Morus. Book review: The Correspondence of Michael Faraday. Isis, 104(2):406–407, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/672191.

Brown:2013:BRP

[1001] Julie K. Brown. Book review: Popular Exhibitions, Science, and Showmanship, 1840–1910. Isis, 104(2):407–408, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/672192.

Holmberg:2013:BRH

[1002] Gustav Holmberg. Book review: History of Astronomy in Finland, 1828– 1918. Isis, 104(2):408–409, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/672193.

Davis:2013:BRF

[1003] Frederick R. Davis. Book review: The Feathery Tribe: Robert Ridgway and the Modern Study of Birds. Isis, 104(2):409–410, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/672194.

Davis:2013:BRD

[1004] Georgiann Davis. Book review: Doubting Sex: Inscriptions, Bodies, and Selves in Nineteenth-Century Hermaphrodite Case Histories. Isis, 104(2): 410–411, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/672195.

Crull:2013:BRH

[1005] Elise Crull. Book review: Hertha Sponer: A Woman’s Life as a Physicist in the Twentieth Century: “So You Won’t Forget Me”, by Marie-Ann Maushart; Annette Vogt; Ralph A. Morris; Brenda P. Winnewisser. Isis, 104(2):411–412, June 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
672196.

Bowler:2013:BRC

[1006] Peter J. Bowler. Book review: A Cultural History of Heredity. Isis, 104 (2):412, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/672197.

Turner:2013:BRA

[1007] Steven Turner. Book review: Acolytes of Nature: Defining Natural Science in Germany, 1770–1850. Isis, 104(2):413, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/672198.

vonSchwerin:2013:BRC

[1008] Alexander von Schwerin. Book review: Creating a Physical Biology: The Three-Man Paper and Early Molecular Biology. Isis, 104(2):414– 415, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/672199.

Hirsh:2013:BRO

[1009] Richard Hirsh. Book review: Overpotential: Fuel Cells, Futurism, and the Making of a Power Panacea. Isis, 104(2):415–416, June 2013. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/672200.

Greene:2013:BRP

[1010] Mott Greene. Book review: The Pseudoscience Wars: Immanuel Velikovsky and the Birth of the Modern Fringe. Isis, 104(2):416–417, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/672201.

Buerki:2013:BRP

[1011] Robert A. Buerki. Book review: Prescribed: Writing, Filling, Using, and Abusing the Prescription in Modern America. Isis, 104(2):417–418, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/672202.

Eghigian:2013:BRC

[1012] Greg Eghigian. Book review: Cold War Social Science: Knowledge Production, Liberal Democracy, and Human Nature. Isis, 104(2):418–419, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/672203.

Johnson:2013:BRB

[1013] Stephen Johnson. Book review: Blue Sky Metropolis: The Aerospace Century in Southern California. Isis, 104(2):419–420, June 2013. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/672204.

Dorries:2013:BRM

[1014] Matthias Dörries. Book review: Michael Polanyi and His Generation: Origins of the Social Construction of Science. Isis, 104(2):420–421, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/672205.

Smyrek:2013:BRO

[1015] Volker Smyrek. Book review: The Oxford Handbook of Sound Studies. Isis, 104(2):421–422, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/672206.

Lynch:2013:BRE

[1016] John M. Lynch. Book review: Evolutionary History: Uniting History and Biology to Understand Life on Earth. Isis, 104(2):422–423, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/672207.

Gattei:2013:BRK

[1017] Stefano Gattei. Book review: K. Brad Wray, Kuhn’s Evolutionary Social Epistemology. Isis, 104(2):423–424, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/672208.

Anonymous:2013:NCb

[1018] Anonymous. Notes on contributors. Isis, 104(2):425–428, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/670958.

Anonymous:2013:Fb

[1019] Anonymous. Frontispiece. Isis, 104(2):vi, June 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/672213.

Crowther:2013:TIE

[1020] Kathleen M. Crowther and Peter Barker. Training the intelligent eye: Understanding illustrations in early modern astronomy texts. Isis, 104 (3):429–470, September 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/673269.pdf.

Endersby:2013:MUE

[1021] Jim Endersby. Mutant utopias: Evening primroses and imagined futures in early Twentieth-Century America. Isis, 104(3):471–503, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/673270.pdf.

vanLunteren:2013:PED

[1022] Frans H. van Lunteren and Marijn J. Hollestelle. Paul Ehrenfest and the dilemmas of modernity. Isis, 104(3):504–536, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/673271.pdf.

Weldon:2013:I

[1023] Stephen P. Weldon. Introduction. Isis, 104(3):537–539, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/673272.pdf.

Weldon:2013:BSO

[1024] Stephen P. Weldon. Bibliography is social: Organizing knowledge in the Isis bibliography from Sarton to the early Twenty-First Century. Isis, 104 (3):540–550, September 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/673273.pdf.

Alfonso-Goldfarb:2013:SCO

[1025] Ana M. Alfonso-Goldfarb, Silvia Waisse, and Márcia H. M. Ferraz. From shelves to cyberspace: Organization of knowledge and the complex identity of history of science. Isis, 104(3):551–560, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/673274.pdf.

Anderson:2013:ODS

[1026] R. Joseph Anderson. The organization and description of science archives in America. Isis, 104(3):561–572, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/pdfplus/10.1086/673275.pdf.

Allen:2013:CCC

[1027] Colin Allen and the InPhO Group. Cross-cutting categorization schemes in the digital humanities. Isis, 104(3):573–583, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/673276.pdf.

Anonymous:2013:HSS

[1028] Anonymous. History of Science Society Annual Meeting, 2012. Isis, 104 (3):584–586, September 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/674458.pdf.

Anonymous:2013:HP

[1029] Anonymous. HSS Prizes for 2012. Isis, 104(3):586–591, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/674459.pdf.

Malley:2013:LE

[1030] Marjorie C. Malley. Letters to the Editor. Isis, 104(3):592, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/673278.pdf.

Anderson:2013:BRR

[1031] Katharine Anderson. Book review: Rima D. Apple; Gregory J. Downey; Stephen L. Vaughan, eds., Science in Print: Essays in the History of Science and the Culture of Print. Isis, 104(3):593–594, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/674460.pdf.

Finlay:2013:BRE

[1032] Mark Finlay. Book review: Edward B. Barbier, Scarcity and Frontiers: How Economies Have Developed through Natural Resource Exploitation. Isis, 104(3):594–595, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/674461.pdf.

Jasanoff:2013:BRM

[1033] Sheila Jasanoff. Book review: Mario Biagioli; Peter Jaszi; Martha Woodmansee, eds., Making and Unmaking Intellectual Property: Creative Production in Legal and Cultural Perspective. Isis, 104(3):595–596, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
674462.pdf.

Nash:2013:BRM

[1034] Linda Nash. Book review: Mark Fiege, The Republic of Nature: An Environmental History of the United States. Isis, 104(3):596–597, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/674463.pdf.

Kimmelman:2013:BRN

[1035] Barbara Kimmelman. Book review: Noel Kingsbury, Hybrid: The History and Science of Plant Breeding. Isis, 104(3):597–598, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/674464.pdf.

Barrow:2013:BRK

[1036] Mark V. Barrow, Jr. Book review: Kristian H. Nielsen; Michael Harbsmeier; Christopher J. Ries, Scientists and Scholars in the Field: Studies in the History of Fieldwork and Expeditions. Isis, 104(3):598–599, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
674465.pdf.

Hanson:2013:BRR

[1037] Elizabeth Hanson. Book review: Rachel Poliquin, The Breathless Zoo: Taxidermy and the Cultures of Longing. Isis, 104(3):599–600, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/674466.pdf.

Carruthers:2013:BRG

[1038] Jane Carruthers. Book review: Gloria Pungetti; Gonzalo Oviedo; Della Hooke, eds., Sacred Species and Sites: Advances in Biocultural Conservation. Isis, 104(3):600–601, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/674467.pdf.

Hunt:2013:BRW

[1039] Bruce J. Hunt. Book review: William Rosen, The Most Powerful Idea in the World: A Story of Steam, Industry, and Invention. Isis, 104(3):601, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
674468.pdf.

Sherman:2013:BRK

[1040] Roger Sherman. Book review: Klaus Staubermann, ed., Reconstructions: Recreating Science and Technology of the Past. Isis, 104(3):601–602, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
674469.pdf.

Williams:2013:BRM

[1041] Gareth Williams. Book review: Marco Beretta; Francesco Citti; Lucia Pasetti, eds., Seneca e le scienze naturali. Isis, 104(3):603–604, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
674470.pdf.

Kessel:2013:BRG

[1042] Grigory Kessel. Book review: Glen M. Cooper, Galen, De diebus decretoriis, from Greek into Arabic: A Critical Edition, with Translation and Commentary, of H .unayn ibn Ish.aq, Kitab ayyam al-buh.ran. Isis, 104(3):604, September 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/674471.pdf.

Pesic:2013:BRD

[1043] Peter Pesic. Book review: Daniel Heller-Roazen, The Fifth Hammer: Pythagoras and the Disharmony of the World. Isis, 104(3):604–605, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
67327913.pdf.

Netz:2013:BRL

[1044] Reviel Netz. Book review: Leonid Zhmud, Pythagoras and the Early Pythagoreans. Isis, 104(3):606–607, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/pdfplus/10.1086/674473.pdf.

Shackelford:2013:BRH

[1045] Jole Shackelford. Book review: Hiro Hirai, Medical Humanism and Natural Philosophy: Renaissance Debates on Matter, Life, and the Soul. Isis, 104(3):607–608, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/674474.pdf.

Neri:2013:BRS

[1046] Janice Neri. Book review: Sachiko Kusukawa, Picturing the Book of Nature: Image, Text, and Argument in Sixteenth-Century Human Anatomy and Medical Botany. Isis, 104(3):608–609, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/674475.pdf.

Becchi:2013:BRA

[1047] Antonio Becchi. Book review: Alina Payne, The Telescope and the Compass: Teofilo Gallaccini and the Dialogue between Architecture and Science in the Age of Galileo Teofilo Gallaccini: Selected Writings and Library. Isis, 104(3):609–610, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/674476.pdf.

Clucas:2013:BRM

[1048] Stephen Clucas. Book review: Matthias Schemmel, The English Galileo: Thomas Harriot’s Work on Motion as an Example of Preclassical Mechanics. Isis, 104(3):610–612, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/674477.pdf.

Hunter:2013:BRP

[1049] Michael Hunter. Book review: Philip Beeley; Christoph J. Scriba, eds., The Correspondence of John Wallis (1616–1703). Isis, 104(3):612–613, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
674478.pdf.

Kusukawa:2013:BRD

[1050] Sachiko Kusukawa. Book review: Daniela Bleichmar, Visible Empire: Botanical Expeditions and Visual Culture in the Hispanic Enlightenment. Isis, 104(3):613–614, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/674479.pdf.

Harper:2013:BRS

[1051] William L. Harper. Book review: Steffen Ducheyne, The Main Business of Natural Philosophy: Isaac Newton’s Natural-Philosophical Methodology. Isis, 104(3):614–615, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/674480.pdf.

Alexander:2013:BRR

[1052] Amir Alexander. Book review: Robert Fox, ed., Thomas Harriot and His World: Mathematics, Exploration, and Natural Philosophy in Early Modern England. Isis, 104(3):615–616, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/pdfplus/10.1086/674481.pdf.

King:2013:BRY

[1053] Helen King. Book review: Yasmin Haskell, ed., Diseases of the Imagination and Imaginary Disease in the Early Modern Period. Isis, 104 (3):616–617, September 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/674482.pdf.

Ducheyne:2013:BRA

[1054] Steffen Ducheyne. Book review: Andrew Janiak; Eric Schliesser, eds., Interpreting Newton: Critical Essays. Isis, 104(3):617–618, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/674483.pdf.

Hankins:2013:BRR

[1055] Thomas L. Hankins. Book review: Russell McCormmach, Weighing the World: The Reverend John Michell of Thornhill. Isis, 104(3):618–619, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
674484.pdf.

Knoeff:2013:BRJ

[1056] Rina Knoeff. Book review: John C. Powers, Inventing Chemistry: Herman Boerhaave and the Reform of the Chemical Arts. Isis, 104(3):619– 620, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/pdfplus/10.
1086/674485.pdf.

Miller:2013:BRG

[1057] David Philip Miller. Book review: George Rousseau, The Notorious Sir John Hill: The Man Destroyed by Ambition in the Era of Celebrity. Isis, 104(3):620–621, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/674486.pdf.

Konoval:2013:BRM

[1058] Brandon Konoval. Book review: Maria Semi, Music as a Science of Mankind in Eighteenth-Century Britain. Isis, 104(3):621–622, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/674487.pdf.

Ruse:2013:BRF

[1059] Michael Ruse. Book review: Frederick Burkhardt; James A. Secord; Janet Browne; Samantha Evans; Shelley Innes; Alison M. Pearn; Paul White, eds., The Correspondence of Charles Darwin. Isis, 104(3):622– 624, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/pdfplus/10.
1086/674488.pdf.

Fleischer:2013:BRD

[1060] Alette Fleischer. Book review: Denis Diagre-Vanderpelen, The Botanic Garden of Brussels (1826–1912): Reflection of a Changing Nation. Isis, 104(3):624–625, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/674489.pdf.

Mazzarello:2013:BRL

[1061] Paolo Mazzarello. Book review: Liborio Dibattista. Il movimento immobile: La fisiologia di E.-J. Marey e C. E. François-Franck (1868–1921). Isis, 104(3):625–626, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/674490.pdf.

Langins:2013:BRR

[1062] Janis Langins. Book review: Robert Fox, The Savant and the State: Science and Cultural Politics in Nineteenth-Century France. Isis, 104 (3):626–627, September 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/674491.pdf.

Maas:2013:BRJ

[1063] Ad Maas. Book review: József Illy, The Practical Einstein: Experiments, Patents, Inventions. Isis, 104(3):627–628, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/674492.pdf.

Reynolds:2013:BRP

[1064] Terry S. Reynolds. Book review: The Papers of Thomas A. Edison, by Paul B. Israel; Louis Carlat; Theresa M. Collins; David Hochfelder, eds. Isis, 104(3):628–629, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/674493.pdf.

Lopez-Denis:2013:BRN

[1065] Adrián López-Denis. Book review: Niklas Thode Jensen, For the Health of the Enslaved: Slaves, Medicine, and Power in the Danish West Indies, 1803–1848. Isis, 104(3):629–630, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/674494.pdf.

Rieppel:2013:BRS

[1066] Lukas Rieppel. Book review: Simon J. Knell, The Great Fossil Enigma: The Search for the Conodont Animal. Isis, 104(3):630–631, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/674495.pdf.

Burton:2013:BRN

[1067] Christopher Burton. Book review: Nikolai Krementsov, A Martian Stranded on Earth: Alexander Bogdanov, Blood Transfusions, and Proletarian Science. Isis, 104(3):631–632, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/pdfplus/10.1086/674496.pdf.

Ferreiro:2013:BRD

[1068] Larrie D. Ferreiro. Book review: Don Leggett; Richard Dunn, eds., Re-inventing the Ship: Science, Technology, and the Maritime World, 1800–1918. Isis, 104(3):632–633, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/674497.pdf.

Bergwik:2013:BRA

[1069] Staffan Bergwik. Book review: Annette Lykknes; Donald L. Opitz; Brigitte Van Tiggelen, eds., For Better or for Worse? Collaborative Couples in the Sciences. Isis, 104(3):633–634, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/674498.pdf.

Conis:2013:BRG

[1070] Elena Conis. Book review: Guenter B. Risse, Plague, Fear, and Politics in San Francisco’s Chinatown. Isis, 104(3):634–635, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/674499.pdf.

Kragh:2013:BRS

[1071] Helge Kragh. Book review: Silvan S. Schweber, Nuclear Forces: The Making of the Physicist Hans Bethe. Isis, 104(3):635–636, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/674500.pdf.

Fiorentini:2013:BRH

[1072] Erna Fiorentini. Book review: Hanna Rose Shell, Hide and Seek: Camouflage, Photography, and the Media of Reconnaissance. Isis, 104(3): 636–637, September 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/674501.pdf.

Finn:2013:BRJ

[1073] Michael Finn. Book review: Jonathan Strauss, Human Remains: Medicine, Death, and Desire in Nineteenth-Century Paris. Isis, 104(3): 637–638, September 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/674502.pdf.

Hardy:2013:BRW

[1074] Anne Hardy. Book review: William C. Summers, The Great Manchurian Plague of 1910–1911: The Geopolitics of an Epidemic Disease. Isis, 104 (3):638–639, September 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/674503.pdf.

Wuthrich:2013:BRA

[1075] Adrian Wüthrich. Book review: Andrea Cappelli; Elena Castellani; Filippo Colomo; Paolo Di Vecchia, eds., The Birth of String Theory. Isis, 104(3):639–640, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/674504.pdf.

Olley:2013:BRP

[1076] Allan Olley. Book review: Paul E. Ceruzzi, Computing: A Concise History. Isis, 104(3):640–641, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/674505.pdf.

Gossin:2013:BRJ

[1077] Pamela Gossin. Book review: John Cheng, Astounding Wonder: Imagining Science and Science Fiction in Interwar America. Alexander C. T. Geppert, ed., Imagining Outer Space: European Astroculture in the Twentieth Century. Isis, 104(3):641–643, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/pdfplus/10.1086/674506.pdf.

Sylvester:2013:BRB

[1078] Roshanna P. Sylvester. Book review: Boris Chertok, ed., Rockets and People. Isis, 104(3):643–644, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/674507.pdf.

Pemberton:2013:BRN

[1079] Stephen Pemberton. Book review: Nathaniel Comfort, The Science of Human Perfection: How Genes Became the Heart of American Medicine. Isis, 104(3):644–645, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/674508.pdf.

Cooter:2013:BRJ

[1080] Robert Cooter. Book review: Joseph Dumit, Drugs for Life: How Pharmaceutical Companies Define Our Health. Isis, 104(3):645–646, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
674509.pdf.

Dick:2013:BRC

[1081] Stephanie Dick. Book review: Christopher A. Fuchs, Coming of Age with Quantum Information: Notes on a Paulian Idea. Isis, 104(3):646–647, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
674510.pdf.

November:2013:BRM

[1082] Joseph November. Book review: Miguel García-Sancho, Biology, Computing, and the History of Molecular Sequencing: From Proteins to DNA, 1945–2000. Isis, 104(3):647–648, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/674511.pdf.

Roland:2013:BRS

[1083] Alex Roland. Book review: Sean F. Johnston, The Neutron’s Children: Nuclear Engineers and the Shaping of Identity. Isis, 104(3):648–649, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
674512.pdf.

Timmermann:2013:BRP

[1084] Carsten Timmermann. Book review: Peter Keating; Alberto Cambrosio, Cancer on Trial: Oncology as a New Style of Practice. Isis, 104(3):649– 650, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/pdfplus/10.
1086/674513.pdf.

Johnston:2013:BRM

[1085] Robert D. Johnston. Book review: Mark A. Largent, Vaccine: The Debate in Modern America. Isis, 104(3):650–651, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/674514.pdf.

Gingras:2013:BRA

[1086] Yves Gingras. Book review: Alfred Nordmann; Hans Radder; Gregor Schiemann, eds., Science Transformed? Debating Claims of an Epochal Break. Isis, 104(3):651–652, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/674515.pdf.

Neswald:2013:BRT

[1087] Elizabeth Neswald. Book review: Theo Plesser; Hans-Ulrich Thamer, eds., Arbeit, Leistung und Ernährung: Vom Kaiser-Wilhelm-Institut für Arbeitsphysiologie in Berlin zum Max-Planck-Institut für molekulare Physiologie und Leibniz Institut für Arbeitsforschung in Dortmund. Isis, 104(3):652–654, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

pdfplus/10.1086/674516.pdf.

Spiro:2013:BRT

[1088] Jonathan Spiro. Book review: Thomas Robertson, The Malthusian Moment: Global Population Growth and the Birth of American Environmentalism. Isis, 104(3):654–655, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/pdfplus/10.1086/674517.pdf.

Smith:2013:BRP

[1089] Robert W. Smith. Book review: Philip J. Stooke, The International Atlas of Mars Exploration: The First Five Decades. Isis, 104(3):655–656, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
674518.pdf.

Moon:2013:BRS

[1090] Suzanne Moon. Book review: Sandra Harding, ed., The Postcolonial Science and Technology Studies Reader. Isis, 104(3):656–657, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/pdfplus/10.1086/
67327960.pdf.

Popa:2013:BRT

[1091] Tiberiu Popa. Book review: Thought Experiments in Methodological and Historical Contexts. Isis, 104(3):657–658, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/674520.pdf.

Isaac:2013:BRT

[1092] Joel Isaac. Book review: Thomas S. Kuhn, The Structure of Scientific Revolutions. Isis, 104(3):658–659, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/pdfplus/10.1086/674521.pdf.

Stanley:2013:BRH

[1093] Matthew Stanley. Book review: Hélène Mialet, Hawking Incorporated: Stephen Hawking and the Anthropology of the Knowing Subject. Isis, 104 (3):659–660, September 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/pdfplus/
10.1086/674522.pdf.

Fuller:2013:BRS

[1094] Steve Fuller. Book review: Sal Restivo, Red, Black, and Objective: Science, Sociology, and Anarchism. Isis, 104(3):660, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/pdfplus/10.1086/674523.pdf.

Anonymous:2013:NCc

[1095] Anonymous. Notes on contributors. Isis, 104(3):661–665, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.1086/673280.pdf.

Anonymous:2013:Fc

[1096] Anonymous. Frontispiece. Isis, 104(3):vi, September 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/pdfplus/10.1086/674568.pdf.

Anonymous:2013:Fd

[1097] Anonymous. Frontispiece. Isis, 104(4):vi, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676391.

DeVos:2013:PMY

[1098] Paula De Vos. The “Prince of Medicine”: Yuh.anna ibn Masawayh and the foundations of the Western pharmaceutical tradition. Isis, 104(4): 667–712, December 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
674940.

Osseo-Asare:2013:SEE

[1099] Abena Dove Osseo-Asare. Scientific equity: Experiments in laboratory education in Ghana. Isis, 104(4):713–741, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/674941.

Zenderland:2013:SSW

[1100] Leila Zenderland. Social science as a “weapon of the weak”: Max Weinreich, the Yiddish Scientific Institute, and the study of culture, personality, and prejudice. Isis, 104(4):742–772, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/674942.

McCook:2013:I

[1101] Stuart McCook. Introduction. Isis, 104(4):773–776, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/674943.

Duarte:2013:BNU

[1102] Regina Horta Duarte. Between the national and the universal: Natural history networks in Latin America in the Nineteenth and Twentieth Centuries. Isis, 104(4):777–787, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/674944.

Prieto:2013:IKS

[1103] Leida Fernández Prieto. Islands of knowledge: Science and agriculture in the history of Latin America and the Caribbean. Isis, 104(4):788–797, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/674945.

Espinosa:2013:GHD

[1104] Mariola Espinosa. Globalizing the history of disease, medicine, and public health in Latin America. Isis, 104(4):798–806, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/674946.

Rodriguez:2013:BPP

[1105] Julia Rodriguez. Beyond prejudice and pride: The human sciences in Nineteenth- and Twentieth-Century Latin America. Isis, 104(4):807–817, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/674947.

Wulff:2013:LE

[1106] Karl Wulff. Letter to the Editor. Isis, 104(4):818, December 2013. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/674948.

Hall:2013:BRR

[1107] Bert S. Hall. Book review: Religion, Technology, and the Great and Little Divergences: China and Europe Compared. Isis, 104(4):819–820, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676337.

Whiteley:2013:BRM

[1108] Louise Whiteley. Book review: Medical Imaging and Philosophy: Challenges, Reflections, Actions. Isis, 104(4):820–821, December 2013. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676338.

Candela:2013:BRF

[1109] Andrea Candela. Book review: Francesco Gerali. L’opera e l’archivio spezzino di Giovanni Capellini, un geologo dell’ottocento. Luca Ciancio.

Vulcan’s Secret Forge: Explorations of the Verona Area by British Aristocrats and Veneto Naturalists during the Eighteenth Century/La fucina segreta di Vulcano: Naturalisti veneti e aristocratici britannici de Settecento alla scoperta del territorio Veronese. Isis, 104(4):821–822, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/676339.

Burian:2013:BRS

[1110] Richard Burian. Book review: Snait B. Gissis; Eva Jablonka, eds. Transformations of Lamarckism: From Subtle Fluids to Molecular Biology. Isis, 104(4):822–824, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676340.

Herring:2013:BRM

[1111] Ann Herring. Book review: Mark Harrison. Contagion: How Commerce Has Spread Disease. Isis, 104(4):824–825, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676341.

Fang:2013:BRJ

[1112] Xiaoping Fang. Book review: T. J. Hinrichs; Linda L. Barnes, eds. Chinese Medicine and Healing: An Illustrated History. Isis, 104(4):825– 826, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676342.

Pettit:2013:BRP

[1113] Michael Pettit. Book review: Peter Lamont. Extraordinary Beliefs: A Historical Approach to a Psychological Problem. Isis, 104(4):826–827, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676343.

Agar:2013:BRMa

[1114] Jon Agar. Book review: Mark Mazower. Governing the World: The History of an Idea. Isis, 104(4):827–828, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/676344.

Bork:2013:BRD

[1115] Kennard B. Bork. Book review: David R. Montgomery. The Rocks Don’t Lie: A Geologist Investigates Noah’s Flood. Isis, 104(4):828–829, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/676345.

Rocke:2013:BRJ

[1116] Alan Rocke. Book review: John Parascandola. King of Poisons: A History of Arsenic. Isis, 104(4):829, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676346.

Wisnioski:2013:BRH

[1117] Matthew Wisnioski. Book review: Henry Petroski. To Forgive Design: Understanding Failure. Isis, 104(4):829–830, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676347.

Pinch:2013:BRH

[1118] Trevor Pinch. Book review: Hillel Schwartz. Making Noise: From Babel to the Big Bang and Beyond. Isis, 104(4):830–831, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676348.

DeVorkin:2013:BRT

[1119] David H. DeVorkin. Book review: Thomas R. Williams; Michael Saladyga. Advancing Variable Star Astronomy: The Centennial History of the American Association of Variable Star Observers. Isis, 104(4):831– 832, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676349.

Rutkin:2013:BRM

[1120] H. Darrel Rutkin. Book review: Monica Azzolini. The Duke and the Stars: Astrology and Politics in Renaissance Milan. Isis, 104(4):832–833, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676350.

Wilks:2013:BRD

[1121] Ian Wilks. Book review: David Bloch. John of Salisbury on Aristotelian Science. Isis, 104(4):833–834, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676351.

VanBrummelen:2013:BRJ

[1122] Glen Van Brummelen. Book review: José Chabás and Bernard R. Goldstein. A Survey of European Astronomical Tables in the Late Middle Ages. Isis, 104(4):834–835, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676352.

Gentilcore:2013:BRD

[1123] David Gentilcore. Book review: Denise Hillard. Traité des eaux artificielles, ou, Vertus des eaux et des herbes: Le texte, ses sources et ses éditions (1483–1625): Édition critique. Isis, 104(4):835–836, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676353.

Kavey:2013:BRF

[1124] Allison Kavey. Book review: Frank Klaassen. The Transformations of Magic: Illicit Learned Magic in the Later Middle Ages and Renaissance. Isis, 104(4):836–837, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676354.

Valleriani:2013:BRM

[1125] Matteo Valleriani. Book review: Mark A. Peterson. Galileo’s Muse: Renaissance Mathematics and the Arts. Isis, 104(4):837–838, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676355.

Panza:2013:BRS

[1126] Marco Panza. Book review: Sabine Rommevaux; Maryvonne Spiesser; Maria Rosa Massa Esteve, eds. Pluralité de l’algèbre à la Renaissance. Isis, 104(4):838–839, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676356.

DeRenzi:2013:BRN

[1127] Silvia De Renzi. Book review: Nancy G. Siraisi. Communities of Learned Experience: Epistolary Medicine in the Renaissance. Isis, 104(4):839– 840, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676357.

Easton:2013:BRR

[1128] Patricia Easton. Book review: Roger Ariew. Descartes among the Scholastics. Isis, 104(4):840–841, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676358.

Deringer:2013:BRR

[1129] William Deringer. Book review: Rüdiger Campe. The Game of Probability: Literature and Calculation from Pascal to Kleist. Isis, 104(4):841– 842, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676359.

Dooley:2013:BRD

[1130] Brendan Dooley. Book review: Dimitrie Cantemir; Vlad Alexandrescu. L’immagine irraffigurabile della scienza sacro-santa. Isis, 104(4):842– 843, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676360.

Harrison:2013:BRS

[1131] Peter Harrison. Book review: Stephen Gaukroger. The Collapse of Mechanism and the Rise of Sensibility: Science and the Shaping of Modernity, 1680–1760. Isis, 104(4):843–845, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676361.

Dym:2013:BRD

[1132] Warren Alexander Dym. Book review: Dagmar Hülsenberg; Ingo Schwarz, eds. Alexander von Humboldt: Gutachten zur Steingutfertigung in Rheinsberg 1792. Isis, 104(4):845–846, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676362.

Willmoth:2013:BRD

[1133] Frances Willmoth. Book review: David Sellers. In Search of William Gascoigne: Seventeenth Century Astronomer. Isis, 104(4):846–847, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/676363.

Zuidervaart:2013:BRH

[1134] Huib J. Zuidervaart. Book review: Heiko Weber. Die Elektrisiermaschinen im 18. Jahrhundert. Isis, 104(4):847–848, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676364.

Brown:2013:BRM

[1135] Michael Brown. Book review: Michelle Allen-Emerson, ed. Sanitary Reform in Victorian Britain. Isis, 104(4):848–849, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676365.

Guttler:2013:BRJ

[1136] Nils Güttler. Book review: Jan Altmann. Zeichnen als Beobachten: Die Bildwerke der Baudin-Expedition (1800–1804). Isis, 104(4):849–850, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/676366.

Shuttleworth:2013:BRG

[1137] Sally Shuttleworth. Book review: Gowan Dawson; Bernard Lightman; Claire Brock; Marwa Elshakry; Sujit Sivasundaram; Ralph O’Connor; Roger Luckhurst; Justin Sausman, eds. Victorian Science and Literature. Isis, 104(4):850–851, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676367.

Marchand:2013:BRM

[1138] Suzanne Marchand. Book review: Modris Eksteins. Solar Dance: Van Gogh, Forgery, and the Eclipse of Certainty. Isis, 104(4):851–852, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/676368.

Buklijas:2013:BRS

[1139] Tatjana Buklijas. Book review: Sarah Ferber; Sally Wilde, eds. The Body Divided: Human Beings and Human “Material” in Modern History. Isis, 104(4):852–853, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676369.

Richards:2013:BRR

[1140] Joan L. Richards. Book review: Raymond Flood; Adrian Rice; Robin Wilson, eds. Mathematics in Victorian Britain. Isis, 104(4):853–855, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676370.

Harper:2013:BRD

[1141] Kristine C. Harper. Book review: Daniel Freund. American Sunshine: Diseases of Darkness and the Quest for Natural Light. Isis, 104(4):855, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676371.

Nyhart:2013:BRK

[1142] Lynn K. Nyhart. Book review: Kristin Johnson. Ordering Life: Karl Jordan and the Naturalist Tradition. Isis, 104(4):855–856, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676372.

Shteir:2013:BRT

[1143] Ann Shteir. Book review: Theresa M. Kelley. Clandestine Marriage: Botany and Romantic Culture. Isis, 104(4):857, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676373.

Reidy:2013:BRD

[1144] Michael S. Reidy. Book review: David N. Livingstone; Charles W. J. Withers, eds. Geographies of Nineteenth-Century Science. Isis, 104(4): 858, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676374.

Hentschel:2013:BRG

[1145] Klaus Hentschel. Book review: Günter Nagel. Wissenschaft für den Krieg: Die geheimen Arbeiten der Abteilung Forschung des Heereswaffenamtes. Isis, 104(4):859–860, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676375.

Farber:2013:BRC

[1146] Paul Lawrence Farber. Book review: Camilo Quintero Toro. Birds of Empire, Birds of Nation: A History of Science, Economy, and Conservation in United States-Colombia Relations. Isis, 104(4):860–861, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676376.

Kohlstedt:2013:BRS

[1147] Sally Gregory Kohlstedt. Book review: Susan Schulten. Mapping the Nation: History and Cartography in Nineteenth-Century America. Isis, 104 (4):861–862, December 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676377.

Hicks:2013:BRJ

[1148] Marie Hicks. Book review: Janet Abbate. Recoding Gender: Women’s Changing Participation in Computing. Isis, 104(4):862–863, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676378.

Wellerstein:2013:BRB

[1149] Alex Wellerstein. Book review: Brian Balmer. Secrecy and Science: A Historical Sociology of Biological and Chemical Warfare. Isis, 104(4): 863–864, December 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676379.

Hersch:2013:BRW

[1150] Matthew H. Hersch. Book review: William J. Clancey. Working on Mars: Voyages of Scientific Discovery with the Mars Exploration Rovers. Isis, 104(4):864, December 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676380.

Schmalzer:2013:BRX

[1151] Sigrid Schmalzer. Book review: Xiaoping Fang. Barefoot Doctors and Western Medicine in China. Isis, 104(4):864–865, December 2013. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676381.

Elcock:2013:BRN

[1152] Chris Elcock. Book review: Nicolas Langlitz. Neuropsychedelia: The Revival of Hallucinogen Research since the Decade of the Brain. Isis, 104 (4):865–866, December 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676382.

Jacyna:2013:BRM

[1153] Stephen Jacyna. Book review: Melissa M. Littlefield; Jenell M. Johnson, eds. The Neuroscientific Turn: Transdisciplinarity in the Age of the Brain. Isis, 104(4):867, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676383.

Agar:2013:BRMb

[1154] Jon Agar. Book review: Michael Sean Mahoney; Thomas Haigh, eds. Histories of Computing. Isis, 104(4):868–869, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676384.

Bromberg:2013:BRD

[1155] Joan Lisa Bromberg. Book review: David P. D. Munns. A Single Sky: How an International Community Forged the Science of Radio Astronomy. Isis, 104(4):869–870, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676385.

Takeshita:2013:BRM

[1156] Chikako Takeshita. Book review: Michelle Murphy. Seizing the Means of Reproduction: Entanglements of Feminism, Health, and Technoscience. Isis, 104(4):870–871, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676386.

Spiller:2013:BRK

[1157] James Spiller. Book review: Kendrick Oliver. To Touch the Face of God: The Sacred, the Profane, and the American Space Program, 1957–1975. Isis, 104(4):871–872, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676387.

Helmstadter:2013:BRH

[1158] Axel Helmstädter. Book review: Heiko Stoff. Wirkstoffe: Eine Wissenschaftsgeschichte der Hormone, Vitamine und Enzyme, 1920–1970. Isis, 104(4):872–873, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676388.

Bud:2013:BRK

[1159] Robert Bud. Book review: Kaushik Sunder Rajan, ed. Lively Capital: Biotechnologies, Ethics, and Governance in Global Markets. Isis, 104 (4):873–874, December 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676389.

Clegg:2013:BRP

[1160] Joshua W. Clegg. Book review: Paul Wouters; Anne Beaulieu; Andrea Scharnhorst; Sally Wyatt, eds. Virtual Knowledge: Experimenting in the Humanities and the Social Sciences. Isis, 104(4):874, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676390.

Anonymous:2013:NCd

[1161] Anonymous. Notes on contributors. Isis, 104(4):875–878, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/674950.

Anonymous:2013:IIV

[1162] Anonymous. Index to Isis, volume 104, 2013. Isis, 104(4):879–889, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/675925.

Anonymous:2013:CIV

[1163] Anonymous. Contents of Isis, volume 104. Isis, 104(4):892–894, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/675924.

Anonymous:2013:CBH

[1164] Anonymous. Current bibliography of the history of science and its cultural influences, 2013. Isis, 104(S1):i–304, December 2013. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/675358.

Anonymous:2014:Fa

[1165] Anonymous. Frontispiece. Isis, 105(1):vi, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/676815.

Heyck:2014:ORH

[1166] Hunter Heyck. The organizational revolution and the human sciences. Isis, 105(1):1–31, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
675549.

Morrison:2014:SIB

[1167] Robert Morrison. A scholarly intermediary between the Ottoman Empire and Renaissance Europe. Isis, 105(1):32–57, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/675550.

Moser:2014:MEK

[1168] Stephanie Moser. Making expert knowledge through the image: Connections between antiquarian and early modern scientific illustration. Isis, 105(1):58–99, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
675551.

Fuller:2014:NNH

[1169] Steve Fuller. Neuroscience, neurohistory, and the history of science: A tale of two brain images. Isis, 105(1):100–109, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/675552.

Smail:2014:NAH

[1170] Daniel Lord Smail. Neurohistory in action: Hoarding and the human past. Isis, 105(1):110–122, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/675553.

Casper:2014:HNI

[1171] Stephen T. Casper. History and neuroscience: An integrative legacy. Isis, 105(1):123–132, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
675554.

Stadler:2014:NBD

[1172] Max Stadler. Neurohistory is bunk?: The not-so-deep history of the postclassical mind. Isis, 105(1):133–144, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/675555.

Cooter:2014:NVW

[1173] Roger Cooter. Neural veils and the will to historical critique: Why historians of science need to take the neuro-turn seriously. Isis, 105 (1):145–154, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
675556.

Olwell:2014:ECW

[1174] Russ Olwell, David Guston, Wade Roush, and Jessica Wang. Eloge: Charles Weiner (1931–2012). Isis, 105(1):155–156, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/675557.

Richards:2014:EEH

[1175] Joan Richards. Eloge: Erwin Hiebert (1919–2012). Isis, 105(1):157– 160, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/675558.

Bordogna:2014:BRF

[1176] Francesca Bordogna. Book review: Free Will and the Human Sciences in Britain, 1870–1910 Between Mind and Nature: A History of Psychology. Isis, 105(1):161–163, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/675559.

Brown:2014:BRS

[1177] Mark B. Brown. Book review: Science, Democracy, and the American University: From the Civil War to the Cold War. Isis, 105(1):164–166, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/675560.

Shank:2014:BRC

[1178] Michael H. Shank. Book review: The Copernican Question: Prognostication, Skepticism, and Celestial Order. Isis, 105(1):167–176, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/675561. See [1179, 1180].

Westman:2014:RMS

[1179] Robert S. Westman. Reply to Michael Shank. Isis, 105(1):177–184, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/675714. See

[1178, 1180].

Shank:2014:R

[1180] Michael H. Shank. Rejoinder. Isis, 105(1):185–187, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/675715. See [1178, 1179].

Wilson:2014:BRRa

[1181] Andrew D. Wilson. Book review: Ørsted Triumphator? Dan Charly Christensen. Hans Christian Ørsted: Reading Nature’s Mind. Isis, 105 (1):188–192, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
675562.

Aldana:2014:BRH

[1182] Gerardo Aldana. Book review: Harvey M. Bricker; Victoria R. Bricker. Astronomy in the Maya Codices. Isis, 105(1):193–195, March 2014. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676751.

Scarborough:2014:BRN

[1183] John Scarborough. Book review: Nicholas Everett. The Alphabet of Galen: Pharmacy from Antiquity to the Middle Ages: A Critical Edition of the Latin Text with English Translation and Commentary. Isis, 105(1):195–196, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676752.

Chimisso:2014:BRM

[1184] Cristina Chimisso. Book review: Miriam Focaccia and Sandra Linguerri, ed. Dizionario biografico delle scienziate italiane (secoli XVIII–XX); Volume 1: Architette, chimiche, fisiche, dottoresse. Dizionario biografico delle scienziate italiane (secoli XVIII–XX). Volume 2: Matematiche, astronome, naturaliste. Isis, 105(1):196–197, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/676753.

Barnes:2014:BRJ

[1185] Trevor Barnes. Book review: Jeanne Haffner. The View from Above: The Science of Social Space. Isis, 105(1):197–198, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676754.

Wazeck:2014:BRM

[1186] Milena Wazeck. Book review: Michael Hagner, ed. Wissenschaft und Demokratie. Isis, 105(1):198–199, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676755.

Parascandola:2014:BRB

[1187] John Parascandola. Book review: Brian B. Hoffman. Adrenaline. Isis, 105(1):199, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676756.

Sime:2014:BRH

[1188] Ruth Lewin Sime. Book review: Horst Kant; Carsten Reinhardt, eds. 100 Jahre Kaiser-Wilhelm-/Max-Planck-Institut für Chemie (Otto-Hahn-Institut): Facetten seiner Geschichte. (Veröffentlichungen aus dem Archiv der Max-Planck-Gesellschaft, 22.). Isis, 105(1):199–200, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676757.

Gossin:2014:BRN

[1189] Pamela Gossin. Book review: Neelam Kumar, ed. Gender and Science: Studies across Cultures. Isis, 105(1):200–202, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676758.

Josephson:2014:BRM

[1190] Paul Josephson. Book review: Milton Leitenberg; Raymond A. Zilinskas; Jens H. Kuhn. The Soviet Biological Weapons Program: A History. Isis, 105(1):202–203, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676759.

Hounshell:2014:BRD

[1191] David A. Hounshell. Book review: David E. Nye. America’s Assembly Line. Isis, 105(1):203–204, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676760.

Coen:2014:BRJ

[1192] Deborah R. Coen. Book review: Jan Plamper; Benjamin Lazier, eds. Fear: Across the Disciplines. Isis, 105(1):204–205, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676761.

Brooke:2014:BRM

[1193] John Hedley Brooke. Book review: Michael Ruse, ed. The Cambridge Encyclopedia of Darwin and Evolutionary Thought. Isis, 105(1):205– 207, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676762.

Hogendijk:2014:BRG

[1194] Jan P. Hogendijk. Book review: Glen Van Brummelen. Heavenly Mathematics: The Forgotten Art of Spherical Trigonometry. Isis, 105(1):207– 208, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676763.

Totelin:2014:BRJ

[1195] Laurence Totelin. Book review: Jacques Jouanna; Neil Allies; Philip van der Eijk, eds. Greek Medicine from Hippocrates to Galen: Selected Papers. Isis, 105(1):208, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676764.

Irby:2014:BRV

[1196] Georgia Irby. Book review: Vivian Nutton. Ancient Medicine. Isis, 105(1):209–210, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676765.

Laird:2014:BRA

[1197] W. R. Laird. Book review: Archimede Latino: Iacopo da San Cassiano e il “Corpus” Archimedeo alla metà del Quattrocento, con edizione della “Circuli dimensio” e della “Quadratura parabolae”, by Paolo d’Alessandro; Pier Daniele Napolitani. Isis, 105(1):210–211, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676766.

Brentjes:2014:BRN

[1198] Sonja Brentjes. Book review: Nader El-Bizri, ed. On Arithmetic and Geometry: An Arabic Critical Edition and English Translation of EPIS-TLES 1 and 2. Isis, 105(1):211–212, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676767.

Ruys:2014:BRM

[1199] Juanita Feros Ruys. Book review: Mary Franklin-Brown. Reading the World: Encyclopedic Writing in the Scholastic Age. Isis, 105(1):212– 213, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676768.

Hayton:2014:BRM

[1200] Darin Hayton. Book review: Michael A. Ryan. A Kingdom of Stargazers: Astrology and Authority in the Late Medieval Crown of Aragon. Isis, 105(1):213–214, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676769.

Olshin:2014:BRG

[1201] Benjamin B. Olshin. Book review: Gerhard Wolf; Joseph Connors; Louis A. Waldman, eds. Colors between Two Worlds: The Florentine Codex of Bernardino de Sahagún. Isis, 105(1):214–215, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676770.

Wilson:2014:BRRb

[1202] Catherine Wilson. Book review: Reid Barbour; David Norbrook; Maria Cristina Zerbino, eds. The Works of Lucy Hutchinson. Isis, 105(1):216– 217, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676771.

Jesseph:2014:BRC

[1203] Douglas Jesseph. Book review: Carlo Borghero; Tomaso Berni Canani. Les cartésiens face à Newton: Philosophie, science et religion dans la première moitié du XVIIIe siècle. Isis, 105(1):217, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676772.

Pugliano:2014:BRP

[1204] Valentina Pugliano. Book review: Paula Findlen, ed. Early Modern Things: Objects and Their Histories, 1500–1800. Isis, 105(1):218–219, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/676773.

Reeves:2014:BRS

[1205] Nicky Reeves. Book review: Stephen J. Hornsby. Surveyors of Empire: Samuel Holland, J. F. W. Des Barres, and the Making of “The Atlantic Neptune”. Isis, 105(1):219–220, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676774.

Gabbey:2014:BRC

[1206] Alan Gabbey. Book review: Christoph Lüthy. David Gorlæus (1591– 1612): An Enigmatic Figure in the History of Philosophy and Science. Isis, 105(1):220–221, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676775.

Emerson:2014:BRI

[1207] Roger Emerson. Book review: Iain McDaniel. Adam Ferguson in the Scottish Enlightenment: The Roman Past and Europe’s Future. Isis, 105 (1):221, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676776.

Luthy:2014:BRG

[1208] Christoph Lüthy. Book review: Gideon Manning, ed. Matter and Form in Early Modern Science and Philosophy. Isis, 105(1):222–223, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676777.

Henry:2014:BRP

[1209] John Henry. Book review: Paul Kléber Monod. Solomon’s Secret Arts: The Occult in the Age of Enlightenment. Isis, 105(1):223–224, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676778.

Hoquet:2014:BRA

[1210] Thierry Hoquet. Book review: Adrien Paschoud; Nathalie Vuillemin, eds.Penser l’ordre naturel, 1680–1810. Isis, 105(1):224–225, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676779.

Stewart:2014:BRK

[1211] Larry Stewart. Book review: Kenneth Quickenden; Sally Baggott; Malcolm Dick, eds. Matthew Boulton: Enterprising Industrialist of the Enlightenment. Isis, 105(1):225–226, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676780.

Soukup:2014:BRL

[1212] Rudolf Werner Soukup. Book review: Lioba Wagner. Alchemie und Naturwissenschaft: Über die Entstehung neuer Ideen an der Reibungsfläche zweier Weltbilder: Gezeigt an Paracelsus, Robert Boyle und Isaac Newton. Isis, 105(1):226, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676781.

Spanagel:2014:BRC

[1213] David Spanagel. Book review: Charles Boewe. The Life of C. S. Rafinesque, a Man of Uncommon Zeal. Isis, 105(1):227–228, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676782.

Farber:2014:BRE

[1214] Paul Lawrence Farber. Book review: Edward H. Burtt; William E. Davis. Alexander Wilson: The Scot Who Founded American Ornithology. Isis, 105(1):228, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676783.

Cohen:2014:BRD

[1215] Claudine Cohen. Book review: Deborah R. Coen. The Earthquake Observers: Disaster Science from Lisbon to Richter. Isis, 105(1):228–229, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/676784.

Sauer:2014:BRJ

[1216] Tilman Sauer. Book review: Jeremy Gray. Henri Poincaré: A Scientific Biography. Isis, 105(1):229–231, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676785.

MacDougall:2014:BRD

[1217] Robert MacDougall. Book review: David Hochfelder. The Telegraph in America, 1832–1920. Isis, 105(1):231, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/676786.

Bijsterveld:2014:BRA

[1218] Karin Bijsterveld. Book review: Alexandra Hui. The Psychophysical Ear: Musical Experiments, Experimental Sounds, 1840–1910. Isis, 105(1):232– 233, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676787.

Rupke:2014:BRA

[1219] Nicolaas Rupke. Book review: Alexander von Humboldt; Vera M. Kutzinski; Ottmar Ette; J. Ryan Poynter, eds. Views of the Cordilleras and Monuments of the Indigenous Peoples of the Americas: A Critical Edition. Isis, 105(1):233–234, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676788.

Scull:2014:BRS

[1220] Andrew Scull. Book review: L. Stephen Jacyna; Stephen T. Casper, eds. The Neurological Patient in History. Isis, 105(1):234–235, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676789.

Novoa:2014:BRT

[1221] Adriana Novoa. Book review: Travis Landry. Subversive Seduction: Darwin, Sexual Selection, and the Spanish Novel. Isis, 105(1):235–236, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676790.

Chakrabarti:2014:BRP

[1222] Pratik Chakrabarti. Book review: Prakash Kumar. Indigo Plantations and Science in Colonial India. Isis, 105(1):236–237, March 2014. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676791.

Evans:2014:BRR

[1223] Sterling Evans. Book review: Russell M. Lawson. Frontier Naturalist: Jean Louis Berlandier and the Exploration of Northern Mexico and Texas. Isis, 105(1):237–238, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/676792.

Stoff:2014:BRC

[1224] Heiko Stoff. Book review: Cheryl A. Logan. Hormones, Heredity, and Race: Spectacular Failure in Interwar Vienna. Isis, 105(1):238–239, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/676793.

Brenner:2014:BRE

[1225] Anastasios Brenner. Book review: Émile Meyerson; Eva Telkes-Klein; Bernadette Bensaude-Vincent, eds. Mélanges: Petites pieces inédites. Isis, 105(1):239–240, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676794.

Heringman:2014:BRJ

[1226] Noah Heringman. Book review: Janelle A. Schwartz. Worm Work: Recasting Romanticism. Isis, 105(1):240–241, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/676795.

Beyler:2014:BRT

[1227] Richard H. Beyler. Book review: Todd H. Weir, ed. Monism: Science, Philosophy, Religion, and the History of a Worldview. Isis, 105(1):241– 242, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676796.

Adcock:2014:BRW

[1228] Tina Adcock. Book review: William F. Althoff. Arctic Mission: 90 North by Airship and Submarine. Isis, 105(1):242–243, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676797.

Hull:2014:BRD

[1229] James Hull. Book review: Donald Avery. Pathogens for War: Biological Weapons, Canadian Life Scientists, and North American Biodefence. Isis, 105(1):243–244, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676798.

Rojas:2014:BRH

[1230] Raul Rojas. Book review: Herbert Bruderer. Konrad Zuse und die Schweiz: Wer hat den Computer erfunden? Isis, 105(1):244–245, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676799.

Gouyon:2014:BRM

[1231] Jean-Baptiste Gouyon. Book review: Marcel Chotkowski LaFollette. Science on American Television: A History. Isis, 105(1):245, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676800.

Smocovitis:2014:BRW

[1232] Vassiliki Betty Smocovitis. Book review: Wendy Dathan. The Reindeer Botanist: Alf Erling Porsild, 1901–1977. Isis, 105(1):246–247, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676801.

Kant:2014:BRS

[1233] Horst Kant. Book review: Silke Fengler; Carola Sachse, eds. Kernforschung in Österreich: Wandlungen eines interdisziplinären Forschungsfeldes 1900–1978. Isis, 105(1):247–248, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676802.

Messeri:2014:BRE

[1234] Lisa Messeri. Book review: Elizabeth A. Kessler. Picturing the Cosmos: Hubble Space Telescope Images and the Astronomical Sublime. Isis, 105 (1):248–249, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676803.

Vertesi:2014:BRW

[1235] Janet Vertesi. Book review: W. Patrick McCray.The Visioneers: How a Group of Elite Scientists Pursued Space Colonies, Nanotechnologies, and a Limitless Future. Isis, 105(1):249–250, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676804.

Ede:2014:BRE

[1236] Andrew Ede. Book review: Edwin A. Martini. Agent Orange: History, Science, and the Politics of Uncertainty. Isis, 105(1):250–251, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676805.

Johnson:2014:BRC

[1237] Ann Johnson. Book review: Cyrus C. M. Mody. Instrumental Community: Probe Microscopy and the Path to Nanotechnology. Isis, 105(1):251– 252, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676806.

Gere:2014:BRN

[1238] Cathy Gere. Book review: Nikolas Rose; Joelle M. Abi-Rached. Neuro: The New Brain Sciences and the Management of the Mind. Isis, 105 (1):252–253, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676807.

Jewett:2014:BRM

[1239] Andrew Jewett. Book review: Mark Solovey. Shaky Foundations: The Politics–Patronage–Social Science Nexus in Cold War America. Isis, 105(1):253–254, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676808.

Egan:2014:BRS

[1240] Michael Egan. Book review: Sarah A. Vogel. Is It Safe? BPA and the Struggle to Define the Safety of Chemicals. Isis, 105(1):254, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676809.

Hamerla:2014:BRJ

[1241] Rich Hamerla. Book review: John R. Walker. Britain and Disarmament: The U.K. and Nuclear, Biological, and Chemical Weapons Arms Control and Programmes, 1956–1975. Isis, 105(1):254–255, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676810.

Balmer:2014:BRA

[1242] Brian Balmer. Book review: Audra J. Wolfe. Competing with the Soviets: Science, Technology, and the State in Cold War America. Isis, 105(1):255–256, March 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676811.

Strasser:2014:BRD

[1243] Bruno J. Strasser. Book review: Dominique Pestre. À contre-science: Politiques et savoirs des sociétés contemporaines. Isis, 105(1):256–257, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/676812.

Kirschenmann:2014:BRP

[1244] Peter P. Kirschenmann. Book review: Paul Hoyningen-Huene. Systematicity: The Nature of Science. Isis, 105(1):257–258, March 2014. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676813.

Anonymous:2014:NCa

[1245] Anonymous. Notes on contributors. Isis, 105(1):259–263, March 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/675564.

Anonymous:2014:Fb

[1246] Anonymous. Frontispiece. Isis, 105(2):vi, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/678036.

Bergwik:2014:ASH

[1247] Staffan Bergwik. An assemblage of science and home: The gendered lifestyle of svante Arrhenius and early Twentieth-century physical chemistry. Isis, 105(2):265–291, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676567.

DalPrete:2014:BWE

[1248] Ivano Dal Prete. “Being the world eternal …: The age of the Earth in Renaissance Italy. Isis, 105(2):292–317, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/676568.

Heggie:2014:WIE

[1249] Vanessa Heggie. Why isn’t exploration a science? Isis, 105(2):318– 334, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/676569.

Rozwadowski:2014:I

[1250] Helen M. Rozwadowski. Introduction. Isis, 105(2):335–337, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676570.

Reidy:2014:SBS

[1251] Michael S. Reidy and Helen M. Rozwadowski. The spaces in between: Science, ocean, empire. Isis, 105(2):338–351, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/676571.

Hamblin:2014:SOS

[1252] Jacob Darwin Hamblin. Seeing the oceans in the shadow of Bergen values. Isis, 105(2):352–363, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/676573.

Hubbard:2014:WPP

[1253] Jennifer Hubbard. In the wake of politics: The political and economic construction of fisheries biology, 1860–1970. Isis, 105(2):364–378, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676572.

Oreskes:2014:SOV

[1254] Naomi Oreskes. Scaling up our vision. Isis, 105(2):379–391, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676574.

Anonymous:2014:HPC

[1255] Anonymous. HSS 2013 prize citations. Isis, 105(2):392–397, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/677205.

Gorelik:2014:LE

[1256] Gennady Gorelik. Letter to the Editor. Isis, 105(2):398–399, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676575.

Ganeri:2014:R

[1257] Jonardon Ganeri. In reply. Isis, 105(2):399–400, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/676587.

Cohen:2014:LE

[1258] H. Floris Cohen. Letter to the Editor. Isis, 105(2):401, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/677658.

Gal:2014:R

[1259] Ofer Gal. In reply. Isis, 105(2):401, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/677659.

Anonymous:2014:E

[1260] Anonymous. Errata. Isis, 105(2):402, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/677204.

Guicciardini:2014:BRN

[1261] Niccolò Guicciardini. Book review: The Newton Project The Chymistry of Isaac Newton The Newton Project Canada. Isis, 105(2):403–409, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/676577.

Hallam:2014:BRC

[1262] Tony Hallam. Book review: The Continental Drift Controversy. Isis, 105(2):410–412, June 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
676580.

Anderson:2014:BRW

[1263] Warwick Anderson. Book review: Writing History in the Age of Biomedicine. Isis, 105(2):413–414, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/677980.

Osborne:2014:BRF

[1264] Michael A. Osborne. Book review: Figures of Medicine: Blood, Face Transplants, Parasites. Isis, 105(2):414, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/677981.

Moore:2014:BRC

[1265] Richard Moore. Book review: Churchill’s Bomb: How the United States Overtook Britain in the First Nuclear Arms Race. Isis, 105(2):415, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/677982.

Shen:2014:BRD

[1266] Grace Yen Shen. Book review: Developmental Fairy Tales: Evolutionary Thinking and Modern Chinese Culture. Isis, 105(2):416, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/677983.

Hestmark:2014:BRD

[1267] Geir Hestmark. Book review: Discovering the Ice Ages: International Reception and Consequences for a Historical Understanding of Climate. Isis, 105(2):416–417, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/677984.

Dowbiggin:2014:BRI

[1268] Ian Dowbiggin. Book review: An Image of God: The Catholic Struggle with Eugenics. Isis, 105(2):417–418, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/677985.

Finkelstein:2014:BRW

[1269] Gabriel Finkelstein. Book review: Wissenschaftsatlas of Heidelberg University: Spatio-Temporal Relations of Academic Knowledge Production. Isis, 105(2):418–419, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/677986.

Stephens:2014:BRA

[1270] Elizabeth Stephens. Book review: Anatomical Dissection in Enlightenment England and Beyond: Autopsy, Pathology, and Display. Isis, 105(2):419–420, June 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
677987.

Barrera-Osorio:2014:BRM

[1271] Antonio Barrera-Osorio. Book review: Missionary Scientists: Jesuit Science in Spanish South America, 1570–1810. Isis, 105(2):420–421, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/677988.

Beretta:2014:BRS

[1272] Marco Beretta. Book review: The Secrets of Alchemy. Isis, 105(2):421– 422, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/677989.

Josefowicz:2014:BRN

[1273] Diane Greco Josefowicz. Book review: Newton, Maxwell, Marx: Spirit, Freedom, and the Scientific Vision. Isis, 105(2):422–423, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/677990.

Fara:2014:BRS

[1274] Patricia Fara. Book review: Spark from the Deep: How Shocking Experiments with Strongly Electric Fish Powered Scientific Discovery. Isis, 105(2):423–424, June 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
677991.

Cohen-Cole:2014:BRM

[1275] Jamie Cohen-Cole. Book review: Memory: Fragments of a Modern History. Isis, 105(2):424–425, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/677992.

Pepper:2014:BRT

[1276] Simon Pepper. Book review: Trattato sulle fortificazioni. Isis, 105(2): 426, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/677993.

Vermeir:2014:BRA

[1277] Koen Vermeir. Book review: The Automaton in English Renaissance Literature. Isis, 105(2):426–427, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/677994.

Smith:2014:BRP

[1278] Pamela H. Smith. Book review: Pliny and the Artistic Culture of the Italian Renaissance. Isis, 105(2):427–429, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/677995.

Olshin:2014:BRW

[1279] Benjamin B. Olshin. Book review: Walter Ralegh’s History of the World and the Historical Culture of the Late Renaissance. Isis, 105(2):429– 430, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/677996.

Grell:2014:BRP

[1280] Ole Peter Grell. Book review: Panaceia’s Daughters: Noblewomen as Healers in Early Modern Germany. Isis, 105(2):430–431, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/677997.

Valleriani:2014:BRT

[1281] Matteo Valleriani. Book review: Il telescopio di Galileo: Una storia europea. Isis, 105(2):431–432, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/677998.

Conforti:2014:BRW

[1282] Maria Conforti. Book review: Watching Vesuvius: A History of Science and Culture in Early Modern Italy. Isis, 105(2):432–433, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/677999.

Hyder:2014:BRK

[1283] David Hyder. Book review: Kant’s Construction of Nature: A Reading of the Metaphysical Foundations of Natural Science. Isis, 105(2):433– 435, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/678000.

Marcaida:2014:BRB

[1284] José Ramón Marcaida. Book review: Baroque Science. Isis, 105(2):435– 436, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/678001.

Jenkins:2014:BRM

[1285] Jane Jenkins. Book review: The Mechanization of Natural Philosophy. Isis, 105(2):436–437, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/678002.

Fraser:2014:BRR

[1286] Craig Fraser. Book review: Roger Hahn, ed. Correspondance de Pierre Simon Laplace (1749–1827). Isis, 105(2):437–438, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/678003.

Gaukroger:2014:BRN

[1287] Stephen Gaukroger. Book review: Newton and the Netherlands: How Isaac Newton Was Fashioned in the Dutch Republic. Isis, 105(2):438– 439, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/678004.

Rusnock:2014:BRW

[1288] Andrea Rusnock. Book review: William Petty on the Order of Nature: An Unpublished Manuscript Treatise. Isis, 105(2):439–440, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/678005.

Peterschmitt:2014:BRP

[1289] Luc Peterschmitt. Book review: La pensée en cornue: Matérialisme, alchimie et saviors secrets à l’âge classique. Isis, 105(2):440–441, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/678006.

Jorink:2014:BRI

[1290] Eric Jorink. Book review: The Insect and the Image: Visualizing Nature in Early Modern Europe, 1500–1700. Isis, 105(2):441–442, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/678007.

Winsor:2014:BRW

[1291] Mary Pickard Winsor. Book review: Web of Nature: Martin Lister (1639–1712), the First Arachnologist. Isis, 105(2):442–443, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/678008.

Waddell:2014:BRE

[1292] Mark A. Waddell. Book review: Egyptian Oedipus: Athanasius Kircher and the Secrets of Antiquity. Isis, 105(2):443–444, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/678009.

Schuster:2014:BRI

[1293] John A. Schuster. Book review: Isaac Beeckman on Matter and Motion: Mechanical Philosophy in the Making. Isis, 105(2):444–445, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/678010.

Prakash:2014:BRE

[1294] Gyan Prakash. Book review: Everyday Technology: Machines and the Making of India’s Modernity. Isis, 105(2):445–446, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/678011.

Anker:2014:BRS

[1295] Peder Anker. Book review: Shaping Ecology: The Life of Arthur Tansley. Isis, 105(2):446–447, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/678012.

Arnold:2014:BRB

[1296] David Arnold. Book review: On Barak, On Time: Technology and Temporality in Modern Egypt. Isis, 105(2):447–448, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/678013.

Otsubo:2014:BRA

[1297] Sumiko Otsubo. Book review: Alexander R. Bay, Beriberi in Modern Japan: The Making of a National Disease. Isis, 105(2):448–449, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/678014.

Ratcliff:2014:BRB

[1298] Jessica Ratcliff. Book review: Barbara J. Becker, Unravelling Starlight: William and Margaret Huggins and the Rise of the New Astronomy. Isis, 105(2):449–450, June 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
678015.

Hale:2014:BRP

[1299] Piers J. Hale. Book review: Peter J. Bowler, Darwin Deleted: Imagining a World without Darwin. Isis, 105(2):450–451, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/678016.

Oldroyd:2014:BRA

[1300] David Oldroyd. Book review: Adelene Buckland, Novel Science: Fiction and the Invention of Nineteenth-Century Geology. Isis, 105(2):451–453, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/678017.

Sauerteig:2014:BRM

[1301] Lutz Sauerteig. Book review: Michael Hagner. Der Hauslehrer: Die Geschichte eines Kriminalfalls: Erziehung, Sexualität und Medien um 1900. Isis, 105(2):453–454, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/678018.

Kojevnikov:2014:BRD

[1302] Alexei Kojevnikov. Book review: Dieter Hoffmann; Mark Walker; Ann M. Hentschel, eds. The German Physical Society in the Third Reich: Physicists between Autonomy and Accommodation. Isis, 105(2):454–455, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/678019.

Muller-Pohl:2014:BRC

[1303] Simone Müller-Pohl. Book review: Christian Holtorf. Der erste Draht zur Neuen Welt: Die Verlegung des transatlantischen Telegrafenkabels. Isis, 105(2):455–456, June 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
678020.

Chiang:2014:BRT

[1304] Howard Chiang. Book review: Tong Lam. A Passion for Facts: Social Surveys and the Construction of the Chinese Nation-State, 1900– 1949. Isis, 105(2):456–457, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/678021.

Kent:2014:BRE

[1305] Deborah Kent. Book review: Elaine McKinnon Riehm; Frances Hoffman. Turbulent Times in Mathematics: The Life of J. C. Fields and the History of the Fields Medal. Isis, 105(2):457–458, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/678022.

Eghigian:2014:BRD

[1306] Greg Eghigian. Book review: David G. Schuster. Neurasthenic Nation: America’s Search for Health, Happiness, and Comfort, 1869–1920. Isis, 105(2):458–459, June 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
678023.

Bankole-Medina:2014:BRM

[1307] Katherine Bankole-Medina. Book review: Marli F. Weiner; Mazie Hough. Sex, Sickness, and Slavery: Illness in the Antebellum South. Isis, 105(2): 459–460, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/678024.

Abbate:2014:BRF

[1308] Janet Abbate. Book review: Finn Brunton. Spam: A Shadow History of the Internet. Isis, 105(2):460–461, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/678025.

Bocking:2014:BRH

[1309] Stephen Bocking. Book review: Hannah Gay. The Silwood Circle: A History of Ecology and the Making of Scientific Careers in Late Twentieth-Century Britain. Isis, 105(2):461–462, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/678026.

Schuring:2014:BRR

[1310] Michael Schüring. Book review: Rolf-Jürgen Gleitsmann; Günther Oetzel. Fortschrittsfeinde im Atomzeitalter? Protest und Innovationsmanagement am Beispiel der frühen Kernenergiepläne der Bundesrepublik Deutschland. Isis, 105(2):462–463, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/678027.

Allen:2014:BRS

[1311] Garland E. Allen. Book review: Sheldon Krimsky; Jeremy Gruber, eds. Genetic Explanations: Sense and Nonsense. Isis, 105(2):463–465, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/678028.

Stahnisch:2014:BRS

[1312] Frank W. Stahnisch. Book review: Stefan Sperling. Reasons of Conscience: The Bioethics Debate in Germany. Isis, 105(2):465–466, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/678029.

vanderHorst:2014:BRM

[1313] Frank van der Horst. Book review: Marga Vicedo. The Nature and Nurture of Love: From Imprinting to Attachment in Cold War America. Isis, 105(2):466–467, June 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
678030.

Rocha:2014:BRC

[1314] Leon Antonio Rocha. Book review: Chunjuan Nancy Wei; Darryl E. Brock (Editors). Mr. Science and Chairman Mao’s Cultural Revolution: Science and Technology in Modern China. xxxii + 391 pp., bibl., index. Lanham, Md.: Lexington Books, 2012. $85 (cloth). Isis, 105(2):467– 468, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/678031.

Anonymous:2014:NCb

[1315] Anonymous. Notes on contributors. Isis, 105(2):469–472, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/676579.

Anonymous:2014:Fc

[1316] Anonymous. Frontispiece. Isis, 105(3):vi, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/679250.

Lightman:2014:FYI

[1317] Bernard Lightman. Farewell from the York Isis Office. Isis, 105(3): 475–476, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
678167.

Crenner:2014:RLN

[1318] Christopher Crenner. Race and laboratory norms: The critical insights of Julian Herman Lewis (1891–1989). Isis, 105(3):477–507, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/678168.

Hoquet:2014:BAB

[1319] Thierry Hoquet. Botanical authority: Benjamin Delessert’s collections between travelers and Candolle’s natural method (1803–1847). Isis, 105 (3):508–539, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
678169.

Phillips:2014:AMM

[1320] Christopher J. Phillips. In accordance with a “more majestic order”: The new math and the nature of mathematics at midcentury. Isis, 105 (3):540–563, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
678170.

Milam:2014:I

[1321] Erika Lorraine Milam. Introduction. Isis, 105(3):564–568, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/678171.

Holmes:2014:GRE

[1322] Brooke Holmes. Greco–Roman ethics and the naturalistic fantasy. Isis, 105(3):569–578, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/678172.

Daston:2014:NFM

[1323] Lorraine Daston. The naturalistic fallacy is modern. Isis, 105(3):579–587, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/678173.

Stanley:2014:OPN

[1324] Matthew Stanley. From ought to is: Physics and the naturalistic fallacy. Isis, 105(3):588–595, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/678174.

Milam:2014:FSC

[1325] Erika Lorraine Milam. A field study of con games. Isis, 105(3):596–605, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/678175.

Anderson:2014:GAO

[1326] Warwick Anderson. Getting ahead of one’s self?: The common culture of immunology and philosophy. Isis, 105(3):606–616, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/678176.

Rocke:2014:EMR

[1327] Alan Rocke. Eloge: Mark Russell Finlay (15 September 1960–6 October 2013). Isis, 105(3):617–620, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/678177.

Sheffield:2014:BRM

[1328] Suzanne Le-May Sheffield. Book review: Marianne Gosztonyi Ainley; Marelene Rayner-Canham; Geoff Rayner-Canham; Alison Prentice, eds. Creating Complicated Lives: Women and Science at English-Canadian Universities, 1880–1980. Isis, 105(3):621, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/679125.

Hentschel:2014:BRO

[1329] Klaus Hentschel. Book review: Olaf Breidbach; Kerrin Klinger; Matthias Müller. Camera Obscura: Die Dunkelkammer in ihrer historischen Entwicklung. Isis, 105(3):622, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/679126.

Olley:2014:BRM

[1330] Allan Olley. Book review: Martin Campbell-Kelly; William Aspray; Nathan Ensmenger; Jeffery R. Yost. Computer: A History of the Information Machine. Isis, 105(3):622–623, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/679127.

Tatarchenko:2014:BRL

[1331] Ksenia Tatarchenko. Book review: Loren Graham. Lonely Ideas: Can Russia Compete? Isis, 105(3):623–624, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/679128.

Lettkemann:2014:BRK

[1332] Eric Lettkemann. Book review: Klaus Hentschel, ed. Zur Geschichte von Forschungstechnologien: Generizität-Interstitialität-Transfer. Isis, 105 (3):624–625, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679129.

Bruno:2014:BRP

[1333] Andy Bruno. Book review: Paul Josephson; Nicolai Dronin; Ruben Mnatsakanian; Aleh Cherp; Dmitry Efremenko; Vladislav Larin. An Environmental History of Russia. Isis, 105(3):625–626, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/679130.

Cittadino:2014:BRM

[1334] Eugene Cittadino. Book review: Michael Ruse. The Gaia Hypothesis: Science on a Pagan Planet. Isis, 105(3):626–627, September 2014. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/679131.

Jenks:2014:BRS

[1335] Andrew Jenks. Book review: Steven A. Usitalo. The Invention of Mikhail Lomonosov: A Russian National Myth. Isis, 105(3):627–628, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/679132.

Delaney:2014:BRM

[1336] Paul Delaney. Book review: Margaret Weitekamp; David DeVorkin; Diane Kidd. Pluto’s Secret: An Icy World’s Tale of Discovery. Isis, 105 (3):628–629, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679133.

Gascoigne:2014:BRG

[1337] John Gascoigne. Book review: Glyn Williams. Naturalists at Sea: Scientific Travellers from Dampier to Darwin. Isis, 105(3):629–630, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/679134.

Gregory:2014:BRF

[1338] Andrew Gregory. Book review: Francesca Alesse; Franco Ferrari, eds. Epinomide: Studi sull’opera e la sua ricezione. Isis, 105(3):630–631, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679135.

Finocchiaro:2014:BRP

[1339] Maurice A. Finocchiaro. Book review: Paul Richard Blum; Peter Henneveld. Giordano Bruno: An Introduction. Isis, 105(3):631–632, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/679136.

Veit:2014:BRN

[1340] Raphaela Veit. Book review: Nahyan Fancy. Science and Religion in Mamluk Egypt: Ibn al-Nafis, Pulmonary Transit, and Bodily Resurrection. Isis, 105(3):632–633, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/679137.

Black:2014:BRJ

[1341] Winston Black. Book reviews: John Flood; James R. Ginther; Joseph W. Goering, eds. Robert Grosseteste and His Intellectual Milieu: New Editions and Studies. Greti Dinkova-Bruun; Giles E. M. Gasper; Michael Huxtable; Tom C. B. McLeish; Cecilia Panti; Hannah Smithson. The Dimensions of Colour: Robert Grosseteste’s De colore. Isis, 105(3):633–635, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679138.

Wallis:2014:BRW

[1342] Faith Wallis. Book review: Winston Black, ed. Anglicanus ortus: A Verse Herbal of the Twelfth Century. Isis, 105(3):635–636, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/679139.

Worthen:2014:BRP

[1343] Shana Worthen. Book review: Pamela O. Long. Artisan/Practitioners and the Rise of the New Sciences, 1400–1600. Isis, 105(3):636–637, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679140.

Biard:2014:BRN

[1344] Joël Biard. Book review: Nicole Oresme; Stefano Caroti; Jean Celeyrette; Stefan Kirschner; Edmond Mazet, eds. Questiones super Physicam (Books I–VII). Isis, 105(3):637–638, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/679141.

Weill-Parot:2014:BRM

[1345] Nicolas Weill-Parot. Book review: Mary Quinlan-McGrath. Influences: Art, Optics, and Astrology in the Italian Renaissance. Isis, 105(3): 638–639, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679142.

Cattaneo:2014:BRS

[1346] Angelo Cattaneo. Book review: Sean Roberts. Printing a Mediterranean World: Florence, Constantinople, and the Renaissance of Geography. Isis, 105(3):639–640, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/679143.

Chen-Morris:2014:BRP

[1347] Raz Chen-Morris. Book review: Patrick J. Boner. Kepler’s Cosmological Synthesis: Astrology, Mechanism, and the Soul. Isis, 105(3):640–641, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679144.

Lindemann:2014:BRK

[1348] Mary Lindemann. Book review: Kristy Wilson Bowers. Plague and Public Health in Early Modern Seville. Isis, 105(3):641–642, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/679145.

Leonhard:2014:BRH

[1349] Karin Leonhard. Book review: Hanneke Grootenboer. Treasuring the Gaze: Intimate Vision in Late Eighteenth-Century Eye Miniatures. Isis, 105(3):642–643, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/679146.

Nickelsen:2014:BRS

[1350] Kärin Nickelsen. Book review: Santiago Madriñán. Nikolaus Joseph Jacquin’s American Plants: Botanical Expedition to the Caribbean (1754– 1759) and the Publication of the Selectarum Stirpium Americanarum Historia. Isis, 105(3):643–644, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/679147.

Findlen:2014:BRT

[1351] Paula Findlen. Book review: Thomas F. Mayer. The Roman Inquisition: A Papal Bureaucracy and Its Laws in the Age of Galileo. Isis, 105(3): 644–645, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679148.

Steinke:2014:BRM

[1352] Hubert Steinke. Book review: Miriam Nicoli. Les savants et les livres: Autour d’Albrecht von Haller (1708–1777) et Samuel-Auguste Tissot (1728–1797). Isis, 105(3):645–646, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/679149.

Terrall:2014:BRC

[1353] Mary Terrall. Book review: E. C. Spary. Eating the Enlightenment: Food and the Sciences in Paris, 1670–1760. Isis, 105(3):646–647, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/679150.

Kucuk:2014:BRG

[1354] B. Harun Küçük. Book review: Gülhan Balsoy. The Politics of Reproduction in Ottoman Society, 1838–1900. Isis, 105(3):647–648, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/679151.

McVicker:2014:BRD

[1355] Donald McVicker. Book review: David L. Browman; Stephen Williams. Anthropology at Harvard: A Biographical History, 1790–1940. Isis, 105 (3):648–649, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679152.

Ernst:2014:BRP

[1356] Waltraud Ernst. Book review: Pratik Chakrabarti. Bacteriology in British India: Laboratory Medicine and the Tropics. Isis, 105(3):649–650, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679153.

Gariboldi:2014:BRI

[1357] Leonardo Gariboldi. Book review: Ileana Chinnici; Antonella Gasperini. Alle origini dell’astrofisica italiana: Il carteggio Secchi–Tacchini 1861– 1877. Isis, 105(3):651, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/679154.

Staum:2014:BRA

[1358] Martin S. Staum. Book review: Alice L. Conklin. In the Museum of Man: Race, Anthropology, and Empire in France, 1850–1950. Isis, 105 (3):651–652, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679155.

Seth:2014:BRM

[1359] Suman Seth. Book review: Michael Eckert. Arnold Sommerfeld: Atomphysiker und Kulturbote 1868–1951: Eine Biografie. Isis, 105(3):652–654, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679156.

Brock:2014:BRM

[1360] Claire Brock. Book review: Michael Hoskin. Caroline Herschel: Priestess of the New Heavens. Isis, 105(3):654, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/679157.

Wimmer:2014:BRB

[1361] Mario Wimmer. Book review: Bernhard Kleeberg, ed. Schlechte Angewohnheiten: Eine Anthologie 1750–1900. Isis, 105(3):654–655, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/679158.

Dawson:2014:BRB

[1362] Gowan Dawson. Book review: Ben Marsden; Hazel Hutchison; Ralph O’Connor, eds. Uncommon Contexts: Encounters between Science and Literature, 1800–1914. Isis, 105(3):655–656, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/679159.

Low:2014:BRI

[1363] Morris Low. Book review: Ian Jared Miller; Harriet Ritvo. The Nature of the Beasts: Empire and Exhibition at the Tokyo Imperial Zoo. Isis, 105 (3):656–657, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679160.

Pandora:2014:BRK

[1364] Katherine Pandora. Book review: Katy Price. Loving Faster than Light: Romance and Readers in Einstein’s Universe. Isis, 105(3):657–658, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679161.

Whitmer:2014:BRA

[1365] Kelly J. Whitmer. Book review: Ayako Sakurai. Science and Societies in Frankfurt am Main. Isis, 105(3):658–659, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/679162.

Lienesch:2014:BRA

[1366] Michael Lienesch. Book review: Adam R. Shapiro. Trying Biology: The Scopes Trial, Textbooks, and the Antievolution Movement in American Schools. Isis, 105(3):659–660, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/679163.

Battimelli:2014:BRR

[1367] Giovanni Battimelli. Book review: Raffaella Simili, ed. Scienziati, patrioti, presidenti: L’Accademia Nazionale dei Lincei (1874–1926). Isis, 105 (3):660–661, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679164.

Hutchison:2014:BRD

[1368] Iain Hutchison. Book review: David Stack. Queen Victoria’s Skull: George Combe and the Mid-Victorian Mind. Isis, 105(3):661–662, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679165.

Waddington:2014:BRJ

[1369] Keir Waddington. Book review: James F. Stark. The Making of Modern Anthrax, 1875–1920: Uniting Local, National, and Global Histories of Disease. Isis, 105(3):662–663, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/679166.

Fichman:2014:BRA

[1370] Martin Fichman. Book review: Alfred Russel Wallace; James T. Costa. On the Organic Law of Change: A Facsimile Edition and Annotated Transcription of Alfred Russel Wallace’s Species Notebook of 1855–1859. Isis, 105(3):663–664, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/679167.

Spanagel:2014:BRR

[1371] David Spanagel. Book review: Rosalind Williams. The Triumph of Human Empire: Verne, Morris, and Stevenson at the End of the World. Isis, 105(3):664–665, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/679168.

Wittje:2014:BRC

[1372] Roland Wittje. Book review: Chen-Pang Yeang. Probing the Sky with Radio Waves: From Wireless Technology to the Development of Atmospheric Science. Isis, 105(3):665–666, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/679169.

Jardine:2014:BRP

[1373] Nicholas Jardine. Book review: Paul Ziche; Petr Rezvykh. Sygkepleriazein: Schelling und die Kepler-Rezeption im 19. Jahrhundert. Isis, 105 (3):666–667, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679170.

Kouw:2014:BRP

[1374] Matthijs Kouw. Book review: Paul M. Leonardi. Car Crashes without Cars: Lessons about Simulation Technology and Organizational Change from Automotive Design. Isis, 105(3):668, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/679172.

Theunissen:2014:BRH

[1375] Bert Theunissen. Book review: Huib J. Zuidervaart; Rob H. van Gent. Between Rhetoric and Reality: Astronomical Practices at the Observatory of the Amsterdam Society “Felix Meritis,” 1786–1889. Isis, 105(3):668, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679171.

MacLeod:2014:BRR

[1376] Roy MacLeod. Book review: Rebecca Priestley. Mad on Radium: New Zealand in the Atomic Age. Isis, 105(3):669–670, September 2014. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/679173.

Turner:2014:BRS

[1377] Roger Turner. Book review: Siobhan Roberts. Wind Wizard: Alan G. Davenport and the Art of Wind Engineering. Isis, 105(3):670–671, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679174.

Roland:2014:BRRa

[1378] Alex Roland. Book review: Rebecca Slayton. Arguments That Count: Physics, Computing, and Missile Defense, 1949–2012. Isis, 105(3): 671–672, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679175.

Staley:2014:BRP

[1379] Kent W. Staley. Book review: Peter Achinstein. Evidence and Method: Scientific Strategies of Isaac Newton and James Clerk Maxwell. Isis, 105 (3):672–673, September 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679176.

Schmidgen:2014:BRB

[1380] Henning Schmidgen. Book review: Bruno Latour; Catherine Porter. An Inquiry into Modes of Existence: An Anthropology of the Moderns. Isis, 105(3):673–675, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/679177.

Anonymous:2014:NCc

[1381] Anonymous. Notes on contributors. Isis, 105(3):676–679, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/678179.

Anonymous:2014:Fd

[1382] Anonymous. Frontispiece. Isis, 105(4):vi, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/680334.

Serjeantson:2014:FBI

[1383] Richard Serjeantson. Francis Bacon and the “interpretation of nature” in the Late Renaissance. Isis, 105(4):681–705, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/679419.

Goodrum:2014:CNS

[1384] Matthew R. Goodrum. Crafting a new science: Defining paleoanthropology and its relationship to prehistoric archaeology, 1860–1890. Isis, 105 (4):706–733, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679420.

Rosner:2014:SIA

[1385] Rachael I. Rosner. The “splendid isolation” of Aaron T. Beck. Isis, 105 (4):734–758, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679421.

Seth:2014:I

[1386] Suman Seth. Introduction. Isis, 105(4):759–763, December 2014. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/679422.

Seth:2014:MSH

[1387] Suman Seth. Materialism, slavery, and the history of Jamaica. Isis, 105 (4):764–772, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679423.

Tilley:2014:RSG

[1388] Helen Tilley. Racial science, geopolitics, and empires: Paradoxes of power. Isis, 105(4):773–781, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/679424.

Anderson:2014:RCG

[1389] Warwick Anderson. Racial conceptions in the global south. Isis, 105 (4):782–792, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679425.

Chung:2014:BSB

[1390] Yuehtsen Juliette Chung. Better science and better race?: Social Darwinism and Chinese eugenics. Isis, 105(4):793–802, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/679426.

Fullwiley:2014:CSW

[1391] Duana Fullwiley. The “contemporary synthesis”: When politically inclusive genomic science relies on biological notions of race. Isis, 105 (4):803–814, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
679427.

Anonymous:2014:EHR

[1392] Anonymous. Eloge: Henrika “Riki’ Kuklick, 1942–2013. Isis, 105(4):815– 818, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679428.

Anonymous:2014:ECJ

[1393] Anonymous. Eloge: Christoph J. Scriba, 1929–2013. Isis, 105(4):819– 824, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679429.

Anonymous:2014:EJP

[1394] Anonymous. Eloge: John A. Popplestone, 1928–2013. Isis, 105(4):825– 826, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/679430.

Guicciardini:2014:LE

[1395] Niccolò Guicciardini. Letter to the Editor. Isis, 105(4):827, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/679431.

Fraser:2014:BRM

[1396] Craig Fraser. Book review: Martin Harwit. In Search of the True Universe: The Tools, Shaping, and Cost of Cosmological Thought. Isis, 105 (4):828–829, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
680263.

Konoval:2014:BRJ

[1397] Brandon Konoval. Book review: James Kennaway. Bad Vibrations: The History of the Idea of Music as a Cause of Disease. Isis, 105(4):829–830, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/680264.

Verboon:2014:BRW

[1398] Annemieke Verboon. Book review: Wibke Larink. Bilder vom Gehirn: Bildwissenschaftliche Zugänge zum Gehirn als Seelenorgan. Isis, 105(4):

830–831, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
680265.

Mandelbrote:2014:BRC

[1399] Scott Mandelbrote. Book review: C. Philipp; E. Nothaft. Dating the Passion: The Life of Jesus and the Emergence of Scientific Chronology (200–1600). Isis, 105(4):831–832, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/680266.

Boantza:2014:BRL

[1400] Victor D. Boantza. Book review: Lawrence M. Principe. The Secrets of Alchemy. Isis, 105(4):832–833, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/680267.

Hardy:2014:BRJ

[1401] Anne Hardy. Book review: Jonathan Reinarz; Leonard Schwarz, eds. Medicine and the Workhouse. Isis, 105(4):833–834, December 2014. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/680268.

Bud:2014:BRA

[1402] Robert Bud. Book review: Alexander von Schwerin; Heiko Stoff; Bettina Wahrig, eds. Biologics: A History of Agents Made from Living Organisms in the Twentieth Century. Isis, 105(4):834–835, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/680269.

Milam:2014:BRL

[1403] Erika Lorraine Milam. Book review: Liv Emma Thorsen; Karen A. Rader; Adam Dodd, eds. Animals on Display: The Creaturely in Museums, Zoos, and Natural History. Isis, 105(4):835–836, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/680270.

Wall:2014:BRK

[1404] Byron E. Wall. Book review: Karine Chemla, ed. The History of Mathematical Proof in Ancient Traditions. Isis, 105(4):836–837, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/680271.

Meeusen:2014:BRY

[1405] Michiel Meeusen. Book review: Yves Lehmann, ed. Aristoteles Romanus: La réception de la science aristotélicienne dans l’Empire gréco-romain. Isis, 105(4):837–838, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/680272.

Olshin:2014:BRP

[1406] Benjamin B. Olshin. Book review: Piero Falchetta. Fra’ Mauro’s World Map: A History. Isis, 105(4):839, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/680273.

Lynn:2014:BRG

[1407] Michael R. Lynn. Book review: G. Matthew Adkins. The Idea of the Sciences in the French Enlightenment: A Reinterpretation. Isis, 105 (4):839–840, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
680274.

Dijksterhuis:2014:BRJ

[1408] Fokko Jan Dijksterhuis. Book review: Jim Bennett; Sofia Talas, eds. Cabinets of Experimental Philosophy in Eighteenth-Century Europe. Isis, 105 (4):840–841, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
680275.

vanBerkel:2014:BRH

[1409] Klaas van Berkel. Book review: Harold J. Cook; Sven Dupré, eds. Translating Knowledge in the Early Modern Low Countries. Isis, 105(4):841– 842, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/680276.

Roos:2014:BRM

[1410] Anna Marie Roos. Book review: Mary Floyd-Wilson. Occult Knowledge, Science, and Gender on the Shakespearean Stage. Isis, 105(4):843, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/680277.

Mandelbrote:2014:BRM

[1411] Scott Mandelbrote. Book review: Matthew C. Hunter. Wicked Intelligence: Visual Art and the Science of Experiment in Restoration London. Isis, 105(4):843–845, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/680278.

Teichmann:2014:BRG

[1412] Jürgen Teichmann. Book review: Georg Christoph Lichtenberg. Vorlesungen zur Naturlehre: Notizen und Materialien zur Astronomie und Physischen Geographie. Isis, 105(4):845–846, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/680279.

Carhart:2014:BRA

[1413] Michael C. Carhart. Book review: Avi Lifschitz. Language and Enlightenment: The Berlin Debates of the Eighteenth Century. Isis, 105 (4):846–847, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
680280.

Coleman:2014:BRC

[1414] Deirdre Coleman. Book review: Cristina Malcolmson. Studies of Skin Color in the Early Royal Society: Boyle, Cavendish, Swift. Isis, 105 (4):847–848, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
680281.

Steigerwald:2014:BRJ

[1415] Joan Steigerwald. Book review: Jennifer Mensch. Kant’s Organicism: Epigenesis and the Development of Critical Philosophy. Isis, 105(4):848– 849, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/680282.

Charmantier:2014:BRP

[1416] Isabelle Charmantier. Book review: P. H. Oswald; C. D. Preston, eds. John Ray’s Cambridge Catalogue (1660). Isis, 105(4):849–850, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/680283.

Tessicini:2014:BRF

[1417] Dario Tessicini. Book review: Federico Tognoni, ed. Le opere di Galileo Galilei. Appendice, Volume 1: Iconografia Galileiana. Isis, 105(4):850– 851, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/680284.

Cutler:2014:BRE

[1418] Brock Cutler. Book review: Ellen J. Amster. Medicine and the Saints: Science, Islam, and the Colonial Encounter in Morocco, 1877–1956. Isis, 105(4):851–852, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/680285.

Bellon:2014:BRK

[1419] Richard Bellon. Book review: Katharine Anderson, ed. The Narrative of the Beagle Voyage, 1831–1836. Charles Darwin; Christiane Bernard; Marie-Thérèse Blanchon. Journal de bord [Diary] du voyage du Beagle [1831–1836]. Isis, 105(4):852–853, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/680286.

Lucier:2014:BRC

[1420] Paul Lucier. Book review: Conevery Bolton Valencius. The Lost History of the New Madrid Earthquakes. Isis, 105(4):854–855, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/680287.

Smith:2014:BRA

[1421] Jonathan Smith. Book review: Anne DeWitt. Moral Authority, Men of Science, and the Victorian Novel. Isis, 105(4):855–856, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/680288.

Mitchell:2014:BRH

[1422] B. D. Mitchell. Book review: Helmut Heit; Lisa Heller, eds. Handbuch Nietzsche und die Wissenschaften. Isis, 105(4):856–857, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/680289.

Preston:2014:BRL

[1423] Christopher Preston. Book review: Leonard Jenyns; Richard C. Preece; Tim H. Sparks, eds. Fauna Cantabrigiensis: The Vertebrate and Molluscan Fauna of Cambridgeshire by the Rev. Leonard Jenyns (1800–1893): Transcript and Commentaries. Isis, 105(4):857–858, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/680290.

Alberti:2014:BRC

[1424] Samuel J. M. M. Alberti. Book review: Claire L. Jones. The Medical Trade Catalogue in Britain, 1870–1914. Isis, 105(4):858–859, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/680291.

Banerjee:2014:BRK

[1425] Somaditya Banerjee. Book review: Kris Manjapra. Age of Entanglement: German and Indian Intellectuals across Empire. Isis, 105(4):859–860, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/680292.

Dick:2014:BRO

[1426] Steven J. Dick. Book review: Omar W. Nasim. Observing by Hand: Sketching the Nebulae in the Nineteenth Century. Isis, 105(4):861–862, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/680293.

Fisher:2014:BRD

[1427] Donald Fisher. Book review: David L. Seim. Rockefeller Philanthropy and Modern Social Science. Isis, 105(4):862–863, December 2014. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/680294.

Stachel:2014:BRD

[1428] John Stachel. Book review: A. Douglas Stone. Einstein and the Quantum: The Quest of the Valiant Swabian. Isis, 105(4):863–864, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/680295.

Gingerich:2014:BRS

[1429] Owen Gingerich. Book review: Stephen J. Dick. Discovery and Classification in Astronomy: Controversy and Consensus. Isis, 105(4):864–865, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/680296.

Doern:2014:BRP

[1430] Bruce Doern. Book review: Philip Enros. Environment for Science: A History of Policy for Science in Environment Canada. Isis, 105(4):865– 866, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/680297.

Roland:2014:BRRb

[1431] Alex Roland. Book review: Robert G. Ferguson. NASA’s First A: Aeronautics from 1958 to 2008. Isis, 105(4):866–867, December 2014. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/680298.

Smith:2014:BRM

[1432] Robert W. Smith. Book review: Malinda K. Goodrich; Alice R. Buchalter; Patrick M. Miller, eds. Toward a History of the Space Shuttle: An Annotated Bibliography, Part 2, 1992–2011. Isis, 105(4):867, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/680299.

Royles:2014:BRV

[1433] Dan Royles. Book review: Victoria A. Harden. AIDS at Thirty: A History. Isis, 105(4):867–868, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/680300.

Widmalm:2014:BRT

[1434] Sven Widmalm. Book review: Thomas Kaiserfeld; Tom O’Dell, eds. Legitimizing ESS: Big Science as a Collaboration across Boundaries. Isis, 105(4):868–869, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/680301.

Hopwood:2014:BRR

[1435] Nick Hopwood. Book review: Renato G. Mazzolini; Hans-Jörg Rheinberger, eds. Differing Routes to Stem Cell Research: Germany and Italy. Isis, 105(4):869–870, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/680302.

Bimm:2014:BRM

[1436] Jordan Bimm. Book review: Michael J. Neufeld, ed. Spacefarers: Images of Astronauts and Cosmonauts in the Heroic Era of Spaceflight. Isis, 105 (4):870–871, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
680303.

deChadarevian:2014:BRD

[1437] Soraya de Chadarevian. Book review: Diane B. Paul; Jeffrey P. Brosco. The PKU Paradox: A Short History of a Genetic Disease. Isis, 105(4):

872, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/680304.

Habib:2014:BRJ

[1438] S. Irfan Habib. Book review: Jahnavi Phalkey. Atomic State: Big Science in Twentieth-Century India. Isis, 105(4):872–873, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/680305.

Sepkoski:2014:BRH

[1439] David Sepkoski. Book review: Hallam Stevens. Life Out of Sequence: A Data-Driven History of Bioinformatics. Isis, 105(4):873–874, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/680306.

Maienschein:2014:BRC

[1440] Jane Maienschein. Book review: Charis Thompson. Good Science: The Ethical Choreography of Stem Cell Research. Isis, 105(4):874–875, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/680307.

Brigandt:2014:BRC

[1441] Ingo Brigandt. Book review: Carl F. Craver; Lindley Darden. In Search of Mechanisms: Discoveries across the Life Sciences. Isis, 105(4):875– 876, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/680308.

Kronfeldner:2014:BRH

[1442] Maria Kronfeldner. Book review: Helen E. Longino. Studying Human Behavior: How Scientists Investigate Aggression and Sexuality. Isis, 105 (4):876–877, December 2014. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
680309.

Anonymous:2014:NCd

[1443] Anonymous. Notes on contributors. Isis, 105(4):878–881, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/679433.

Anonymous:2014:IIV

[1444] Anonymous. Index to Isis, volume 105, 2014. Isis, 105(4):882–892, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/679732.

Anonymous:2014:CIV

[1445] Anonymous. Contents of Isis, volume 105. Isis, 105(4):895–897, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/679731.

Anonymous:2014:CBH

[1446] Anonymous. Current bibliography of the history of science and its cultural influences, 2014. Isis, 105(S1):i–301, December 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/680345.

Carman:2015:TEE

[1447] Christián Carlos Carman and James Evans. The two Earths of Eratosthenes. Isis, 106(1):1–16, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/681034.

Keller:2015:ASD

[1448] Vera Keller and Leigh T. I. Penman. From the archives of scientific diplomacy: Science and the shared interests of Samuel Hartlib’s London and Frederick Clodius’s Gottorf. Isis, 106(1):17–42, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/681035.

Jackson:2015:WPG

[1449] Catherine M. Jackson. The “wonderful properties of glass”: Liebig’s Kaliapparat and the practice of chemistry in glass. Isis, 106(1):43–69, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/681036.

Clarke:2015:GMP

[1450] Imogen Clarke. The gatekeepers of modern physics: Periodicals and peer review in 1920s Britain. Isis, 106(1):70–93, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/681037.

Lassman:2015:PMB

[1451] Thomas C. Lassman. Putting the military back into the history of the military–industrial complex: The management of technological innovation in the U.S. Army, 1945–1960. Isis, 106(1):94–120, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/681038.

Pettit:2015:MNV

[1452] Michael Pettit, Darya Serykh, and Christopher D. Green. Multispecies networks: Visualizing the psychological research of the Committee for Research in Problems of Sex. Isis, 106(1):121–149, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/681039.

Anonymous:2015:ERA

[1453] Anonymous. Eloge: Richard Adrian Jarrell (1946–2013). Isis, 106 (1):150–151, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681040.

Anonymous:2015:EJV

[1454] Anonymous. Eloge: John V. Pickstone (1944–2014). Isis, 106(1):152– 155, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/681041.

Navarro:2015:BRJ

[1455] Jaume Navarro. Book review: Jed Z. Buchwald; Robert Fox, eds., The Oxford Handbook of the History of Physics. Isis, 106(1):156–157, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/681824.

Konoval:2015:BRH

[1456] Brandon Konoval. Book review: Howard P. Segal. Utopias: A Brief History from Ancient Writings to Virtual Communities. Isis, 106(1):157– 158, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/681825.

Kumar:2015:BRD

[1457] Deepak Kumar. Book review: Dagmar Wujastyk, Well-Mannered Medicine: Medical Ethics and Etiquette in Classical Ayurveda. Isis, 106 (1):158–159, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681826.

Salazar:2015:BRA

[1458] C. F. Salazar. Book review: Anne Van Arsdall; Timothy Graham, eds., Herbs and Healers from the Ancient Mediterranean through the Medieval West: Essays in Honor of John M. Riddle. Isis, 106(1):159–160, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/681827.

Fraser:2015:BRS

[1459] Craig Fraser. Book review: Suzanne Féry, ed., Aventures de l’analyse de Fermat à Borel: Mélanges en l’honneur de Christian Gilain. Isis, 106(1):160–161, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681828.

Burns:2015:BRO

[1460] Conor Burns. Book review: Ola Wolfhechel Jensen, ed.m Histories of Archaeological Practices: Reflections on Methods, Strategies, and Social Organisation in Past Fieldwork. David L. Browman, Cultural Negotiations: The Role of Women in the Founding of Americanist Archaeology. Isis, 106(1):162–163, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/681829.

Gerali:2015:BRC

[1461] Francesco Gerali. Book review: Curzio Cipriani; Luciana Fantoni; Luisa Poggi; Alba Scarpellini. Le collezioni mineralogiche del Museo di Storia Naturale dell’Universita di Firenze dalle origini a oggi. Isis, 106(1):163– 164, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/681830.

Gauvin:2015:BRP

[1462] Jean-François Gauvin. Book review: Patrizio Barbieri. Physics of Wind Instruments and Organ Pipes, 1100–2010: New and Extended Writings. Isis, 106(1):164–165, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/681831.

Zuccolin:2015:BRM

[1463] Gabriella Zuccolin. Book review: Maria Pia Donato; Luc Berlivet; Sara Cabibbo; Raimondo Michetti; Marilyn Nicoud, eds. Médecine et religion: Compétitions, collaborations, conflits (XII–XX siècles). Isis, 106(1):165– 166, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/681832.

Evans:2015:BRE

[1464] James Evans. Book review: Elly Dekker, Illustrating the Phaenomena: Celestial Cartography in Antiquity and the Middle Ages. Isis, 106(1):166– 167, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/681833.

VanBrummelen:2015:BRD

[1465] Glen Van Brummelen. Book review: Daniel W. Graham. Science before Socrates: Parmenides, Anaxagoras, and the New Astronomy. Isis, 106 (1):167–168, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681834.

Taub:2015:BRE

[1466] Liba Taub. Book review: Emma Gee. Aratus and the Astronomical Tradition. Isis, 106(1):168–169, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/681835.

Leith:2015:BRA

[1467] David Leith. Book review: Andrea Falcon. Aristotelianism in the First Century B.C.E.: Xenarchus of Seleucia. Isis, 106(1):169–170, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/681836.

Valleriani:2015:BRA

[1468] Matteo Valleriani. Book review: Alan C. Bowen. Simplicius on the Planets and Their Motions: In Defense of a Heresy. Isis, 106(1):170–171, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/681837.

Whittington:2015:BRC

[1469] Karl Whittington. Book review: Carole Rawcliffe. Urban Bodies: Communal Health in Late Medieval English Towns and Cities. Isis, 106 (1):171–172, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681838.

Dooley:2015:BRJ

[1470] Brendan Dooley. Book review: Jean-Patrice Boudet; Franck Collard; Nicolas Weill-Parot, eds., Médecine, astrologie et magie entre Moyen Âge et Renaissance: Autour de Pietro d’Abano. Isis, 106(1):172–173, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/681839.

Shank:2015:BRF

[1471] Michael H. Shank. Book review: Fabrizio Bònoli; Giuseppe Bezza; Salvo De Meis; Cinzia Colavita, eds., I pronostici di Domenico Maria da Novara. Isis, 106(1):173–174, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/681840.

Touber:2015:BRB

[1472] Jetze Touber. Book review: Bernd Roling. Physica sacra: Wunder, Naturwissenschaft und historischer Schriftsinn zwischen Mittelalter und Früher Neuzeit. Isis, 106(1):174–175, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/681841.

Saliba:2015:BRA

[1473] George Saliba. Book review: Arun Bala, ed., Asia, Europe, and the Emergence of Modern Science: Knowledge Crossing Boundaries. Isis, 106(1):175–176, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681842.

Dupre:2015:BRT

[1474] Sven Dupré. Book review: Timothy McCall; Sean Roberts; Giancarlo Fiorenza, eds., Visual Cultures of Secrecy in Early Modern Europe. Isis, 106(1):176–177, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681843.

Stoff:2015:BRS

[1475] Heiko Stoff. Book review: Sabine Anagnostou. Missionspharmazie: Konzepte, Praxis, Organisation und wissenschaftliche Ausstrahlung. Isis, 106(1):177–178, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681844.

Walton:2015:BRM

[1476] Steven A. Walton. Book review: Matteo Valleriani; Matteo Valleriani; Lindy Divarci; Anna Siebold. Metallurgy, Ballistics, and Epistemic Instruments: The Nova scientia of Nicolò Tartaglia: A New Edition. Isis, 106(1):178–179, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681845.

Jacquart:2015:BRE

[1477] Danielle Jacquart. Book review: Elisa Andretta. Roma medica: Anatomie d’un système médical au XVIe siècle. Isis, 106(1):179–180, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/681846.

Henry:2015:BRG

[1478] John Henry. Book review: Gareth Roberts; Fenny Smith, eds., Robert Recorde: The Life and Times of a Tudor Mathematician. Isis, 106 (1):180–181, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681847.

VandenBroecke:2015:BRMa

[1479] Steven Vanden Broecke. Book review: Miguel Ángel Granada, ed., Novas y cometas entre 1572 y 1618: Revolución cosmológica y renovación política y religiosa. Isis, 106(1):181–182, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/681848.

Hall:2015:BRM

[1480] Crystal Hall. Book review: Maurice A. Finocchiaro. The Routledge Guidebook to Galileo’s Dialogue. Isis, 106(1):182–183, March 2015. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/681849.

Bucciantini:2015:BRC

[1481] Massimo Bucciantini. Book review: Chantal Grell; Sylvie Taussig, eds., La lune aux XVIIe et XVIIIe siècles. Isis, 106(1):183–184, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/681850.

Jenner:2015:BRP

[1482] Mark Jenner. Book review: Peter Elmer. The Miraculous Conformist: Valentine Greatrakes, the Body Politic, and the Politics of Healing in Restoration Britain. Isis, 106(1):184–185, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/681851.

Cook:2015:BRD

[1483] Harold J. Cook. Book review: Daniela Prögler. English Students at Leiden University, 1575–1650: “Advancing Your Abilities in Learning and Bettering Your Understanding of the World and State Affairs”. Isis, 106(1):185–186, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681852.

Nutton:2015:BRR

[1484] Vivian Nutton. Book review: Reid Barbour. Sir Thomas Browne: A Life. Isis, 106(1):186–187, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/681853.

Farber:2015:BRO

[1485] Paul Lawrence Farber. Book review: Oeuvres complètes. Volume 5: Histoire naturelle, générale et particulière, avec la description du Cabinet du Roi. Oeuvres complètes. Volume 7: Histoire naturelle, générale et particulière, avec la description du Cabinet du Roi by Georges-Louis Leclerc de Buffon; Stéphane Schmitt; Cédric Crémière; Georges-Louis Leclerc de Buffon; Stéphane Schmitt; Cédric Crémière. Isis, 106(1):187–188, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/681854.

Alexander:2015:BRA

[1486] Jennifer Karns Alexander. Book review: Adelheid Voskuhl. Androids in the Enlightenment: Mechanics, Artisans, and Cultures of the Self. Isis, 106(1):188–189, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681855.

Evans:2015:BRM

[1487] James Evans. Book review: Michael Hoskin; David Dewhirst; Wolfgang Steinicke. The Construction of the Heavens: William Herschel’s Cosmology. Isis, 106(1):189–190, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/681856.

Madison:2015:BRC

[1488] Mark Madison. Book review: Christopher P. Iannini. Fatal Revolutions: Natural History, West Indian Slavery, and the Routes of American Literature. Isis, 106(1):190–191, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/681857.

Whitmer:2015:BRA

[1489] Kelly J. Whitmer. Book review: Anna Echterhölter. Schattengefechte: Genealogische Praktiken in Nachrufen auf Naturwissenschaftler (1710– 1860). Isis, 106(1):191–192, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/681858.

Pancaldi:2015:BRA

[1490] Giuliano Pancaldi. Book review: Angela Bandinelli. Le origini chimiche della vita: Legami tra la rivoluzione di Lavoisier e la biologia di Lamarck. Isis, 106(1):192–193, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/681859.

Saavedra:2015:BRP

[1491] Monica Saavedra. Book review: Poonam Bala, ed. Contesting Colonial Authority: Medicine and Indigenous Responses in Nineteenth- and Twentieth-Century India. Isis, 106(1):193–194, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/681860.

Buckland:2015:BRR

[1492] Adelene Buckland. Book review: Roger S. Wotton, Walking with Gosse: Natural History, Creation, and Religious Conflicts. Isis, 106(1):194–195, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/681861.

Smith:2015:BRC

[1493] Elise Juzda Smith. Book review: Claude Blanckaert, ed. La Vénus hottentote: Entre Barnum et Muséum. Isis, 106(1):195–196, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/681862.

Tresch:2015:BRR

[1494] John Tresch. Book review: Robert Mitchell. Experimental Life: Vitalism in Romantic Science and Literature. Isis, 106(1):196–197, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/681863.

Gomez:2015:BRA

[1495] Leila Gómez. Book review: Alex Levine; Adriana Novoa. ‘!Darwinistas! The Construction of Evolutionary Thought in Nineteenth Century Argentina. Isis, 106(1):197–198, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/681864.

Imber:2015:BRO

[1496] Jonathan B. Imber. Book review: Owen Whooley. Knowledge in the Time of Cholera: The Struggle over American Medicine in the Nineteenth Century. Isis, 106(1):198–199, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/681865.

James:2015:BRG

[1497] Frank A. J. L. James. Book review: Geoffrey Cantor, ed. The Great Exhibition: A Documentary History. Isis, 106(1):199–200, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/681866.

Geisthovel:2015:BRF

[1498] Alexa Geisthövel. Book review: Florian Mildenberger. Medizinische Belehrung für das Bürgertum: Medikale Kulturen in der Zeitschrift “Die Gartenlaube” (1853–1944). Isis, 106(1):200–201, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/681867.

Bigg:2015:BRJ

[1499] Charlotte Bigg. Book review: Jean-Claude Pont, ed. Le destin douloureux de Walther Ritz (1878–1909): Physicien théoricien de génie. Isis, 106 (1):201–202, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681868.

Nye:2015:BRH

[1500] Mary Jo Nye. Book review: Hermann Kopp; Alan J. Rocke. From the Molecular World: A Nineteenth-Century Science Fantasy. Isis, 106 (1):202–203, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681869.

England:2015:BRD

[1501] Richard England. Book review: Don O’Leary. Irish Catholicism and Science: From “Godless Colleges” to the “Celtic Tiger”. Isis, 106(1):203– 204, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/681870.

Arabatzis:2015:BRJ

[1502] Theodore Arabatzis. Book review: Jaume Navarro. A History of the Electron: J. J. and G. P. Thomson. Isis, 106(1):204–205, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/681871.

Tyrrell:2015:BRR

[1503] Ian Tyrrell. Book review: Robert B. Townsend, History’s Babel: Scholarship, Professionalization, and the Historical Enterprise in the United States, 1880–1940. Isis, 106(1):205–206, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/10.1086/681872.

Krebs:2015:BRS

[1504] Stefan Krebs. Book review: Susan Schmidt Horning. Chasing Sound: Technology, Culture, and the Art of Studio Recording from Edison to the LP. Isis, 106(1):206–207, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/681873.

Foster:2015:BRA

[1505] Amy E. Foster. Book review: Amy Sue Bix. Girls Coming to Tech! A History of American Engineering Education for Women. Isis, 106 (1):207–208, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681874.

Maerker:2015:BRN

[1506] Anna Maerker. Book review: Nancy Anderson; Michael Dietrich, eds. The Educated Eye: Visual Culture and Pedagogy in the Life Sciences. Isis, 106(1):208–209, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/681875.

Badino:2015:BRA

[1507] Massimiliano Badino. Book review: Albert Einstein; Diana Kormos Buchwald; Ze’ev Rosenkranz; Tilman Sauer; József Illy; Virginia Iris Holmes, eds. The Collected Papers of Albert Einstein. Volume 12: The Berlin Years: Correspondence, January December 1921. Albert Einstein; Diana Kormos Buchwald; József Illy; Ze’ev Rosenkranz; Tilman Sauer, eds. The Collected Papers of Albert Einstein. Volume 13: The Berlin Years: Writings and Correspondence, January 1922 March 1923. Isis, 106(1):209–211, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681876.

Staley:2015:BRD

[1508] Richard Staley. Book review: Dieter Hoffmann. Einstein’s Berlin: In the Footsteps of a Genius. Isis, 106(1):211–212, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/681877.

Kaji:2015:BRE

[1509] Masanori Kaji. Book review: Eric Scerri. A Tale of Seven Elements. Isis, 106(1):212–213, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681878.

Sethi:2015:BRS

[1510] Megan K. Sethi. Book review: Sevan G. Terzian, Science Education and Citizenship: Fairs, Clubs, and Talent Searches for American Youth, 1918–1958. Isis, 106(1):213–214, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/681879.

Bracco:2015:BRS

[1511] Christian Bracco. Book review: Shaul Katzir; Christoph Lehner; Jürgen Renn, eds. Traditions and Transformations in the History of Quantum Physics. Isis, 106(1):214–215, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/681880.

Lessl:2015:BRL

[1512] Thomas Lessl. Book review: Leah Ceccarelli. On the Frontier of Science: An American Rhetoric of Exploration and Exploitation. Isis, 106(1):215– 216, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/681881.

Kohlstedt:2015:BRR

[1513] Sally Gregory Kohlstedt. Book review: Robert C. Post, Who Owns America’s Past? The Smithsonian and the Problem of History. Isis, 106(1):216–217, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681882.

Ankeny:2015:BRA

[1514] Rachel A. Ankeny. Book review: Alexandra Minna Stern. Telling Genes: The Story of Genetic Counseling in America. Isis, 106(1):217–218, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/681883.

Agar:2015:BRH

[1515] Jon Agar. Book review: Harry Collins. Gravity’s Ghost and Big Dog: Scientific Discovery and Social Analysis in the Twenty-first Century. Isis, 106(1):218–219, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681884.

Cowan:2015:BRM

[1516] Ruth Schwartz Cowan. Book review: Matthew Wisnioski. Engineers for Change: Competing Visions of Technology in 1960s America. Isis, 106(1):219–220, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681885.

Lehner:2015:BRH

[1517] Christoph Lehner. Book review: Hugh Everett; Jeffrey A. Barrett; Peter Byrne, eds., The Everett Interpretation of Quantum Mechanics: Collected Works, 1955–1980, with Commentary. Isis, 106(1):220–221, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/681886.

Kroker:2015:BRA

[1518] Kenton Kroker. Book review: Angela N. H. Creager, Life Atomic: A History of Radioisotopes in Science and Medicine. Isis, 106(1):221–222, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/681887.

Princehouse:2015:BRD

[1519] Patricia Princehouse. Book review: David Sepkoski. Rereading the Fossil Record: The Growth of Paleobiology as an Evolutionary Discipline. Isis, 106(1):222–223, March 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681888.

Anonymous:2015:NCa

[1520] Anonymous. Notes on contributors. Isis, 106(1):224–228, March 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/681042.

Cohen:2015:VT

[1521] H. Floris Cohen. The vision thing. Isis, 106(2):229–234, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/681976.

Silver:2015:PSB

[1522] Sean Silver. The prehistory of serendipity, from Bacon to Walpole. Isis, 106(2):235–256, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681977.

Achbari:2015:BNS

[1523] Azadeh Achbari. Building networks for science: Conflict and cooperation in nineteenth-century global marine studies. Isis, 106(2):257–282, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/682020.

Eghigian:2015:DCU

[1524] Greg Eghigian. A drifting concept for an unruly menace: A history of psychopathy in Germany. Isis, 106(2):283–309, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/681994.

Way:2015:IIV

[1525] Albert G. Way. The invisible and indeterminable value of ecology: From malaria control to ecological research in the American South. Isis, 106(2): 310–336, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/681978.

Bod:2015:IHS

[1526] Rens Bod and Julia Kursell. Introduction: The humanities and the sciences. Isis, 106(2):337–340, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/681993.

Bouterse:2015:DDT

[1527] Jeroen Bouterse and Bart Karstens. A diversity of divisions: Tracing the history of the demarcation between the sciences and the humanities. Isis, 106(2):341–352, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681995.

Kursell:2015:TNH

[1528] Julia Kursell. A third note: Helmholtz, Palestrina, and the early history of musicology. Isis, 106(2):353–366, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682003.

Bod:2015:CFS

[1529] Rens Bod. A comparative framework for studying the histories of the humanities and science. Isis, 106(2):367–377, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/681979.

Daston:2015:HSH

[1530] Lorraine Daston and Glenn W. Most. History of science and history of philologies. Isis, 106(2):378–390, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/681980.

Anonymous:2015:EDC

[1531] Anonymous. Eloge: David C. Lindberg 1935–2015. Isis, 106(2):391– 393, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/681981.

Anonymous:2015:HSS

[1532] Anonymous. History of Science Society annual meeting, 2014. Isis, 106(2):394–400, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
681982.

Brain:2015:ERT

[1533] Robert Brain. Essay review: Theater of machines: The Romantic Machine: Utopian Science and Technology after Napoleon. Isis, 106(2):401– 405, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682064.

Israel:2015:ERI

[1534] Paul Israel. Essay review: Ideal, illusion, and invention: W. Bernard Carlson, Tesla: Inventor of the Electrical Age, xiii + 500 pp., illus., index. Princeton, N.J./Oxford: Princeton University Press, 2013. $29.95 (cloth). Isis, 106(2):406–409, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682065.

Duffin:2015:BRG

[1535] Jacalyn Duffin. Book review: Gary B. Ferngren, Medicine and Religion: A Historical Introduction, xii + 241 pp., index. Baltimore: Johns Hopkins University Press, 2014. $24.95 (paper). Isis, 106(2):410–411, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/682751.

Romano:2015:BRM

[1536] Antonella Romano. Book review: Mark Bradley, ed. Rome, Pollution, and Propriety: Dirt, Disease, and Hygiene in the Eternal City from Antiquity to Modernity, With Kenneth Stow. (British School at Rome Studies.) xx + 320 pp., illus., bibl., index. Cambridge: Cambridge University Press, 2012. $99 (cloth). Isis, 106(2):411–412, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682752.

Gouk:2015:BRP

[1537] Penelope Gouk. Book review: Peter Pesic. Music and the Making of Modern Science, viii + 347 pp., illus., bibl., index. Cambridge, Mass.: MIT Press, 2014. $40 (cloth). Isis, 106(2):412–413, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682753.

VandenBroecke:2015:BRMb

[1538] Steven Vanden Broecke. Book review: Michel Blay, Dieu, la nature et l’homme: L’originalité de l’Occident. (Le Temps des Idées.) 360 pp., index. Paris: Armand Colin, 2013. EUR 26 (paper). Isis, 106(2):413– 415, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682754.

Low:2015:BRY

[1539] Morris Low. Book review: Yung Sik Kim, Questioning Science in East Asian Contexts: Essays on Science, Confucianism, and the Comparative History of Science. (Science and Religion in East Asia, 1.) vii + 284 pp., bibl., index. Leiden: Brill, 2014. EUR 103 (cloth). Isis, 106(2):415– 416, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682755.

Thomas:2015:BRJ

[1540] Marion Thomas and Jean-Yves Bart. Book review: Jean-Marc Drouin, Philosophie de l’insecte. (Science Ouverte.) 249 pp., illus., bibl., index. Paris: Éditions du Seuil, 2014. EUR 19.50 (paper). Isis, 106(2):416– 417, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682756.

Archibald:2015:BRD

[1541] Tom Archibald. Book review: Dominique Flament; Philippe Nabonnand, eds. Justifier en mathématiques. 371 pp., illus., bibls. Paris: Éditions de la Maison des Sciences de l’Homme, 2011. EUR 29 (paper). Isis, 106(2):

417–419, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682757.

Sarkar:2015:BRP

[1542] Sahotra Sarkar. Book review: Paul Griffiths; Karola Stotz, eds. Genetics and Philosophy: An Introduction. viii + 270 pp., illus., bibl., index. Cambridge: Cambridge University Press, 2013. $30.95 (paper). Isis, 106(2): 419–420, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682758.

Carusi:2015:BRA

[1543] Annamaria Carusi. Book review: Alan G. Gross; Joseph E. Harmon. Science from Sight to Insight: How Scientists Illustrate Meaning, vii + 332 pp., illus., bibl., index. Chicago/London: University of Chicago Press, 2013. $30 (paper). Isis, 106(2):420–421, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682759.

Wilder:2015:BRS

[1544] Kelley Wilder. Book review: Shawn Michelle Smith, At the Edge of Sight: Photography and the Unseen, xvii + 293 pp., illus., bibl., index. Durham, N.C./London: Duke University Press, 2013. $28.95 (paper). Isis, 106(2): 421–422, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682760.

Lehmann:2015:BRP

[1545] Ann-Sophie Lehmann. Book review: Pamela H. Smith; Amy R. W. Meyers; Harold J. Cook, eds. Ways of Making and Knowing: The Material Culture of Empirical Knowledge. (Bard Graduate Center Cultural Histories of the Material World.) xi + 430 pp., illus., table, bibl., index. Ann Arbor: University of Michigan Press, 2014. $60 (cloth). Isis, 106(2):423– 424, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682761.

Mosshammer:2015:BRS

[1546] Alden Mosshammer. Book review: Sacha Stern; Charles Burnett, eds. Time, Astronomy, and Calendars in the Jewish Tradition, (Time, Astronomy, and Calendars, 3.) xxi + 365 pp., illus., tables, bibl., index. Leiden/Boston: Brill, 2014. $220 (cloth). Isis, 106(2):424–425, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/682762.

Grattan-Guinness:2015:BRJ

[1547] I. Grattan-Guinness. Book review: Joseph Mazur, Enlightening Symbols: A Short History of Mathematical Notation and Its Hidden Powers, xxiii + 285 pp., illus., tables, apps., bibls., index. Princeton, N.J./Oxford: Princeton University Press, 2014. $29.95 (cloth). Isis, 106(2):425–426, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/682763.

Green:2015:BRJ

[1548] Christopher D. Green. Book review: Jerome McGann, A New Republic of Letters: Memory and Scholarship in the Age of Digital Reproduction, + 238 pp., illus., index. Cambridge, Mass./London: Harvard University Press, 2014. $39.95 (cloth). Isis, 106(2):426–427, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682764.

Savage-Smith:2015:BRH

[1549] Emilie Savage-Smith. Book review: Housni Alkhateeb Shehada, Mamluks and Animals: Veterinary Medicine in Medieval Islam, (Sir Henry Wellcome Asian Series, 11.) xxii + 537 pp., illus., bibl., index. Leiden: Brill, 2013. $245, EUR 176 (cloth). Isis, 106(2):428–429, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/682765.

Page:2015:BRM

[1550] Sophie Page. Book review: Michael D. Bailey, Fearful Spirits, Reasoned Follies: The Boundaries of Superstition in Late Medieval Europe. Isis, 106(2):429–430, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
682766.

Robert:2015:BRN

[1551] Aurélien Robert. Book review: Nicolas Weill-Parot, Points aveugles de la nature: La rationalité scientifique médiévale face à l’occulte, l’attraction magnétique et l’horreur du vide (XIIIe-milieu du XVe siècle), (Histoire, 120.) 652 pp., bibl., index. Paris: Les Belles Lettres, 2013. EUR 55 (paper). Isis, 106(2):430–432, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/682767.

Leibenguth:2015:BRA

[1552] Erik Leibenguth. Book review: Anke Timmermann, Verse and Transmutation: A Corpus of Middle English Alchemical Poetry (Critical Editions and Studies), (History of Science and Medicine Library, 42; Medieval and Early Modern Science, 21.) xiv + 374 pp., bibl., index. Leiden: Brill, 2013. $171 (cloth). . Isis, 106(2):432–433, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682768.

Weill-Parot:2015:BRV

[1553] Nicolas Weill-Parot. Book review: Violaine Giacomotto-Charra; Christine Silvi, eds. Lire, choisir, écrire: La vulgarisation des savoirs du Moyen Âge à la Renaissance, (Etudes et Rencontres de l’Ecole des Chartes, 43.) 276 pp., illus., index. Paris: École Nationale des Chartes, 2014. EUR 24 (paper). Isis, 106(2):433–434, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/682769.

Bilak:2015:BRM

[1554] Donna Bilak. Book review: Marco Beretta; Maria Conforti, eds. Fakes!? Hoaxes, Counterfeits, and Deceptions in Early Modern Science, xv + 280 pp., illus., figs., tables, apps., index. Sagamore Beach, Mass.: Science History Publications, 2014. $47.96 (paper). Isis, 106(2):434–435, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/682770.

Marcus:2015:BRT

[1555] Hannah Marcus. Book review: Thomas F. Mayer, The Roman Inquisition on the Stage of Italy, c. 1590–1640, (Haney Foundation Series.) 392 pp., bibl., index. Philadelphia: University of Pennsylvania Press, 2014. $79.95 (cloth). Isis, 106(2):436–437, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682771.

Henderson:2015:BRD

[1556] Felicity Henderson. Book review: Dirk van Miert, ed. Communicating Observations in Early Modern Letters (1500–1675): Epistolography and Epistemology in the Age of the Scientific Revolution, (Warburg Institute Colloquia, 23.) ix + 268 pp., illus., index. London: Warburg Institute/ Turin: Nino Aragno Editore, 2013. EUR 50 (paper). Isis, 106(2):437– 438, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682772.

Ragland:2015:BRO

[1557] Evan Ragland. Book review: Ohad Nachtomy; Justin E. H. Smith, eds. The Life Sciences in Early Modern Philosophy, xiii + 256 pp., illus., indexes. Oxford: Oxford University Press, 2014. $74 (cloth). Isis, 106(2): 438–439, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682773.

Kim:2015:BRV

[1558] Mi Gyung Kim. Book review: Victor D. Boantza, Matter and Method in the Long Chemical Revolution: Laws of Another Order, (Science, Technology, and Culture, 1700–1945). xiv + 266 pp., illus., bibl., index. Farnham, Surrey: Ashgate, 2013. $124.95 (cloth). Isis, 106(2):439–440, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/682774.

Rabin:2015:BRK

[1559] Sheila J. Rabin. Book review: Ken Kurihara, Celestial Wonders in Reformation Germany, (Religious Cultures in the Early Modern World, 13). xi + 210 pp., bibl., index. London: Pickering & Chatto, 2014. $99 (cloth). Isis, 106(2):441, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
682775.

Vermij:2015:BRD

[1560] Rienk H. Vermij. Book review: Dario Tessicini; Patrick J. Boner, eds. Celestial Novelties on the Eve of the Scientific Revolution, 1540–1630, (Biblioteca di Galilaeana, 3.) xvi + 282 pp., illus., index. Florence: Leo S. Olschki, 2013. EUR 32 (cloth). Isis, 106(2):441–442, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682776.

Pesic:2015:BRW

[1561] Peter Pesic. Book review: Walter Bühler,Musikalische Skalen bei Naturwissenschaftlern der frühen Neuzeit: Eine elementarmathematische Analyse, 263 pp., illus., bibl., index. Frankfurt am Main: Peter Lang Academic Research, 2013. $56.54 (cloth). Isis, 106(2):442–444, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/682777.

Blume:2015:BRU

[1562] Dieter Blume and H. Floris Cohen. Book review: Ulrike Feist, Sonne, Mond und Venus: Visualisierungen astronomischen Wissens im frühneuzeitlichen Rom, (Actus et Imago, 10.) viii + 259 pp., illus., bibl. Berlin: Akademie Verlag, 2013. EUR 74.58 (cloth). Isis, 106(2):444, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/682778.

Hatch:2015:BRR

[1563] Robert Alan Hatch. Book review: Richard L. Kremer; Jarosław Włodarczyk, eds. Johannes Hevelius and His World: Astronomer, Cartographer, Philosopher, and Correspondent, (Studia Copernicana, 44.) viii + 235 pp., illus., tables, bibls. Warsaw: Polish Academy of Sciences; Institute for the History of Science, Copernicus Center for Interdisciplinary Studies, 2013. EUR 42 (cloth). Isis, 106(2):445–446, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/682779.

VanDyck:2015:BRJ

[1564] Maarten Van Dyck. Book review: Joella G. Yoder, ed. A Catalogue of the Manuscripts of Christiaan Huygens, Including a Concordance with His Œuvres Complètes, (History of Science and Medicine Library, 35.) xii + 338 pp., illus., bibl. Leiden/Boston: Brill, 2013. $216 (cloth). Isis, 106(2):446–447, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
682780.

Hoquet:2015:BRM

[1565] Thierry Hoquet. Book review: Mary Terrall, Catching Nature in the Act: Réaumur and the Practice of Natural History in the Eighteenth Century, 264 pp., illus., bibl., index. Chicago/London: University of Chicago Press, 2013. $40 (cloth). Isis, 106(2):448–449, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/682781.

Davids:2015:BRT

[1566] Karel Davids. Book review: Ted Binnema, Enlightened Zeal: The Hudson’s Bay Company and Scientific Networks, 1670–1870, xvi + 458 pp., illus., bibl., index. Toronto: University of Toronto Press, 2014. $37.95 (paper). Isis, 106(2):449–450, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682782.

Neufeld:2015:BRP

[1567] Matthew Neufeld. Book review: Paul Kopperman, ed. “Regimental Practice,” by John Buchanan, M.D., (History of Medicine in Context.) xiv + 231 pp., bibl., index. Farnham, Surrey: Ashgate, 2014. £70 (cloth). Isis, 106(2):450–451, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
682783.

Bocking:2015:BRF

[1568] Stephen Bocking. Book review: Fredrik Albritton Jonsson, Enlightenment’s Frontier: The Scottish Highlands and the Origins of Environmentalism, (Lewis Walpole Series in Eighteenth-Century Culture and History.) ix + 344 pp., illus., bibl., index. New Haven, Conn./London: Yale University Press, 2013. $50 (cloth). Isis, 106(2):452–453, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/682784.

Bennett:2015:BRB

[1569] Jim Bennett. Book review: Brain Gee, Francis Watkins and the Dollond Telescope Patent Controversy, Edited by Anita McConnell and A. D. Morrison-Low. (Science, Technology, and Culture, 1700–1945.) xxvi + 392 pp., illus., figs., tables, apps., bibl., index. Farnham, Surrey: Ashgate, 2014. £85 (cloth). Isis, 106(2):453–454, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682785.

Thrush:2015:BRJ

[1570] Coll Thrush. Book review: John Gascoigne, Encountering the Pacific in the Age of the Enlightenment, xxxv + 542 pp., illus., bibl., index. Cambridge: Cambridge University Press, 2014. $120 (cloth). Isis, 106(2): 454–455, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682786.

Roberts:2015:BRM

[1571] Lissa Roberts. Book review: Margaret C. Jacob, The First Knowledge Economy: Human Capital and the European Economy, 1750–1850, ix + 257 pp., illus., bibl., index. Cambridge: Cambridge University Press, 2014. £19.99 (paper). Isis, 106(2):456–457, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682787.

Cook:2015:BRS

[1572] Alexandra Cook. Book review: Sue Ann Prince, ed. Of Elephants and Roses: French Natural History, 1790–1830, (Memoirs of the American Philosophical Society, 267.) xxvi + 268 pp., illus., index. Philadelphia: American Philosophical Society, 2013. $50 (paper). Isis, 106(2):457– 458, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682788.

Naylor:2015:BRD

[1573] Simon Naylor. Book review: David Lambert, Mastering the Niger: James MacQueen’s African Geography and the Struggle over Atlantic Slavery, viii + 309 pp., illus., bibl., index. Chicago/London: University of Chicago Press, 2013. $50 (cloth). Isis, 106(2):458–459, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/682789.

Weber:2015:BRN

[1574] Andreas Weber. Book review: Nils Güttler, Das Kosmoskop: Karten und ihre Benutzer in der Pflanzengeographie des 19. Jahrhunderts, 545 pp., illus., bibl., index. Göttingen: Wallstein Verlag, 2014. EUR 64.90 (cloth). Isis, 106(2):459–460, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/682790.

Chang:2015:BRP

[1575] Ku ming (Kevin) Chang. Book review: Pascale Rabault-Feuerhahn, Archives of Origins: Sanskrit, Philology, Anthropology in Nineteenth Century Germany, Translated by Dominique Bach and Richard Willet. (Studies in Cultural and Social Sciences, 9.) 336 pp., app., bibl., index. Wiesbaden: Harrassowitz Verlag, 2013. EUR 68 (cloth). Isis, 106(2): 461–462, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682791.

James:2015:BRN

[1576] Frank A. J. L. James. Book review: Nancy Forbes and Basil Mahon, Faraday, Maxwell, and the Electromagnetic Field: How Two Men Revolutionized Physics, 320 pp., illus., bibl., index. Amherst, N.Y.: Prometheus Books, 2014. $25.95 (cloth). Isis, 106(2):462–463, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682792.

Cantor:2015:BRB

[1577] Geoffrey Cantor. Book review: Bernard Lightman and Michael S. Reidy, eds. The Age of Scientific Naturalism: Tyndall and His Contemporaries, (Science and Culture in the Nineteenth Century, 24.) xv + 256 pp., illus., index. London: Pickering & Chatto, 2014. £60 (cloth). Isis, 106(2):463– 464, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682793.

Canadelli:2015:BRE

[1578] Elena Canadelli. Book review: Elena Zanoni, Scienza, patria, religione: Antonio Stoppani e la cultura italiana dell’Ottocento, (Storia, 431.) 301 pp., illus., bibl., index. Milan: Franco Angeli, 2014. EUR 39 (paper). Isis, 106(2):464–465, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
682794.

Dauben:2015:BRS

[1579] Joseph W. Dauben. Book review: Steve Nadis and Shing-Tung Yau, A History in Sum: 150 Years of Mathematics at Harvard (1825–1975), xviii + 249 pp., illus., index. Cambridge, Mass./London: Harvard University Press, 2013. $39.95 (cloth). Isis, 106(2):466–467, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682795.

Echterholter:2015:BRG

[1580] Anna Echterhölter. Book review: Gabriel Finkelstein, Emil du Bois-Reymond: Neuroscience, Self, and Society in Nineteenth-Century Germany, (Transformations: Studies in the History of Science and Technology.) 362 pp., illus., bibl., index. Cambridge, Mass./London: MIT Press, 2013. $38 (cloth). Isis, 106(2):467–468, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682796.

Jankovic:2015:BRP

[1581] Vladimir Janković. Book review: Peter Reed, Acid Rain and the Rise of the Environmental Chemist in Nineteenth-Century Britain: The Life and Work of Robert Angus Smith, (Science, Technology, and Culture, 1700– 1945) xv + 209 pp., illus., bibls., index. Farnham, Surrey: Ashgate, 2014. £70 (cloth). Isis, 106(2):468–469, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682797.

Wilder:2015:BRJ

[1582] Kelley Wilder. Book review: Josh Ellenbogen, Reasoned and Unreasoned Images: The Photography of Bertillon, Galton, and Marey, xi + 265 pp., illus., bibl., index. University Park: Pennsylvania State University Press, 2012. $74.95 (cloth). Isis, 106(2):469–471, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682798.

Guagnini:2015:BRS

[1583] Anna Guagnini. Book review: Stathis Arapostathis and Graeme Gooday, Patently Contestable: Electrical Technologies and Inventor Identities on Trial in Britain, (Inside Technology.) xv + 294 pp., illus., bibl., index. Cambridge, Mass./London: MIT Press, 2013. $40 (cloth). Isis, 106(2): 471–472, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682799.

Schlunder:2015:BRL

[1584] Martina Schlünder. Book review: Lynne Fallwell, Modern German Midwifery, 1885–1960, xiii + 263 pp., bibl., index. London: Pickering & Chatto, 2013. $99 (cloth). Isis, 106(2):472–474, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682800.

Adams:2015:BRA

[1585] Mark B. Adams. Book review: Anindita Banerjee, We Modern People: Science Fiction and the Making of Russian Modernity, (Wesleyan Early Classics of Science Fiction.) viii + 206 pp., illus., index. Middletown, Conn.: Wesleyan University Press, 2012. $24.95 (paper). Isis, 106(2): 474–475, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682801.

Bunn:2015:BRM

[1586] Geoff Bunn. Book review: Michael Pettit, The Science of Deception: Psychology and Commerce in America, ix + 311 pp., bibl., index. Chicago/London: University of Chicago Press, 2013. $50 (cloth). Isis, 106(2):475–477, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
682802.

Schoepflin:2015:BRP

[1587] Rennie B. Schoepflin. Book review: Paul Eli Ivey, Radiance from Halcyon: A Utopian Experiment in Religion and Science, xiii + 312 pp., illus., index. Minneapolis/London: University of Minnesota Press, 2013. $25 (paper). Isis, 106(2):477–478, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682803.

Allen:2015:BRE

[1588] Garland E. Allen. Book review: Erika Dyck, Facing Eugenics: Reproduction, Sterilization, and the Politics of Choice, xi + 304 pp., illus., bibl., index. Toronto: University of Toronto Press, 2013. $29.95 (paper). Isis, 106(2):478–479, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
682804.

Stoffers:2015:BRT

[1589] Manuel Stoffers. Book review: Tony Hadland and Hans-Erhard Lessing, Bicycle Design: An Illustrated History, xiii + 564 pp., illus., tables, apps., bibls., index. Cambridge, Mass./London: MIT Press, 2014. $34.95 (cloth). Isis, 106(2):480–481, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682805.

Archibald:2015:BRW

[1590] Tom Archibald. Book review: Wilfried Sieg. Hilbert’s Programs and Beyond. xii + 440 pp., illus., bibl., index. Oxford: Oxford University Press, 2013. $85 (cloth).; William Ewald; Wilfried Sieg (Editors). Michael Hallett (Associate Editor). David Hilbert’s Lectures on the Foundations of Arithmetic and Logic, 1917–1933. (David Hilbert’s Lectures on the Foundations of Mathematics and Physics, 1891–1933, 3.) xxv + 1,062 pp., tables, bibl., indexes. Berlin: Springer, 2013. $139 (cloth). Isis, 106(2):481–483, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
682806.

Heymann:2015:BRR

[1591] Matthias Heymann. Book review: Rebecka Lettevall, Geert Somsen, and Sven Widmalm, eds. Neutrality in Twentieth-Century Europe: Intersections of Science, Culture, and Politics after the First World War, (Routledge Studies in Cultural History.) ix + 351 pp., illus., index. New York/London: Routledge, 2012. $130 (cloth). Isis, 106(2):483–485, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/682807.

Rossiter:2015:BRR

[1592] Margaret W. Rossiter. Book review: Renate Tobies; Annette B. Vogt, eds. Women in Industrial Research, xv + 258 pp., illus., bibl., index. Stuttgart: Franz Steiner Verlag, 2014. EUR 52 (cloth). Isis, 106(2):485– 486, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682808.

Smith:2015:BRH

[1593] David Smith. Book review: Helen Zoe Veit, Modern Food, Moral Food: Self-Control, Science, and the Rise of Modern American Eating in the Early Twentieth Century, xiii + 300 pp., bibl., index. Chapel Hill: University of North Carolina Press, 2013. $39.95 (cloth). Isis, 106(2):486– 487, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682809.

Borrelli:2015:BRC

[1594] Arianna Borrelli. Book review: Christophe Eckes, Les groupes de Lie dans l’oeuvre de Hermann Weyl, 401 pp., illus., bibl., index. Nancy: Presses Universitaires de Nancy, 2013. EUR 19 (paper). Isis, 106(2): 487–488, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682810.

Weindling:2015:BRR

[1595] Paul Weindling. Book review: Robert J. Richards, Was Hitler a Darwinian? Disputed Questions in the History of Evolutionary Theory, 269 pp., illus., bibl., index. Chicago/London: University of Chicago Press, 2013. $27.50 (paper). Isis, 106(2):488–490, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682811.

Herran:2015:BRL

[1596] Néstor Herran. Book review: Lino Camprubí, Engineers and the Making of the Francoist Regime, xiii + 298 pp., illus., bibl., index. Cambridge, Mass./London: MIT Press, 2014. $40 (cloth). Isis, 106(2):490–491, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/682812.

Kursell:2015:BRV

[1597] Julia Kursell. Book review: Volker Smyrek, Die Geschichte des Tonmischpults: Die technische Entwicklung der Mischpulte und der Wandel der medialen Produktionsverfahren im Tonstudio von den 1920er-Jahren bis heute, xxix + 539 pp., illus., bibl., indexes. Berlin: Logos Verlag, 2013. EUR 57 (paper). Isis, 106(2):491–492, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682813.

Abir-Am:2015:BRO

[1598] Pnina G. Abir-Am. Book review: Oren Harman; Michael R. Dietrich, eds. Outsider Scientists: Routes to Innovation in Biology, xi + 374 pp., illus., bibl., index. Chicago/London: University of Chicago Press, 2013. $35 (paper). Isis, 106(2):492–494, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682814.

Meikle:2015:BRM

[1599] Jeffrey Meikle. Book review: Mimi Sheller, Aluminum Dreams: The Making of Light Modernity, (Mobility Studies.) xii + 367 pp., illus., bibl., index. Cambridge, Mass./London: MIT Press, 2014. $29.95 (cloth). Isis, 106(2):494–495, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
682815.

Darnell:2015:BRJ

[1600] Regna Darnell. Book review: Joshua A. Bell, Alison K. Brown, and Robert J. Gordon, eds. Recreating First Contact: Expeditions, Anthropology, and Popular Culture, xii + 261 pp., illus., bibls., index. Washington, D.C.: Smithsonian Institution Scholarly Press, 2013. $49.95 (cloth). Isis, 106(2):495–496, June 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
682816.

Richmond:2015:BRS

[1601] Marsha L. Richmond. Book review: Sarah S. Richardson, Sex Itself: The Search for Male and Female in the Human Genome, vii + 311 pp., illus., tables, bibl., index. Chicago/London: University of Chicago Press, 2013. $45 (cloth). Isis, 106(2):496–497, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682817.

Jenks:2015:BRP

[1602] Andrew L. Jenks. Book review: Paul R. Josephson, The Conquest of the Russian Arctic, viii + 441 pp., illus., index. Cambridge, Mass./London: Harvard University Press, 2014. $55 (cloth). Isis, 106(2):497–498, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/682818.

Wiame:2015:BRJ

[1603] Aline Wiame. Book review: John Protevi, Life, War, Earth: Deleuze and the Sciences, ix + 256 pp., bibl., index. Minneapolis/London: University of Minnesota Press, 2013. $25 (paper). Isis, 106(2):499–500, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/682819.

Smith:2015:BRT

[1604] Robert W. Smith. Book review: Thomas Gold, Taking the Back Off the Watch: A Personal Memoir, (Astrophysics and Space Science Library, 381.) Edited by Simon Mitton. xii + 233 pp., illus., index. Berlin:

Springer, 2012. $119 (cloth). Isis, 106(2):500–501, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/682820.

Isaac:2015:BRP

[1605] Joel Isaac. Book review: Paul Erickson, Judy L. Klein, Lorraine Daston, Rebecca Lemov, Thomas Sturm, and Michael D. Gordin, How Reason Almost Lost Its Mind: The Strange Career of Cold War Rationality, viii + 259 pp., illus., bibl., index. Chicago/London: University of Chicago Press, 2013. $35 (cloth). Isis, 106(2):501–502, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/682821.

Karaca:2015:BRA

[1606] Koray Karaca. Book review: Allan Franklin, Shifting Standards: Experiments in Particle Physics in the Twentieth Century, lv + 302 pp., illus., tables, bibl., index. Pittsburgh, Pa.: University of Pittsburgh Press, 2013. $50 (cloth). Isis, 106(2):502–503, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682822.

Randalls:2015:BRJ

[1607] Samuel Randalls. Book review: Joshua P. Howe, Behind the Curve: Science and the Politics of Global Warming, xviii + 290 pp., illus., bibl., index. Seattle/London: University of Washington Press, 2014. $34.95 (cloth). Isis, 106(2):503–505, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682823.

Wallis:2015:BRJ

[1608] Faith Wallis. Book review: Jacalyn Duffin, Medical Saints: Cosmas and Damian in a Postmodern World, ix + 229 pp., illus., bibl., index. Oxford: Oxford University Press, 2013. £18 (cloth). Isis, 106(2):505–506, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/682824.

David:2015:BRE

[1609] Michael Z. David. Book review: Erin Koch, Free Market Tuberculosis: Managing Epidemics in Post-Soviet Georgia, xiv + 231 pp., illus., bibl., index. Nashville, Tenn.: Vanderbilt University Press, 2013. $59.95 (cloth). Isis, 106(2):506–507, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/682825.

Misa:2015:BRL

[1610] Thomas J. Misa. Book review: Laura DeNardis, The Global War for Internet Governance, viii + 288 pp., illus., tables, index. New Haven, Conn./London: Yale University Press, 2014. $38 (cloth). Isis, 106(2): 507–508, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/682826.

Anonymous:2015:NCb

[1611] Anonymous. Notes on contributors. Isis, 106(2):509–515, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/681984.

Nicolai:2015:POP

[1612] Roel Nicolai. The premedieval origin of Portolan charts: New geodetic evidence. Isis, 106(3):517–543, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/683532.

White:2015:UAC

[1613] Sam White. Unpuzzling American climate: New world experience and the foundations of a new science. Isis, 106(3):544–566, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683166.

Morawski:2015:EDP

[1614] Jill Morawski. Epistemological dizziness in the psychology laboratory: Lively subjects, anxious experimenters, and experimental relations, 1950–1970. Isis, 106(3):567–597, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/683411.

Blum:2015:RGR

[1615] Alexander Blum, Roberto Lalli, and Jürgen Renn. The reinvention of general relativity: A historiographical framework for assessing one hundred years of curved space-time. Isis, 106(3):598–620, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683425.

Cowles:2015:I

[1616] Henry M. Cowles, William Deringer, Stephanie Dick, and Colin Webster. Introduction. Isis, 106(3):621–622, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/10.1086/683196.

Dick:2015:MMI

[1617] Stephanie Dick. Of models and machines: Implementing bounded rationality. Isis, 106(3):623–634, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/683527.

Cowles:2015:HBT

[1618] Henry M. Cowles. Hypothesis bound: Trial and error in the Nineteenth Century. Isis, 106(3):635–645, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/683528.

Deringer:2015:WWH

[1619] William Deringer. For what it’s worth: Historical financial bubbles and the boundaries of economic rationality. Isis, 106(3):646–656, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/683529.

Webster:2015:HMM

[1620] Colin Webster. Heuristic medicine: The methodists and metalepsis. Isis, 106(3):657–668, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/683530.

Daston:2015:SSC

[1621] Lorraine Daston. Simon and the sirens: A commentary. Isis, 106(3): 669–676, September 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
683531.

Sawday:2015:BRF

[1622] Jonathan Sawday. Book review: The Fabric of the Human Body: An Annotated Translation of the 1543 and 1555 Editions of “De Humani Corporis Fabrica Libri Septem”, by Andreas Vesalius, edited by Daniel H. Garrison and Malcolm H. Hast. Isis, 106(3):677–683, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683457.

vanDongen:2015:ERA

[1623] Jeroen van Dongen. Essay review: Albert Einstein and the History and Philosophy of Science, The Cambridge Companion to Einstein, edited by Michel Janssen and Christoph Lehner. Isis, 106(3):684–689, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/683458.

Fehige:2015:BRJ

[1624] Yiftach Fehige and Adam Richter. Book review: John Doody; Adam Goldstein; Kim Paffenroth, eds., Augustine and Science. Isis, 106(3): 690–691, September 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
683167.

Daniels:2015:BRA

[1625] Tim Daniels. Book review: Ann Hibner Koblitz, Sex and Herbs and Birth Control: Women and Fertility Regulation through the Ages. Isis, 106 (3):691–692, September 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
683405.

Keblusek:2015:BRK

[1626] Marika Keblusek. Book review: Kristie Macrakis, Prisoners, Lovers, and Spies: The Story of Invisible Ink from Herodotus to Al-Qaeda. Isis, 106 (3):692–693, September 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
683195.

Santing:2015:BRH

[1627] Catrien Santing. Book review: Helen King, The One-Sex Body on Trial: The Classical and Early Modern Evidence. Isis, 106(3):693–695, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/683168.

Hutchins:2015:BRE

[1628] Barnaby R. Hutchins. Book review: Eric Watkins, ed., The Divine Order, the Human Order, and the Order of Nature: Historical Perspectives. Isis, 106(3):695–696, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/683414.

Huistra:2015:BRJ

[1629] Hieke Huistra. Book review: Jane Maienschein, Embryos under the Microscope: The Diverging Meanings of Life. Isis, 106(3):696–697, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/683248.

Worthen:2015:BRL

[1630] Shana Worthen. Book review: Lino Conti, ed., Natura e scienza nella rivoluzione francescana. Isis, 106(3):697–698, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/683459.

Fransen:2015:BRU

[1631] Sietske Fransen. Book review: Ute Frietsch, Häresie und Wissenschaft: Eine Genealogie der paracelsischen Alchemie. Isis, 106(3):699–700, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/683460.

Land:2015:BRM

[1632] Karine van ‘t Land. Book review: Marilyn Nicoud; Danielle Jacquart, Le prince et les médecins: Pensée et pratiques médicales à Milan (1402– 1476) Être médecin à la cour (Italie, France, Espagne, XIIIe–XVIIIe siècle). Isis, 106(3):700–702, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/683423.

Henderson:2015:BRS

[1633] Felicity Henderson. Book review: Sarah Dry, The Newton Papers: The Strange and True Odyssey of Isaac Newton’s Manuscripts. Isis, 106(3): 702–703, September 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
683247.

Brenni:2015:BRS

[1634] Paolo Brenni. Book review: Silke Ackermann; Richard L. Kremer; Mara Miniati, eds., Scientific Instruments on Display. Isis, 106(3):703–704, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/683169.

Stewart:2015:BRS

[1635] Keith Stewart. Book review: Susan P. Mattern, The Prince of Medicine: Galen in the Roman Empire. Isis, 106(3):704–705, September 2015. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683249.

Kremer:2015:BRD

[1636] Richard L. Kremer. Book review: Dieter Blume; Mechthild Haffner; Wolfgang Metzger. Sternbilder des Mittelalters: Der gemalte Himmel zwischen Wissenschaft und Phantasie. Isis, 106(3):705–707, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/683250.

Sylla:2015:BRJ

[1637] Edith Dudley Sylla. Book review: Joel Kaye, A History of Balance, 1250–1375: The Emergence of a New Model of Equilibrium and Its Impact on Thought. Isis, 106(3):707–708, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/683246.

Guerrini:2015:BRK

[1638] Anita Guerrini. Book review: Karen Raber, Animal Bodies, Renaissance Culture. Isis, 106(3):708–709, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/683190.

Ryan:2015:BRB

[1639] Michael A. Ryan. Book review: Brendan Dooley, ed., A Companion to Astrology in the Renaissance. Isis, 106(3):709–710, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683245.

Felfe:2015:BRK

[1640] Robert Felfe. Book review: Katharina Pilaski Kaliardos, The Munich Kunstkammer: Art, Nature, and the Representation of Knowledge in Courtly Contexts. Isis, 106(3):710–712, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/683409.

Bresadola:2015:BRL

[1641] Marco Bresadola. Book review: Lino Conti, ed., Itinerari della medicina nel Seicento. Frate Evangelista Quattrami, Breve trattato intorno alla preservatione, et cura della peste, Lino Conti; Paolo Capitanucci, eds. Isis, 106(3):712–713, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/683402.

Hsia:2015:BRR

[1642] Florence C. Hsia. Book review: Roger Hart, Imagined Civilizations: China, the West, and Their First Encounter. Isis, 106(3):713–716, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/683413.

Clarke:2015:BRR

[1643] Desmond M. Clarke. Book review: Raffaella De Rosa, Descartes and the Puzzle of Sensory Representation. Isis, 106(3):716–717, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683404.

Hansen:2015:BRT

[1644] Jens Morten Hansen. Book review: Troels Kardel; Paul Maquet, eds., Nicolaus Steno: Biography and Original Papers of a Seventeenth-Century Scientist. Isis, 106(3):717–718, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/683410.

Nadler:2015:BRN

[1645] Steven Nadler. Book review: Nicholas Jolley, Causality and Mind: Essays on Early Modern Philosophy. Isis, 106(3):718–719, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683420.

Dross:2015:BRS

[1646] Fritz Dross. Book review: Sabine Herrmann, Giacomo Casanova und die Medizin des 18. Jahrhunderts. Isis, 106(3):719–720, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683407.

Eddy:2015:BRR

[1647] Matthew Daniel Eddy. Book review: Robert G. W. Anderson; Jean Jones, eds., The Correspondence of Joseph Black. Isis, 106(3):720–722, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/683189.

Klaver:2015:BRI

[1648] Jan Marten Ivo Klaver. Book review: Ib Friis; Michael Harbsmeier; Jørgen Bæk Simonsen, eds., Early Scientific Expeditions and Local Encounters: New Perspectives on Carsten Niebuhr and “The Arabian Journey”: Proceedings of a Symposium on the Occasion of the 250th Anniversary of the Royal Danish Expedition to Arabia Felix. Isis, 106(3):722–723, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/683416.

Andrews:2015:BRH

[1649] Jonathan Andrews. Book review: Heather R. Beatty, Nervous Disease in Late Eighteenth-Century Britain: The Reality of a Fashionable Disorder.

Isis, 106(3):723–724, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/683188.

Beretta:2015:BRP

[1650] Marco Beretta. Book review: Patrice Bret; Henri Kagan, eds., Oeuvres de Lavoisier: Correspondance. Isis, 106(3):724–726, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683400.

Reidy:2015:BRP

[1651] Michael S. Reidy. Book review: Peter H. Hansen, The Summits of Modern Man: Mountaineering after the Enlightenment. Isis, 106(3):726–727, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/683421.

Dibattista:2015:BRW

[1652] Liborio Dibattista. Book review: Walter Tega, Tradizione e rivoluzione: Scienza e potere in Francia (1815–1840). Isis, 106(3):727–728, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/683406.

Babb:2015:BRG

[1653] Genie Babb. Book review: Gowan Dawson; Bernard Lightman, eds., Victorian Scientific Naturalism: Community, Identity, Continuity. Isis, 106 (3):728–729, September 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
683399.

Casado:2015:BRA

[1654] Santos Casado. Book review: Augusto González de Linares; Carlos Nieto Blanco, eds., Obra completa. Isis, 106(3):730–731, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683403.

vanDelft:2015:BRR

[1655] Dirk van Delft. Book review: S. Rowlinson, Sir James Dewar, 1842– 1923: A Ruthless Chemist. Isis, 106(3):731–732, September 2015. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683422.

England:2015:BRB

[1656] Richard England. Book review: Bradley Gundlach, Process and Providence: The Evolution Question at Princeton, 1845–1929. Isis, 106(3):

732–733, September 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
683408.

Wolff:2015:BRW

[1657] Stefan L. Wolff. Book review: Wenchao Li; Hartmut Rudolph, eds., “Leibniz” in der Zeit des Nationalsozialismus. Isis, 106(3):733–734, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/683424.

Kline:2015:BRW

[1658] Wendy Kline. Book review: Wendy Mitchinson, Body Failure: Medical Views of Women, 1900–1950. Isis, 106(3):735–736, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/683417.

Borossa:2015:BRJ

[1659] Julia Borossa. Book review: John Burnham, ed., After Freud Left: A Century of Psychoanalysis in America. Isis, 106(3):736–737, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/683401.

McMichael:2015:BRA

[1660] Anthony McMichael. Book review: Alison Bashford, Global Population: History, Geopolitics, and Life on Earth. Isis, 106(3):737–738, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/683419.

Aronova:2015:BRJ

[1661] Elena Aronova. Book review: Jacob Darwin Hamblin, Arming Mother Nature: The Birth of Catastrophic Environmentalism. Isis, 106(3): 738–739, September 2015. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.1086/
683398.

Lemov:2015:BRJ

[1662] Rebecca Lemov. Book review: Jamie Cohen-Cole, The Open Mind: Cold War Politics and the Sciences of Human Nature. Isis, 106(3):739–741, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/683418.

Heyck:2015:BRJ

[1663] Hunter Heyck. Book review: Joel Isaac; Duncan Bell, eds., Uncertain Empire: American History and the Idea of the Cold War Armed with Expertise: The Militarization of American Social Research during the Cold War. Isis, 106(3):741–742, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/683412.

Junod:2015:BRS

[1664] Suzanne White Junod. Book review: Soraya Boudia; Nathalie Jas, eds., Toxicants, Health, and Regulation since 1945. Isis, 106(3):742–743, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/683415.

Anonymous:2015:NCc

[1665] Anonymous. Notes on contributors. Isis, 106(3):744–748, September 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/683656.

Watts:2015:PIL

[1666] Iain P. Watts. Philosophical intelligence: Letters, print, and experiment during Napoleon’s continental blockade. Isis, 106(4):749–770, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Naylor:2015:LBL

[1667] Simon Naylor. Log books and the law of storms: Maritime meteorology and the British Admiralty in the Nineteenth Century. Isis, 106(4):771– 797, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Raby:2015:AAM

[1668] Megan Raby. Ark and archive: Making a place for long-term research on Barro Colorado Island, Panama. Isis, 106(4):798–824, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Haagensen:2015:MRC

[1669] Erling Haagensen and Niels C. Lind. Medieval round churches and the shape of the Earth. Isis, 106(4):825–834, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Davids:2015:IBC

[1670] Karel Davids. Introduction: Bridging concepts. Isis, 106(4):835–839, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Long:2015:TZE

[1671] Pamela O. Long. Trading zones in Early Modern Europe. Isis, 106 (4):840–847, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Popplow:2015:FIT

[1672] Marcus Popplow. Formalization and interaction: Toward a comprehensive history of technology-related knowledge in Early Modern Europe. Isis, 106(4):848–856, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roberts:2015:PEA

[1673] Lissa Roberts. Producing (in) Europe and Asia, 1750–1850. Isis, 106 (4):857–865, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Davids:2015:MSO

[1674] Karel Davids. On machines, self-organization, and the global traveling of knowledge, circa 1500–1900. Isis, 106(4):866–874, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2015:EIG

[1675] Anonymous. Eloge: Ivor Grattan-Guinness 1941–2014. Isis, 106(4):875– 879, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Snobelen:2015:ERI

[1676] Stephen D. Snobelen. Essay review: Isaac Newton, historian: Redivivus: Jed Z. Buchwald; Mordechai Feingold. Newton and the Origin of Civilization. ix + 528 pp., illus., tables, index. Princeton, N.J.: Princeton University Press, 2013. $49.50 (cloth). Isis, 106(4):880–888, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kuhl:2015:BRZ

[1677] Kevin Kuhl. Book review: Zvi Artstein; Alan Hercberg. Mathematics and the Real World: The Remarkable Role of Evolution in the Making of Mathematics. Translated by Alan Hercberg. 426 pp., illus., figs., tables, bibl., index. Amherst, N.Y.: Prometheus Books, 2014. $26 (cloth). Isis, 106(4):889–890, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hardy:2015:BRC

[1678] Anne Hardy. Book review: Christopher Hamlin. More Than Hot: a Short History of Fever. Isis, 106(4):890–891, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wear:2015:BRM

[1679] Andrew Wear. Book review: Manfred Horstmanshoff; Helen King; Claus Zittel, eds. Blood, Sweat and Tears: The Changing Concepts of Physiology from Antiquity into Early Modern Europe. Isis, 106(4):891–892, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Spary:2015:BRJ

[1680] Emma C. Spary. Book review: Jonathan Reinarz. Past Scents: Historical Perspectives on Smell. Isis, 106(4):892–894, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

LeRoux:2015:BRJ

[1681] Thomas Le Roux. Book review: James Rodger Fleming; Ann Johnson, eds. Toxic Airs: Body, Place, Planet in Historical Perspective. Isis, 106 (4):894–895, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Henry:2015:BRR

[1682] John Henry. Book review: Rens Bod. A New History of the Humanities: The Search for Principles and Patterns from Antiquity to the Present. Isis, 106(4):895–896, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gingras:2015:BRA

[1683] Yves Gingras. Book review: Anne-Lise Rey, ed. Méthode et histoire: Quelle histoire font les historiens des sciences et des techniques? Isis, 106 (4):896–897, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barwich:2015:BRA

[1684] Ann-Sophie Barwich. Book review: Astrid Schwarz. Experiments in Practice. Isis, 106(4):898–899, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hopwood:2015:BRC

[1685] Nick Hopwood. Book review: Catelijne Coopmans; Janet Vertesi; Michael Lynch; Steve Woolgar, eds. Representation in Scientific Practice Revisited. Isis, 106(4):899–901, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Woods:2015:BRA

[1686] Rebecca J. H. Woods. Book review: Andrea S. Wiley. Cultures of Milk: The Biology and Meaning of Dairy Products in the United States and India. Isis, 106(4):901–902, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tousignant:2015:BRA

[1687] Noémi Tousignant. Book review: Abena Dove Osseo-Asare. Bitter Roots: The Search for Healing Plants in Africa. Isis, 106(4):902–903, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weidman:2015:BRJ

[1688] Nadine Weidman. Book review: James Tabery. Beyond Versus: The Struggle to Understand the Interaction of Nature and Nurture. Isis, 106 (4):903–904, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dean:2015:BRP

[1689] Carolyn Dean. Book review: Prudence M. Rice. Space-Time Perspectives on Early Colonial Moquegua. Isis, 106(4):905–906, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mellyn:2015:BRT

[1690] Elizabeth W. Mellyn. Book review: Timothy S. Miller; John W. Nesbitt. Walking Corpses: Leprosy in Byzantium and the Medieval West. Isis, 106 (4):906–907, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Verboon:2015:BRT

[1691] Annemieke Verboon. Book review: Tanja Klemm. Bildphysiologie: Wahrnehmung und Körper in Mittelalter und Renaissance. Isis, 106(4): 907–908, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nijdam:2015:BRA

[1692] Han Nijdam. Book review: Anne Kirkham; Cordelia Warr, eds. Wounds in the Middle Ages. Isis, 106(4):908–909, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mosshammer:2015:BRC

[1693] Alden Mosshammer. Book review: C. Philipp E. Nothaft. Medieval Latin Christian Texts on the Jewish Calendar: a Study with Five Editions and Translations. Isis, 106(4):910–911, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Andretta:2015:BRE

[1694] Elisa Andretta. Book review: Elizabeth W. Mellyn. Mad Tuscans and Their Families: a History of Mental Disorder in Early Modern Italy. Isis, 106(4):911–912, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hall:2015:BRJ

[1695] Bert Hall. Book review: Johannes Fontana; Horst Kranz, eds. Liber instrumentorum iconographicus: Ein illustriertes Maschinenbuch. Isis, 106 (4):912–913, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hirai:2015:BRC

[1696] Hiro Hirai. Book review: Craig Martin. Subverting Aristotle: Religion, History, and Philosophy in Early Modern Science. Isis, 106(4):913–914, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blumenthal:2015:BRA

[1697] Debra Blumenthal. Book review: Adrian Wilson. Ritual and Conflict: The Social Relations of Childbirth in Early Modern England. Isis, 106 (4):915–916, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Christianson:2015:BRM

[1698] J. R. Christianson. Book review: Miguel A. Granada; Adam Mosley; Nicholas Jardine, eds. Christoph Rothmann’s Discourse on the Comet of 1585: an Edition and Translation with Accompanying Essays. Isis, 106 (4):916–917, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gal:2015:BRE

[1699] Ofer Gal. Book review: Eileen Reeves. Evening News: Optics, Astronomy, and Journalism in Early Modern Europe. Isis, 106(4):917–918, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Valleriani:2015:BRP

[1700] Matteo Valleriani. Book review: Paolo Galluzzi, ed. Libertà di filosofare in naturalibus: I mondi paralleli di Cesi e Galileo. Isis, 106(4):919–920, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hunter:2015:BRD

[1701] Michael Hunter. Book review: Douglas D. C. Chambers; David Galbraith, eds. The Letterbooks of John Evelyn. Isis, 106(4):920–921, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roos:2015:BRK

[1702] Anna Marie Roos. Book review: Karl A. E. Enenkel; Paul J. Smith, eds. Zoology in Early Modern Culture: Intersections of Science, Theology, Philology, and Political and Religious Education. Isis, 106(4):921–922, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanNetten:2015:BRD

[1703] Djoeke van Netten. Book review: Dániel Margócsy. Commercial Visions: Science, Trade, and Visual Culture in the Dutch Golden Age. Isis, 106 (4):922–927, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Miller:2015:BRN

[1704] David Philip Miller. Book review: Neil Chambers, ed. The Indian and Pacific Correspondence of Sir Joseph Banks, 1768–1820. Isis, 106(4): 924–925, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Yenter:2015:BRJ

[1705] Timothy Yenter. Book review: Jamie C. Kassler. Seeking Truth: Roger North’s Notes on Newton and Correspondence with Samuel Clarke, c. 1704–1713. Isis, 106(4):925–926, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Meli:2015:BRM

[1706] Domenico Bertoloni Meli. Book review: Maria Pia Donato. Sudden Death: Medicine and Religion in Eighteenth-Century Rome. Isis, 106 (4):926–927, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reeves:2015:BRR

[1707] Nicky Reeves. Book review: Richard Sorrenson. Perfect Mechanics: Instrument Makers at the Royal Society of London in the Eighteenth Century. Isis, 106(4):927–929, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bigelow:2015:BRE

[1708] Allison Margaret Bigelow. Book review: Emily Berquist Soule. The Bishop’s Utopia: Envisioning Improvement in Colonial Peru. Isis, 106 (4):929–930, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Klaver:2015:BRL

[1709] Jan Marten Ivo Klaver. Book review: Lawrence J. Baack. Undying Curiosity: Carsten Niebuhr and the Royal Danish Expedition to Arabia, 1761–1767. Isis, 106(4):930–931, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Duchesneau:2015:BRI

[1710] François Duchesneau. Book review: Ina Goy; Eric Watkins, eds. Kant’s Theory of Biology. Isis, 106(4):931–932, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Williams:2015:BRE

[1711] Elizabeth A. Williams. Book review: Emma C. Spary. Feeding France: New Sciences of Food, 1760–1815. Isis, 106(4):933–934, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Greene:2015:BRD

[1712] Mott T. Greene. Book review: David I. Spanagel. DeWitt Clinton and Amos Eaton: Geology and Power in Early New York. Isis, 106(4):934– 935, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ehrhardt:2015:BRY

[1713] Caroline Ehrhardt. Book review: Yvette Kosmann-Schwarzbach, ed. Siméon-Denis Poisson: Les mathématiques au service de la science. Isis, 106(4):935–936, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gregory:2015:BRS

[1714] Frederick Gregory. Book review: Susanne Lettow, ed. Reproduction, Race, and Gender in Philosophy and the Early Life Sciences. Isis, 106 (4):936–938, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fayter:2015:BRC

[1715] Paul Fayter. Book review: Charles Mollan, ed. William Parsons, Third Earl of Rosse: Astronomy and the Castle in Nineteenth-Century Ireland. Isis, 106(4):938–939, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Martin:2015:BRA

[1716] Alison E. Martin. Book review: Alexander von Humboldt; Stephen T. Jackson; Laura Dassow Walls; Mark W. Person, eds. Views of Nature. Isis, 106(4):939–940, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hays:2015:BRK

[1717] J. N. Hays. Book review: Krista Maglen. The English System: Quarantine, Immigration and the Making of a Port Sanitary Zone. Isis, 106 (4):940–941, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barton:2015:BRJ

[1718] Ruth Barton. Book review: James A. Secord. Visions of Science: Books and Readers at the Dawn of the Victorian Age. Isis, 106(4):942–943, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Clark:2015:BRT

[1719] J. F. M. Clark. Book review: Takashi Ito. London Zoo and the Victorians, 1828–1859. Isis, 106(4):943–944, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ghiselin:2015:BRF

[1720] Michael Ghiselin. Book review: Frederick Burkhardt; James Secord, eds. The Correspondence of Charles Darwin. Isis, 106(4):944–946, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Meneghello:2015:BRJ

[1721] Laura Meneghello. Book review: Johannes Zachhuber. Theology as Science in Nineteenth-Century Germany: From F. C. Baur to Ernst Troeltsch. Isis, 106(4):946–948, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dryer:2015:BRG

[1722] Theodora Dryer. Book review: George Steinmetz, ed. Sociology and Empire: The Imperial Entanglements of a Discipline. Isis, 106(4):948– 949, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kerr:2015:BRP

[1723] Ashley Kerr. Book review: Patience A. Schell. The Sociable Sciences: Darwin and His Contemporaries in Chile. Isis, 106(4):949–950, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dean:2015:BRJ

[1724] Sheila Ann Dean. Book review: J. David Pleins. In Praise of Darwin: George Romanes and the Evolution of a Darwinian Believer. Isis, 106 (4):951–952, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pollock:2015:BRL

[1725] Anne Pollock. Book review: Lundy Braun. Breathing Race into the Machine: The Surprising Career of the Spirometer from Plantation to Genetics. Isis, 106(4):952–953, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Becker:2015:BRJ

[1726] Barbara J. Becker. Book review: John B. Hearnshaw. The Analysis of Starlight: Two Centuries of Astronomical Spectroscopy. Isis, 106(4):953– 954, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hale:2015:BRB

[1727] Piers J. Hale. Book review: Bernard Lightman; Bennett Zon, eds. Evolution and Victorian Culture. Isis, 106(4):954–956, December 2015. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Richards:2015:BRK

[1728] Evelleen Richards. Book review: Kimberly A. Hamlin. From Eve to Evolution: Darwin, Science, and Women’s Rights in Gilded Age America. Isis, 106(4):956–957, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mildenberger:2015:BRA

[1729] Florian G. Mildenberger. Book review: Avi Sharma. We Lived for the Body: Natural Medicine and Public Health in Imperial Germany. Isis, 106(4):957–958, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bravo:2015:BRL

[1730] Michael Bravo. Book review: Ludger Müller-Wille. The Franz Boas Enigma: Inuit, Arctic, and Sciences. Isis, 106(4):958–959, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mergoupi-Savaidou:2015:BRW

[1731] Eirini Mergoupi-Savaidou. Book review: W. Boyd Rayward, ed. Information beyond Borders: International Cultural and Intellectual Exchange in the Belle Époque. Isis, 106(4):959–960, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanderMeer:2015:BRF

[1732] Jitse van der Meer. Book review: Ab Flipse. Christelijke wetenschap: Nederlandse rooms-katholieken en gereformeerden over de natuurwetenschap, 1880–1940. Isis, 106(4):960–962, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Grunden:2015:BRJ

[1733] Walter E. Grunden. Book review: Jing Tsu; Benjamin A. Elman, eds. Science and Technology in Modern China, 1880s–1940s. Isis, 106(4): 962–963, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hacker:2015:BRK

[1734] Barton C. Hacker. Book review: Katherine C. Epstein. Torpedo: Inventing the Military-Industrial Complex in the United States and Great Britain. Isis, 106(4):963–964, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gilman:2015:BRL

[1735] Sander L. Gilman. Book review: Luise Hirsch. From the Shtetl to the Lecture Hall: Jewish Women and Cultural Exchange. Isis, 106(4):964– 965, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kevles:2015:BRA

[1736] Bettyann Holtzmann Kevles. Book review: Adrian M. K. Thomas; Arpan K. Banerjee. The History of Radiology. Isis, 106(4):966–967, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Grout:2015:BRP

[1737] Holly Grout. Book review: Paula Martin. Suzanne Noël: Cosmetic Surgery, Feminism and Beauty in Early Twentieth-Century France. Isis, 106(4):967–968, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gordin:2015:BRN

[1738] Michael D. Gordin. Book review: Noah J. Efron. A Chosen Calling: Jews in Science in the Twentieth Century. Isis, 106(4):968–969, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hardy:2015:BRG

[1739] Anne Hardy. Book review: George Weisz. Chronic Disease in the Twentieth Century: a History. Isis, 106(4):969–970, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Travis:2015:BRK

[1740] Anthony S. Travis. Book review: Kathryn Steen. The American Synthetic Organic Chemicals Industry: War and Politics, 1910–1930. Isis, 106(4):970–972, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rentetzi:2015:BRF

[1741] Maria Rentetzi. Book review: Finn Aaserud; J. L. Heilbron. Love, Literature, and the Quantum Atom: Niels Bohr’s 1913 Trilogy Revisited.

Isis, 106(4):972–973, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Zitarelli:2015:BRR

[1742] David E. Zitarelli. Book review: Roman Duda; Daniel Davies. Pearls from a Lost City: The Lvov School of Mathematics. Isis, 106(4):973– 974, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Thompson:2015:BRG

[1743] Roger R. Thompson. Book review: Grace Yen Shen. Unearthing the Nation: Modern Geology and Nationalism in Republican China. Isis, 106 (4):974–976, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Azzouni:2015:BRA

[1744] Safia Azzouni. Book review: Anke te Heesen; Lori Lantz. The Newspaper Clipping: a Modern Paper Object. Isis, 106(4):976–977, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Puig:2015:BRN

[1745] Albert Presas Puig. Book review: Néstor Herran; Xavier Roqué, eds. La física en la dictadura: Físicos, cultura y poder en Enspaña 1939–1975. Isis, 106(4):977–979, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Meldrum:2015:BRC

[1746] Marcia Meldrum. Book review: Carsten Timmermann. A History of Lung Cancer: The Recalcitrant Disease. Isis, 106(4):979–980, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Redman:2015:BRC

[1747] Emily Redman. Book review: Christopher Hollings. Mathematics across the Iron Curtain: a History of the Algebraic Theory of Semigroups. Isis, 106(4):980–981, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Smith:2015:BRA

[1748] Robert W. Smith. Book review: Andrew J. Butrica. The Navigators: a History of NASA’s Deep-Space Navigation. Isis, 106(4):981–982, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Winsor:2015:BRA

[1749] Mary P. Winsor. Book review: Andrew Hamilton, ed. The Evolution of Phylogenetic Systematics. Isis, 106(4):982–983, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mosby:2015:BRL

[1750] Ian Mosby. Book review: Laura Dawes. Childhood Obesity in America: Biography of an Epidemic. Isis, 106(4):983–984, December 2015. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lungu:2015:BRA

[1751] Dov Lungu. Book review: Andrew L. Russell. Open Standards and the Digital Age: History, Ideology, and Networks. Isis, 106(4):985–986, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rodriguez:2015:BRP

[1752] Julia Rodriguez. Book review: Peter Wade; Carlos López Beltrán; Eduardo Restrepo; Ricardo Ventura Santos, eds. Mestizo Genomics: Race Mixture, Nation, and Science in Latin America. Isis, 106(4):986–987, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Matthews:2015:BRJ

[1753] J. Rosser Matthews. Book review: John S. Haller. Shadow Medicine: The Placebo in Conventional and Alternative Therapies. Isis, 106(4): 987–988, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stapleford:2015:BRT

[1754] Thomas Stapleford. Book review: Till Düppe; E. Roy Weintraub. Finding Equilibrium: Arrow, Debreu, McKenzie and the Problem of Scientific Credit. Isis, 106(4):988–990, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blum:2015:BRD

[1755] Alexander Blum. Book review: Dean Rickles. A Brief History of String Theory: From Dual Models to M-Theory. Isis, 106(4):990–991, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Logsdon:2015:BRW

[1756] John M. Logsdon. Book review: W. Henry Lambright. Why Mars: NASA and the Politics of Space Exploration. Isis, 106(4):991–992, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Elmore:2015:BRP

[1757] Bart Elmore. Book review: Paul Sabin. The Bet: Paul Ehrlich, Julian Simon, and Our Gamble over Earth’s Future. Isis, 106(4):992–993, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harrington:2015:BRE

[1758] Anne Harrington. Book review: Elizabeth Lunbeck. The Americanization of Narcissism. Isis, 106(4):993–995, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stingl:2015:BRA

[1759] Alexander I. Stingl. Book review: Amit Prasad. Imperial Technoscience: Transnational Histories of MRI in the United States, Britain, and India. Isis, 106(4):995–996, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2015:NC

[1760] Anonymous. Notes on contributors. Isis, 106(4):997–1004, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2015:BRI

[1761] Anonymous. Book review: Index to Isis, volume 106, 2015. Isis, 106(4): 1005–1024, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2015:BRC

[1762] Anonymous. Book review: Contents of Isis, volume 106. Isis, 106(4): 1025–1026, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2015:CBH

[1763] Anonymous. Current bibliography of the history of science and its cultural influences, 2015. Isis, 106(S1):i–300, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2016:Ma

[1764] Anonymous. Masthead. Isis, 107(1):??, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Statman:2016:FES

[1765] Alexander Statman. Fusang: The enlightenment story of the Chinese discovery of America. Isis, 107(1):1–25, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Toal:2016:PCM

[1766] Ciaran Toal. Protestants, Catholics, and Masonic conspiracies: The British Association in Montreal (1884). Isis, 107(1):26–48, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Radick:2016:UMS

[1767] Gregory Radick. The unmaking of a modern synthesis: Noam Chomsky, Charles Hockett, and the politics of behaviorism, 1955–1965. Isis, 107(1): 49–73, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Yale:2016:ICA

[1768] Elizabeth Yale. Introduction: Consider the archive. Isis, 107(1):74– 76, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ogilvie:2016:SAA

[1769] Brian Ogilvie. Scientific archives in the age of digitization. Isis, 107(1): 77–85, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Popper:2016:ABE

[1770] Nicholas Popper. Archives and the boundaries of early modern science. Isis, 107(1):86–94, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Portuondo:2016:FSA

[1771] María M. Portuondo. Finding “ science” in the archives of the Spanish monarchy. Isis, 107(1):95–105, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Yale:2016:BAH

[1772] Elizabeth Yale. The book and the archive in the history of science. Isis, 107(1):106–115, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

James:2016:HAH

[1773] Kathryn James. The history of archives and the history of science: Comment. Isis, 107(1):116–120, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Porter:2016:ECC

[1774] Theodore M. Porter. Eloge: Charles Coulston Gillispie (1918–2015). Isis, 107(1):121–126, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kingsland:2016:EJL

[1775] Sharon Kingsland. Eloge: Joan Lisa Bromberg (1929–2015). Isis, 107 (1):127–129, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nicolai:2016:E

[1776] Roel Nicolai and Sander Nicolai. To the editor. Isis, 107(1):130–131, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Haagensen:2016:R

[1777] Erling Haagensen and Niels C. Lind. In reply. Isis, 107(1):132–133, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Secord:2016:E

[1778] Jim Secord. To the editor. Isis, 107(1):??, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marcaida:2016:PCH

[1779] José Ramón Marcaida. Pictures and conversation: How to study the visual cultures of science. Isis, 107(1):134–139, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fara:2016:BRL

[1780] Patricia Fara. Book review: Lynda Walsh. Scientists as Prophets: a Rhetorical Genealogy. Isis, 107(1):140–141, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Johns:2016:BRM

[1781] Adrian Johns. Book review: Michael F. Suarez; S. J.; H. R. Woudhuysen, eds. The Book: a Global History. Isis, 107(1):141–142, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Geyer-Kordesch:2016:BRM

[1782] Joanna Geyer-Kordesch. Book review: Mariacarla Gadebusch Bondio, ed. Medical Ethics: Premodern Negotiations between Medicine and Philosophy. Isis, 107(1):142–143, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chang:2016:BRU

[1783] Hasok Chang. Book review: Ursula Klein; Carsten Reinhardt, eds. Objects of Chemical Inquiry. Isis, 107(1):143–144, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Klose:2016:BRS

[1784] Joshua Klose. Book review: Stefaan Blancke; Hans Henrik Hjermitslev; Peter C. Kjærgaard; Ronald L. Numbers, eds. Creationism in Europe. Isis, 107(1):144–146, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Misra:2016:BRI

[1785] Anuj Misra. Book review: Isabelle Pingree; John M. Steele, eds. Pathways into the Study of Ancient Sciences: Selected Essays by David Pingree. Isis, 107(1):146–147, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kneeland:2016:BRS

[1786] Timothy W. Kneeland. Book review: Sharon Packer. Neuroscience in Science Fiction Films. Isis, 107(1):147–148, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sabernig:2016:BRJ

[1787] Katharina Sabernig. Book review: Janet Gyatso. Being Human in a Buddhist World: an Intellectual History of Medicine in Early Modern Tibet. Isis, 107(1):148–149, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cook:2016:BRM

[1788] Harold J. Cook. Book review: Marieke M. A. Hendriksen. Elegant Anatomy: The Eighteenth-Century Leiden Anatomical Collections; Rina Knoeff; Robert Zwijnenberg, eds. The Fate of Anatomical Collections. Isis, 107(1):149–152, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Morrison:2016:BRG

[1789] Robert Morrison. Book review: Gad Freudenthal, ed. Science in Medieval Jewish Cultures. Isis, 107(1):152–153, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gee:2016:BRJ

[1790] Emma Gee. Book review: Juanita Feros Ruys; John O. Ward; Melanie Heyworth, eds. The Classics in the Medieval and Renaissance Classroom: The Role of Ancient Texts in the Arts Curriculum as Revealed by Surviving Manuscripts and Early Printed Books. Isis, 107(1):153–155, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Falchetta:2016:BRB

[1791] Piero Falchetta. Book review: Benjamin B. Olshin. The Mysteries of the Marco Polo Maps. Isis, 107(1):155–156, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McVaugh:2016:BRA

[1792] Michael McVaugh. Book review: Azélina Jaboulet-Vercherre. The Physician, the Drinker, and the Drunk: Wine’s Uses and Abuses in Late Medieval Natural Philosophy. Isis, 107(1):156–157, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Savage-Smith:2016:BRF

[1793] Emilie Savage-Smith. Book review: F. Jamil Ragep; Faith Wallis; Pamela Miller; Adam Gacek, eds. The Herbal of al-Ghāfiqī: a Facsimile Edition with Critical Essays. Isis, 107(1):157–158, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Oosterhoff:2016:BRD

[1794] Richard J. Oosterhoff. Book review: David Albertson. Mathematical Theologies: Nicholas of Cusa and the Legacy of Thierry of Chartres. Isis, 107(1):158–160, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Valleriani:2016:BRU

[1795] Matteo Valleriani. Book review: Ulrich Alertz; Frank Haster; Thomas Kreft; Dietrich Lohrmann. Electronic Commented Edition of Leonardo da Vinci, Codex Madrid I, Madrid Biblioteca Nacional. Isis, 107(1):160– 161, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Swan:2016:BRE

[1796] Claudia Swan. Book review: Eric Jorink; Bart Ramakers, eds. Art and Science in the Early Modern Netherlands. Isis, 107(1):161–163, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wilson:2016:BRR

[1797] Lindsay Wilson. Book review: Robert Weston. Medical Consulting by Letter in France, 1665–1789; Joël Coste. Les écrits de la souffrance: La consultation médicale en France (1550–1825). Isis, 107(1):163–165, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blay:2016:BRJ

[1798] Michel Blay. Book review: John Schuster. Descartes-Agonistes: Physicomathematics, Method, Corpuscular-Mechanism, 1618–33; Mariafranca Spallanzani. Descartes: La règle de la raison. Isis, 107(1):165–166, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Omodeo:2016:BRM

[1799] Pietro Daniel Omodeo. Book review: Marian Turek, ed. Johannes Hevelius and His Gdańsk. Isis, 107(1):166–167, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Segre:2016:BRJ

[1800] Michael Segre. Book review: Joseph Agassi. The Very Idea of Modern Science: Francis Bacon and Robert Boyle. Isis, 107(1):167–168, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hutton:2016:BRB

[1801] Sarah Hutton. Book review: Brandie R. Siegfried; Lisa T. Sarasohn, eds. God and Nature in the Thought of Margaret Cavendish. Isis, 107(1):168– 170, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Findlen:2016:BRM

[1802] Paula Findlen. Book review: Marco Guardo; Raniero Orioli, eds. Cronache e statuti della Prima Accademia dei Lincei: Gesta Lynceorum, “Ristretto”delle costituzioni, Praescriptiones Lynceae Academiae.

Isis, 107(1):170–171, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marcacci:2016:BRF

[1803] Flavia Marcacci. Book review: Francesco Crapanzano. Koyré, Galileo e il “Vecchio Sogno”di Platone. Isis, 107(1):171–172, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bellis:2016:BRD

[1804] Delphine Bellis. Book review: David Marshall Miller. Representing Space in the Scientific Revolution. Isis, 107(1):172–173, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rabinovitch:2016:BRH

[1805] Oded Rabinovitch. Book review: Hiro Hirai, ed. Jacques Gaffarel: Between Magic and Science. Isis, 107(1):173–174, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Orrje:2016:BRJ

[1806] Jacob Orrje. Book review: Johan C.-E. Stén. A Comet of the Enlightenment: Anders Johan Lexell’s Life and Discoveries. Isis, 107(1):174–175, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Margocsy:2016:BRF

[1807] Dániel Margócsy. Book review: Fabian Krämer. Ein Zentaur in London: Lektüre und Beobachtung in der frühneuzeitlichen Naturforschung. Isis, 107(1):175–177, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shea:2016:BRL

[1808] William R. Shea. Book review: Lawrence Lipking. What Galileo Saw: Imagining the Scientific Revolution. Isis, 107(1):177–178, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schilt:2016:BRA

[1809] Cornelis J. Schilt. Book review: Andrew Janiak. Newton. Isis, 107(1): 178–179, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nadler:2016:BRY

[1810] Steven Nadler. Book review: Yitzhak Y. Melamed, ed. The Young Spinoza: a Metaphysician in the Making. Isis, 107(1):179–180, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ricciardo:2016:BRM

[1811] Salvatore Ricciardo. Book review: Michael Hunter. Boyle Studies: Aspects of the Life and Thought of Robert Boyle (1627–91). Isis, 107(1): 180–181, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Carlton:2016:BRC

[1812] Genevieve Carlton. Book review: Christine Marie Petto. Mapping and Charting in Early Modern England and France: Power, Patronage, and Production. Isis, 107(1):181–182, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Boschiero:2016:BRN

[1813] Luciano Boschiero. Book review: Nick Wilding. Galileo’s Idol: Gianfrancesco Sagredo and the Politics of Knowledge. Isis, 107(1):182–184, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Muller-Ahrndt:2016:BRR

[1814] Henriette Müller-Ahrndt. Book review: Robert Felfe. Naturform und bildnerische Prozesse: Elemente einer Wissensgeschichte in der Kunst des 16. und 17. Jahrhunderts. Isis, 107(1):184–185, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tessicini:2016:BRW

[1815] Dario Tessicini. Book review: Wolfgang Neuber; Thomas Rahn; Claus Zittel, eds. The Making of Copernicus: Early Modern Transformations of a Scientist and His Science. Isis, 107(1):185–186, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Outram:2016:BRT

[1816] Dorinda Outram. Book review: Theodore W. Pietsch, ed. Cuvier’s History of the Natural Sciences: Twenty-four Lessons from Antiquity to the Renaissance. Isis, 107(1):186–187, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Katzir:2016:BRA

[1817] Shaul Katzir. Book review: André Authier. Early Days of X-ray Crystallography. Isis, 107(1):187–188, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dawson:2016:BRA

[1818] Gowan Dawson. Book review: Aaron Worth. Imperial Media: Colonial Networks and Information Technologies in the British Literary Imagination, 1857–1918. Isis, 107(1):188–189, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dolan:2016:BRR

[1819] Brian Dolan. Book review: Robert J. Mayhew. Malthus: The Life and Legacies of an Untimely Prophet. Isis, 107(1):189–191, March 2016. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Otis:2016:BRH

[1820] Laura Otis. Book review: Henning Schmidgen; Nils F. Schott. The Helmholtz Curves: Tracing Lost Time. Isis, 107(1):191–192, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Klaver:2016:BRJ

[1821] Jan Marten Ivo Klaver. Book review: Jonathan Conlin. Evolution and the Victorians: Science, Culture, and Politics in Darwin’s Britain. Isis, 107(1):192–193, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hentschel:2016:BRM

[1822] Klaus Hentschel. Book review: Milena Wazeck; Geoffrey S. Koby. Einstein’s Opponents: The Public Controversy about the Theory of Relativity in the 1920s. Isis, 107(1):193–195, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Seidel:2016:BRK

[1823] Robert W. Seidel. Book review: Kenneth D. McRae. Nuclear Dawn: F. E. Simon and the Race for Atomic Weapons in World War II. Isis, 107(1):195–196, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Akhtar:2016:BRD

[1824] Leena Akhtar. Book review: Donna J. Drucker. The Classification of Sex: Alfred Kinsey and the Organization of Knowledge. Isis, 107(1):196– 197, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vetter:2016:BRC

[1825] Jeremy Vetter. Book review: Christopher F. Jones. Routes of Power: Energy and Modern America. Isis, 107(1):197–198, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Valencius:2016:BRG

[1826] Conevery Bolton Valencius. Book review: Gillen D’Arcy Wood. Tambora: The Eruption That Changed the World. Isis, 107(1):198–199, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stachel:2016:BRA

[1827] John Stachel. Book review: Arthur S. Eddington; H. G. Callaway. The Nature of the Physical World: Gifford Lectures of 1927: an Annotated Edition. Isis, 107(1):199–201, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kojevnikov:2016:BRN

[1828] Alexei Kojevnikov. Book review: Nikolai Krementsov. Revolutionary Experiments: The Quest for Immortality in Bolshevik Science and Fiction. Isis, 107(1):201–202, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Knight:2016:BRK

[1829] William Knight. Book review: Karen A. Rader; Victoria E. M. Cain. Life on Display: Revolutionizing U.S. Museums of Science and Natural History in the Twentieth Century. Isis, 107(1):202–204, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schalenberg:2016:BRH

[1830] Marc Schalenberg. Book review: Heather Ellis; Ulrike Kirchberger, eds. Anglo-German Scholarly Networks in the Long Nineteenth Century. Isis, 107(1):204–205, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pleins:2016:BRD

[1831] J. David Pleins. Book review: David N. Livingstone. Dealing with Darwin: Place, Politics, and Rhetoric in Religious Engagements with Evolution. Isis, 107(1):205–206, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

November:2016:BRN

[1832] Joseph November. Book review: Nicolas Rasmussen. Gene Jockeys: Life Science and the Rise of Biotech Enterprise. Isis, 107(1):206–208, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lanzoni:2016:BRF

[1833] Susan Lanzoni. Book review: Frank Biess; Daniel M. Gross, eds. Science and Emotions after 1945: a Transatlantic Perspective; David Cantor; Edmund Ramsden, eds. Stress, Shock, and Adaptation in the Twentieth Century. Isis, 107(1):208–210, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kerr:2016:BRE

[1834] Ashley Kerr. Book review: Eden Medina; Ivan da Costa Marques; Christina Holmes, eds. Beyond Imported Magic: Essays on Science, Technology, and Society in Latin America. Isis, 107(1):210–211, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Edelblute:2016:BRM

[1835] Heather Edelblute. Book review: Melinda Cooper; Catherine Waldby. Clinical Labor: Tissue Donors and Research Subjects in the Global Bioeconomy. Isis, 107(1):211–212, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Long:2016:BRT

[1836] Vicky Long. Book review: Teresa Ortiz-Gómez; María Jesús Santesmases, eds. Gendered Drugs and Medicine: Historical and Socio-Cultural Perspectives. Isis, 107(1):212–214, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bangham:2016:BRK

[1837] Jenny Bangham. Book review: Kara W. Swanson. Banking on the Body: The Market in Blood, Milk, and Sperm in Modern America. Isis, 107 (1):214–215, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Olby:2016:BRS

[1838] Robert Olby. Book review: Suzanne Bourgeois. Genesis of the Salk Institute: The Epic of Its Founders. Isis, 107(1):215–216, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Golan:2016:BRA

[1839] Tal Golan. Book review: Ari Barell. Engineer-King: David Ben-Gurion, Science, and Nation Building. Isis, 107(1):216–217, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2016:NC

[1840] Anonymous. Notes on contributors. Isis, 107(1):218–223, March 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2016:Mb

[1841] Anonymous. Masthead. Isis, 107(2):??, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Goddu:2016:LAB

[1842] André Goddu. Ludwik Antoni Birkenmajer and Curtis Wilson on the origin of Nicholas Copernicus’s heliocentrism. Isis, 107(2):225–253, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sponsel:2016:ABH

[1843] Alistair Sponsel. An amphibious being: How maritime surveying reshaped Darwin’s approach to natural history. Isis, 107(2):254–281, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hamlin:2016:PRH

[1844] Christopher Hamlin. The pedagogical roots of the history of science: Revisiting the vision of James Bryant Conant. Isis, 107(2):282–308, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cohen:2016:EIa

[1845] H. Floris Cohen. Editor’s introduction. Isis, 107(2):309–310, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jacobs:2016:SHM

[1846] Noortje Jacobs. Summary of The History Manifesto. Isis, 107(2):311– 314, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kevles:2016:WMH

[1847] Daniel J. Kevles. What’s manifest in the history of SciTech: Reflections on The History Manifesto. Isis, 107(2):315–323, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chemla:2016:RHM

[1848] Karine Chemla. Reading The History Manifesto as a historian of mathematics in ancient china. Isis, 107(2):324–333, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Suarez-Diaz:2016:HMR

[1849] Edna Suárez-Díaz. The History Manifesto as read from Latin America. Isis, 107(2):334–335, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dun:2016:SQA

[1850] Liu Dun. Sima Qian’s ambition and Needham’s vision: a comment on The History Manifesto from China’s perspective. Isis, 107(2):336–337, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Romano:2016:HMH

[1851] Antonella Romano. The History Manifesto, history of science, and big narratives: Some pending questions. Isis, 107(2):338–340, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gaukroger:2016:UOH

[1852] Stephen Gaukroger. Undercontextualization and overcontextualization in the history of science. Isis, 107(2):340–342, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Soderqvist:2016:MP

[1853] Thomas Söderqvist. The muse(um) is political. Isis, 107(2):342–344, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Maienschein:2016:TIN

[1854] Jane Maienschein. Time, impact, and the need for digital history and philosophy of science. Isis, 107(2):344–345, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Flis:2016:DHT

[1855] Ivan Flis, Evina Steinová, and Paul Wouters. Digital humanities are a two-way street. Isis, 107(2):346–348, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Oreskes:2016:LMH

[1856] Naomi Oreskes. Let’s make history more welcoming. Isis, 107(2):348– 350, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heilbron:2016:HFR

[1857] J. L. Heilbron. Are historians fit to rule? Isis, 107(2):350–352, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Armitage:2016:LLD

[1858] David Armitage and Jo Guldi. Longing for the Longue Durée. Isis, 107(2):353–357, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2016:HSS

[1859] Anonymous. History of Science Society annual meeting, 2015: HSS prizes for 2015. Isis, 107(2):358–364, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hunter:2016:EEM

[1860] Michael Hunter. Editing early modern scientific correspondence: The way forward. Isis, 107(2):365–372, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Timmermann:2016:BRR

[1861] Carsten Timmermann. Book review: Robert Jütte, ed. Medical Pluralism: Past — Present — Future. Isis, 107(2):373–374, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jullien:2016:BRP

[1862] Vincent Jullien. Book review: A. Papadopoulos; G. Théret, eds. La théorie des lignes parallèles de Johann Heinrich Lambert. Isis, 107(2): 374–375, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Casado:2016:BRS

[1863] Santos Casado. Book review: Sebastian Vincent Grevsmühl. La Terre vue d’en haut: L’invention de l’environnement global. Isis, 107(2):375– 376, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Condrau:2016:BRJ

[1864] Flurin Condrau. Book review: Joanna Bourke. The Story of Pain: From Prayer to Painkillers; Keith Wailoo. Pain: a Political History. Isis, 107 (2):377–379, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Szulakowska:2016:BRR

[1865] Urszula Szulakowska. Book review: Rafał T. Prinke. Zwodniczy ogród błedów: Piśmiennictwo alchemiczne do końca XVIII wieku [The Deceptive Garden of Errors: Alchemical Writings until the End of the Eighteenth Century]. Isis, 107(2):379–380, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brandt:2016:BRA

[1866] Reinhard Brandt. Book review: Alix Cohen, ed. Kant’s Lectures on Anthropology. Isis, 107(2):380–382, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tobies:2016:BRM

[1867] Renate Tobies. Book review: Magdolna Hargittai. Women Scientists: Reflections, Challenges, and Breaking Boundaries; Paola Govoni; Zelda Alice Franceschi, eds. Writing about Lives in Science: (Auto)Biography, Gender, and Genre. Isis, 107(2):382–383, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dijksterhuis:2016:BRV

[1868] Fokko Jan Dijksterhuis. Book review: Vincenzo De Risi, ed. Mathematizing Space: The Objects of Geometry from Antiquity to the Early Modern Age. Isis, 107(2):383–384, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bensaude-Vincent:2016:BRP

[1869] Bernadette Bensaude-Vincent. Book review: Peter J. T. Morris. The Matter Factory: a History of the Chemistry Laboratory. Isis, 107(2): 384–386, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wilson:2016:BRJ

[1870] Malcolm Wilson. Book review: Jean De Groot. Aristotle’s Empiricism: Experience and Mechanics in the Fourth Century B.C. Isis, 107(2):386– 387, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stevens:2016:BRA

[1871] Wesley M. Stevens. Book review: Alfred Lohr. Der Computus Gerlandi: Edition, übersetzung, und Erlaüterung. Isis, 107(2):387–388, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Whitney:2016:BRA

[1872] Elspeth Whitney. Book review: Adam Lucas. Ecclesiastical Lordship, Seigneurial Power, and the Commercialization of Milling in Medieval England. Isis, 107(2):388–389, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Carolino:2016:BRB

[1873] Luís Miguel Carolino. Book review: Bernardino Telesio. Sobre los cometas y la Vía Láctea/De cometis et lacteo circulo. Isis, 107(2):390– 391, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Spary:2016:BRC

[1874] Emma Spary. Book review: Corinna Wagner. Pathological Bodies: Medicine and Political Culture. Isis, 107(2):391–392, June 2016. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanNetten:2016:BRT

[1875] Djoeke van Netten. Book review: Timothy Brook. Mr. Selden’s Map of China: The Spice Trade, a Lost Chart, and the South China Sea. Isis, 107(2):392–393, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mandelbrote:2016:BRZ

[1876] Scott Mandelbrote. Book review: Zvi Biener; Eric Schliesser, eds. Newton and Empiricism. Isis, 107(2):393–394, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Aranda:2016:BRV

[1877] Marcelo Aranda. Book review: Víctor Navarro Brotons. Disciplinas, saberes y prácticas: Filosofía natural, matemáticas y astronomía en la sociedad española de la época moderna. Isis, 107(2):395–396, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schuyler:2016:BRM

[1878] Alfred E. Schuyler. Book review: Matthias Schönhofer. Letters from an American Botanist: The Correspondences of Gotthilf Heinrich Ernst Mühlenberg (1753–1815). Isis, 107(2):396–397, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Morus:2016:BRB

[1879] Iwan Rhys Morus. Book review: Ben Russell. James Watt: Making the World Anew. Isis, 107(2):397–398, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Andrault:2016:BRR

[1880] Raphaële Andrault. Book review: Roger Ariew. Descartes and the First Cartesians; Alexander X. Douglas. Spinoza and Dutch Cartesianism: Philosophy and Theology. Isis, 107(2):398–399, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cram:2016:BRW

[1881] David Cram. Book review: Wenchao Li, ed. Einheit der Vernunft und Vielfalt der Sprachen: Beitraäge zu Leibniz’ Sprachforschung und Zeichentheorie. Isis, 107(2):400–401, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guerrini:2016:BRPa

[1882] Anita Guerrini. Book review: Pierre-Yves Lacour. La République naturaliste: Collections d’histoire naturelle et Révolution française (1789– 1804). Isis, 107(2):401–402, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rodrigues:2016:BRB

[1883] Ana Duarte Rodrigues. Book review: Bjørn Okholm Skaarup. Anatomy and Anatomists in Early Modern Spain; Enrique Fernández. Anxieties of Interiority and Dissection in Early Modern Spain. Isis, 107(2):402– 404, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rankin:2016:BRM

[1884] Alisha Rankin. Book review: Meredith K. Ray. Daughters of Alchemy: Women and Scientific Culture in Early Modern Italy. Isis, 107(2):404– 405, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Margocsy:2016:BRB

[1885] Dániel Margócsy. Book review: Benjamin Schmidt. Inventing Exoticism: Geography, Globalism, and Europe’s Early Modern World. Isis, 107(2):

405–407, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Golinski:2016:BRH

[1886] Jan Golinski. Book review: Hjalmar Fors. The Limits of Matter: Chemistry, Mining, and Enlightenment. Isis, 107(2):407–408, June 2016. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lynch:2016:BRJ

[1887] John M. Lynch. Book review: Jeffrey P. Moran. American Genesis: The Evolution Controversies from Scopes to Creation Science. Isis, 107(2): 408–409, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Zetti:2016:BRC

[1888] Daniela Zetti. Book review: Colin B. Burke. Information and Intrigue: From Index Cards to Dewey Decimals to Alger Hiss. Isis, 107(2):409– 411, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ochs:2016:BRE

[1889] Kathleen Ochs. Book review: Eric C. Nystrom. Seeing Underground: Maps, Models, and Mining Engineering in America. Isis, 107(2):411– 412, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Banerjee:2016:BRJ

[1890] Somaditya Banerjee. Book review: Joydeep Sen. Astronomy in India, 1784–1876. Isis, 107(2):412–414, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schubring:2016:BRP

[1891] Gert Schubring. Book review: Pieter Josephson; Thomas Karlsohn; Johan östling, eds. The Humboldtian Tradition: Origins and Legacies. Isis, 107(2):414–415, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Moynahan:2016:BRM

[1892] Gregory Moynahan. Book review: Michael Kaasch; Joachim Kaasch, eds. Ordnung — Organisation — Organismus: Beitrage zur 20. Jahrestagung der DGGTB in Bonn 2011. Isis, 107(2):416–417, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mauck:2016:BRG

[1893] Aaron Pascal Mauck. Book review: Gerald N. Grob. Aging Bones: a Short History of Osteoporosis. Isis, 107(2):417–418, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pearce:2016:BRN

[1894] Trevor Pearce. Book review: Naomi Beck. La gauche évolutionniste: Spencer et ses lecteurs en France et en Italie. Isis, 107(2):418–419, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ash:2016:BRP

[1895] Mitchell G. Ash. Book review: Philip Ball. Serving the Reich: The Struggle for the Soul of Physics under Hitler. Isis, 107(2):419–420, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Deichmann:2016:BRR

[1896] Ute Deichmann. Book review: Robert Wald Sussman. The Myth of Race: The Troubling Persistence of an Unscientific Idea; Michael Yudell. Race Unmasked: Biology and Race in the Twentieth Century. Isis, 107(2): 421–424, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blum:2016:BRA

[1897] Alexander Blum. Book review: Amit Hagar. Discrete or Continuous? The Quest for Fundamental Length in Modern Physics. Isis, 107(2):424– 425, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hornby:2016:BRC

[1898] Louise Hornby. Book review: Christina Walter. Optical Impersonality: Science, Images, and Literary Modernism. Isis, 107(2):426–427, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sauer:2016:BRP

[1899] Tilman Sauer. Book review: Paul Halpern. Einstein’s Dice and Schrödinger’s Cat: How Two Great Minds Battled Quantum Randomness to Create a Unified Theory of Physics. Isis, 107(2):427–428, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bradley:2016:BRA

[1900] Ben Bradley. Book review: Alexis Harley. Autobiologies: Charles Darwin and the Natural History of the Self. Isis, 107(2):428–429, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dear:2016:BRM

[1901] Peter Dear. Book review: Matthew Stanley. Huxley’s Church and Maxwell’s Demon: From Theistic Science to Naturalistic Science. Isis, 107(2):429–431, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reed:2016:BRL

[1902] Cameron Reed. Book review: Luis A. Campos. Radium and the Secret of Life. Isis, 107(2):431–432, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bassi:2016:BRD

[1903] Joe Bassi. Book review: David Meerman Scott; Richard Jurek. Marketing the Moon: The Selling of the Apollo Lunar Program. Isis, 107(2):432– 433, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hegarty:2016:BRD

[1904] Peter Hegarty. Book review: Donna J. Drucker. The Machines of Sex Research: Technology and the Politics of Identity, 1945–1985. Isis, 107 (2):433–435, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Abma:2016:BRR

[1905] Ruud Abma. Book review: Roger E. Backhouse; Philippe Fontaine, eds. A Historiography of the Modern Social Sciences. Isis, 107(2):435– 436, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tatarchenko:2016:BRG

[1906] Ksenia Tatarchenko. Book review: Gerard Alberts; Ruth Oldenziel, eds. Hacking Europe: From Computing Cultures to Demoscenes. Isis, 107(2): 436–437, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roque:2016:BRA

[1907] Xavier Roqué. Book review: Angelo Baracca; Jürgen Renn; Helge Wendt, eds. The History of Physics in Cuba. Isis, 107(2):438–439, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harper:2016:BRN

[1908] Kristine C. Harper. Book review: Naomi Oreskes; John Krige, eds. Science and Technology in the Global Cold War. Isis, 107(2):439–440, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Coopersmith:2016:BRS

[1909] Jonathan Coopersmith. Book review: Slava Gerovitch. Soviet Space Mythologies: Public Images, Private Memories, and the Making of a Cultural Identity; Slava Gerovitch. Voices of the Soviet Space Program: Cosmonauts, Soldiers, and Engineers Who Took the USSR into Space; Doug Millard, ed. Cosmonauts: Birth of the Space Age. Isis, 107(2):440– 442, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brinkman:2016:BRN

[1910] Paul D. Brinkman. Book review: Niles Eldredge. Eternal Ephemera: Adaptation and the Origin of Species from the Nineteenth Century through Punctuated Equilibria and Beyond. Isis, 107(2):442–443, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Podolsky:2016:BRJ

[1911] Scott Podolsky. Book review: Jean-Paul Gaudillière; Ulrike Thoms, eds. The Development of Scientific Marketing in the Twentieth Century. Isis, 107(2):443–444, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kjeldsen:2016:BRC

[1912] Tinne Hoff Kjeldsen. Book review: Christopher J. Phillips. The New Math: a Political History. Isis, 107(2):445–446, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wuthrich:2016:BRO

[1913] Adrian Wüthrich. Book review: Olival Freire, Jr. The Quantum Dissidents: Rebuilding the Foundations of Quantum Mechanics (1950–1990). Isis, 107(2):446–447, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Scheffler:2016:BRD

[1914] Robin Wolfe Scheffler. Book review: Doogab Yi. The Recombinant University: Genetic Engineering and the Emergence of Stanford Biotechnology. Isis, 107(2):447–448, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2016:Mc

[1915] Anonymous. Masthead. Isis, 107(3):??, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hicks:2016:VSP

[1916] Daniel J. Hicks and Thomas A. Stapleford. The virtues of scientific practice: MacIntyre, virtue ethics, and the historiography of science. Isis, 107(3):449–472, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fors:2016:MMC

[1917] Hjalmar Fors. Medicine and the making of a city: Spaces of pharmacy and scholarly medicine in seventeenth-century Stockholm. Isis, 107(3): 473–494, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ratcliff:2016:EIC

[1918] Jessica Ratcliff. The East India Company, the Company’s museum, and the political economy of natural history in the early nineteenth century. Isis, 107(3):495–517, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bruno:2016:EMA

[1919] Andy Bruno. A Eurasian mineralogy: Aleksandr Fersman’s conception of the natural world. Isis, 107(3):518–539, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cohen:2016:EIb

[1920] H. Floris Cohen. Editor’s introduction. Isis, 107(3):??, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nicolaidis:2016:SOC

[1921] Efthymios Nicolaidis, Eudoxie Delli, Nikolaos Livanos, Kostas Tampakis, and George Vlahakis. Science and orthodox Christianity: an overview. Isis, 107(3):542–566, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Luthy:2016:TSO

[1922] Christoph Lüthy. Theology and science in the orthodox world: Some doubts from a Latin perspective. Isis, 107(3):567–572, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hall:2016:SRO

[1923] Karl Hall and Dimitri Bayuk. Science and Russian Orthodox scholarship. Isis, 107(3):573–578, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Morrison:2016:RSE

[1924] Robert Morrison. Religion and science in the Eastern Mediterranean. Isis, 107(3):579–582, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rabkin:2016:TOS

[1925] Yakov M. Rabkin. Two orthodoxies and science: Comparative reflections. Isis, 107(3):583–586, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harrison:2016:SEO

[1926] Peter Harrison. Science, Eastern Orthodoxy, and Protestantism. Isis, 107 (3):587–591, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brooke:2016:SEO

[1927] John Hedley Brooke and Ronald L. Numbers. Science, Eastern Orthodoxy, and world religions. Isis, 107(3):592–596, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wood:2016:EVT

[1928] Roger J. Wood. Eloge: Vítêzslav Orel 1926–2015. Isis, 107(3):597–600, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Westman:2016:CR

[1929] Robert S. Westman. A Copernican renaissance? Isis, 107(3):601–607, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Casper:2016:BRA

[1930] Stephen T. Casper. Book review: Andrew Scull. Madness in Civilization: a Cultural History of Insanity from the Bible to Freud, from the Madhouse to Modern Medicine. Isis, 107(3):608–609, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Finnegan:2016:BRG

[1931] Diarmid A. Finnegan. Book review: Greg Goodale. The Rhetorical Invention of Man: a History of Distinguishing Humans from Other Animals. Isis, 107(3):609–610, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Arens:2016:BRR

[1932] Katherine Arens. Book review: Rens Bod; Jaap Maat; Thijs Weststeijn, eds. The Making of the Humanities, Volume 3: The Modern Humanities. Isis, 107(3):611–612, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jonsson:2016:BRA

[1933] Fredrik Albritton Jonsson. Book review: Andrew C. Isenberg, ed. The Oxford Handbook of Environmental History. Isis, 107(3):612–613, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Laerke:2016:BRP

[1934] Mogens Lærke. Book review: Peter Harrison. The Territories of Science and Religion. Isis, 107(3):613–614, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McCrea:2016:BRM

[1935] Heather L. McCrea. Book review: Marcos Cueto; Steven Palmer. Medicine and Public Health in Latin America: a History. Isis, 107(3): 614–616, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vanagt:2016:BRM

[1936] Katrien Vanagt. Book review: A. Mark Smith. From Sight to Light: The Passage from Ancient to Modern Optics. Isis, 107(3):616–617, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Saletta:2016:BRA

[1937] Morgan Saletta. Book review: Anne S. Dowd; Susan Milbrath, eds. Cosmology, Calendars, and Horizon-Based Astronomy in Ancient Mesoamerica; Jeffrey L. Cooley. Poetic Astronomy in the Ancient Near East: The Reflexes of Celestial Science in Ancient Mesopotamian, Ugaritic, and Israelite Narrative. Isis, 107(3):617–620, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Totelin:2016:BRI

[1938] Laurence Totelin. Book review: Ido Israelowich. Patients and Healers in the High Roman Empire. Isis, 107(3):620–621, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Eamon:2016:BRA

[1939] William Eamon. Book review: Antonio Sánchez. La espada, la cruz y el Padrón: Soberanía, fe y representación cartográfica en el mundo ibérico bajo la Monarquía Hispánica, 1503–1598. Isis, 107(3):621–622, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Albertson:2016:BRT

[1940] David Albertson. Book review: Thierry of Chartres. The Commentary on the De arithmetica of Boethius. Isis, 107(3):622–624, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kursell:2016:BRE

[1941] Julia Kursell. Book review: Erwin Hiebert. The Helmholtz Legacy in Physiological Acoustics. Isis, 107(3):624–625, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bilak:2016:BRS

[1942] Donna Bilak. Book review: Sven Dupré, ed. Laboratories of Art: Alchemy and Art Technology from Antiquity to the Eighteenth Century; Sven Dupré; Dedo von Kerssenbrock-Krosigk; Beat Wismer, eds. Art and Alchemy: The Mystery of Transformation. Isis, 107(3):625–627, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gambaro:2016:BRC

[1943] Ivana Gambaro. Book review: Christopher M. Graney. Setting Aside All Authority: Giovanni Battista Riccioli and the Science against Copernicus in the Age of Galileo. Isis, 107(3):627–629, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kardel:2016:BRT

[1944] Troels Kardel. Book review: Thomas Bartholin. The Anatomy House in Copenhagen. Isis, 107(3):629–630, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hsia:2016:BRM

[1945] Ronnie Po chia Hsia. Book review: Mark A. Waddell. Jesuit Science and the End of Nature’s Secrets. Isis, 107(3):630–631, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kraye:2016:BRA

[1946] Jill Kraye. Book review: Andrew D. Berns. The Bible and Natural Philosophy in Renaissance Italy: Jewish and Christian Physicians in Search of Truth. Isis, 107(3):631–632, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Trambaiolo:2016:BRF

[1947] Daniel Trambaiolo. Book review: Federico Marcon. The Knowledge of Nature and the Nature of Knowledge in Early Modern Japan. Isis, 107 (3):632–634, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Verwaal:2016:BRR

[1948] Ruben E. Verwaal. Book review: Robert G. W. Anderson, ed. Cradle of Chemistry: The Early Years of Chemistry at the University of Edinburgh. Isis, 107(3):634–635, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Davies:2016:BRA

[1949] Surekha Davies. Book review: Anita Guerrini. The Courtiers’ Anatomists: Animals and Humans in Louis XIV’s Paris. Isis, 107(3): 635–636, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Levitt:2016:BRM

[1950] Theresa Levitt. Book review: Melvyn C. Usselman. Pure Intelligence: The Life of William Hyde Wollaston. Isis, 107(3):637–638, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ossa-Richardson:2016:BRS

[1951] Anthony Ossa-Richardson. Book review: Sarah Hutton. British Philosophy in the Seventeenth Century. Isis, 107(3):638–639, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kilburn:2016:BRW

[1952] Kevin J. Kilburn. Book review: William Huggins. Selected Correspondence of William Huggins. Isis, 107(3):639–640, September 2016. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hollestelle:2016:BRG

[1953] Marijn Hollestelle. Book review: Gary Patterson. A Prehistory of Polymer Science; Gary Patterson. Polymer Science from 1935–1953: Consolidating the Paradigm. Isis, 107(3):640–642, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cat:2016:BRP

[1954] Jordi Cat. Book reviews: Peter Achinstein. Evidence and Method: Scientific Strategies of Isaac Newton and James Clerk Maxwell; Raymond Flood; Mark McCartney; Andrew Whitaker, eds. James Clerk Maxwell: Perspectives on His Life and Work. Isis, 107(3):642–644, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weiss:2016:BRA

[1955] Martin Weiss. Book review: Anke te Heesen; Margarete Vöhringer, eds. Wissenschaft im Museum: Ausstellung im Labor. Isis, 107(3):644–646, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Treitel:2016:BRE

[1956] Corinna Treitel. Book review: Egil Asprem. The Problem of Disenchantment: Scientific Naturalism and Esoteric Discourse, 1900–1939. Isis, 107 (3):646–647, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Funke:2016:BRK

[1957] Jana Funke. Book review: Kathleen Frederickson. The Ploy of Instinct: Victorian Sciences of Nature and Sexuality in Liberal Governance. Isis, 107(3):647–648, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Smocovitis:2016:BRR

[1958] Vassiliki Betty Smocovitis. Book review: Retha Edens-Meier; Peter Bernhardt, eds. Darwin’s Orchids: Then and Now. Isis, 107(3):648–650, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harwood:2016:BRM

[1959] Jonathan Harwood. Book review: Margaret E. Derry. Masterminding Nature: The Breeding of Animals, 1750–2010. Isis, 107(3):650–651, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Eckert:2016:BRM

[1960] Michael Eckert. Book review: Martina Kölbl-Ebert. From Local Patriotism to a Planetary Perspective: Impact Crater Research in Germany, 1930s–1970s. Isis, 107(3):651–652, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wang:2016:BRL

[1961] Zuoyue Wang. Book review: Lawrence R. Sullivan; Nancy Y. Liu. Historical Dictionary of Science and Technology in Modern China. Isis, 107 (3):652–654, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Champion:2016:BRD

[1962] Timothy Champion. Book review: Dorothy U. Seyler. The Obelisk and the Englishman: The Pioneering Discoveries of Egyptologist William Bankes. Isis, 107(3):654–655, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

ZaitsevaBaum:2016:BRM

[1963] Elena Zaitseva (Baum). Book review: Mary R. S. Creese (with Thomas M. Creese). Ladies in the Laboratory IV: Imperial Russia’s Women in Science, 1800–1900: a Survey of Their Contributions to Research. Isis, 107(3):655–656, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anderson:2016:BRA

[1964] Stuart Anderson. Book review: Anthony C. Cartwright. The British Pharmacopoeia, 1864 to 2014: Medicines, International Standards and the State. Isis, 107(3):656–657, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Zarimis:2016:BRC

[1965] Maria Zarimis. Book review: Curtis Johnson. Darwin’s Dice: The Idea of Chance in the Thought of Charles Darwin. Isis, 107(3):658–659, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gradmann:2016:BRC

[1966] Christoph Gradmann. Book review: Christian W. McMillen. Discovering Tuberculosis: a Global History, 1900 to the Present. Isis, 107(3):659–660, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gouyon:2016:BRO

[1967] Jean-Baptiste Gouyon. Book review: Oliver Gaycken. Devices of Curiosity: Early Cinema and Popular Science. Isis, 107(3):660–661, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Haalboom:2016:BRA

[1968] Floor Haalboom. Book review: Anne Hardy. Salmonella Infections, Networks of Knowledge, and Public Health in Britain, 1880–1975. Isis, 107 (3):661–663, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Clarke:2016:BRM

[1969] Imogen Clarke. Book review: Melinda Baldwin. Making “Nature”: The History of a Scientific Journal. Isis, 107(3):663–664, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Campos:2016:BRE

[1970] Luis Campos. Book review: Eva Hemmungs Wirtén. Making Marie Curie: Intellectual Property and Celebrity Culture in an Age of Information. Isis, 107(3):664–665, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rayward:2016:BRM

[1971] W. Boyd Rayward. Book review: Michael D. Gordin. Scientific Babel: How Science Was Done Before and After Global English. Isis, 107(3): 665–666, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wittje:2016:BRS

[1972] Roland Wittje. Book review: Silke Fengler; Carola Sachse, eds. Kernforschung in Österreich: Wandlungen eines interdisziplinären Forschungsfeldes, 1900–1978; Silke Fengler. Kerne, Kooperation und Konkurrenz: Kernforschung in Österreich im internationalen Kontext (1900–1950). Isis, 107(3):666–668, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mauskopf:2016:BRM

[1973] Seymour Mauskopf. Book review: Masanori Kaji; Helge Kragh; Gábor Palló, eds. Early Responses to the Periodic System. Isis, 107(3):669–670, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Meneghello:2016:BRW

[1974] Laura Meneghello. Book review: William R. Woodward. Hermann Lotze: an Intellectual Biography. Isis, 107(3):670–671, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Borrelli:2016:BRF

[1975] Arianna Borrelli. Book review: Finn Aaserud; Helge Kragh, eds. One Hundred Years of the Bohr Atom: Proceedings from a Conference. Isis, 107(3):671–672, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stoff:2016:BRG

[1976] Heiko Stoff. Book review: Greg Eghigian. The Corrigible and the Incorrigible: Science, Medicine, and the Convict in Twentieth-Century Germany. Isis, 107(3):672–673, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fuller:2016:BRH

[1977] Steve Fuller. Book review: Harvey J. Graff. Undisciplining Knowledge: Interdisciplinarity in the Twentieth Century. Isis, 107(3):674–675, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mirmalek:2016:BRJ

[1978] Zara Mirmalek. Book review: Janet Vertesi. Seeing Like a Rover: How Robots, Teams, and Images Craft Knowledge of Mars. Isis, 107(3):675– 676, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stingl:2016:BRM

[1979] Alexander I. Stingl. Book review: M. Joan Dawson. Paul Lauterbur and the Invention of MRI. Isis, 107(3):676–677, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Crease:2016:BRH

[1980] Robert P. Crease. Book review: Henning Schmidgen. Bruno Latour in Pieces: an Intellectual Biography. Isis, 107(3):677–679, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brain:2016:BRE

[1981] Stephen Brain. Book review: Eduard I. Kolchinskii. Tak vspominaetsja … [So I Remember …]. Isis, 107(3):679–680, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Decoteau:2016:BRP

[1982] Claire Laurier Decoteau. Book review: Peter Piot. AIDS between Science and Politics. Isis, 107(3):680–681, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Camprubi:2016:BRA

[1983] Lino Camprubí. Book review: Amparo Gómez; Antonio Fco. Canales; Brian Balmer, eds. Science Policies and Twentieth-Century Dictatorships: Spain, Italy, and Argentina. Isis, 107(3):681–682, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bourgeois:2016:BRC

[1984] Suzanne Bourgeois. Book review: Charlotte DeCroes Jacobs. Jonas Salk: a Life. Isis, 107(3):683–684, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Richmond:2016:BRK

[1985] Marsha L. Richmond. Book review: Kersten T. Hall. The Man in the Monkeynut Coat: William Astbury and the Forgotten Road to the Double-Helix; Matthew Cobb. Life’s Greatest Secret: The Race to Crack the Genetic Code. Isis, 107(3):684–685, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shcheglov:2016:AAC

[1986] Dmitry A. Shcheglov. The accuracy of ancient cartography reassessed: The longitude error in Ptolemy’s map. Isis, 107(4):687–706, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wang:2016:FBM

[1987] Xiaona Wang. Francis Bacon and magnetical cosmology. Isis, 107(4): 707–721, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cowles:2016:AMW

[1988] Henry M. Cowles. The age of methods: William Whewell, Charles Peirce, and scientific kinds. Isis, 107(4):722–737, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Alagona:2016:SCC

[1989] Peter S. Alagona. Species complex: Classification and conservation in American environmental history. Isis, 107(4):738–761, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanLunteren:2016:CCM

[1990] Frans van Lunteren. Clocks to computers: a machine-based “big picture” of the history of modern science. Isis, 107(4):762–776, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bensaude-Vincent:2016:MMT

[1991] Bernadette Bensaude-Vincent. Machines as “mental tools”. Isis, 107 (4):777–780, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wise:2016:A

[1992] M. Norton Wise. Agency. Isis, 107(4):781–784, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fara:2016:OL

[1993] Patricia Fara. Object lessons. Isis, 107(4):785–788, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chang:2016:PVP

[1994] Hasok Chang. Pluralism versus periodization. Isis, 107(4):789–792, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kucuk:2016:DCB

[1995] B. Harun Küçük. Darnton’s cats, Bacon’s rifle, and History of Science 101. Isis, 107(4):793–795, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeBont:2016:MMS

[1996] Raf De Bont. Machine metaphors: Some reflections. Isis, 107(4):796– 799, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanLunteren:2016:CCS

[1997] Frans van Lunteren. Clocks to computers: Some clarifications. Isis, 107 (4):800–804, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gavroglu:2016:EFP

[1998] Kostas Gavroglu. Eloge: Faidra Papanelopoulou (1978–2016). Isis, 107 (4):805–808, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Henry:2016:SRF

[1999] John Henry. The scientific revolution: Five books about it. Isis, 107 (4):809–817, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hamlin:2016:BRM

[2000] Christopher Hamlin. Book review: Melvin Santer, Confronting Contagion: Our Evolving Understanding of Disease. Isis, 107(4):818–819, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brazelton:2016:BRH

[2001] Mary Augusta Brazelton. Book review: Howard Chiang, ed. Historical Epistemology and the Making of Modern Chinese Medicine. Isis, 107 (4):819–820, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reinhart:2016:BRH

[2002] Katherine M. Reinhart. Book review: Horst Bredekamp; Vera Dünkel; Birgit Schneider, eds. The Technical Image: a History of Styles in Scientific Imagery. Isis, 107(4):820–821, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Henry:2016:BRO

[2003] John Henry. Book review: Owen Gingerich. God’s Planet. Isis, 107(4): 822, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Channell:2016:BRR

[2004] David Channell. Book review: Raffaele Pisano, ed. A Bridge between Conceptual Frameworks: Sciences, Society, and Technology Studies. Isis, 107(4):823–824, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Levine:2016:BRJ

[2005] Deborah Levine. Book review: John C. Burnham. Health Care in America: a History. Isis, 107(4):824–825, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lehner:2016:BRO

[2006] Christoph Lehner. Book review: Olivier Darrigol. Physics and Necessity: Rationalist Pursuits from the Cartesian Past to the Quantum Present. Isis, 107(4):825–826, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Carluccio:2016:BRB

[2007] Dana Carluccio. Book review: Barbara Larson; Sabine Flach, eds. Darwin and Theories of Aesthetics and Cultural History. Isis, 107(4):826– 827, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vandendriessche:2016:BRR

[2008] Joris Vandendriessche. Book review: Robert Halleux; Geneviève Xhayet; Pascal Pirot; Rik Raedschelders; Jan Vandersmissen. Tant qu’il y aura des chercheurs: Science et politique en Belgique de 1772 à 2015. Isis, 107 (4):827–829, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schenkel:2016:BRP

[2009] Elmar Schenkel. Book review: Philip Ball. Invisible: The Dangerous Allure of the Unseen. Isis, 107(4):829–830, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Zhao:2016:BRR

[2010] Wenrui Zhao. Book review: E. R. Truitt. Medieval Robots: Mechanism, Magic, Nature, and Art. Isis, 107(4):830–831, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Keller:2016:BRK

[2011] Vera Keller. Book review: Katherine Eggert. Disknowledge: Literature, Alchemy, and the End of Humanism in Renaissance England. Isis, 107 (4):831–832, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vartija:2016:BRJ

[2012] Devin Vartija. Book review: Justin E. H. Smith. Nature, Human Nature, and Human Difference: Race in Early Modern Philosophy. Isis, 107(4): 832–833, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hunt:2016:BRJ

[2013] John Dixon Hunt. Book review: Juliette Ferdinand, ed. From Art to Science: Experiencing Nature in the European Garden, 1500–1700. Isis, 107 (4):834–835, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guerrini:2016:BRPb

[2014] Anita Guerrini. Book review: Pascal Duris. Quelle révolution scientifique? Les sciences de la vie dans la querelle des Anciens et des Modernes (XVIe–XVIIIe siècles). Isis, 107(4):835–836, December 2016. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Neri:2016:BRB

[2015] Janice Neri. Book review: Brent Elliott; Luigi Guerrini; David Pegler. Flora: Federico Cesi’s Botanical Manuscripts. Isis, 107(4):836–837, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanMiert:2016:BRD

[2016] Dirk van Miert. Book review: Dmitri Levitin. Ancient Wisdom in the Age of the New Science: Histories of Philosophy in England, c. 1640–1700. Isis, 107(4):837–839, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dupre:2016:BRR

[2017] Sven Dupré. Book review: Richard Yeo. Notebooks, English Virtuosi, and Early Modern Science. Isis, 107(4):839–840, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weisser:2016:BRS

[2018] Olivia Weisser. Book review: Sabine Arnaud. On Hysteria: The Invention of a Medical Category between 1670 and 1820. Isis, 107(4):840–841, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cobb:2016:BRR

[2019] Matthew Cobb. Book review: Raymond Stephanson; Darren N. Wagner, eds. The Secrets of Generation: Reproduction in the Long Eighteenth Century. Isis, 107(4):841–842, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Durr:2016:BRK

[2020] Renate Dürr. Book review: Kelly Joan Whitmer. The Halle Orphanage as Scientific Community: Observation, Eclecticism, and Pietism in the Early Enlightenment. Isis, 107(4):842–843, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Whitmer:2016:BRK

[2021] Kelly J. Whitmer. Book review: Klaas van Berkel; Bart Ramakers, eds. Petrus Camper in Context: Science, the Arts, and Society in the Eighteenth-Century Dutch Republic. Isis, 107(4):844–845, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Home:2016:BRR

[2022] R. W. Home. Book review: Russell McCormmach. The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities. Isis, 107(4):845–846, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Papadopoulos:2016:BRC

[2023] Athanase Papadopoulos. Book review: Christian Gilain; Alexandre Guilbaud, eds. Sciences mathématiques 1750–1850: Continuités et ruptures. Isis, 107(4):846–848, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fischel:2016:BRH

[2024] Angela Fischel. Book review: Haru Hamanaka. Erkenntnis und Bild: Wissenschaftsgeschichte der Lichtenbergischen Figuren um 1800. Isis, 107(4):848–849, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mijnhardt:2016:BRK

[2025] Wijnand Mijnhardt. Book review: Kathleen Hardesty Doig. From Encyclopédie to Encyclopédie méthodique: Revision and Expansion. Isis, 107 (4):849–850, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Myklebost:2016:BRR

[2026] Kari Aga Myklebost. Book review: Ryan Tucker Jones. Empire of Extinction: Russians and the North Pacific’s Strange Beasts of the Sea, 1741–1867. Isis, 107(4):850–851, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kutzinski:2016:BRU

[2027] Vera M. Kutzinski. Book review: Ursula Klein. Humboldts Preußen: Wissenschaft und Technik im Aufbruch. Isis, 107(4):852–853, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Secord:2016:BRI

[2028] Jim Secord. Book review: Innes M. Keighren; Charles W. J. Withers; Bill Bell. Travels into Print: Exploration, Writing, and Publishing with John Murray, 1773–1859. Isis, 107(4):853–854, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rumberg:2016:BRB

[2029] Antje Rumberg. Book review: Bernard Bolzano; Paul Rusnock; Rolf George. Theory of Science. Isis, 107(4):854–856, December 2016. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Aubin:2016:BRJ

[2030] David Aubin. Book review: James Lequeux. François Arago: a 19th Century French Humanist and Pioneer in Astrophysics. Isis, 107(4): 856–857, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hausman:2016:BRK

[2031] Bernice L. Hausman. Book review: Karen L. Walloch. The Antivaccine Heresy: Jacobson v. Massachusetts and the Troubled History of Compulsory Vaccination in the United States. Isis, 107(4):857–858, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brock:2016:BRT

[2032] William H. Brock. Book review: Tamson Pietsch. Empire of Scholars: Universities, Networks, and the British Academic World, 1850–1939; William C. Lubenow. “Only Connect”: Learned Societies in Nineteenth-Century Britain. Isis, 107(4):858–860, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Olson:2016:BRD

[2033] Richard Olson. Book review: Dana Simmons. Vital Minimum: Need, Science, and Politics in Modern France. Isis, 107(4):860–861, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Palfreyman:2016:BRC

[2034] Harriet Palfreyman. Book review: Carin Berkowitz. Charles Bell and the Anatomy of Reform. Isis, 107(4):861–862, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Voskuhl:2016:BRW

[2035] Adelheid Voskuhl. Book review: Wolfgang König. Der Gelehrte und der Manager: Franz Reuleaux (1829–1905) und Alois Riedler (1850–1936) in Technik, Wissenschaft und Gesellschaft. Isis, 107(4):862–864, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Olszewski:2016:BRW

[2036] Todd M. Olszewski. Book review: W. Bruce Fye. Caring for the Heart: Mayo Clinic and the Rise of Specialization. Isis, 107(4):864–865, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Richards:2016:BRN

[2037] Robert J. Richards. Book review: Nick Hopwood. Haeckel’s Embryos: Images, Evolution, and Fraud. Isis, 107(4):865–867, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jardine:2016:BRR

[2038] Boris Jardine. Book review: Robert Michael Brain. The Pulse of Modernism: Physiological Aesthetics in Fin-de–Siècle Europe. Isis, 107(4): 867–868, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hesketh:2016:BRS

[2039] Ian Hesketh. Book review: Sarah C. Alexander. Victorian Literature and the Physics of the Imponderable. Isis, 107(4):868–869, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gates:2016:BRM

[2040] Barbara T. Gates. Book review: Melanie Keene. Science in Wonderland: The Scientific Fairy Tales of Victorian Britain. Isis, 107(4):869–870, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Powell:2016:BRM

[2041] James Lawrence Powell. Book review: Mott T. Greene. Alfred Wegener: Science, Exploration, and the Theory of Continental Drift. Isis, 107(4): 871–872, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Greenblatt:2016:BRS

[2042] Samuel H. Greenblatt. Book review: Stephen T. Casper. The Neurologists: a History of a Medical Speciality in Modern Britain, c. 1789–2000. Isis, 107(4):872–873, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mayer:2016:BRJ

[2043] Andreas Mayer. Book review: James E. Strick. Wilhelm Reich, Biologist. Isis, 107(4):873–874, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Widmalm:2016:BRA

[2044] Sven Widmalm. Book review: Alex Wright. Cataloging the World: Paul Otlet and the Birth of the Information Age. Isis, 107(4):874–876, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Aaserud:2016:BRM

[2045] Finn Aaserud. Book review: A. M. Glazer; Patience Thomson, eds. Crystal Clear: The Autobiographies of Sir Lawrence and Lady Bragg. Isis, 107(4):876–877, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rynasiewicz:2016:BRH

[2046] Robert Rynasiewicz. Book review: Howard J. Fisher. Maxwell’s “Treatise on Electricity and Magnetism”: The Central Argument. Isis, 107(4):877– 878, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bynum:2016:BRA

[2047] William Bynum. Book review: Anna Greenwood; Harshad Topiwala. Indian Doctors in Kenya, 1895–1940: The Forgotten History. Isis, 107 (4):878–879, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Turda:2016:BRG

[2048] Marius Turda. Book review: Gerald V. O’Brien. Framing the Moron: The Social Construction of Feeble-Mindedness in the American Eugenic Era. Isis, 107(4):879–881, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lalli:2016:BRM

[2049] Roberto Lalli. Book review: Marcia Bartusiak. Black Hole: How an Idea Abandoned by Newtonians, Hated by Einstein, and Gambled on by Hawking Became Loved; Andrew Robinson; Diana Kormos Buchwald. Einstein: a Hundred Years of Relativity. Isis, 107(4):881–883, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Core:2016:BRS

[2050] Rachel Core. Book review: Sean Hsiang-Lin Lei. Neither Donkey nor Horse: Medicine in the Struggle over China’s Modernity. Isis, 107(4): 883–884, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeBont:2016:BRG

[2051] Raf De Bont. Book review: Georgina M. Montgomery. Primates in the Real World: Escaping Primate Folklore and Creating Primate Science. Isis, 107(4):884–885, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Goenner:2016:BRH

[2052] Hubert Goenner. Book review: Helge Kragh. Varying Gravity: Dirac’s Legacy in Cosmology and Geophysics. Isis, 107(4):885–887, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Forsbach:2016:BRM

[2053] Ralf Forsbach. Book review: Matthis Krischel. Urologie und Nationalsozialismus: Eine Studie zu Medizin und Politik als Ressourcen füreinander. Isis, 107(4):887–888, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ankeny:2016:BRM

[2054] Rachel A. Ankeny. Book review: Matthew Smith. Another Person’s Poison: a History of Food Allergy. Isis, 107(4):888–889, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sankaran:2016:BRA

[2055] Neeraja Sankaran. Book review: Anthony R. Rees. The Antibody Molecule: From Antitoxins to Therapeutic Antibodies. Isis, 107(4):889– 890, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dorsey:2016:BRU

[2056] M. Girard Dorsey. Book review: Ulf Schmidt. Secret Science: a Century of Poison Warfare and Human Experiments. Isis, 107(4):890–891, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kessel:2016:BRJ

[2057] Nils Kessel. Book review: Jeremy A. Greene. Generic: The Unbranding of Modern Medicine. Isis, 107(4):892–893, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Backhouse:2016:BRP

[2058] Roger E. Backhouse. Book review: Paul Erickson. The World the Game Theorists Made. Isis, 107(4):893–894, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chemla:2016:BRJ

[2059] Karine Chemla. Book review: Jiri Hudecek. Reviving Ancient Chinese Mathematics: Mathematics, History, and Politics in the Work of Wu Wen–Tsun. Isis, 107(4):894–896, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Laszlo:2016:BRI

[2060] Pierre Laszlo. Book review: István Hargittai; Magdolna Hargittai. Budapest Scientific: a Guidebook; George A. Olah; Thomas Mathew. A Life of Magic Chemistry: Autobiographical Reflections Including Post-Nobel Prize Years and the Methanol Economy. Isis, 107(4):896–898, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Eghigian:2016:BRD

[2061] Greg Eghigian. Book review: Douglas A. Vakoch; Matthew F. Dowd, eds. The Drake Equation: Estimating the Prevalence of Extraterrestrial Life through the Ages. Isis, 107(4):898–899, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rand:2016:BRB

[2062] Lisa Ruth Rand. Book review: Bruce Clarke, ed. Earth, Life, and System: Evolution and Ecology on a Gaian Planet. Isis, 107(4):900–901, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dick:2016:BRJ

[2063] Steven J. Dick. Book review: John M. Logsdon. After Apollo? Richard Nixon and the American Space Program. Isis, 107(4):901–902, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bursten:2016:BRA

[2064] Julia R. Bursten. Book review: Anne Marcovich; Terry Shinn. Toward a New Dimension: Exploring the Nanoscale. Isis, 107(4):902–903, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rasmussen:2016:BRM

[2065] Nicolas Rasmussen. Book review: Myles Jackson. The Genealogy of a Gene: Patents, HIV/AIDS, and Race. Isis, 107(4):903–905, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Larson:2016:BRP

[2066] Allan Larson. Book review: Phillip R. Sloan; Gerald McKenny; Kathleen Eggleson, eds. Darwin in the Twenty-First Century: Nature, Humanity, and God. Isis, 107(4):905–906, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2016:IIV

[2067] Anonymous. Index to isis, volume 107, 2016. Isis, 107(4):907–925, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2016:CIV

[2068] Anonymous. Contents of Isis, volume 107. Isis, 107(4):926–928, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2016:M

[2069] Anonymous. Masthead. Isis, 107(4):??, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2016:CBH

[2070] Anonymous. Current bibliography of the history of science and its cultural influences, 2016. Isis, 107(S1):1–240, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hanson:2017:MFE

[2071] Marta Hanson and Gianna Pomata. Medicinal formulas and experiential knowledge in the seventeenth-century epistemic exchange between China and Europe. Isis, 108(1):1–25, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lehner:2017:MQA

[2072] Christoph Lehner and Helge Wendt. Mechanics in the Querelle des Anciens et des Modernes. Isis, 108(1):26–39, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Swanson:2017:REB

[2073] Kara W. Swanson. Rubbing elbows and blowing smoke: Gender, class, and science in the nineteenth-century Patent Office. Isis, 108(1):40– 61, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kerr:2017:SSI

[2074] Ashley Kerr. From savagery to sovereignty: Identity, politics, and international expositions of Argentine anthropology (1878–1892). Isis, 108(1): 62–81, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Backhouse:2017:RTE

[2075] Roger E. Backhouse and Harro Maas. A road not taken: Economists, historians of science, and the making of the Bowman Report. Isis, 108 (1):82–106, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Achbari:2017:RLA

[2076] Azadeh Achbari. The reviews of Leviathan and the Air-Pump: a Survey. Isis, 108(1):108–116, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hankins:2017:CRR

[2077] Thomas L. Hankins. Comment: Reviewing a review. Isis, 108(1):117– 118, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pestre:2017:CMR

[2078] Dominique Pestre. Comment: The making of a reference book. Isis, 108(1):119–121, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heilbron:2017:CLJ

[2079] J. L. Heilbron. Comment: a last judgment. Isis, 108(1):122–123, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wood:2017:CBV

[2080] Paul Wood. Comment: Behemoth v. the sceptical chymist, revisited. Isis, 108(1):124–126, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pinch:2017:CAP

[2081] Trevor Pinch. Comment: All pumped up about the sociology of scientific knowledge. Isis, 108(1):127–129, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bouterse:2017:CRC

[2082] Jeroen Bouterse. Comment: On rereading a classic. Isis, 108(1):130– 132, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jiang:2017:CTE

[2083] Lijing Jiang. Comment: Taking experiments afresh again. Isis, 108(1): 133–135, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Garcia-Deister:2017:CLB

[2084] Vivette García-Deister. Comment: El Leviathan y la bomba de vacío, an enduring epistemological intervention. Isis, 108(1):136–139, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wellerstein:2017:CEC

[2085] Alex Wellerstein. Comment: The epistemology of civility and the civility of epistemology. Isis, 108(1):140–142, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schaffer:2017:FCA

[2086] Simon Schaffer and Steven Shapin. Final comment: Authors’ response. Isis, 108(1):143–144, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Olesko:2017:EPW

[2087] Kathryn M. Olesko. Eloge: L. Pearce Williams (1927–2015). Isis, 108 (1):145–148, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Martin:2017:WOT

[2088] Joseph D. Martin. Who owns the Twentieth Century? (And is it worth owning?). Isis, 108(1):149–157, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wegener:2017:BRW

[2089] Daan Wegener. Book review: William H. Brock. The History of Chemistry: a Very Short Introduction; John L. Heilbron. Physics: a Short History from Quintessence to Quarks. Isis, 108(1):158–160, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Redding:2017:BRA

[2090] Gordon Redding. Book review: Arun Bala; Prasenjit Duara, eds. The Bright Dark Ages: Comparative and Connective Perspectives. Isis, 108 (1):160–161, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Despeux:2017:BRM

[2091] Catherine Despeux. Book review: Miranda Brown. The Art of Medicine in Early China: The Ancient and Medieval Origins of a Modern Archive. Isis, 108(1):161–163, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stewart:2017:BRY

[2092] Ian G. Stewart. Book review: Yves Gingras. L’impossible dialogue: Sciences et religions. Isis, 108(1):163–164, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kodera:2017:BRM

[2093] Sergius Kodera. Book review: Miranda Anderson. The Renaissance Extended Mind. Isis, 108(1):164–165, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hahn:2017:BRJ

[2094] Andre Michael Hahn. Book review: Jessica Riskin. The Restless Clock: a History of the Centuries-Long Argument over What Makes Living Things Tick. Isis, 108(1):165–167, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hendriksen:2017:BRK

[2095] Marieke M. A. Hendriksen. Book review: Kimberly Anne Coles; Ralph Bauer; Zita Nunes; Carla L. Peterson, eds. The Cultural Politics of Blood, 1500–1900. Isis, 108(1):167–168, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Messner:2017:BRB

[2096] Angelika C. Messner. Book review: Benjamin A. Elman, ed. Antiquarianism, Language, and Medical Philology: From Early Modern to Modern Sino-Japanese Medical Discourses. Isis, 108(1):168–169, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Spiller:2017:BRH

[2097] James Spiller. Book review: Hugh Richard Slotten, ed. The Oxford Encyclopedia of the History of American Science, Medicine, and Technology. Isis, 108(1):169–171, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Thumiger:2017:BRH

[2098] Chiara Thumiger. Book review: Hynek Bartos. Philosophy and Dietetics in the Hippocratic. On Regimen: A Delicate Balance of Health. Isis, 108(1):171–172, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nicolaidis:2017:BRJ

[2099] Efthymios Nicolaidis. Book review: Jian-Liang Lin; Hong-Sen Yan. Decoding the Mechanisms of Antikythera Astronomical Device. Isis, 108(1): 172–173, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Luthy:2017:BRM

[2100] Christoph Lüthy. Book review: Marco Beretta. La rivoluzione culturale di Lucrezio: Filosofia e scienza nell’antica, Roma; Lucrezio; Marco Beretta, eds. De rerum natura: Editio princeps (1472–73). Isis, 108(1): 174–176, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Leaman:2017:BRM

[2101] Oliver Leaman. Book review: Moses Maimonides. Medical Aphorisms, Treatises 16–21: a Parallel Arabic-English Edition. Isis, 108(1):176– 177, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Levy:2017:BRR

[2102] Tony Lévy. Book review: Ruth Glasner. Gersonides: a Portrait of a Fourteenth-Century Philosopher–Scientist. Isis, 108(1):177–178, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Caroti:2017:BRJ

[2103] Stefano Caroti. Book review: John Buridan. Quaestiones super octo libros Physicorum Aristotelis (secundum ultimam lecturam): Libri I–II. Isis, 108(1):179–180, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Berns:2017:BRM

[2104] Andrew D. Berns. Book review: Maud Kozodoy. The Secret Faith of Maestre Honoratus: Profayt Duran and Jewish Identity in Late Medieval Iberia. Isis, 108(1):180–181, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Juste:2017:BRD

[2105] David Juste. Book review: Darin Hayton. The Crown and the Cosmos: Astrology and the Politics of Maximilian I. Isis, 108(1):181–182, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rossetter:2017:BRK

[2106] Thomas Rossetter. Book review: Karen Hunger Parshall; Michael T. Walton; Bruce T. Moran, eds. Bridging Traditions: Alchemy, Chemistry, and Paracelsian Practices in the Early Modern Era. Isis, 108(1):184– 185, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roos:2017:BRP

[2107] Anna Marie Roos. Book review: Pietro Daniel Omodeo, ed.; with Karin Friedrich. Duncan Liddel (1561–1613): Networks of Polymathy and the Northern European Renaissance. Isis, 108(1):185–186, March 2017. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gouk:2017:BRR

[2108] Penelope Gouk. Book review: Rebecca Cypess. Curious and Modern Inventions: Instrumental Music as Discovery in Galileo’s Italy. Isis, 108 (1):186–187, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stewart:2017:BRV

[2109] Larry Stewart. Book review: Vera Keller. Knowledge and the Public Interest, 1575–1725. Isis, 108(1):187–189, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bleichmar:2017:BRJ

[2110] Daniela Bleichmar. Book review: José Ramón Marcaida López. Arte y ciencia en el barroco español: Historia natural, coleccionismo y cultura visual. Isis, 108(1):189–190, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Murphy:2017:BRU

[2111] Hannah Murphy. Book review: Ulinka Rublack. The Astronomer and the Witch: Johannes Kepler’s Fight for His Mother. Isis, 108(1):190– 191, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Canizares-Esguerra:2017:BRM

[2112] Jorge Cañizares-Esguerra. Book review: Matthew James Crawford. The Andean Wonder Drug: Cinchona Bark and Imperial Science in the Spanish Atlantic, 1630–1800. Isis, 108(1):192–193, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pickering:2017:BRS

[2113] Victoria Pickering. Book review: Susannah Gibson. Animal, Vegetable, Mineral? How Eighteenth-Century Science Disrupted the Natural Order. Isis, 108(1):193–194, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gray:2017:BRR

[2114] Jeremy Gray. Book review: Ronald S. Calinger. Leonhard Euler: Mathematical Genius in the Enlightenment; Leonhard Euler. Correspondence. Isis, 108(1):194–197, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Leitch:2017:BRH

[2115] Stephanie Leitch. Book review: Han F. Vermeulen. Before Boas: The Genesis of Ethnography and Ethnology in the German Enlightenment. Isis, 108(1):197–198, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Abou-Nemeh:2017:BRS

[2116] Catherine Abou-Nemeh. Book review: Stephen Gaukroger. The Natural and the Human: Science and the Shaping of Modernity, 1739–1841. Isis, 108(1):198–199, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Finkelstein:2017:BRJ

[2117] Gabriel Finkelstein. Book review: John A. McCarthy; Stephanie M. Hilger; Heather I. Sullivan; Nicholas Saul, eds. The Early History of Embodied Cognition, 1740–1920: The Lebenskraft-Debate and Radical Reality in German Science, Music, and Literature. Isis, 108(1):200– 201, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Miller:2017:BRJ

[2118] David Philip Miller. Book review: Jan Golinski. The Experimental Self: Humphry Davy and the Making of a Man of Science. Isis, 108(1):201– 202, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bonah:2017:BRJ

[2119] Christian Bonah. Book review: Joseph M. Gabriel. Medical Monopoly: Intellectual Property Rights and the Origins of the Modern Pharmaceutical Industry. Isis, 108(1):202–203, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Donnelly:2017:BRP

[2120] James Donnelly. Book review: Peter Reed. Entrepreneurial Ventures in Chemistry: The Muspratts of Liverpool, 1793–1934. Isis, 108(1):203– 204, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pols:2017:BRD

[2121] Hans Pols. Book review: S. D. Lamb. Pathologist of the Mind: Adolf Meyer and the Origins of American Psychiatry. Isis, 108(1):205–206, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Matthews:2017:BRD

[2122] J. Rosser Matthews. Book review: Dan Bouk. How Our Days Became Numbered: Risk and the Rise of the Statistical Individual. Isis, 108(1): 206–207, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rowe:2017:BRS

[2123] David E. Rowe. Book review: Steven Gimbel. Einstein: His Space and Times. Isis, 108(1):207–208, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Staubermann:2017:BRK

[2124] Klaus B. Staubermann. Book review: Klaus Hentschel; Josef Webel, eds. Geschichte und Praxis der Materialforschung an den Beispielen: Materialprüfung und Materialprüfungsanstalt (MPA) Stuttgart, Flüssigkristalle und Bildschirmtechnik sowie Supraleitung. Isis, 108(1):208–209, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roll-Hansen:2017:BRJ

[2125] Nils Roll-Hansen. Book review: Jonathan D. Oldfield; Denis J. B. Shaw. The Development of Russian Environmental Thought: Scientific and Geographical Perspectives on the Natural Environment. Isis, 108(1):209– 211, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Huisman:2017:BRP

[2126] Frank Huisman. Book review: Peter Jan Knegtmans. Geld, IJdelheid en Hormonen: Ernst Laqueur, Hoogleraar en Ondernemer. Isis, 108(1):211– 212, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Davie:2017:BRK

[2127] Grace Davie. Book review: Keith Snedegar. Mission, Science, and Race in South Africa: A. W. Roberts of Lovedale, 1883–1938. Isis, 108(1):212– 214, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anduaga:2017:BRJ

[2128] Aitor Anduaga. Book review: James Rodger Fleming. Inventing Atmospheric Science: Bjerknes, Rossby, Wexler, and the Foundations of Modern Meteorology. Isis, 108(1):214–215, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Williams:2017:BRE

[2129] Alison J. Williams. Book review: Ed Regis. Monsters: The Hindenburg Disaster and the Birth of Pathological Technology. Isis, 108(1):215–216, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rohde:2017:BRR

[2130] Joy Rohde. Book review: Richard G. Olson. Scientism and Technocracy in the Twentieth Century: The Legacy of Scientific Management. Isis, 108(1):217–218, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

MacLeod:2017:BRD

[2131] Roy MacLeod. Book review: David Aubin; Catherine Goldstein, eds. The War of Guns and Mathematics: Mathematical Practices and Communities in France and Its Western Allies around World War I. Isis, 108(1):218–219, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Janssen:2017:BRH

[2132] Michel Janssen. Book review: Helge Kragh. Niels Bohr and the Quantum Atom: The Bohr Model of Atomic Structures, 1913–1925. Isis, 108(1): 219–220, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kohout:2017:BRR

[2133] Amy Kohout. Book review: Robert H. Kargon; Karen Fiss; Morris Low; Arthur P. Molella. World’s Fairs on the Eve of War: Science, Technology, and Modernity, 1937–1942. Isis, 108(1):221–222, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Grausam:2017:BRJ

[2134] Daniel Grausam. Book review: Jonathan Hogg. British Nuclear Culture: Official and Unofficial Narratives in the Long Twentieth Century. Isis, 108(1):222–223, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stoll:2017:BRJ

[2135] Mark Stoll. Book review: Janet Biehl. Ecology or Catastrophe: The Life of Murray Bookchin. Isis, 108(1):223–224, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Oldfield:2017:BRL

[2136] Jonathan Oldfield. Book review: Loren Graham. Lysenko’s Ghost: Epigenetics and Russia. Isis, 108(1):224–225, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mills:2017:BRJ

[2137] Eric L. Mills. Book review: Jason Sean Ridler. Maestro of Science: Omond McKillop Solandt and Government Science in War and Hostile Peace, 1939–1956. Isis, 108(1):225–226, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Herfeld:2017:BRW

[2138] Catherine Herfeld. Book review: William Thomas. Rational Action: The Sciences of Policy in Britain and America, 1940–1960. Isis, 108(1):227– 228, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lykknes:2017:BRL

[2139] Annette Lykknes. Book review: Laura Micheletti Puaca. Searching for Scientific Womanpower: Technocratic Feminism and the Politics of National Security, 1940–1980. Isis, 108(1):228–229, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wellerstein:2017:BRS

[2140] Alex Wellerstein. Book review: Sonja D. Schmid. Producing Power: The Pre-Chernobyl History of the Soviet Nuclear Industry. Isis, 108(1):229– 230, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hogg:2017:BRP

[2141] Jonathan Hogg. Book review: Peter Middleton. Physics Envy: American Poetry and Science in the Cold War and After. Isis, 108(1):231–232, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dick:2017:BRJ

[2142] Steven J. Dick. Book review: James Spiller. Frontiers for the American Century: Outer Space, Antarctica, and Cold War Nationalism. Isis, 108 (1):232–233, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Daum:2017:BRI

[2143] Andreas W. Daum. Book review: Ina Heumann. Gegenstücke: Populäres Wissen im transatlantischen Vergleich (1948–1984). Isis, 108(1):233– 234, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fisher:2017:BRM

[2144] Amy Fisher. Book review: Michael Riordan; Lillian Hoddeson; Adrienne W. Kolb. Tunnel Visions: The Rise and Fall of the Superconducting Super Collider. Isis, 108(1):234–236, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Macauley:2017:BRE

[2145] William R. Macauley. Book review: Erik M. Conway. Exploration and Engineering: The Jet Propulsion Laboratory and the Quest for Mars. Isis, 108(1):236–237, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hannaway:2017:BRR

[2146] Caroline Hannaway. Book review: Richard C. Keller. Fatal Isolation: The Devastating Paris Heat Wave of 2003. Isis, 108(1):237–238, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shavit:2017:BRJ

[2147] Ayelet Shavit. Book review: Jörg Matthias Determann. Researching Biology and Evolution in the Gulf States: Networks of Science in the Middle East. Isis, 108(1):238–240, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2017:Ma

[2148] Anonymous. Masthead. Isis, 108(1):??, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cohen:2017:EIa

[2149] H. Floris Cohen. Editor’s introduction. Isis, 108(1):??, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Elmer:2017:BRS

[2150] Peter Elmer. Book review: Stephen Brogan. The Royal Touch in Early Modern England: Politics, Medicine, and Sin. Isis, 108(1):??, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lehoux:2017:OCE

[2151] Daryn Lehoux. Observation claims and epistemic confidence in Aristotle’s biology. Isis, 108(2):241–258, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mendelsohn:2017:CKB

[2152] J. Andrew Mendelsohn and Annemarie Kinzelbach. Common knowledge: Bodies, evidence, and expertise in Early Modern Germany. Isis, 108(2): 259–279, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nall:2017:CCM

[2153] Joshua Nall. Constructing canals on Mars: Event astronomy and the transmission of international telegraphic news. Isis, 108(2):280–306, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Manias:2017:JSM

[2154] Chris Manias. Jesuit scientists and Mongolian fossils: The French paleontological missions in China, 1923–1928. Isis, 108(2):307–332, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cohen:2017:EIb

[2155] H. Floris Cohen. Editor’s introduction. Isis, 108(2):333–334, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Albaladejo:2017:CSC

[2156] Carolina Martín Albaladejo and Borja Sanchiz. Consequences of the Spanish Civil War for entomology: a quantitative example of abrupt alteration in scientific research dynamics. Isis, 108(2):335–352, June 2017.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). See response [2157] and reply [2158].

Johnson:2017:QCR

[2157] Kristin Johnson. Qualifying consequences: a response to “Consequences of the Spanish Civil War for Entomology”. Isis, 108(2):353–357, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). See [2156].

Albaladejo:2017:BKC

[2158] Carolina Martín Albaladejo and Borja Sanchiz. On building knowledge: Comments on “Qualifying Consequences”. Isis, 108(2):358–359, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). See [2156, 2157].

Maas:2017:IHS

[2159] Ad Maas. Introduction: History of science museums between academics and audiences. Isis, 108(2):360–365, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bergers:2017:SMP

[2160] Lara Bergers and Didi van Trijp. Science museums: a panoramic view. Isis, 108(2):366–370, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Higgitt:2017:CTG

[2161] Rebekah Higgitt. Challenging tropes: Genius, heroic invention, and the longitude problem in the museum. Isis, 108(2):371–380, June 2017. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cain:2017:PTL

[2162] Victoria Cain. Present tense: Locating history in Boston’s museums of science. Isis, 108(2):381–389, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lourenco:2017:TCS

[2163] Marta C. Lourenço and José Pedro Sousa Dias. “time capsules” of science: Museums, collections, and scientific heritage in Portugal. Isis, 108(2):390–398, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heering:2017:SMS

[2164] Peter Heering. Science museums and science education. Isis, 108(2): 399–406, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2017:HSS

[2165] Anonymous. History of Science Society annual meeting, 2016. Isis, 108 (2):407–414, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sidoli:2017:BRK

[2166] Nathan Sidoli. Book review: Karine Chemla; Jacques Virbel, eds. Texts, Textual Acts, and the History of Science. Isis, 108(2):417–418, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lund:2017:BRM

[2167] Mary Ann Lund. Book review: Matthew Bell. Melancholia: The Western Malady; Stephanie Shirilan. Robert Burton and the Transformative Powers of Melancholy. Isis, 108(2):418–420, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vermij:2017:BRB

[2168] Rienk Vermij. Book review: Brian P. Copenhaver. Magic in Western Culture: From Antiquity to the Enlightenment. Isis, 108(2):420–421, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kremer:2017:BRP

[2169] Richard L. Kremer. Book review: Peter Barthel; George H. van Kooten, eds. The Star of Bethlehem and the Magi: Interdisciplinary Perspectives from Experts on the Ancient Near East, the Greco-Roman World, and Modern Astronomy. Isis, 108(2):421–423, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Giannetto:2017:BRM

[2170] Enrico Giannetto. Book review: Maurice Clavelin. Galilée, cosmologie et science du mouvement, suivi de Regards sur l’empirisme au XXe siècle. Isis, 108(2):423–424, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jacobs:2017:BRE

[2171] Noortje Jacobs. Book review: Erika Dyck; Larry Stewart, eds. The Uses of Humans in Experiment: Perspectives from the Seventeenth to the Twentieth Century. Isis, 108(2):424–426, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bennett:2017:BRF

[2172] Jim Bennett. Book review: Fraser MacDonald; Charles W. J. Withers, eds. Geography, Technology, and Instruments of Exploration. Isis, 108 (2):426–427, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reaume:2017:BRW

[2173] Geoffrey Reaume. Book review: Waltraud Ernst, ed. Work, Psychiatry, and Society, c. 1750–2015. Isis, 108(2):428–429, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bocking:2017:BRR

[2174] Stephen Bocking. Book review: Rebecca Priestley. Dispatches from Continent Seven: an Anthology of Antarctic Science. Isis, 108(2):429–430, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kindi:2017:BRR

[2175] Vasso Kindi. Book review: Robert J. Richards; Lorraine Daston, eds. Kuhn’s “Structure of Scientific Revolutions” at Fifty: Reflections on a Science Classic; William J. Devlin; Alisa Bokulich, eds. Kuhn’s “Structure of Scientific Revolutions” — Fifty Years On. Isis, 108(2):430–432, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Piergiacomi:2017:BRM

[2176] Enrico Piergiacomi. Book review: Marco Beretta; Francesco Citti; Alessandro Iannucci, eds. Il culto di Epicuro: Testi, iconografia e paesaggio. Isis, 108(2):433–434, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Evans:2017:BRH

[2177] James Evans. Book review: Héron d’Alexandrie. Metrica: Introduction, texte critique, traduction française et notes de commentaire. Isis, 108(2): 434–435, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Irby:2017:BRS

[2178] Georgia L. Irby. Book review: Serena Bianchetti; Michele Cataudella; Hans-Joachim Gehrke, eds. Brill’s Companion to Ancient Geography:

The Inhabited World in Greek and Roman Tradition. Isis, 108(2):435– 437, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Crisciani:2017:BRD

[2179] Chiara Crisciani. Book review: Danielle Jacquart. Recherches médiévales sur la nature humaine: Essais sur la réflexion médicale (XIIe–XVe s.). Isis, 108(2):437–438, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Glasner:2017:BRP

[2180] Ruth Glasner. Book review: Paul J. J. M. Bakker, ed. Averroes’ Natural Philosophy and Its Reception in the Latin West. Isis, 108(2):438–439, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weill-Parot:2017:BRF

[2181] Nicolas Weill-Parot. Book review: Franck Collard. Les écrits sur les poisons. Isis, 108(2):439–441, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Burnett:2017:BRT

[2182] Charles Burnett. Book review: Tadeusz Grzesik, ed. Disputationes “Physicorum” dictae Magistri Serpentis “Excercitium” (editio critica). Isis, 108(2):441–442, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vonStuckrad:2017:BRA

[2183] Kocku von Stuckrad. Book review: Andreas Lerch. Scientia astrologiae: Der Diskurs über die Wissenschaftlichkeit der Astrologie und die lateinischen Lehrbücher 1470–1610. Isis, 108(2):442–443, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Peterschmitt:2017:BRD

[2184] Luc Peterschmitt. Book review: Didier Kahn. Le fixe et le volatile: Chimie et alchimie, de Paracelse à Lavoisier. Isis, 108(2):443–444, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Maffioli:2017:BRF

[2185] Cesare S. Maffioli. Book review: Federica Favino. La filosofia naturale di Giovanni Ciampoli. Isis, 108(2):444–446, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kaplan:2017:BRG

[2186] Abram Kaplan. Book review: Geoffrey Gorham; Benjamin Hill; Edward Slowik; C. Kenneth Waters, eds. The Language of Nature: Reassessing the Mathematization of Natural Philosophy in the Seventeenth Century. Isis, 108(2):446–447, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nothaft:2017:BRJ

[2187] C. Philipp E. Nothaft. Book review: Johannes Kepler. Vom wahren Geburtsjahr Christi. Isis, 108(2):447–448, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeLacy:2017:BRC

[2188] Margaret DeLacy. Book review: Courtney Weiss Smith. Empiricist Devotions: Science, Religion, and Poetry in Early Eighteenth-Century England. Isis, 108(2):448–449, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tannenbaum:2017:BRT

[2189] Rebecca J. Tannenbaum. Book review: Thomas A. Apel. Feverish Bodies, Enlightened Minds: Science and the Yellow Fever Controversy in the Early American Republic. Isis, 108(2):449–450, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Linnemann:2017:BRD

[2190] Ulf Linnemann. Book review: Dezsö Gurka, ed. Deutsche und ungarische Mineralogen in Jena: Wissenstransfer an der Wende des 18–19. Jahrhunderts im Rahmen der “Societät für die gesammte Mineralogie zu Jena”. Isis, 108(2):450–451, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bergland:2017:BRT

[2191] Renée Bergland. Book review: Tamara Plakins Thornton. Nathaniel Bowditch and the Power of Numbers: How a Nineteenth-Century Man of Business, Science, and the Sea Changed American Life. Isis, 108(2): 451–453, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weber:2017:BRD

[2192] Andreas Weber. Book review: Diarmid A. Finnegan; Jonathan Jeffrey Wright, eds. Spaces of Global Knowledge: Exhibition, Encounter, and Exchange in an Age of Empire. Isis, 108(2):453–454, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Braun:2017:BRS

[2193] Marta Braun. Book review: Sarah Gordon. Indecent Exposures: Eadweard Muybridge’s “Animal Locomotion” Nudes. Isis, 108(2):454–455, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fraser:2017:BRJ

[2194] Craig Fraser. Book review: Jeremy Gray. The Real and the Complex: a History of Analysis in the Nineteenth Century. Isis, 108(2):455–456, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wachelder:2017:BRA

[2195] Joseph Wachelder. Book review: Ana Simões; Maria Paula Diogo; Kostas Gavroglu, eds. Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes. Isis, 108(2):457–458, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tamborini:2017:BRG

[2196] Marco Tamborini. Book review: Gowan Dawson. Show Me the Bone: Reconstructing Prehistoric Monsters in Nineteenth-Century Britain and America; Martin J. S. Rudwick. Earth’s Deep History: How It Was Discovered and Why It Matters. Isis, 108(2):458–460, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mildenberger:2017:BRR

[2197] Florian G. Mildenberger. Book review: Ralph M. Leck. Vita Sexualis: Karl Ulrichs and the Origins of Sexual Science. Isis, 108(2):460–461, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Conlin:2017:BRS

[2198] Jonathan Conlin. Book review: Sarah Amato. Beastly Possessions: Animals in Victorian Consumer Culture. Isis, 108(2):461–462, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bellon:2017:BRJ

[2199] Richard Bellon. Book review: James T. Costa. Wallace, Darwin, and the Origin of Species. Isis, 108(2):463–464, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Huijnen:2017:BRH

[2200] Pim Huijnen. Book review: Heiko Stoff. Gift in der Nahrung: Zur Genese der Verbraucherpolitik Mitte des 20. Jahrhunderts. Isis, 108(2):464–465, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hartmann:2017:BRJ

[2201] Heinrich Hartmann. Book review: Joris Vandendriessche; Evert Peeters; Kaat Wils, eds. Scientists’ Expertise as Performance: Between State and Society, 1860–1960. Isis, 108(2):465–467, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rader:2017:BRS

[2202] Karen A. Rader. Book review: Samuel J. Redman. Bone Rooms: From Scientific Racism to Human Prehistory in Museums. Isis, 108(2):467– 468, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Freyberg:2017:BRG

[2203] Sascha Freyberg. Book review: Gregory B. Moynahan. Ernst Cassirer and the Critical Science of Germany, 1899–1919. Isis, 108(2):468–470, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Killen:2017:BRS

[2204] Andreas Killen. Book review: Scott Curtis. The Shape of Spectatorship: Art, Science, and Early Cinema in Germany. Isis, 108(2):470–471, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Walker:2017:BRT

[2205] J. Samuel Walker. Book review: Timothy J. Jorgensen. Strange Glow: The Story of Radiation. Isis, 108(2):471–472, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Matthews:2017:BRH

[2206] J. Rosser Matthews. Book review: Hyung Wook Park. Old Age, New Science: Gerontologists and Their Biosocial Visions, 1900–1960. Isis, 108(2):472–473, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Talas:2017:BRG

[2207] Sofia Talas. Book review: Giuseppe Bruzzaniti. Enrico Fermi: The Obedient Genius. Isis, 108(2):473–475, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lustig:2017:BRT

[2208] A. J. Lustig. Book review: Tania Munz. The Dancing Bees: Karl von Frisch and the Discovery of the Honeybee Language. Isis, 108(2):475– 476, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Walker:2017:BRC

[2209] Mark Walker. Book review: Christian Fleck. Etablierung in der Fremde: Vertriebene Wissenschaftler in den USA nach 1933. Isis, 108(2):476– 477, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bishop:2017:BRC

[2210] Paul Bishop. Book review: C. G. Jung; Erich Neumann. Analytical Psychology in Exile: The Correspondence of C. G. Jung and Erich Neumann. Isis, 108(2):477–479, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gerovitch:2017:BRS

[2211] Slava Gerovitch. Book review: Sergei S. Demidov; Boris V. Lëvshin, eds. The Case of Academician Nikolai Nikolaevich Luzin. Isis, 108(2):479– 480, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Surman:2017:BRA

[2212] Jan Surman. Book review: Axel Jansen; Andreas Franzmann; Peter Münte, eds. Legitimizing Science: National and Global Publics (1800– 2010). Isis, 108(2):480–482, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dean:2017:BRA

[2213] Katrina Dean. Book review: Aitor Anduaga. Geophysics, Realism, and Industry: How Commercial Interests Shaped Geophysical Conceptions, 1900–1960. Isis, 108(2):482–483, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hays:2017:BRN

[2214] J. N. Hays. Book review: Neel Ahuja. Bioinsecurities: Disease Interventions, Empire, and the Government of Species. Isis, 108(2):483–484, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sieg:2017:BRJ

[2215] Wilfried Sieg. Book review: Juliette Kennedy, ed. Interpreting Gödel: Critical Essays. Isis, 108(2):484–486, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nofre:2017:BRG

[2216] David Nofre. Book review: Giovanni Sommaruga; Thomas Strahm, eds. Turing’s Revolution: The Impact of His Ideas about Computability. Isis, 108(2):486–487, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Murciano-Goroff:2017:BRR

[2217] Yonina R. Murciano-Goroff. Book review: Robert Gilmore McKinnell. The Understanding, Prevention and Control of Human Cancer: The Historic Work and Lives of Elizabeth Cavert Miller and James A. Miller. Isis, 108(2):487–488, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Greenwood:2017:BRM

[2218] Anna Greenwood. Book review: Melissa Graboyes. The Experiment Must Continue: Medical Research and Ethics in East Africa, 1940–2014. Isis, 108(2):489–490, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shapira:2017:BRR

[2219] Michal Shapira. Book review: Rebecca Lemov. Database of Dreams: The Lost Quest to Catalog Humanity. Isis, 108(2):490–491, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Doel:2017:BRS

[2220] Ronald E. Doel. Book review: Sarah Bridger. Scientists at War: The Ethics of Cold War Weapons Research. Isis, 108(2):491–493, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Orr:2017:BRO

[2221] Jackie Orr. Book review: Orit Halpern. Beautiful Data: a History of Vision and Reason since 1945. Isis, 108(2):493–494, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Robinson:2017:BRJ

[2222] Sam Robinson. Book review: Jeroen van Dongen, ed. Cold War Science and the Transatlantic Circulation of Knowledge. Isis, 108(2):494–496, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ensmenger:2017:BRR

[2223] Nathan Ensmenger. Book review: Ronald R. Kline. The Cybernetics Moment: Or Why We Call Our Age the Information Age. Isis, 108(2): 496–497, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dearing:2017:BRS

[2224] John Dearing. Book review: Sigrid Schmalzer. Red Revolution, Green Revolution: Scientific Farming in Socialist China. Isis, 108(2):497–498, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McLennan:2017:BRK

[2225] Sarah McLennan. Book review: Karen Bush Gibson. Women in Space: 23 Stories of First Flights, Scientific Missions, and Gravity-Breaking Adventures. Isis, 108(2):498–499, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Beyler:2017:BRM

[2226] Richard H. Beyler. Book review: Michael Schüring. “Bekennen gegen den Atomstaat”: Die evangelischen Kirchen in der Bundesrepublik Deutschland und die Konflikte um die Atomenergie, 1970–1990. Isis, 108(2): 499–501, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Howe:2017:BRC

[2227] Joshua P. Howe. Book review: Christophe Bonneuil; Jean-Baptiste Fressoz. The Shock of the Anthropocene: The Earth, History and Us. Isis, 108(2):501–502, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2017:Mb

[2228] Anonymous. Masthead. Isis, 108(2):??, June 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mayer:2017:LE

[2229] Andreas Mayer. Letters to the Editor. Isis, 108(2):??, June 2017. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Strick:2017:LE

[2230] James E. Strick. Letters to the Editor. Isis, 108(2):??, June 2017. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ragland:2017:MTS

[2231] Evan R. Ragland. “making trials” in sixteenth- and early seventeenthcentury European academic medicine. Isis, 108(3):503–528, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Buchanan:2017:DDR

[2232] Roderick D. Buchanan and James Bradley. “Darwin’s delay”: a reassessment of the evidence. Isis, 108(3):529–552, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Yeang:2017:MCL

[2233] Chen-Pang Yeang. From modernizing the Chinese language to information science: Chao Yuen Ren’s route to cybernetics. Isis, 108(3):553–580, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Porter:2017:PMC

[2234] Jim Wynter Porter. A “precious minority”: Constructing the “gifted” and “academically talented” student in the era of Brown v. Board of Education and the National Defense Education Act. Isis, 108(3):581–605, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gordin:2017:IHL

[2235] Michael D. Gordin. Introduction: Hegemonic languages and science. Isis, 108(3):606–611, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ragab:2017:CAT

[2236] Ahmed Ragab. “In a clear Arabic tongue”: Arabic and the making of a science-language regime. Isis, 108(3):612–620, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schafer:2017:TMT

[2237] Dagmar Schäfer. Thinking in many tongues: Language(s) and late Imperial China’s science. Isis, 108(3):621–628, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fransen:2017:LTC

[2238] Sietske Fransen. Latin in a time of change: The choice of language as signifier of a new science? Isis, 108(3):629–635, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Terrall:2017:FSL

[2239] Mary Terrall. French in the Siècle des Lumières: a universal language? Isis, 108(3):636–642, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Aronova:2017:RMW

[2240] Elena Aronova. Russian and the making of world languages during the Cold War. Isis, 108(3):643–650, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harvey:2017:JM

[2241] Joy Harvey and Margaret W. Rossiter. Joan T. Mark (1937–2015). Isis, 108(3):651–653, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bengtsson:2017:RR

[2242] Keith R. Bengtsson. Ronald Rainger (1949–2016). Isis, 108(3):654–656, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Westman:2017:R

[2243] Robert S. Westman. In reply. Isis, 108(3):658–659, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heilbron:2017:CC

[2244] J. L. Heilbron. A choice of companions. Isis, 108(3):660–663, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stahnisch:2017:FRH

[2245] Frank W. Stahnisch. False reflexes in the history of physiology. Isis, 108 (3):664–671, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Subramaniam:2017:BRF

[2246] Banu Subramaniam. Book review: Fay Bound Alberti. This Mortal Coil: The Human Body in History and Culture. Isis, 108(3):672–673, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Turner:2017:BRJ

[2247] Laura E. Turner. Book review: John Stillwell. Elements of Mathematics: From Euclid to Gödel. Isis, 108(3):673–674, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Long:2017:BRY

[2248] Jeffery D. Long. Book review: Yiftach Fehige, ed. Science and Religion: East and West. Isis, 108(3):674–675, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Arnold:2017:BRA

[2249] David Arnold. Book review: Anna Winterbottom; Facil Tesfaye, eds. Histories of Medicine and Healing in the Indian Ocean World. Isis, 108 (3):676–678, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kuiper:2017:BRR

[2250] Timothy Kuiper. Book review: Robert B. Campenot. Animal Electricity: How We Learned That the Body and Brain Are Electric Machines. Isis, 108(3):678–679, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sha:2017:BRC

[2251] Richard C. Sha. Book review: Christine Lehleiter, ed. Fact and Fiction: Literary and Scientific Cultures in Germany and Britain. Isis, 108(3): 680–681, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cowan:2017:BRA

[2252] Ruth Schwartz Cowan. Book review: Angela Willey. Undoing Monogamy: The Politics of Science and the Possibilities of Biology.

Isis, 108(3):681–682, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sturdy:2017:BRM

[2253] Steve Sturdy. Book review: Miriam Solomon. Making Medical Knowledge. Isis, 108(3):682–683, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Feser:2017:BRA

[2254] Edward Feser. Book review: Armand Marie Leroi. The Lagoon: How Aristotle Invented Science. Isis, 108(3):684–685, September 2017. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Koetsier:2017:ENC

[2255] Teun Koetsier. Efthymios Nicolaidis; Constantine Skordoulis, eds. The Inscriptions of the Antikythera Mechanism. Isis, 108(3):685–686, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sobiech:2017:BRS

[2256] Frank Sobiech. Book review: Steven P. Marrone. A History of Science, Magic, and Belief: From Medieval to Early Modern Europe. Isis, 108 (3):686–687, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Martelli:2017:BRC

[2257] Matteo Martelli. Book review: Christopher Braun. Das Kitab Sidrat almuntaha des Pseudo-Ibn Wah.siya: Einleitung, Edition und Übersetzung eines hermetisch-allegorischen Traktats zur Alchemie. Isis, 108(3):687– 688, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Spruyt:2017:BRP

[2258] Joke Spruyt. Book review: Peter of Spain. Questiones super libro “De Animalibus” Aristotelis: Critical Edition with Introduction. Isis, 108(3): 688–689, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Goddard:2017:BRS

[2259] Peter A. Goddard. Book review: Surekha Davies. Renaissance Ethnography and the Invention of the Human: New Worlds, Maps, and Monsters. Isis, 108(3):690–691, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schmiedel:2017:BRC

[2260] Irina Schmiedel. Book review: Cristina Bellorini. The World of Plants in Renaissance Tuscany: Medicine and Botany. Isis, 108(3):691–692, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Klein:2017:BRS

[2261] Wouter Klein. Book review: Samir Boumediene. La colonisation du savoir: Une histoire de plantes médicinales du “Nouveau Monde” (1492– 1750). Isis, 108(3):692–694, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brimblecombe:2017:BRW

[2262] Peter Brimblecombe. Book review: William M. Cavert. The Smoke of London: Energy and Environment in the Early Modern City. Isis, 108 (3):694–695, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jesseph:2017:BRV

[2263] Douglas M. Jesseph. Book review: Vincenzo De Risi. Leibniz on the Parallel Postulate and the Foundations of Geometry: The Unpublished Manuscripts. Isis, 108(3):695–696, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lawson:2017:BRRa

[2264] Ian Lawson. Book review: Raz Chen-Morris. Measuring Shadows: Kepler’s Optics of Invisibility. Isis, 108(3):697–698, September 2017. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Slater:2017:BRC

[2265] John Slater. Book review: Cristian Berco. From Body to Community: Venereal Disease and Society in Baroque Spain. Isis, 108(3):698–699, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Charmantier:2017:BRE

[2266] Isabelle Charmantier. Book review: Elizabeth Yale. Sociable Knowledge: Natural History and the Nation in Early Modern Britain. Isis, 108(3): 699–700, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jordanova:2017:BRG

[2267] Ludmilla Jordanova. Book review: E. Geoffrey Hancock; Nick Pearce; Mungo Campbell, eds. William Hunter’s World: The Art and Science of Eighteenth-Century Collecting. Isis, 108(3):700–701, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

daCosta:2017:BRJ

[2268] Palmira Fontes da Costa. Book review: José Pinto de Azeredo. Essays on Some Maladies of Angola (1799). Isis, 108(3):702–703, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gauvin:2017:BRS

[2269] Jean-François Gauvin. Book review: Sofie Lachapelle. Conjuring Science: a History of Scientific Entertainment and Stage Magic in Modern France. Isis, 108(3):703–705, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Palmieri:2017:BRR

[2270] Kristine Palmieri. Book review: Roger Paulin. The Life of August Wilhelm Schlegel, Cosmopolitan of Art and Poetry. Isis, 108(3):705–706, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harley:2017:BRD

[2271] Alexis Harley. Book review: Duncan M. Porter; Peter W. Graham. Darwin’s Sciences.. Isis, 108(3):706–707, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hahn:2017:BRR

[2272] Andre Michael Hahn. Book review: Robert J. Richards; Michael Ruse. Debating Darwin. Isis, 108(3):707–708, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Staubermann:2017:BRS

[2273] Klaus B. Staubermann. Book review: Stephan Paetrow; Wolfgang Wimmer. Carl Zeiss: Eine Biografie, 1816–1888. Isis, 108(3):708–709, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mitchell:2017:BRF

[2274] Daniel Jon Mitchell. Book review: Friedrich Steinle. Exploratory Experiments: Ampère, Faraday, and the Origins of Electrodynamics. Isis, 108 (3):709–711, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barberis:2017:BRD

[2275] Daniela S. Barberis. Book review: David A. West. Darwin’s Man in Brazil: The Evolving Science of Fritz Müller. Isis, 108(3):711–712, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gambaro:2017:BRM

[2276] Ivana Gambaro. Book review: Miriam Focaccia. Uno Scienziato Galantuomo a Via Panisperna: Pietro Blaserna e la nascita dell’Istituto fisico di Roma. Isis, 108(3):712–714, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Patterson:2017:BRH

[2277] Gary Patterson. Book review: Hannah Gay; William P. Griffith. The Chemistry Department at Imperial College London: a History, 1845-2000. Isis, 108(3):714–715, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schroder:2017:BRJ

[2278] Iris Schröder. Book review: Jason D. Hansen. Mapping the Germans: Statistical Science, Cartography, and the Visualization of the German Nation, 1848-1914. Isis, 108(3):715–716, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Casado:2017:BRA

[2279] Santos Casado. Book review: Annette Mülberger, ed. Los límites de la ciencia: espiritismo, hipnotismo y el estudio de los fenómenos paranormales (1850–1930). Isis, 108(3):716–718, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hossfeld:2017:BRH

[2280] Uwe Hoßfeld. Book review: Hanna Engelmeier. Der Mensch, der Affe: Anthropologie und Darwin-Rezeption in Deutschland, 1850-1900. Isis, 108(3):718–719, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Aberdein:2017:BRF

[2281] Andrew Aberdein. Book review: Francine F. Abeles; Mark E. Fuller, eds. Modern Logic, 1850-1950, East and West. Isis, 108(3):719–720, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Osborne:2017:BRR

[2282] Michael A. Osborne. Book review: Rémi Luglia. Des savants pour protéger la nature: La Société d’acclimatation (1854-1960). Isis, 108 (3):720–721, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Grunden:2017:BRS

[2283] Walter Grunden. Book review: Shellen Xiao Wu. Empires of Coal: Fueling China’s Entry into the Modern World Order, 1860-1920. Isis, 108 (3):722–723, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schwarz:2017:BRR

[2284] Astrid Schwarz. Book review: Raf de Bont. Stations in the Field: a History of Place-Based Animal Research, 1870-1930. Isis, 108(3):723– 724, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Yeomans:2017:BRA

[2285] Rory Yeomans. Book review: Amos Morris-Reich. Race and Photography: Racial Photography as Scientific Evidence, 1876-1980. Isis, 108(3):724– 725, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Droscher:2017:BRS

[2286] Ariane Dröscher. Book review: Sona Strbánová. Holding Hands with Bacteria: The Life and Work of Marjory Stephenson. Isis, 108(3):725– 726, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stoltzner:2017:BRJ

[2287] Michael Stöltzner. Book review: Josef Georg Huber. Walther Gerlach (1889-1979) und sein Weg zum erfolgreichen Experimentalphysiker bis etwa 1925. Isis, 108(3):727–728, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sharma:2017:BRS

[2288] Avi Sharma. Book review: Susan E. Cayleff. Nature’s Path: a History of Naturopathic Healing in America. Isis, 108(3):728–729, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Beretta:2017:BRP

[2289] Marco Beretta. Book review: Paola Zambelli. Alexandre Koyré in Incognito. Isis, 108(3):729–731, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weiss:2017:BRP

[2290] Martin P. M. Weiss. Book review: Pascal Schillings. Der letzte weiße Flecken: Europäische Antarktisreisen um 1900. Isis, 108(3):731–732, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Giovanelli:2017:BRJ

[2291] Marco Giovanelli. Book review: Jimena Canales. The Physicist and the Philosopher: Einstein, Bergson, and the Debate That Changed Our Understanding of Time. Isis, 108(3):732–733, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Staudenmaier:2017:BRF

[2292] Peter Staudenmaier. Book review: Florian G. Mildenberger. Der Deutsche Zentralverein homöopathischer Ärzte im Nationalsozialismus: Bestandsaufnahme, Kritik, Interpretation. Isis, 108(3):734–735, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mody:2017:BRR

[2293] Cyrus C. M. Mody. Book review: Rebecca Onion. Innocent Experiments: Childhood and the Culture of Popular Science in the United States. Isis, 108(3):735–736, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Muka:2017:BRJ

[2294] Samantha Muka. Book review: Jennifer Hubbard; David J. Wildish; Robert L. Stephenson, eds. A Century of Maritime Science: The St. Andrews Biological Station. Isis, 108(3):736–737, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fontaine:2017:BRC

[2295] Philippe Fontaine. Book review: Corine Defrance; Anne Kwaschik, eds. La guerre froide et l’internationalisation des sciences: Acteurs, réseaux et institutions. Isis, 108(3):738–739, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Holmes:2017:BRS

[2296] Marcia Holmes. Book review: Scott Selisker. Human Programming: Brainwashing, Automatons, and American Unfreedom. Isis, 108(3):739– 740, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Augustine:2017:BRJ

[2297] Dolores L. Augustine. Book review: John Krige. Sharing Knowledge, Shaping Europe: U.S. Technological Collaboration and Nonproliferation. Isis, 108(3):740–742, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Raby:2017:BRA

[2298] Megan Raby. Book review: Angelo Baracca; Rosella Franconi. Subalternity vs. Hegemony, Cuba’s Outstanding Achievements in Science and Biotechnology, 1959-2014. Isis, 108(3):742–743, September 2017. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brock:2017:BRJ

[2299] David C. Brock. Book review: John Orton; Tom Foxon. Molecular Beam Epitaxy: a Short History. Isis, 108(3):743–744, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hollings:2017:BRK

[2300] Christopher Hollings. Book review: Ken Ono; Amir D. Aczel. My Search for Ramanujan: How I Learned to Count. Isis, 108(3):744–745, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Park:2017:BRL

[2301] Buhm Soon Park. Book review: Lara V. Marks. The Lock and Key of Medicine: Monoclonal Antibodies and the Transformation of Healthcare. Isis, 108(3):746–747, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Peirson:2017:BRA

[2302] B. R. Erick Peirson. Book review: Alan C. Love, ed. Conceptual Change in Biology: Scientific and Philosophical Perspectives on Evolution and Development. Isis, 108(3):747–748, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2017:BRM

[2303] Anonymous. Book review: Masthead. Isis, 108(3):??, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vesel:2017:LE

[2304] Matjaz Vesel. Letters to the Editor. Isis, 108(3):??, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pastorino:2017:PCF

[2305] Cesare Pastorino. The philosopher and the craftsman: Francis Bacon’s notion of experiment and its debt to early Stuart inventors. Isis, 108 (4):749–768, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Somsen:2017:SFF

[2306] Geert Somsen. Science, fascism, and foreign policy: The exhibition “Scienza Universale” at the 1942 Rome World’s Fair. Isis, 108(4):769– 791, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rohde:2017:PTC

[2307] Joy Rohde. Pax technologica: Computers, international affairs, and human reason in the Cold War. Isis, 108(4):792–813, December 2017. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wilson:2017:SIP

[2308] Adrian Wilson. Science’s imagined pasts. Isis, 108(4):814–826, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blum:2017:LRI

[2309] Alexander Blum. The literature review as imagined past. Isis, 108(4): 827–829, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Thomas:2017:SIP

[2310] William Thomas. Scientists’ imagined pasts and historians’ appreciation of scientific thought. Isis, 108(4):830–835, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). See response

Badino:2017:UAP

[2311] Massimiliano Badino. Using, abusing, and perusing the past. Isis, 108 (4):836–840, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Olesko:2017:CTP

[2312] Kathryn M. Olesko. Coming to terms with the past: The great transition. Isis, 108(4):841–845, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wilson:2017:RCS

[2313] Adrian Wilson. Response to comments on “Science’s Imagined Pasts”. Isis, 108(4):846–851, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). See [2310].

Giglioni:2017:GE

[2314] Guido Giglioni. Germana Ernst (1943–2016). Isis, 108(4):852–854, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Smocovitis:2017:WBP

[2315] Vassiliki Betty Smocovitis. William B. Provine (1942–2015). Isis, 108 (4):855–860, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kohlstedt:2017:MCA

[2316] Sally Gregory Kohlstedt. Michele La Clergue Aldrich (1942–2016). Isis, 108(4):861–864, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Demetriou:2017:BRA

[2317] Tania Demetriou. Book review: Anthony Grafton; Glenn W. Most, eds. Canonical Texts and Scholarly Practices: a Global Comparative Approach. Isis, 108(4):867–868, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hughes:2017:BRC

[2318] J. Donald Hughes. Book review: Carolyn Merchant. Autonomous Nature: Problems of Prediction and Control from Ancient Times to the Scientific Revolution. Isis, 108(4):868–869, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Priest:2017:BRM

[2319] Greg Priest. Book review: Marco Solinas. From Aristotle’s Teleology to Darwin’s Genealogy: The Stamp of Inutility. Isis, 108(4):869–870, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bottazzini:2017:BRL

[2320] Umberto Bottazzini. Book review: Leo Corry. A Brief History of Numbers. Isis, 108(4):870–871, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rowe:2017:BRK

[2321] David E. Rowe. Book review: Karine Chemla; Renaud Chorlay; David Rabouin, eds. The Oxford Handbook of Generality in Mathematics and the Sciences. Isis, 108(4):872–873, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dwyer:2017:BRM

[2322] Ellen Dwyer. Book review: Michael R. Trimble. The Intentional Brain: Motion, Emotion, and the Development of Modern Neuropsychiatry. Isis, 108(4):873–874, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Frevert:2017:BRS

[2323] Ute Frevert. Book review: Susan Broomhall, ed. Ordering Emotions in Europe, 1100–1800. Isis, 108(4):874–875, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Crawshaw:2017:BRC

[2324] Jane Stevens Crawshaw. Book review: Claudia Pancino. La natura dei bambini: Cura del corpo, malattie e medicina della prima infanzia fra Cinquecento e Settecento. Isis, 108(4):875–876, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kwa:2017:BRD

[2325] Chunglin Kwa. Book review: Dengjian Jin. The Great Knowledge Transcendence: The Rise of Western Science and Technology Reframed; James E. McClellan; Harold Dorn. Science and Technology in World History: an Introduction. Isis, 108(4):877–879, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Teissier:2017:BRD

[2326] Pierre Teissier. Book review: Donald L. Opitz; Staffan Bergwik; Brigitte Van Tiggelen, eds. Domesticity in the Making of Modern Science. Isis, 108(4):879–880, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Randalls:2017:BRM

[2327] Samuel Randalls. Book review: Marcel Boumans. Science Outside the Laboratory: Measurement in Field Science and Economics. Isis, 108 (4):880–881, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeVorkin:2017:BRL

[2328] David H. DeVorkin. Book review: Lawrence Squeri. Waiting for Contact: The Search for Extraterrestrial Intelligence. Isis, 108(4):882–883, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Berryman:2017:BRC

[2329] Sylvia Berryman. Book review: Courtney Roby. Technical Ekphrasis in Greek and Roman Science and Literature: The Written Machine between Alexandria and Rome. Isis, 108(4):883–884, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cronier:2017:BRA

[2330] Marie Cronier. Book review: Alain Touwaide. A Census of Greek Medical Manuscripts: From Byzantium to the Renaissance. Isis, 108(4):884–885, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roby:2017:BRG

[2331] Courtney Roby. Book review: Giuseppina Ferriello; Maurizio Gatto; Romano Gatto, eds. The “Baroulkos” and the “Mechanics” of Heron. Isis, 108(4):887–888, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Oosterhoff:2017:BRA

[2332] Richard Oosterhoff. Book review: Angela Axworthy. Le mathématicien renaissant et son savoir: Le statut des mathématiques selon Oronce Fine. Isis, 108(4):888–889, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rodrigues:2017:BRU

[2333] Ana Duarte Rodrigues. Book review: Urszula Sowina. Water, Towns, and People: Polish Lands against a European Background until the Mid-Sixteenth Century. Isis, 108(4):889–890, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Manzo:2017:BRJ

[2334] Silvia Manzo. Book review: John M. Forrester, ed. The “De Subtilitate” of Girolamo Cardano. Isis, 108(4):891–892, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ragland:2017:BRM

[2335] Evan R. Ragland. Book review: Michael Stolberg. Uroscopy in Early Modern Europe. Isis, 108(4):892–893, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Thompson:2017:BRC

[2336] Helen Thompson. Book review: Claire Preston. The Poetics of Scientific Investigation in Seventeenth-Century England. Isis, 108(4):893–894, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanGroesen:2017:BRD

[2337] Michiel van Groesen. Book review: Dorothee Schmidt. Reisen in das Orientalische Indien: Wissen über fremde Welten um 1600. Isis, 108 (4):895–896, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fransen:2017:BRM

[2338] Sietske Fransen. Book review: Margaret DeLacy. The Germ of an Idea: Contagionism, Religion, and Society in Britain, 1660–1730. Isis, 108 (4):896–897, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Baldassarri:2017:BRH

[2339] Fabrizio Baldassarri. Book review: Hubertus Fischer; Volker R. Remmert; Joachim Wolschke-Bulmahn, eds. Gardens, Knowledge, and the Sciences in the Early Modern Period. Isis, 108(4):897–898, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Giglioni:2017:BRG

[2340] Guido Giglioni. Book review: Georgiana D. Hedesan. An Alchemical Quest for Universal Knowledge: The ”Christian Philosophy” of Jan Baptist Van Helmont (1579–1644). Isis, 108(4):900–901, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Yale:2017:BRT

[2341] Elizabeth Yale. Book review: Tim Birkhead, ed. Virtuoso by Nature: The Scientific Worlds of Francis Willughby FRS (1635–1672). Isis, 108 (4):901–902, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vermander:2017:BRA

[2342] Benoît Vermander. Book review: Antonella Romano. Impressions de Chine: L’Europe et l’englobement du monde (XVIe–XVIIe siècle). Isis, 108(4):902–903, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lauterbach:2017:BRI

[2343] Iris Lauterbach. Book review: Irina Schmiedel. Pompa e intelletto: Formen der Ordnung und Inszenierung botanischen Wissens im späten Großherzogtum der Medici. Isis, 108(4):904–905, December 2017. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gilje:2017:BRJ

[2344] Paul A. Gilje. Book review: John M. Dixon. The Enlightenment of Cadwallader Colden: Empire, Science, and Intellectual Culture in British New York. Isis, 108(4):905–906, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ratcliff:2017:BRA

[2345] Jessica Ratcliff. Book review: Anna Winterbottom. Hybrid Knowledge in the Early East India Company World. Isis, 108(4):906–907, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Eriksen:2017:BRK

[2346] Christoffer Basse Eriksen. Book review: Keith Michael Baker; Jenna M. Gibbs, eds. Life Forms in the Thinking of the Long Eighteenth Century. Isis, 108(4):907–909, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Seising:2017:BRU

[2347] Rudolf Seising. Book review: Ursula Klein. Nützliches Wissen: Die Erfindung der Technikwissenschaften. Isis, 108(4):909–910, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tarkka-Robinson:2017:BRA

[2348] Laura Tarkka-Robinson. Book review: Anne Mariss. ”A World of New Things”: Praktiken der Naturgeschichte bei Johann Reinhold Forster. Isis, 108(4):910–911, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Oestmann:2017:BRR

[2349] Günther Oestmann. Book review: Richard Dunn; Rebekah Higgitt, eds. Navigational Enterprises in Europe and Its Empires, 1730–1850. Isis, 108 (4):911–912, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Moruno:2017:BRM

[2350] Dolores Martín Moruno. Book review: Marc J. Ratcliff. Genèse d’une découverte: La division des infusoires (1765–1766). Isis, 108(4):913–914, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jankovic:2017:BRA

[2351] Vladimir Janković. Book review: Anya Zilberstein. A Temperate Empire: Making Climate Change in Early America. Isis, 108(4):914–915, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Clement:2017:BRD

[2352] Albert Clement. Book review: Deirdre Loughridge. Haydn’s Sunrise, Beethoven’s Shadow: Audiovisual Culture and the Emergence of Musical Romanticism. Isis, 108(4):915–916, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sparavigna:2017:BRA

[2353] Amelia Carolina Sparavigna. Book review: Amedeo Avogadro. Relazioni accademiche; Amedeo Avogadro. Lettere. Isis, 108(4):917–918, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanMiert:2017:BRM

[2354] Dirk van Miert. Book review: Marion Gindhart; Hanspeter Marti; Robert Seidel, eds. Frühneuzeitliche Disputationen: Polyvalente Produktionsapparate gelehrten Wissens. Isis, 108(4):918–919, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mandrino:2017:BRM

[2355] Agnese Mandrino. Book review: Mauro Gargano; Emilia Olostro Cirella; Massimo Della Valle, eds. Che il diavolo benedica i Pulcinella! Cronache napoletane, scientifiche e non, dell’astronomo von Zach. Isis, 108(4): 919–920, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brooking:2017:BRJ

[2356] Tom Brooking. Book review: James Beattie; Edward Melillo; Emily O’Gorman, eds. Eco-Cultural Networks and the British Empire: New Views on Environmental History. Isis, 108(4):921–922, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mathew:2017:BRD

[2357] John Mathew. Book review: David Arnold. Toxic Histories: Poison and Pollution in Modern India. Isis, 108(4):922–923, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lightman:2017:BRM

[2358] Bernard Lightman. Book review: Michael Ruse. Darwinism as Religion: What Literature Tells Us about Evolution. Isis, 108(4):923–924, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shorter:2017:BRJ

[2359] Edward Shorter. Book review: Jennifer L. Lambe. Madhouse: Psychiatry and Politics in Cuban History. Isis, 108(4):925–926, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Burstyn:2017:BRP

[2360] Harold L. Burstyn. Book review: Peter Moore. The Weather Experiment: The Pioneers Who Sought to See the Future. Isis, 108(4):926–927, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Arvidsson:2017:BRM

[2361] Stefan Arvidsson. Book review: Michael F. Robinson. The Lost White Tribe: Explorers, Scientists, and the Theory That Changed a Continent. Isis, 108(4):927–928, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Frederickson:2017:BRM

[2362] Kathleen Frederickson. Book review: Marc Flandreau. Anthropologists in the Stock Exchange: a Financial History of Victorian Science. Isis, 108 (4):928–929, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cahan:2017:BRC

[2363] David Cahan. Book review: Charles E. McClelland. Berlin, the Mother of All Research Universities, 1860–1918. Isis, 108(4):930–931, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Somsen:2017:BRR

[2364] Geert Somsen. Book review: Robert Fox. Science without Frontiers: Cosmopolitanism and National Interests in the World of Learning, 1870– 1940. Isis, 108(4):931–932, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Israel:2017:BRM

[2365] Paul Israel. Book review: Marc Raboy. Marconi: The Man Who Networked the World. Isis, 108(4):932–933, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Walloch:2017:BRD

[2366] Karen Walloch. Book review: David Wright. SickKids: The History of the Hospital for Sick Children. Isis, 108(4):934–935, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sarukkai:2017:BRP

[2367] Sundar Sarukkai. Book review: Projit Bihari Mukharji. Doctoring Traditions: Ayurveda, Small Technologies, and Braided Sciences. Isis, 108 (4):935–936, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Davis:2017:BRR

[2368] Peter Davis. Book review: Regina Horta Duarte. Activist Biology: The National Museum, Politics, and Nation Building in Brazil. Isis, 108(4):

936–937, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lachmund:2017:BRO

[2369] Jens Lachmund. Book review: Oliver Hochadel; Agustí Nieto-Galan, eds. Barcelona: an Urban History of Science and Modernity, 1888–1929. Isis, 108(4):938–939, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shamdasani:2017:BRK

[2370] Sonu Shamdasani. Book review: Katja Guenther. Localization and Its Discontents: a Genealogy of Psychoanalysis and the Neuro Disciplines. Isis, 108(4):939–940, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Priestley:2017:BRA

[2371] Rebecca Priestley. Book review: Andrew Stuhl. Unfreezing the Arctic: Science, Colonialism, and the Transformation of Inuit Lands. Isis, 108 (4):940–941, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kline:2017:BRT

[2372] Ronald R. Kline. Book review: Tara H. Abraham. Rebel Genius: Warren S. McCulloch’s Transdisciplinary Life in Science. Isis, 108(4):942–943, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DAntonio:2017:BRE

[2373] Patricia D’Antonio. Book review: Elizabeth Fee. Disease and Discovery: a History of the Johns Hopkins School of Hygiene and Public Health, 1916–1939; Karen Kruse Thomas. Health and Humanity: The Story of the Johns Hopkins Bloomberg School of Public Health. Isis, 108(4):943– 945, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schmidt:2017:BRT

[2374] Susanne Schmidt. Book review: Timothy Aubry; Trysh Travis, eds. Rethinking Therapeutic Culture; David Kaiser; W. Patrick McCray, eds. Groovy Science: Knowledge, Innovation, and American Counterculture. Isis, 108(4):946–948, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Washington:2017:BRA

[2375] Harriet A. Washington. Book review: Anthony Ryan Hatch. Blood Sugar: Racial Pharmacology and Food Justice in Black America. Isis, 108(4): 948–949, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2017:IIV

[2376] Anonymous. Index to Isis, volume 108, 2017. Isis, 108(4):950–967, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2017:CIV

[2377] Anonymous. Contents of Isis, volume 108. Isis, 108(4):968–970, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2017:M

[2378] Anonymous. Masthead. Isis, 108(4):??, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heilbron:2017:R

[2379] J. L. Heilbron. In reply. Isis, 108(4):??, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Largent:2017:LE

[2380] Mark Largent and Georgina Montgomery. Letters to the Editor. Isis, 108(4):??, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lawson:2017:BRRb

[2381] Russell Lawson. Book review: Richard Carrier. Science Education in the Early Roman Empire. Isis, 108(4):??, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lightman:2017:LE

[2382] Bernard Lightman. Letters to the Editor. Isis, 108(4):??, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Long:2017:BRS

[2383] Pamela O. Long. Book review: Simone Testa. Italian Academies and Their Networks, 1525–1700: From Local to Global. Isis, 108(4):??, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2017:CBH

[2384] Anonymous. Current bibliography of the history of science and its cultural influences, 2017. Isis, 108(S1):??, December 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heaney:2018:HMI

[2385] Christopher Heaney. How to make an Inca mummy: Andean embalming, Peruvian science, and the collection of empire. Isis, 109(1):1–27, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anthony:2018:MWW

[2386] Patrick Anthony. Mining as the working world of Alexander von Humboldt’s plant geography and vertical cartography. Isis, 109(1):28–55, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gordin:2018:LUS

[2387] Michael D. Gordin. Lysenko unemployed: Soviet genetics after the aftermath. Isis, 109(1):56–78, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Langermann:2018:CSS

[2388] Y. Tzvi Langermann. Capturing the “spirit” of stoic physics. Isis, 109(1): 80–83, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lewis:2018:FHS

[2389] Orly Lewis. Fragments of history and science. Isis, 109(1):83–86, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Engler:2018:CRH

[2390] Fynn Ole Engler, Jürgen Renn, and Matthias Schemmel. Creating room for historical rationality. Isis, 109(1):87–91, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rietmann:2018:NEF

[2391] Felix E. Rietmann. No escape from Fleck. Isis, 109(1):91–94, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeBont:2018:HTS

[2392] Raf De Bont. Humor in a time of science wars: Rereading Isabelle Stengers. Isis, 109(1):95–98, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Leclercq:2018:TSM

[2393] Valérie Leclercq and Joris Vandendriessche. Taking scientists (more) seriously. Isis, 109(1):98–101, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Badino:2018:DCO

[2394] Massimiliano Badino. The discreet charm of ontology. Isis, 109(1):102– 105, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guzzardi:2018:LGE

[2395] Luca Guzzardi. The logic that governs each step of scientific research. Isis, 109(1):105–108, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Portuondo:2018:BTP

[2396] María M. Portuondo. Between two poles: Positivism and historicism in the history of contemporary geography. Isis, 109(1):109–112, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lois:2018:HGN

[2397] Carla Lois. The history of geography: Narratives to discipline the past and reinvent geography as a modern discipline. Isis, 109(1):112–115, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mayer:2018:TRH

[2398] Andreas Mayer. Toward a reflexive history of modern subjectivity. Isis, 109(1):116–118, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Scarfone:2018:PSM

[2399] Marianna Scarfone. The place of subjecthood in madness: Toward an intellectual history of psychiatry on a philosophical basis. Isis, 109(1): 119–121, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanLunteren:2018:GC

[2400] Frans van Lunteren. Geology and Christianity. Isis, 109(1):122–126, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Flipse:2018:FBP

[2401] Abraham (Ab) C. Flipse. Four books for the price of one. Isis, 109(1): 126–129, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schwach:2018:ASW

[2402] Vera Schwach. Appropriating science for a weathered area. Isis, 109(1): 130–133, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roberts:2018:FCH

[2403] Peder Roberts. A forecasting classic in hindsight. Isis, 109(1):133–136, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kevles:2018:AW

[2404] Daniel J. Kevles. Alison Winter (1965–2016). Isis, 109(1):137–139, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Casper:2018:JB

[2405] Stephen T. Casper. John Burnham (1929–2017). Isis, 109(1):140–142, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harrison:2018:AJ

[2406] Carol E. Harrison. Ann Johnson (1965–2016). Isis, 109(1):143–144, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wright:2018:BRL

[2407] Alex Wright. Book review: Lorraine Daston, ed. Science in the Archives: Pasts, Presents, Futures. Isis, 109(1):145–146, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Arabatzis:2018:BRA

[2408] Theodore Arabatzis. Book review: Allan Franklin. What Makes a Good Experiment? Reasons and Roles in Science. Isis, 109(1):146–147, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wragge-Morley:2018:BRJ

[2409] Alexander Wragge-Morley. Book review: Judy A. Hayden, ed. Literature in the Age of Celestial Discovery: From Copernicus to Flamsteed. Isis, 109(1):148–149, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marsh:2018:BRR

[2410] Margaret Marsh. Book review: Robin E. Jensen. Infertility: Tracing the History of a Transformative Term. Isis, 109(1):149–150, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hansen:2018:BRM

[2411] Jason Hansen. Book review: Miles A. Kimball; Charles Kostelnick, eds. Visible Numbers: Essays on the History of Statistical Graphics. Isis, 109(1):150–151, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Davids:2018:BRH

[2412] Karel Davids. Book review: Helge Wendt, ed. The Globalization of Knowledge in the Iberian Colonial World. Isis, 109(1):151–152, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

duPlessis:2018:BRL

[2413] Rory du Plessis. Book review: Lawrence D. Longo; Lawrence P. Reynolds. Wombs with a View: Illustrations of the Gravid Uterus from the Renaissance through the Nineteenth Century. Isis, 109(1):153–154, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hicks:2018:BRM

[2414] Robert D. Hicks. Book review: Marco Beretta; Maria Conforti; Paolo Mazzarello, eds. Savant Relics: Brains and Remains of Scientists. Isis, 109(1):154–155, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Clericuzio:2018:BRJ

[2415] Antonio Clericuzio. Book review: J. D. Trout. Wondrous Truths: The Improbable Triumph of Modern Science. Isis, 109(1):155–156, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shank:2018:BRD

[2416] J. B. Shank. Book review: Danilo Capecchi. The Problem of the Motion of Bodies: a Historical View of the Development of Classical Mechanics; Agamemnon R. E. Oliveira. A History of the Concept of Work: From Physics to Economics. Isis, 109(1):156–160, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

LeLay:2018:BRE

[2417] Colette Le Lay. Book review: Édouard Mehl; Nicolas Roudet, eds. Le temps des astronomes: L’astronomie et la décompte du temps de Pierre d’Ailly à Newton. Isis, 109(1):160–161, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stevens:2018:BRF

[2418] Kathryn Stevens. Book review: Francesca Rochberg. Before Nature: Cuneiform Knowledge and the History of Science. Isis, 109(1):162–163, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Danino:2018:BRJ

[2419] Michel Danino. Book review: Jean-Michel Delire. Les mathématiques de l’autel védique: Le Baudhayana ’Sulbasutra et son commentaire ’Sulbadipika. Isis, 109(1):163–164, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeGroot:2018:BRJ

[2420] Jean De Groot. Book review: Joyce van Leeuwen. The Aristotelian Mechanics: Text and Diagrams. Isis, 109(1):164–165, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanDalen:2018:BRC

[2421] Benno van Dalen. Book review: Christopher Cullen. The Foundations of Celestial Reckoning: Three Ancient Chinese Astronomical Systems. Isis, 109(1):166–167, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Webster:2018:BRH

[2422] Colin Webster. Book review: Harald Siebert. Die ptolemäische ‘Optik” in Spätantike und byzantinischer Zeit: Historiographische Dekonstruktion, textliche Neuerschließung, Rekontextualisierung. Isis, 109(1):167– 168, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ragep:2018:BRA

[2423] F. Jamil Ragep. Book review: Agostino Paravicini Bagliani, ed. The Impact of Arabic Sciences in Europe and Asia. Isis, 109(1):168–170, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Berns:2018:BRD

[2424] Andrew Berns. Book review: David I. Shyovitz. A Remembrance of His Wonders: Nature and the Supernatural in Medieval Ashkenaz. Isis, 109 (1):170–171, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Neuenschwander:2018:BRV

[2425] Erwin Neuenschwander. Book review: Victor J. Katz; Menso Folkerts; Barnabas Hughes; Roi Wagner; J. Lennart Berggren, eds. Sourcebook in the Mathematics of Medieval Europe and North Africa. Isis, 109(1):171– 173, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Galle:2018:BRD

[2426] Karl Galle. Book review: Dag Nikolaus Hasse. Success and Suppression: Arabic Sciences and Philosophy in the Renaissance. Isis, 109(1):173– 174, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Laird:2018:BRR

[2427] W. R. Laird. Book review: Raffaele Pisano; Danilo Capecchi. Tartaglia’s Science of Weights and Mechanics in the Sixteenth Century: Selections from “Quesiti et inventioni diverse”: Books VII–VIII. Isis, 109(1):174– 176, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanBerkel:2018:BRD

[2428] Klaas van Berkel. Book review: Delphine Antoine-Mahut; Stephen Gaukroger, eds. Descartes’ Treatise on Man and Its Reception. Isis, 109(1):177–178, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lynn:2018:BRM

[2429] Michael R. Lynn. Book review: Mi Gyung Kim. The Imagined Empire: Balloon Enlightenments in Revolutionary Europe. Isis, 109(1):178–179, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Paskins:2018:BRA

[2430] Matthew Paskins. Book review: Al Coppola. The Theater of Experiment: Staging Natural Philosophy in Eighteenth-Century Britain. Isis, 109(1): 179–180, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanDamme:2018:BRM

[2431] Stéphane van Damme. Book review: Meghan K. Roberts. Sentimental Savants: Philosophical Families in Enlightenment France. Isis, 109(1): 181–182, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Crippa:2018:BRA

[2432] Davide Crippa. Book review: Amir Alexander. Infinitesimal: How a Dangerous Mathematical Theory Shaped the Modern World. Isis, 109(1): 183–184, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Margocsy:2018:BRJ

[2433] Dániel Margócsy. Book review: Juan Pimentel; Peter Mason. The Rhinoceros and the Megatherium: an Essay in Natural History. Isis, 109(1):184–185, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wardhaugh:2018:BRJ

[2434] Benjamin Wardhaugh. Book review: James Q. Davies; Ellen Lockhart, eds. Sound Knowledge: Music and Science in London, 1789–1851. Isis, 109(1):186–187, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jensen:2018:BRH

[2435] William B. Jensen. Book review: Helge Kragh. Julius Thomsen: a Life in Chemistry and Beyond. Isis, 109(1):187–188, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tattersdill:2018:BRS

[2436] Will Tattersdill. Book review: Stella Pratt-Smith. Transformations of Electricity in Nineteenth-Century Literature and Science. Isis, 109(1): 188–189, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hesketh:2018:BRG

[2437] Ian Hesketh. Book review: Gillian Beer. Alice in Space: The Sideways Victorian World of Lewis Carroll. Isis, 109(1):189–191, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ziche:2018:BRS

[2438] Paul Ziche. Book review: Saulo de Freitas Araujo. Wundt and the Philosophical Foundations of Psychology: a Reappraisal. Isis, 109(1):191–192, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blowers:2018:BRB

[2439] Geoffrey Blowers. Book review: Brian J. McVeigh. The History of Japanese Psychology: Global Perspectives, 1875–1950. Isis, 109(1):192– 193, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barrow:2018:BRM

[2440] Mark V. Barrow, Jr. Book review: Marc E. Epstein. Moths, Myths, and Mosquitoes: The Eccentric Life of Harrison G. Dyar, Jr. Isis, 109(1): 194–195, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stump:2018:BRM

[2441] David J. Stump. Book review: Michel Fichant; Sophie Roux, eds. Louis Couturat (1868–1914): Mathématiques, langage, philosophie. Isis, 109 (1):195–196, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Flipse:2018:BRL

[2442] Abraham C. Flipse. Book review: Lyvia Diser. Wetenschap op de proef: Laboratoria in het Belgisch overheidsbeleid, 1870–1940. Isis, 109(1):196– 197, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Howell:2018:BRJ

[2443] Joel D. Howell. Book review: Jeanne Kisacky. Rise of the Modern Hospital: an Architectural History of Health and Healing, 1870–1940. Isis, 109(1):197–198, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Matthews:2018:BRA

[2444] J. Rosser Matthews. Book review: Andreas-Holger Maehle. Contesting Medical Confidentiality: Origins of the Debate in the United States, Britain, and Germany. Isis, 109(1):199–200, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mills:2018:BRB

[2445] Eric L. Mills. Book review: Bo Poulsen. Global Marine Science and Carlsberg: The Golden Connections of Johannes Schmidt (1877–1933). Isis, 109(1):200–201, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Eckert:2018:BRK

[2446] Michael Eckert. Book review: Klaus Hentschel; Ning Yan Zhu, eds. Gustav Robert Kirchhoff’s Treatise “On the Theory of Light Rays” (1882): English Translation, Analysis, and Commentary. Isis, 109(1):201–202, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Broere:2018:BRL

[2447] Sebastiaan Broere. Book review: Leslie Topp. Freedom and the Cage: Modern Architecture and Psychiatry in Central Europe, 1890–1914. Isis, 109(1):202–203, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ruberg:2018:BRI

[2448] Willemijn Ruberg. Book review: Ian Burney; Neil Pemberton. Murder and the Making of English CSI. Isis, 109(1):204–205, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Metraux:2018:BRJ

[2449] Alexandre Métraux. Book review: Johann Georg Reicheneder. Freuds Traum von einer neuen Wissenschaft: Chemie und Bakteriologie im Traum von Irmas Injektion. Isis, 109(1):205–206, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Libell:2018:BRJ

[2450] Monica Libell. Book review: James Carleton Paget; Michael J. Thate, eds. Albert Schweitzer in Thought and Action: a Life in Parts. Isis, 109(1):206–207, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hoeneveld:2018:BRM

[2451] Friso Hoeneveld. Book review: Martijn van Calmthout. Sam Goudsmit: Zijn jacht op de atoombom van Hitler. Isis, 109(1):207–208, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Largent:2018:BRA

[2452] Mark A. Largent. Book review: Angus McLaren. Reproduction by Design: Sex, Robots, Trees, and Test-Tube Babies in Interwar Britain. Isis, 109(1):209–210, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hopwood:2018:BRK

[2453] Nick Hopwood. Book review: Klaus Taschwer. Der Fall Paul Kammerer: Das abenteurliche Leben des umstrittensten Biologen seiner Zeit. Isis, 109(1):210–211, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Beyler:2018:BRS

[2454] Richard H. Beyler. Book review: Sören Flachowsky; Rüdiger Hachtmann; Florian Schmaltz, eds. Ressourcenmobilisierung: Wissenschaftspolitik und Forschungspraxis im NS-Herrschaftssystem. Isis, 109(1): 211–213, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Thomas:2018:BRC

[2455] Wolfgang Thomas. Book review: Chris Bernhardt. Turing’s Vision: The Birth of Computer Science. Isis, 109(1):213–214, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kirsh:2018:BRS

[2456] Nurit Kirsh. Book review: Shelley Z. Reuter. Testing Fate: Tay–Sachs Disease and the Right to Be Responsible. Isis, 109(1):214–215, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Esposito:2018:BRE

[2457] Maurizio Esposito. Book review: Erik L. Peterson. The Life Organic: The Theoretical Biology Club and the Roots of Epigenetics. Isis, 109(1): 216–217, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Evans:2018:BRK

[2458] Bonnie Evans. Book review: Kieran McNally. A Critical History of Schizophrenia. Isis, 109(1):217–218, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guttler:2018:BRW

[2459] Nils Güttler. Book review: William Rankin. After the Map: Cartography, Navigation, and the Transformation of Territory in the Twentieth Century. Isis, 109(1):219–220, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hogan:2018:BRF

[2460] Andrew J. Hogan. Book review: Franklin H. Portugal. The Least Likely Man: Marshall Nirenberg and the Discovery of the Genetic Code. Isis, 109(1):220–221, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Erickson:2018:BRH

[2461] Paul Erickson. Book review: Hunter Heyck. Age of System: Understanding the Development of Modern Social Science. Isis, 109(1):221– 222, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Isaac:2018:BRP

[2462] Joel Isaac. Book review: Paul Rubinson. Redefining Science: Scientists, the National Security State, and Nuclear Weapons in Cold War America. Isis, 109(1):223–224, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dvorak:2018:BRJ

[2463] Anna Elizabeth Dvorak. Book review: James W. Feldman, ed. Nuclear Reactions: Documenting American Encounters with Nuclear Energy. Isis, 109(1):224–225, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lambright:2018:BRD

[2464] W. Henry Lambright. Book review: Doug Millard. Satellite: Innovation in Orbit; David J. Shayler; David M. Harland. The Hubble Space Telescope: From Concept to Success. Isis, 109(1):225–226, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vertesi:2018:BRM

[2465] Janet Vertesi. Book review: Melissa L. Sevigny. Under Desert Skies: How Tucson Mapped the Way to the Moon and Planets. Isis, 109(1): 227–228, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Alexander:2018:BRM

[2466] Amir Alexander. Book review: Mark Coeckelbergh. New Romantic Cyborgs: Romanticism, Information Technology, and the End of the Machine. Isis, 109(1):228–229, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Spiller:2018:BRA

[2467] James Spiller. Book review: Adrian Howkins. Frozen Empires: an Environmental History of the Antarctic Peninsula. Isis, 109(1):229–230, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gregory:2018:BRP

[2468] Jane Gregory. Book review: Peter Collins. The Royal Society and the Promotion of Science since 1960. Isis, 109(1):231–232, March 2018. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Slayton:2018:BRB

[2469] Rebecca Slayton. Book review: Book Review: Cyrus C. M. Mody. The Long Arm of Moore’s Law: Microelectronics and American Science. Isis, 109(1):232–233, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2018:Ma

[2470] Anonymous. Masthead. Isis, 109(1):??, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cohen:2018:EI

[2471] H. Floris Cohen. Editor’s introduction. Isis, 109(1):??, March 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hendriksen:2018:CCA

[2472] Marieke M. A. Hendriksen. Criticizing Chrysopoeia? Alchemy, chemistry, academics, and satire in the northern Netherlands, 1650–1750. Isis, 109(2):235–253, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rocke:2018:TVP

[2473] Alan J. Rocke. Theory versus practice in German chemistry: Erlenmeyer beyond the flask. Isis, 109(2):254–275, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ramsden:2018:NDL

[2474] Edmund Ramsden. A neurotic dog’s life: Experimental psychiatry and the conditional reflex method in the work of W. Horsley Gantt. Isis, 109(2):276–301, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dupre:2018:ISP

[2475] Sven Dupré. Introduction: Science and practices of translation. Isis, 109(2):302–307, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lehoux:2018:LUM

[2476] Daryn Lehoux. Let us make the effort: Science into Latin in antiquity. Isis, 109(2):308–312, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Montgomery:2018:MSE

[2477] Scott L. Montgomery. Mobilities of science: The era of translation into Arabic. Isis, 109(2):313–319, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kucuk:2018:ATS

[2478] B. Harun Küçük. Arabic into Turkish in the Seventeenth Century. Isis, 109(2):320–325, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Frumer:2018:TWB

[2479] Yulia Frumer. Translating words, building worlds: Meteorology in Japanese, Dutch, and Chinese. Isis, 109(2):326–332, June 2018. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vermij:2018:TAM

[2480] Rienk Vermij. Translating, adapting, mutilating: Or, how to make an Enlightenment classic. Isis, 109(2):333–338, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kursell:2018:AET

[2481] Julia Kursell. Alexander Ellis’s translation of Helmholtz’s Sensations of Tone. Isis, 109(2):339–345, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2018:HSS

[2482] Anonymous. History of Science Society annual meeting, 2017. Isis, 109 (2):346–353, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Farber:2018:FBC

[2483] Paul Lawrence Farber. Frederick Barton Churchill (1932–2017). Isis, 109(2):354–355, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heilbron:2018:TF

[2484] J. L. Heilbron. Tore Frängsmyr (1938–2017). Isis, 109(2):356–359, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Karstens:2018:KPT

[2485] Bart Karstens. Karl Popper and Thomas Kuhn: Historical figures with future significance. Isis, 109(2):360–364, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Beckers:2018:BRB

[2486] Danny Beckers. Book review: Brendan Larvor, ed. Mathematical Cultures: The London Meetings, 2012–2014. Isis, 109(2):365–366, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marrone:2018:BRE

[2487] Steven P. Marrone. Book review: Edward Bever; Randall Styers, eds. Magic in the Modern World: Strategies of Repression and Legitimization. Isis, 109(2):366–367, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sorensen:2018:BRP

[2488] Henrik Kragh Sørensen. Book review: Patrick Popescu-Pampu. What Is the Genus? Isis, 109(2):367–368, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ogilvie:2018:BRP

[2489] Brian W. Ogilvie. Book review: Peter Marren. Rainbow Dust: Three Centuries of Butterfly Delight. Isis, 109(2):369–370, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Archibald:2018:BRV

[2490] Tom Archibald. Book review: Volker R. Remmert; Martina R. Schneider; Henrik Kragh Sørensen, eds. Historiography of Mathematics in the Nineteenth and Twentieth Centuries. Isis, 109(2):370–372, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Belknap:2018:BRG

[2491] Geoffrey Belknap. Book review: Gregg Mitman; Kelley Wilder, eds. Documenting the World: Film, Photography, and the Scientific Record. Isis, 109(2):372–373, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Seidel:2018:BRD

[2492] Robert W. Seidel. Book review: David F. Channell. A History of Technoscience: Erasing the Boundaries between Science and Technology. Isis, 109(2):373–374, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vargha:2018:BRR

[2493] Dora Vargha. Book review: Randall M. Packard. A History of Global Health: Interventions into the Lives of Other Peoples. Isis, 109(2):374– 375, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

King:2018:BRA

[2494] Colin Guthrie King. Book review: Alexander Jones, ed. Time and Cosmos in Greco-Roman Antiquity. Isis, 109(2):376–377, June 2018. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mayhew:2018:BRD

[2495] Robert J. Mayhew. Book review: Daniela Dueck, ed. The Routledge Companion to Strabo. Isis, 109(2):377–378, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nutton:2018:BRO

[2496] Vivian Nutton. Book review: Orly Lewis. Praxagoras of Cos on Arteries, Pulse, and Pneuma. Isis, 109(2):378–379, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lamberton:2018:BRM

[2497] Robert Lamberton. Book review: Michiel Meeusen. Plutarch’s Science of Natural Problems: a Study with Commentary on “Quaestiones Naturales”. Isis, 109(2):379–380, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Oliva:2018:BRK

[2498] Pilar Herráiz Oliva. Book review: Kellie Robertson. Nature Speaks: Medieval Literature and Aristotelian Philosophy. Isis, 109(2):380–381, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blanco:2018:BRJ

[2499] Mónica Blanco. Book review: Jacques Sesiano. Magic Squares in the Tenth Century: Two Arabic Treatises by Ant.aki and Buzjani. Isis, 109 (2):381–382, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McVaugh:2018:BRG

[2500] Michael McVaugh. Book review: Gilles de Corbeil; Mireille Ausécache, eds. Liber de uirtutibus et laudibus compositorum medicaminum. Isis, 109(2):383–384, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hannam:2018:BRZ

[2501] James Hannam. Book review: Zachary A. Matus. Franciscans and the Elixir of Life: Religion and Science in the Later Middle Ages. Isis, 109 (2):384–385, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cifoletti:2018:BRG

[2502] Giovanna C. Cifoletti. Book review: Guillaume Gosselin; Odile Le Guillou-Kouteynikoff. De arte magna libri IV: Traité d’algèbre, suivi de Praelectio / Leçon sur la mathématique. Isis, 109(2):385–386, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Draxler:2018:BRR

[2503] Sonja Draxler. Book review: Rivka Feldhay; F. Jamil Ragep, eds. Before Copernicus: The Cultures and Contexts of Scientific Learning in the Fifteenth Century. Isis, 109(2):386–387, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lammel:2018:BRJ

[2504] Hans-Uwe Lammel. Book review: Jana Madlen Schütte. Medizin im Konflikt: Fakultäten, Märkte und Experten in deutschen Universitätsstädten des 14. bis 16. Jahrhunderts. Isis, 109(2):387–388, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Horst:2018:BRB

[2505] Thomas Horst. Book review: Battista Agnese; Ingrid Baumgärtner, eds. Der Portulan-Atlas des Battista Agnese: Das Kasseler Prachtexemplar von 1542. Isis, 109(2):389–390, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Regier:2018:BRR

[2506] Jonathan Regier. Book review: Robert S. Westman. Copernicus and the Astrologers: Dibner Library Lecture, December 12, 2013. Isis, 109(2): 390–391, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Phaf-Rheinberger:2018:BRPa

[2507] Ineke Phaf-Rheinberger. Book review: Pablo F. Gómez. The Experiential Caribbean: Creating Knowledge and Healing in the Early Modern Atlantic. Isis, 109(2):392–393, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Knoeff:2018:BRO

[2508] Rina Knoeff. Book review: Ole Peter Grell; Andrew Cunningham, eds. Medicine, Natural Philosophy, and Religion in Post-Reformation Scandinavia. Isis, 109(2):393–394, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Giudice:2018:BRM

[2509] Franco Giudice. Book review: Miguel Á. Granada; Patrick J. Boner; Dario Tessicini, eds. Unifying Heaven and Earth: Essays in the History of Early Modern Cosmology. Isis, 109(2):394–395, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bellhouse:2018:BRB

[2510] David Bellhouse. Book review: Brian Rice; Enrique González–Velasco; Alexander Corrigan. The Life and Works of John Napier. Isis, 109(2): 396–397, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kaplan:2018:BRR

[2511] Abram Kaplan. Book review: Renée Raphael. Reading Galileo: Scribal Technology and the “Two New Sciences”. Isis, 109(2):397–398, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Abou-Nemeh:2018:BRL

[2512] Catherine Abou-Nemeh. Book review: Lesley B. Cormack; Steven A. Walton; John A. Schuster, eds. Mathematical Practitioners and the Transformation of Natural Knowledge in Early Modern Europe. Isis, 109(2):398–400, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Frilli:2018:BRG

[2513] Guido Frilli. Book review: Gregorio Baldin, ed. Hobbes e Galileo: Metodo, materia e scienza del moto. Isis, 109(2):400–401, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Canizares-Esguerra:2018:BRA

[2514] Jorge Cañizares-Esguerra. Book review: Amber Brian. Alva Ixtlilxochitl’s Native Archive and the Circulation of Knowledge in Colonial Mexico. Isis, 109(2):401–402, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nasim:2018:BRC

[2515] Omar W. Nasim. Book review: Clifford J. Cunningham. Bode’s Law and the Discovery of Juno: Historical Studies in Asteroid Research. Isis, 109(2):403–404, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pyenson:2018:BRA

[2516] Lewis Pyenson. Book review: Alan Bewell. Natures in Translation: Romanticism and Colonial Natural History. Isis, 109(2):404–405, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kant:2018:BRJ

[2517] Horst Kant. Book review: Jürgen Teichmann. Der Geheimcode der Sterne: Eine neue Landschaft des Himmels und die Geburt der Astrophysik. Isis, 109(2):405–406, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kingsland:2018:BRC

[2518] Sharon Kingsland. Book review: Courtney Fullilove. The Profit of the Earth: The Global Seeds of American Agriculture; Helen Anne Curry. Evolution Made to Order: Plant Breeding and Technological Innovation in Twentieth-Century America. Isis, 109(2):406–409, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Duffin:2018:BRT

[2519] Jacalyn Duffin. Book review: Tine van Osselaer; Henk de Smaele; Kaat Wils, eds. Sign or Symptom? Exceptional Corporeal Phenomena in Religion and Medicine in the Nineteenth and Twentieth Centuries. Isis, 109(2):409–410, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sommer:2018:BRK

[2520] Andreas Sommer. Book review: Krister Dylan Knapp. William James: Psychical Research and the Challenge of Modernity. Isis, 109(2):410– 411, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Yudell:2018:BRS

[2521] Michael Yudell. Book review: Staffan Müller–Wille; Christina Brandt, eds. Heredity Explored: Between Public Domain and Experimental Science, 1850–1930. Isis, 109(2):412–413, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Engelmeier:2018:BRU

[2522] Hanna Engelmeier. Book review: Uwe Hoßfeld. Geschichte der biologischen Anthropologie in Deutschland: Von den Anfängen bis in die Nachkriegszeit. Isis, 109(2):413–414, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Burnett:2018:BRG

[2523] D. Graham Burnett. Book review: Gisela Parak. Photographs of Environmental Phenomena: Scientific Images in the Wake of Environmental Awareness, USA 1860s-1970s. Isis, 109(2):414–415, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chiang:2018:BRE

[2524] Howard Chiang. Book review: Hoi-Eun Kim. Doctors of Empire: Medical and Cultural Encounters between Imperial Germany and Meiji Japan. Isis, 109(2):416–417, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pichel:2018:BRG

[2525] Beatriz Pichel. Book review: Geoffrey Belknap. From a Photograph: Authenticity, Science, and the Periodical Press, 1870-1890. Isis, 109(2): 417–418, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gossin:2018:BRD

[2526] Pamela Gossin. Book review: David Baron. American Eclipse: a Nation’s Epic Race to Catch the Shadow of the Moon and Win the Glory of the World. Isis, 109(2):418–419, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Arnold:2018:BRT

[2527] David Arnold. Book review: Thuy Linh Nguyen. Childbirth, Maternity, and Medical Pluralism in French Colonial Vietnam, 1880-1945. Isis, 109 (2):420–421, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Durant:2018:BRM

[2528] John Durant. Book review: Marianne Sommer. History Within: The Science, Culture, and Politics of Bones, Organisms, and Molecules. Isis, 109(2):421–422, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brenni:2018:BRR

[2529] Paolo Brenni. Book review: Roland Wittje. The Age of Electroacoustics: Transforming Science and Sound. Isis, 109(2):423–424, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cayleff:2018:BRH

[2530] Susan E. Cayleff. Book review: Holly Folk. The Religion of Chiropractic: Populist Healing from the American Heartland. Isis, 109(2):424–425, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gariboldi:2018:BRD

[2531] Leonardo Gariboldi. Book review: David N. Schwartz. The Last Man Who Knew Everything: The Life and Times of Enrico Fermi, Father of the Nuclear Age. Isis, 109(2):425–427, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fichman:2018:BRM

[2532] Martin Fichman. Book review: Matthew J. James. Collecting Evolution: The Galápagos Expedition That Vindicated Darwin. Isis, 109(2):427– 428, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Borck:2018:BRM

[2533] Cornelius Borck. Book review: Michael Sappol. Body Modern: Fritz Kahn, Scientific Illustration, and the Homuncular Subject. Isis, 109(2): 428–429, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stoff:2018:BRT

[2534] Heiko Stoff. Book review: Tiago Saraiva. Fascist Pigs: Technoscientific Organisms and the History of Fascism. Isis, 109(2):430–431, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanBerkel:2018:BRJ

[2535] Klaas van Berkel. Book review: Johannes Feichtinger; Herbert Matis; Stefan Sienell; Heidemarie Uhl; Nick Somers; Cynthia Peck-Kubaczek, eds. The Academy of Sciences in Vienna, 1938 to 1945. Isis, 109(2): 431–432, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ito:2018:BRB

[2536] Takashi Ito. Book review: Book Review: Daniel E. Bender. The Animal Game: Searching for Wildness at the American Zoo. Isis, 109(2):432– 433, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shindell:2018:BRK

[2537] Matthew Shindell. Book review: Kristine C. Harper. Make It Rain: State Control of the Atmosphere in Twentieth-Century America. Isis, 109(2): 434–435, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Phillips:2018:BRM

[2538] Christopher J. Phillips. Book review: Margot Lee Shetterly, ???? Isis, 109(2):435–436, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Grier:2018:BRM

[2539] David Alan Grier. Book review: Marie Hicks. Programmed Inequality: How Britain Discarded Women Technologists and Lost Its Edge in Computing. Isis, 109(2):436–437, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Halpern:2018:BRR

[2540] Paul Halpern. Book review: Roberto Lalli. Building the General Relativity and Gravitation Community during the Cold War. Isis, 109(2): 437–438, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harper:2018:BRE

[2541] Kristine C. Harper. Book review: Egle Rindzeviciute. The Power of Systems: How Policy Sciences Opened Up the Cold War World. Isis, 109(2):439–440, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Townsend:2018:BRS

[2542] Robert B. Townsend. Book review: Steven J. Diner. Universities and Their Cities: Urban Higher Education in America. Isis, 109(2):440– 441, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wilson:2018:BRB

[2543] Duncan Wilson. Book review: Book Review: J. Benjamin Hurlbut. Experiments in Democracy: Human Embryo Research and the Politics of Bioethics. Isis, 109(2):441–442, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rothman:2018:BRB

[2544] David J. Rothman. Book review: Book Reviews: Duncan Wilson. The Making of British Bioethics; Sarah Ferber. Bioethics in Historical Perspective. Isis, 109(2):442–443, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2018:Mb

[2545] Anonymous. Masthead. Isis, 109(2):??, June 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Zhmud:2018:AGM

[2546] Leonid Zhmud and Alexei Kouprianov. Ancient Greek Mathemata from a sociological perspective: a quantitative analysis. Isis, 109(3):445–472, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hess:2018:PMC

[2547] Volker Hess. A paper machine of clinical research in the Early Twentieth Century. Isis, 109(3):473–493, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Carvajalino:2018:EBW

[2548] Juan Carvajalino. Edwin Bidwell Wilson and mathematics as a language. Isis, 109(3):494–514, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Inkpen:2018:SLA

[2549] Danielle K. Inkpen. The scientific life in the alpine: Recreation and moral life in the field. Isis, 109(3):515–537, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Baldwin:2018:SAP

[2550] Melinda Baldwin. Scientific autonomy, public accountability, and the rise of “peer review” in the Cold War United States. Isis, 109(3):538–558, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reinhardt:2018:I

[2551] Carsten Reinhardt. Introduction. Isis, 109(3):559–564, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Homburg:2018:CIT

[2552] Ernst Homburg. Chemistry and industry: a tale of two moving targets. Isis, 109(3):565–576, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rocke:2018:ICP

[2553] Alan J. Rocke. Ideas in chemistry: The pure and the impure. Isis, 109 (3):577–586, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nye:2018:BTH

[2554] Mary Jo Nye. Boundaries, transformations, historiography: Physics in chemistry from the 1920s to the 1960s. Isis, 109(3):587–596, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bensaude-Vincent:2018:CB

[2555] Bernadette Bensaude-Vincent. Chemists without borders. Isis, 109(3): 597–607, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Smocovitis:2018:EHC

[2556] Vassiliki Betty Smocovitis. Eloge: Hamilton Cravens (1938–2015). Isis, 109(3):608–611, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Deuchler:2018:BRS

[2557] Martina Deuchler. Book review: Soyoung Suh. Naming the Local: Medicine, Language, and Identity in Korea since the Fifteenth Century. Isis, 109(3):612–613, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanBesouw:2018:BRJ

[2558] Jip van Besouw. Book review: Jutta Schickore. About Method: Experimenters, Snake Venom, and the History of Writing Scientifically. Isis, 109(3):613–614, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mishra:2018:BRJ

[2559] Saurabh Mishra. Book review: Jonathan Lamb. Scurvy: The Disease of Discovery. Isis, 109(3):614–615, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cornish:2018:BRS

[2560] Caroline Cornish. Book review: Steven Lubar. Inside the Lost Museum: Curating, Past and Present. Isis, 109(3):616–617, September 2018. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Zachary:2018:BRB

[2561] G. Pascal Zachary. Book review: Benoît Godin. Models of Innovation: The History of an Idea. Isis, 109(3):617–618, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bertoni:2018:BRG

[2562] Daniel Bertoni. Book review: Gavin Hardy; Laurence Totelin. Ancient Botany. Isis, 109(3):618–619, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Touwaide:2018:BRB

[2563] Alain Touwaide. Book review: Barbara S. Bowers; Linda Migl Keyser, eds. The Sacred and the Secular in Medieval Healing: Sites, Objects, and Texts. Isis, 109(3):619–620, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mimura:2018:BRD

[2564] Taro Mimura. Book review: Dominique Raynaud. A Critical Edition of Ibn al-Haytham’s On the Shape of the Eclipse: The First Experimental Study of the Camera Obscura. Isis, 109(3):621–622, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Trigg:2018:BRS

[2565] Scott Trigg. Book review: Sally P. Ragep. Jaghmini’s Mulakhkhas.: An Islamic Introduction to Ptolemaic Astronomy. Isis, 109(3):622–623, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tomas:2018:BRB

[2566] Natalie Tomas. Book review: Brendan Dooley. Angelica’s Book and the World of Reading in Late Renaissance Italy. Isis, 109(3):623–624, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Arrizabalaga:2018:BRE

[2567] Jon Arrizabalaga. Book review: Elena del Rio Parra. Materia médica: Rareza, singularidad y accidente en la España temprano-moderna. Isis, 109(3):624–625, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schechner:2018:BRM

[2568] Sara J. Schechner. Book review: Matteo Valleriani, ed. The Structures of Practical Knowledge. Isis, 109(3):626–627, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rey:2018:BRG

[2569] Anne-Lise Rey. Book review: Guillaume Carnino; Liliane Hilaire–Pérez; Aleksandra Kobiljski, eds. Histoire des techniques: Mondes, sociétés, cultures (XVIe–XVIIIe siècle); Liliane Hilaire–Pérez; Fabien Simon; Marie Thébaud–Sorger, eds. L’Europe des sciences et des techniques: Un dialogue des savoirs, Xve–XVIIIe siècle. Isis, 109(3):627–629, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cavallo:2018:BRS

[2570] Sandra Cavallo. Book review: Sabrina Minuzzi. Sul filo dei segreti: Farmacopea, libri e pratiche terapeutiche a Venezia in età moderna. Isis, 109 (3):629–630, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bruyns:2018:BRG

[2571] W. F. J. Mörzer Bruyns. Book review: Günter Schilder. Early Dutch Maritime Cartography: The North Holland School of Cartography (c. 1580–c. 1620). Isis, 109(3):632–633, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pennell:2018:BRA

[2572] Sara Pennell. Book review: Anne Stobart. Household Medicine in Seventeenth-Century England. Isis, 109(3):633–634, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Agar:2018:BRM

[2573] Jon Agar. Book review: Matthew L. Jones. Reckoning with Matter: Calculating Machines, Innovation, and Thinking about Thinking from Pascal to Babbage. Isis, 109(3):634–635, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Moxham:2018:BRD

[2574] Noah Moxham. Book review: David Banks. The Birth of the Academic Article: Le Journal des Sçavans and the Philosophical Transactions, 1665–1700. Isis, 109(3):636–637, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jin:2018:BRC

[2575] Dengjian Jin. Book review: Clifford Siskin. System: The Shaping of Modern Knowledge. Isis, 109(3):637–638, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wichner:2018:BRC

[2576] Jessika Wichner. Book review: Clare Brant. Balloon Madness: Flights of Imagination in Britain, 1783–1786. Isis, 109(3):638–639, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harley:2018:BRD

[2577] Alexis Harley. Book review: Devin Griffiths. The Age of Analogy: Science and Literature between the Darwins. Isis, 109(3):640–641, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Beltran:2018:BRM

[2578] José Beltrán. Book review: Marie–Noëlle Bourguet. Le monde dans un carnet: Alexander von Humboldt en Italie (1805). Isis, 109(3):641–642, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Elffers:2018:BRR

[2579] Els Elffers. Book review: Richard Leblanc. Fearful Asymmetry: Bouillaud, Dax, Broca, and the Localization of Language, Paris, 1825–1879. Isis, 109(3):642–643, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Conlin:2018:BREa

[2580] Jonathan Conlin. Book review: Evelleen Richards. Darwin and the Making of Sexual Selection. Isis, 109(3):644–645, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sibeud:2018:BRJ

[2581] Emmanuelle Sibeud. Book review: Jehanne-Emmanuelle Monnier; Évelyne Combeau-Mari. Profession explorateur: Alfred Grandidier 1836– 1921. Isis, 109(3):645–646, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Morgan:2018:BRB

[2582] Monique R. Morgan. Book review: Benjamin Morgan. The Outward Mind: Materialist Aesthetics in Victorian Science and Literature. Isis, 109(3):646–647, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Naylor:2018:BRJ

[2583] Simon Naylor. Book review: Jeremy Vetter. Field Life: Science in the American West during the Railroad Era. Isis, 109(3):647–649, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lefkaditou:2018:BRM

[2584] Ageliki Lefkaditou. Book review: Marius Turda; Aaron Gillette. Latin Eugenics in Comparative Perspective. Marius Turda, ed. The History of East-Central European Eugenics, 1900–1945: Sources and Commentaries. Isis, 109(3):649–652, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Castelao-Lawless:2018:BRB

[2585] Teresa Castelão-Lawless. Book review: Bernadette Bensaude-Vincent; Eva Telkes-Klein. Les identités multiples d’Émile Meyerson. Isis, 109 (3):652–653, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Abma:2018:BRA

[2586] Ruud Abma. Book review: Aldon D. Morris. The Scholar Denied: W. E. B. Du Bois and the Birth of Modern Sociology. Isis, 109(3):653–654, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Verburgt:2018:BRJ

[2587] Lukas M. Verburgt. Book review: Jan von Plato. The Great Formal Machinery Works: Theories of Deduction and Computation at the Origins of the Digital Age. Isis, 109(3):655–656, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Domingues:2018:BRE

[2588] Heloisa M. Bertol Domingues. Book review: Eve E. Buckley. Technocrats and the Politics of Drought and Development in Twentieth-Century Brazil. Isis, 109(3):656–657, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rupke:2018:BRK

[2589] Nicolaas Rupke. Book review: Krishna Dronamraju. Popularizing Science: The Life and Work of J. B. S. Haldane. Isis, 109(3):657–658, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Krementsov:2018:BRW

[2590] Nikolai Krementsov. Book review: Wladimir Velminski. Homo Sovieticus: Brain Waves, Mind Control, and Telepathic Destiny. Isis, 109(3): 659–660, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Feichtinger:2018:BRW

[2591] Johannes Feichtinger. Book review: Wolfgang U. Eckart; Rainer Gödel, eds. “Krieg der Gelehrten” und die Welt der Akademien 1914–1924. Isis, 109(3):660–661, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Devine:2018:BRS

[2592] Shauna Devine. Book review: Susan L. Smith. Toxic Exposures: Mustard Gas and the Health Consequences of World War II in the United States. Isis, 109(3):661–663, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Walker:2018:BRM

[2593] Mark Walker. Book review: Monique Laney. German Rocketeers in the Heart of Dixie: Making Sense of the Nazi Past during the Civil Rights Era. Isis, 109(3):663–664, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Haalboom:2018:BRS

[2594] Floor Haalboom. Book review: Stuart Blume. Immunization: How Vaccines Became Controversial. Isis, 109(3):664–665, September 2018. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bynum:2018:BRA

[2595] William Bynum. Book review: Aya Homei; Michael Worboys. Fungal Disease in Britain and the United States, 1850–2000: Mycoses and Modernity. Isis, 109(3):666–667, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Peterson:2018:BRS

[2596] Erik L. Peterson. Book review: Susan Merrill Squier. Epigenetic Landscapes: Drawings as Metaphor. Isis, 109(3):667–668, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lovell:2018:BRB

[2597] Stephen Lovell. Book review: Benjamin Peters. How Not to Network a Nation: The Uneasy History of the Soviet Internet. Isis, 109(3):668–669, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blume:2018:BRE

[2598] Stuart Blume. Book review: Elena Conis. Vaccine Nation: America’s Changing Relationship with Immunization. Isis, 109(3):670–671, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Carlson:2018:BRA

[2599] Elof Axel Carlson. Book review: Andrew J. Hogan. Life Histories of Genetic Disease: Patterns and Prevention in Postwar Medical Genetics. Isis, 109(3):671–672, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2018:FCa

[2600] Anonymous. Front cover. Isis, 109(3):??, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2018:Mc

[2601] Anonymous. Masthead. Isis, 109(3):??, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sandman:2018:BRC

[2602] Alison Sandman. Book review: Chet Van Duzer. The World for a King: Pierre Desceliers’ World Map of 1550. Isis, 109(3):??, September 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Morar:2018:RQG

[2603] Florin-Stefan Morar. Relocating the Qing in the global history of science: The Manchu translation of the 1603 world map by Li Yingshi and Matteo Ricci. Isis, 109(4):673–694, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hogan:2018:TCC

[2604] Andrew J. Hogan. The “two cultures” in clinical psychology: Constructing disciplinary divides in the management of mental retardation. Isis, 109(4):695–719, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cittadino:2018:BCP

[2605] Eugene Cittadino. Barry Commoner and Paul Sears on Project Chariot: Epiphany, ecology, and the Atomic Energy Commission. Isis, 109(4): 720–743, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cohen:2018:I

[2606] H. Floris Cohen. Introduction. Isis, 109(4):744–746, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guerrini:2018:TWL

[2607] Anita Guerrini. Translation as a way of life. Isis, 109(4):747–752, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rocke:2018:EK

[2608] Alan J. Rocke. Englishing Kopp. Isis, 109(4):753–759, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hentschel:2018:THS

[2609] Ann M. Hentschel and Klaus Hentschel. Translating in the history of science: a concerted effort. Isis, 109(4):760–766, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reeves:2018:IIT

[2610] Eileen Reeves and Albert Van Helden. Idiom and image: Translating the Letters on Sunspots. Isis, 109(4):767–773, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blaasjo:2018:TM

[2611] Viktor Blåsjö and Jan P. Hogendijk. On translating mathematics. Isis, 109(4):774–781, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Butian:2018:THS

[2612] Zhang Butian. Translating history of science books into Chinese: Why? Which ones? How? Isis, 109(4):782–788, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tsu:2018:HST

[2613] Jing Tsu. Historians of science translating the history of science: Blur versus grit. Isis, 109(4):789–795, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Post:2018:NRL

[2614] Anna-Luna Post and Andreas Weber. Notes on the reviewing of learned websites, digital resources, and tools. Isis, 109(4):796–800, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Endersby:2018:AL

[2615] Jim Endersby. Acknowledging limits. Isis, 109(4):801–805, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wellerstein:2018:MPC

[2616] Alex Wellerstein. The many places of Cold War science. Isis, 109(4): 806–808, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Baronov:2018:SAC

[2617] David Baronov. Shifting agendas and competing interests within public health, science and technology, and medicine in Africa. Isis, 109(4):809– 816, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Levao:2018:BRR

[2618] Ronald Levao. Book review: Roslynn D. Haynes. From Madman to Crime Fighter: The Scientist in Western Culture. Isis, 109(4):817–818, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sysling:2018:BRC

[2619] Fenneke Sysling. Book review: Cathy Gere. Pain, Pleasure, and the Greater Good: From Panopticon to the Skinner Box and Beyond. Isis, 109(4):818–819, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Abi-Rached:2018:BRF

[2620] Joelle M. Abi-Rached. Book review: Fernando Vidal; Francisco Ortega. Being Brains: Making the Cerebral Subject. Isis, 109(4):819–820, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Huisman:2018:BRJ

[2621] Frank Huisman. Book review: Joel Peter Eigen. Mad-Doctors in the Dock: Defending the Diagnosis, 1760–1913. Isis, 109(4):821–822, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bruyns:2018:BRC

[2622] W. F. J. Mörzer Bruyns. Book review: Charles W. J. Withers. Zero Degrees: Geographies of the Prime Meridian. Isis, 109(4):822–823, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Huijnen:2018:BRE

[2623] Pim Huijnen. Book review: Elizabeth Neswald; David F. Smith; Ulrike Thoms, eds. Setting Nutritional Standards: Theory, Policies, Practices. Isis, 109(4):823–824, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Omodeo:2018:BRB

[2624] Pietro Daniel Omodeo. Book review: Bojana Mladenović. Kuhn’s Legacy: Epistemology, Metaphilosophy, and Pragmatism. Isis, 109(4):824–826, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tversky:2018:BRM

[2625] Barbara Tversky. Book review: Matthias Schemmel, ed. Spatial Thinking and External Representation: Towards a Historical Epistemology of Space. Isis, 109(4):826–827, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Yan:2018:BRA

[2626] Hong-Sen Yan. Book review: Alexander Jones. A Portable Cosmos: Revealing the Antikythera Mechanism, Scientific Wonder of the Ancient World. Isis, 109(4):827–828, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Spinelli:2018:BRJ

[2627] Emidio Spinelli. Book review: Jürgen Hammerstaedt; Pierre-Marie Morel; Refik Güremen, eds. Diogenes of Oinoanda: Epicureanism and Philosophical Debates / Diogène d’OEnoanda: Épicurisme et controverses. Isis, 109(4):829–830, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hudecek:2018:BRT

[2628] Jirí Hudecek. Book review: Tina Su Lyn Lim; Donald B. Wagner. The Continuation of Ancient Mathematics: Wang Xiaotong’s Jigu suanjing, Algebra, and Geometry in Seventh-Century China. Isis, 109(4):830–832, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Irannejad:2018:BRS

[2629] Shahrzad Irannejad. Book review: Sonja Brentjes; Jürgen Renn, eds. Globalization of Knowledge in the Post-Antique Mediterranean, 700-1500. Isis, 109(4):832–833, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heering:2018:BRT

[2630] Peter Heering. Book review: Thomas Horst; Marília dos Santos Lopes; Henrique Leitão, eds. Renaissance Craftsmen and Humanistic Scholars: Circulation of Knowledge between Portugal and Germany. Isis, 109(4): 833–834, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Biow:2018:BRA

[2631] Douglas Biow. Book review: Arielle Saiber. Measured Words: Computation and Writing in Renaissance Italy. Isis, 109(4):834–835, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guerrini:2018:BRV

[2632] Anita Guerrini. Book review: Vivian Nutton, ed. Principles of Anatomy according to the Opinion of Galen by Johann Guinter and Andreas Vesalius. Isis, 109(4):836–837, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reeves:2018:BRS

[2633] Eileen Reeves. Book review: Susanna Berger. The Art of Philosophy: Visual Thinking in Europe from the Late Renaissance to the Early Enlightenment. Isis, 109(4):837–838, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Valleriani:2018:BRL

[2634] Matteo Valleriani. Book review: Leslie Tomory. The History of the London Water Industry. Isis, 109(4):838–839, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Phaf-Rheinberger:2018:BRPb

[2635] Ineke Phaf-Rheinberger. Book review: Peter C. Mancall. Nature and Culture in the Early Modern Atlantic. Isis, 109(4):839–840, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mota:2018:BRE

[2636] Bernardo Machado Mota. Book review: Enrique García Santo–Tomás. The Refracted Muse: Literature and Optics in Early Modern Spain. Isis, 109(4):841–842, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hunter:2018:BRW

[2637] Michael Hunter. Book review: William Poole, ed. John Wilkins (1614-1672): New Essays. Isis, 109(4):842–843, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Henry:2018:BRR

[2638] John Henry. Book review: Rob Iliffe. Priest of Nature: The Religious Worlds of Isaac Newton. Isis, 109(4):843–844, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kim:2018:BRY

[2639] Julie Chun Kim. Book review: Yota Batsaki; Sarah Burke Cahalan; Anatole Tchikine, eds. The Botany of Empire in the Long Eighteenth Century. Isis, 109(4):844–845, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brockliss:2018:BRM

[2640] Laurence Brockliss. Book review: Margaret DeLacy. Contagionism Catches On: Medical Ideology in Britain, 1730-1800. Isis, 109(4):845– 846, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Oosterhoff:2018:BRP

[2641] Richard J. Oosterhoff. Book review: Paola Bertucci. Artisanal Enlightenment: Science and the Mechanical Arts in Old Regime France. Isis, 109 (4):847–848, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Talas:2018:BRE

[2642] Sofia Talas. Book review: Ethan W. Lasser, ed. The Philosophy Chamber: Art and Science in Harvard’s Teaching Cabinet, 1766-1820. Isis, 109(4):

848–849, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bennett:2018:BRR

[2643] Jim Bennett. Book review: Roderick W. Home; Isabel M. Malaquias; Manuel F. Thomaz, eds. For the Love of Science: The Correspondence of J. H. de Magellan (1722-1790). Isis, 109(4):849–850, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Richards:2018:BRC

[2644] Robert J. Richards. Book review: Christine Lehleiter. Romanticism, Origins, and the History of Heredity. Isis, 109(4):850–851, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jones:2018:BRA

[2645] Matthew L. Jones. Book review: Alberto Cevolini, ed. Forgetting Machines: Knowledge Management Evolution in Early Modern Europe. Isis, 109(4):852–853, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Maerker:2018:BRL

[2646] Anna Maerker. Book review: Lucia Dacome. Malleable Anatomies: Models, Makers, and Material Culture in Eighteenth-Century Italy. Isis, 109 (4):853–854, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Terrall:2018:BRS

[2647] Mary Terrall. Book review: Sarah Easterby-Smith. Cultivating Commerce: Cultures of Botany in Britain and France, 1760-1815. Isis, 109 (4):854–855, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Plunkett:2018:BRC

[2648] John Plunkett. Book review: Carin Berkowitz; Bernard Lightman, eds. Science Museums in Transition: Cultures of Display in Nineteenth-Century Britain and America. Isis, 109(4):855–857, December 2018. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dean:2018:BRJ

[2649] Sheila Ann Dean. Book review: J. David Archibald. Origins of Darwin’s Evolution: Solving the Species Puzzle through Time and Place. Isis, 109 (4):857–858, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Conlin:2018:BREb

[2650] Jonathan Conlin. Book review: Edward J. Gillin. The Victorian Palace of Science: Scientific Knowledge and the Building of the Houses of Parliament. Isis, 109(4):858–859, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hopwood:2018:BRE

[2651] Nick Hopwood. Book review: Ernst Haeckel; Ausgewählte Briefwechsel: Historisch-Kritische Ausgabe, Vol. 1: Familienkorrespondenz, Februar 1839-April 1854. Isis, 109(4):860–861, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hollin:2018:BRS

[2652] Greg Hollin. Book review: Stephen T. Casper; Delia Gavrus, eds. The History of the Brain and Mind Sciences: Technique, Technology, Therapy. Isis, 109(4):861–862, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Smith:2018:BRD

[2653] Kerry Smith. Book review: David G. Wittner; Philip C. Brown, eds. Science, Technology, and Medicine in the Modern Japanese Empire. Isis, 109 (4):862–863, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schnadelbach:2018:BRA

[2654] Sandra Schnädelbach. Book review: Andreas Killen. Homo Cinematicus: Science, Motion Pictures, and the Making of Modern Germany. Isis, 109 (4):864–865, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mormann:2018:BRK

[2655] Thomas Mormann. Book review: Karl Sigmund. Exact Thinking in Demented Times: The Vienna Circle and the Epic Quest for the Foundations of Science. Isis, 109(4):865–866, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hogan:2018:BRK

[2656] Andrew J. Hogan. Book review: Karl S. Matlin; Jane Maienschein; Manfred D. Laubichler, eds. Visions of Cell Biology: Reflections Inspired by Cowdry’s ”General Cytology.”. Isis, 109(4):866–868, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Riess:2018:BRW

[2657] Falk Rieß. Book review: Wilhelm Füßl; Andrea Lucas; Matthias Röschner. Wirklichkeit und Illusion: Die Dioramen im Deutschen Museum. Isis, 109(4):868–869, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Walloch:2018:BRJ

[2658] Karen Walloch. Book review: Janet Weston. Medicine, the Penal System, and Sexual Crimes in England, 1919-1960s: Diagnosing Deviance. Isis, 109(4):869–870, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hochadel:2018:BRR

[2659] Oliver Hochadel. Book review: Ryan A. Davis; Alicia Cerezo Paredes, eds. Modernity and Epistemology in Nineteenth-Century Spain: Fringe Discourses. Isis, 109(4):870–871, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barberis:2018:BRM

[2660] Daniela S. Barberis. Book review: Michael R. Finn. Figures of the Pre-Freudian Unconscious from Flaubert to Proust. Isis, 109(4):872–873, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Meneghello:2018:BRI

[2661] Laura Meneghello. Book review: Ian Hesketh. Victorian Jesus: J. R. Seeley, Religion, and the Cultural Significance of Anonymity. Isis, 109 (4):873–874, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Finkelstein:2018:BRL

[2662] Gabriel Finkelstein. Book review: Laura Meneghello. Jacob Moleschott, a Transnational Biography: Science, Politics, and Popularization in Nineteenth-Century Europe. Isis, 109(4):874–875, December 2018. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harman:2018:BRJ

[2663] Oren Harman. Book review: Jeroen van Dongen; Herman Paul, eds. Epistemic Virtues in the Sciences and the Humanities. Isis, 109(4):875– 876, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vennen:2018:BRB

[2664] Mareike Vennen. Book review: Brian Noble. Articulating Dinosaurs: a Political Anthropology. Isis, 109(4):876–878, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Allen:2018:BRM

[2665] Garland E. Allen. Book review: Michelle Murphy. The Economization of Life. Isis, 109(4):878–879, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Raposo:2018:BRW

[2666] Pedro M. P. Raposo. Book review: William Firebrace. Star Theatre: The Story of the Planetarium. Isis, 109(4):879–880, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rankin:2018:BRJ

[2667] William Rankin. Book review: Jess Bier. Mapping Israel, Mapping Palestine: How Occupied Landscapes Shape Scientific Knowledge. Isis, 109 (4):880–882, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schalenberg:2018:BRS

[2668] Marc Schalenberg. Book review: Svea Koischwitz. Der Bund Freiheit der Wissenschaft in den Jahren 1970-1976: Ein Interessenverband zwischen Studentenbewegung und Hochschulreform. Isis, 109(4):882–883, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Carlson:2018:BRS

[2669] Elof Axel Carlson. Book review: Stephen Hilgartner. Reordering Life: Knowledge and Control in the Genomics Revolution. Isis, 109(4):883– 884, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hilgartner:2018:BRS

[2670] Stephen Hilgartner. Book review: Sabina Leonelli. Data-Centric Biology: a Philosophical Study. Isis, 109(4):884–885, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lalli:2018:BRH

[2671] Roberto Lalli. Book review: Harry Collins. Gravity’s Kiss: The Detection of Gravitational Waves. Isis, 109(4):885–887, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Healey:2018:BRI

[2672] Jenna Healey. Book review: Ilana Löwy. Imperfect Pregnancies: a History of Birth Defects and Prenatal Diagnosis. Isis, 109(4):887–888, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2018:IIV

[2673] Anonymous. Index to Isis, volume 109, 2018. Isis, 109(4):889–905, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2018:CIV

[2674] Anonymous. Contents of Isis, volume 109. Isis, 109(4):906–908, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2018:FCb

[2675] Anonymous. Front cover. Isis, 109(4):??, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2018:Md

[2676] Anonymous. Masthead. Isis, 109(4):??, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stolzenberg:2019:HOR

[2677] Daniel Stolzenberg. The Holy Office in the Republic of Letters: Roman censorship, Dutch atlases, and the European information order, circa 1660. Isis, 110(1):1–23, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lempert:2019:FGA

[2678] Michael Lempert. Fine-grained analysis: Talk therapy, media, and the microscopic science of the face-to-face. Isis, 110(1):24–47, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rouphail:2019:CES

[2679] Robert M. Rouphail. Cyclonic ecology: Sugar, cyclone science, and the limits of empire in Mauritius and the Indian Ocean world, 1870s–1930s. Isis, 110(1):48–67, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dahn:2019:BSN

[2680] Ryan Dahn. Big science, nazified? Pascual Jordan, Adolf Meyer-Abich, and the abortive scientific journal Physis. Isis, 110(1):68–90, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cohen:2019:EIa

[2681] H. Floris Cohen. Editor’s introduction. Isis, 110(1):91–93, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hsia:2019:HST

[2682] Florence Hsia and Dagmar Schäfer. History of science, technology, and medicine: a second look at Joseph Needham. Isis, 110(1):94–99, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nappi:2019:RNN

[2683] Carla Nappi and McKenzie Wark. Reading Needham now. Isis, 110(1): 100–108, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chemla:2019:NIC

[2684] Karine Chemla. Needham and the issue of Chinese as a language for science: Taking a linguistic turn materially. Isis, 110(1):109–115, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kuo:2019:EMY

[2685] Wen-Hua Kuo. An ecumenical medicine yet to come: Reflections on Needham on medicine. Isis, 110(1):116–121, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chen:2019:NMF

[2686] BuYun Chen. Needham, matter, form, and us. Isis, 110(1):122–128, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tilley:2019:GSD

[2687] Helen Tilley. A great (scientific) divergence: Synergies and fault lines in global histories of science. Isis, 110(1):129–136, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Golinski:2019:IT

[2688] Jan Golinski. The industrious Tyndall. Isis, 110(1):137–141, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Browman:2019:BRG

[2689] David L. Browman. Book review: Glynn Custred. A History of Anthropology as a Holistic Science. Isis, 110(1):142–143, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nacol:2019:BRL

[2690] Emily C. Nacol. Book review: Laura Ephraim. Who Speaks for Nature? On the Politics of Science. Isis, 110(1):143–144, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Williams:2019:BRN

[2691] Rosalind Williams. Book review: Nasser Zakariya. A Final Story: Science, Myth, and Beginnings. Isis, 110(1):144–145, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kaiser:2019:BRS

[2692] Marie I. Kaiser. Book review: Scott Lidgard; Lynn K. Nyhart, eds. Biological Individuality: Integrating Scientific, Philosophical, and Historical Perspectives. Isis, 110(1):146–147, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Abma:2019:BRM

[2693] Ruud Abma. Book review: Maarten Derksen. Histories of Human Engineering: Tact and Technology. Isis, 110(1):147–148, March 2019. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Talbot:2019:BRC

[2694] Ann Talbot. Book review: Caspar Henderson. A New Map of Wonders: a Journey in Search of Modern Marvels. Isis, 110(1):148–149, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Quan:2019:BRL

[2695] Tang Quan. Book review: Liu Ciyuan [Chinese text]. A Critical Examination of the Astronomical Records in the Dynastic Histories. Isis, 110(1):150–151, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jacquart:2019:BRJ

[2696] Danielle Jacquart. Book review: John C. Lamoreaux, ed. and trans. H .unayn ibn Ish.aq on His Galen Translations: a Parallel English-Arabic Text. Isis, 110(1):151–152, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Grell:2019:BRJ

[2697] Chantal Grell. Book review: Jean-Patrice Boudet; Martine Ostorero; Agostino Paravicini Bagliani, eds. De Frédéric II à Rodolphe II: Astrologie, divination et magie dans les cours (XIIIe–XVIIe siècle). Isis, 110 (1):152–154, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Goulding:2019:BRM

[2698] Robert Goulding. Book review: Marcia Kupfer. Art and Optics in the Hereford Map: an English Mappa Mundi, c. 1300. Isis, 110(1):154– 155, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Frijhoff:2019:BRJ

[2699] Willem Frijhoff. Book review: Jan-Hendryk de Boer; Marian Füssel; Maximilian Schuh, eds. Universitäre Gelehrtenkultur vom 13.-16. Jahrhundert: Ein interdisziplinäres Quellen- und Methodenhandbuch. Isis, 110(1):155–156, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blum:2019:BRG

[2700] Paul Richard Blum. Book review: Giordano Bruno. The Ash Wednesday Supper. Isis, 110(1):156–157, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rizvi:2019:BRP

[2701] Sajjad Rizvi. Book review: Peter Adamson; Peter E. Pormann, eds. Philosophy and Medicine in the Formative Period of Islam. Isis, 110(1): 157–158, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Prins:2019:BRA

[2702] Jacomien Prins. Book review: Andrew Hicks. Composing the World: Harmony in the Medieval Platonic Cosmos. Isis, 110(1):159–160, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vanpaemel:2019:BRG

[2703] Geert Vanpaemel. Book reviews: Guillaume Carnino; Liliane Hilaire– Pérez; Jochen Hoock, eds. La technologie générale: Johann Beckmann; Stéphane Blond; Liliane Hilaire–Pérez; Michel Virol, eds. Mobilités d’ingénieurs en Europe, XVe–XVIIIe siècle: Mélanges en l’honneur d’Hélène Vérin. Isis, 110(1):160–161, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wiesner-Hanks:2019:BRG

[2704] Merry Wiesner-Hanks. Book review: Gabrielle Robilliard. Tending Mothers and the Fruits of the Womb: The Work of the Midwife in the Early Modern German City. Isis, 110(1):161–162, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rocknak:2019:BRT

[2705] Stefanie Rocknak. Book review: Tamás Demeter. David Hume and the Culture of Scottish Newtonianism: Methodology and Ideology in Enlightenment Inquiry. Isis, 110(1):163–164, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sylla:2019:BRD

[2706] Edith Dudley Sylla. Book review: Deborah Boyle. The Well-Ordered Universe: The Philosophy of Margaret Cavendish. Isis, 110(1):164–165, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Grandi:2019:BRT

[2707] Giovanni B. Grandi. Book review: Thomas Reid. Thomas Reid on Mathematics and Natural Philosophy. Isis, 110(1):165–166, March 2019. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vermij:2019:BRD

[2708] Rienk Vermij. Book review: Dirk van Miert; Henk Nellen; Piet Steenbakkers; Jetze Touber, eds. Scriptural Authority and Biblical Criticism in the Dutch Golden Age: God’s Word Questioned. Isis, 110(1):166–167, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Henry:2019:BRE

[2709] John Henry. Book review: Elizabethanne Boran; Mordechai Feingold, eds. Reading Newton in Early Modern Europe. Isis, 110(1):168–169, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Block:2019:BRR

[2710] Kristen Jean Block. Book review: Rana A. Hogarth. Medicalizing Blackness: Making Racial Difference in the Atlantic World, 1780-1840. Isis, 110(1):169–170, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weststeijn:2019:BRG

[2711] Thijs Weststeijn. Book review: Gero Seelig. Medusa’s Menagerie: Otto Marseus van Schrieck and the Scholars. Isis, 110(1):170–171, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bos:2019:BRH

[2712] Erik-Jan Bos. Book review: Harold J. Cook. The Young Descartes: Nobility, Rumor, and War. Isis, 110(1):172–173, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shea:2019:BRA

[2713] William R. Shea. Book review: Alberto A. Martínez. Burned Alive: Giordano Bruno, Galileo, and the Inquisition. Isis, 110(1):173–174, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

deLeeuw:2019:BRK

[2714] Karl de Leeuw. Book review: Katherine Ellison. A Cultural History of Early Modern English Cryptography Manuals. Isis, 110(1):174–175, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lowther:2019:BRJ

[2715] David Lowther. Book review: James Delbourgo. Collecting the World: The Life and Curiosity of Hans Sloane. Isis, 110(1):175–176, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Timberlake:2019:BRJ

[2716] Todd Timberlake. Book review: Johannes Kepler. Nova stereometria doliorum vinariorum / New Solid Geometry of Wine Barrels. Isis, 110 (1):177–178, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Paradis:2019:BRP

[2717] James Paradis. Book reviews: Piers J. Hale. Political Descent: Malthus, Mutualism, and the Politics of Evolution in Victorian England; Rob Boddice. The Science of Sympathy: Morality, Evolution, and Victorian Civilization. Isis, 110(1):178–180, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Seifer:2019:BRJ

[2718] Marc J. Seifer. Book review: Jennifer L. Lieberman. Power Lines: Electricity in American Life and Letters, 1882-1952. Isis, 110(1):181–182, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gambaro:2019:BRR

[2719] Ivana Gambaro. Book review: Raffaella Simili, ed. Umanisti e presidenti: L’Accademia Nazionale dei Lincei (1900-1933). Isis, 110(1):182– 183, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kosmann-Schwarzbach:2019:BRL

[2720] Yvette Kosmann-Schwarzbach. Book review: Lizhen Ji; Athanase Papadopoulos; Sumio Yamada, eds. From Riemann to Differential Geometry and Relativity. Isis, 110(1):183–184, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mayer:2019:BRP

[2721] Andreas Mayer. Book review: Philip Kuhn. Psychoanalysis in Britain, 1893-1913: Histories and Historiography. Isis, 110(1):185–186, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Simon:2019:BRJ

[2722] Rodolfo Gutiérrez Simón. Book review: Juan Vicente Mayoral. Thomas S. Kuhn: La búsqueda de la estructura. Isis, 110(1):186–187, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeVorkin:2019:BRR

[2723] David H. DeVorkin. Book review: Richard Taibi. Charles Olivier and the Rise of Meteor Science. Isis, 110(1):187–188, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wilson:2019:BRJ

[2724] Evan Wilson. Book review: Jason W. Smith. To Master the Boundless Sea: The U.S. Navy, the Marine Environment, and the Cartography of Empire. Isis, 110(1):188–189, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Richmond:2019:BRP

[2725] Marsha L. Richmond. Book review: Patricia Fara. A Lab of One’s Own: Science and Suffrage in the First World War. Isis, 110(1):189– 190, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hale:2019:BRM

[2726] Piers J. Hale. Book review: Michael Boulter. Bloomsbury Scientists: Science and Art in the Wake of Darwin. Isis, 110(1):191–192, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Coen:2019:BRS

[2727] Deborah R. Coen. Book review: Stefani Engelstein. Sibling Action: The Genealogical Structure of Modernity. Isis, 110(1):192–193, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barrow:2019:BRH

[2728] Mark V. Barrow, Jr. Book review: Henry A. McGhie. Henry Dresser and Victorian Ornithology: Birds, Books, and Business. Isis, 110(1): 193–194, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mahony:2019:BRJ

[2729] Martin Mahony. Book review: Jamie L. Pietruska. Looking Forward: Prediction and Uncertainty in Modern America. Isis, 110(1):194–196, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Close-Koenig:2019:BRK

[2730] Tricia Close-Koenig. Book review: Kaat Wils; Raf de Bont; Sokhieng Au, eds. Bodies beyond Borders: Moving Anatomies, 1750-1950. Isis, 110(1):196–197, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Forstner:2019:BRP

[2731] Christian Forstner. Book review: Paul Halpern. The Quantum Labyrinth: How Richard Feynman and John Wheeler Revolutionized Time and Reality. Isis, 110(1):197–198, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Theunissen:2019:BRR

[2732] Bert Theunissen. Book review: Rebecca J. H. Woods. The Herds Shot Round the World: Native Breeds and the British Empire, 1800-1900. Isis, 110(1):198–199, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ruse:2019:BRP

[2733] Michael Ruse. Book review: Peter J. Woodford. The Moral Meaning of Nature: Nietzsche’s Darwinian Religion and Its Critics. Isis, 110(1):200– 201, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rowe:2019:BRH

[2734] David E. Rowe. Book reviews: Hanoch Gutfreund; Jürgen Renn. The Formative Years of Relativity: The History and Meaning of Einstein’s Princeton Lectures; Galina Weinstein. Einstein’s Pathway to the Special Theory of Relativity. Isis, 110(1):201–204, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bauer:2019:BRV

[2735] Heike Bauer. Book review: Veronika Fuechtner; Douglas E. Haynes; Ryan M. Jones, eds. A Global History of Sexual Science, 1880–1960. Isis, 110(1):204–205, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Williams:2019:BRR

[2736] Amrys O. Williams. Book review: Roger L. Williams. Evan Pugh’s Penn State: America’s Model Agricultural College. Isis, 110(1):205–206, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ishibashi:2019:BRY

[2737] Yuto Ishibashi. Book review: Yulia Frumer. Making Time: Astronomical Time Measurement in Tokugawa Japan. Isis, 110(1):207–208, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roll-Hansen:2019:BRH

[2738] Nils Roll-Hansen. Book review: Hans-Jörg Rheinberger; Staffan Müller– Wille. The Gene from Genetics to Postgenomics. Isis, 110(1):208–209, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Launius:2019:BRN

[2739] Roger D. Launius. Book review: Nathalia Holt. Rise of the Rocket Girls: The Women Who Propelled Us, from Missiles to the Moon to Mars. Isis, 110(1):209–210, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hepler-Smith:2019:BRA

[2740] Evan Hepler-Smith. Book review: Anthony Chaney. Runaway: Gregory Bateson, the Double Bind, and the Rise of Ecological Consciousness. Isis, 110(1):211–212, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Snelders:2019:BRV

[2741] Stephen Snelders. Book review: Vivek Neelakantan. Science, Public Health, and Nation-Building in Soekarno-Era Indonesia. Isis, 110(1): 212–213, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barrett:2019:BRA

[2742] Jeffrey A. Barrett. Book review: Adam Becker. What Is Real? The Unfinished Quest for the Meaning of Quantum Physics. Isis, 110(1):213– 214, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Townsend:2019:BRD

[2743] Robert B. Townsend. Book review: Donald B. Kraybill. Eastern Mennonite University: a Century of Countercultural Education. Isis, 110(1): 216–217, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barth:2019:BRH

[2744] Howard G. Barth. Book review: Helge Kragh. From Transuranic to Superheavy Elements: a Story of Dispute and Creation. Isis, 110(1):217– 218, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sheldon:2019:BRS

[2745] Myrna Perez Sheldon. Book review: Susan L. Trollinger; William Vance Trollinger. Righting America at the Creation Museum. Isis, 110(1):218– 219, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gilleran:2019:BRS

[2746] Sean Gilleran. Book review: Sigrid Schmalzer; Daniel S. Chard; Alyssa Botelho, eds. Science for the People: Documents from America’s Movement of Radical Scientists. Isis, 110(1):220–221, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2019:FCa

[2747] Anonymous. Front cover. Isis, 110(1):??, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2019:Ma

[2748] Anonymous. Masthead. Isis, 110(1):??, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Freire:2019:BRJ

[2749] Olival Freire Jr. Book review: Jennifer Burwell. Quantum Language and the Migration of Scientific Concepts. Isis, 110(1):??, March 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cohen:2019:EFR

[2750] H. Floris Cohen. Editorial: Farewell reflections. Isis, 110(2):223–227, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lea:2019:CDK

[2751] Andrew Lea. Computerizing diagnosis: Keeve Brodman and the medical data screen. Isis, 110(2):228–249, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Burton:2019:RCR

[2752] Elise K. Burton. Red crescents: Race, genetics, and sickle cell disease in the Middle East. Isis, 110(2):250–269, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Katsos:2019:CLW

[2753] Isidoros Katsos. Chasing the light: What happened to the ancient theories? Isis, 110(2):270–282, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). See response [2754].

Smith:2019:SLR

[2754] A. Mark Smith. Seeing the light: a response to “Chasing the Light”. Isis, 110(2):283–289, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). See [2753].

Cohen:2019:EIb

[2755] H. Floris Cohen. Editor’s introduction. Isis, 110(2):290–291, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Steele:2019:EBA

[2756] John Steele. Explaining Babylonian astronomy. Isis, 110(2):292–295, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Westman:2019:HDC

[2757] Robert S. Westman. How did Copernicus become a Copernican? Isis, 110(2):296–301, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Portuondo:2019:EEM

[2758] María M. Portuondo. Explaining an early modern Magnum Opus. Isis, 110(2):302–304, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heilbron:2019:EA

[2759] J. L. Heilbron. Have I explained anything? Isis, 110(2):305–307, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schuster:2019:POC

[2760] John A. Schuster. Pitfalls and opportunities of contextual explanation: The case of Isaac Beeckman’s invention of the mechanical philosophy. Isis, 110(2):308–311, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shank:2019:HCB

[2761] J. B. Shank. How calculus-based mathematical physics arose in France after 1700: a historicized actor–network narrative as explanation. Isis, 110(2):312–316, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bycroft:2019:HTN

[2762] Michael Bycroft. Hindsight, tradition, and national styles in the prehistory of the Leyden jar. Isis, 110(2):317–320, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanLunteren:2019:HEC

[2763] Frans van Lunteren. Historical explanation and causality. Isis, 110(2): 321–324, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vetter:2019:ESC

[2764] Jeremy Vetter. Explaining structural constraints on lay participation in field science. Isis, 110(2):325–327, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weidman:2019:CRB

[2765] Nadine Weidman. Cultural relativism and biological determinism: a problem in historical explanation. Isis, 110(2):328–331, June 2019. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stark:2019:E

[2766] Laura Stark. Emergence. Isis, 110(2):332–336, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Curry:2019:WSS

[2767] Helen Anne Curry. Why save a seed? Isis, 110(2):337–340, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Theunissen:2019:OF

[2768] Bert Theunissen. The Oostvaardersplassen Fiasco. Isis, 110(2):341– 345, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guldi:2019:MPE

[2769] Jo Guldi. The modern paradigms of explanation: Significance, agency, and writing history in the era of climate change. Isis, 110(2):346–353, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Arabatzis:2019:ESH

[2770] Theodore Arabatzis. Explaining science historically. Isis, 110(2):354– 359, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2019:HSS

[2771] Anonymous. History of Science Society Annual Meeting, 2018. Isis, 110(2):360–369, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sigurdsson:2019:ESS

[2772] Skúli Sigurdsson. Eloge: Silvan Samuel Schweber (1928–2017). Isis, 110(2):370–373, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gibbs:2019:IFR

[2773] Fred Gibbs. Interfacing facsimiles: a review of the Berlin State Library’s Alexander von Humboldt Portal. Isis, 110(2):375–376, June 2019. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gaukroger:2019:LDH

[2774] Stephen Gaukroger. Longue durée history: Science to the rescue? Isis, 110(2):377–380, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Yeo:2019:BRJ

[2775] Richard Yeo. Book review: John Willinsky. The Intellectual Properties of Learning: a Prehistory from Saint Jerome to John Locke. Isis, 110(2): 381–382, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bresadola:2019:BRD

[2776] Marco Bresadola. Book review: Domenico Bertoloni Meli. Visualizing Disease: The Art and History of Pathological Illustrations. Isis, 110(2): 382–383, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wachelder:2019:BRK

[2777] Joseph Wachelder. Book reviews: Klaas van Berkel. Universiteit van het Noorden: Vier Eeuwen Academisch Leven in Groningen. Volume 1: De Oude Universiteit, 1614–1876; Klaas van Berkel. Universiteit van het Noorden: Vier Eeuwen Academisch Leven in Groningen. Volume 2: De Klassieke Universiteit, 1876–1945. Isis, 110(2):383–386, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mildenberger:2019:BRA

[2778] Florian G. Mildenberger. Book review: Alexa Geisthövel; Volker Hess, eds. Medizinisches Gutachten: Geschichte einer neuzeitlichen Praxis. Isis, 110(2):386–387, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Karstens:2019:BRE

[2779] Bart Karstens. Book review: Eric Scerri. A Tale of Seven Scientists and a New Philosophy of Science. Isis, 110(2):387–388, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mayhew:2019:BRA

[2780] Robert Mayhew. Book review: Andy Merrills. Roman Geographies of the Nile: From the Late Republic to the Early Empire. Isis, 110(2):389– 390, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Leszl:2019:BRJ

[2781] Walter Leszl. Book review: Julius Rocca, ed. Teleology in the Ancient World: Philosophical and Medical Approaches. Isis, 110(2):390–391, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wakeling:2019:BRC

[2782] Will Wakeling. Book review: Christopher Cullen. Heavenly Numbers: Astronomy and Authority in Early Imperial China. Isis, 110(2):391– 392, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brown:2019:BRD

[2783] Miranda Brown. Book review: Daniel Patrick Morgan. Astral Sciences in Early Imperial China. Isis, 110(2):392–393, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cormack:2019:BRR

[2784] Lesley B. Cormack. Book review: Roel Nicolai. The Enigma of the Origin of Portolan Charts: a Geodetic Analysis of the Hypothesis of a Medieval Origin. Isis, 110(2):394–395, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vermij:2019:BRH

[2785] Rienk Vermij. Book review: Hans Bots. De Republiek der Letteren: De Europese intellectuele wereld, 1500–1760. Isis, 110(2):395–396, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Long:2019:BRM

[2786] Pamela O. Long. Book review: Michael Wintroub. The Voyage of Thought: Navigating Knowledge across the Sixteenth-Century World. Isis, 110(2):396–397, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Baker:2019:BRA

[2787] Tawrin Baker. Book review: Arianna Borrelli; Giora Hon; Yaakov Zik, eds. The Optics of Giambattista Della Porta (ca. 1535–1615): a Reassessment. Isis, 110(2):397–399, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Romeiras:2019:BRB

[2788] Francisco Malta Romeiras. Book review: Bradford A. Bouley. Pious Postmortems: Anatomy, Sanctity, and the Catholic Church in Early Modern Europe. Isis, 110(2):399–400, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barker:2019:BRA

[2789] Peter Barker. Book review: Aviva Rothman. The Pursuit of Harmony: Kepler on Cosmos, Confession, and Community. Isis, 110(2):400–401, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Boyle:2019:BRH

[2790] Jen Boyle. Book review: Howard Marchitello; Evelyn Tribble, eds. The Palgrave Handbook of Early Modern Literature and Science. Isis, 110(2): 401–403, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Osterhage:2019:BRP

[2791] Wolfgang Osterhage. Book review: Paolo Galluzzi. The Lynx and the Telescope: The Parallel Worlds of Federico Cesi and Galileo. Isis, 110 (2):403–404, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

VanHelden:2019:BRC

[2792] Albert Van Helden. Book review: Christopher M. Graney. Mathematical Disquisitions: The Booklet of Theses Immortalized by Galileo. Isis, 110 (2):404–405, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanBerkel:2019:BRD

[2793] Klaas van Berkel. Book review: David S. Sytsma. Richard Baxter and the Mechanical Philosophers. Isis, 110(2):405–406, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ross:2019:BRP

[2794] Alan Ross. Book review: Peter Sahlins. 1668: The Year of the Animal in France. Isis, 110(2):406–408, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Arthur:2019:BRG

[2795] Richard T. W. Arthur. Book review: G. W. Leibniz. The Leibniz– Arnauld Correspondence: With Selections from the Correspondence with Ernst, Landgrave of Hessen–Rheinfels; G. W. Leibniz. The Leibniz–Stahl Controversy. Isis, 110(2):408–410, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weiss:2019:BRD

[2796] Martin P. M. Weiss. Book review: A. D. Morrison-Low; Sara J. Schechner; Paolo Brenni, eds. How Scientific Instruments Have Changed Hands. Isis, 110(2):410–411, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stockhorst:2019:BRS

[2797] Stefanie Stockhorst. Book review: Sietske Fransen; Niall Hodson; Karl A. E. Enenkel, eds. Translating Early Modern Science. Isis, 110(2):411– 412, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Demeter:2019:BRJ

[2798] Tamás Demeter and Gábor Áron Zemplén. Book review: John H. Zammito. The Gestation of German Biology: Philosophy and Physiology from Stahl to Schelling. Isis, 110(2):413–414, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anthony:2019:BRJ

[2799] Patrick Anthony. Book review: Julian Drews; Ottmar Ette; Tobias Kraft; Barbara Schneider-Kempf; Jutta Weber, eds. Forster-Humboldt-Chamisso: Weltreisende im Spannungsfeld der Kulturen. Isis, 110(2): 414–415, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hess:2019:BRP

[2800] Volker Hess. Book review: Peter Cryle; Elizabeth Stephens. Normality: a Critical Genealogy. Isis, 110(2):416–417, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Goldsmith:2019:BRG

[2801] Mike Goldsmith. Book review: Graeme Gooday; Karen Sayer. Managing the Experience of Hearing Loss in Britain, 1830–1930. Isis, 110(2):417– 418, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barnes:2019:BRM

[2802] David S. Barnes. Book review: Melanie A. Kiechle. Smell Detectives: an Olfactory History of Nineteenth-Century Urban America. Isis, 110 (2):418–420, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Talairach:2019:BRD

[2803] Laurence Talairach. Book review: David N. Stamos. Edgar Allan Poe, “Eureka”, and Scientific Imagination. Isis, 110(2):420–421, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Simpson:2019:BRD

[2804] Julian M. Simpson. Book review: Douglas M. Haynes. Fit to Practice: Empire, Race, Gender, and the Making of British Medicine, 1850-1980. Isis, 110(2):421–422, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Home:2019:BRT

[2805] R. W. Home. Book review: Timothy P. Barnard. Nature’s Colony: Empire, Nation, and Environment in the Singapore Botanic Gardens. Isis, 110(2):422–423, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wiseman:2019:BRA

[2806] Matthew S. Wiseman. Book reviews: Alan I. Marcus, ed. Science as Service: Establishing and Reformulating American Land-Grant Universities, 1865–1930; Alan I. Marcus, ed. Service as Mandate: How American Land-Grant Universities Shaped the Modern World, 1920–2015. Isis, 110(2):424–425, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Efron:2019:BRP

[2807] Noah Efron. Book review: Patrizia Guarnieri. Italian Psychology and Jewish Emigration under Fascism: From Florence to Jerusalem and New York. Isis, 110(2):426–427, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kennedy:2019:BRJ

[2808] Juliette Kennedy. Book review: Janet L. Beery; Sarah J. Greenwald; Jacqueline A. Jensen-Vallin; Maura B. Mast, eds. Women in Mathematics: Celebrating the Centennial of the Mathematical Association of America. Isis, 110(2):427–428, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Moran:2019:BRB

[2809] Jeffrey P. Moran. Book review: Bernard Lightman, ed. Global Spencerism: The Communication and Appropriation of a British Evolutionist. Isis, 110(2):428–429, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Calinger:2019:BRG

[2810] Ronald Calinger. Book review: Gertraud Sturm. David Heinrich Müller und die südarabische Expedition der Kaiserlichen Akademie der Wissenschaften 1898/99. Isis, 110(2):429–431, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guttler:2019:BRT

[2811] Nils Güttler. Book review: Tobias Scheidegger. “Petite Science”: Außeruniversitäre Naturforschung in der Schweiz um 1900. Isis, 110(2): 431–432, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kinney:2019:BRT

[2812] Jeremy R. Kinney. Book review: Timothy P. Schultz. The Problem with Pilots: How Physicians, Engineers, and Airpower Enthusiasts Redefined Flight. Isis, 110(2):432–433, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cittadino:2019:BRN

[2813] Eugene Cittadino. Book review: Nathan F. Sayre. The Politics of Scale: a History of Rangeland Science. Isis, 110(2):433–434, June 2019. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Boddice:2019:BRS

[2814] Rob Boddice. Book review: Stefanos Geroulanos; Todd Meyers. The Human Body in the Age of Catastrophe: Brittleness, Integration, Science, and the Great War. Isis, 110(2):435–436, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brain:2019:BRM

[2815] Stephen Brain. Book review: Maria A. Rogacheva. The Private World of Soviet Scientists from Stalin to Gorbachev. Isis, 110(2):436–437, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Westfall:2019:BRC

[2816] Catherine Westfall. Book review: C. Bruce Tarter. The American Lab: an Insider’s History of the Lawrence Livermore National Laboratory. Isis, 110(2):437–438, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mazzeo:2019:BRN

[2817] Agata Mazzeo. Book review: Noémi Tousignant. Edges of Exposure: Toxicology and the Problem of Capacity in Postcolonial Senegal. Isis, 110(2):438–440, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2019:FCb

[2818] Anonymous. Front cover. Isis, 110(2):??, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2019:Mb

[2819] Anonymous. Masthead. Isis, 110(2):??, June 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schmalzer:2019:LE

[2820] Sigrid Schmalzer. Letter to the Editor. Isis, 110(2):??, June 2019. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Virastau:2019:CRW

[2821] Nicolae Virastau. Calendar reform and world chronology: Pierre De Lille’s Tria Calendaria Parva (1529). Isis, 110(3):441–459, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hesketh:2019:TSS

[2822] Ian Hesketh. Technologies of the scientific self: John Tyndall and his journal. Isis, 110(3):460–482, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bod:2019:FCG

[2823] Rens Bod, Jeroen van Dongen, Sjang L. ten Hagen, Bart Karstens, and Emma Mojet. The flow of cognitive goods: a historiographical framework for the study of epistemic transfer. Isis, 110(3):483–496, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gibson:2019:I

[2824] Abraham Gibson, Manfred D. Laubichler, and Jane Maienschein. Introduction. Isis, 110(3):497–501, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Laubichler:2019:CHK

[2825] Manfred D. Laubichler, Jane Maienschein, and Jürgen Renn. Computational history of knowledge: Challenges and opportunities. Isis, 110 (3):502–512, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Damerow:2019:HGD

[2826] Julia Damerow and Dirk Wintergrün. The hitchhiker’s guide to data in the history of science. Isis, 110(3):513–521, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Aiello:2019:THU

[2827] Kenneth D. Aiello and Michael Simeone. Triangulation of history using textual data. Isis, 110(3):522–537, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Painter:2019:NAD

[2828] Deryc T. Painter, Bryan C. Daniels, and Jürgen Jost. Network analysis for the digital humanities: Principles, problems, extensions. Isis, 110 (3):538–554, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gibson:2019:HSS

[2829] Abraham Gibson and Cindy Ermus. The history of science and the science of history: Computational methods, algorithms, and the future of the field. Isis, 110(3):555–566, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Uchacz:2019:EML

[2830] Tianna Helena Uchacz. Early Modern Letters Online (EMLO). Isis, 110 (3):567–569, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rozenblatt:2019:BRP

[2831] Daphne Rozenblatt. Book review: Petteri Pietikäinen. Madness: a History. Isis, 110(3):570–571, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Haalboom:2019:BRI

[2832] Floor Haalboom. Book review: Iwan Rhys Morus, ed. The Oxford Illustrated History of Science. Isis, 110(3):571–572, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lauzon:2019:BRJ

[2833] Matthew Lauzon. Book review: John E. Joseph. Language, Mind, and Body: a Conceptual History. Isis, 110(3):572–574, September 2019. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lockhart:2019:PPP

[2834] Ellen Lockhart. Peter pesic. Polyphonic Minds: Music of the Hemispheres.. Isis, 110(3):574–575, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Alberti:2019:BRA

[2835] Samuel J. M. M. Alberti. Book review: Arthur MacGregor, ed. Naturalists in the Field: Collecting, Recording, and Preserving the Natural World from the Fifteenth to the Twenty-First Century. Isis, 110(3): 576–577, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Por:2019:BRM

[2836] Anne Por. Book review: Michael Friedman. A History of Folding in Mathematics: Mathematizing the Margins. Isis, 110(3):577–578, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeBont:2019:BRB

[2837] Raf De Bont. Book review: Ben A. Minteer; Jane Maienschein; James P. Collins, eds. The Ark and Beyond: The Evolution of Zoo and Aquarium Conservation. Isis, 110(3):578–579, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brunet:2019:BRM

[2838] T. D. P. Brunet. Book review: Michael Ruse. On Purpose. Isis, 110 (3):580–581, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Arnold:2019:BRR

[2839] David Arnold. Book review: Robert A. Voeks. The Ethnobotany of Eden: Rethinking the Jungle Medicine Narrative. Isis, 110(3):581–582, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wujastyk:2019:CPS

[2840] Dominik Wujastyk. C. Pierce salguero, ed. Buddhism and Medicine: an Anthology of Premodern Sources. Isis, 110(3):582–584, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

David:2019:BRJ

[2841] Rosalie David. Book review: John Z. Wee, ed. The Comparable Body: Analogy and Metaphor in Ancient Mesopotamian, Egyptian, and Greco-Roman Medicine. Isis, 110(3):584–585, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tolsa:2019:BRR

[2842] Cristian Tolsa. Book review: Richard Carrier. The Scientist in the Early Roman Empire. Isis, 110(3):585–586, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Polloni:2019:BRA

[2843] Nicola Polloni. Book review: Andreas Lammer. The Elements of Avicenna’s Physics: Greek Sources and Arabic Innovations. Isis, 110(3): 586–587, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kaplan:2019:BRR

[2844] Abram Kaplan. Book review: Richard J. Oosterhoff. Making Mathematical Culture: University and Print in the Circle of Lefèvre d’Étaples. Isis, 110(3):588–589, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vai:2019:BRG

[2845] Gian Battista Vai. Book review: Giuseppe Olmi; Fulvio Simoni, eds. Ulisse Aldrovandi: Libri e immagini di Storia naturale nella prima Età moderna. Isis, 110(3):589–590, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Habermann:2019:BRC

[2846] Katharina Habermann. Book review: C. Philipp E. Nothaft. Scandalous Error: Calendar Reform and Calendrical Astronomy in Medieval Europe. Isis, 110(3):590–592, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hendriksen:2019:BRR

[2847] Marieke M. A. Hendriksen. Book review: Raymond Clemens, ed. The Voynich Manuscript. Isis, 110(3):592–593, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bennett:2019:BRN

[2848] Jim Bennett. Book review: Nicolàs de Hilster. Navigation on Wood: Wooden Navigational Instruments, 1590-1731: an Analysis of Early Modern Western Instruments for Celestial Navigation, Their Origins, Mathematical Concepts, and Accuracies. Isis, 110(3):593–594, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wallis:2019:BRO

[2849] Patrick Wallis. Book review: Ofer Hadass. Medicine, Religion, and Magic in Early Stuart England: Richard Napier’s Medical Practice. Isis, 110 (3):595–596, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shea:2019:BRF

[2850] William R. Shea. Book review: Flavia Marcacci. Cieli in contraddizione: Giovanni Battista Riccioli e il terzo sistema del mondo. Isis, 110(3): 596–597, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tateo:2019:BRV

[2851] Luca Tateo. Book review: Vittorio Hösle. Vico’s New Science of the Intersubjective World. Isis, 110(3):597–598, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blay:2019:BRJ

[2852] Michel Blay. Book review: J. B. Shank. Before Voltaire: The French Origins of “Newtonian” Mechanics, 1680–1715. Isis, 110(3):598–599, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Henry:2019:BRN

[2853] John Henry. Book review: Niccolò Guicciardini. Isaac Newton and Natural Philosophy. Isis, 110(3):599–600, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pauw:2019:BRT

[2854] Sylvia Pauw. Book review: Thomas Sonar. The History of the Priority Dispute between Newton and Leibniz: Mathematics in History and Culture. Isis, 110(3):601–602, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hofmann:2019:BRO

[2855] Martin Hofmann. Book review: Ori Sela. China’s Philological Turn: Scholars, Textualism, and the Dao in the Eighteenth Century. Isis, 110 (3):602–603, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vartija:2019:BRA

[2856] Devin Vartija. Book review: Anne C. Vila. Suffering Scholars: Pathologies of the Intellectual in Enlightenment France. Isis, 110(3):603–604, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schilt:2019:BRL

[2857] Cornelis J. Schilt. Book review: Laura Miller. Reading Popular Newtonianism: Print, the Principia, and the Dissemination of Newtonian Science. Isis, 110(3):604–606, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bruneau:2019:BRB

[2858] Olivier Bruneau. Book review: Benjamin Wardhaugh, ed. The Correspondence of Charles Hutton (1737-1823): Mathematical Networks in Georgian Britain. Isis, 110(3):606–607, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anthony:2019:BRL

[2859] Patrick Anthony. Book review: Lissa L. Roberts; Simon Werrett, eds. Compound Histories: Materials, Governance and Production, 1760– 1840. Isis, 110(3):607–608, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Casado:2019:BRS

[2860] Santos Casado. Book review: Stefan Bargheer. Moral Entanglements: Conserving Birds in Britain and Germany. Isis, 110(3):608–610, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Monroe:2019:BRE

[2861] John Warne Monroe. Book review: Emily Ogden. Credulity: a Cultural History of US Mesmerism. Isis, 110(3):610–611, September 2019. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lorimer:2019:BRE

[2862] Douglas A. Lorimer. Book review: Efram Sera-Shriar, ed. Historicizing Humans: Deep Time, Evolution, and Race in Nineteenth-Century British Sciences. Isis, 110(3):611–613, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Howkins:2019:BRJ

[2863] Adrian Howkins. Book review: John V. H. Dippel. To the Ends of the Earth: The Truth behind the Glory of Polar Exploration. Isis, 110(3): 613–614, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kraft:2019:BRM

[2864] Tobias Kraft. Book review: Myron Echenberg. Humboldt’s Mexico: In the Footsteps of the Illustrious German Scientific Traveller. Isis, 110 (3):614–616, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schwach:2019:BRK

[2865] Vera Schwach. Book review: Kim Helsvig; Jan Eivind Myhre, eds. Making a Modern University: The University of Oslo, 1811–2018. Isis, 110 (3):616–617, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Glass:2019:BRS

[2866] Ian S. Glass. Book review: Stephen Case. Making Stars Physical: The Astronomy of Sir John Herschel. Isis, 110(3):617–618, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heaton:2019:BRM

[2867] Matthew M. Heaton. Book review: Manuella Meyer. Reasoning against Madness: Psychiatry and the State in Rio de Janeiro, 1830–1944. Isis, 110(3):618–620, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Laubichler:2019:BRF

[2868] Manfred D. Laubichler. Book review: Frederick B. Churchill. August Weismann: Development, Heredity, and Evolution. Isis, 110(3):620–621, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mitchell:2019:BRL

[2869] Daniel Jon Mitchell. Book review: Lee T. Macdonald. Kew Observatory and the Evolution of Victorian Science, 1840–1910. Isis, 110(3):621–622, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barth:2019:BRA

[2870] Howard G. Barth. Book review: Alexander Pechenkin. The History of Research on Chemical Periodic Processes. Isis, 110(3):623–624, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heymann:2019:BRR

[2871] Matthias Heymann. Book review: Ralph Jewell. The Weather’s Face: Features of Science in the Story of Vilhelm Bjerknes and the Bergen School of Meteorology. Isis, 110(3):624–625, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Turda:2019:BRB

[2872] Marius Turda. Book review: Book Review: Dennis L. Durst. Eugenics and Protestant Social Reform: Hereditary Science and Religion in America, 1860-1940.; Karin Alejandra Rosemblatt. The Science and Politics of Race in Mexico and the United States, 1910-1950. Isis, 110(3):625–627, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hahn:2019:BRM

[2873] Andre Michael Hahn. Book review: Mark Blacklock. The Emergence of the Fourth Dimension: Higher Spatial Thinking in the Fin de Siècle. Isis, 110(3):628–629, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kenny:2019:BRE

[2874] Robert Kenny. Book review: Élisabeth Roudinesco. Freud in His Time and Ours. Isis, 110(3):629–630, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gouvea:2019:BRJ

[2875] Fernando Q. Gouvêa. Book review: Jean-Marc Ginoux. History of Nonlinear Oscillations Theory in France (1880-1940). Isis, 110(3):630–631, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ogilvie:2019:BRJ

[2876] Marilyn Bailey Ogilvie. Book review: Joeri Bruyninckx. Listening in the Field: Recording and the Science of Birdsong. Isis, 110(3):632–633, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mestyan:2019:BRD

[2877] Adam Mestyan. Book review: Daniel A. Stolz. The Lighthouse and the Observatory: Islam, Science, and Empire in Late Ottoman Egypt. Isis, 110(3):633–634, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Evans:2019:BRC

[2878] James Evans. Book review: Christian Bracco. Quand Albert devient Einstein. Isis, 110(3):634–636, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Abt:2019:BRK

[2879] Jeffrey Abt. Book review: Kathleen L. Sheppard, ed. My Dear Miss Ransom ...”: Letters between Caroline Ransom Williams and James Henry Breasted, 1898-1935. Isis, 110(3):636–637, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Whitworth:2019:BRM

[2880] Michael H. Whitworth. Book review: Mark S. Morrisson. Modernism, Science, and Technology. Isis, 110(3):637–638, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ariga:2019:BRY

[2881] Nobumichi Ariga. Book review: Yasu Furukawa. Chemists’ Kyoto School: Gen-itsu Kita and Japan’s Chemistry. Isis, 110(3):639–640, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Killen:2019:BRC

[2882] Andreas Killen. Book review: Christian Bonah; David Cantor; Anja Laukötter, eds. Health Education Films in the Twentieth Century. Isis, 110(3):640–641, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dunlop:2019:HPS

[2883] Katherine Dunlop. Henri Poincaré. Science and Hypothesis: The Complete Text. Isis, 110(3):641–642, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Golan:2019:BRD

[2884] Tal Golan. Book review: Diana Dolev. The Planning and Building of the Hebrew University, 1919-1948: Facing the Temple Mount. Isis, 110 (3):643–644, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Crockatt:2019:BRZ

[2885] Richard Crockatt. Book review: Ze’ev Rosenkranz, ed. The Travel Diaries of Albert Einstein: The Far East, Palestine, and Spain, 1922-1923. Isis, 110(3):644–645, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Laubichler:2019:BRM

[2886] Manfred D. Laubichler. Book review: Maurizio Meloni. Political Biology: Science and Social Values in Human Heredity from Eugenics to Epigenetics. Isis, 110(3):645–646, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hoeneveld:2019:BRD

[2887] Friso Hoeneveld. Book review: Dirk van Delft; Ton van Helvoort. Beelden zonder weerga: De elektronenmicroscoop van Ernst Ruska tot Ben Feringa. Isis, 110(3):646–648, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Walker:2019:BRB

[2888] Mark Walker. Book review: Brian E. Crim. Our Germans: Project Paperclip and the National Security State. Isis, 110(3):648–649, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Siddiqi:2019:BRJ

[2889] Asif Siddiqi. Book review: Jörg Matthias Determann. Space Science and the Arab World: Astronauts, Observatories, and Nationalism in the Middle East. Isis, 110(3):649–651, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Olszewski:2019:SMA

[2890] Todd M. Olszewski. Shelley McKellar. Artificial Hearts: The Allure and Ambivalence of a Controversial Medical Technology. Isis, 110(3):651–652, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vicedo:2019:BRL

[2891] Marga Vicedo. Book review: Lee Alan Dugatkin; Lyudmila Trut. How to Tame a Fox (and Build a Dog): Visionary Scientists and a Siberian Tale of Jump-Started Evolution. Isis, 110(3):652–654, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Healey:2019:BRR

[2892] Dan Healey. Book review: Rebecca Reich. State of Madness: Psychiatry, Literature, and Dissent after Stalin. Isis, 110(3):654–655, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lambright:2019:BRI

[2893] W. Henry Lambright. Book review: Iina Kohonen. Picturing the Cosmos: a Visual History of Early Soviet Space Endeavor. Isis, 110(3):655–656, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Baneke:2019:BRA

[2894] David Baneke. Book review: Alexander C. T. Geppert, ed. Limiting Outer Space: Astroculture after Apollo. Isis, 110(3):656–657, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blume:2019:BRB

[2895] Stuart Blume. Book review: Book Review: Paul Richards. Ebola: How A People’s Science Helped End an Epidemic.; Debora Diniz. Zika: From the Brazilian Backlands to Global Threat. Isis, 110(3):658–660, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2019:FCc

[2896] Anonymous. Front cover. Isis, 110(3):??, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2019:Mc

[2897] Anonymous. Masthead. Isis, 110(3):??, September 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Regier:2019:RCR

[2898] Jonathan Regier. Reading Cardano with the Roman inquisition: Astrology, celestial physics, and the force of heresy. Isis, 110(4):661–679, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pancaldi:2019:DTL

[2899] Giuliano Pancaldi. Darwin’s technology of life. Isis, 110(4):680–700, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kranakis:2019:TTI

[2900] Eda Kranakis. A tale of two inventions: Monsanto, biotechnology, and the geography of postmodern science. Isis, 110(4):701–725, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pelegrin:2019:ADM

[2901] Diego Pelegrin and Christián C. Carman. Accurata delineatio motuum stellae martis: How accurate is Kepler’s pretzel diagram? Isis, 110(4): 726–741, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Zeng:2019:CHY

[2902] Dian Zeng, Jian Yang, and Lewis Pyenson. Compatible humanists: Yuen ren chao meets George sarton. Isis, 110(4):742–753, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

LaFollette:2019:TCU

[2903] Marcel Chotkowski LaFollette. Track conditions: Upon revisiting How Superstition Won and Science Lost. Isis, 110(4):755–757, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weidman:2019:BPS

[2904] Nadine Weidman. Burnham, popular science, and popularization. Isis, 110(4):758–761, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schmalzer:2019:LAF

[2905] Sigrid Schmalzer. Look from afar: Fragmented authority in China and the United States. Isis, 110(4):762–765, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gordin:2019:ODP

[2906] Michael D. Gordin. The other demarcation problem. Isis, 110(4):766– 769, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hsiung:2019:WSW

[2907] Hansun Hsiung. Whose science wins or loses? (and What’s left for reason after?). Isis, 110(4):770–774, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lewenstein:2019:PVH

[2908] Bruce V. Lewenstein. Polemic versus history: Reflections on John c. Burnham’s How Superstition Won and Science Lost. Isis, 110(4):775– 778, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rasmussen:2019:MMJ

[2909] Nicolas Rasmussen and Robert N. Proctor. From maverick to mole: John c. Burnham, tobacco consultant. Isis, 110(4):779–783, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pandora:2019:PMS

[2910] Katherine Pandora. Popularizing, moralizing, and the soul of American science. Isis, 110(4):784–787, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tomes:2019:AJJ

[2911] Nancy Tomes. An American jeremiad: John c. Burnham and the history of science popularization. Isis, 110(4):788–791, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Casper:2019:HSW

[2912] Stephen T. Casper. “With hindsight, I see that I was right”: John C. Burnham’s final words, as recounted by a trickster. Isis, 110(4):792–795, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chen:2019:TC

[2913] Jessie Wei-Hsuan Chen. Time capsule. Isis, 110(4):796–797, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ragep:2019:BRS

[2914] Sally P. Ragep. Book review: Sonja Brentjes. Teaching and Learning the Sciences in Islamicate Societies (800-1700). Isis, 110(4):798–800, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Illy:2019:BRJ

[2915] József Illy. Book review: John Milsom. The Hunt for Earth Gravity: a History of Gravity Measurement from Galileo to the Twenty-First Century. Isis, 110(4):800–801, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Zieme:2019:BRC

[2916] Stefan Zieme. Book review: Christoph Lüthy; Claudia Swan; Paul Bakker; Claus Zittel, eds. Image, Imagination, and Cognition: Medieval and Early Modern Theory and Practice. Isis, 110(4):801–802, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Muka:2019:BRO

[2917] Sam Muka. Book review: Oren Harman; Michael R. Dietrich, eds. Dreamers, Visionaries, and Revolutionaries in the Life Sciences. Isis, 110 (4):802–804, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kiverstein:2019:BRJ

[2918] Julian Kiverstein. Book review: Jeff Kochan. Science as Social Existence: Heidegger and the Sociology of Scientific Knowledge. Isis, 110(4):804– 805, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nye:2019:BRA

[2919] David E. Nye. Book review: Alan G. Gross. The Scientific Sublime: Popular Science Unravels the Mysteries of the Universe. Isis, 110(4): 805–806, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bian:2019:GMS

[2920] He Bian. Georges Métailié. Science and Civilisation in China. volume 6: Biology and Biological Technology.Part 4: Traditional Botany: an Ethnobotanical Approach. Isis, 110(4):806–807, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shcheglov:2019:BRP

[2921] Dmitry A. Shcheglov. Book review: Paul T. Keyser; John Scarborough, eds. The Oxford Handbook of Science and Medicine in the Classical World. Isis, 110(4):808–809, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sidoli:2019:BRM

[2922] Nathan Sidoli. Book review: Michalis Sialaros, ed. Revolutions and Continuity in Greek Mathematics. Isis, 110(4):809–810, December 2019. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Koetsier:2019:BRA

[2923] Teun Koetsier. Book review: Adrienne Mayor. Gods and Robots: Myths, Machines, and Ancient Dreams of Technology. Isis, 110(4):810–811, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Panaino:2019:BRS

[2924] Antonio Panaino. Book review: Stefanie Rudolf. Syrische Astrologie und das Syrische Medizinbuch. Isis, 110(4):812–813, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Clement:2019:BRK

[2925] Albert Clement. Book review: Karsten Mackensen. Musik und die Ordnung der Dinge im ausgehenden Mittelalter und in der Frühen Neuzeit. Isis, 110(4):813–814, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bruzzone:2019:BRS

[2926] Raffaella Bruzzone. Book review: Sarah R. Kyle. Medicine and Humanism in Late Medieval Italy: The ”Carrara Herbal” in Padua. Isis, 110 (4):814–815, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nutton:2019:BRB

[2927] Vivian Nutton. Book review: Book Review: Erika Gielen; Michèle Goyens, eds. Towards the Authority of Vesalius: Studies on Medicine and the Human Body from Antiquity to the Renaissance and Beyond.; Rinaldo Fernando Canalis; Massimo Ciavolella, eds. Andreas Vesalius and the ”Fabrica” in the Age of Printing: Art, Anatomy, and Printing in the Italian Renaissance. Isis, 110(4):816–817, December 2019. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McClure:2019:BRP

[2928] George McClure. Book review: Pamela O. Long. Engineering the Eternal City: Infrastructure, Topography, and the Culture of Knowledge in Late Sixteenth-Century Rome. Isis, 110(4):817–819, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Klestinec:2019:BRB

[2929] Cynthia Klestinec. Book review: Barbara Tramelli. Giovanni Paolo Lomazzo’s ”Trattato dell’Arte della Pittura”: Color, Perspective, and Anatomy. Isis, 110(4):819–820, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Madruga:2019:BRH

[2930] Catarina Madruga. Book review: Hugh Cagle. Assembling the Tropics: Science and Medicine in Portugal’s Empire, 1450-1700. Isis, 110(4):820– 821, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Long:2019:BRA

[2931] Pamela O. Long. Book review: Alexander Marr; Raphaële Garrod; José Ramón Marcaida; Richard J. Oosterhoff. Logodaedalus: Word Histories of Ingenuity in Early Modern Europe. Isis, 110(4):822–823, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Keblusek:2019:BRK

[2932] Marika Keblusek. Book review: K. Hoogendoorn. Bibliography of the Exact Sciences in the Low Countries from ca. 1470 to the Golden Age (1700). Isis, 110(4):823–824, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanGelder:2019:AMC

[2933] Roelof van Gelder. Arthur MacGregor. Company Curiosities: Nature, Culture, and the East India Company, 1600-1874. Isis, 110(4):824–825, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jonker:2019:BRH

[2934] Matthijs Jonker. Book review: Henrietta McBurney; Paula Findlen; Caterina Napoleone; Ian Rolfe, eds. Birds, Other Animals and Natural Curiosities. Isis, 110(4):825–827, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Scott-Warren:2019:BRL

[2935] Jason Scott-Warren. Book review: Lianne Habinek. The Subtle Knot: Early Modern English Literature and the Birth of Neuroscience. Isis, 110 (4):827–828, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hackmann:2019:BRM

[2936] Willem D. Hackmann. Book review: Michael Hunter. The Image of Restoration Science: The Frontispiece to Thomas Sprat’s History of the Royal Society (1667). Isis, 110(4):828–829, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pantaloni:2019:BRF

[2937] Marco Pantaloni. Book review: Francesco Luzzini. Theory, Practice, and Nature In-between: Antonio Vallisneri’s ”Primi Itineris Specimen.”. Isis, 110(4):829–831, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Osborne:2019:BRA

[2938] Michael A. Osborne. Book review: Anton Serdeczny. Du Tabac pour le Mort: Une Histoire de la Réanimation. Isis, 110(4):831–832, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rioux-Beaulne:2019:BRG

[2939] Mitia Rioux-Beaulne. Book review: Gerhardt Stenger, ed. Les singularités de la nature. Isis, 110(4):832–833, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wils:2019:BRB

[2940] Kaat Wils. Book review: Book Review: Michel Bourdeau; Mary Pickering; Warren Schmaus, eds. Love, Order, and Progress: The Science, Philosophy, and Politics of Auguste Comte.; Johannes Feichtinger; Franz L. Fillafer; Jan Surman, eds. The Worlds of Positivism: a Global Intellectual History, 1770-1930. Isis, 110(4):833–835, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hopwood:2019:BRM

[2941] Nick Hopwood. Book review: Mareike Vennen. Das Aquarium: Praktiken, Techniken und Medien der Wissensproduktion (1840-1910). Isis, 110(4):837–838, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gillin:2019:BRR

[2942] Edward J. Gillin. Book review: Ruth Barton. The X Club: Power and Authority in Victorian Science. Isis, 110(4):838–839, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Crook:2019:BRM

[2943] Tom Crook. Book review: Matthew L. Newsom Kerr. Contagion, Isolation, and Biopolitics in Victorian London. Isis, 110(4):839–840, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Theunissen:2019:ESN

[2944] Bert Theunissen. Ellinor Schweighöfer. Vom Neandertal nach Afrika: Der Streit um den Ursprung der Menschheit im 19. und 20. Jahrhundert. Isis, 110(4):841–842, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Durst:2019:BRD

[2945] Dennis L. Durst. Book review: Diane B. Paul; John Stenhouse; Hamish G. Spencer, eds. Eugenics at the Edges of Empire: New Zealand, Australia, Canada, and South Africa. Isis, 110(4):842–843, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Darrigol:2019:BRB

[2946] Olivier Darrigol. Book review: Book Review: A. J. Kox, ed. The Scientific Correspondence of H. A. Lorentz.; A. J. Kox, ed. The Scientific Correspondence of H. A. Lorentz. volume 2: The Dutch Correspondents. Isis, 110(4):843–844, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Badino:2019:BRJ

[2947] Massimiliano Badino. Book review: Judith R. Goodstein. Einstein’s Italian Mathematicians: Ricci, Levi-Civita, and the Birth of General Relativity. Isis, 110(4):845–846, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lightman:2019:BRM

[2948] Bernard Lightman. Book review: Michael J. Crowe. The Gestalt Shift in Conan Doyle’s Sherlock Holmes Stories. Isis, 110(4):846–847, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mayer:2019:BRU

[2949] Andreas Mayer. Book review: Ulrike May; trans. by Daniela Haller; Bettina Mathes; Michael Molnar; Philip Slotkin; Deirdre Winter. Freud at Work: On the History of Psychoanalytic Theory and Practice, with an Analysis of Freud’s Patient Record Books. Isis, 110(4):847–848, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sysling:2019:BRH

[2950] Fenneke Sysling. Book review: Hans Pols. Nurturing Indonesia: Medicine and Decolonisation in the Dutch East Indies. Isis, 110(4): 849–850, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lannoo:2019:BRM

[2951] Michael J. Lannoo. Book review: Megan Raby. American Tropics: The Caribbean Roots of Biodiversity Science. Isis, 110(4):850–851, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ginoux:2019:BRJ

[2952] Jean-Marc Ginoux. Book review: Juliet Floyd; Alisa Bokulich, eds. Philosophical Explorations of the Legacy of Alan Turing: Turing 100. Isis, 110(4):851–852, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gooday:2019:BRD

[2953] Graeme Gooday. Book review: David Kaldewey; Désirée Schauz, eds. Basic and Applied Research: The Language of Science Policy in the Twentieth Century. Isis, 110(4):853–854, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nye:2019:BRB

[2954] Mary Jo Nye. Book review: Book Review: Roger H. Stuewer. The Age of Innocence: Nuclear Physics between the First and Second World Wars.; Joseph D. Martin. Solid State Insurrection: How the Science of Substance Made American Physics Matter. Isis, 110(4):854–857, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hamilton:2019:BRK

[2955] Emily Hamilton. Book review: Keith M. Parsons; Robert A. Zaballa. Bombing the Marshall Islands: a Cold War Tragedy. Isis, 110(4):857– 858, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Adler:2019:BRS

[2956] Antony Adler. Book review: Samuel A. Robinson. Ocean Science and the British Cold War State. Isis, 110(4):858–859, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Piper:2019:BRD

[2957] Liza Piper. Book review: Dóra Vargha. Polio across the Iron Curtain: Hungary’s Cold War with an Epidemic. Isis, 110(4):860–861, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Launius:2019:BRM

[2958] Roger D. Launius. Book review: Michael J. Neufeld. Spaceflight: a Concise History. Isis, 110(4):861–862, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Spiller:2019:DHD

[2959] James Spiller. David h. DeVorkin. Fred Whipple’s Empire: The Smithsonian Astrophysical Observatory, 1955-1973. Isis, 110(4):862–863, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barth:2019:BRD

[2960] Howard G. Barth. Book review: David Quammen. The Tangled Tree: a Radical New History of Life. Isis, 110(4):864–865, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Slagstad:2019:BRL

[2961] Ketil Slagstad. Book review: Lukas Engelmann. Mapping AIDS: Visual Histories of an Enduring Epidemic. Isis, 110(4):865–866, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2019:IIV

[2962] Anonymous. Index to Isis, volume 110, 2019. Isis, 110(4):867–883, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2019:CIV

[2963] Anonymous. Contents of Isis, volume 110. Isis, 110(4):884–886, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2019:FCd

[2964] Anonymous. Front cover. Isis, 110(4):??, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2019:Md

[2965] Anonymous. Masthead. Isis, 110(4):??, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fayter:2019:BRI

[2966] Paul Fayter. Book review: Iwan Rhys Morus. William Robert Grove: Victorian Gentleman of Science. Isis, 110(4):??, December 2019. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lavine:2019:BRB

[2967] Matthew Lavine and Alexandra Hui. Book review: Book Review: John Burnham’s How Superstition Won and Science Lost: Editors’ introduction. Isis, 110(4):??, December 2019. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kline:1970:BLI

[2968] Marcia B. Kline. Beyond the land itself: views of nature in Canada and the United States. Essays in history and literature. Harvard University Press, Cambridge, MA, USA, 1970. ISBN 0-674-06915-3. 75 pp. LCCN PR143 .K5.

Launius:1992:THS

[2969] Roger D. Launius and Aaron K. Gillette. Toward a history of the Space Shuttle: an annotated bibliography, volume 1 (part 1) of Monographs in aerospace history. National Aeronautics and Space Administration, Office of Communications, Public Outreach Division, History Program Office, Washington, DC, USA, November 1992. ISBN ???? vi + 97 pp. LCCN ????

Koterski:2002:NLH

[2970] Joseph W. Koterski. Natural law and human nature. The great courses. Teaching Co., Chantilly, VA, 2002. LCCN RZA 7379, etc. 12 sound cassettes.

Lichtenberg:2005:VNG

[2971] Georg Christoph Lichtenberg. Vorlesungen zur Naturlehre. (German) [Lectures on Natural Philosophy]. Gesammelte Schriften / Georg Christoph Lichtenberg. Wallstein, Göttingen, Germany, 2005–2013. ISBN 3-89244-891-4 (vol. 1), 3-8353-0147-0 (vol. 2). ???? pp. LCCN Q157 .L48 2008.

Falchetta:2006:FMW

[2972] Piero Falchetta. Fra Mauro’s world map: with a commentary and translations of the inscriptions, volume 5 of Terrarvm orbis: histoire des representations de l’espace: textes, images; history of the representation of space in text and image; 5. Brepols, Turnhout, Belgium, 2006. ISBN 2-503-51726-9 (hardcover). 829 pp. LCCN GA221 .F35 2006.

Stack:2008:QVS

[2973] David Stack. Queen Victoria’s skull: [George Combe and the mid-Victorian mind]. Hambledon Continuum, London, UK, 2008. ISBN 1-84725-233-8 (hardcover). xvii + 350 pp. LCCN BF869.C6 S73 2008.

DeRosa:2010:DPS

[2974] Raffaella De Rosa. Descartes and the puzzle of sensory representation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-957037-X (hardcover). xiii + 190 pp. LCCN B1878.R44 D47 2010.

Beatty:2011:NDL

[2975] Heather Beatty. Nervous disease in late eighteenth-century Britain: the reality of a fashionable disorder, volume 6 of Studies for the Society for the Social History of Medicine. Pickering and Chatto, London, UK, 2011. ISBN 1-84893-309-6 (Pickering and Chatto), 1-283-29204-1 (MyiLibrary), 1-84893-308-8. vii + 241 pp. LCCN RC506 .B43 2011eb.

Epstein:2011:IMI

[2976] Katherine Cranston Epstein. Inventing the military-industrial complex: torpedo development, property rights, and naval warfare in the United States and Great Britain before World War I. Ph.D. thesis, Ohio State University, Columbus, Ohio, 2011. 452 (est) pp. URL http://search.

proquest.com/pqdtglobal/docview/906278137.

Feist:2011:SMV

[2977] Ulrike Feist. Sonne, Mond und Venus Visualisierungen astronomischen Wissens im frühneuzeitlichen Rom. (German) [Sun, Moon and Venus visualizations of astronomical Knowledge in Early Modern Rome], volume 10 of Actus et Imago. Akademie-Verlag, Berlin, Germany, 2011. ISBN 3-05-009541-5, 3-05-006365-3. viii + 259 pp. LCCN QB33.I8 F457 2013.

Larink:2011:BGB

[2978] Wibke Larink. Bilder vom Gehirn: bildwissenschaftliche Zugänge zum Gehirn als Seelenorgan. (German) [Images of the Brain: Image science approaches to the brain as a soul organ]. Akademie-Verlag, Berlin, Germany, 2011. ISBN 3-05-005118-3, 3-05-006242-8 (e-book), 1-306-48821-4 (e-book). xx + 521 pp. LCCN QP385 .L257 2011.

Long:2011:APR

[2979] Pamela O. Long. Artisan/practitioners and the rise of the new sciences, 1400–1600. The OSU Press Horning visiting scholars publication series. Oregon State University Press, Corvallis, OR, USA, 2011. ISBN 0-87071-609-3 (paperback), 0-87071-647-6 (e-book). xii + 196 pp. LCCN Q124.97 .L66 2011.

Martinez:2011:SST

[2980] Alberto Antonio Martínez. Science secrets: the truth about Darwin’s finches, Einstein’s wife, and other myths. University of Pittsburgh Press, Pittsburgh, PA, USA, 2011. ISBN 0-8229-4407-3 (hardcover). xviii + 324 pp. LCCN Q173 .M316 2011.

Nothaft:2011:DPL

[2981] C. Philipp E. Nothaft. Dating the Passion: the life of Jesus and the emergence of scientific chronology (200–1600), volume 1 of Time, astronomy, and calendars. Brill, Leiden, The Netherlands, 2011. ISBN 90-04-21219-1 (hardcover). ISSN 2211-632X. vi + 319 pp. LCCN CE6 .N68 2011.

Oswald:2011:JRC

[2982] Philip H. Oswald and Christopher David Preston, editors. John Ray’s Cambridge catalogue (1660). Ray Society, London, UK, 2011. ISBN 0-903874-43-1. ix + 612 pp. LCCN QK306 .R28 2011.

Prince:2011:ERE

[2983] Sue Ann Prince, editor. Of elephants and roses: encounters with French natural history, 1790–1830: an exhibition guide. American Philosophical Society Museum, Philadelphia, PA, USA, 2011. ISBN 0-87169-267-8 (paperback). xxvi + 268 pp. LCCN QH147 .O33 2011.

Ainley:2012:CCL

[2984] Marianne Gosztonyi Ainley, Marelene Rayner-Canham, and Marelene F. Rayner-Canham. Creating complicated lives: women and science at English–Canadian universities, 1880–1980. McGill-Queen’s University Press, Montréal, Québec, Canada, 2012. ISBN 0-7735-4066-0 (hardcover), 0-7735-4067-9 (paperback). xi + 200 pp. LCCN HQ1397 .A45 2012.

Alesse:2012:ESS

[2985] Francesca Alesse, Franco Ferrari, and Maria Cristina Dalfino. Epinomide: studi sull’opera e la sua ricezione. (Italian) [Epinomis: Studies on the work and its reception], volume 60.1 of Elenchos. Bibliopolis, Naples, Italy, 2012. ISBN 88-7088-619-0. 539 pp. LCCN B391.Z5 E625 2012.

Banerjee:2012:WMP

[2986] Anindita Banerjee. We modern people: science fiction and the making of Russian modernity. The Wesleyan early classics of science fiction series. Wesleyan University Press, Middletown, CT, USA, 2012. ISBN 0-8195-7333-7 (hardcover), 0-8195-7334-5 (paperback), 0-8195-7335-3 (e-book). viii + 206 pp. LCCN PG3098.S5 B36 2012eb.

Black:2012:AOV

[2987] Winston E. Black, editor. Anglicanus ortus: a verse herbal of the twelfth century, volume 180; 3 of Studies and texts; British writers of the Middle Ages and the early modern period. Pontifical Institute of Mediaeval Studies, Toronto, ON, Canada, 2012. ISBN 0-88844-180-0 (PIMS), 1-85124-284-8 (Bodleian Library). ISSN 0082-5328. xiv + 561 pp. LCCN RM666.H33 H456 2012.

Black:2012:CJB

[2988] Joseph Black, R. G. W. (Robert Geoffrey William) Anderson, and Jean Jones, editors. The correspondence of Joseph Black. Ashgate, Farnham, Surrey, England, 2012. ISBN 0-7546-0131-5 (hardcover). xiv + 1564 (two volumes) pp. LCCN QD22.B549 A4 2012.

Blum:2012:GBI

[2989] Paul Richard Blum. Giordano Bruno: An Introduction. Value Inquiry Book S. Editions Rodopi B.V., Amsterdam, The Netherlands, 2012. ISBN 90-420-3555-2 (paperback), 94-012-0829-8 (e-book). viii + 128 pp. LCCN B783.Z7.

Blume:2012:SMG

[2990] Dieter Blume, Mechthild Haffner, and Wolfgang Metzger. Sternbilder des Mittelalters: der gemalte Himmel zwischen Wissenschaft und Phantasie. (German) [Constellations of the Middle Ages: the painted sky between science and fantasy]. Akademie-Verlag, Berlin, Germany, 2012. ISBN 3-05-005664-9 (vol. 1). 651 + 403 pp. LCCN QB23 .B65 2012.

Bret:2012:OLC

[2991] Patrice Bret and Henri B. Kagan. Oeuvres de Lavoisier, correspondance. Volume VII, 1792–1794. Éditions Hermann, Paris, France, 2012. ISBN 2-7056-8288-0. xv + 487 pp. LCCN QD3 L414 1965 7.

Burnham:2012:AFL

[2992] John C. (John Chynoweth) Burnham, editor. After Freud left: a century of psychoanalysis in America. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-08137-0 (hardcover). 274 pp. LCCN RC503 .A38 2012.

Campe:2012:GPL

[2993] Rüdiger Campe. The game of probability: literature and calculation from Pascal to Kleist. Cultural memory in the present. Stanford University Press, Stanford, CA, USA, 2012. ISBN 0-8047-6864-1 (hardcover), 0-8047-6865-X (paperback). viii + 486 pp. LCCN PN744 .C3613 2012. Translated by Ellwood H. Wiggins, Jr.

Chakrabarti:2012:BBI

[2994] Pratik Chakrabarti. Bacteriology in British India: laboratory medicine and the tropics, volume 22 of Rochester studies in medical history. University of Rochester Press, Rochester, NY, 2012. ISBN 1-58046-408-4 (hardcover). ISSN 1526-2715. x + 304 pp. LCCN QR46 .C43 2012.

Chemla:2012:HMP

[2995] Karine Chemla, editor. The history of mathematical proof in ancient traditions. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-01221-X (hardcover). xv + 596 pp. LCCN QA22 .H588 2012. URL http:

//www.loc.gov/catdir/enhancements/fy1305/2012472526-b.html;
 http://www.loc.gov/catdir/enhancements/fy1305/2012472526-d.
html;   http://www.loc.gov/catdir/enhancements/fy1305/2012472526-
t.html.

Cook:2012:TKE

[2996] Harold John Cook and Sven Dupré, editors. Translating knowledge in the Early Modern Low Countries, volume 3 of Low Countries studies on the circulation of natural knowledge. LIT Verlag, Zürich, Switzerland, 2012. ISBN 3-643-90246-8 (paperback). ii + 466 pp. LCCN Q124 .T74 2012.

Darwin:2012:JBD

[2997] Charles Darwin, Patrick Tort, Michel Prum, Christiane Bernard, Marie-Thérèse Blanchon, and Claude Rouquette. Journal de bord du voyage du Beagle [1831–1836]. (French) [Diary of the voyage of the Beagle [1831– 1836]], volume 15 of Champion classiques. Essais. H. Champion, Paris, France, 2012. ISBN 2-7453-2449-7 (paperback). 822 pp. LCCN ????

Ellenbogen:2012:RUI

[2998] Josh Ellenbogen. Reasoned and unreasoned images: the photography of Bertillon, Galton, and Marey. Pennsylvania State University Press, University Park, PA, USA, 2012. ISBN 0-271-05259-7 (hardcover). xi + 265 pp. LCCN TR139 .E44 2012.

Gold:2012:TBW

[2999] Thomas Gold and Simon Mitton, editors. Taking the back off the watch: a personal memoir, volume 381 of Astrophysics and space science library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-27587-7, 3-642-27588-5 (e-book). ISSN 0067-0057. xvii + 232 pp. LCCN QB36.G653 G65 2012.

Goodrich:2012:THS

[3000] Malinda K. Goodrich, Alice R. Buchalter, and Patrick M. Miller. Toward a history of the Space Shuttle: an annotated bibliography, volume 49 (part

2) of Monographs in aerospace history. National Aeronautics and Space Administration, Office of Communications, Public Outreach Division, History Program Office, Washington, DC, USA, 2012. ISBN ???? v + 141 pp. LCCN ???? NASA SP-2012-4549.

Grootenboer:2012:TGI

[3001] Hanneke Grootenboer. Treasuring the gaze: intimate vision in late eighteenth-century eye miniatures. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-30966-5 (hardcover), 0-226-30971-1 (e-book). xvii + 223 pp. LCCN N8217.E9 G76 2012.

Harden:2012:AH

[3002] Victoria Angela Harden. AIDS at 30: a history. Potomac Books, Washington, DC, USA, 2012. ISBN 1-59797-294-0. xvi + 324 pp. LCCN RA643.8 .H37 2012.

Herrmann:2012:GCM

[3003] Sabine Herrmann. Giacomo Casanova und die Medizin des 18. Jahrhunderts. (German) [Giacomo Casanova and medicine in the 18th Century], volume 44 of Medizin, Gesellschaft und Geschichte. Beiheft. Franz Steiner Verlag, Stuttgart, Germany, 2012. ISBN 3-515-10175-6. 214 pp. LCCN D285.8.C4 H477 2012.

Isaac:2012:UEA

[3004] Joel Isaac and Duncan Bell, editors. Uncertain empire: American history and the idea of the Cold War. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-982612-9 (hardcover), 0-19-982614-5 (paperback). x + 302 pp. LCCN E169.12 .U478 2012.

Jenyns:2012:FCV

[3005] Leonard Jenyns, R. C. Preece, and Tim Sparks, editors. Fauna Cantabrigiensis: the vertebrate and molluscan fauna of Cambridgeshire by the Rev. Leonard Jenyns (1800–1893): transcript and commentaries, volume 174 of Ray Society. Ray Society, London, UK, 2012. ISBN 0-903874-44-X. vii + 226 pp. LCCN QL256 .J46 2012.

Jonsson:2012:EFS

[3006] Fredrik Albritton Jonsson. Enlightenment’s frontier the Scottish Highlands and the origins of environmentalism. Lewis Walpole series in eighteenth-century culture and history. Yale University Press, New Haven, CT, USA, 2012. ISBN 0-300-16254-5 (hardcover). LCCN GE199.G7 J66 2013.

Kennaway:2012:BVH

[3007] James Gordon Kennaway. Bad vibrations: the history of the idea of music as cause of disease. History of medicine in context. Ashgate, Farnham, Surrey, 2012. ISBN 1-4094-2642-4 (hardcover), 1-4094-2643-2 (e-book), 1-280-68951-X (MyiLibrary). xii + 213 pp. LCCN ML3830 .K43 2012.

Kleeberg:2012:SAA

[3008] Bernhard Kleeberg. Schlechte Angewohnheiten: eine Anthologie 1750– 1900. (German) [Bad habits: An Anthology 1750–1900], volume 2002 of Suhrkamp Taschenbuch Wissenschaft. Suhrkamp, Berlin, Germany, 2012. ISBN 3-518-29602-7 (paperback). 446 pp. LCCN BJ1534 .S35 2012.

Koch:2012:FMT

[3009] Erin Koch. Free market tuberculosis: managing epidemics in post-Soviet Georgia. Vanderbilt University Press, Nashville, TN, USA, 2012. ISBN 0-8265-1892-3, 0-8265-1893-1, 0-8265-1894-X. xiv + 231 pp. LCCN RA644.T7 K63 2013.

Kopperman:2012:RPJ

[3010] Paul Kopperman. “Regimental Practice’ by John Buchanan, M.D.: an Eighteenth-Century medical diary and manual. The history of medicine in context. Ashgate, Farnham, UK, 2012. ISBN 0-7546-6877-0 (hardcover), 1-4094-4724-3 (e-book). xiii + 231 pp. LCCN UH258.3 .B83 2012.

Kragh:2012:NBQ

[3011] Helge Kragh. Niels Bohr and the Quantum Atom: The Bohr Model of Atomic Structure 1913–1925. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-965498-0 (hardcover), 0-19-163045-4 (e-book), 0-19-174169-8 (e-book). vi + 410 pp. LCCN QC774.A2.

Leonardi:2012:CCC

[3012] Paul M. Leonardi. Car crashes without cars: lessons about simulation technology and organizational change from automotive design. Acting with technology. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01784-9 (hardcover). x + 334 pp. LCCN TL240 .L426 2012.

Lifschitz:2012:LEB

[3013] Avi Lifschitz. Language and Enlightenment: the Berlin debates of the eighteenth century. Oxford historical monographs. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-966166-9. x + 231 pp. LCCN PF3083 .L54 2012. URL http://www.loc.

gov/catdir/enhancements/fy1403/2012464899-b.html;   http://
www.loc.gov/catdir/enhancements/fy1403/2012464899-d.html;
http://www.loc.gov/catdir/enhancements/fy1403/2012464899-t.
html.

Price:2012:LFT

[3014] Katy Price. Loving faster than light: romance and readers in Einstein’s universe. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-68073-8 (hardcover), 0-226-68075-4 (e-book). xi + 261 pp. LCCN PR478.S26 P75 2012.

Priestley:2012:MRN

[3015] Rebecca Priestley. Mad on radium: New Zealand in the Atomic Age. Auckland University Press, Auckland, NZ, 2012. ISBN 1-86940-727-X (paperback), 1-77558-586-7 (e-book). xii + 284 pp. LCCN TK9122.5 .P75 2012.

Principe:2012:SA

[3016] Lawrence M. Principe. The Secrets of Alchemy. Synthesis. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-68295-1 (hardcover), 0-226-10379-X (paperback), 0-226-92378-9 (e-book). LCCN QD13 .P75 2013.

Roberts:2012:PMW

[3017] Sean Roberts. Printing a Mediterranean world: Florence, Constantinople, and the renaissance of geography. I Tatti Studies in Italian Renaissance History. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-06648-0. xiii + 293 + 25 pp. LCCN G1005 1482 .B5 2012.

URL http://www.jstor.org/stable/10.2307/j.ctt2jbx98.

Rossiter:2012:WSA

[3018] Margaret W. Rossiter. Women scientists in America: forging a new world since 1972. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0233-5 (hardcover), 1-4214-0363-3 (paperback). xx + 426 pp. LCCN Q130 .R6835 2012.

Rowlinson:2012:SJD

[3019] J. S. (John Shipley) Rowlinson. Sir James Dewar, 1842–1923: a ruthless chemist. Science, technology and culture. Ashgate, Farnham, Surrey, England, 2012. ISBN 1-4094-0613-X (hardcover), 1-4094-0614-8 (ebook). ISSN 1700-1945. xviii + 236 pp. LCCN QD22.D345 R69 2012.

Simili:2012:SPP

[3020] Raffaella Simili. Scienziati, patrioti, presidenti: l’Accademia nazionale dei Lincei (1874–1926), volume 150 of Percorsi; Storia delle istituzioni scientifiche in Italia e in Europa. GLF editori Laterza, Roma, Italy, 2012. ISBN 88-420-9966-X (paperback). x + 197 pp. LCCN AS222.R69 S35 2012. 20.00 EUR. URL http://www.loc.gov/catdir/

toc/casalini10/12380776.pdf.

Spary:2012:EEF

[3021] E. C. (Emma C.) Spary. Eating the Enlightenment: food and the sciences in Paris. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-76886-4 (hardcover), 0-226-76888-0 (e-book). xi + 366 pp. LCCN GT2853.F8 S737 2012.

Tresch:2012:RMU

[3022] John Tresch. The Romantic Machine: Utopian Science and Technology after Napoleon. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-81220-0 (hardcover). xvii + 449 pp. LCCN T26.F8 T74 2012.

Achinstein:2013:EMS

[3023] Peter Achinstein. Evidence and method: scientific strategies of Isaac Newton and James Clerk Maxwell. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-992185-7 (hardcover), 0-19-933289-4 (e-book). xv + 177 pp. LCCN Q174.8 .A238 2013.

Amster:2013:MSS

[3024] Ellen Amster. Medicine and the saints: science, Islam, and the colonial encounter in Morocco, 1877–1956. University of Texas Press, Austin, TX, USA, 2013. ISBN 0-292-74544-3 (hardcover). xiv + 334 pp. LCCN R653.M8 A47 2013. Foreword by Rajae El Aoued.

Arapostathis:2013:PCE

[3025] Stathis Arapostathis and Graeme Gooday. Patently contestable: Electrical technologies and inventor identities on trial in Britain. Inside technology. MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-01903-5. xv + 294 pp. LCCN KD1384 .A973 2013.

Bailey:2013:FSR

[3026] Michael David Bailey. Fearful spirits, reasoned follies: the boundaries of superstition in late medieval Europe. Cornell University Press, Ithaca, NY, USA, 2013. ISBN 0-8014-5144-2 (hardcover). xv + 295 pp. LCCN GR135 .B35 2013.

Balsoy:2013:PRO

[3027] Gülhan Balsoy. The politics of reproduction in Ottoman society, 1838– 1900, volume Number 12 of The Body, Gender and Culture. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-325-8, 1-84893-326-6 (ebook). x + 180 pp. LCCN HB901 .B35 2013.

Bell:2013:RFC

[3028] Joshua A. (Joshua Alexander) Bell, Alison K. (Alison Kay) Brown, and Robert J. Gordon, editors. Recreating first contact: expeditions, anthropology, and popular culture. Smithsonian Institution Scholarly Press, Washington, DC, USA, 2013. ISBN 1-935623-14-1 (hardcover), 1-935623-24-9 (e-book). xii + 261 pp. LCCN GN345 .R43 2013.

Bennett:2013:CEP

[3029] J. A. (James A.) Bennett and Sofia Talas, editors. Cabinets of experimental philosophy in eighteenth-century Europe, volume 40 of History of science and medicine library, ; Scientific instruments and collections. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-25296-7 (hardcover), 90-04-25297-5 (e-book). ISSN 1872-0684. xxiii + 253 pp. LCCN Q185 .C27 2013.

Blay:2013:DNH

[3030] Michel Blay. Dieu, la nature et l’homme: l’originalité de l’Occident. (French) [God, nature and man: the originality of the West]. Le Temps des idées. A. Colin, Paris, France, 2013. ISBN 2-200-27772-5 (paperback). ISSN 1958-9980. 358 pp. LCCN BD581 .B5436 2013.

Boantza:2013:MML

[3031] Victor D. Boantza. Matter and method in the long chemical revolution: laws of another order. Science, technology and culture. Ashgate, Farnham, Surrey, UK, 2013. ISBN 1-4094-1867-7 (hardcover), 1-4094-1868-5 (e-book), 1-4724-0398-3 (e-pub). xiv + 266 pp. LCCN QD14 .B63 2013.

Boner:2013:KCS

[3032] Patrick J. Boner. Kepler’s cosmological synthesis: astrology, mechanism and the soul, volume 39/20 of History of science and medicine library. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-24608-8 (hardcover), 90-04-24609-6 (e-book), 1-299-69098-X (e-book). ISSN 1872-0684. x + 187 pp. LCCN QB36.K4 .B638 2013. URL http://booksandjournals.

brillonline.com/content/9789004246096.

Bowers:2013:PPH

[3033] Kristy Wilson Bowers. Plague and public health in early modern Seville. Rochester studies in medical history volume 26. University of Rochester Press, Rochester, NY, 2013. ISBN 1-58046-451-3 (hardcover). ISSN 1526-2715. x + 139 pp. LCCN RC178.S72 B69 2013.

Breidbach:2013:COD

[3034] Olaf Breidbach, Kerrin Klinger, and Matthias Müller. Camera obscura: die Dunkelkammer in ihrer historischen Entwicklung. (German) [Camera obscura: the darkroom and its historical development]. Wissenschaftsgeschichte. Franz Steiner Verlag, Stuttgart, Germany, 2013. ISBN 3-515-10005-9 (paperback). 227 pp. LCCN TR287.

Browman:2013:AHB

[3035] David L. Browman and Stephen Williams. Anthropology at Harvard: a biographical history, 1790–1940, volume 11 of Peabody museum monographs. Peabody Museum Press, Cambridge, MA, USA, 2013. ISBN 0-87365-913-9 (hardcover). xi + 589 + 64 pp. LCCN GN43.2.M4 B76 2013.

Buhler:2013:MSN

[3036] Walter Bühler. Musikalische Skalen bei Naturwissenschaftlern der frühen Neuzeit: eine elementarmathematische Analyse. (German) [Musical scales of scientists of the Early Modern Period: an elementary mathematical analysis]. PL Academic Research, Imprint der Peter Lang GmbH, Frankfurt am Main, 2013. ISBN 3-631-64430-2 (hardcover), 3-653-03056-0 (e-book), 1-299-87848-2 (e-book). 281 pp. LCCN ML3809 .B84 2013.

Carlson:2013:TIE

[3037] W. Bernard Carlson. Tesla: Inventor of the Electrical Age. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-05776-1 (hardcover), 1-4008-9912-5. xiii + 500 pp. LCCN Z8868.4 .C384 2012. URL

http://press.princeton.edu/chapters/i9941.pdf.

Caroti:2013:NOQ

[3038] Stefano Caroti, J. (Jean) Celeyrette, Stefan Kirschner, and E. Mazet, editors. Nicole Oresme, questiones super physicam (books I–VII), volume 112 of Studien und Texte zur Geistesgeschichte des Mittelalters. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-23637-6 (hardcover), 90-04-25025-5 (e-book). ISSN 0169-8028. xliv + 853 pp. LCCN Q151 .O75 2013.

Chinnici:2013:AOD

[3039] Ileana Chinnici and Antonella Gasperini. Alle origini dell’astrofisica italiana: il carteggio Secchi–Tacchini, 1861–1877. (Italian) [The Italian origins of astrophysics: the Secchi–Tacchini letters, 1861–1877], volume 116 of Fondazione Giorgio Ronchi. Fondazione Giorgio Ronchi, Firenze, Italia, 2013. ISBN 88-88649-38-7 (paperback). 496 pp. LCCN ????

Conklin:2013:MMR

[3040] Alice L. Conklin. In the museum of man: race, anthropology, and empire in France, 1850–1950. Cornell University Press, Ithaca, NY, USA, 2013. ISBN 0-8014-3755-5 (hardcover), 0-8014-7878-2 (paperback). xii + 374 pp. LCCN GN17.3.F8 C66 2013.

Conti:2013:NSN

[3041] Lino Conti, editor. Natura e scienza nella rivoluzione francescana. (Italian) [Nature and science in the Franciscan revolution]. Edizioni Centro Stampa, Perugia, Italia, 2013. ISBN ???? 316 (est.) pp. LCCN ????

Craver:2013:SMD

[3042] Carl F. Craver and Lindley Darden. In search of mechanisms: discoveries across the life sciences. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-03965-X (hardcover), 0-226-03979-X (paperback), 0-226-03979-X (e-book). xxii + 228 pp. LCCN QH331 .C898 2013.

Davids:2013:RTG

[3043] C. A. Davids. Religion, technology, and the great and little divergences: China and Europe compared, c. 700–1800, volume 32 of History of science and medicine library, 1872-0684 Knowledge infrastructure and knowledge economy; volume 2. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-23388-1 (hardcover), 90-04-23695-3 (e-book). xii + 278 pp. LCCN T26.A1 D38 2013.

DeNardis:2013:GWI

[3044] Laura DeNardis. The global war for Internet governance. Yale University Press, New Haven, CT, USA, 2013. ISBN 0-300-18135-3 (hardcover), 0-300-18211-2 (e-book). viii + 288 pp. LCCN K4345 .D46 2014.

DeWitt:2013:MAM

[3045] Anne DeWitt. Moral authority, men of science, and the Victorian novel, volume 84 of Cambridge studies in Nineteenth-century literature and culture. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-03617-8 (hardcover). ix + 273 pp. LCCN PR868.S34 D49 2013.

Dick:2013:DCA

[3046] Steven J. Dick. Discovery and classification in astronomy: controversy and consensus. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-03361-6 (hardcover). xvi + 458 pp. LCCN QB61 .D53 2013.

Duffin:2013:MSC

[3047] Jacalyn Duffin. Medical Saints: Cosmas and Damian in a Postmodern World. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-974317-7, 0-19-987673-8 (e-book). LCCN BT732.5 .D835 2013.

Dyck:2013:FER

[3048] Erika Dyck. Facing eugenics: reproduction, sterilization, and the politics of choice. University of Toronto Press, Toronto, ON, Canada, 2013. ISBN 1-4426-4416-8 (hardcover), 1-4426-1255-X (paper). 304 + 12 pp. LCCN HQ767.7.

Eckert:2013:ASA

[3049] Michael Eckert. Arnold Sommerfeld: Atomphysiker und Kulturbote, 1868–1951: eine Biografie. (German) [Arnold Sommerfeld: atomic physicist and cultural messenger 1868–1951: a biography], volume 29 of Abhandlungen und Berichte / Deutsches Museum (Germany). Wallstein, Göttingen, Germany, 2013. ISBN 3-8353-1206-5 (hardcover). 604 pp. LCCN QC16.S76 E25 2013.

Enros:2013:ESH

[3050] Philip Charles Enros. Environment for science: a history of policy for science in Environment Canada. [Published by the Author], Toronto, ON, Canada, 2013. ISBN 0-9920704-0-6 (paperback). x + 255 pp. LCCN ????

Erickson:2013:HRA

[3051] Paul Erickson, Judy L. Klein, Lorraine Daston, Rebecca M. (Rebecca Maura) Lemov, Thomas Sturm, and Michael D. Gordin. How reason almost lost its mind: the strange career of Cold War rationality. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-04663-X (hardcover), 0-226-04677-X (e-book). viii + 259 pp. LCCN D843 .E69 2013.

Fallwell:2013:MGM

[3052] Lynne Anne Fallwell. Modern German midwifery, 1885–1960, volume 13 of Studies for the Society for the Social History of Medicine. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-428-9 (hardcover), 1-78144-053-0 (e-book). xiii + 263 pp. LCCN RG950 .F35 2013.

Fancy:2013:SRM

[3053] Nahyan A. G. Fancy. Science and religion in Mamluk Egypt: Ibn al-Nafis, pulmonary transit and bodily resurrection, volume 37 of Culture and civilization in the Middle East. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2013. ISBN 0-415-62200-X (hardcover), 0-203-63161-7 (e-book), 1-299-64122-9 (MyiLibrary). xii + 186 pp. LCCN R143 .F36 2013.

Ferguson:2013:NFA

[3054] Robert G. Ferguson. NASA’s first A: aeronautics from 1958 to 2008, volume 4412 of NASA history series. National Aeronautics and Space Administration, Office of Communications, Public Outreach Division, History Program Office, Washington, DC, USA, 2013. ISBN ???? viii + 293 pp. LCCN ????

Finkelstein:2013:EDB

[3055] Gabriel Ward Finkelstein. Emil du Bois-Reymond: neuroscience, self, and society in nineteenth-century Germany. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-01950-7 (hardcover), 0-262-31484-3 (e-book), 0-262-31485-1 (e-book), 1-306-09112-8 (e-book), 1-4619-5032-5 (e-book). xviii + 362 pp. LCCN QP26.D8 F56 2013.

Floyd-Wilson:2013:OKS

[3056] Mary Floyd-Wilson. Occult knowledge, science, and gender on the Shakespearean stage. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-03632-1. xi + 236 pp. LCCN PR2965 .F56 2013. URL http://

www.loc.gov/catdir/enhancements/fy1307/2012050275-b.html;
http://www.loc.gov/catdir/enhancements/fy1307/2012050275-d.
html;   http://www.loc.gov/catdir/enhancements/fy1307/2012050275-
t.html.

Franklin:2013:SSE

[3057] Allan Franklin. Shifting standards: experiments in particle physics in the twentieth century. University of Pittsburgh Press, Pittsburgh, PA, USA, 2013. ISBN 0-8229-4430-8 (hardcover). lv + 302 pp. LCCN QC793.412 .F73 2013.

Frietsch:2013:HWG

[3058] Ute Frietsch. Häresie und Wissenschaft: eine Genealogie der paracelsischen Alchemie. (German) [Heresy and science: a genealogy of Paracelsus alchemy]. Wilhelm Fink, München, Germany, 2013. ISBN 3-7705-5433-7 (hardcover). 474 pp. LCCN QD13 .F75 2013.

Graham:2013:LIC

[3059] Loren R. Graham. Lonely ideas: can Russia compete? MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-01979-5 (hardcover), 1-4619-4297-7. xi + 204 + 12 pp. LCCN T173.5.R9 G73 2013.

Griffiths:2013:GPI

[3060] Paul Griffiths and Karola Stotz. Genetics and philosophy: an introduction. Cambridge introductions to philosophy and biology. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-00212-5 (hardcover), 0-521-17390-6 (paperback). viii + 270 pp. LCCN QH447 .G75 2013.

Gundlach:2013:PPE

[3061] Bradley J. Gundlach. Process and providence: the evolution question at Princeton, 1845–1929. William B. Eerdmans Publishing Company, Grand Rapids, MI, USA, 2013. ISBN 0-8028-6898-3 (paperback), 1-4674-3896-0 (e-book). xiv + 374 pp. LCCN BT712 .G85 2013.

Hamblin:2013:AMN

[3062] Jacob Darwin Hamblin. Arming Mother Nature: the birth of catastrophic environmentalism. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-974005-4. x + 298 pp. LCCN GE180 .H35 2013.

Hansen:2013:SMM

[3063] Peter H. Hansen. The summits of modern man: mountaineering after the enlightenment. Harvard University Press, Cambridge, MA, USA, 2013. ISBN 0-674-04799-0. x + 380 pp. LCCN GV199.89 .H36 2013.

Hart:2013:ICC

[3064] Roger (Roger Preston) Hart. Imagined civilizations: China, the West, and their first encounter. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-0606-3. vii + 374 pp. LCCN BV3417 .H37 2013.

Harwit:2013:STU

[3065] Martin Harwit. In search of the true universe: the tools, shaping, and cost of cosmological thought. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-04406-5 (hardcover), 1-107-62025-2 (paperback). xvii + 393 pp. LCCN QB32 .H27 2013.

Hilbert:2013:DHL

[3066] David Hilbert, William Ewald, Wilfried Sieg, Michael Hallett, Ulrich Majer, and Dirk Schlimm. David Hilbert’s lectures on the foundations of arithmetic and logic 1917–19. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-540-20578-0. xxv + 1062 pp. LCCN QA248 .H54 2013.

Hoskin:2013:CHP

[3067] Michael A. Hoskin. Caroline Herschel: priestess of the new heavens. Science History Publications/USA, Sagamore Beach, MA, USA, 2013. ISBN 0-88135-436-8 (hardcover), 0-88135-412-0. xiii + 256 pp. LCCN QB36.H58 H67 2013.

Hunter:2013:WIV

[3068] Matthew C. Hunter. Wicked intelligence: visual art and the science of experiment in Restoration London. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-01729-X (hardcover), 0-226-01732-X (ebook). xv + 329 + 10 pp. LCCN Q222 .H86 2013.

Ivey:2013:RHU

[3069] Paul Eli Ivey. Radiance from Halcyon: a Utopian experiment in religion and science. University of Minnesota Press, Minneapolis, MN, USA, 2013. ISBN 0-8166-8050-7 (hardcover), 0-8166-8051-5 (paperback). xiii + 312 pp. LCCN BP565.I94 R33 2013.

Jacob:2013:FKE

[3070] Margaret C. Jacob. The first knowledge economy: Human capital and the European economy, 1750–1850. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-61983-1 (paperback), 1-107-04401-4 (hardcover). ix + 257 pp. LCCN HC240.

Jolley:2013:CME

[3071] Nicholas Jolley. Causality and mind: essays on early modern philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-966955-4. ix + 279 pp. LCCN B801 .J65 2013. URL http://

www.loc.gov/catdir/enhancements/fy1506/2013953478-b.html;
http://www.loc.gov/catdir/enhancements/fy1506/2013953478-d.
html;   http://www.loc.gov/catdir/enhancements/fy1506/2013953478-
t.html.

Jones:2013:MTC

[3072] Claire L. Jones. The medical trade catalogue in Britain, 1870–1914. Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-443-2 (hardcover), 1-78144-067-0 (ebook). xii + 264 pp. LCCN R487 .J66 2013.

Josephson:2013:EHR

[3073] Paul Josephson, Nicolai Dronin, Ruben Mnatsakanian, Aleh Cherp, Dmitry Efremenko, and Vladislav Larin. An environmental history of Russia. Studies in environment and history. Cambridge University Press, Cambridge, UK, 2013. ISBN 0-521-86958-7 (hardcover), 0-521-68972-4 (paperback), 1-139-02104-4 (e-book). vii + 340 pp. LCCN GF602.7 .J67 2013.

Kaiserfeld:2013:LEB

[3074] Thomas Kaiserfeld and Tom O’Dell, editors. Legitimizing ESS: big science as a collaboration across boundaries. Nordic Academic Press, Lund, Sweden, 2013. ISBN 91-87351-10-2. 240 + 16 pp. LCCN QC794.6.S3 L44 2013.

Kardel:2013:NSB

[3075] Troels Kardel and Paul Maquet. Nicolaus Steno: Biography and Original Papers of a Seventeenth-Century Scientist. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. xxxviii + 739 pp.

King:2013:OSB

[3076] Helen King. The one-sex body on trial: the classical and early modern evidence. The history of medicine in context. Ashgate, Farnham, Surrey, UK, 2013. ISBN 1-4094-6335-4 (hardcover), 1-4094-6336-2 (e-PDF), 1-4094-6337-0 (e-PUB). xii + 274 pp. LCCN QP278 .K56 2013.

Kremer:2013:JHH

[3077] Richard L. (Richard Lynn) Kremer and Jarosław Włodarczyk. Johannes Hevelius and his world: astronomer, cartographer, philosopher and correspondent, volume 44 of Studia Copernicana. In-t Historii Nauki PAN, Warsaw, Poland, 2013. ISBN 83-86062-94-0. ISSN 0081-6701. vii + 1 + 235 pp. LCCN QB36.H65 J64 2013.

Lambert:2013:MNJ

[3078] David (Historian) Lambert. Mastering the Niger: James MacQueen’s African geography and the struggle over Atlantic slavery. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-07806-X (hardcover). viii + 309 pp. LCCN DT360 .L24 2013.

Latour:2013:IME

[3079] Bruno Latour. An inquiry into modes of existence: an anthropology of the moderns. Harvard University Press, Cambridge, MA, USA, 2013. ISBN 0-674-72499-2, 0-674-72855-6 (e-book). xxvii + 486 pp. LCCN

CB358 .L27813 2013. URL http://www.modesofexistence.org/.

Longino:2013:SHB

[3080] Helen E. Longino. Studying human behavior: how scientists investigate aggression and sexuality. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-49287-7 (hardcover), 0-226-49288-5 (paperback), 0-226-92182-4 (e-book). xi + 249 pp. LCCN BF199 .L58 2013.

Madrinan:2013:BRB

[3081] Santiago Madriñán. Book Review: Book Review: Nikolaus Joseph Jacquin’s American plants: botanical expedition to the Caribbean (1754– 1759) and the publication of the Selectarum stirpium Americanarum historia. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-23410-1 (hardcover), 90-04-23411-X (e-book). xi + 425 + 1 pp. LCCN QK225 .M33 2012.

Malcolmson:2013:SSC

[3082] Cristina Malcolmson. Studies of skin color in the early Royal Society: Boyle, Cavendish, Swift. Literary and Scientific Cultures of Early Modernity. Ashgate, Burlington, VT, USA, 2013. ISBN 0-7546-3778-6 (hardcover), 1-4094-6216-1 (e-book), 1-4724-0520-X (e-pub). xii + 233 + 1 pp. LCCN Q41.L86 M35 2013.

Marsden:2013:UCE

[3083] Ben Marsden, Hazel Hutchison, and Ralph O’Connor, editors. Uncommon contexts: encounters between science and literature, 1800–1914, volume 23 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-362-2 (hardcover). xiii + 239 pp. LCCN PR468.S34 U53 2013.

Mattern:2013:PMG

[3084] Susan P. Mattern. The prince of medicine: Galen in the Roman Empire. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-960545-9 (hardcover). xx + 334 + 8 pp. LCCN B577.G24 M38 2013.

Mayer:2013:RIP

[3085] Thomas F. (Thomas Frederick) Mayer. The Roman Inquisition: a papal bureaucracy and its laws in the age of Galileo. Haney Foundation series.

University of Pennsylvania Press, Philadelphia, 2013. ISBN 0-8122-4473-7 (hardcover), 0-8122-0764-5 (e-book). 381 pp. LCCN BX1723 .M38 2013.

McGrath:2013:IAO

[3086] Mary Quinlan McGrath. Influences: Art, Optics and Astrology in the Italian Renaissance. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-92284-7 (hardcover), 0-226-92285-5 (e-book). xi + 284 + 12 pp. LCCN N6915 .M42 2013.

Mensch:2013:KOE

[3087] Jennifer Mensch. Kant’s organicism: epigenesis and the development of critical philosophy. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-02198-X (hardcover), 0-226-02203-X (e-book). xi + 246 pp. LCCN B2799.N37 M46 2013.

Miller:2013:NBE

[3088] Ian Jared Miller. The nature of the beasts: empire and exhibition at the Tokyo Imperial Zoo. Asia–local studies/global themes. University of California Press, Berkeley, CA, USA, 2013. ISBN 0-520-27186-6 (hardcover), 0-520-95210-3 (e-book). xxvii + 322 pp. LCCN QL76.5.J32 M55 2013.

Mitchinson:2013:BFM

[3089] Wendy Mitchinson. Body failure: medical views of women, 1900–1950. University of Toronto Press, Toronto, ON, Canada, 2013. ISBN 1-4426-1431-5 (paperback), 1-4426-4626-8 (hardcover). xiii + 414 + 24 pp. LCCN RA564.85.

Nadis:2013:HSY

[3090] Steve Nadis and Shing-Tung Yau. A history in sum: 150 years of mathematics at Harvard (1825–1975). Harvard University Press, Cambridge, MA, USA, 2013. ISBN 0-674-72500-X, 0-674-72655-3 (e-book). xviii + 249 + 10 pp. LCCN QA13.5.M43 H376 2013eb.

Nasim:2013:OHS

[3091] Omar W. Nasim. Observing by hand: sketching the nebulae in the nineteenth century. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-08437-X (hardcover), 0-226-08440-X (e-book). xii + 304 pp. LCCN QB32 .N37 2013.

Nicoli:2013:SLA

[3092] Miriam Nicoli. Les savants et les livres: autour d’Albrecht von Haller (1708–1777) et Samuel-Auguste Tissot (1728–1797), volume 16 of Travaux sur la Suisse des Lumières. Slatkine, Genève, Switzerland, 2013. ISBN 2-05-102509-6 (hardcover). 365 pp. LCCN Q127.S9 N53 2013.

Panti:2013:BRB

[3093] Cecilia Panti, Tom McLeish, Greti Dinkova-Bruun, Giles E. M. (Giles Edward Murray) Gasper, M. (Michael) Huxtable, and Hannah Smithson. Book Review: Book Review: The dimensions of colour: Robert Grosseteste’s De colore, volume 4 of Durham medieval and renaissance texts. Pontifical Institute of Mediaeval Studies, Toronto, ON, Canada, 2013. ISBN 0-88844-564-4 (paperback). x + 94 pp. LCCN B765.G73 D4313 2013.

Paul:2013:PPS

[3094] Diane B. Paul and Jeffrey P. (Jeffrey Paul) Brosco. The PKU paradox: a short history of a genetic disease. Johns Hopkins biographies of disease. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-1131-8 (paperback), 1-4214-1132-6 (electronic). xxiv + 289 pp. LCCN RC632.H87 P38 2013.

Pettit:2013:SDP

[3095] Michael John Pettit. The science of deception: psychology and commerce in America. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-92374-6 (hardcover), 0-226-92375-4 (e-book). ix + 311 pp. LCCN BF637.D42 P48 2013.

Phalkey:2013:ASB

[3096] Jahnavi Phalkey. Atomic state: big science in Twentieth-Century India. The Indian century. Permanent Black, Ranikhet, India, 2013. ISBN 81-7824-376-8. xvii + 335 pp. LCCN QC773.3.I4 P43 2013.

Pilaski:2013:MKA

[3097] Katharina Pilaski. The Munich Kunstkammer: art, nature, and the representation of knowledge in courtly contexts, volume 73 of Spätmittelalter, Humanismus, Reformation, Studies in the late middle ages, humanism and the reformation; 73. Mohr Siebeck, Tübingen, Germany, 2013. ISBN 3-16-152188-9 (hardcover). ISSN 1865–2840. viii + 212 pp. LCCN N5267.A43 P55 2013.

Protevi:2013:LWE

[3098] John Protevi. Life, war, earth: Deleuze and the sciences. University of Minnesota Press, Minneapolis, MN, USA, 2013. ISBN 0-8166-8101-5 (hardcover), 0-8166-8102-3 (paperback), 0-8166-8449-9 (e-book), 1-4619-4294-2 (e-book). ix + 256 pp. LCCN B2430.D454 P76 2013.

Raber:2013:ABR

[3099] Karen Raber. Animal bodies, Renaissance culture. Haney Foundation series. University of Pennsylvania Press, Philadelphia, 2013. ISBN 0-8122-4536-9 (hardcover). 234 pp. LCCN B105.A55 R33 2013.

Rice:2013:STP

[3100] Prudence M. Rice. Space-time perspectives on early colonial Moquegua. University Press of Colorado, Boulder, Colorado, 2013. ISBN 1-60732-275-7 (hardcover), 1-60732-276-5 (e-book). xx + 378 pp. LCCN F3429.1.M65 R53 2013.

Richards:2013:WHD

[3101] Robert J. (Robert John) Richards. Was Hitler a Darwinian?: disputed questions in the history of evolutionary theory. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-05876-X (hardcover), 0-226-05893-X (paperback), 0-226-05909-X (e-book). 269 pp. LCCN QH361 .R534 2013.

Richardson:2013:SIS

[3102] Sarah S. Richardson. Sex itself: the search for male and female in the human genome. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-08468-X (hardcover), 0-226-08471-X (e-book). vii + 311 pp. LCCN QH600.5 .R53 2013.

Roberts:2013:WWA

[3103] Siobhan Roberts. Wind wizard: Alan G. Davenport and the art of wind engineering. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-15153-9 (hardcover), 1-4008-4470-3. 278 pp. LCCN TA654.5 .R636 2013.

Ruse:2013:GHS

[3104] Michaël Ruse. The Gaia hypothesis: science on a pagan planet. Science.culture. University of Chicago Press, Chicago, IL, USA, online-ausg. edition, 2013. ISBN 0-226-73170-7 (hardcover), 0-226-06039-X (e-book), 1-299-80362-8 (e-book). xiv + 251 pp. LCCN QH331 .R8785 2013.

Sakurai:2013:SSF

[3105] Ayako Sakurai. Science and societies in Frankfurt am Main, volume 20 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-397-5 (hardcover), 1-78144-030-1 (ebook). xi + 235 pp. LCCN Q49 .S36 2013.

Seim:2013:RPM

[3106] David L. Seim. Rockefeller philanthropy and modern social science. Studies in business history. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-391-6 (hardcover), 1-78144-026-3 (e-book). ix + 265 pp. LCCN HV97.R6 S44 2013.

Shehada:2013:MAV

[3107] Housni Alkhateeb Shehada. Mamluks and animals: veterinary medicine in medieval Islam. Sir Henry Wellcome Asian series; volume 11. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-23405-5 (hardcover), 90-04-23422-5 (e-book). xxii + 237 + 32 pp. LCCN SF713.I742 A44 2013.

Sieg:2013:HPB

[3108] Wilfried Sieg. Hilbert’s programs and beyond. Logic and computation in philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-537222-0 (hardcover), 0-19-970715-4 (e-book). xii + 439 pp. LCCN QA8.4 .S545 2013.

Slayton:2013:ACP

[3109] Rebecca Slayton. Arguments that count: physics, computing, and missile defense, 1949–2012. Inside technology. MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-01944-2 (hardcover). xi + 325 pp. LCCN UA22 .S57 2013.

Smith:2013:ESP

[3110] Shawn Michelle Smith. At the edge of sight: photography and the unseen. Duke University Press, Durham, NC, USA, 2013. ISBN 0-8223-5486-1 (hardcover), 0-8223-5502-7 (paperback). xvii + 293 pp. LCCN TR23 .S626 2013.

Smyrek:2013:GTT

[3111] Volker Smyrek. Die Geschichte des Tonmischpults: die technische Entwicklung der Mischpulte und der Wandel der medialen Produktionsverfahren im Tonstudio von den 1920er-Jahren bis heute. (German) [The history of the audio mixer: the technical development of mixing consoles and the transformation of media production in the studio of the 1920s until today]. Logos, Berlin, Germany, 2013. ISBN 3-8325-3402-4 (paperback). xxix + 539 pp. LCCN TK7881.4 .S69 2013.

Stark:2013:MMA

[3112] James F. Stark. The making of modern anthrax, 1875–1920: uniting local, national and global histories of disease, volume 21 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-446-7 (hardcover), 1-78144-070-0 (e-book). xi + 251 pp. LCCN QR201.A6 S83 2013.

Stevens:2013:LSD

[3113] Hallam Stevens. Life out of sequence: a data-driven history of bioinformatics. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-08017-X (hardcover), 0-226-08020-X (paperback), 0-226-08034-X (e-book). 294 pp. LCCN QH324.2 .S726 2013.

Stone:2013:EQQ

[3114] A. Douglas Stone. Einstein and the quantum: the quest of the valiant Swabian. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-13968-7 (hardcover), 1-4008-4834-2 (e-book), 1-299-87423-1 (e-book). x + 332 pp. LCCN QC16.E5 S76 2013.

URL http://www.npr.org/2013/11/01/242356997/einsteins-real-
breakthrough-quantum-theory.

Tega:2013:TRS

[3115] Walter Tega. Tradizione e rivoluzione: scienza e potere in Francia, 1815– 1840. (Italian) [Tradition and Revolution: Science and power in France, 1815–1840], volume 11 of Pansophia. L. S. Olschki, Firenze, Italia, 2013. ISBN 88-222-6296-4 (paperback). ix + 346 pp. LCCN AE.

Thompson:2013:GSE

[3116] Charis Thompson. Good science: the ethical choreography of stem cell research. MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-02699-6 (hardcover). x + 343 pp. LCCN QH588.S83 T48 2013.

Timmermann:2013:VTC

[3117] Anke Timmermann. Verse and transmutation: a corpus of Middle English alchemical poetry (critical editions and studies), volume 42 (21) of History of science and medicine library, ; Medieval and early modern science. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-25484-6 (hardcover), 90-04-25483-8 (e-book). ISSN 1872-0684. xiv + 374 pp. LCCN QD26 .T63 2013.

Tognoni:2013:OGG

[3118] Federico Tognoni. Le opere di Galileo Galilei. Volume 1, Appendice: edizione nazionale: iconografia galileiana. Giunti Editore, Firenze, Italia, 2013. ISBN 88-09-77618-6 (paperback). 603 + 16 pp. LCCN QB3 .G32

Usitalo:2013:IML

[3119] Steven A. Usitalo. The invention of Mikhail Lomonosov: a Russian national myth. Imperial encounters in Russian history. Academic Studies Press, Boston, MA, USA, 2013. ISBN 1-61811-173-6 (hardcover), 1-61811-195-7 (e-book), 1-299-95425-1 (e-book). 298 pp. LCCN PG3316.Z6 U85 2013.

Valencius:2013:LHN

[3120] Conevery Bolton Valenčius. The lost history of the New Madrid earthquakes. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-05389-X (hardcover), 0-226-05392-X (e-book), 1-299-82003-4 (ebook). ix + 460 pp. LCCN QE535.2.U6 V354 2013.

Veit:2013:MFM

[3121] Helen Zoe Veit. Modern food, moral food: self-control, science, and the rise of modern American eating in the early twentieth century. The University of North Carolina Press, Chapel Hill, NC, USA, 2013. ISBN 1-4696-0770-0 (hardcover), 1-4696-1275-5 (e-book). xiii + 300 pp. LCCN

RA418; TX360.U6 V45 2013. URL http://www.jstor.org/stable/
10.5149/9781469607719_Veit.

vonSchwerin:2013:BHA

[3122] Alexander von Schwerin, Heiko Stoff, and Bettina Wahrig, editors. Biologics: a history of agents made from living organisms in the twentieth century. Studies for the Society for the Social History of Medicine. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-430-0 (hardcover), 1-78144-055-7 (e-book). xvii + 260 pp. LCCN RS162 .B56 2013.

Wallace:2013:OLC

[3123] Alfred Russel Wallace and James T. Costa. On the organic law of change: a facsimile edition and annotated transcription of Alfred Russel Wallace’s Species notebook of 1855–1859. Harvard University Press, Cambridge, MA, USA, 2013. ISBN 0-674-72488-7. xii + 573 pp. LCCN QH375 .W35 2013.

Weill-Parot:2013:PAN

[3124] Nicolas Weill-Parot. Points aveugles de la nature. La rationalité scientifique médiévale face à l’occulte, l’attraction magnétique et l’horreur du vide (XIIIe-milieu du XVe siècle). Histoire. Editions Belles Lettres, Le Kremlin Bicêtre, France, 2013. ISBN 2-251-38120-1. 656 pp. LCCN D900.

Weitekamp:2013:PSI

[3125] Margaret A. Weitekamp, David H. DeVorkin, and Diane Kidd. Pluto’s secret: an icy world’s tale of discovery. Abrams Books for Young Readers, New York, NY, USA, 2013. ISBN 1-4197-0423-0 (hardcover). 39 pp. LCCN QB701 .W45 2013.

Williams:2013:NSS

[3126] Glyndwr Williams. Naturalists at sea: scientific travellers from Dampier to Darwin. Yale University Press, New Haven, CT, USA, 2013. ISBN 0-300-18073-X (hardcover), 0-300-18220-1 (e-book), 1-299-84127-9 (MyiLibrary). xv + 309 + 24 pp. LCCN QH11 .W45 2013.

Williams:2013:THE

[3127] Rosalind H. Williams. The triumph of human empire: Verne, Morris, and Stevenson at the end of the world. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-89955-1 (hardcover), 0-226-89958-6 (e-book). xi + 416 pp. LCCN PN3435 .W495 2013.

Yeang:2013:PSR

[3128] Chen-Pang Yeang. Probing the sky with radio waves: from wireless technology to the development of atmospheric science. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-01519-X (hardcover), 0-226-03481-X (e-book). xv + 361 pp. LCCN QC676.4 .Y43 2013.

Yoder:2013:BRB

[3129] Joella Gerstmeyer Yoder, editor. Book Review: Book Review: Catalogue of the manuscripts of Christiaan Huygens including a concordance with his Œuvres complètes, volume 35 of History of science and medicine library. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-23564-7, 90-04-23565-5. xii + 338 pp. LCCN QA33 .C38 2013eb.

Ziche:2013:SSK

[3130] Paul Ziche, Petr Rezvykh, and Daniel A. Di Liscia. Sygkepleriazein: Schelling und die Kepler-Rezeption im 19. Jahrhundert. (German) [Sygkepleriazein: Schelling and the Kepler reception in the 19th century], volume 21 of Schellingiana. Frommann-Holzboog, Stuttgart, Germany, 2013. ISBN 3-7728-2441-2 (paperback), 3-7728-3052-8 (e-book). vi + 299 pp. LCCN B2899 .N3 Z53 2013.

Zuidervaart:2013:BRR

[3131] H. J. Zuidervaart and Robert H. van Gent. Between rhetoric and reality: instrumental practices at the astronomical observatory of the Amsterdam society ‘Felix Meritis’, 1786–1889. Verloren, Hilversum, The Netherlands, 2013. ISBN 90-8704-363-5. 152 pp. LCCN QB82.N4 Z853 2013.

Adkins:2014:ISF

[3132] G. Matthew (Gregory Matthew) Adkins. The idea of the sciences in the French Enlightenment: a reinterpretation. University of Delaware Press, Newark, DE, USA, 2014. ISBN 1-61149-474-5 (hardcover), 1-61149-475-3 (electronic). x + 163 pp. LCCN Q127.F8 A35 2014.

Bashford:2014:GPH

[3133] Alison Bashford. Global population: history, geopolitics, and life on earth. Columbia studies in international and global history. Columbia University Press, New York, NY, USA, 2014. ISBN 0-231-14766-X (hardcover). xii + 466 pp. LCCN HB849.44 .B37 2014.

Beretta:2014:FHC

[3134] Marco Beretta and Maria Conforti, editors. Fakes!?: hoaxes, counterfeits, and deception in early modern science. Science History Publications/USA, Sagamore Beach, MA, USA, 2014. ISBN 0-88135-495-3. xv + 280 pp. LCCN Q175.37 .F35 2014.

Binnema:2014:EZH

[3135] Theodore Binnema. “Enlightened zeal”: the Hudson’s Bay Company and scientific networks, 1670–1870. University of Toronto Press, Toronto, ON, Canada, 2014. ISBN 1-4426-1475-7, 1-4426-4697-7. xvi + 458 pp. LCCN F1060.7.S35 B55 2014.

Camprubi:2014:EMF

[3136] Lino Camprubí. Engineers and the making of the Francoist regime. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02717-8. xiii + 298 pp. LCCN TA87 .C35 2014.

Cohen-Cole:2014:OMC

[3137] Jamie Nace Cohen-Cole. The open mind: Cold War politics and the sciences of human nature. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-09216-X, 0-226-09233-X. 406 pp. LCCN ????

Dawson:2014:VSN

[3138] Gowan Dawson and Bernard V. Lightman, editors. Victorian scientific naturalism: community, identity, continuity. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-10950-X (hardcover), 0-226-10964-X (e-book). viii + 345 pp. LCCN Q127.G4 V454 2014.

Drouin:2014:PIF

[3139] Jean-Marc Drouin. Philosophie de l’insecte. (French) [Philosophy of the insect]. Science ouverte. Éditions du Seuil, Paris, France, 2014. ISBN 2-02-111889-4. 249 pp. LCCN QL497 .D76 2014.

Dry:2014:NPS

[3140] Sarah Dry. The Newton Papers: the Strange and True Odyssey of Isaac Newton’s Manuscripts. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-995104-7 (hardcover). xi + 238 pp. LCCN QC16.N7 D79 2014.

Eckes:2014:BRB

[3141] Christophe Eckes, Amaury Thuillier, and Erhard Scholz. Book Review: Book Review: Les groupes de Lie dans l’oeuvre de Hermann Weyl: traduction et commentaire de l’article Théorie de la représentation des groupes continus semi-simples par des transformations linéaires (1925– 1926). Histoires de géométries. Presses universitaires de Nancy, Nancy, France, 2014. ISBN 2-8143-0180-2 (paperback). ISSN 2115-6492. 401 pp. LCCN ????

Ferngren:2014:MRH

[3142] Gary B. Ferngren. Medicine and religion: a historical introduction. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1215-2 (hardcover), 1-4214-1216-0 (paperback), 1-4214-1217-9 (electronic). xii + 241 pp. LCCN BL65.M4 F47 2014.

Foltz:2014:NNE

[3143] Bruce V. Foltz. The noetics of nature: environmental philosophy and the holy beauty of the visible. Groundworks: ecological issues in philosophy and theology. Fordham University Press, New York, NY, USA, 2014. ISBN 0-8232-5464-X (hardcover), 0-8232-5465-8 (paper). xix + 296 pp. LCCN BD581 .F65 2014.

Forbes:2014:FME

[3144] Nancy Forbes and Basil Mahon. Faraday, Maxwell, and the electromagnetic field: how two men revolutionized physics. Prometheus Books, Amherst, NY, USA, 2014. ISBN 1-61614-942-6 (hardcover), 1-61614-943-4 (e-book). 320 pp. LCCN QC16.F2 F67 2014.

Gascoigne:2014:EPA

[3145] John Gascoigne. Encountering the Pacific in the Age of Enlightenment. Cambridge University Press, Cambridge, UK, 2014. ISBN 0-521-87959-0 (hardcover). xxxv + 542 pp. LCCN DU28 .G37 2014; GN368.

Gee:2014:FWD

[3146] Brian Gee, Anita McConnell, and A. D. Morrison-Low, editors. Francis Watkins and the Dollond telescope patent controversy. Science, Technology and Culture. Ashgate, Surrey, England, 2014. ISBN 1-4094-6643-4 (hardcover), 1-4094-6644-2 (e-book), 1-4094-6645-0 (e-book). xxvi + 392 + 12 pp. LCCN QB88 .G44 2014eb.

GonzalezdeLinares:2014:OC

[3147] Augusto González de Linares and Carlos Nieto, editors. Obra completa, volume 2 of Coedciones. Universidad de Cantabria, Servicio de Publicaciones, Santander, Spain, 2014. ISBN 84-8102-697-2. 591 pp. LCCN ????

Grim:2014:ER

[3148] John Grim and Mary Evelyn Tucker. Ecology and religion. Foundations of contemporary environmental studies. Island Press, Washington, DC, USA, 2014. ISBN 1-59726-707-4 (hardcover), 1-59726-708-2 (paperback). x + 265 pp. LCCN BL65.E36 G75 2014.

Gross:2014:SSI

[3149] Alan G. Gross and Joseph E. Harmon. Science from sight to insight: how scientists illustrate meaning. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-06820-X (hardcover), 0-226-06848-X (paperback). vii + 332 pp. LCCN Q223 .G772 2014.

Guttler:2014:KKI

[3150] Nils Güttler. Das Kosmoskop: Karten und ihre Benützer in der Pflanzengeographie des 19. Jahrhunderts. (German) [The Cosmoscope: cards and their users in 19th Century plant geography]. Wallstein, Göttingen, Germany, 2014. ISBN 3-8353-1429-7 (hardcover). 545 pp. LCCN QK101 G88 2014x.

Heit:2014:HNW

[3151] Helmut Heit and Lisa Heller. Handbuch Nietzsche und die Wissenschaften: Natur-, geistes- und sozialwissenschaftliche Kontexte. (German) [Handbook of Nietzsche and the sciences: natural sciences, humanities and social science contexts]. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 2014. ISBN 3-11-028578-9 (hardcover), 3-11-028569-X, 3-11-028568-1 (e-book). ix + 395 pp. LCCN B3317 .H255

Howe:2014:BCS

[3152] Joshua P. Howe. Behind the curve: science and the politics of global warming. Weyerhaeuser environmental classics. University of Washington Press, Seattle, WA, USA, 2014. ISBN 0-295-99368-5, 0-295-80509-9 (e-book). xviii + 290 pp. LCCN QC981.8.G56 H69 2014. URL http://

muse.jhu.edu/books/9780295805092/.

Josephson:2014:CRA

[3153] Paul R. Josephson. The Conquest of the Russian Arctic. Harvard University Press, Cambridge, MA, USA, 2014. ISBN 0-674-41982-0. viii + 441 pp. LCCN G615 .J67 2014.

Kaye:2014:HBE

[3154] Joel Kaye. A history of balance, 1250–1375: the emergence of a new model of equilibrium and its impact on thought. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-02845-0 (paperback). ix + 519 pp. LCCN B721 .K35 2014. URL http://www.loc.

gov/catdir/enhancements/fy1408/2013034173-b.html;   http://
www.loc.gov/catdir/enhancements/fy1408/2013034173-d.html;
http://www.loc.gov/catdir/enhancements/fy1408/2013034173-t.
html.

Kim:2014:QSE

[3155] Yong sik Kim. Questioning science in East Asian contexts: essays on science, Confucianism, and the comparative history of science, volume 1 of Science and religion. Brill, Leiden, The Netherlands, 2014. ISBN 90-04-26509-0 (hardcover), 90-04-26531-7 (e-book). ISSN 2214-8329. vi + 284 pp. LCCN Q127.E26 K56 2014.

Koblitz:2014:SHB

[3156] Ann Hibner Koblitz. Sex and herbs and birth control: women and fertility regulation through the ages. Kovalevskaia Fund, Seattle, WA, USA, 2014. ISBN 0-9896655-0-X. viii + 345 pp. LCCN RM666.H33 K62 2014.

Kurihara:2014:CWR

[3157] Ken Kurihara. Celestial wonders in Reformation Germany, volume 13 of Religious cultures in the early modern world. Pickering and Chatto, London, UK, 2014. ISBN 1-84893-444-0 (hardcover). xi + 211 pp. LCCN BR855 .K87 2014.

Macrakis:2014:PLS

[3158] Kristie Macrakis. Prisoners, lovers, and spies: the story of invisible ink from Herodotus to al-Qaeda. Yale University Press, New Haven, CT, USA, 2014. ISBN 0-300-17925-1 (hardcover). xiv + 377 pp. LCCN Z104.5 .M33 2014.

Maienschein:2014:EUM

[3159] Jane Maienschein. Embryos under the microscope: the diverging meanings of life. Harvard University Press, Cambridge, MA, USA, 2014. ISBN 0-674-72555-7 (hardcover). x + 336 pp. LCCN QM603 .M35 2014.

Manjapra:2014:AEG

[3160] Kris Manjapra. Age of entanglement: German and Indian intellectuals across empire, volume 183 of Harvard historical studies. Harvard University Press, Cambridge, MA, USA, 2014. ISBN 0-674-72514-X (hardcover). 442 pp. LCCN DS428 .M36 2014.

Mayer:2014:RIS

[3161] Thomas F. (Thomas Frederick) Mayer. The Roman Inquisition on the stage of Italy, c. 1590–1640. University of Pennsylvania Press, Philadelphia, 2014. ISBN 0-8122-4573-3 (hardcover). 361 pp. LCCN BX1723 .M385 2014.

Mazur:2014:ESS

[3162] Joseph Mazur. Enlightening symbols: a short history of mathematical notation and its hidden powers. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-15463-5 (hardcover). xxiii + 285 pp. LCCN QA41 .M39 2014.

McGann:2014:NRL

[3163] Jerome J. McGann. A new republic of letters: memory and scholarship in the age of digital reproduction. Harvard University Press, Cambridge, MA, USA, 2014. ISBN 0-674-72869-6. x + 238 pp. LCCN AZ186 .M35 2014.

Nicoud:2014:PMP

[3164] Marilyn Nicoud and Danielle Jacquart. Le prince et les médecins: pensée et pratiques médicales à Milan (1402–1476), volume 488 of Collection de l’École française de Rome. École française de Rome, Rome, Italy, 2014. ISBN 2-7283-0955-8 (paperback). ISSN 0223-5099. xi + 804 pp. LCCN R519.M5 N53 2014.

Pesic:2014:MMM

[3165] Peter Pesic. Music and the making of modern science. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02727-5 (hardcover). viii + 347 pp. LCCN Q172.5.M87 P47 2014.

Rabault-Feuerhahn:2014:AOS

[3166] Pascale Rabault-Feuerhahn, Dominique Bach, and Rick Willet. Archives of Origins: Sanskrit, Philology, Anthropology in 19th Century Germany, volume 9 of Kultur- und sozialwissenschaftliche Studien /Studies in Cultural and Social Sciences. Harrassowitz Verlag, Wiesbaden, Germany, 2014. ISBN 3-447-19136-8. 337 pp. LCCN PK411.G3; PQ283.

Reed:2014:ARR

[3167] Peter Reed. Acid rain and the rise of the environmental chemist in nineteenth-century Britain: [the life and work of Robert Angus Smith]. Science, technology and culture. Ashgate, Farnham, UK, 2014. ISBN 1-4094-5775-3 (hardcover), 1-4724-2720-3 (PDF e-book), 1-4724-2721-1 (ePub e-book). xv + 209 pp. LCCN QD22.S615 R33 2014.

Schwarz:2014:EP

[3168] Astrid E. Schwarz. Experiments in practice, volume 2 of History and philosophy of technoscience. Pickering and Chatto, London, UK, 2014. ISBN 1-84893-485-8, 1-78144-450-1 (e-book). vii + 257 pp. LCCN Q175 .S39 2014.

Sheller:2014:ADM

[3169] Mimi Sheller. Aluminum dreams. The making of light modernity. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02682-1. xii + 367 pp. LCCN TA480.A6 S48 2014.

Terrall:2014:CNA

[3170] Mary Terrall. Catching nature in the act: Réaumur and the practice of natural history in the Eighteenth Century. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-08860-X (hardcover). viii + 275 pp. LCCN QH15 .T47 2014.

vonHumboldt:2014:VN

[3171] Alexander von Humboldt. Views of nature. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-92318-5 (hardcover), 0-226-92319-3 (e-book). viii + 313 pp. LCCN Q171 .H8613 2014. Edited by Stephen T. Jackson, Laura Dassow Walls, and Mark W. Person.

Zanoni:2014:SPR

[3172] Elena Zanoni. Scienza patria religione: Antonio Stoppani e la cultura italiana dell’Ottocento. (Italian) [Science, homeland, religion: Antonio Stoppani and Italian culture of the Ninth Century], volume 431 of Studi e ricerche storiche. Franco Angeli, Milano, Italia, 2014. ISBN 88-917-0541-1. 301 pp. LCCN QE22.S8 Z36 2014.

Bartusiak:2015:BHH

[3173] Marcia Bartusiak. Black hole: how an idea abandoned by Newtonians, hated by Einstein, and gambled on by Hawking became loved. Yale University Press, New Haven, CT, USA, 2015. ISBN 0-300-21085-X (hardcover). xii + 237 pp. LCCN QB843.B55 B37 2015.

Campos:2015:RSL

[3174] Luis A. Campos. Radium and the secret of life. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-23827-X (hardcover), 0-226-23830-X (e-book). 378 pp. LCCN QD181.R1 C36 2015.

Gimbel:2015:EHS

[3175] Steven Gimbel. Einstein: his space and times. Jewish lives. Yale University Press, New Haven, CT, USA, 2015. ISBN 0-300-19671-7 (hardcover), 0-300-21361-1 (e-book). ix + 191 pp. LCCN QC16.E5 G53 2015.

Bruzzaniti:2016:EFO

[3176] Giuseppe Bruzzaniti. Enrico Fermi: the Obedient Genius. Springer Biographies. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 1-4939-3531-3 (hardcover), 1-4939-3533-X (e-book). xiii + 348 + 54 pp. LCCN QC774.F4. URL http://link.

springer.com/book/10.1007/978-1-4939-3533-8.

Kragh:2016:VGD

[3177] Helge Kragh, editor. Varying gravity: Dirac’s legacy in cosmology and geophysics, volume 54 of Science networks. Historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2016. ISBN 3-319-24377-2 (hardcover), 3-319-24379-9 (e-book). xi + 185 pp. LCCN QB981.

Pfeiffer:1967:HRG

[3178] John Pfeiffer and Donald R. Rooney. The history of radiology in Georgia. Georgia Radiological Society, Atlanta, GA, USA, 1967. 185 pp. LCCN RM849 .P47.

vonHumboldt:1975:VN

[3179] Alexander von Humboldt. Views of nature. History, philosophy, and sociology of science. Arno Press, New York, NY, USA, 1975. ISBN 0-405-06626-0. xxx + 452 pp. LCCN Q161 .H8313 1975.

Rouner:1984:N

[3180] Leroy S. Rouner, editor. On nature, volume 6 of Boston University studies in philosophy and religion. University of Notre Dame Press, Notre Dame, IN, USA, 1984. ISBN 0-268-01499-X. xvi + 188 pp. LCCN BD581 .O5 1984. US$22.95.

Burkhardt:1985:CCC

[3181] Frederick Burkhardt, Sydney Smith, David Kohn, William Montgomery, Charles Darwin, Correspondence, et al., editors. A calendar of the correspondence of Charles Darwin, 1821–1882, volume 369 of Garland reference library of the humanities. Garland Pub., New York, NY, USA, 1985. ISBN 0-8240-9224-4. 690 + 1 pp. LCCN QH31.D2 C15 1985.

Burkhardt:1985:CCD

[3182] Frederick Burkhardt and Sydney Smith, editors. The correspondence of Charles Darwin. Cambridge University Press, Cambridge, UK, 1985– 2010. ISBN 0-521-25587-2 (vol. 1), 0-521-25588-0 (vol. 2), 0-521-25589-9 (vol. 3), 0-521-25590-2 (vol. 4), 0-521-25591-0 (vol. 5), 0-521-25586-4 (vol. 6), 0-521-84459-2 (vol. 14), 0-521-85931-X (vol. 15 : hardback), 0-521-88195-1 (vol. 16, pt. 1 : hardback), 0-521-88196-X (vol. 16, pt. 2 : hardback), 0-521-51836-9 (vol. 16, 2 vol. hardback set), 0-521-19030-4 (vol. 17), 0-521-76889-6 (vol. 18), 0-521-44241-9. ???? pp. LCCN QH31.D2 A4 1985.

Schmitt:1990:AOT

[3183] Natalie Crohn Schmitt. Actors and onlookers: theater and twentiethcentury scientific views of nature. Northwestern University Press, Evanston, Ill., 1990. ISBN 0-8101-0836-4, 0-8101-0837-2 (paperback). 163 pp. LCCN PN2266.5 .S36 1990.

Salisbury:1993:MWN

[3184] Joyce E. Salisbury, editor. The medieval world of nature: a book of essays, volume 1550 of Garland reference library of the humanities; Garland medieval casebooks; v. 5. Garland, New York, NY, USA, 1993. ISBN 0-8153-0752-7. xvi + 265 pp. LCCN PN682.N3 M43 1993.

Burkhardt:1994:CCC

[3185] Frederick Burkhardt, Sydney Smith, David Kohn, William Montgomery, Charles Darwin, Correspondence, et al., editors. A calendar of the correspondence of Charles Darwin, 1821–1882, with supplement. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-43423-8 (hardcover). viii + 690 + 46 pp. LCCN QH31.D2 C15 1994. URL

http://www.loc.gov/catdir/description/cam026/94191109.html;
 http://www.loc.gov/catdir/toc/cam029/94191109.html.

Calder:2001:HRS

[3186] John F. Calder. The history of radiology in Scotland, 1896–2000. Dunedin Academic Press, Edinburgh, Scotland, 2001. ISBN 1-903765-05-6. xii + 127 pp. LCCN R895.6.G7 C34 2001.

Schut:2002:FFJ

[3187] Michael Schut, Diane Ackerman, et al., editors. Food and faith: justice, joy and daily bread. Living the Good News, Denver, CO, 2002. ISBN 1-889108-90-1. 296 pp. LCCN BR115.N87 F65 2002.

Selin:2003:NAC

[3188] Helaine Selin and Arne Kalland, editors. Nature across cultures: views of nature and the environment in non-western cultures, volume 4 of Science across cultures. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1235-7 (hardcover). xxiii + 481 pp. LCCN GF41 .N38 2003. URL http://www.loc.gov/

catdir/enhancements/fy0814/2003048865-d.html;   http://www.
loc.gov/catdir/toc/fy043/2003048865.html.

Chambers:2008:IPC

[3189] Neil Chambers, editor. The Indian and Pacific correspondence of Sir Joseph Banks, 1768–1820. The Pickering masters. Pickering and Chatto, London, UK, 2008–2014. ISBN 1-85196-835-0 (vol. 1.), 1-85196-836-9 (vol. 2.), 1-85196-837-7 (vol. 3.), 1-85196-838-5 (vol. 4.), 1-85196-839-3 (vol. 5.), 1-85196-840-7 (vol. 6.), 1-85196-634-X (vol. 7), 1-84893-380-0 (vol. 8). ???? pp. LCCN QH31.B19 A4 2008.

Yuasa:2008:OMS

[3190] Yasuo Yuasa, Shigenori Nagatomo, and John Wesley Megumu Krummel. Overcoming modernity: synchronicity and image-thinking. State University of New York Press, Albany, 2008. ISBN 0-7914-7401-1 (hardcover), 0-7914-7402-X (paperback). vii + 247 pp. LCCN B799 .Y79 2008. URL

http://www.loc.gov/catdir/toc/ecip0726/2007035479.html.

Burkhardt:2009:CCD

[3191] Frederick Burkhardt, editor. The correspondence of Charles Darwin: 1821–1860. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-12154-X (paperback anniversary set), 0-521-12191-4 (vol. 1: paperback), 0-521-12192-2 (vol. 2: paperback), 0-521-12193-0 (vol. 3: paperback), 0-521-12194-9 (vol. 4: paperback), 0-521-12195-7 (vol. 5: paperback), 0-521-12196-5 (vol. 6: paperback), 0-521-12197-3 (vol. 7: paperback), 0-521-12198-1 (vol. 8: paperback). ???? pp. LCCN QH31.D2 A4 2009.

Anderson:2012:NBV

[3192] Katharine Anderson, editor. The narrative of the Beagle voyage, 1831– 1836, The Pickering masters. Pickering and Chatto, London, UK, 2012. ISBN 1-85196-844-X (hardcover). LCCN F2936 .F58 2012. 4 volumes.

Bradley:2012:RPP

[3193] Mark Bradley and Kenneth R. Stow, editors. Rome, pollution, and propriety: dirt, disease, and hygiene in the eternal city from antiquity to modernity. British School at Rome studies. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-01443-3 (hardcover). xx + 320 pp. LCCN TD80.R65 R66 2012.

Fangerau:2012:MIP

[3194] Heiner Fangerau, Rethy Chhem, Irmgard Müller, and Shih-Chang Wang, editors. Medical imaging and philosophy: challenges, reflections and actions: conference proceedings, volume 3 of KulturAnamnesen; Wissenschaftsforschung. Steiner, Stuttgart, Germany, 2012. ISBN 3-515-10046-6 (paperback). LCCN R836 .M435 2012eb.

Herran:2012:FSD

[3195] Néstor Herrán and Xavier Roqué. La física en la dictadura: físicos, cultura y poder en España, 1939–1975. Universitat Autònoma de Barcelona, Bellaterra, Spain, 2012. ISBN 84-490-3238-5. 310 pp. LCCN QC9.S7 F58 2012.

Horstmanshoff:2012:BST

[3196] H. F. J. Horstmanshoff, Helen King, and Claus Zittel, editors. Blood, sweat, and tears: the changing concepts of physiology from antiquity into early modern Europe, volume 25 of Intersections. Brill, Leiden, The Netherlands, 2012. ISBN 90-04-22918-3 (hardcover), 90-04-22920-5 (ebook). xxvi + 772 pp. LCCN R135 .B56 2012.

Kosmann-Schwarzbach:2012:SDP

[3197] Yvette Kosmann-Schwarzbach. Siméon-Denis Poisson: les mathématiques au service de la science. (French) [Siméon-Denis Poisson: mathematics in the service of science]. Histoire des mathématiques et des sciences physiques. École polytechnique, Palaiseau, France, 2012. ISBN 2-7302-1584-0 (paperback). xiii + 522 pp. LCCN QA29.P64 S55 2012.

Lettevall:2012:NTC

[3198] Rebecka Lettevall, Geert Somsen, and Sven Widmalm, editors. Neutrality in Twentieth-Century Europe: intersections of science, culture, and politics after the First World War. Routledge studies in cultural history. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2012. ISBN 0-415-89377-1 (hardcover). ix + 351 pp. LCCN D723 .N49 2012eb.

Mazzolini:2012:DRS

[3199] Renato G. (Renato Giuseppe) Mazzolini and Hans-Jörg Rheinberger, editors. Differing routes to stem cell research: Germany and Italy, volume 26 of Annali dell’Istituto storico italo-germanico in Trento. Contributi. Societa editrice il Mulino, Bologna, Italia, 2012. ISBN 3-428-13849-X (Duncker and Humblot: paperback), 88-15-23878-6 (Societa editrice il Mulino: paperback). LCCN QH588.S83 D57 2012.

Aaserud:2013:LLQ

[3200] Finn Aaserud and J. L. Heilbron. Love, Literature, and the Quantum Atom: Niels Bohr’s 1913 Trilogy Revisited. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-968028-0 (hardcover). viii + 284 pp. LCCN QC774.B64 A19 2013.

Boudia:2013:THR

[3201] Soraya Boudia and Nathalie Jas, editors. Toxicants, health and regulation since 1945, volume 9 of Studies for the Society for the Social History of Medicine. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-403-3 (hardcover), 1-78144-035-2 (e-book). xiii + 193 pp. LCCN RA1219.3 .T69 2013.

Buchwald:2013:NOC

[3202] Jed Z. Buchwald and Mordechai Feingold. Newton and the origin of civilization. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-15478-3 (hardcover). ix + 528 pp. LCCN QC16.N7 B93 2013.

Conti:2013:BTI

[3203] Lino Conti and Paolo Capitanucci, editors. Breve trattato intorno alla preservatione, et cura della peste. (Italian) [Short treatise on the preservation and care of plague]. Edizioni Centro Stampa, Città di Castello, Italia, 2013. ISBN ???? x + 83 pp. LCCN ????

Conti:2013:IDM

[3204] Lino Conti, editor. Itinerari della medicina nel Seicento. (Italian) [Routes of medicine in the Seventeenth Century], volume 11 of Ricerche Filosofiche. Edizioni Centro Stampa, Città di Castello, Italia, 2013. ISBN ???? 144 (est.) pp. LCCN ????

Doody:2013:AS

[3205] John Doody, Adam Goldstein, and Kim Paffenroth, editors. Augustine and science. Augustine in conversation. Lexington Books, Lexington, MA, USA, 2013. ISBN 0-7391-7434-7 (hardcover), 0-7391-7435-5 (e-book). vi + 233 pp. LCCN BR65.A9 A844 2013.

Flood:2013:RGH

[3206] John Flood, James R. Ginther, and Joseph Ward Goering, editors. Robert Grosseteste and his intellectual milieu. New editions and studies, volume 24 of Papers in Mediaeval Studies (PMS). Pontifical Institute of Mediavel Studies, Toronto, ON, Canada, 2013. ISBN 0-88844-824-4 (hardcover), 1-77110-353-1 (e-book). LCCN B765.G74 I58 2003.

Friis:2013:ESE

[3207] Ib Friis, Michael Harbsmeier, and Jurgen Bøk Simonsen, editors. Early scientific expeditions and local encounters: new perspectives on Carsten Niebuhr and The Arabian Journey: proceedings of a Symposium on the occasion of the 250. anniversary of the the Royal Danish Expedition to Arabia Felix. Det Kongelige Danske Videnskabernes Selskab, Copenhagen, Denmark, 2013. ISBN 87-7304-375-3. LCCN ????

Harman:2013:OSR

[3208] Oren Solomon Harman and Michael R.. Dietrich, editors. Outsider scientists: routes to innovation in biology. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-07837-X (hardcover), 0-226-07840-X (paperback), 0-226-07854-X (e-book). 374 pp. LCCN QH26 .O98 2013.

Hirsch:2013:SLH

[3209] Luise Hirsch. From the shtetl to the lecture hall: Jewish women and cultural exchange. Studies in Judaism. University Press of America, Inc., Lanham, MD, USA, 2013. ISBN 0-7618-5992-6 (paperback). xiv + 319 pp. LCCN LC1571.G3 H57 2013.

Klemm:2013:BWK

[3210] Tanja Klemm. Bildphysiologie: Wahrnehmung und Körper in Mittelalter und Renaissance. (German) [Picture Physiology: perception and body in the Middle Ages and the Renaissance]. Akademie-Verlag, Berlin, Germany, 2013. ISBN 3-05-006478-1. xii + 323 pp. LCCN B828.45 .K64

Lehmann:2013:ARR

[3211] Yves Lehmann, editor. Aristoteles romanus: la réception de la science aristotélicienne dans l’Empire gréco-romain, volume 17 of Recherches sur les rhétoriques religieuses. Brepols, Turnhout, Belgium, 2013. ISBN 2-503-54913-6 (paperback). LCCN B491.K6 A75 2013; B485 .A75 2013.

Li:2013:LZN

[3212] Wenchao Li and Hartmut Rudolph, editors. “Leibniz” in der Zeit des Nationalsozialismus, volume 42 of Studia leibnitiana. Sonderhefte. Steiner, Stuttgart, Germany, 2013. ISBN 3-515-10308-2 (paperback). LCCN B2598 .L452 2013.

Neufeld:2013:SIA

[3213] Michael J. Neufeld, editor. Spacefarers: images of astronauts and cosmonauts in the heroic era of spaceflight, A Smithsonian contribution to knowledge. Smithsonian Institution Scholarly Press, Washington, DC, USA, 2013. ISBN 1-935623-19-2 (hardcover). LCCN P96.A792 S63 2013.

Reinarz:2013:MW

[3214] Jonathan Reinarz and Leonard Schwarz, editors. Medicine and the workhouse, volume 27 of Rochester studies in medical history. University of Rochester Press, Rochester, NY, 2013. ISBN 1-58046-448-3 (hardcover). ISSN 1526-2715. LCCN RA418.5.P6 M433 2013.

Rey-Courtel:2013:MHQ

[3215] Anne-Lise Rey-Courtel. Méthode et histoire: quelle histoire font les historiens des sciences et des techniques?, volume 6 of Histoire et philosophie des sciences. Classiques Garnier, Paris, France, 2013. ISBN 2-8124-1419-7. ISSN 2117-3508. 513 pp. LCCN Q126.9 .M45 2014.

Sabin:2013:BPE

[3216] Paul Sabin. The bet: Paul Ehrlich, Julian Simon, and our gamble over Earth’s future. Yale University Press, New Haven, CT, USA, 2013. ISBN 0-300-17648-1 (hardcover). xv + 304 pp. LCCN HC79.E5 S213 2013.

Schell:2013:SSD

[3217] Patience A. Schell. The Sociable Sciences: Darwin and His Contemporaries in Chile. Palgrave studies in the history of science and technology. Palgrave Macmillan, Basingstoke, 2013. ISBN 1-299-54768-0, 1-137-28605-9, 1-137-28606-7 (e-book). LCCN QH119 .S34 2013.

Secord:2013:VSB

[3218] James A. Secord. Visions of science: books and readers at the Dawn of the Victorian Age: Sandars lectures, University of Cambridge, 25–27 February 2013. Sandars lectures. Cambridge University Press, Cambridge, UK, 2013. 52 + 10 pp. LCCN ????

Shapiro:2013:TBS

[3219] Adam R. Shapiro, editor. Trying biology: the Scopes trial, textbooks, and the antievolution movement in American schools. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-02945-X (hardcover), 0-226-02959-X (e-book). LCCN QH362 .S53 2013.

Sorrenson:2013:PMI

[3220] Richard John Sorrenson. Perfect mechanics: instrument makers at the Royal Society of London in the Eighteenth Century. Docent Press, Boston, MA, USA, 2013. ISBN 0-9887449-2-9. ix + 240 pp. LCCN Q185 .S67 2013.

Steinmetz:2013:SEI

[3221] George Steinmetz. Sociology and empire: the imperial entanglements of a discipline. Politics, history, and culture. Duke University Press, Durham, NC, USA, 2013. ISBN 0-8223-5258-3 (hardcover), 0-8223-5279-6 (paperback), 0-8223-9540-1 (e-book). xvi + 610 pp. LCCN JC359 .S693 2013.

Thomas:2013:HR

[3222] Adrian Thomas and Arpan K. Banerjee. The history of radiology. Oxford medical histories. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-963997-3 (hardcover). xiii + 222 pp. LCCN R895.5 .T46 2013.

Thorsen:2013:ADC

[3223] Liv Emma Thorsen, Karen A. (Karen Ann) Rader, and Adam Dodd, editors. Animals on display: the creaturely in museums, zoos, and natural history. Animalibus: of animals and cultures. The Pennsylvania State University Press, University Park, Pennsylvania, 2013. ISBN 0-271-06070-0 (hardcover). 222 pp. LCCN QL71.A1 A55 2013.

vanMiert:2013:COE

[3224] Dirk van Miert, editor. Communicating observations in early modern letters (1500–1675): epistolography and epistemology in the age of the scientific revolution, volume 23 of Warburg Institute Colloquia. Warburg Institute, London, UK, 2013. ISBN 1-908590-46-7 (paperback). ISSN 1352-9986. ix + 268 pp. LCCN Q223.2 .C66 2013.

Watkins:2013:DOH

[3225] Eric Watkins, editor. The divine order, the human order, and the order of nature: historical perspectives. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-993440-1. xxviii + 240 pp. LCCN B105.O7 D58 2013.

Wilson:2013:RCS

[3226] Adrian Wilson. Ritual and conflict: the social relations of childbirth in early modern England. The history of medicine in context. Ashgate, Farnham, Surrey, England, 2013. ISBN 1-4094-6812-7 (hardcover), 1-4094-6813-5 (e-book). vi + 261 pp. LCCN RG518.G7 W55 2014.

Zachhuber:2013:TSN

[3227] Johannes Zachhuber. Theology as science in nineteenth century Germany: from F. C. Baur to Ernst Troeltsch. Changing paradigms in historical and systematic theology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-964191-9, 0-19-175249-5 (ebook). xii + 318 pp. LCCN BV4140.G3 Z33 2013.

Ackermann:2014:SID

[3228] Silke Ackermann, Richard L. (Richard Lynn) Kremer, and Mara Miniati, editors. Scientific instruments on display, volume 46(4) of History of science and medicine library; Scientific instruments and collections. Brill, Leiden, The Netherlands, 2014. ISBN 90-04-26439-6 (hardcover), 90-04-26440-X (e-book). ISSN 1872-0684. xxxiv + 231 pp. LCCN Q185 .S39 2014.

Artstein:2014:MRW

[3229] Zvi Artstein and Aland Hercberg. Mathematics and the real world: the remarkable role of evolution in the making of mathematics. Prometheus Books, Amherst, NY, USA, 2014. ISBN 1-61614-091-7 (hardcover), 1-61614-546-3 (e-book). 426 pp. LCCN QC6 .A6913 2014.

Baack:2014:UCC

[3230] Lawrence J. Baack. Undying curiosity: Carsten Niebuhr and the Royal Danish Expedition to Arabia (1761–1767), volume 22 of Oriens et occidens: Studien zu antiken Kulturkontakten und ihrem Nachleben; History. Franz Steiner Verlag, Stuttgart, Germany, 2014. ISBN 3-515-10768-1 (hardcover). 443 pp. LCCN DS206 .B33 2014.

BerquistSoule:2014:BUE

[3231] Emily Berquist Soule. The Bishop’s Utopia: Envisioning Improvement in Colonial Peru. University of Pennsylvania Press, Philadelphia, 2014. ISBN 0-8122-4591-1. LCCN F3611.T8 S68 2014.

Bod:2014:NHH

[3232] Rens Bod. A new history of the humanities: the search for principles and patterns from antiquity to the present. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-175747-0. xv + 384 pp. LCCN AZ201.B64 2013.

Braun:2014:BRM

[3233] Lundy Braun. Breathing race into the machine: the surprising career of the spirometer from plantation to genetics. University of Minnesota Press, Minneapolis, MN, USA, 2014. ISBN 0-8166-8357-3 (hardcover), 0-8166-8359-X (paperback), 1-4529-4099-1. 304 pp. LCCN RC773 2014.

Butrica:2014:NHN

[3234] Andrew J. Butrica. The navigators: a history of NASA’s deep-space navigation. Andrew J. Butrica, San Bernardino, CA, UYSA, 2014. ISBN 1-4927-7783-8. iii + 367 pp. LCCN TL3026 B88 2014.

Chambers:2014:LJE

[3235] D. D. C. Chambers and David Ian Galbraith, editors. The Letterbooks of John Evelyn. University of Toronto Press, Toronto, ON, Canada, 2014. ISBN 1-4426-4786-8 (bound: set). lxiii + 1236 (2 volumes) pp. LCCN PR3433 E93 A4 C444 2014.

Coopmans:2014:RSP

[3236] Catelijne Coopmans, editor. Representation in scientific practice revisited. Inside technology. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-52538-0 (paperback), 0-262-31915-2 (e-book). ix + 366 pp. LCCN Q180.55.M4 R455 2014.

Dawes:2014:COA

[3237] Laura Dawes. Childhood obesity in America: biography of an epidemic. Harvard University Press, Cambridge, MA, USA, 2014. ISBN 0-674-28144-6 (hardcover), 0-674-36957-2. viii + 305 pp. LCCN RJ399.C6 D39

2014. URL http://www.degruyter.com/view/product/429828.

Donato:2014:SDM

[3238] Maria Pia Donato and Valentina Mazzei. Sudden death: medicine and religion in eighteenth-century Rome. The history of medicine in context.

Ashgate, Farnham, Surrey, UK, 2014. ISBN 1-4724-1873-5 (hardcover). vi + 229 pp. LCCN RB150.S84 D6613 2014.

Dooley:2014:CAR

[3239] Brendan Maurice Dooley, editor. A companion to astrology in the Renaissance, volume 49 of Brill’s companions to the Christian tradition. Brill, Leiden, The Netherlands, 2014. ISBN 90-04-18352-3 (hardcover), 90-04-26230-X (e-book). ISSN 1871-6377. xvi + 453 pp. LCCN BF1676 .C66 2014.

Duda:2014:PLC

[3240] Roman Duda, editor. Pearls from a lost city: the Lvov school of mathematics, volume 40 of History of Mathematics. American Mathematical Society, Providence, RI, USA, 2014. ISBN 1-4704-1076-1. LCCN QA14.U382 L85313 2014. Translation from Polish to English by Daniel Davies.

Duppe:2014:FEA

[3241] Till. Düppe and E. Roy Weintraub. Finding equilibrium: Arrow, Debreu, McKenzie and the problem of scientific credit. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-15664-6 (hardcover). xxv + 1 + 276 pp. LCCN HB145 .D87 2014.

Efron:2014:CCJ

[3242] Noah J. Efron. A chosen calling: Jews in science in the twentieth century. Medicine, science, and religion in historical context. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1381-7 (hardcover), 1-4214-1382-5 (electronic). xv + 149 pp. LCCN Q180.55.S62 E37 2014.

Enenkel:2014:ZEM

[3243] K. A. E. Enenkel and P. J. (Paul J.) Smith, editors. Zoology in early modern culture: intersections of science, theology, philology, and political and religious education, volume 32 of Intersections. Brill, Leiden, The Netherlands, 2014. ISBN 90-04-26823-5 (hardcover), 90-04-27917-2 (ebook). ISSN 1568-1181. xxiv + 522 pp. LCCN QL15 .E25 2014.

Epstein:2014:TIM

[3244] Katherine C. Epstein. Torpedo: Inventing the military-industrial complex in the United States and Great Britain. Harvard University Press, Cambridge, MA, USA, 2014. ISBN 0-674-72526-3, 0-674-72628-6 (e-book). 305 pp. LCCN V850 .E67 2014.

Fleming:2014:TAB

[3245] James Rodger Fleming, Ann Johnson, and Peter Brimblecombe, editors. Toxic airs: body, place, planet in historical perspective. University of Pittsburgh Press, Pittsburgh, PA, USA, 2014. ISBN 0-8229-6290-X (paperback), 0-8229-7952-7 (e-book). xiv + 284 pp. LCCN TD883 .T66 2014.

Flipse:2014:CWN

[3246] Ab (Abraham Christiaan) Flipse. Christelijke wetenschap: Nederlandse rooms-katholieken en gereformeerden over de natuurwetenschap, 1880– 1940. (Dutch) [Christian Science: Dutch Catholics and Calvinists on science, 1880–1940], volume 44 of Passage reeks. Verloren, Hilversum, The Netherlands, 2014. ISBN 90-8704-415-1 (paperback). 327 pp. LCCN BL240.3 .F54 2014.

Fontana:2014:LII

[3247] Johannes Fontana and Horst Kranz. Liber instrumentorum iconographicus = Ein illustriertes Maschinenbuch, volume Band 66 of Boethius, ; Wissenschaftsgeschichte. Franz Steiner Verlag, Stuttgart, Germany, 2014. ISBN 3-515-10660-X (hardcover). ISSN 0523-8226. 192 pp. LCCN TJ230 .F65 2014.

Galluzzi:2014:LFN

[3248] Paolo Galluzzi. “Libertà di filosofare in naturalibus”: i mondi paralleli di Cesi e Galileo. (Italian) [“Freedom of philosophizing in nature”: the parallel worlds of Cesi and Galileo]. Scienze e lettere editore commerciale, Roma, Italy, 2014. ISBN 88-218-1092-5. 599 pp. LCCN Q143.C38 G35 2014.

Garrison:2014:FHB

[3249] Daniel H. Garrison and Malcolm Howard Hast, editors. The fabric of the human body: an annotated translation of the 1543 and 1555 editions of “De humani corporis fabrica libri septem” [by Andreas Vesalius]. Karger, Basel, Switzerland, 2014. ISBN 3-318-02246-2 (hardcover). cxix + 1337 + v (two volumes) pp. LCCN QM25.

Giacomotto-Charra:2014:LCE

[3250] Violaine Giacomotto-Charra and Christine Silvi, editors. Lire, choisir, écrire: la vulgarisation des savoirs du Moyen Âge à la Renaissance, volume 43 of Études et rencontres de l’École des chartes. École des chartes, Paris, France, 2014. ISBN 2-35723-041-X (paperback). ISSN 1289-7566. LCCN Q225.2.E85 L57 2014.

Goy:2014:KTB

[3251] Ina Goy and Eric Watkins, editors. Kant’s Theory of Biology. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 2014. ISBN 3-11-022578-6 (hardcover), 1-306-52936-0 (MyiLibrary). x + 321 pp. LCCN

QH331 .K273 2014. URL http://www.degruyter.com/viewbooktoc/
product/43450.

Granada:2014:CRD

[3252] Miguel Angel Granada, Adam Mosley, and Nicholas Jardine. Christoph Rothmann’s Discourse on the comet of 1585: an edition and translation with accompanying essays, volume 44. Medieval and early modern science of History of science and medicine library; volume 22. Brill, Leiden, The Netherlands, 2014. ISBN 90-04-26034-X (hardcover). ISSN 1567-8393. xx + 376 pp. LCCN QB721 .R6813 2014.

Haller:2014:SMP

[3253] John S. Haller, Jr. Shadow medicine: the placebo in conventional and alternative therapies. Columbia University Press, New York, NY, USA, 2014. ISBN 0-231-16904-3 (hardcover), 0-231-53770-0 (e-book). xxix + 255 pp. LCCN R733 .H35 2014.

Hamilton:2014:EPS

[3254] Andrew Hamilton, editor. The evolution of phylogenetic systematics, volume 5 of Species and systematics. University of California Press, Berkeley, CA, USA, 2014. ISBN 0-520-27658-2 (hardcover), 0-520-95675-3 (e-book). viii + 311 pp. LCCN QH83 .E96 2014.

Hamlin:2014:EED

[3255] Kimberly A. Hamlin. From Eve to evolution: Darwin, science, and women’s rights in Gilded Age America. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-13461-X (hardcover), 0-226-13475-X (e-book). vii + 238 pp. LCCN HQ1191.U6 H36 2014.

Hamlin:2014:MTH

[3256] Christopher Hamlin. More than hot: a short history of fever. Johns Hopkins biographies of disease. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1502-X (paperback), 1-4214-1503-8 (electronic). xii + 383 pp. LCCN RB129 .H36 2014.

Hearnshaw:2014:AST

[3257] J. B. Hearnshaw. The analysis of starlight: two centuries of astronomical spectroscopy. Cambridge University Press, Cambridge, UK, second edition, 2014. ISBN 1-107-03174-5 (hardcover), 1-139-38277-2 (e-book). xvi + 367 pp. LCCN QB465 .H43 2014.

Hollings:2014:MAI

[3258] Christopher Hollings. Mathematics across the Iron Curtain: a history of the algebraic theory of semigroups, volume 41 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2014. ISBN 1-4704-1493-7. xi + 441 pp. LCCN QA182 .H65 2014.

Ito:2014:LZVa

[3259] Takashi Ito. London Zoo and the Victorians, 1828–1859. Studies in history new series. The Royal Historical Society, London, UK, 2014. ISBN 0-86193-321-4. ISSN 0269-2244. xi + 204 pp. LCCN QL76.5.G72 L6648 2014.

Ito:2014:LZVb

[3260] Takashi Ito. London Zoo and the Victorians, 1828–1859. Royal Historical Society studies in history. New series. The Royal Historical Society, Woodbridge, Suffolk, UK, 2014. ISBN 0-86193-321-4 (hardcover), 1-78204-261-X (e-book). xi + 204 pp. LCCN QL76.5.G72 L669 2014.

Ito:2014:LZVc

[3261] Takashi Ito. London Zoo and the Victorians, 1828–1859, volume 87 of Royal Historical Society Studies in History New Series. Boydell and Brewer Group Ltd, New York, NY, USA, 2014. ISBN 0-86193-321-4, 1-78204-282-2. 429 pp. LCCN QL76.5.G72 L6648 2014.

Janssen:2014:CCE

[3262] Michel Janssen and Christopher Lehner, editors. The Cambridge Companion to Einstein. Cambridge University Press, Cambridge, UK, 2014. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). xvi + 562 pp. LCCN QC16.E5 C36 2014.

Kassler:2014:STR

[3263] Jamie Croy Kassler. Seeking truth: Roger North’s notes on Newton and correspondence with Samuel Clarke c. 1704–1713. Ashgate, Surrey, England, 2014. ISBN 1-4094-4921-1, 1-4094-4922-X (e-book), 1-4724-0646-X (e-pub). 387 pp. LCCN DA437.N87 .K38 2014.

Kirkham:2014:WMA

[3264] Anne (Deborah Anne) Kirkham and Cordelia Warr, editors. Wounds in the Middle Ages. The history of medicine in context. Ashgate, Farnham, Surrey, UK, 2014. ISBN 1-4094-6569-1 (hardcover), 1-4094-6570-5 (e-PDF), 1-4094-6571-3 (e-PUB). xiv + 254 pp. LCCN RD93 .W6882 2014.

Lambright:2014:WMN

[3265] W. Henry Lambright. Why Mars: NASA and the politics of space exploration. New series in NASA history. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1279-9 (hardcover), 1-4214-1280-2 (e-book). x + 320 pp. LCCN TL799.M3 L36 2014.

Lettow:2014:RRG

[3266] Susanne Lettow, Florence Vienne, Peter Hanns Reill, Jocelyn Holland, Joan Steigerwald, Renato G. (Renato Giuseppe) Mazzolini, Sara Eigen Figal, Penelope Deutscher, Staffan Müller-Wille, Robert Bernasconi, and Alison Stone, editors. Reproduction, race, and gender in philosophy and the early life sciences. SUNY series, Philosophy and race. State University of New York Press, Albany, 2014. ISBN 1-4384-4949-6 (hardcover), 1-4384-4950-X (e-book). vi + 294 pp. LCCN HT1521 .R456 2014.

Lightman:2014:ASN

[3267] Bernard V. Lightman and Michael S. Reidy, editors. The age of scientific naturalism: Tyndall and his contemporaries, volume 24 of Science and culture in the nineteenth century. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 1-84893-463-7 (hardcover), 1-306-39773-1 (e-book), 1-78144-085-9 (e-book). xv + 256 pp. LCCN QC16 .T9; Q175.32.N38.

Lightman:2014:EVC

[3268] Bernard V. Lightman, Bennett Zon, and Cannon Schmitt, editors. Evolution and Victorian culture. Cambridge studies in nineteenth-century literature and culture. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-02842-6 (hardcover). xvii + 320 pp. LCCN HM626 .E945 2014.

Lunbeck:2014:AN

[3269] Elizabeth Lunbeck. The Americanization of narcissism. Harvard University Press, Cambridge, MA, USA, 2014. ISBN 0-674-72486-0. 367 pp. LCCN BF575.N35 L86 2014.

Maglen:2014:ESQ

[3270] Krista Maglen. The English system: quarantine, immigration and the making of a port sanitary zone. Manchester University Press, Manchester, UK, 2014. ISBN 0-7190-8965-4. x + 240 pp. LCCN ????

Margocsy:2014:CVS

[3271] Dániel Margócsy. Commercial visions: science, trade, and visual culture in the Dutch Golden Age. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-11774-X (hardcover). x + 319 pp. LCCN Q127.N2 M37 2014.

Martin:2014:SAR

[3272] Craig Martin. Subverting Aristotle: religion, history, and philosophy in early modern science. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1316-7 (hardcover), 1-4214-1317-5 (electronic). 262 pp. LCCN B485 .M267 2014.

Martin:2014:SNC

[3273] Paula J. Martin. Suzanne Noël: cosmetic surgery, feminism and beauty in early twentieth-century France. The history of medicine in context. Ashgate, Farnham, Surrey, England, 2014. ISBN 1-4724-1188-9 (hardcover), 1-4724-1189-7 (e-book), 1-4724-1190-0 (e-pub). 157 pp. LCCN RD119 .M387 2014.

Mellyn:2014:MTT

[3274] Elizabeth Walker Mellyn. Mad Tuscans and their families: a history of mental disorder in early modern Italy. University of Pennsylvania Press, Philadelphia, 2014. ISBN 0-8122-4612-8 (hardcover). 290 pp. LCCN RC450.I82 M45 2014.

Miller:2014:WCL

[3275] Timothy S. Miller and John W. Nesbitt. Walking corpses: leprosy in Byzantium and the Medieval West. Cornell University Press, Ithaca, NY, USA, 2014. ISBN 0-8014-5135-3 (hardcover). xiv + 243 pp. LCCN RC154.3 .M55 2014.

Mollan:2014:WPE

[3276] Charles Mollan. William Parsons, 3rd Earl of Rosse: astronomy and the castle in Nineteenth-Century Ireland. Manchester University Press, Manchester, UK, 2014. ISBN 0-7190-9144-6. xxii + 368 pp. LCCN QB36.R77 W55 2014.

Muller-Wille:2014:FBE

[3277] Ludger Müller-Wille and Rosemary Lévy Zumwalt. The Franz Boas enigma: Inuit, Arctic, and sciences. Baraka Books, Montréal, Québec, Canada, 2014. ISBN 1-77186-001-4 (paperback), 1-77186-014-6 (epub), 1-77186-015-4 (pdf), 1-77186-016-2 (mobi). 186 pp. LCCN GN21.B56 M85 2014.

Nachtomy:2014:LSE

[3278] Ohad Nachtomy and Justin E. H. Smith, editors. The life sciences in early modern philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-998731-9 (hardcover). xiii + 256 pp. LCCN QH331 .L538 2014.

Osseo-Asare:2014:BRS

[3279] Abena Dove Agyepoma Osseo-Asare. Bitter roots: the search for healing plants in Africa. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-08552-X (hardcover), 0-226-08602-X (paperback). vii + 300 pp. LCCN RS181 .O87 2014.

Pleins:2014:PDG

[3280] J. David Pleins. In praise of Darwin: Georges Romanes and the evolution of a Darwinian believer. Bloomsbury, New York, NY, USA, 2014. ISBN 1-62356-594-4 (paperback), 1-62356-554-5 (hardcover). xvii + 397 pp. LCCN QH31.R7 P54 2014. URL http://www.loc.

gov/catdir/enhancements/fy1512/2014434386-b.html;   http://
www.loc.gov/catdir/enhancements/fy1512/2014434386-d.html;
http://www.loc.gov/catdir/enhancements/fy1512/2014434386-t.
html.

Prasad:2014:ITT

[3281] Amit Prasad. Imperial technoscience: transnational histories of MRI in the United States, Britain, and India. Inside technology series. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02695-3 (hardcover), 0-262-32206-4 (e-book), 1-306-53160-8 (e-book). xi + 219 pp.

LCCN RC78.7.N83 P737 2014. URL http://www.jstor.org/stable/
10.2307/j.ctt9qf6n7.

Rayward:2014:IBB

[3282] W. Boyd Rayward. Information Beyond Borders: International Cultural and Intellectual Exchange in the Belle Époque. Ashgate, Farnham, UK, 2014. ISBN 1-4094-4225-X, 1-306-54983-3 (e-book). 336 pp. LCCN ZA3157.

Reeves:2014:ENO

[3283] Eileen Adair Reeves. Evening news: optics, astronomy, and journalism in early modern Europe. Material texts. University of Pennsylvania Press, Philadelphia, 2014. ISBN 0-8122-4574-1 (hardcover). 308 pp. LCCN PN5110 .R44 2014.

Reinarz:2014:PSH

[3284] Jonathan Reinarz. Past scents: historical perspectives on smell. Studies in sensory history. University of Illinois Press, Urbana, IL, USA, 2014. ISBN 0-252-03494-5 (hardcover), 0-252-07979-5 (paperback). xi + 279 pp. LCCN QP458 .R45 2014.

Rickles:2014:BHS

[3285] Dean Rickles. A Brief History of String Theory From Dual Models to M-Theory. The Frontiers Collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-642-45128-4. xix + 264 pp. LCCN QC794.6 .S85. URL http://www.springerlink.

com/content/978-3-642-45128-7.

Russell:2014:OSD

[3286] Russell. Open Standards and the Digital Age: History, Ideology, and Networks. Cambridge studies in the emergence of global enterprise. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-03919-3 (hardcover), 1-107-61204-7 (paperback), 1-139-85655-3 (e-book). 316 pp. LCCN T59.2.U6 R87 2014.

Secord:2014:VSB

[3287] James A. Secord. Visions of science: books and readers at the dawn of the Victorian age. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-967526-0 (hardcover), 0-19-166275-5 (e-book). xiii + 306 + 8 pp. LCCN PR468.S34 S43 2014.

Sharma:2014:WLB

[3288] Avinash Sharma. We lived for the body natural medicine and public health in Imperial Germany. NIU Press, DeKalb, IL, USA, 2014. ISBN 0-87580-704-6 (paperback), 1-60909-154-X (e-book). ix + 210 pp. LCCN RZ440 .S468 2014.

Shen:2014:UNM

[3289] Grace Yen Shen. Unearthing the nation: modern geology and nationalism in republican China. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-09040-X (hardcover), 0-226-09054-X (e-book), 1-306-42853-X (MyiLibrary). 309 pp. LCCN QE13.C6 S54 2014.

Smith:2014:WMK

[3290] Pamela H. Smith, Amy R. W. Meyers, and Harold J. Cook, editors. Ways of making and knowing: the material culture of empirical knowledge. The Bard Graduate Center cultural histories of the material world. University of Michigan Press, Ann Arbor, MI, USA, 2014. ISBN 0-472-11927-3, 0-472-12018-2 (hardcover). xi + 430 pp. LCCN AZ101 .W39 2014.

Spanagel:2014:DCA

[3291] David I. Spanagel. DeWitt Clinton and Amos Eaton: geology and power in early New York. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1104-0 (hardcover), 1-4214-1105-9 (e-book). xii + 270 pp. LCCN Q143.E3 S65 2014.

Spary:2014:FFN

[3292] E. C. (Emma C.) Spary. Feeding France: new sciences of food, 1760– 1815, volume 19 of Cambridge social and cultural histories. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-03105-2 (hardcover), 1-139-95981-6 (e-book), 1-139-38118-0 (e-book), 1-139-95770-8 (e-book), 1-139-95875-5 (e-book). xii + 418 pp. LCCN HD6012.5 .S65 2014.

Steen:2014:ASO

[3293] Kathryn Steen. The American synthetic organic chemicals industry: war and politics, 1910–1930. The University of North Carolina Press, Chapel Hill, NC, USA, 2014. ISBN 1-4696-1290-9 (paperback), 1-4696-1291-7 (e-book), 1-4696-1445-6 (e-book). xii + 403 pp. LCCN HD9651.5 .S73 2014.

Stern:2014:TAC

[3294] Sacha Stern and Charles Burnett, editors. Time, astronomy, and calendars in the Jewish tradition, volume 3 of Time, astronomy, and calendars. Brill, Leiden, The Netherlands, 2014. ISBN 90-04-25965-1 (hardcover), 90-04-25966-X (e-book). ISSN 2211-632X. xxi + 365 pp. LCCN CE35 .T46 2014.

Tabery:2014:BVS

[3295] James Tabery. Beyond versus: the struggle to understand the interaction of nature and nurture. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02737-2. 304 pp. LCCN BF341.

teHeesen:2014:NCM

[3296] Anke te Heesen. The newspaper clipping. A modern paper object. Manchester University Press, Manchester, UK, 2014. ISBN 0-7190-8702-3. 320 pp. LCCN N6480-6495.

Timmermann:2014:HLC

[3297] Carsten Timmermann. A history of lung cancer: the recalcitrant disease. Science, technology and medicine in modern history. Palgrave Macmillan, Houndsmills, Basingstoke, Hampshire, UK, 2014. ISBN 1-4039-8802-1 (hardcover), 1-137-38423-9 (e-book). x + 244 pp. LCCN RC280.L8 T56 2014.

Tobies:2014:WIR

[3298] Renate Tobies, Annette Vogt, and Valentine Pakis, editors. Women in industrial research, volume 8 of Wissenschaftskultur um 1900, ; Wissenschaftsgeschichte. Franz Steiner, Stuttgart, Germany, 2014. ISBN 3-515-10670-7 (hardcover). ISSN 1613-673X. xv + 258 pp. LCCN Q130 .W6535 2014.

Tsu:2014:STM

[3299] Jing Tsu and Benjamin A. Ellman, editors. Science and technology in modern China, 1880s–1940s, volume 27 of China studies. Brill, Leiden, The Netherlands, 2014. ISBN 90-04-25853-1 (hardcover), 90-04-26878-2 (e-book). ISSN 1570-1344. xii + 347 pp. LCCN Q127.C5 S332 2014.

Wade:2014:MGR

[3300] Peter Wade, Carlos López Beltrán, Eduardo Restrepo, and Ricardo Ventura Santos, editors. Mestizo genomics: race mixture, nation, and science in Latin America. Duke University Press, Durham, NC, USA, 2014. ISBN 0-8223-5648-1 (hardcover), 0-8223-5659-7 (paperback), 0-8223-7672-5 (ebook). xii + 304 pp. LCCN F1419.A1 M48 2014.

Weisz:2014:CDT

[3301] George Weisz. Chronic disease in the twentieth century: a history. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1302-7 (hardcover), 1-4214-1303-5 (paperback), 1-4214-1304-3 (electronic). xvi + 307 pp. LCCN RA418 .W425 2014.

Wiley:2014:CMB

[3302] Andrea S. Wiley. Cultures of milk: the biology and meaning of dairy products in the United States and India. Harvard University Press, Cambridge, MA, USA, 2014. ISBN 0-674-72905-6, 0-674-36969-6. xi + 193 pp. LCCN GT2920.M55 W54 2014. URL http://www.degruyter.com/

view/product/429829.

Sommaruga:2015:TRI

[3303] Giovanni Sommaruga and Thomas (Thomas Adrian) Strahm, editors. Turing’s revolution: the impact of his ideas about computability. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2015. ISBN 3-319-22155-8, 3-319-22156-6 (ebook). xxiv + 329 pp. LCCN QA29.T8 T87 2015. URL

http://link.springer.com/10.1007/978-3-319-22156-4;   http:
//swbplus.bsz-bw.de/bsz44537344xcov.htm.

Skoldberg:2016:AIO

[3304] Ulla Johansson Skoldberg, Jill Woodilla, and Ariane Berthoin Antal, editors. Artistic interventions in organizations: research, theory and practice. Routledge research in creative and cultural industries management. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2016. ISBN 1-138-82113-6 (hardcover), 1-315-74348-5 (e-book). xxi + 257 pp. LCCN HD58.8 .A78 2016.